A Complete Bibliography of Publications in the Journal of Fish Biology (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

31 August 2024 Version 1.08

Title word cross-reference

3 [1665]. $54.95 [1772]. $69.00 [1651]. $69.95 [1405]. $80.00 [1493]. > [2283]. ^{+} [3160, 399, 2330, 476, 804]. ^{13} [2555, 2783, 2833, 1144, 2586, 2489]. 15 [2555, 2783, 2164, 1144, 2586, 2489]. ^{2+} [1044]. ^{31} [3319]. ^{85} [1331]. ^{st} [1824]. ^{th} [1428, 1706]. _{2} [2384]. _{4} [3346]. α [99, 2677, 678, 2559, 2467, 2636]. b [34, 840, 2719, 886, 3341, 2981, 96, 1401, 2706]. β [1387, 466, 2413, 489, 2677, 678, 530, 2559, 2190, 2467, 451, 1659]. β_{2} [1483]. c [2331]. : [214, 2202, 2096, 2942, 1870, 2210, 3074]. δ [2783, 2164, 2833, 1144, 2586, 2489]. × [2691, 2216].

-adrenergic [1483]. -adrenoceptors [1483]. -ATPase [3160, 399, 476, 804]. -D [1665]. -glucan [1659]. -K [3160]. -MSH [99]. -oestradiol [466, 2413, 489, 530]. -Oxidation [2190]. -spectrins [2467].

/K [2330].

0 [1403, 1496, 1406, 1654, 1402, 1773, 1774, 1493, 1652, 1494]. 0-12-615185-7 [1494]. 0-415-25463-9 [1406]. 0-632-05412-3 [1652]. 0-632-05733-5 [1402]. 0-632-06385-8 [1773]. 0-632-06389-0 [1496]. 0-632-06457-9 [1774]. 0-691-11297-5 [1403]. 0-85199-443-1 [1654]. 0-group [218]. 0^{o} [660].

1 [1654, 1772, 1651, 1860, 1493, 828, 1405, 1497]. 1-884125-98-0 [1493]. 1-888569-57-3 [1651]. 1-930665-74-1 [1405]. 1-930665-98-9 [1772]. 10 [1706]. 10-14 [1963]. 11-13 [2077]. 11-ketotestosterone [1387, 172, 3311]. 11^{o} [3326]. 12S [292]. 18^{o} [2623]. 1st [2077].

2 [2283, 1775]. 2* [828]. 2-7099-1490-5 [1775]. 2-7355-0501-4 [1405]. 2-phenoxyethanol [2198]. 2005 [1965]. 2006 [1706, 1989, 2100]. 2008 [3478]. 21-Trihydroxy-4-pregnen-3-one [678]. 21st [614]. 23^{o} [3326]. 27-year [236]. 28S [32]. 2nd [1772, 1493].

3 [1651, 1652]. 3-89104-668-5 [1497]. 3-89104-669-3 [1497].

4 [1068].

5 [1403, 1402, 1775, 1495]. 5S [1049]. 5th [3478].

6 [1860].

7 [1428, 1404, 1494]. 7. [1653]. 7th [2788].

8 [1773]. 85 [1495]. 85-7430-361-5 [1495].

9 [1406, 1772, 1774]. 92 [1653]. 92-807-2059-7. [1653]. 981 [1404]. 981-02-4899-7 [1404].

A. [1400]. Aals [3411]. abaster [759, 1970, 2301]. Abbottina [2431]. abdominalis [2474]. Abel [3425]. Aberdeen [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. aberrant [1799, 1253]. aberrations [3441]. abilities [202, 520, 2147, 710, 3100]. Ability [462, 583, 2192, 2877, 2932, 30, 3371, 1725, 2248, 2405, 526, 1828, 2485, 1466]. abiotic [967, 1553, 3253]. ablation [1127]. Abnormal [2649, 1848, 3206]. abnormalities [2304, 3250, 2604]. above [241, 2027]. above- [2027]. Abramis [2315]. Abrolhos [2275]. abrupt [1363]. Absence [16, 1343, 1447, 3331, 1819, 2856]. absent [2137]. absolute [1603]. absorbance [2072, 2280]. absorption [540, 2203, 210, 2407, 1794]. abstract [285]. Abstracts [924, 925, 1959, 1960, 2349, 2350]. Abundance [915, 47, 138, 162, 1863, 1683, 913, 3249, 1861, 2273, 2682, 1971, 70, 2024, 655, 703, 1953, 1291, 2067, 2662, 2418, 119, 2256, 1691, 2299, 1548, 2669, 337, 2236, 2343, 1050, 3200]. abundances [3148, 5, 1743, 675]. abundant [2167]. abuse [457]. abyssal [1065]. Academic [1494]. acanthias [2815, 3047]. Acanthocephala [449]. Acanthochromis [1281, 719]. Acanthodraco [283]. Acantholingua [2129]. Acanthopagrus [2660, 3376, 8, 3115, 3040]. Acanthuridae [1520]. Acanthurus [1520]. acanthus [2851]. acceleration [2818, 3384, 1296]. access [393, 2583, 1260, 1097]. accessibility [890]. accessory [415, 401, 2196]. acclimated [1321]. Acclimation [2437, 625, 33, 3198, 819, 1452, 626, 2614]. according [871]. account [3475, 2833]. accretion [493]. accumulation [1420, 2040, 2665]. Accuracy [457, 3315, 1631]. Acentrogobius [3468]. acetate [1555]. Acheilognathus [1831, 3309]. achieve [1944]. achne [3136]. Achoerodus [2669]. Acid [487, 930, 459, 1633, 466, 1524, 2197, 128, 442, 2419, 170, 2242, 2942, 658, 818, 426, 2917, 2083, 1794, 197, 3089, 3078, 804, 45, 1866, 367]. acid-base [1524, 128, 818, 45]. acidic [425, 1154]. acidification [656]. acidophilic [425]. acids [214, 263, 2030, 370, 3327, 469, 1472, 1556, 2028]. Acipenser [3094, 1643, 1321, 2440, 2740, 3137, 1252, 1591, 2409, 3046]. Acoustic [859, 3001, 2946, 209, 2872, 2283, 3333, 227, 1436, 1254, 3394, 3096, 1226, 2637]. acquired [174]. acquisition [75]. Acrochordonichthys [314]. acronotus [1412, 2389]. Acropomatidae [1465]. across [2717, 692, 1860, 1765, 304, 3432, 2147, 1110, 3239]. Acrossocheilus [2141, 2854]. ACTH [983]. actinopterygians [197]. Actinopterygii [2697, 1977, 2975, 2846, 2858, 2618]. action [1339, 1790, 2728]. activated [1176, 105]. activation [2044, 773]. active [1613, 1616]. activities [1054, 1488, 1015]. activity [746, 1578, 1643, 3243, 2534, 230, 3235, 445, 2897, 2611, 1555, 909, 297, 327, 498, 260, 1189, 1353, 3160, 399, 427, 3031, 103, 3189, 14, 2999, 1527, 503, 723, 3196, 1426, 1530, 1639, 2936, 1128, 2217, 3383, 3203, 2691, 2258, 3005, 1257, 48, 1065, 298, 20, 330, 964, 843, 476, 2698, 2038, 1917, 2673, 2628, 1500, 1268, 750, 709, 2831, 3336]. aculeatus [2121, 3254, 3084, 3422, 2460, 3424, 3416, 2089, 3413, 2226, 3472, 3423, 1253, 3071, 3421, 2413, 3425, 3418, 2950, 2323, 3420, 2668, 3426, 1234, 1235, 3400, 3412, 3306, 3419, 3417]. acuta [831]. acute [350, 2993, 1350, 2126, 3337, 1836, 1296, 2838, 2399, 2474]. acutus [2105]. adaptability [885]. adaptation [610, 3097, 511, 2909, 11, 1259, 463, 2476, 447, 2064]. adaptations [3418, 748, 2560]. Adaptive [753, 867, 2168, 1225, 1262, 1130, 153, 2336, 1186, 3080]. adding [1029]. addition [1694, 2611]. Additional [1628, 1521, 2921]. additive [3434]. adenomus [2911]. adhesion [431]. Adhesive [1701]. Adhesiveness [812]. adipose [3271]. adjacent [1613, 916, 2544, 605]. adjust [1188]. adjustments [231]. administration [339, 1617, 1089, 1386, 1325]. adopting [1248]. adoption [480, 1140]. adrenergic [1483]. adrenoceptor [1645]. adrenoceptor-mediated [1645]. adrenoceptors [1483]. Adriatic [509, 115, 3437, 982, 1318, 1132, 2967, 2409, 157, 452, 203, 2048]. adspersa [1684]. adult [637, 384, 953, 466, 977, 2888, 931, 2015, 1126, 2016, 2772, 560, 2627, 1514, 2042, 2764, 239, 2659, 985, 1388, 929, 3139, 323, 1391, 2320, 2688, 1125, 1365, 3184, 1148, 2172, 2556, 1573, 2236, 181, 2196, 2969, 650]. adults [2337]. Advanced [1093, 464, 1232]. Advancements [1901]. Advances [715]. Advantage [2583, 2877, 1018]. advantages [1197]. advection [1264]. Advisory [537]. Aegean [2865, 2180]. aeglefinus [2773, 2264, 2090]. aeneus [1701, 3382]. aerial [2938, 2488]. Aerobic [3189, 48, 2814]. Aeromonas [2511]. aethiopicus [2324]. aff [2505, 3459]. aff. [2232]. affect [2404, 1003, 1194, 2157, 1771, 3024, 1740, 2229, 1576]. affected [2590, 1216, 2401, 930, 1957, 1884, 2922]. affecting [2994, 1249, 1561, 1135, 2627, 1566, 3376, 2468, 706, 300]. affects [1859, 2227, 3317, 2006, 2867, 3067, 416, 2680]. affinis [552, 1370]. affinities [586]. affinity [1791]. afforested [299]. Africa [354, 334, 1383, 728, 2657, 1336, 1537, 1374, 488]. African [1579, 1242, 1512, 2020, 912, 1731, 1463, 1282, 1799, 2935, 3082, 1016, 671, 1509, 2200, 2961, 3016, 2156, 3330, 1800, 3395, 3101, 3266, 1802, 2243, 2969, 1050]. after [54, 2026, 3270, 1606, 631, 696, 1378, 818, 649, 2769, 2997, 344, 1296, 2665, 118, 2299, 150, 2205]. against [792, 2917, 1220, 472, 3476, 145]. Agassiz [3168, 2282, 1917]. Age [2642, 667, 2836, 2931, 943, 1968, 1426, 3467, 2398, 2002, 1912, 3392, 195, 2035, 2666, 1803, 80, 654, 2578, 2490, 2357, 3470, 1888, 2453, 3107, 827, 1672, 3359, 1453, 2905, 457, 1116, 1111, 729, 3386, 2670, 101, 2610, 1029, 661, 541, 1637, 3189, 845, 2037, 1460, 2164, 1290, 2565, 3248, 328, 3028, 3019, 1053, 1109, 2119, 1047, 959, 865, 2470, 1916, 36, 647, 388, 1440, 1448, 2, 1910, 2268, 257, 2604, 968, 1385]. age-0 [2610, 2037, 2, 257]. age-1 [328]. age-and-growth [1453]. Age-specific [2642]. age-structured [3248]. ageing [2617, 2727, 1770, 1366]. agents [1641]. ages [2122]. aggregata [869]. aggregated [867]. aggregates [173]. aggregation [2910, 2642, 1732, 2612, 376, 3147, 1713]. aggregation-spawning [2612]. aggregations [2227, 2877, 1785, 854, 2205, 2290]. aggregative [1389]. Aggression [735, 736, 2901, 2867, 75, 3425, 524, 811]. aggressiveness [1571, 3129]. aglomerular [476]. Agnatha [3437]. Agonistic [3333, 1390]. agonists [1483]. aguavina [106]. aiereba [3160]. aims [911]. Ain [111]. Air [1800, 2683, 286, 2582, 2570, 757, 454, 1295]. Air-breathing [1800, 2582, 2570, 757]. air-sac [1295]. akallopisos [3333]. akashiwo [117]. akysid [314]. Alarm [510, 1718, 1330, 1154, 499]. Alaska [3254, 2610, 3416, 1139, 1845, 3124, 1811, 3041, 2320, 3361, 2331, 799, 1352, 807]. albacares [1895, 3194]. albacore [195, 1755]. albisella [418]. Alboran [469]. Albula [2996, 2919]. Albulidae [2919]. Albuliformes [2919]. Alburnoides [2069, 2735, 1667, 2148]. Alburnus [3341]. albuvnoides [679]. alcian [2376]. Alegro [1495]. Alejandro [1901]. Aleknagik [3041]. Alestes [1537]. Alestidae [1537]. Aleutian [3124]. aleutica [3124]. alevins [737, 2590, 3469]. alewife [3075]. Alexander [3168]. Alexandrium [2774]. alga [753, 117]. algae [544]. algae-induced [544]. alien [607, 234, 2697]. Alimentary [1175, 2134, 2116, 785, 1619, 2158]. alizarin [3376, 1900, 3109, 501]. alkaline [2399]. all-fish [3371, 2384]. all-male [679]. allele [1601]. Alleles [828]. Allelic [2519, 249, 1941]. Allenbatrachus [1536]. allettevatus [629]. alligator [3151]. allis [2823, 3028, 2405, 2841]. allocation [2394, 146, 324, 288, 29, 1338, 669, 551]. Allometric [1156, 958, 1809, 1977, 2185, 864, 2292, 1703]. allometries [2832]. allometry [668, 2361, 742]. Alloparental [550]. allopatric [3056, 3197]. allopolyploids [2471]. Allothunnus [2498, 2906]. allotransgenic [2763]. allow [1793]. Allozyme [1936, 41, 950, 450, 276, 78, 2725, 1933, 1067, 734, 166, 275, 1932]. allozymes [160, 2147, 127]. alone [2460]. Along [3415, 3191, 944, 2362, 2445, 1682, 2640, 468, 2827, 529, 2585, 675, 2158, 1974, 3048, 2047, 3177, 3112, 1911, 2478, 2108]. Alongshore [1474]. Alopias [2174]. Alosa [2328, 2857, 3075, 2978, 3028, 2405, 2841, 2738, 52]. alpha [2982]. alpine [2625, 2844, 659, 1037, 2604]. alpinus [3290, 2932, 2198, 3460, 3045, 3029, 2909, 3372, 2726, 3370, 3128, 1854, 3058, 3056, 3197, 1324]. Alps [2319]. already [1880]. alta [1241]. alter [454, 1024, 1424]. Alteration [1068, 1799, 2095]. Alterations [746, 1295, 1915, 1853]. altered [878]. alternate [2814, 1066]. alternation [3391]. Alternative [1254, 777, 2824, 79, 326, 2014, 2295, 389, 120, 352, 1042, 475, 778, 2651, 1268]. alters [2409, 1455]. altitude [659, 3055]. altitudinal [3258]. altivelis [1754, 1914, 2531]. Altukhov [2382]. aluminium [367]. Alvarez [2804]. alveoli [1666]. Alwyne [1965]. Amarillo [1251]. amarine [277]. amarus [3211]. Amazon [2314, 1446, 2197, 1489, 2813, 3160, 2638, 2034, 747, 2232, 1096, 2707]. Amazonian [764, 2883, 2712, 3208, 1022, 2325, 3449]. amberjack [2391]. ambient [2202, 818, 2881, 2221]. ambigua [1662]. amblops [3044]. Amblyraja [2010, 2331, 2268]. amboinensis [2046, 2594]. Ambon [2594]. Ameiurus [2247, 902, 2903, 2896, 2004]. ameliorate [1958]. America [2658, 3348, 193, 1931, 2806, 7, 1796, 1286, 2750, 3419, 1495]. American [1651, 1911, 1932, 2857, 2368, 466, 233, 3231, 840, 902, 765, 190, 1394, 959, 361, 2316, 1720, 1343, 379, 1896, 3232, 2637, 1702, 2365, 1169]. americana [1736]. americanus [2443, 1609, 1561, 2826, 2869, 3200]. amino [2942, 2030, 370, 3327, 1472, 1556, 2028]. ammocoetes [1635]. Ammodytes [2295]. Ammonia [1904, 350, 3131, 3155, 995, 1536]. amoebic [219, 930, 2922]. among [2872, 412, 1848, 1557, 2956, 2857, 1111, 1488, 783, 2532, 2130, 1933, 1203, 1731, 289, 1083, 2353, 249, 465, 2793, 652, 1459, 2679, 2846, 3068, 2314, 3028, 2868, 1759, 3041, 608, 1176, 1109, 1539, 374, 734, 711, 3350, 1123, 1399, 1370, 1425, 1796, 2522, 1819, 1305, 3037, 2530, 1839, 3172, 1602, 1385, 2107, 3222, 52, 2095]. amongst [8]. Amphibious [343, 2552]. amphidromous [1115, 521, 3079, 2531]. Amphiliidae [1282]. amphioxus [3217, 2859]. amphipod [559]. Amphiprion [3304, 2497, 3333, 3433]. amplification [2517, 1698]. amplified [1754]. amplitude [869]. anabantoid [1800]. Anabarilius [3038]. Anabas [2972, 3020]. Anableps [2388]. anadromous [2461, 2340, 2383, 2670, 2347, 1689, 40, 1634, 2793, 2978, 520, 3029, 2218, 3133, 3443, 483, 1278, 579, 1728, 3140, 1462, 143, 1636, 1319, 1575, 2545, 1939, 2738]. anadromy [496, 2501]. Anaemia [1364]. anaerobic [1725, 48, 1346]. anaesthesia [1371, 3403]. Anaesthetic [2272, 255, 650, 3403]. Anaesthetization [2198]. Anal [961, 1648, 3027]. analogs [1056]. analogue [649]. analyse [3008]. analyser [3085]. analyses [1277, 2866, 2988, 91, 303, 3433, 429, 2303, 1067, 2284, 41, 627, 820, 90, 3227, 2795, 1432]. Analysis [1858, 114, 100, 2718, 1735, 294, 267, 2529, 3185, 960, 3002, 3290, 1131, 2203, 355, 1712, 759, 1758, 1933, 2714, 606, 79, 730, 3075, 1890, 1419, 2719, 3431, 588, 3289, 2291, 1754, 2960, 3047, 1591, 3019, 1254, 3392, 2150, 2528, 1539, 2246, 734, 2140, 612, 1978, 3121, 2321, 1370, 634, 2855, 564, 2057, 1819, 775, 2114, 598, 2489, 2984, 1024, 2533, 1692, 1129, 3065, 423, 2313]. Anarhichas [2062, 2998, 1737, 2058, 2791, 2186]. Anatomical [315, 2161, 3103]. anatomically [888]. anatomy [3166, 369]. ancestors [98]. ancestry [790]. Anchoa [2618]. anchor [2962]. anchor-ice [2962]. anchovy [977, 3379, 2496, 3126, 630, 2355, 1882, 2291, 3106, 1367, 2618, 1294, 1057, 2048, 3251]. ancient [1753]. Ancistrus [3429, 2832, 1996, 3329]. Anderson [1772]. Andersonia [1282]. Andes [2053]. Androgen [358, 1177, 1498, 311, 870]. androgen-inducible [311]. androgenesis [800, 1408]. androgenetic [1049, 3441]. androgens [1507]. anemonefish [2497]. anemonefishes [2951]. angelfish [3451, 3263]. angiotensin [1155]. angle [1675]. angled [3103]. angled-and-released [3103]. anglerfish [223, 2071]. anglerfishes [2439]. angling [3444, 302, 3463, 460, 2212, 1973]. angorae [1262]. Anguilla [3234, 1906, 2739, 3246, 3229, 3243, 1856, 3235, 3242, 2481, 3231, 3241, 1715, 3233, 3230, 2509, 3236, 2398, 3226, 1312, 2690, 3461, 3272, 3238, 823, 2035, 3209, 1880, 1400, 3237, 3227, 1093, 1813, 1548, 2530, 3109, 3244, 2236, 3232, 3239, 2454, 3411, 3154, 3228]. anguillicaudatus [3294, 1674]. Anguillicola [961, 563, 1312, 3244]. Anguillicoloides [3061]. Anguillid [3224, 3240, 3245, 395, 3225]. anguilliform [2696]. angustata [1869]. animal [1195]. animals [40]. animation [3318]. Anion [1016]. Anisakis [1464, 2068]. Anne [2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513]. annectens [3330]. Annecy [3089]. Announcement [834, 2464, 2255]. Announcements [39, 62, 86, 109, 135, 159, 184, 225, 247, 270, 332, 357, 381, 506, 568, 596, 644, 714]. Annual [1406, 1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2381, 2436, 2465, 2492, 632, 146, 2411, 239, 254, 955, 2596, 2513, 1888, 1738, 1162, 278, 2619, 2900, 3447, 1153, 421]. annular [3461]. annulatus [1821]. annuli [1447]. annulus [2401, 2969]. anomalies [2392, 93]. Anonymous [2864]. Anostomidae [315, 1896]. anoxia [1369]. antagonistic [1917]. antagonists [339]. Antarctic [1777, 210, 3127, 274, 1750, 277, 110, 1836, 1666, 2118, 279, 273, 3018, 476, 380, 2984, 2467]. Antarctica [443, 272, 283]. antarcticus [2386]. antenna [1847, 1540, 3212]. antennae [1847]. antennas [2374]. anterior [2332]. Anthias [2458]. Anthropogenic [607, 3422, 2473, 3297]. Anti [553, 1330, 532, 1197, 1208, 2922]. anti-Neoparamoeba [2922]. Anti-Predator [553, 1330, 532, 1197, 1208]. Antibacterial [434]. antibiotic [3130, 3071]. antibodies [2922]. antibody [1681]. anticipate [884]. anticipatory [297, 298]. antifreeze [1404, 745]. Antimicrobial [2897, 3031]. antioxidant [3153, 2611]. antioxidants [1915]. Antipredator [1551, 3302, 3072, 1587, 3181, 1565, 1393, 1967]. antoniae [3222]. Apagesoma [502]. Apareiodon [3435]. aper [2181]. Aphanius [2353, 2855]. Aphia [335, 691]. Aphyonidae [2620]. apical [2060]. apodus [2930]. Apogon [2250, 2885, 345, 813]. Apogonidae [2621, 2250, 2885]. apoptosis [691, 2358, 2293, 1896, 2729]. apparatus [415, 2676, 431, 592, 2370]. Apparent [2479, 2571]. appear [2951]. appearance [972, 2196]. Application [303, 2321, 1894, 2138, 1035, 519, 791, 1213, 366, 980, 2425, 2748, 422, 2933, 1787]. applied [957, 3323]. applying [1893, 1746, 1264, 1713]. Approach [1989, 3410, 97, 907, 3255, 1758, 1247, 2263, 272, 1949, 1746, 280, 3272, 2841, 2422, 2774, 1026, 1229, 3281, 2315, 2689]. approached [1133]. Approaches [3480, 3444, 1947, 3008]. April [1901]. Apristurus [1664]. Apure [1083]. aqua [1247]. aquaculture [1605, 3475, 1615, 1591, 2763, 3252, 1597]. aquaporins [3427]. aquaria [1238, 3203, 1693]. aquarium [2034, 2408]. Aquatic [537, 2296, 2100, 1105, 186, 3408, 2683, 3403]. AQUI [3403]. AQUI-S(R) [3403]. Arabian [3207, 3152]. aragonite [1161]. Aral [3174]. Arapaima [2883]. Arauca [1083]. archer [2488]. arches [71]. Arching [2939]. Archipelago [2910, 2275, 2851, 1253, 2961, 564, 2118]. architecture [2303, 1109, 1575]. archival [641, 3194]. archived [107]. Archocentrus [2248]. Arctic [1324, 2635, 1913, 3161, 668, 960, 3290, 1292, 2907, 158, 1038, 735, 2932, 2198, 659, 3460, 737, 2571, 1725, 1547, 3045, 651, 2635, 993, 3029, 333, 1039, 702, 1019, 2909, 75, 3372, 1931, 1539, 2726, 251, 282, 1117, 3370, 143, 1796, 232, 29, 663, 980, 1724, 2423, 2506, 3128, 1711, 1469, 2288, 760, 3058, 1522, 3056]. arcticus [1850]. Arctogadus [1039]. ardelio [1758, 2407]. Area [604, 1137, 852, 1332, 2827, 126, 1486, 2170, 1150, 611, 3278, 509, 1132, 2708, 361, 2992, 1996, 3227, 3277, 1638, 605, 754, 492]. area-restricted [126]. areas [262, 2034, 1639, 2284, 1503, 1308, 2338, 921]. argentea [3060, 2964, 3467]. argenteus [1917]. argentilineatus [2552]. Argentina [211, 106, 1922, 2501, 1034]. Argentine [3176, 3104, 1543, 1057, 2870]. arginine [1386]. Argulus [180]. argus [2847]. Argyropelecus [1413]. Argyrosomus [2287, 1723]. Arhynchobatinae [2882]. Ariake [2007, 2285]. ariakensis [3289]. arid [1908, 1105]. armatus [2666, 2063]. armour [3418]. armoured [2832, 388, 1996]. Arndt [3479]. aromatase [677]. Arousa [228]. array [1540]. arrays [1176]. arrival [1372]. artedi [1717]. Artedius [1795]. artefacts [767]. Artemia [12, 879]. arteriosus [280]. Arthur [618]. artifacts [3201]. Artificial [784, 1400, 416, 2680, 2204, 1947, 189, 2499, 636, 1563, 1957, 395, 76, 2615, 1153]. artificially [3305, 12]. aruanus [2189]. As51 [3292]. ascend [2240]. ascending [80]. Ascorbic [2197, 658, 2917, 197]. asexual [3213, 1862]. Asia [2521, 1036]. Asian [790, 653, 2613, 2759, 2797, 2853]. Asiatic [851, 1538, 789, 2278, 3099]. Aspects [115, 1404, 2766, 860, 1349, 909, 2440, 3146, 556]. asper [1086, 3316]. assay [1525, 1993, 722, 348]. assemblage [3414, 1805, 2263, 621, 546, 2093, 510, 2194, 1025, 1953, 1316, 1022, 3281, 3192]. assemblages [1487, 699, 1085, 1632, 2994, 2777, 2404, 2813, 1732, 2827, 2007, 397, 863, 1735, 2276, 1096, 917, 1974, 1909, 364, 2701, 822, 2680]. assess [1852, 1478, 2946, 1879, 1723]. assessed [1489]. Assessing [1937, 1689, 2222, 820, 1760, 158, 42, 2273, 2414, 3365]. Assessment [2739, 3359, 317, 1054, 2319, 2004, 2266, 3014, 3257, 3336, 973, 2714, 3075, 272, 1882, 519, 1769, 633, 2986, 366, 258, 364, 2489, 1542, 1424, 2529, 2568, 1388]. assessments [1889]. Assimilation [685, 1019]. Asso [2444]. associated [2622, 3148, 764, 1947, 3424, 196, 1159, 3100, 3402, 1735, 776, 3209, 1380, 926, 3384, 1519, 778, 2734, 801, 3090]. Association [532, 1434, 2664, 1604, 2679, 2006, 797, 1235]. Associations [397, 2276, 2757, 1712, 1936, 2495, 2606, 2091, 1142]. assortative [2768, 782]. assumption [587]. assumptions [2930]. Asturias [1937]. Astyanax [2486, 3043, 3292, 3439]. asymmetric [2721]. asymmetrical [3118, 1979]. asymmetries [694]. Asymmetry [1373, 167, 2608, 2770, 3462, 75, 2959, 487, 526, 2604]. asymptotic [64]. Asystasia [1463]. atavistic [1799]. Atherina [808, 2987]. Atherinella [2737]. Atherinidae [1497, 1684, 599]. Atheriniformes [2987, 2737]. Atherinomorpha [3267]. Atherinops [552, 1370]. Atherinopsidae [1497, 552, 1370]. Atherton [375]. athletic [2582]. Atkinson [1406]. Atlantic [15, 1290, 2174, 2961, 557, 3364, 2106, 57, 2907, 424, 412, 1195, 1946, 34, 188, 2055, 620, 2011, 637, 2265, 1031, 3193, 1398, 1891, 1748, 2227, 3191, 2771, 1023, 2043, 1937, 3229, 1983, 2911, 2934, 2642, 3069, 1127, 844, 1192, 1108, 1513, 1171, 2897, 1559, 3465, 2543, 2532, 1481, 979, 3000, 187, 1935, 1933, 712, 3450, 1332, 1560, 302, 894, 2974, 3389, 177, 1248, 2924, 1382, 829, 2551, 1733, 2367, 260, 1793, 1984, 2579, 692, 2572, 1207, 1584, 1646, 2157, 2392, 2275, 2770, 106, 2836, 1851, 1634, 92, 1422, 2620, 893, 3062, 859, 1890, 3469, 2060, 2115]. Atlantic [1810, 1765, 652, 1459, 1504, 2899, 148, 2719, 3189, 700, 3328, 1562, 1861, 623, 2635, 1288, 394, 749, 503, 223, 2016, 2031, 587, 1612, 3066, 2682, 3202, 1426, 1642, 2663, 850, 2126, 1530, 1997, 25, 974, 1290, 2627, 2565, 3443, 1212, 1936, 1138, 1434, 1376, 2083, 3131, 2858, 3181, 190, 1062, 2366, 641, 1690, 611, 2153, 66, 58, 3261, 1669, 1673, 1502, 1590, 3019, 933, 929, 508, 2181, 1520, 703, 1468, 393, 1377, 1931, 1744, 1566, 1464, 885, 154, 950, 2791, 1567, 617, 633, 1789, 2428, 2427, 1918, 627, 2650, 1930, 1333, 1940, 1938, 53, 681]. Atlantic [1835, 1755, 3081, 2895, 3285, 439, 994, 80, 485, 2240, 2067, 3121, 2986, 1568, 1179, 140, 2923, 831, 2882, 516, 1508, 1425, 2593, 416, 2688, 1310, 1569, 2648, 344, 1361, 2838, 2586, 2605, 2963, 2567, 244, 316, 219, 930, 3237, 288, 2109, 1819, 1686, 1691, 1178, 253, 1448, 2500, 1421, 2172, 976, 1145, 1510, 843, 463, 1226, 2802, 35, 43, 287, 1228, 1033, 1808, 2860, 2220, 3129, 870, 2819, 2489, 1934, 2539, 1596, 603, 2609, 879, 2269, 69, 1124, 464, 2228, 2628, 1598, 1021, 2918, 3186, 2962, 2835, 1910, 2190, 811, 1709, 815, 181]. Atlantic [1057, 1911, 2545, 206, 3083, 1602, 1130, 1941, 1939, 1929, 1932, 2922, 920, 1747, 1012, 1416, 1500, 2685, 3374, 493, 2519, 1556, 2379, 1881, 367, 2180, 486, 1396, 1586, 507, 841]. atlanticus [2543, 725]. Atlanto [50]. Atlantoraja [2882]. atlas [627, 3322]. Atol [2079]. Atoll [3308]. ATP [301]. ATPase [3160, 399, 2330, 1860, 476, 804]. Atractosteus [3036]. atresia [24, 2834, 253]. atrial [447]. atripectoralis [2796]. atromaculatus [1968]. attached [1003, 533]. attaching [1293]. attachment [3002, 2832, 2045]. attachments [826]. attack [186, 1024, 1675]. Attraction [806, 384, 1947, 901, 2015]. attractive [2241]. attractiveness [1188]. attracts [220]. attributes [1978]. attributing [1947]. aucha [1438]. Auchenipteridae [2370]. August [2424]. AULA [1497]. AULA-Verlag [1497]. Aulopiformes [3326]. aurata [1604, 1608, 2781, 1650, 3223, 2170, 1919, 3383, 3150, 2965, 2602]. auratus [2983, 2861, 1692, 3305, 2694, 2459, 1867, 3011, 3275, 1492]. aureus [2691]. aurita [2657]. auritus [3205]. Australia [1453, 1347, 2098, 1684, 2052, 3286, 2147, 375, 3446, 2964, 2968, 1110, 3375, 1662, 3119, 3182, 3447, 2468, 2144, 2530, 1306, 294, 2357, 2231, 1385, 2107, 762]. Australian [2092, 3017, 2261, 462, 839, 3260, 2630, 1663, 3463, 2660, 3446, 3021, 3467, 1550, 2766, 599, 2795, 1801, 2839, 2856, 2378, 2957, 1061, 2108, 2519]. australiensis [2839]. australis [3379, 3446, 823, 742, 3119, 2530, 3403]. austriacus [3373]. Austrian [956, 2223, 609]. Austrolebias [1843]. Austrolethops [2851]. autapomorphies [458]. autochthonous [2223, 896]. Automated [1129]. Autoradiographic [2407]. Autotransgenic [2763]. Autotriploid [2617]. autumn [1025, 1831, 2767, 3284, 2567, 3204]. autumn-spawning [1831]. Autumnal [2999]. autumnalis [616]. availability [2877, 500, 2387, 1952, 2772, 518, 1304, 1193, 2873, 2354, 948, 1028, 1130, 1881]. Aveiro [918]. avoid [462, 1892, 3106]. avoidance [1172, 953, 555, 1443, 161, 3059, 271, 738, 1967]. avoided [1563]. Awassa [2929]. Awe [2317]. axelrodi [2676, 2712, 2437]. axes [536]. axis [1283]. ayu [324, 396, 1914, 2531]. Azorean [252]. Azores [1521, 1055].

B [3321, 1652, 3406, 3407, 3320, 355, 1914, 1490]. Back [536, 572]. back-calculated [572]. Back-calculation [536]. backcalculation [2490]. backcalculations [1029]. backed [1336]. Background [495]. Backiel [3006]. backpack [171]. backscatter [227]. bacteria [3130]. bacterin [2511]. bad [2460]. baeri [2740]. Baffin [3372]. baganensis [123]. bagrid [2201]. Bagridae [2135]. Bahamas [1291]. Bahía [2562]. Baikal [705]. bait [1024]. baited [1542]. balance [818, 2815]. Balancing [3476, 1220]. Balaton [727]. Balitoridae [3471]. Balkan [1318]. Balkanian [2779, 3199]. Balkans [2544]. Baltic [2507, 746, 2424, 1863, 387, 429, 1997, 1434, 611, 112, 2834, 496, 3461, 1549, 938, 3285, 2986, 1074, 955, 2068, 2057, 287, 3244, 945, 1602, 903, 709, 2831]. bamboo [2864, 1853]. band [3447]. band-pair [3447]. banded [2326, 384, 953, 465]. banding [534]. Banggai [1188]. Bangladesh [74, 2267, 2574]. Bank [3364, 1686, 2587, 667, 241, 2587, 2495]. barb [2929, 1799]. Barbadian [222]. barbatus [3073]. barbel [1616, 3026, 1576, 512]. barbels [1784]. Barbonymus [2850]. barbouri [1088]. barbs [2263, 1408]. barbus [1203, 607, 582, 1204, 2929, 1799, 1616, 2547, 3118, 1509, 3026, 1105]. barcode [3114]. barcodes [3116]. barcoding [2798, 2797]. bare [1672, 2484]. Barents [2877, 2295, 2162, 968]. barges [3404]. Baringo [1998, 2324, 2527]. barium [3268]. Barkley [847]. Barnard [2650]. baro [142]. baro- [142]. barometric [1159]. barracuda [2224, 2224]. barramundi [3093, 165, 2182, 2797]. barred [513, 1110, 2141]. barrelfish [2836]. Barrier [947, 2052, 3050, 1785, 1699, 65, 1011, 762, 2481, 2027, 1427, 2325]. barriers [2856, 3222]. basal [2975, 2498, 2906]. base [1524, 128, 818, 930, 45]. based [3185, 200, 2387, 2735, 1194, 1206, 1525, 2616, 931, 1029, 3171, 830, 1419, 1335, 886, 2469, 1186, 723, 1529, 2342, 3341, 3047, 1753, 1070, 3392, 292, 166, 2152, 127, 1122, 3237, 1401, 1393, 1957, 1229, 2540, 1294, 1601, 2706, 1235, 1267, 2689, 2238, 2724, 1471, 2869, 2331]. Baseline [364]. bases [1247]. Basilichthys [742]. Basin [3286, 3214, 3366, 600, 1137, 234, 1833, 2647, 1667, 2472, 156, 3401, 2525, 2083, 1514, 2528, 2718, 1537, 2522, 1326, 3015, 3459, 1846, 1802, 2732, 3160, 790, 3043, 1896, 3177]. basins [2735, 2638, 1860, 3454]. basis [104, 2057]. basket [1993]. basophils [2173]. bass [2396, 2910, 263, 97, 2956, 2271, 311, 900, 555, 2502, 2387, 1608, 2270, 327, 610, 1003, 1041, 928, 291, 795, 3457, 1056, 2523, 2373, 46, 1663, 3463, 1378, 2007, 701, 3183, 2607, 1565, 649, 374, 3391, 28, 2111, 517, 999, 2216, 417, 748, 2795, 2354, 2212, 2824, 3362, 1973, 1354, 2533, 2797, 318, 670, 2313]. bassensis [2689]. Batasio [2135]. Batch [1723, 3315, 2259, 3306, 421]. batch-spawning [3306]. batches [1315, 1780]. Batesian [3454]. bathymetric [3284]. bathypelagic [1685, 2362, 2445]. bathypevimensis [889]. Bathyraja [1682, 3124, 3105, 2649, 2984, 2331, 464]. Batoidea [1745, 2585, 2503]. batrachoid [415]. Batrachoididae [415]. batrachus [214]. Bavaria [304]. BAY [1405, 2941, 2098, 3132, 2199, 2369, 2284, 1909, 35, 1348, 3115, 621, 2846, 2397, 2618, 1013, 3126, 2510, 2355, 2258, 2259, 1736, 2285, 2937, 2377, 1317, 3347]. Bay-scale [35]. Bayesian [2795, 1601]. BC [2812]. be [179, 1563, 1711, 3426]. Beach [2931, 445, 3267]. Bean [2876, 3124, 2348, 2867]. Beard [3478]. beat [2165, 869, 991]. beaugregory [95, 1225, 1001]. beautiful [972]. Beaver [1777, 1750]. bed [1743, 1458]. beds [522, 2851, 2284, 1635]. been [2744]. before [1697, 2504, 696, 1159, 1093]. beginning [614]. Behaviour [1647, 3137, 518, 763, 3094, 1859, 1172, 528, 1195, 1238, 1239, 3148, 525, 192, 583, 2972, 932, 2600, 2043, 1240, 2121, 844, 1192, 3141, 3235, 445, 965, 482, 2812, 544, 2780, 1003, 1889, 532, 3389, 1248, 55, 1194, 784, 1207, 1646, 1975, 1342, 3424, 859, 1667, 773, 1091, 2484, 2030, 2006, 1186, 3024, 1209, 2184, 1564, 673, 426, 553, 1435, 1822, 1063, 3299, 3181, 1214, 3180, 3472, 1669, 1919, 794, 2297, 1252, 3139, 205, 1254, 1565, 1255, 671, 265, 2664, 2200, 1285, 3391, 3150, 2570, 926, 2232, 3142, 3067, 343, 439, 1256, 516, 252, 416, 2063]. behaviour [1741, 797, 2648, 142, 2216, 1800, 1224, 663, 1510, 2562, 1226, 1887, 1393, 1228, 3147, 2301, 1229, 1182, 3090, 2556, 2695, 1551, 3186, 124, 1232, 2475, 1986, 121, 3009, 2545, 2584, 1454, 1266, 76, 1130, 3426, 52, 1268, 400, 3194, 1119, 1270, 3320]. Behavioural [2970, 1501, 1038, 404, 3368, 1826, 3021, 941, 1275, 3355, 1080, 1503, 1231, 3204, 1233, 1349, 1244, 307, 1610, 104, 3413, 1077, 2126, 1002, 1254, 2817, 988, 2111, 3184, 411, 3178, 1432, 2343, 2568]. behaviours [1792, 2226, 2764, 3402, 3252, 1367, 1390]. behind [3051]. Behnke [1358]. Being [1220, 2045, 1018]. belcheri [3217]. Belgium [2752, 1849]. Belize [1785]. bellied [350, 2638, 1337, 2474]. belly [2995, 1221]. below [2027]. below-dam [2027]. beneath [2205]. Benedenia [856]. Benefits [2026, 1201, 2571, 1770, 1220, 1416]. Benguela [3106, 2537]. Beni [1974]. Benin [354]. Bennett [2864]. bent [1223]. Benth [1463]. benthic [682, 12, 1043, 913, 242, 2826, 403, 138]. benthicola [2132]. benthivory [2260]. Benthochromis [2811]. benthopelagic [502]. Bering [729, 1682, 2640, 3105, 1352]. Bermuda [604]. bernacchii [2467]. Berounka [2556]. Bertalanffy [1746]. Beta [1858, 3392]. Bethesda [1651]. Betta [2138]. better [1217, 1229, 1967]. between [1292, 600, 2721, 1558, 1704, 2084, 1339, 1908, 1239, 569, 2675, 3168, 3206, 860, 1534, 900, 2678, 2897, 500, 2667, 1242, 1923, 233, 1162, 2902, 302, 2756, 532, 286, 862, 1204, 349, 234, 1014, 2510, 805, 1584, 256, 2213, 3326, 2052, 1773, 893, 2061, 1126, 1020, 103, 174, 575, 1667, 3045, 541, 290, 1683, 529, 2484, 2332, 1663, 3050, 2099, 429, 1563, 1861, 3464, 2657, 1209, 1037, 3357, 1460, 2845, 791, 2226, 1435, 1081, 1588, 1743, 1589, 1434, 1040, 2027, 3021, 2858, 3122, 2909, 1067, 403, 397, 2143, 1590, 2690, 508, 2005, 1075, 1254, 1565, 717]. between [916, 2718, 3355, 599, 3445, 3349, 1618, 2441, 1333, 2276, 536, 579, 705, 1755, 1080, 363, 550, 2040, 1400, 2421, 1569, 1717, 2662, 169, 1595, 2954, 1096, 2605, 3387, 2152, 9, 2429, 2794, 2415, 1450, 896, 1343, 2802, 598, 2602, 2955, 804, 1808, 2038, 1934, 2375, 879, 2269, 2628, 1598, 1231, 2835, 201, 422, 3115, 782, 137, 1601, 1289, 697, 1233, 2732, 1522, 2564, 406, 1119, 493, 2315, 2133, 2958, 1099, 971, 3346]. between-morph [2564]. between-population [1292]. bi [937, 2402, 2494]. bi-directional [937, 2402, 2494]. biaculeatus [3151]. Bianchi [3410]. Biandantang [207]. bias [3069, 2833, 1678]. biased [2710, 2679, 1211, 2094, 514]. bicarbonate [720]. bicolor [2398, 1093]. bicoloratus [2476, 1099]. bidirectional [3311]. Bidyanus [2378]. bifasciatum [1274]. Big [972, 350, 1337]. big-bellied [1337]. bigeye [689, 3044, 2105, 1410, 3334, 240]. bighead [2457, 3221]. Bight [346, 3200]. biguttatus [3380]. Bile [342, 1446, 341, 2509, 368]. biliverdin [2120]. Bimini [1291]. bimodal [960, 2249]. binding [1555, 2728, 1791]. bioassay [1002]. Biochemical [1349, 3198, 2480, 390, 3177, 2180, 2387, 839, 1905, 3369, 2581, 1853, 1101]. bioconcentration [2413]. biodiversity [608]. bioelectrogenesis [410]. bioenergetic [2892]. Bioenergetics [2864, 957, 216, 693, 2422, 1125, 2933]. biogeographic [2731, 586]. biogeographical [245]. biogeography [3435, 2540]. Bioinformatics [3323]. Biological [909, 2079, 635, 1493, 3480, 1901, 2934, 179, 1732, 3217, 3339, 419, 9, 3369, 471, 1656, 3313, 3154]. biologically [82]. Biology [1406, 2279, 2309, 2381, 2436, 2465, 2492, 1404, 193, 2411, 1115, 1652, 695, 2513, 2800, 786, 1168, 952, 211, 1412, 1682, 3152, 2694, 394, 2105, 3401, 3413, 2964, 2671, 509, 1593, 1468, 2666, 1055, 1410, 2612, 2953, 1351, 1337, 2871, 461, 1026, 1021, 1630, 2672, 2503, 1991, 1493]. Biology/An [1406]. bioluminescence [2995]. biomanipulation [963]. biomarkers [2920, 3089, 2591]. Biomass [912, 1165, 65, 3379, 215]. Biomechanics [1403]. Biometric [759]. biometry [218, 3123]. biopsy [63]. Biosynthesis [197]. biosynthetic [983]. Biotic [1553, 967, 3253]. biparental [2248, 266]. biphenyl [2873]. Biplot [79]. bipolarity [1882]. bipunctatus [2069]. bird [3458]. Birds [1773]. birth [1803]. Biscay [1013, 3126, 2510, 2355, 2258, 2259]. bisexual [2983]. Bitterling [1199, 1831, 2168, 2008, 3211, 1184, 1222, 336, 3309, 1259, 2415, 1230, 2695]. bitterlings [292]. bivalves [126]. Biwa [1753]. black [2271, 2394, 2660, 3025, 2903, 2896, 1336, 3268, 3376, 2301]. black-backed [1336]. black-spot [3025]. black-striped [2301]. Blackburn [1772, 1405]. blackeye [1699]. blackmouth [2543]. blacknose [2389]. blackspot [3149, 1618, 1622, 1623, 3091]. blackspotted [3210]. blacktail [1342]. blacktip [2941, 3003, 1159, 3308]. blackwaters [3160]. Blackwell [1496, 1402, 1774, 1652]. blacodes [1477]. bladder [1817, 347]. blastulas [1112]. Bläzka [84]. Bläzka-type [84]. bleaching [2951]. bleaks [3341]. bled [721]. Bleeker [2185, 2850, 314, 2377]. blennies [156, 3, 65]. blenniid [1476, 1250, 475, 2163]. Blenniidae [1497, 1733, 3]. blennioid [1733, 1735]. blennioids [3070]. blenny [2535, 1733, 2554, 694, 773, 1091, 2297, 3396, 343, 252, 358, 3147, 3027, 2768, 1431, 2444, 1442]. Blepsias [1411]. blind [1680, 2486]. Bloch [2451, 2983, 1915, 2826, 1980, 2784, 2292, 2274, 2861, 2505, 2182, 2144]. block [3083, 2692]. Blood [2188, 2433, 1276, 2198, 1555, 3085, 535, 2714, 707, 2302, 2993, 722, 1377, 2266, 999, 1033, 3319, 367]. Blood-chemistry [2188]. bloom [117]. Blue [3010, 153, 814, 2796, 2120, 538, 394, 241, 2304, 3182, 3447, 2669, 2376, 1451, 3276]. blue-eye [1451]. blue-spotted [3182, 3447]. blue-throated [2120]. bluebanded [2494]. bluefin [2599, 1058, 2616, 3381, 633, 2119, 640, 2113, 348, 2584, 3353, 3356]. Bluegill [2456, 962, 1423, 3317, 1378, 3337, 2354, 406]. bluespotted [129, 383, 3009]. bluestripe [3065]. boarfish [2181]. boat [1247]. boat- [1247]. bocagei [3026]. bodied [3003, 3249]. bodies [84, 8]. Body [1852, 1414, 248, 3418, 1123, 463, 1140, 1888, 954, 200, 1038, 2608, 495, 523, 2558, 2283, 2678, 3351, 729, 532, 1382, 191, 89, 522, 2899, 803, 198, 969, 1040, 305, 85, 3311, 655, 3013, 2257, 1327, 2437, 2961, 252, 1190, 1708, 376, 2229, 781, 2109, 843, 2955, 8, 1445, 1623, 1466, 2, 1338, 2477, 1675, 697, 2939, 1747, 2224, 3080, 1368, 365, 841]. body-shape [2961]. boenak [1505]. bogaraveo [3149, 1618, 1622, 1623, 3091]. BOL [3114]. bold [1004]. boldness [2678, 1269]. Boleophthalmus [2809]. Bolivia [1974, 2314]. Bolivian [1489, 2638, 2232, 1022]. bomb [1116, 1366]. Bonaparte [2624]. bonapartii [2408]. bonapavtii [703]. bonariensis [2430, 2677]. bonasus [2641]. bone [2060, 213, 2490]. bonefish [2996, 36, 590, 2919]. bones [627]. bonito [147, 450, 2180]. bony [2975, 572, 556]. Book [2017, 3479, 1403, 21, 37, 38, 108, 133, 134, 183, 224, 269, 331, 407, 435, 455, 504, 643, 713, 766, 875, 876, 904, 905, 1010, 1094, 1496, 1497, 1406, 1654, 2175, 2252, 2307, 2434, 2575, 2652, 2785, 2928, 2976, 3032, 3143, 3163, 3410, 1875, 1555, 1301, 1402, 1772, 1924, 3323, 1651, 1404, 3322, 1298, 1773, 1774, 1134, 1493, 1775, 3478, 3405, 3321, 3311, 236, 1665, 1926, 1925, 1299, 1405, 1652, 3406, 1495, 1494, 3407, 1300, 3480, 3408, 1653, 3320, 3411, 2074, 3409]. Books [408, 1095, 1302, 1776, 1876, 1927, 2018, 2075, 2176, 2253, 2308, 2435, 2576, 2653, 2786, 1773]. borchgrevinki [1060, 273]. boreal [2021, 2999, 2843]. Borne [3170]. boscii [2244]. Bosnia [2522]. Boston [1824]. both [1761, 447]. Bothnian [287]. bottlenecks [1322]. Bottom [1265, 307, 2039, 2493, 917, 921]. bottom-dwelling [307]. bottom-feeding [2493]. Boulenger [2910, 1802]. boundaries [2796, 2147, 95]. bourgeois [1254]. bovine [1681]. boyeri [808, 2987]. brachiusculus [1092]. Brachyplatystoma [1489, 3449]. Brachyrhaphis [2678]. brachysoma [2201]. brackish [540, 2346, 2082, 329, 2246, 2479, 2485, 669]. Brain [2005, 746, 2073, 191, 18, 426, 390, 1386, 2498]. Braithwaite [3320]. Brama [2088, 2315]. Branchial [128, 1814, 71, 2437]. branchialis [2257]. Branchiostoma [3217]. Brandt [2740]. brandtii [3401]. brasilianus [2870, 3176]. brasiliensis [2934, 733, 899, 2737]. Brazil [546, 1495, 2800, 1137, 2079, 2827, 2472, 2846, 3401, 3019, 1520, 2174, 2304, 2528, 1123, 2329, 2169, 1996, 1326, 364, 2618, 3436, 2737, 201, 3454]. Brazilian [2800, 908, 2934, 2388, 1087, 909, 3326, 2166, 3048, 2870, 3176]. breadth [500, 538]. break [539]. break-up [539]. Breakdown [1558]. breaksea [2666]. bream [1054, 1112, 2781, 1015, 326, 465, 2041, 529, 658, 2660, 3275, 2170, 23, 2949, 173, 3150, 1055, 3376, 2393, 574, 31, 983, 2602, 1675, 1981, 2211, 2315, 3257]. breams [401, 586]. breathing [2582, 2570, 757, 1800, 2683]. bred [2408]. breeders [1611]. Breeding [2297, 2232, 2825, 1511, 91, 1162, 2849, 2631, 3401, 1898, 358, 1177, 1170, 1432, 313, 3380]. Brest [1775]. breviceps [2163, 1827]. Brevimyrus [2570]. brevipinnis [3299]. brevirostris [2079, 3283]. brevirostrum [1643, 2440, 3137]. Brevoortia [2721, 2526, 2770]. Brief [559, 556]. briefly [3285]. bristlenose [3329]. Britain [2247]. British [1628, 1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2381, 2436, 2465, 2492, 847, 2411, 3184, 2513, 1, 1456, 180, 3175, 1936, 2686, 3334, 598]. britoi [2009]. Brittany [3077, 461]. broad [1983, 303]. broadstripe [222]. brood [888, 480, 1053, 3303]. brooder [2163]. broodfish [2590]. brooding [2044, 2970, 2020, 3398, 705, 797, 2873, 2768]. broodline [653]. Broods [2241, 3047, 514]. broodstock [1559, 1618]. brook [966, 484, 525, 2807, 2905, 3437, 1111, 2667, 2383, 30, 442, 1243, 2756, 1689, 6, 861, 1999, 1104, 806, 1333, 130, 1140]. brotula [2912]. Broussonet [2683]. Brown [2080, 1687, 600, 625, 1172, 1558, 1320, 1577, 2131, 1840, 78, 736, 1240, 855, 687, 268, 1457, 1111, 2667, 2710, 2319, 3170, 1581, 535, 712, 2756, 532, 1825, 682, 2551, 956, 2317, 2499, 77, 299, 2893, 1634, 636, 1461, 3097, 2590, 382, 103, 575, 1375, 3179, 1197, 3174, 623, 2635, 656, 942, 2273, 2371, 1037, 1363, 1971, 2032, 2810, 164, 2218, 3133, 3443, 3248, 2083, 3122, 483, 1921, 236, 2004, 949, 559, 828, 2718, 626, 612, 460, 676, 978, 3067, 1143, 113, 970, 585, 581, 1433, 3258, 2001, 1296, 2953, 1103, 2589, 744, 1260]. brown [1636, 1272, 2003, 2835, 1455, 2012, 1535, 1846, 1319, 3332, 2843, 185, 76, 892, 3336, 2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513]. Brownell [157]. bruneus [3108]. Brunnich [1623]. Bruno [1654]. Brycon [1353, 1077, 1714]. bryosalmonae [3055]. Bubble [391]. budget [2384]. budgets [1104]. buds [1476, 674]. buffer [1236]. buffering [891]. Bufo [462]. building [2950]. bulb [1556]. bulbus [280]. bulldog [2243]. bullhead [1888, 2453, 639, 1037, 1427, 2903, 2896, 2004, 3369, 131]. bullheads [2506]. bullies [2204]. bully [1100, 1330]. Bump [1010]. bundles [1696]. Bunodera [2262]. buoyancy [3418]. burbot [1152, 1285, 874, 1125, 1730, 3098, 706, 774, 2365]. burden [2449]. Bureau [3321]. buried [126]. Burn [2481]. Burrishoole [1171]. burrow [300]. burrows [1906, 2851]. Burst [2579, 3301, 2681]. burtoni [2020]. Bush [302]. butcheri [2660, 3376, 8]. butterfly [3446]. butterflyfish [3373]. butterflyfishes [3350, 2256, 3022, 1224]. by-catch [2766]. Bythitidae [2912].

C [3479, 1404, 3405, 1405, 1495, 3320, 2555, 2783, 2623, 658, 2833, 1144, 2586, 660, 2489, 2155, 1149]. c-myc [1149]. C-start [2155]. C. [44]. Ca [2202, 1044, 2210, 3074]. CABI [1654]. cabrilla [374]. Cadmium [527, 2203, 155, 2945]. cadmium-supplemented [2203]. caeca [2909, 17]. caeruleum [2165, 3431, 2794]. Cage [1654, 2205]. cages [309, 416]. Cain [3425]. Cais [1909]. calabaricus [2975]. calcarifer [3093, 2182, 2797]. calcification [2489]. calcified [2645]. calcium [2203, 1617, 2437, 2238]. calcium- [2203]. calculated [572]. calculation [536]. Caldwell [1772, 1405]. Caledonia [1712]. Calibration [3085, 830]. California [1975, 1487, 1488, 3307, 864, 1280, 2931, 2369, 3216, 3267, 2953, 36, 2919, 1812, 2230, 2875, 638]. Californian [1518]. callarias [2834]. Callichthyidae [1701, 3382, 2370]. Callionymidae [1742]. callipterus [238]. calls [2991]. calpain [1624]. calpastatin [1624]. calpastatin/calpain [1624]. Camallanus [3061]. Camargue [2035, 731]. camera [1665]. camouflage [2995]. campaign [3114]. Campbell [2812]. campechanus [2022]. Campeche [1989, 1671]. Can [2207, 2446, 431, 2973, 2746, 1196, 884, 1563, 1188, 2806, 1816, 3018, 3363, 2745]. Canada [2812, 651, 403, 3372, 3280, 598, 35, 2637]. Canadian [2360, 854, 2625]. canadum [2400, 2684]. canal [882, 2134, 2116, 785, 1897]. Canaria [642]. Canary [642]. Canestrini [3402, 433]. canestrinii [3402]. canicula [549, 2888, 2330, 2878, 2993, 2385, 1090, 2483, 1721]. Cannibalism [1711, 216, 217, 1053, 1225, 1001, 2187, 1313, 2162, 813, 492]. cannibals [849, 1711]. cannulation [2572]. Cantabrian [3364, 2483, 1721]. Cantor [2847, 2606]. canyon [2362, 2445, 2894]. canyon-shaped [2362, 2445, 2894]. Cap [1079]. capability [102]. capacities [3460]. Capacity [413, 1414, 3060, 1725, 951, 2437, 891, 48, 2814, 983, 2190, 2668]. Cape [541, 1336, 2961, 2537]. capelin [2877, 2295, 271, 2162]. capensis [1284, 2537]. Capros [2181]. capsalid [856]. capsule [2487]. capsules [2808, 464]. Captive [3451, 2631, 1559, 1699, 3247, 1620, 383, 640, 1170]. captive-raised [640]. captive-reared [3247]. captivity [1821, 2487, 2408, 2769, 1365, 2280]. capture [2996, 168, 2607, 3474, 3476, 2425, 826, 1454, 1806]. captured [2800, 2011, 1647, 1504, 3142, 1354]. captures [223]. capturing [257]. Carangidae [2988]. carapid [369]. Carassius [3305, 2915, 1867, 3011, 2983, 2525, 2784, 2861, 3219, 1692]. carbohydrases [724]. carbohydrate [191, 2116]. carbon [764, 927, 2493, 2833, 2473, 3089, 279, 754]. carbonate [780]. carcass [2214]. Carcharhinidae [2304, 2672]. Carcharhiniformes [2672]. Carcharhinus [2941, 3003, 1412, 1803, 3308, 2389]. carcharias [2468]. Carcharodon [2468]. cardiac [317, 2623, 626, 1362]. cardinal [2712, 934, 2437]. cardinalfish [2885, 1188, 345, 2479]. cardinalfishes [265]. Cardiorespiratory [625]. Cardiovascular [438]. care [1792, 1181, 2241, 1003, 1752, 1227, 1482, 550, 1795, 2163]. Careproctus [44]. Caribbean [1639, 1115, 2397, 3117, 750, 1236]. carnivores [1488]. Carolina [2420, 1412, 2730, 3479]. carotene [2611]. Carotenoid [200, 2051]. Carotenoid-based [200]. carotenoids [3466]. Carothers [3405]. carp [3366, 2092, 2374, 132, 2457, 2758, 539, 199, 170, 3371, 2915, 790, 2716, 2384, 259, 2327, 3286, 2983, 572, 2632, 3327, 3221, 3100, 490, 2245, 3355, 2480, 981, 2861, 20, 330, 3252, 118, 437, 1572, 1306, 45, 208, 447, 2681, 2636, 3080, 365, 3438, 3346, 231]. Carpathian [2547]. carpathicus [2547, 3118]. carpenteri [244]. carpio [2758, 199, 3371, 3286, 2632, 3327, 3100, 1753, 2245, 3355, 2480, 3252, 2681, 2636, 3080, 3346]. carps [2759, 3219, 1839]. cartilage [2060]. Caruso [306]. Carvão [1909]. case [2789, 1847, 3076, 1605, 1041, 2751, 2300, 3124, 2644, 612, 1184, 1160, 980, 1958, 3342, 3257, 3454]. cases [2782, 831]. Caspases [3146, 2729]. Caspian [3336]. caspius [3153, 3336]. catabolic [2728]. catadromy [3240]. Cataetyx [2912]. Catak [1846]. Catalina [1975]. Catch [2672, 302, 2766, 80, 1678, 1576]. catchability [531, 1264]. catches [1663, 1146, 857]. catchment [3234, 299, 3431, 2069]. catchments [1385, 2107]. catecholamines [279, 398]. catfish [2356, 2247, 1242, 1512, 2053, 2658, 721, 707, 2944, 902, 2832, 3010, 2935, 1378, 2613, 1394, 671, 2200, 2370, 3016, 711, 542, 314, 491, 266, 388, 3015, 3262, 2354, 82, 2556, 1325, 2201, 2581, 3449, 2969, 3329, 32, 356]. catfishes [2134, 1489, 3429, 1083, 686, 1282, 897, 1036, 550, 127, 1996, 201]. Catla [2327]. Católica [1495]. catostomid [2540]. Catostomidae [886]. catostomids [2420]. Catostomus [530]. catshark [402, 2330, 2878, 47]. catsharks [2543]. catus [2247]. caudal [534, 1797, 861, 2051, 2206, 2628]. caught [1821, 1608, 1762, 566, 2672, 2503, 2831]. Caulerpa [753, 1275]. cause [1245, 1734, 1492, 60]. caused [1060, 1068]. Causes [2577, 1214, 2097, 9]. cave [2153, 2648, 1996, 1781]. cave-dwelling [2153, 2648]. cavefish [1088]. Cavtat [2255]. Cay [1846]. Cayman [2225]. cDNA [132]. cease [3230]. Cell [2977, 263, 1848, 1112, 2980, 2358, 1060, 707, 3430, 3033, 1737, 1817, 3123, 1179, 2593, 1369, 2293, 1742, 3027, 660, 60, 454, 2318, 3046]. cells [362, 752, 1513, 2611, 2599, 1321, 1681, 2145, 535, 1068, 2330, 2195, 3328, 23, 1027, 1757, 1438, 2385, 547, 1473, 745, 2542, 825, 1090, 2312, 344, 573, 744, 2433, 3319, 810]. Cellular [325, 2593]. Celtic [2952, 1193]. census [3023, 1909, 561]. Central [1137, 2507, 1495, 3107, 883, 115, 607, 2139, 1457, 1058, 1733, 751, 1139, 1845, 1698, 2082, 1516, 3211, 236, 1753, 916, 1719, 1549, 81, 542, 1399, 695, 1065, 1326, 2871, 1812, 2144, 1386, 3064, 1352, 2230, 1358, 2088, 1782, 101, 1702]. centrarchid [2747]. Centrarchidae [3205]. centre [2936]. centred [3444]. centres [1639, 2665, 2196]. centromere [3293]. Centrophorus [667, 2904, 264]. Centropomidae [3048]. Centropomus [3048]. Centropyge [3451]. Centroscymnus [2904, 264]. century [1824, 614, 918]. cepedianum [3095]. cephalic [1141, 458]. Cephalopholis [2597, 67, 1505, 2961, 1855]. Cephalopod [2839, 1816]. cephalus [2693, 1276, 1077, 3291, 3387, 3112, 1707]. cercariae [3059]. cerebellar [1263]. cernuus [1545, 2933, 3061]. certain [2366]. cessation [2205]. Cestoda [2262, 138]. cestode [1414, 2586]. Cetina [2913]. cf [2832, 1088]. chabelardi [175]. chaberti [1088]. Chaetodon [3373, 2256]. chaetodontid [1712]. Chaetodontidae [3350, 3022]. chain [338, 3343, 1744, 2715, 2774, 32]. Chairs [1901]. Chalcalburnus [1707]. chalcogramma [1815, 2610, 3278, 2422]. chalcoides [1707]. chalk [2012]. challenge [14, 1350, 2828, 2838]. challenges [906, 199]. chamber [2425]. chambered [227]. Champlain [3012, 3192]. change [2643, 1239, 2271, 932, 997, 3175, 519, 700, 3443, 1212, 2700, 2008, 937, 2508, 3337, 1836, 2161, 2402, 1955, 1916, 2401, 1065, 288, 1262, 1432, 2661, 3312, 493]. changeover [1085]. Changes [761, 554, 1512, 1606, 2122, 2974, 125, 399, 2195, 1637, 1151, 3066, 1391, 1179, 1741, 797, 590, 918, 1974, 2205, 2431, 438, 2788, 3153, 1871, 2622, 607, 1947, 632, 2940, 500, 1658, 1524, 30, 187, 1935, 3095, 743, 1353, 1545, 2363, 2616, 372, 191, 1547, 2294, 2832, 2191, 989, 14, 3082, 871, 5, 673, 861, 1063, 3131, 2936, 23, 1673, 1757, 309, 2161, 102, 2292, 489, 151, 81, 732, 323, 1760, 254, 1717, 1703, 1316, 1144, 2169, 288, 964, 1671, 2500, 663, 2602, 1033, 1957, 2163, 870, 3090, 121, 2957, 451, 1676, 922, 2454, 2211, 3080]. changes [347, 2689, 1153]. changing [172, 3311, 1220, 2494]. Channa [1945, 1915, 2847]. Channel [3251, 1320, 721, 707, 529, 1378, 1550, 2707, 1958, 3262, 2354, 1325, 2581, 32, 356]. channels [860, 2585]. Channichthyidae [2467]. Channidae [2847]. Chanos [3271]. chaos [1213]. characid [1353]. Characidae [2314, 3043, 3292, 2862, 1714, 3454]. Characidium [3296]. characiform [1537]. Characiforme [2292]. Characiformes [573, 315, 3043, 2493, 3296, 1326, 3459, 1896, 3454]. characin [2486]. Characteristics [2915, 1288, 215, 2589, 383, 1885, 3151, 3077, 2481, 1972, 606, 682, 3429, 299, 3430, 2744, 1845, 656, 391, 1150, 708, 895, 1067, 1467, 3280, 3391, 938, 252, 430, 1795, 1720, 1258, 1985, 271, 3227, 1636, 390, 471, 82, 1124, 257, 945, 822, 3449, 3214, 1883, 2123, 3380, 3154, 2080]. Characterization [2396, 3450, 3343, 427, 1350, 1355, 1666, 1369, 3064, 2918, 2854, 2516, 3039, 93, 2312, 880, 1279, 348, 2129, 2636, 2318, 3438]. Characterizing [597, 1165]. characters [1733, 93, 1591, 3195, 3245, 176, 3115, 3468]. Charleston [1010]. Charr [1324, 1913, 3161, 668, 966, 960, 3290, 1292, 484, 158, 525, 1038, 2807, 2905, 735, 2932, 2198, 659, 2383, 30, 442, 3460, 737, 2571, 1725, 1547, 3045, 651, 3102, 2635, 861, 993, 3029, 333, 3122, 702, 1019, 2909, 75, 3372, 1104, 1539, 806, 2726, 536, 579, 865, 1740, 1117, 3370, 143, 1796, 232, 29, 663, 980, 1724, 2423, 3128, 1711, 1469, 1140, 760, 3058, 1522, 3056, 2133, 2958]. charrs [281]. chaunacid [2439]. Chaunax [306]. checkered [3005]. checklist [601]. checks [2845]. Cheilopogon [1868]. Cheirodontinae [2862]. chelator [344]. Chemical [2100, 2165, 773, 2730, 3181, 1330, 499, 806]. chemically [3072]. chemicals [2028]. chemistry [2207, 2125, 2188, 2261, 2633, 2342, 2959, 2266, 3453, 2344, 367]. chemoreception [3339]. CHESAPEAKE [1405, 1736, 2937]. chevron [1945]. Chikugo [2285]. Chile [3107, 2912, 2585, 81, 419, 1284, 2943, 3064, 240]. Chilean [1518, 2478]. chilensis [2002, 2585]. chilling [463]. Chilomycterus [3215]. Chiloscyllium [2864, 2631, 1853]. chimaerid [1835]. Chimaeridae [1835]. Chimaeriformes [3359]. chimaeroid [3359]. chimpanzees [1217]. China [1025, 1912, 3257, 3294, 1680, 3038, 237, 3040, 207, 428, 3042, 2141, 3438, 2989]. Chinese [3046]. Chinook [589, 192, 2283, 399, 2901, 1126, 117, 414, 1589, 1514, 2042, 2764, 1892, 2320, 3361, 3204, 212, 2615, 3404]. Chionodraco [2467]. Chirostoma [2114]. Chloride [1321, 761, 2195, 2945, 60, 454]. chlorpheniramine [738]. Choerodon [3025]. choice [779, 192, 1241, 1200, 494, 100, 2554, 1185, 1771, 1335, 1993, 722, 3423, 2699, 1180, 1795, 1304, 2375, 514, 1268, 2615, 406]. choices [3243]. Cholecystokinin [843, 339, 3328]. cholecystokinin-immunoreactive [3328]. chondrichthyan [1664]. Chondrichthyes [2487, 3000, 2408, 2304, 2882, 2649, 2503]. chondrocranium [2000]. Chondroid [2060]. Chondrostoma [1847, 1606, 1949, 1615, 3159, 3214]. choriogenins [150]. Chorion [2621, 1552, 1907, 1765, 1164]. Chromis [2966, 2796]. Chromogobius [2009]. Chromosomal [3292, 32, 3429, 2920]. chromosome [3296, 3441, 1692]. chromosomes [3435, 2926]. chronic [350, 658, 1836, 999, 3128, 2305, 2474]. chronology [57, 1366]. chrysops [2216]. chrysotaenia [452]. chub [3044, 2693, 1068, 604, 1106, 1447, 396, 3387, 3112, 3380]. chum [932, 619, 871, 2686, 2403, 2990]. Chwaka [2284]. chymotrypsin [2502]. Cichla [2232, 2329]. Cichlasoma [3178, 946]. cichlid [1533, 2872, 3318, 3388, 1511, 494, 2020, 1383, 480, 2248, 997, 2199, 3398, 2849, 1177, 3395, 2747, 3178, 3266, 1432, 2142, 2811, 488, 1702, 946]. Cichlidae [334, 1098, 1526, 2325, 1458, 2142, 2132, 2811]. cichlids [1491, 2072, 1993, 1388, 7, 176, 2475, 1059]. Cienegas [946]. ciliaris [3263]. ciliary [1696]. ciliata [3103]. Ciliated [1090]. Circadian [2391]. circle [737]. circulating [1578, 1757, 2965]. circuli [2845]. circulus [3066]. Cirrhinus [3198]. cirrhosus [1411]. cirrosa [880, 1644]. ciscoes [2136]. Citharichthys [2369]. City [1493]. Cladistia [2975]. cladoceran [1000, 548]. Clarias [214, 1859, 728, 2935, 3016, 711, 2969]. clariid [2981]. Clark [12]. clarkii [3279, 3304, 3279]. class [1858, 355, 890, 2446, 3434, 366, 2323, 2918, 1064, 2519]. classes [632, 1111, 2469, 612, 3078]. Classical [2737, 2920]. classification [620, 1715]. clathratus [1975]. clay [1765]. cleaner [856, 1771, 3050]. cleanerfish [1202]. Cleaning [67, 750, 222, 1201, 1245, 1246, 1236]. clear [2232]. cleavage [1848]. Cleisthenes [3136]. Clha [2918]. Clha-DAA [2918]. client [67, 1771]. Climate [1486, 318, 3175, 3443, 2963, 2401, 3312]. Climate-related [318]. climates [2393]. climatic [1983, 2047]. climbing [2972, 543]. cline [2547]. clingfish [94]. Clinostomus [3335]. Clinton [2591]. clones [2784]. Cloning [2677, 2533, 2982, 132]. close [1923, 1188, 3014]. closely [2328, 1533, 1842, 1488, 2394, 2314, 1818, 2955]. closely-related [1818]. cloud [1216]. Clove [650, 699, 2272, 255, 1371]. clown [2497]. clownfish [3333, 3433]. Clues [3227]. Clupea [3132, 3110, 2295, 2406, 1794, 2421, 2918, 1499]. clupeid [3158, 2213, 2978, 2738]. Clupeidae [2857, 2082, 2298]. clupeids [2328]. clupeiform [1818]. cluster [274, 3344]. clustering [1691]. Clutch [1845, 1001]. Co [1901, 1404, 3057, 482, 2849, 2964, 1184, 1177, 3014, 1432, 1231, 1941]. Co-Chairs [1901]. co-dominant [1941]. co-evolution [1184]. Co-evolutionary [3057]. co-existence [3014, 1231]. co-existing [482]. co-occurring [2964]. co-operatively [2849, 1177, 1432]. coal [360]. coast [935, 2009, 944, 3017, 1397, 1982, 796, 2510, 2579, 1650, 106, 630, 193, 1318, 2082, 2634, 1931, 3280, 734, 854, 282, 2806, 1425, 917, 3048, 461, 2737, 2230, 2236, 2054, 2377, 3455, 2478, 921, 2108, 3276]. Coastal [1989, 3234, 2800, 2424, 2139, 1948, 1332, 2781, 1412, 619, 465, 1563, 2105, 1669, 2024, 916, 617, 1123, 1287, 2992, 35, 1393, 8, 1124, 920, 754, 1385, 2107, 3279]. coasts [1911]. cobia [632, 2400, 2684]. Cobitidae [601, 3471, 2471]. Cobitis [601, 2733, 2471, 654]. Cobitoidea [3428]. cocktail [430]. Cod [1686, 2343, 1266, 34, 2227, 2771, 2897, 284, 979, 187, 1332, 2924, 2579, 2561, 1206, 2446, 2235, 2848, 3189, 700, 3328, 2031, 2663, 2126, 2342, 511, 1039, 1588, 3273, 112, 2834, 66, 58, 1673, 2024, 2257, 1377, 41, 1789, 2666, 3081, 938, 3285, 2321, 2541, 1425, 3284, 3300, 1361, 2838, 2401, 288, 253, 1448, 1994, 476, 1226, 2802, 2264, 35, 1228, 1261, 1033, 1808, 3014, 2433, 2539, 603, 2609, 2228, 2628, 294, 2288, 945, 968, 2345, 1556, 1881, 486, 507, 841]. coded [2251, 1900]. coding [2921]. coefficient [2292]. Coelacanths [615]. coelolepis [2904, 264]. coercion [2699]. Coeur [12]. coevolve [1752]. coexistence [500, 2755, 2508, 321]. coexisting [3213, 1333, 1883]. cognition [1196]. cohesion [3298]. Cohesive [3094]. Coho [1524, 2812, 293, 1113, 985, 1017, 967, 3195, 83, 1393, 1573, 13, 1761]. cohort [2057]. cohorts [363, 3008, 2992]. coI [2331]. Coilia [2238]. coincident [2759]. coioides [3039, 3152]. Cold [3007, 208, 447, 546, 1363, 511, 2826, 2687, 774]. cold-water [2826]. collagen [2303]. collagens [1694, 82]. collapse [403]. collected [187, 3326]. Collection [1405, 2951, 395]. collections [2202]. collective [673]. collectors [3313]. Cololabis [2217]. Colombia [1805]. Colombian [2053]. colonial [1423]. colonization [2966, 2745, 2547, 1515, 828, 1931]. colonizations [3348]. colonized [212]. colonizing [2885]. Colorado [2097]. coloration [200]. Colossoma [2707]. Colour [3263, 495, 523, 1546, 480, 665, 1150, 2881, 374, 2206, 3101, 576, 811, 2564]. colouration [753, 495, 2120, 3210, 2967, 2364, 1854]. colourful [1517]. Columbia [847, 1456, 1514, 2686, 3184, 598, 3279]. column [1310]. columnare [3130]. combine [3084]. Combined [3181, 2424, 142]. combining [2866, 2202]. combtoothed [1431]. comes [2606, 695]. Comments [2978, 3217, 1780, 2516, 1141, 2658, 1520, 2458, 458, 3454]. commercial [3242, 3473, 2766, 854]. commercially [2357]. commerson [2108]. commersoni [530]. commersonii [852, 2372]. Commission [537]. Common [1, 1628, 2374, 3068, 1292, 2723, 132, 539, 3351, 199, 3371, 571, 790, 2716, 2384, 1100, 751, 98, 3286, 2632, 1330, 3100, 68, 490, 2174, 2245, 3355, 2395, 2480, 981, 2414, 3117, 3252, 118, 437, 1306, 1597, 838, 1675, 1028, 208, 313, 451, 556, 1603, 2636, 1630, 3080, 793, 3346, 231]. commonly [213]. communal [1567]. communication [1246, 1226]. Communications [559]. communities [569, 1534, 682, 1165, 1422, 1043, 242, 2027, 992, 2553, 919, 228, 1656, 1862, 2231, 163]. Community [908, 1805, 2275, 2866, 912, 3249, 1209, 2397, 675, 2118]. compact [2582]. Compagno [2839]. Comparative [699, 262, 1382, 291, 2394, 1860, 1611, 444, 2472, 2471, 1591, 51, 2708, 612, 517, 7, 2715, 3, 1167, 2134, 888, 1476, 1593, 2150, 2246, 999, 1596, 2673, 2859, 1403]. compare [105]. compared [40, 553, 2045]. Comparing [1851, 227, 2837, 2539, 1578, 2153]. comparision [2831]. Comparison [1842, 1534, 327, 2510, 256, 541, 914, 2633, 2517, 2124, 646, 1589, 2027, 2200, 2119, 1554, 232, 1393, 2792, 2288, 1862, 1099, 377, 966, 827, 3304, 1577, 3255, 3351, 2897, 952, 1413, 2813, 1141, 3157, 2096, 725, 2920, 1832, 2273, 389, 572, 756, 3221, 2005, 2587, 1618, 536, 579, 169, 1686, 1764, 2090, 2950, 824, 461, 1352, 2129, 1661, 2545, 2454, 1119, 1806]. Comparisons [2818, 1488, 33, 1966, 3267, 3445, 750]. compartments [721]. compatible [3423]. compensation [413, 27, 446, 2826, 1958]. Compensatory [20, 1602, 365, 940, 2215, 2716, 998, 393, 1640, 377]. compensatory-growing [998]. compete [3371]. competing [1234]. Competition [1459, 2835, 2044, 2907, 779, 1199, 1807, 868, 295, 2059, 1037, 2759, 705, 1569, 1348, 3014, 1809, 1231, 1389, 1451, 1237, 2379]. Competitive [583, 942, 526, 209, 1439, 2192, 2932, 3302, 500, 1725, 3102, 202]. competitor [3102]. competitors [1177]. Complementary [718]. Complete [3174, 956, 3221, 2593, 2148]. Complex [3295, 1494, 1858, 3290, 679, 2941, 1872, 1718, 1667, 1648, 1098, 2733, 2164, 2983, 1213, 1176, 2664, 2765, 3439, 2544, 2148, 3342, 378]. complexing [1641]. complexity [1683, 1209, 2418, 948, 1313]. complexone [3376, 501]. complexone-stained [3376]. component [965]. Components [1227, 3068, 1757, 3350]. composite [2268]. Composition [721, 871, 1287, 1917, 459, 2227, 764, 2214, 1633, 1241, 735, 890, 1407, 2419, 1147, 79, 1382, 170, 92, 2242, 2450, 3068, 943, 2273, 1151, 2865, 1844, 2847, 2968, 1620, 1717, 2186, 3370, 3237, 2057, 624, 1623, 82, 3177, 1161, 2231, 2672, 365, 2180]. compositions [2666, 3078, 8, 1866, 3353, 2503]. compressed [84, 1445]. compression [3110]. Computer [3318, 1176]. computer-activated [1176]. computerized [1822]. concentrate [3149]. concentration [293, 2929, 1044, 310, 3106, 1836, 717, 3268, 1973, 2365]. concentrations [821, 404, 783, 2202, 1547, 3456, 3275, 3311, 3155, 1736, 1392, 2036, 451, 367]. concept [3399, 1568]. Concepts [2601]. concerns [3476]. conchifera [2697]. concolor [521]. concordance [286]. Condition [2461, 1105, 1277, 1848, 1414, 200, 719, 2430, 2387, 979, 187, 2996, 1072, 532, 730, 2640, 2770, 2294, 1280, 519, 3463, 2848, 1478, 1951, 2016, 2657, 2759, 2634, 2769, 366, 2884, 1717, 3284, 3387, 878, 1813, 3232, 3358, 2604, 423, 1101, 3200]. Condition-dependent [2461]. Conditioning [1359, 1967]. conditions [1888, 402, 216, 217, 192, 3231, 55, 2579, 387, 2486, 652, 1637, 2623, 1564, 164, 2728, 3021, 708, 722, 1154, 3423, 2466, 716, 1567, 3300, 3252, 2182, 2221, 1736, 2003, 2, 1693, 991, 815, 486]. conductivity [1072, 2025]. Cone [1157]. Conference [1428, 1989, 2077, 2601]. confers [1197, 2379]. confined [2111]. confinement [3325, 168, 1526, 930, 437, 43, 1536]. confirm [2535, 651]. confirmation [2407, 3043]. confirmed [1727, 1522]. confirms [3019]. conflict [3168]. conflicts [3067]. congener [3118]. congeneric [3304, 3057]. conger [825, 228, 461, 793, 1972]. Congress [3478, 2255]. conirostris [1394]. conjunction [2114]. connected [2956, 2416, 2417]. Connecticut [1430]. connection [529]. connectivity [3367, 1244, 3192]. Conorhynchus [1394]. consequence [1617, 205]. Consequences [2179, 1590, 2825, 2693, 2021, 1647, 2753, 2577, 685, 3049, 3259, 3303]. Conservation [1324, 1902, 1963, 1824, 1990, 2019, 2049, 2078, 2101, 2127, 611, 1946, 601, 1198, 1087, 3170, 1582, 3000, 2336, 610, 1429, 3344, 616, 1615, 1616, 3288, 2458, 1954, 1944, 1600, 2444, 3409]. conserved [2728, 3295]. consideration [2480]. considerations [722, 2488]. considered [1282]. considering [2743]. consistency [57, 916]. consistent [1693]. conspecific [1501, 866, 2165, 1718, 3398, 2509, 1077, 1154]. conspecifics [3094, 1826]. Conspicuous [3148]. Constance [1285, 2558, 3098]. constant [153, 2073, 2818]. constellations [1793]. constituents [824]. Constraints [1466, 958, 3380]. construction [1158]. consumed [1471, 3453]. consumer [2164, 2403]. Consumption [1548, 240, 1480, 350, 2907, 954, 1384, 2282, 783, 1863, 1631, 2944, 2032, 511, 3131, 985, 169, 2505, 517, 232, 976, 727, 2228, 2628, 939]. contact [2326, 2718, 2732]. containing [825, 441]. contains [1694]. contaminated [12, 1968]. contamination [2825, 3408]. contemporary [3051]. content [2866, 554, 1633, 729, 2359, 1545, 301, 342, 2284, 1666, 1304, 1441, 469, 775, 1820, 1120, 804, 1354, 1395]. Contents [2310, 2333, 2549, 2742, 2515, 2704, 3086, 3264, 3482, 545, 1894, 79, 1631, 427, 2783, 1952, 508, 557, 1873, 566, 709]. contest [584]. context [3302]. continental [667, 2510, 277, 1682, 2904, 3326, 1783, 1034, 3200]. contingents [2202]. continuity [599]. Continuous [498, 2893, 2347, 3212]. continuum [3415]. contractility [447]. contraction [2816]. contrast [1345, 653]. Contrasting [3444, 3041, 2320, 3284, 56, 1374, 558, 3170, 3379, 2551, 2007, 2156, 2038]. Contrasts [520]. contribute [2060, 497]. Contribution [127, 2210]. contributions [1689]. Control [751, 457, 2713, 712, 956, 1860, 1611, 635, 3174, 3291, 3167, 2004, 2963, 3037]. controlled [716]. controlling [808]. Controls [3480]. Convenor [2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513, 1593]. Convenors [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. conventional [583, 3368, 3276]. convergence [3324]. conversion [540, 954, 1815, 470, 235, 473, 25, 288]. convict [480, 2248, 1388]. cookei [883]. cool [2642, 2640]. cooling [1357, 2209]. Cooper [2112]. cooperate [1234]. cooperation [1232, 1234]. cooperatively [1511]. Copadichromis [488]. copepod [2007, 1979]. coping [3024]. copper [2691, 173, 2997]. copulating [1792]. copulatory [1411]. Coral [1107, 453, 418, 2675, 1712, 3314, 1175, 2052, 168, 2727, 3399, 3311, 1515, 2397, 265, 2046, 1005, 2395, 3005, 2256, 3022, 1258, 3117, 2852, 1229, 1158, 2701, 1431, 2629, 2930, 1494]. coral-dwelling [1005]. Coral-reef [1107, 418, 3314, 3399, 3117, 2852, 2701, 2629]. corallivorous [3373, 1224]. corals [3373]. cord [1980]. cord-like [1980]. coregonid [2223, 275]. Coregonidae [1833, 2136]. Coregonus [3206, 1833, 616, 1311, 1969, 1717, 3089]. coriiceps [274, 660]. corkwing [137]. cormorants [3242]. cornetfish [129]. Coronary [3393]. correction [1815]. correlate [1290]. correlated [1072]. Correlates [1078, 15, 151, 3184, 318]. Correlation [2678, 1076, 857]. Correlations [2061, 692, 3462, 1109]. correlative [309]. correspond [3461]. Correspondence [2602]. Corrigendum [2254, 2463]. Cortez [2919]. cortical [1666]. Cortisol [636, 2828, 1913, 589, 1525, 1814, 2157, 2392, 168, 1350, 3203, 2441, 983, 437, 1820, 1325, 1981, 1838, 1099]. Corydoras [1701, 3382]. Corynosoma [179]. Corythoichthys [2595]. Cost [3059, 1202, 1584, 898, 3337, 2257, 2425]. cost-effective [898]. Costa [1353]. Costly [1395]. Costs [1201, 2218, 3299, 986, 1220, 345]. cotti [2506]. Cottidae [2110, 1795]. cottoid [1411]. Cottus [1807, 2110, 1987, 3369, 2506, 1849]. Council [456, 477]. counter [2995]. counter-illumination [2995]. Countergradient [235, 1228]. counterparts [2240]. countries [3461]. Coupling [1889]. courting [1533]. Courtship [3458, 3402, 975, 2872, 177, 665, 773, 2226, 2219, 1395, 1451]. cover [1892, 328, 76]. coviiceps [752]. cownose [2641]. Cowx [1773]. Cox [2630]. coxii [2630]. Cranial [2498, 3382, 2959, 1148]. craniofacial [1355]. crappie [2504]. crassum [2586]. crassus [3161, 961, 563, 1312, 3244, 3061]. Craterocephalus [1684, 599]. crayfish [1459]. Creagrutus [1326]. creatine [3080]. Creating [608]. crebripunctata [3324]. Credit [2080]. Creediidae [1649]. creeks [908]. Crenicichla [1241]. Crenuchidae [3296]. Creplin [138]. cretense [2822, 2158]. crimson [3247]. cristata [3396, 1042]. cristatus [2139]. criteria [740]. criterion [2206]. Critical [311, 2215, 737, 2818, 3026, 1568, 2748, 2681]. critically [3440]. croaker [1345, 3287, 2982, 1782]. Croatia [2255, 6, 2913]. Croatian [3417]. crocea [1864, 2982]. Cromarty [3186]. Cross [1698, 2517, 1540, 756, 1270]. cross-amplification [2517]. cross-population [1270]. cross-river [1540]. Cross-species [1698]. crosses [1692]. crucian [2983, 2861, 3219, 447]. crude [3370]. cruentata [1855]. cruentatus [1467]. Cruz [343, 2501]. cryopelagic [273]. Cryopreservation [1087, 336, 672, 1448, 628, 2805]. cryopreserved [672]. Cryptic [2725, 3350, 1726, 3348, 561, 971]. cryptobenthic [2580]. crystal [952]. crystalline [891]. Crystallodytes [1649]. ‘Ctenochromi [2132]. Ctenopharyngodon [1838]. Ctenopoma [1800]. CTHS [1405]. Cu [1668, 2277]. Cuatro [946]. cues [2991, 2460, 1718, 773, 1330, 1734, 499, 806, 1235, 1267]. Culaea [1499]. cultivated [1558, 1564]. culture [263, 2192, 1582, 1217, 2274, 2312, 2720, 1574, 1654]. cultured [821, 2897, 2419, 1561, 1585, 2590, 1593, 1507, 147, 1450, 1622, 2090, 1576, 1600, 2989, 1586]. cultures [1513]. Curcumin [3020]. currani [2151]. Current [1992, 3034, 2794, 2562, 3307, 3106, 2537, 3228]. curves [1760]. Cutaneous [674, 1765]. cuticola [741]. cutthroat [12, 1452, 598, 3279]. Cuvier [662, 2023, 1785, 2865, 2707, 2870, 2689]. cyanobacteria [685]. cyanobacterium [2098, 2801]. cyanopterus [1785]. cycle [459, 97, 2265, 821, 1409, 1058, 2358, 278, 301, 2191, 146, 2657, 16, 1051, 1741, 964, 1788, 3090, 2389, 2596, 1059]. cycles [413, 935, 445, 2297, 939]. Cyclic [2215, 2073]. cyclical [1153]. cyclically [2214]. cycling [3210, 802]. Cyclopteridae [2711, 2171]. Cyclopterus [2171]. Cymatogaster [869]. Cynoglossus [3293]. Cynolebias [1738]. Cynoscion [1156, 819, 2937, 1911, 3347, 3312]. CYP1A1 [2739]. Cyprichromis [2142]. Cyprinella [3096]. cyprinid [3185, 851, 1538, 1731, 789, 3357, 2845, 1509, 2884, 741, 1257, 2097, 2278, 853, 524, 3057, 727, 2141, 3099, 2853]. Cyprinidae [582, 3220, 2676, 1203, 1799, 1667, 1831, 3341, 2547, 3345, 1519, 2520, 3219, 2148, 1374, 2854, 1707, 2095, 3303, 3042, 2913]. cyprinids [3255, 840, 160, 1698, 2708, 675, 1878, 2209, 726, 2844]. Cypriniformes [3044, 3428, 3471, 1519, 3042, 2913]. Cyprinodon [3348]. cyprinodontid [423]. Cyprinodontidae [3125, 2353, 2855]. Cyprinodontiformes [2619, 1843]. cyprinoid [3038]. cyprinoides [2683]. Cyprinus [2758, 199, 3371, 3286, 2632, 3327, 3100, 1753, 2245, 3355, 2480, 3252, 2681, 2636, 3080, 3346]. Cyrino [3321]. Cytochemical [2173, 1720, 2117]. cytochrome [34, 840, 2719, 886, 3341, 2981, 96, 1401, 2331, 2706]. Cytogenetic [2516, 2528]. cytogenetical [2740]. cytogenetics [2737]. cytometry [2920]. cytoplasmic [2542]. Cytotaxonomy [3043, 3459, 3435]. Czech [1572, 2556, 3214, 2861].

D [1654, 1774, 3478, 3321, 1665, 1652, 3407, 3480, 2140]. D-loop [2140]. DAA [2918]. dab [346, 1952, 2634]. dace [3190, 3335, 524, 1233]. Dactylogyrus [3057]. dactylopterus [2596, 1026, 2596]. daggertooth [400]. dahomeyensis [354]. Daily [1865, 3243, 977, 884, 261, 698, 3004, 2298, 1899, 3269, 169, 1679, 3159, 2595]. d’Alene [12]. dalli [172, 2494]. dam [2027]. damage [2920, 771]. Damalichthys [869]. dammed [7]. dams [230, 2962]. damsel [2166]. damselfish [418, 1281, 95, 794, 2594, 1225, 1001, 2066, 2280]. damselfishes [2796, 2189]. Danguy [3322]. Daniel [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. Danio [3120, 2980, 2892, 3218, 3033, 3217, 3378, 2193, 1887, 2208, 738, 2155, 1269]. danionin [2193]. Danois [3364]. Danube [2069, 790, 304, 3214, 1531, 1532]. Dapeng [3115]. Daphnia [1048, 1000, 548]. dares [1270]. dark [2486]. darkness [1820]. Darling [3286]. darter [510, 3431, 2794]. darters [256, 2165]. Darwinian [3414, 3412]. Dascyllus [418, 1722, 2189]. Dasyatidae [3182]. Dasyatis [3009]. data [966, 2535, 545, 2985, 79, 1745, 64, 790, 886, 3341, 1753, 3023, 166, 3439, 3008, 1994, 2045, 1721, 1622, 1358, 488, 2108, 2685]. datalogger [1869]. Date [3070, 2610, 2390, 950, 1450, 1768]. date-dependent [1768]. day [3391, 2231]. days [829, 1066]. de-calcification [2489]. dearborni [3348]. death [2593, 1702]. debris [2299]. debt [2683]. decade [2752]. decades [1600]. Deceptive [2695]. decision [1230]. decisions [2104, 3423]. decline [2760, 613, 3280, 2299]. Declines [2256, 2109]. declining [1637, 3236]. decoupling [392]. decreasing [2614]. decussatus [3005]. deduced [2771, 2108]. Dee [2122]. Deep [3141, 2439, 1502, 664, 1413, 730, 264, 2840, 84, 1540, 3216, 857, 502, 1918, 1835, 3078, 3364, 464, 2036, 2244, 240, 889, 793]. Deep-diving [3141]. deep-sea [664, 1413, 730, 2840, 3078, 3364, 2244, 240, 793]. Deep-water [2439, 264, 3216, 1835, 2036]. deepwater [667, 3000, 2425]. defence [3153, 1002, 1222, 2166]. defend [2248]. defensive [2010]. deficiency [746]. deficiency-related [746]. deformations [3206]. deformed [1161]. degraded [641]. degree [3096, 1294]. dehydrogenase [1509]. delay [2379]. delayed [1647, 1660, 1280, 1828]. delays [527]. delineated [622]. delineating [1719]. delineatus [2760, 1956, 2746, 2485]. delivery [1176, 649]. delosommatus [502]. delta [1121, 2647, 846]. demand [717, 1173]. demersal [2994, 913, 1918, 169, 488, 2205]. demineralization [1138]. demographic [374, 3448, 2919, 848]. demographical [3288]. demography [2348, 229]. demoiselles [797]. demonstrate [1720]. demonstrates [684]. Demonstration [830, 2730]. Denmark [1570]. dense [202]. densities [1577]. Density [412, 74, 3248, 704, 376, 3262, 1535, 1578, 1748, 2043, 736, 758, 2367, 3241, 3126, 805, 46, 1950, 3102, 1350, 942, 342, 1642, 2772, 1744, 485, 1569, 955, 467, 1450, 755, 72, 804, 1551, 1424, 1576, 2048, 1454, 892, 561, 492]. density-dependence [892]. Density-dependent [412, 74, 3248, 704, 1535, 736, 758, 46, 1950]. Density-independent [3262]. density-separated [804]. dentary [2060]. dentatus [2730]. Dentex [1657, 3257]. Dentition [1282]. depend [2375]. dependence [1437, 113, 970, 726, 892]. dependency [929]. dependent [2461, 1893, 412, 736, 3132, 758, 1807, 1014, 3110, 2235, 3469, 46, 1950, 74, 342, 341, 1530, 3248, 704, 66, 58, 2784, 2335, 1033, 10, 1264, 1535, 1289, 2155, 1768, 492, 3423]. depensation [2216]. depleted [3399]. depletion [2687, 3278, 1617, 2256]. deposition [1481, 261, 1526, 1309, 2298, 1294, 365]. depressed [855]. depression [3128, 1625]. deprivation [54, 413, 191, 140, 31, 20, 772, 2665, 1640, 939, 365, 377, 940]. Depth [445, 3110, 2184, 3154, 1084, 549, 725, 2755, 1150, 2553, 1327, 2844, 2231, 2615]. Depth-dependent [3110]. depths [2974]. depuration [1641, 1420]. Derivation [1112]. derivatives [3020]. derived [956, 18, 572, 2277, 1575]. dermal [1293]. Dermatolepis [2910]. descending [2812, 1059]. describe [1893]. Description [44, 123, 2142, 2811, 226, 1453, 1684, 2840, 2733, 283, 491, 1036, 2135, 296, 1802, 1465]. desert [2091]. design [1523, 3305, 3335, 1957, 2425, 2680]. Designing [2338]. destroying [608]. destructive [1770, 1905]. detect [820, 2315]. Detecting [1205, 3076, 1601]. Detection [1154, 3023, 2922, 1127, 2878, 1540, 3394, 785, 2321, 1049, 1440, 1024, 405]. Determinants [1976, 1174, 2239]. Determination [515, 2114, 1354, 1747, 814, 457, 2733, 994, 2490]. determine [2207, 632, 2929, 872, 2900, 1436, 1258, 2268, 2344]. determined [3127, 325, 120, 3384, 1909, 2556, 2637, 3194, 3276]. determines [2494, 238]. Determining [2777, 2660, 2726]. determinism [3234]. deterrence [2098]. detritivorous [764, 3177]. developed [2924]. developing [3328, 843]. Development [928, 341, 693, 3212, 2405, 3016, 825, 1125, 837, 3245, 1981, 556, 1101, 2318, 3438, 507, 1738, 620, 418, 1722, 1387, 1239, 554, 947, 3104, 122, 87, 1856, 2443, 1995, 1054, 1349, 783, 297, 2185, 2995, 1889, 2419, 1825, 1733, 165, 2499, 261, 858, 2157, 326, 795, 427, 2294, 2235, 3433, 1890, 864, 2472, 3463, 2631, 1076, 987, 426, 1063, 2151, 527, 3275, 440, 250, 2912, 2071, 1505, 2945, 2778, 1999, 996, 2607, 1864, 1756, 1482, 2656, 147, 2428, 2927, 151, 2898, 2594, 366, 970, 1310, 1828, 1956, 530, 1373, 221, 1441, 1448, 2000, 657, 2842, 1622, 417]. development [1657, 1958, 1886, 3362, 246, 1307, 1714, 1830, 1061, 2243, 1381, 1767, 1768, 3336]. Developmental [519, 2208, 338, 3217, 3124, 3024, 2662, 320, 2491, 2155, 2305, 2859]. device [395]. devices [1732, 901]. DeVries [3480]. dewatering [1231, 2835]. dewitti [283]. dexamethasone [547]. dextral [3136]. D’Hondt [1405]. di- [90]. diadromous [2772, 3299, 2027, 2841]. diadromy [2261]. diagnostic [2996, 828]. Dialommus [343]. diameter [3135, 2288]. Diana [1493]. diandric [1722, 2508]. diapause [1831]. diaphanus [2326, 3349]. Diaphus [2911, 1397, 1899]. Dicentrarchus [1923, 2387, 1608, 3457, 2195, 3183, 3391, 1829, 1710, 2824, 3362, 2533]. dichromatisms [2290]. did [3051]. didactylus [2306, 2991]. dieffenbachii [823]. Diego [1494]. Diel [1085, 1685, 1147, 698, 2022, 1530, 1311, 2960, 780, 2378, 1329, 2777, 860, 1948, 859, 1426, 3025, 1257, 3098, 3252, 1389, 1782]. Diet [222, 2092, 2426, 3314, 1086, 1204, 2551, 538, 2630, 700, 202, 1971, 3446, 1516, 2865, 2968, 3396, 1786, 724, 287, 2183, 2574, 1885, 459, 1084, 217, 3073, 2807, 3388, 1023, 1407, 2780, 3095, 79, 862, 1175, 2640, 2275, 139, 1667, 259, 2450, 698, 429, 1019, 2766, 68, 3125, 2837, 2193, 2292, 81, 897, 2416, 2884, 1477, 1117, 2186, 2954, 1193, 7, 47, 232, 72, 2003, 1264, 2201, 1455, 1866, 2560, 1322, 1883, 1235, 2564, 1142]. diet-based [1235]. Dietary [1758, 1083, 3196, 1668, 474, 1022, 2244, 428, 214, 2265, 284, 3149, 784, 2942, 658, 426, 2007, 3350, 775, 824, 8, 1623, 1325, 2839]. Dietary-morphological [428]. diets [1908, 2203, 2866, 3149, 349, 5, 173, 891, 1620, 3022, 3014]. differ [1435]. Difference [3073, 2302, 1571]. Differences [1533, 1584, 387, 3213, 174, 575, 2909, 1565, 2040, 983, 1346, 237, 668, 1292, 2872, 2667, 1417, 284, 2610, 686, 2745, 1610, 465, 2849, 3464, 2089, 791, 2226, 702, 2690, 1783, 205, 3355, 1080, 439, 1916, 1361, 1421, 2873, 1742, 1934, 1004, 2269, 697, 423, 1268, 1522, 1236, 2564, 1768]. Different [2963, 879, 484, 1577, 2339, 1485, 3084, 888, 783, 2996, 2358, 1248, 3190, 256, 249, 194, 2025, 656, 1186, 473, 1435, 2056, 685, 440, 3137, 3397, 3378, 1252, 2690, 2553, 3208, 1327, 716, 885, 1735, 3418, 3384, 178, 2965, 3258, 2393, 919, 2182, 2491, 748, 2442, 69, 1675, 386, 2584, 2211, 2180, 486, 507]. Differential [392, 2876, 1489, 3231, 1188, 1104, 1567, 499, 83, 1844, 2987, 2320, 2628, 1499, 1173]. differentially [2034]. differentially-fished [2034]. Differentiated [3435]. differentiation [1052, 2643, 3290, 2328, 2516, 1031, 153, 2339, 2723, 311, 2340, 1923, 2125, 2347, 2317, 571, 2440, 1020, 259, 304, 2978, 749, 1969, 1980, 2246, 1829, 1143, 1710, 3048, 450, 281, 3027, 2530, 1827, 2856, 1602, 1932, 2738, 2518, 52]. differently [768]. difficult [1503]. Diffusion [1243, 2302, 1264]. digenean [3148]. digesta [170]. Digestion [210, 1794, 2876, 12, 898, 2783, 2450, 2406, 3453, 843]. Digestive [1488, 3460, 3321, 3208, 2158, 2186, 315, 1054, 1015, 1353, 2400, 196, 951, 2691, 1864, 3362, 3336]. dihydropyridine [1744]. dilute [128, 476]. dimension [2803]. dimensions [73, 2170]. dimidiatus [856]. dimorphic [137]. dimorphism [2676, 584, 2097, 430, 2618, 238, 2142, 2661, 1061]. dinoflagellate [2774]. Diodon [1521]. Diodontidae [3215]. dioxide [927, 279]. Diphyllobothrium [138]. Diplodus [586]. diploid [1995, 1560, 2735, 291, 795, 2302, 2901, 3293, 43, 2722]. Diplomystes [420]. Diplostomum [3058]. Dipturus [2766, 2002, 2585]. Direct [1081, 2764, 1993, 2896]. direction [2608, 1183, 2547]. directional [937, 2402, 2494]. Directions [1901]. disadvantageous [1169]. discard [1647]. discarded [3142]. discharge [1195, 1244, 2093, 587, 791, 3211, 1730]. discharge-specific [791]. discharges [3092, 2243]. discharging [1648]. discontinuous [2347]. discovered [903]. Discovery [1664, 1753, 6, 2748]. Discrepancies [233]. discrepancy [2213]. discriminate [2791]. Discriminating [2341]. Discrimination [293, 823, 2068, 2934, 1110, 2943, 1543, 3009, 1057, 1911, 2870, 3176, 2478, 3206]. Discus [2325]. discussion [2709]. discussions [3476]. Disease [1113, 1595, 1060, 219, 930, 2922, 2280]. disease-affected [2922]. Diseases [1654]. Disentangling [3423, 2340]. dishonest [1395]. disintegration [344]. disjunct [2523, 251]. disjunctivus [3010]. disorder [746]. Disorders [1654]. disparity [2216]. dispensing [1176]. dispermic [1408]. Dispersal [2279, 2309, 2381, 2436, 2465, 2492, 3255, 2411, 1432, 2513, 1474, 1194, 2147, 375, 1211, 2946, 3134, 1257, 321, 1258, 3117, 2485, 1229]. disperse [2746]. dispersed [867]. Dispersion [756, 3069]. displacement [525, 1825, 942]. display [338, 1395, 1390]. displaying [2959, 1395]. disrupting [2121]. Disruption [1617, 1323]. disruptive [1451]. dissemination [1920]. dissimilar [2466]. dissipation [105]. Dissolved [310, 2920, 3106, 3155, 717, 83, 748]. Dissostichus [2207, 2033]. distal [3083]. distance [3141, 1484, 2302, 126, 2613, 2042, 2973, 2705]. distinct [1842, 3174, 3432, 3043, 2223, 273, 3459]. Distinction [2152]. distinctness [1458]. distinguish [3245]. Distinguishing [1759, 3115]. distributed [2793]. Distribution [1538, 2362, 2780, 12, 468, 3328, 156, 1318, 2007, 2033, 1285, 3308, 2894, 1621, 2887, 3101, 2483, 1979, 2669, 2386, 569, 932, 132, 549, 2710, 2445, 2979, 3450, 2496, 2640, 110, 1476, 3307, 803, 1186, 241, 2733, 120, 3292, 2249, 1817, 2585, 3349, 2987, 251, 1728, 2232, 1074, 361, 741, 416, 2118, 142, 1644, 1287, 47, 2162, 1264, 2267, 2844, 807, 1379, 138, 1064, 65, 3055, 3312, 1329, 3154]. Distributional [2397, 938]. distributions [1165, 2610, 303, 2603, 2160, 2455, 2187, 3022, 69]. District [15]. disturbance [2404, 3422, 2021, 2550, 3384, 1362]. disturbances [1105]. disturbed [1712]. disulphides [2639]. dithizone [344]. ditremum [138]. ditropis [2948, 1811]. Diurnal [2556, 2999, 1530, 1532]. Divergence [1401, 2038, 212, 2675, 3440, 2717, 2543, 2336, 1860, 3045, 2099, 3041, 1462, 1123, 166, 229, 1808, 3324, 3172, 2545, 971]. Divergent [2956, 2718, 2732].

Diverse [2110, 3439]. diversification [3044]. Diversity [2733, 3349, 1494, 78, 312, 1457, 2517, 190, 2775, 3345, 2150, 617, 3121, 2329, 917, 1122, 229, 2792, 603, 3192, 2701, 2545, 3116, 597, 561, 2519, 3438, 2989]. Dividing [599]. diving [3141]. DNA [107, 689, 34, 484, 2131, 2988, 3440, 3437, 2543, 2713, 2521, 91, 1489, 2096, 2740, 2793, 243, 519, 3431, 2727, 2147, 646, 1614, 1471, 2983, 3292, 190, 2736, 59, 622, 2140, 2869, 2987, 2981, 1755, 1399, 3121, 2321, 292, 2855, 564, 2919, 1518, 127, 2520, 2325, 1343, 3037, 276, 1873, 2798, 2331, 2540, 267, 1470, 422, 2529, 3116, 2797, 3114, 378, 1522, 971, 2990]. DNA-based [1471, 2869, 2331]. DNAs [2854]. Do [1951, 3029, 1752, 3461, 2260, 1258, 8, 1234, 2215, 1196, 3084, 1003, 2571, 1587, 1495, 1909, 1599, 1711]. docosahexaenoic [426]. Döderlein [1465]. doederleini [813]. Does [2404, 494, 3190, 1771, 2682, 1435, 2449, 1388, 1740, 321, 1193, 1229, 60, 1576, 3368, 295, 3230, 1668, 873, 474, 3453, 1571, 2277, 3426]. dogfish [761, 984, 2888, 2904, 2815, 3047, 2385, 1268, 1155]. dogfishes [549]. doll [3453]. dollfusi [2967, 157]. dolloi [3330, 2938]. Dolly [2094, 2133]. dolomieu [2956]. domensticated [382]. domestic [1581, 1611]. domesticated [1579, 1584, 3355, 2895, 1934, 2519]. domesticates [1557]. Domestication [1593, 1603, 1172, 1583, 1375, 1564, 2720]. dominance [2758, 1242, 1512, 1249, 1209, 2371, 1566, 3150, 1117, 1260, 767, 1346, 3259]. dominant [2571, 1710, 3177, 1941]. dominated [1102, 3308]. domination [673]. Domingos [1137, 1996]. donor [1589]. donors [2429]. don’t [1219]. dopaminergic [746]. dorada [216, 217, 987]. Dordogne [2841]. Dorosoma [3095]. dorsal [3359, 401, 1770, 838]. dorso [497]. dorso-lateral [497]. dosage [3376]. Dose [2155]. Dose-dependent [2155]. Doto [3278, 492]. Douro [600]. down [1246, 2767]. down-regulated [2767]. Downeast [1933, 190]. Downstream [1825, 2426, 2777, 2070, 2481, 2042, 1744, 1257, 2688, 2567]. drag [141]. dragonet [1742]. dragonfishes [2960]. drainage [860, 442, 2147, 2148]. drainages [3185, 304]. Drake [44]. Dramatic [2299]. Dreissena [3012]. drift [2723, 103, 853, 524, 1260, 1531, 1532]. drift-feeding [524, 1260]. drifting [624, 2378]. drinking [926]. drive [576]. driven [3251]. drives [3211, 45]. driving [2715]. droughts [77]. drum [1328]. dry [1908, 1620, 2231, 3263]. DSc [645]. dual [1513, 227]. dual-chambered [227]. duboulayi [3247]. Dubrovnik [2255]. duct [2643, 3031, 401, 1379]. due [1558, 1336]. Duero [2718]. Duméril [2487]. dumerili [2391]. dumerilii [1397]. dummy [3111, 2228]. Dunaliella [2611]. Dunbar [792]. duperrey [3458]. duplication [3035, 2916]. Durance [1086]. duration [3282, 2065, 2580, 3229, 2390, 1526, 959, 298, 2681]. during [362, 438, 2643, 459, 152, 625, 209, 975, 1439, 528, 1387, 554, 122, 2642, 3141, 2368, 445, 1054, 1349, 1524, 3427, 783, 128, 1972, 2358, 2419, 718, 125, 829, 743, 165, 77, 2157, 2392, 399, 12, 2640, 372, 427, 301, 546, 2373, 1375, 529, 1637, 3432, 2623, 117, 14, 1151, 1612, 987, 693, 1025, 527, 2412, 1138, 2815, 3275, 2767, 2826, 1278, 2217, 2223, 1515, 1919, 1017, 309, 2161, 3256, 1756, 116, 2686, 633, 2427, 3150, 2276, 1829, 938, 366, 140, 2923, 1620, 1283, 1144, 590, 574, 1956, 918, 1544, 319, 663, 1813, 463, 437, 83]. during [141, 1657, 150, 1124, 1231, 2835, 2190, 2476, 2533, 313, 451, 1640, 1234, 2518, 1059, 493, 347, 3336, 231]. dusky [944, 1922, 1174, 359, 2578]. DVD [1775]. dwarf [238]. dwelling [525, 900, 1347, 307, 1950, 2059, 1037, 1491, 3248, 2094, 2153, 933, 1005, 585, 2648, 232, 1458]. Dyje [2861]. dynamic [1339, 3302, 1658, 1790, 623, 2962]. Dynamics [546, 1494, 1945, 1946, 216, 2296, 3379, 1206, 3456, 1185, 723, 389, 3397, 236, 1592, 3238, 1075, 732, 366, 2335, 130, 624, 1367, 2562, 3079, 1306, 3169, 1237]. Dysomma [1912].

E.v. [456, 477]. ear [1696, 2385, 2113]. earlier [2240]. Early [1901, 1738, 2905, 2995, 360, 1463, 2079, 619, 1949, 98, 250, 929, 1505, 2778, 2428, 322, 1341, 2000, 1970, 3362, 1714, 115, 1848, 217, 1891, 2424, 3044, 2443, 1349, 2809, 3132, 1935, 1332, 165, 2496, 189, 3158, 261, 858, 2684, 387, 2373, 2472, 3033, 1191, 3024, 440, 2826, 1980, 2945, 2841, 2769, 1756, 53, 366, 1391, 581, 1310, 634, 3089, 1956, 2087, 3387, 1258, 3117, 1334, 1120, 417, 1294, 1575, 2537, 131, 1767, 1768]. Early-life [1901]. Early-season [929]. East [1383, 913, 334, 1347, 2098, 2481, 1684, 3000, 3017, 730, 2079, 1504, 2184, 2181, 1719, 599, 2428, 1835, 854, 2963, 3119, 3182, 3447, 3364, 364, 281, 1808, 2183, 2539, 3186, 2236, 2288, 2108, 3276, 1025, 511, 1912, 1537, 1036, 2530, 3257, 394, 223]. Easter [787, 1649]. eastern [2800, 3077, 2439, 1805, 1923, 3450, 2447, 1430, 839, 1682, 2640, 630, 2948, 728, 688, 2630, 2846, 156, 2660, 3105, 375, 3446, 2964, 3472, 2398, 1502, 2949, 3019, 2002, 3013, 1520, 3280, 2961, 1918, 1123, 3375, 2135, 1662, 1796, 896, 1305, 1401, 2943, 2144, 2618, 3436, 3324, 1352, 464, 1021, 2357, 2343, 2377, 2701, 2244, 1385, 2107, 2503, 3374, 2146, 2180, 2853, 1396, 101]. Eastern-Central [101]. eating [2947, 2475]. Ebel [3411]. Ebro [2647]. Echeneid [1142]. echinocephalus [2675]. Echinorhinus [883]. echo [1436, 411]. ECI [2255]. eco [725, 369]. eco-morphological [725, 369]. ecocline [2994]. Ecohydrological [2601]. Ecohydrology [1653]. Ecological [2601, 1146, 3372, 3340, 1598, 2564, 960, 1647, 3416, 765, 1467, 2181, 1144, 1018, 1073, 1431, 2668, 145, 3353, 922]. ecologically [3464]. Ecology [1314, 1706, 2788, 621, 1493, 282, 346, 1672, 409, 436, 877, 2941, 262, 2053, 1585, 1429, 630, 2100, 1949, 651, 1139, 394, 2361, 193, 1670, 1115, 2552, 2775, 703, 1549, 343, 2507, 3309, 1956, 1686, 1944, 1736, 2609, 1542, 2875, 1034, 1630, 3390, 207, 3409]. ecomorphological [2263]. Economics [1772, 3323]. Ecosystem [1824, 1989, 3410, 3049, 882, 1658, 700, 1333, 3308, 775, 1431, 1494]. Ecosystems [1989, 3138, 512]. ecotoxicological [272]. ecotype [3127, 1739]. ecotypes [2909]. ectenes [2238]. ectoparasite [168, 2662]. Ectoparasites [1245]. Ecuador [343]. ed [1772, 1493]. Edge [1774]. EDIPUCRS [1495]. Edited [3479, 1403, 1497, 1406, 1654, 3410, 1402, 3323, 1651, 1404, 3322, 1773, 1774, 1775, 3478, 3405, 3321, 1405, 1652, 3406, 1494, 3407, 3480, 1653, 3320, 1496]. Editora [1495]. Editorial [3088, 1271, 2128, 2977]. Edward [2118]. Edwards [2458]. Edwin [645]. eel [3234, 2789, 19, 2739, 2346, 3246, 3229, 1856, 2481, 233, 3231, 839, 3233, 180, 563, 2242, 2724, 259, 765, 818, 2509, 1063, 2690, 3272, 3238, 2246, 3209, 1880, 258, 1621, 3237, 228, 461, 1472, 2530, 2210, 1900, 3109, 2236, 3232, 720, 628, 2637, 1317, 1449, 1802, 1061, 2454, 347, 3154, 3228, 1496]. Eels [3224, 2825, 122, 3243, 373, 558, 890, 3235, 3240, 965, 466, 3231, 961, 963, 1189, 3230, 194, 2674, 340, 1135, 3236, 3461, 823, 926, 149, 361, 3245, 395, 2376, 3244, 1279, 3225, 237, 1385, 2107]. Effect [350, 953, 3130, 1821, 3377, 1407, 3093, 1583, 3060, 1631, 1790, 1798, 2191, 2438, 2484, 2783, 1797, 2727, 658, 198, 1062, 701, 2607, 717, 1256, 1569, 3385, 2965, 1828, 1510, 2491, 3159, 1325, 2598, 3313, 2488, 1866, 810, 76, 3306, 939, 1368, 2396, 214, 1604, 2138, 3282, 1891, 523, 735, 1856, 2710, 297, 1907, 2367, 260, 2251, 580, 2213, 1029, 2332, 856, 2849, 3011, 2058, 426, 2879, 396, 2027, 432, 819, 1528, 3378, 1919, 3383, 1452, 3256, 1377, 547, 1053, 2833, 361, 2063, 2186, 1365, 1001, 999, 31, 298, 2159, 1441, 853, 1691, 1448, 2500, 1994]. effect [809, 1887, 3147, 1551, 566, 69, 2228, 720, 2545, 2939, 1129, 754, 1234, 1236, 2029, 1030, 570, 2365, 217, 2608]. effective [3171, 3045, 898, 2348, 2531]. effectiveness [2414, 1958]. Effects [3002, 540, 263, 2265, 1238, 1641, 2262, 2192, 1240, 1276, 230, 2283, 2611, 172, 3242, 758, 1041, 2996, 606, 2073, 177, 1660, 2499, 191, 2878, 470, 1547, 2034, 1043, 1280, 2295, 259, 3463, 3102, 1356, 818, 1363, 2810, 481, 2056, 1831, 1048, 1987, 75, 3203, 933, 2945, 2881, 3111, 1744, 1756, 716, 1526, 885, 3103, 1082, 2696, 531, 2050, 1160, 2593, 1190, 1730, 3300, 1708, 1089, 2505, 3370, 2418, 219, 3252, 3369, 976, 320, 3205, 2774, 379, 1386, 1170, 2285, 2212, 3244, 26, 2844, 1542, 768, 826, 2036, 2591, 181, 1659, 3058, 1838, 1881, 3352, 835]. effects [2453, 3094, 459, 954, 392, 569, 932, 2340, 2121, 578, 1511, 1645, 2443, 2481, 2197, 1581, 979, 544, 1608, 3231, 1484, 2272, 33, 1825, 3389, 2496, 1609, 1904, 2392, 2743, 3360, 3434, 2592, 2093, 2942, 168, 2719, 2848, 2059, 1371, 1426, 2663, 1113, 164, 2218, 3443, 2281, 3181, 305, 2051, 2705, 2691, 3337, 4, 1482, 2422, 1789, 2427, 3268, 3391, 2833, 3150, 2403, 3301, 3081, 1105, 252, 48, 2801, 3108, 517, 3297, 2216, 781, 930, 530, 878, 1334, 772, 995, 1228, 3178, 767, 2, 1338, 811, 385, 386, 770, 533, 1454, 2604, 1364, 1536, 1235, 2155, 2615, 2305, 3156]. effects [1806, 2689, 1340, 2280, 3228, 841]. Efferent [2643]. Efficacy [1559, 3282, 3403]. efficiencies [288]. Efficiency [1601, 699, 540, 954, 210, 1200, 470, 1186, 235, 473, 2058, 25, 2903, 2452, 3394, 2393, 2598]. efficient [2714]. effluent [1002, 4]. effluents [2499]. effort [1118]. effuent [82]. Egg [94, 2487, 2619, 831, 2162, 1888, 1848, 1871, 3130, 1023, 3427, 1552, 42, 1984, 2499, 73, 2550, 2460, 2947, 3179, 3124, 3105, 3401, 1589, 1062, 1188, 1752, 2817, 1482, 967, 2441, 3140, 2259, 1620, 1310, 3138, 3300, 2808, 383, 2500, 1870, 2873, 2651, 464, 3466, 313, 486]. Egg-cases [831]. egg-eating [2947]. egg-tending [2817]. egg-to-fry [1984, 2499]. eggs [2621, 2571, 2617, 2590, 2460, 2242, 2235, 431, 1765, 1701, 1993, 934, 51, 806, 1074, 186, 574, 755, 812, 1472, 2651, 175, 2378, 2722]. eggspots [684]. eggvinii [2711]. EIFAC [537]. Eigenmann [2832, 3454]. eight [582, 3208, 2360, 2193, 2455, 2593]. Eighth [409, 436]. Ejaculate [2908, 3013, 3011]. ejaculation [2695]. elasmobranch [2993, 2645]. Elasmobranchii [1745, 2631, 2766]. elasmobranchs [2837]. elastomer [639, 1100, 2674, 2376, 2012]. Elbe [304, 2525, 1572, 1119]. Electric [2025, 2789, 171, 2878, 1648, 3298, 3092, 2243]. Electric-induced [2025]. electric-organ [2243]. electrical [171]. electricus [2789]. Electroencephalogram [1556]. Electrofishing [1362, 1340, 1320, 1276, 2273, 1554, 1678]. Electrofishing-induced [1362]. electrolyte [3378]. electromyogram [1503, 185]. electromyograms [1126]. electron [94, 543]. Electronic [3142, 1869, 1879]. electrophoretic [1128]. Electrophorus [2789]. electrophysiological [1501, 941]. electroreception [410]. Electrosensory [161]. electroshocking [1378]. eleginoides [2207, 2033]. Eleginops [1716]. Elemental [2361, 2959, 1763, 3353]. elements [1127, 465, 824, 2036]. Eleotridae [1684]. Eleutheronema [3276]. Elevated [3371, 2684, 2906, 286, 2179, 818, 1062, 2366, 1441, 3018, 2485, 367]. elevation [373, 1484, 503]. elevational [803]. eliciting [3181]. elicits [2695]. Elimination [362]. ELISA [1981]. Elizabeth [1901]. Ellesmere [340]. Elliott [1402]. ellipes [2768]. Ellopostoma [3428]. elongate [1445]. elongatus [2964, 3467, 2229, 3335]. eluates [768]. elver [1813]. elvers [122, 3230, 2454]. embayment [822]. embed [1679]. embedded [2918]. embeddedness [1240]. Embiotoca [3168, 2282, 869]. Embryo [3124, 1573, 2419, 353, 1381]. embryogeny [2392]. Embryonic [3038, 3267, 2808, 554, 2443, 1112, 2157, 2631, 1831, 1756, 970, 2842]. Embryos [2174, 1384, 1907, 360, 1463, 2592, 733, 1378, 1076, 1081, 693, 1027, 2223, 3378, 2071, 3445, 1164, 1090, 1441, 2559, 320, 2491, 464, 768, 3362, 2843]. emerged [3368]. emergence [1243, 15, 310, 2843, 1030]. emergent [1825]. emerging [3476]. emersion [454]. emissions [209]. emperor [947, 294, 1011]. emphasis [1382, 1737, 190, 1770, 3373, 2425, 2875, 1638, 2028]. Empirical [64, 2907, 2710, 1771]. employing [3305]. empty [1676]. Ems [914]. enable [1107]. Encheloclarias [182]. encircled [204]. enclosed [3367]. Enclosing [3405]. enclosure [2958]. encoding [2593, 1441, 769]. encountering [1114]. encrasicolus [1982, 2496, 3126, 2355, 3106, 3462, 2048]. end [1781]. endangered [1086, 1948, 73, 3344, 2517, 2978, 1563, 3066, 2671, 1067, 1770, 1122, 1170, 2201, 2804, 1849, 423, 2444, 2738, 3336]. endemic [1203, 2947, 2766, 3118, 2585, 703, 2708, 2129, 2804, 1286, 2913]. endemism [1810]. Endocrine [1281, 1323, 2121, 1438, 1089]. endocrinology [3457, 481]. endocytosis [1681]. endogenous [724]. endoparasite [175]. endoparasitism [2369]. endoscopic [1787]. endostyle [2887]. endothermy [2498]. Endurance [1560, 2245, 2, 1449, 113]. endure [1466]. energetic [1584, 2016, 3184]. Energetics [1484, 858, 2600, 2668, 3058]. Energy [3126, 2686, 755, 1338, 2048, 1640, 540, 188, 3235, 729, 2394, 2384, 830, 2942, 146, 845, 685, 440, 2687, 1365, 1361, 467, 288, 29, 1498, 1386, 1033, 2]. England [19, 2751, 796, 789, 623, 1955, 2884, 2567]. English [1323, 3255, 2455, 2012, 1603]. Engraulidae [2618]. Engraulididae [123]. Engraulis [1982, 3379, 2496, 3126, 2355, 1882, 2291, 3106, 3462, 2048, 2478, 3251]. Enhanced [956, 53, 118, 1761, 670, 1490]. enhancement [844, 1582, 3399, 704, 1592, 1958, 1723, 1801, 1555]. enhances [760]. enigmatic [3428, 2520]. Enisey [2732]. enlargement [3128]. Ennell [3154]. Enos [403]. enriched [583, 173, 3310]. enrichment [1192, 2555]. ENSO [546]. enterocytes [3083]. entière [1405]. Entomacrodus [3210]. Entosphenus [2793]. entrainment [1430]. entry [2627, 3261, 929, 3461, 2454, 1169]. envelope [150]. Environment [2638, 3478, 1653, 1292, 1578, 583, 3368, 894, 784, 2213, 2878, 1949, 1952, 723, 1936, 3256, 1367, 1166, 2962, 1884, 537]. Environment-related [2638]. Environmental [3234, 808, 1192, 1482, 143, 2468, 300, 1888, 761, 2788, 1171, 30, 979, 1608, 3389, 2496, 2592, 2294, 2692, 1478, 242, 1435, 1434, 397, 929, 1075, 2276, 254, 1316, 1096, 781, 1974, 1228, 1261, 379, 1542, 1231, 1973, 815, 1853, 512, 1625, 922]. Environmentally [899, 919]. environments [2346, 327, 3379, 444, 2246, 1274, 3052, 3408, 3407]. Enzymatic [2938, 3149, 1938, 1164]. Enzyme [1015, 1513, 1054, 1488, 1162, 1353, 427, 898, 14, 2691, 1688, 1940, 48, 348]. enzyme-linked [348]. enzyme-staining [1688]. enzymes [2400, 3208, 2158, 2186, 3336]. eos [1872, 3190]. eosinophilic [2542, 344]. epaulette [988]. epibiota [2776]. epidemic [3161]. epidermal [362, 3456, 825]. epidermis [2639, 1179]. Epigean [1137, 1088]. Epinephelides [2666]. epinepheline [1505]. Epinephelus [2643, 1552, 3039, 3152, 1922, 1688, 2161, 3108, 2612, 775, 2578, 1178, 1713]. epinephrine [2073]. episcopi [2678]. episodes [546]. epithelia [1321, 88]. epithelial [1847, 1513]. epithelium [632, 2330, 1737, 1817]. equal [178]. equal-mass [178]. equations [1815]. Equatorial [3228, 3062, 1021]. Equivalent [3311]. erectus [3374]. Erethistes [1141]. Erethistidae [1141]. erethistids [1141]. Ericht [1539]. erinacea [3386]. Erne [563]. erodes [1292]. Erpetoichthys [2975]. Errata [2410, 2512, 2829, 2830]. Erratum [136, 308, 505, 593, 594, 595, 567, 666, 832, 833, 1705, 1778, 1779, 1823, 1877, 2076, 2177, 2462, 2538, 2626, 2654, 2702, 2799, 2863, 2889, 3144, 3164, 3187, 3338, 3442, 3481, 3145, 3165, 3188, 2178]. error [957]. Erythrinidae [2292, 573, 3459, 3436]. erythrinus [1619, 1597]. erythrocyte [1836]. erythrocytes [771, 1016, 279, 804, 3250, 2117]. erythroid [1757]. erythrophthalmus [2315]. erythropterus [2144]. erythrospilus [3311]. Escape [2566, 2262, 1919, 2063, 2155]. escaped [3139, 1598]. Escapees [1605]. esculentus [71]. Esox [2973, 2936, 2903, 1757, 2143, 1399, 3253, 2187, 903]. essential [1243, 2557].

EST [3237]. EST-linked [3237]. established [607, 2660, 3061]. establishing [3025]. Establishment [2279, 2309, 2381, 2436, 2465, 2492, 2411, 3150, 2513, 3046, 1513, 125, 2743, 2749]. esterified [466]. esters [785]. estimate [545, 1894, 2497, 64, 215, 727, 3313]. estimated [1998, 58, 469]. Estimates [3229, 585, 3315, 3386, 327, 912, 3379, 256, 1631, 3171, 857, 2159, 1721, 726]. Estimating [2031, 1515, 1678, 1264, 2685, 2582, 3394, 627, 388, 3232, 2701]. Estimation [2406, 1444, 827, 3359, 667, 698, 980]. estimations [1430, 541]. estor [2114, 2114]. estuaries [906, 1904, 1478, 3082, 2603, 3467, 2156, 228, 8, 2236, 2680, 922, 1402]. Estuarine [1007, 808, 1407, 1347, 2812, 2725, 2419, 3289, 2007, 3029, 1759, 1357, 282, 2567, 918, 1148, 919, 2182, 1736, 1124, 1801, 2454, 1099, 3468]. estuarine-dwelling [1347]. Estuary [1314, 913, 670, 1632, 2994, 262, 909, 2641, 912, 2496, 3233, 2096, 546, 2660, 2222, 397, 2287, 2167, 119, 2285, 3232, 1709, 2615, 1050, 2238, 1153, 907, 914]. ethanol [1356]. Etheostoma [2165, 3431, 2794]. Ethical [1964, 3476]. ethics [3477]. Ethiopia [2929]. Ethologische [456, 477]. Ethologists [456, 477]. Etmopteridae [2995, 3000]. Etmopterus [2995, 3000]. Etrumeus [2087]. Etruscan [2713]. Eubothrium [2586]. Eucinostomus [2151]. eulachon [2670]. Eumicrotremus [2711]. Eupallasella [2056]. EUR [1497, 1775]. Eurasia [2146]. Eurasian [840, 2731]. Europe [1497, 2535, 2717, 3243, 245, 3432, 3211, 828, 461]. European [2978, 2255, 3234, 761, 984, 537, 115, 263, 97, 2739, 2009, 2346, 3246, 3206, 311, 601, 3229, 3243, 3235, 977, 2481, 1608, 961, 2496, 180, 291, 3457, 3344, 2355, 259, 2679, 304, 2341, 218, 160, 818, 2509, 2674, 2733, 1063, 2525, 190, 3211, 3183, 59, 2690, 3461, 3238, 1565, 111, 649, 1464, 2246, 1222, 3209, 1940, 1938, 2258, 2259, 531, 732, 978, 1399, 323, 740, 258, 517, 1621, 3237, 624, 2415, 417, 1571, 2210, 1900, 2376, 3109, 3244, 2556, 1230, 2695, 1372, 3362, 609, 1449, 512, 2738, 1155, 939, 1490, 3154, 3099]. euros [1405]. euryhaline [540, 128]. eurystoma [2804]. eurytopic [3395]. euspermatozoa [1411]. Euthynnus [629]. eutrophic [764, 3422, 3242, 675, 2399, 1322]. eutrophication [3426]. evacuated [2563]. Evacuation [1136, 402, 545, 1894, 1893, 2876, 3377, 3157, 1437, 2700, 305, 3397, 178, 1444, 726, 385, 386, 770, 1480]. evaluate [1534, 1984, 64]. evaluated [2996, 1331]. Evaluating [2273, 1691]. Evaluation [3193, 1681, 1100, 2674, 2644, 1905, 2376, 1878, 1830, 3400, 957, 171, 639, 1972, 1611, 2920, 1427, 1357, 1994, 411, 1900, 501, 1981, 3239]. evaluations [196]. even [653]. even-broodline [653]. event [1170]. events [2642, 3124]. Everglades [2747, 3178]. Evidence [2825, 1704, 2558, 965, 3170, 712, 503, 2314, 560, 3122, 702, 3047, 1309, 116, 460, 1380, 2289, 2473, 2567, 2136, 2148, 782, 2290, 1941, 3065, 1155, 2345, 2721, 2424, 947, 3388, 3069, 500, 27, 3017, 2756, 3095, 686, 2015, 2793, 3066, 2759, 1818, 2858, 1217, 2896, 2402, 2837, 599, 2708, 705, 2329, 254, 3287, 574, 2325, 843, 1781, 1543, 2066, 566, 1432, 267, 1911, 3176, 1602, 1939, 892, 1381, 3236]. evidenced [2547, 1358]. Evolution [3096, 410, 3035, 2979, 958, 3429, 1194, 277, 3166, 3295, 3345, 2150, 1184, 2524, 1781, 2477, 1232, 1600, 2916]. evolutionarily [2728, 608]. Evolutionary [3169, 1558, 153, 2796, 3043, 2143, 2415, 3146, 3057, 2544, 3324, 3342, 3412, 762]. Ewart [1404]. examination [484, 1857, 149, 1679, 3404]. examine [2261]. examined [1567]. examining [1781]. example [2877, 1983, 3344, 2701]. excess [985]. exchange [1765, 1016]. exchanging [2330]. exclusion [2418]. exclusive [2527]. excretion [2356, 2684, 2942, 818, 850, 3131, 279, 1536]. excursion [2973]. exercise [1524, 2623, 985, 2245, 1082, 1365, 1973, 2683]. Exeter [2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513]. exhausted [1363]. exhibits [3031]. existence [3014, 1231, 1458]. existing [482]. exogenous [3093, 441, 879]. exopolysaccharide [2902]. exotic [3010, 2647, 290, 765, 1000, 2896, 548, 1372, 1638]. expanded [1894]. Expanding [2866]. expansion [3044, 2717, 1070, 2372, 1572, 2854]. Expedition [282]. expenditure [2908]. expenditures [2686]. experience [1172, 1192, 2046, 1806]. experienced [1041, 553, 1330, 2452]. experiencing [413]. experiment [2086, 170, 202, 2958]. Experimental [1581, 910, 684, 142, 472, 216, 217, 525, 192, 1204, 2052, 1118, 396, 3021, 1994, 1313, 991, 3056]. Experiments [3197, 98, 1568, 411, 1597, 1599, 2379]. explain [2745, 321, 1229]. Explaining [1246]. explanation [725]. explanations [803]. explicit [1746, 2422]. exploit [2011]. exploitation [1935, 236]. exploited [667, 1783, 143, 2578, 2629]. exploration [2755]. Exploring [3220, 2348, 3413]. exposed [3282, 1485, 189, 1461, 2195, 2165, 2920, 1002, 2257, 2413, 2801, 2614, 476, 83, 3128, 367]. exposing [2571]. Exposure [1292, 3155, 350, 438, 2739, 2265, 284, 128, 2157, 2392, 272, 818, 3059, 2917, 4, 2997, 1296, 2873, 3018, 995, 2399, 2938, 1416, 2155, 2211]. Expressed [18, 2924]. Expression [987, 1821, 752, 3274, 33, 3343, 3039, 325, 2373, 3217, 1612, 527, 1212, 3038, 3256, 1744, 2677, 745, 990, 2923, 474, 1441, 2559, 2715, 2533, 1941, 1149, 2859]. extant [197]. extenders [490]. extension [1398]. Extensions [245]. Extensive [873, 1827, 2328, 2143, 2256, 2148]. extent [2518]. extenta [786]. External [2045, 3130, 1847, 2374, 842, 2069, 3111, 1005, 149, 2544, 2228, 533]. Externally [1003, 2044, 1293, 533]. extinct [1833, 2517]. extracellular [1694]. extract [1077]. extraction [2727]. extracts [1463, 3196, 768]. extraordinary [2497]. extreme [2234, 1669, 238, 1638]. Extremely [2712, 2437]. extremes [1231]. eydouxii [1521]. Eye [811, 718, 3038, 2607, 565, 1914, 3058, 1451]. eye-specific [3038]. eyed [3130, 2388, 2486, 2441, 343]. eyed-egg [2441]. eyelid [3271].

F [3479, 1496, 1651, 3322, 1405, 1494, 1463, 2384, 941, 676, 3346]. F-prostaglandins [941]. F-series [676]. F.-W [1496]. Fabricius [2711]. facial [370]. facilitates [52]. facilitation [2632]. facility [1605]. facing [906]. factor [3274, 1547, 2759, 2728, 978, 2533, 2288, 2982]. factor-binding [2728]. factor-I [3274]. Factors [162, 2994, 1982, 1561, 1135, 2627, 3180, 3376, 706, 1319, 822, 808, 3034, 3220, 30, 100, 3389, 2592, 1434, 397, 929, 967, 2276, 434, 1316, 2963, 2715, 1354, 300, 3055]. facts [614]. facultative [2655]. FADs [289]. fails [454]. Failure [800, 446]. Falkland [1084, 664, 2033, 1477]. fall [3204]. fallai [2498, 2906]. fallax [2328, 2738, 2978]. falling [1159]. familiar [1297, 3247]. Familiarity [1197, 2679, 2026, 1342, 1234]. Familiarity-biased [2679]. families [2816, 3471, 1175, 444, 1375, 3293, 2370]. family [2710, 293, 1733, 3326, 2517, 1596, 3172, 356]. family-specific [293]. Fangzheng [2983]. far [2744, 281]. farm [2499, 2115, 2845, 708, 2205]. farm-reared [2845]. farmed [805, 1584, 2157, 2392, 3223, 1617, 1590, 1744, 1618, 1595, 1621, 1691, 463, 2269, 2581, 206, 1601]. farming [1945, 1693, 2205]. Faroe [2587]. Fas [2293]. Fas/FasL [2293]. fasciatum [1489]. fasciatus [2353, 2486, 2855]. fasdatus [697]. FasL [2293]. Fast [2716, 2262, 2296, 2319, 2684, 2899, 1540, 3100, 68, 2587, 2245, 2895, 3363, 121]. fast-flowing [2319]. Fast-growing [2716, 2684, 3100, 2587]. fast-start [2262, 2296]. fast-starts [121]. Fasting [1913, 398, 1547, 1798, 2533]. fat [3048]. Fate [2561]. fathead [404, 3072, 2021, 3059, 481, 2277, 1307]. father [3168]. Fatigue [2111, 1160]. Fatty [2083, 459, 263, 1633, 466, 2419, 170, 2242, 1794, 3089, 3078, 1866]. fatty-acid [3089]. fauna [607, 92, 2827, 2169, 918]. favours [3070]. Feasibility [1613]. features [315, 3333, 839, 1980, 2689]. fecund [602]. Fecundity [112, 24, 3315, 2692, 2062, 2998, 1589, 2767, 2259, 1796, 2500, 2288, 1713, 670, 3257, 421]. fed [1885, 2203, 2214, 12, 519, 3328, 342, 173, 891, 879, 2190]. feed [540, 1242, 894, 339, 1890, 2058, 2200, 3150, 20, 1136, 2598, 939, 365, 940]. feed-deprivation [940]. Feeding [346, 57, 1672, 424, 2941, 752, 2098, 2053, 482, 55, 743, 2030, 2085, 2896, 1549, 2891, 2507, 2393, 1686, 1166, 1445, 2003, 3281, 2054, 3390, 2569, 209, 1439, 188, 907, 2426, 2215, 1418, 2932, 2388, 24, 262, 3093, 1658, 1512, 297, 884, 2391, 544, 894, 1147, 1203, 1660, 580, 1207, 1646, 1790, 3230, 2555, 3469, 1280, 444, 1375, 71, 2013, 2630, 2484, 2406, 2999, 241, 1952, 1437, 3202, 693, 560, 2056, 2700, 1278, 1048, 2552, 1017, 552, 2493, 3208, 1483, 2284, 717, 950, 1789, 3150, 343, 1740, 2416, 2417, 3385, 2801, 3300, 1708, 169, 1828, 517, 2216]. feeding [1066, 1421, 1724, 2423, 524, 2665, 2114, 1260, 2602, 591, 1736, 2212, 879, 1693, 1542, 1424, 2875, 2957, 754, 3058, 2088, 2064, 400, 921, 207, 1782, 1099, 2028, 2386, 3352, 486, 507, 841, 3321]. feeding- [2028]. feeds [1242, 1512]. Feigning [1702]. Female [2104, 192, 1517, 2554, 2571, 2094, 3363, 3466, 514, 1267, 1888, 2643, 459, 975, 97, 779, 1501, 1133, 3254, 2504, 2599, 494, 631, 684, 692, 3424, 1826, 773, 2998, 3398, 2899, 414, 3011, 1811, 340, 1993, 1062, 2366, 2008, 2936, 3311, 4, 986, 1221, 633, 499, 2289, 1710, 574, 1337, 56, 2695, 1352, 348, 1319, 3332, 945, 451, 237, 368, 1713, 220, 570, 1881, 3356]. Female-biased [2094]. female-dominant [1710]. female-to-male [2643]. Females [838, 1633, 2438, 1188, 345, 321, 238]. feminization [677]. feral [1921, 3355]. ferox [2317]. Ferraris [1495]. Ferreira [1843]. ferritin [562]. ferruginea [1848]. Fertilisation [1237]. Fertilization [42, 389, 1795, 2617, 2370, 1365, 2429, 1870, 2379]. fertilizing [2664]. fibre [1417, 110, 2899, 2245, 1622]. fibres [2245]. fibroblast [1369]. fibroblast-like [1369]. fibroblastic [1513]. fidelity [2771, 1474, 2936, 265]. Field [3465, 931, 2052, 1357, 2402, 545, 2424, 3351, 187, 2996, 2878, 1718, 98, 164, 3248, 705, 371, 1994, 1444, 411, 3159, 650]. Field-based [931]. field-collected [187]. fifteen [1163]. fifteen-spined [1163]. fight [894]. fighting [391, 584, 1388]. figures [614]. filament [3471, 431]. filamentosus [3375]. filaments [2300]. filial [1053, 1225, 813]. filling [17]. films [1834]. filter [71]. fin [3359, 2780, 2929, 1648, 1797, 861, 2051, 2644, 869, 2206, 3064, 1723, 2456, 3046]. fin-beat [869]. final [1387, 2899, 319, 313]. Fine [2322, 3401, 2728, 671, 1739, 774]. Finfish [1654, 1801]. fingerprints [1763]. finite [1834]. Finland [1516, 675, 2506, 3162]. Finlets [1895]. Finn [3406]. finned [2150]. Finnish [138]. finnmarchica [2985, 1729]. fins [89, 2089, 838]. firebrow [2911]. First [3294, 2597, 1727, 1906, 1521, 1398, 2009, 2911, 2247, 2457, 851, 2599, 662, 2383, 1397, 2171, 2886, 2250, 180, 449, 3215, 733, 902, 2850, 2015, 2647, 129, 2525, 2822, 787, 949, 283, 2914, 3375, 452, 3063, 2255, 203, 464, 1092, 3374, 3161, 1777, 24, 1750, 829, 580, 291, 1674, 64, 1890, 3469, 1280, 2896, 3461, 1594, 950, 2470, 266, 1620, 28, 3253, 1828, 591, 2003, 879, 838, 3214, 2379, 1881, 1396]. first-feed [1890]. first-male [2379]. first-time [24, 1881]. First-year [949]. Firth [3186]. Fish [1487, 1428, 1902, 1963, 1901, 1990, 2019, 2049, 2078, 2101, 2127, 1991, 882, 906, 1523, 1196, 2790, 911, 1244, 894, 1404, 3322, 1773, 1775, 1025, 1219, 2337, 1652, 675, 1316, 919, 1909, 2091, 1424, 423, 921, 1531, 1532, 362, 699, 827, 410, 2563, 1777, 1558, 1738, 3444, 1085, 2081, 3282, 1908, 1238, 1632, 2994, 1415, 2777, 3073, 569, 764, 2404, 607, 3388, 2043, 311, 2877, 1712, 3035, 2388, 1534, 230, 478, 555, 2897, 1409, 3270, 172, 1087, 3427, 2197, 2952, 2621, 3034, 3220, 2975, 3085, 2725, 274, 758, 1935, 1041, 1072, 1933, 2813, 2714]. fish [606, 912, 817, 2735, 1297, 1750, 2023, 101, 498, 79, 1165, 2263, 2179, 1732, 2929, 289, 3007, 1014, 1189, 1353, 3371, 2499, 1360, 1904, 1383, 1463, 2886, 1175, 2202, 614, 1975, 3477, 295, 2582, 326, 2684, 3053, 3457, 387, 307, 1585, 92, 1422, 3360, 621, 2384, 2014, 2294, 546, 2061, 725, 635, 3344, 997, 2034, 1250, 2827, 1185, 2093, 2755, 144, 2041, 1251, 425, 2993, 696, 1044, 1187, 1683, 2947, 898, 913, 1648, 168, 856, 1771, 3050, 2783, 519, 914, 2406, 2633, 1478, 3082, 2840, 3249, 1411, 3399, 2194, 1146, 248, 588, 723, 242, 227, 3289, 2085]. fish [1992, 3475, 1211, 2303, 2482, 1971, 2845, 324, 2058, 2759, 391, 584, 693, 1743, 3299, 3473, 1769, 70, 2027, 1915, 2281, 3298, 708, 1213, 1214, 1215, 1844, 1617, 397, 722, 3049, 934, 1217, 1759, 1734, 1107, 2896, 51, 68, 1218, 3030, 3295, 1783, 2841, 2959, 2553, 2493, 1593, 1696, 4, 3038, 1357, 863, 2587, 1255, 2402, 2656, 2046, 599, 2968, 857, 734, 502, 1918, 2639, 3199, 3391, 2677, 1899, 3150, 2276, 724, 1978, 1537, 1954, 2763, 1223, 3142, 2274, 3373, 1160, 1955, 2455, 2884, 2593, 2118, 157, 2156, 142, 1595, 1096, 2169, 3453, 917, 1519, 479, 2167, 647]. fish [275, 918, 1022, 1974, 9, 2047, 3117, 3201, 213, 2697, 441, 3365, 1800, 1944, 2715, 755, 980, 772, 2182, 1367, 3018, 364, 72, 411, 1166, 2774, 824, 3436, 1656, 2651, 26, 2798, 2748, 2655, 3281, 3177, 1274, 2338, 2628, 1542, 3250, 2190, 3192, 1599, 2036, 2142, 2811, 2729, 2054, 2173, 3313, 3135, 2488, 2591, 2856, 774, 257, 2701, 822, 2231, 2205, 810, 2939, 1603, 2680, 1883, 3012, 3390, 2243, 2969, 1050, 2629, 3156, 2930, 2531, 3303, 2689, 3409, 3257, 3403, 2141, 3352, 405, 2874, 507, 835, 3114, 152, 1780, 1404, 3406, 3407, 3320, 3311]. FISH-BOL [3114]. fish-community [3249]. fish-gut [2783]. fish-habitat [1712]. fish-specific [3035]. fished [2034, 1770, 605]. Fisher [1651]. fisheri [3451]. Fisheries [537, 1010, 1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 1989, 2381, 2436, 2465, 2492, 1772, 3323, 1651, 3478, 3405, 2411, 1006, 2513, 906, 1198, 911, 3473, 611, 1592, 3474, 3310, 1954, 3476, 2795, 603, 1576, 2672, 2503, 3410, 1651, 1652, 3478]. fishers [890]. fishery [2758, 635, 2766, 854, 1957]. FISHES [1405, 3479, 1181, 1, 1628, 1497, 1706, 2279, 2309, 2381, 2436, 2465, 2492, 1947, 277, 1007, 3326, 2411, 1495, 1494, 922, 2513, 315, 540, 2563, 409, 436, 877, 1964, 753, 907, 1557, 808, 935, 2426, 2065, 2580, 1605, 2757, 3255, 860, 1712, 3367, 664, 1580, 2890, 3127, 3314, 1805, 1413, 1488, 2532, 27, 1948, 2336, 2751, 462, 2359, 958, 2234, 1731, 40, 862, 963, 3260, 730, 1194, 449, 2582, 2619, 110, 2275, 468, 1476, 64, 2900, 3452, 2744, 2199, 2165, 3175, 290, 522, 255, 3166, 3050, 2517, 2749, 3068, 674, 3464, 2361, 2588, 2025]. fishes [2753, 3339, 1371, 84, 2147, 1564, 2577, 613, 771, 2603, 3212, 2412, 2879, 1040, 2027, 1550, 1752, 941, 2868, 1515, 1000, 608, 508, 2005, 2466, 1836, 1879, 2284, 1520, 2775, 1075, 2587, 2397, 2150, 2360, 2160, 2193, 2776, 1046, 2696, 3310, 2981, 2395, 820, 3296, 282, 2416, 2417, 2473, 475, 741, 1257, 369, 1554, 2814, 467, 577, 2335, 1193, 1679, 2414, 3297, 279, 2894, 1258, 3340, 2520, 2325, 853, 624, 2852, 1781, 2114, 2955, 1120, 602, 3476, 2971, 2220, 2399, 1229, 2425, 1873, 1809, 2540, 1661, 1574, 2804, 2701, 1286, 769, 1600, 3116, 1129, 2752, 3055, 3114, 1101, 561]. fishes [1329, 2569, 207, 428, 2916, 802, 163, 1493, 3321, 1402]. Fishes/Dynamics [1494]. Fishing [1824, 1773, 171, 3242, 909, 613, 1978, 647, 602, 1542, 240]. fishway [3184]. fishy [233]. Fistularia [1521, 852, 2372]. Fistulariidae [2372]. Fitness [164, 2104, 3434, 2759, 3462, 1590, 1122]. Fitness-related [164, 1590]. five [2872, 659, 2621, 1733, 2846, 3292, 916, 734, 1985, 1934, 2673, 727, 726, 1883]. five-spotted [1985]. fjord [1672, 3367, 1634, 66]. fjords [1690, 2912, 2585, 1421, 1479]. flagfish [2867]. Flanders [2752]. flatbed [3212]. Flatfish [2788, 346, 2940, 2846, 3397, 2495, 2244]. flatfishes [2106, 2516, 87, 2500]. flathead [2934, 3291, 1352, 2689]. flats [915, 2284]. flavicans [1489]. Flavobacterium [3130]. flavolineatus [2152]. flee [1133]. Flemish [1079]. flesus [1672, 2086, 2821, 2484, 1460, 2507]. Flexibility [3303, 2084, 1130]. Flexible [496, 3074]. flexion [2387]. floating [2813]. flock [1203, 277, 378]. flood [1908, 2445, 2299, 1958]. flooded [2813]. flooding [539]. Floodplain [2601, 1085, 1908, 2092, 860, 2813, 529, 74, 2588, 1550, 1000, 2968, 2861, 1096, 1022, 3262, 3192, 1059]. floodplains [2883]. Florence [2100]. Florida [3010, 162, 2721, 2641, 3263, 1348, 2747, 3178, 337]. floridae [2867, 1482]. floridanus [1806]. flounder [2608, 2086, 99, 1645, 2443, 297, 1430, 2073, 55, 743, 1249, 1660, 1609, 2821, 325, 3136, 2484, 519, 218, 1151, 704, 2730, 1507, 1719, 2869, 151, 745, 990, 2542, 2507, 3385, 298, 2792, 1392, 2476, 3455, 3277, 1173, 240, 2064, 3200, 507]. Flow [2920, 792, 1320, 22, 2723, 855, 2340, 1183, 3235, 1347, 1207, 1584, 1646, 256, 636, 2793, 2341, 861, 3443, 1743, 2699, 1510, 2182, 3037, 2299, 2285, 3319, 76, 3222, 1050, 1153]. flowing [2319, 1540, 2950]. flows [2234, 1663, 3205]. Fluctuating [2770, 694, 3462, 153, 636, 1461, 1447, 2604, 76]. fluctuation [3277]. fluctuations [803, 3251]. fluid [1163, 1069, 2509, 760]. fluidity [804]. fluke [3058]. flume [1675]. flumes [2681, 76]. fluorochrome [2969]. Fluvial [1706, 409, 436, 877, 1382, 148, 1769, 2817, 2470, 485]. fluviatilis [1633, 2558, 1545, 2438, 2013, 3004, 2297, 1080, 2159, 1551, 798, 2444, 2123]. flying [958]. foetus [2304]. fokiensis [2521]. foliaceus [180]. folklore [3409]. follicular [2834, 1841]. follower [2220]. following [1643, 2265, 525, 2534, 1242, 1512, 2821, 2993, 1350, 2828, 2256, 20, 1500, 2454, 365, 1153, 377]. Foltz [3479]. fontinalis [2905, 2756]. Food [2907, 719, 3452, 139, 3357, 483, 1333, 3373, 169, 2335, 1304, 1059, 54, 413, 209, 545, 954, 1894, 392, 2972, 2070, 500, 2387, 297, 3371, 1863, 191, 1631, 470, 522, 2469, 1152, 235, 473, 25, 1311, 685, 440, 2767, 1019, 2051, 75, 2181, 393, 1176, 1114, 716, 2968, 1740, 2416, 2417, 140, 2167, 31, 288, 298, 878, 3252, 232, 809, 772, 2665, 976, 2774, 1348, 1466, 727, 3281, 690, 1761, 1028, 2957, 1640, 1130, 406, 939, 1881, 377]. food-chain [2774]. Food-dependent [2335]. food-stocking [2972]. forage [2877]. forager [2039]. Foraging [165, 3005, 663, 2475, 3009, 1454, 2866, 844, 1192, 1200, 2780, 758, 784, 2893, 1667, 1197, 126, 867, 1811, 993, 1118, 2466, 1219, 2160, 2954, 2432, 1440, 1224, 1182, 811]. forced [1643, 1500]. forebrain [3354]. foreign [362, 2460, 562]. Forest [1955, 2813, 1353]. Foreword [1928]. Fork [12, 536]. forktail [964]. form [1447, 2249, 3472, 3071, 2243]. Formation [3441, 2106, 977, 1383, 2782, 2845, 3004, 1899, 3081, 2401, 660, 2969, 2722]. forms [3305, 601, 1056, 1020, 1667, 202, 2226, 1311, 1969, 1104, 2987, 1462, 1401]. Forsgren [3320]. forskolin [1645]. Forsskål [621, 2612]. Forth [913]. fossil [3414]. fossilis [2717, 246, 1295]. fouling [3313, 533]. found [1810, 3372]. founder [2719, 1170]. Four [1799, 356, 3305, 2388, 3316, 1488, 2532, 91, 274, 291, 1860, 2724, 1102, 559, 3125, 3280, 2495, 1046, 711, 343, 8, 2117, 1499, 3417]. four-eyed [2388, 343]. four-spined [3417]. Fourth [502]. Fowler [3042]. Foyle [1983]. fraction [2920]. Fractionation [2586, 479]. fragment [1754]. fragmentation [2961]. fragments [1158]. framework [769]. France [1901, 3170, 1775, 3234, 637, 3077, 78, 2756, 1360, 2287, 3238, 111, 1656, 461, 669]. Francis [1406]. Francisco [2472, 3401, 1896, 2222]. Fraser [2726]. Free [589, 1472, 1525, 1814, 588, 1831, 3327, 1794, 1182, 2628]. free-embryonic [1831]. free-ranging [588, 1182]. free-swimming [2628]. freezing [2821, 1046, 271]. Frémur [3234]. frenatus [3433]. French [7, 944, 2319, 659, 2143, 167, 3065]. frequency [64, 1790, 869, 1760, 3308, 991, 1601, 1676]. frequency-based [1601]. fresh [1069, 850, 329, 1296, 782]. Freshwater [1497, 614, 282, 1495, 2476, 1653, 540, 1704, 2084, 569, 2346, 607, 2890, 965, 2713, 3220, 2557, 2751, 302, 1689, 2419, 2359, 40, 3260, 2496, 3160, 2893, 2034, 575, 2013, 2361, 2588, 2753, 242, 520, 2147, 560, 1915, 2921, 23, 1528, 3461, 2360, 599, 2246, 3199, 3310, 3418, 433, 2040, 2524, 2414, 3340, 2415, 812, 2182, 3361, 1656, 3159, 82, 838, 2117, 3135, 1575, 2545, 3074, 782, 2322, 3116, 2752, 1416, 131, 3409, 2805, 3403, 793, 3479]. freshwater-feeding [2013]. freshwater-resident [1689, 2040]. friedrichsthalii [1702]. fright [2442]. frog [3046]. front [2634, 3287]. frontal [3014]. FRSC [645]. fructose [474]. frugivore [1353]. fry [3193, 1685, 2362, 2445, 1984, 2499, 15, 480, 429, 1562, 310, 2772, 1002, 1310, 1393, 2003, 501, 1500, 3056, 3336]. Fryken [3139]. FSH [2677]. FSH- [2677]. fuel [2920]. Fukushima [3455]. fullness [809]. fulva [67]. fulvescens [3094]. Function [1250, 1349, 2979, 2332, 3033, 1746, 1209, 85, 401, 552, 1180, 3335, 1263, 948, 2595, 1390]. Functional [888, 1360, 2403, 2842, 3380, 3304, 418, 1523, 3305, 3367, 2790, 3283, 2114, 2475, 556, 2916]. Functionally [273]. Functions [3321, 3271, 791, 1473, 3146, 1289, 1659, 1838]. fundic [1644]. Fundulus [2326, 3148, 2647, 1817, 3167, 3331, 3349]. fundyense [2774]. fungal [2459]. Fungi [2590, 186]. funieri [920]. furcatus [1868]. furnieri [1977, 3287]. Further [3420, 1920, 1279]. furthers [1260]. furzeri [3101]. fuscescens [2098, 1304]. fuscoguttatus [2612]. fuscus [688, 343, 2166]. Future [1901, 911, 922]. Fuxian [3038, 2431]. fuzzy [1978]. fyke [3246, 2004].

G [3479, 3410, 1772, 1773, 3321, 3406, 3407, 3320, 3411]. Gabès [2871]. Gabrielle [1159]. Gadidae [2985, 1729, 2820]. gadoid [674, 3397]. gadoids [545, 1894, 1893, 2877, 90]. Gadus [1804, 2227, 2771, 1694, 2897, 2924, 2579, 2446, 2235, 2848, 3189, 3328, 2663, 2126, 2342, 1966, 1588, 2834, 3081, 3285, 2321, 2541, 3284, 3300, 2838, 2401, 2264, 3014, 2539, 2343, 2288, 1830, 1837]. gag [2876, 1168, 2348]. gaimardianus [2709]. gaining [2583]. Gait [2282, 869, 2814]. galanin [2134]. galanin-like [2134]. Galapagos [343]. galaxias [2694, 2459, 2204, 2419, 2694, 2459, 3299, 776, 1160, 378]. galaxiid [2772, 608]. Galaxiidae [2459]. Galaxioidea [1379]. galectin [825]. Galeus [549, 2543]. Galicia [228]. gall [1817]. gallbladder [1446]. Gallego [1901]. Gambusia [3077, 3021, 3013, 2804]. gamete [2908]. gametes [42]. gametic [414]. gametogenesis [1609, 2440, 4]. gammarid [1987]. Gander [3045]. Gape [2264, 1638, 2417]. gar [3036]. garden [3351]. gariepinus [1859, 728, 2935, 3016, 2969]. Garman [2840]. Garonne [2841]. Garra [1519]. Gaspé [2637]. Gasterosteus [1842, 2121, 3254, 3084, 3422, 2460, 3424, 3416, 2089, 3413, 2226, 3472, 3423, 1253, 3071, 3421, 2413, 3425, 3418, 2950, 2323, 3420, 2668, 3426, 1234, 1235, 3400, 3412, 3306, 3419, 3417].

Gastric [402, 1894, 385, 386, 770, 545, 1893, 3377, 1749, 325, 1437, 2700, 305, 3397, 178, 1644]. gastrically [2172]. gastrointestinal [648, 2400, 427, 1438, 2593, 351, 709]. gates [307]. gather [1665]. gear [171, 1542]. Gender [1223, 1289, 1787]. Gender- [1289]. Gene [22, 1612, 1604, 2723, 855, 2340, 1347, 3274, 256, 2384, 2793, 2341, 3423, 3038, 2699, 2593, 441, 2559, 1401, 3037, 2918, 3222, 32, 356, 2916]. genera [2532, 51]. generalist [3388]. generating [3120]. generation [2720, 1573]. genes [1858, 355, 312, 3428, 274, 325, 3166, 1700, 1212, 2734, 2323, 514, 3051, 793, 1149, 2916]. Genetc [166]. Genetic [600, 2721, 689, 3036, 1872, 2723, 1726, 1579, 1457, 1923, 1933, 3017, 2735, 872, 2523, 2324, 3045, 304, 117, 2858, 2736, 1464, 599, 617, 374, 734, 2140, 1930, 1940, 1755, 1462, 439, 2986, 2731, 2329, 2855, 3287, 3119, 848, 3237, 2522, 2144, 281, 2530, 1934, 380, 1846, 2129, 1575, 2856, 1692, 1939, 1522, 49, 2925, 2526, 1031, 2340, 3220, 1935, 2347, 1430, 2536, 2317, 3171, 1610, 3344, 728, 1882, 2341, 2517, 2327, 3431, 1588, 2348, 1754, 972, 702, 1427, 1067, 2547, 677, 2143, 2705, 3041, 3238, 1109, 1719, 2927, 2898, 1938, 820, 994, 3121, 275, 2720]. genetic [1122, 3048, 1621, 229, 1228, 2707, 1170, 3064, 2792, 2539, 1827, 3172, 2540, 1358, 1374, 1458, 2545, 2854, 3115, 2322, 1286, 810, 2750, 1011, 1602, 1929, 2537, 52, 3040, 2531, 2990, 3438]. genetic-based [3171]. genetically [91, 1393]. genetics [679, 2749, 3286, 1180, 1600]. Geneva [2191, 2438, 1144]. genital [172]. genome [2789, 3035, 2921]. genomes [1933, 3221]. genomic [873, 2150]. genomics [2790, 2715]. Genotoxicity [2714, 2917]. genotype [2137, 2795]. Genotyping [641]. Genten [3322]. genus [3185, 1858, 226, 788, 601, 582, 2053, 443, 2658, 1882, 1976, 2733, 1039, 3292, 41, 1918, 1835, 711, 314, 491, 2325, 3172, 586, 2142, 2811, 1862, 1465, 3042, 2853]. Genypterus [1477]. Geographic [3107, 2675, 1843, 25, 2585, 3416, 1251, 473, 2733, 2541, 564, 2047, 1651]. Geographical [3082, 1421, 3185, 382, 751, 1098, 887, 423, 3055]. geographically [3451, 3041]. geography [2340, 1936]. Geometric [697, 1370, 2939]. Geophagus [899]. George [392, 1380]. Georges [2495]. Georgia [557, 2420, 1410, 322, 1916, 3205]. Germ [2358, 2599, 2293]. German [2061]. Germany [2070, 1106, 1674, 1285, 2136, 1119]. germinal [632]. Gerreidae [2532]. Gesellschaft [456, 477, 1009]. getting [1220]. ggrayling [609]. GH [1913, 1761, 3346]. GH-transgenic [1761]. GHR1 [2559]. GHtransgene [1761]. Giant [238, 2523, 1016, 660, 881]. gibbosus [2339, 2179, 2745, 2560]. gibel [2861, 20, 330, 365]. gibelio [2915, 2983, 2784, 2861, 1692]. Gibraltar [633]. Gibson [1406]. gigas [2883, 2523]. Gilbert [3124]. Gilchrist [2250]. Gill [3471, 2170, 3148, 1321, 399, 2302, 2257, 88, 2542, 219, 930, 3054, 45, 2054, 638, 2922, 418, 3036, 1701, 3347]. Gill-filament [3471]. gillnet [857]. gillnets [1354]. gills [3153, 2622, 1513, 2557, 3160, 2195, 218, 1999, 2997, 476, 748, 1979, 454]. gilthead [459, 2546, 1608, 2781, 1015, 3275, 2170, 173, 3150, 2965, 574, 2602, 1981].

Girard [1975, 2875]. Girardinichthys [1251, 2914]. girdle [2676, 1141, 458]. Girella [2047, 3037]. Girnock [2481]. Gironde [2287, 2841]. gizzard [162, 3095, 1043]. glacial [2326, 3044, 828]. glaciation [3432]. gladius [2358, 1763, 1816]. gland [761, 984, 3031]. glands [3027]. glandular [1742]. glanis [59]. glass [122, 1856, 3235, 3231, 3233, 3230, 1135, 3236, 823, 1880, 3245, 395, 3109, 2236, 2454]. glauca [2304]. Glaucosoma [680]. glide [2681]. Global [3478, 3312, 3161]. globally [1833]. globin [274]. glochidia [206]. Glomeruli [1148]. glossopharyngeal [370]. glucagon [648]. glucan [1659]. glucokinase [474]. glucose [2773, 1276, 636, 168, 1836]. Glugea [1677]. glycerin [1116]. glycerin-preserved [1116]. Glycoconjugate [880, 1644]. Glycoconjugates [769]. glycolysis [1725]. glycoprotein [1552, 2677, 1279]. glycoproteins [2480]. glycosylase [2321]. glyptosternine [542]. Glyptothorax [2521]. Gmelin [1727]. Gnathia [1699, 2395]. gnathiid [3050]. Gnathonemus [1648, 1390]. GnRH [2979, 1056, 2677, 2313]. go [152]. goals [611]. goatfish [2039, 2220]. gobies [222, 1863, 2394, 1132, 1005, 383, 3466, 3074, 750, 2805, 3468]. Gobiesocidae [1810]. Gobiesociformes [1742]. gobiid [2725, 1476, 1953, 2402, 3199, 955, 300]. Gobiidae [1497, 2009, 2713, 2145, 2725, 1115, 2851, 2967, 1221, 2779, 3199, 2731]. Gobio [2662, 3063, 1807, 1987, 3369, 2506, 1849]. Gobiodon [937, 3311]. gobioid [1661]. Gobioidei [2886, 1661]. Gobiomorphus [2204, 2630, 1827]. Gobiosoma [1348]. Gobius [2394, 1467, 777, 2898, 778]. goby [779, 335, 952, 2713, 2394, 3031, 431, 1976, 2921, 509, 3311, 787, 1254, 3402, 2817, 592, 2779, 433, 2731, 521, 2494, 3079, 203, 838, 124, 1028, 1395, 1236]. going [3397]. golden [2694, 2459, 3383, 221]. goldfish [3305, 3011, 547, 780, 1129, 1381, 1149, 398]. Gomphosus [3458]. Gonad [947, 401, 1052, 3206, 1633, 335, 278, 2294, 259, 2161, 1710, 657, 1661, 2604]. Gonadal [418, 1722, 749, 1618, 475, 246, 214, 1387, 2504, 1995, 795, 3463, 1769, 1062, 1980, 1505, 309, 63, 797, 1392, 945, 451, 1307, 1061]. Gonadotropin [319, 263, 649, 797, 2636, 2313]. Gonadotropin-induced [319]. gonadotropin-releasing [649, 2313]. gonads [680, 1045, 996, 1373, 1306]. gonochoristic [2666]. good [3423, 514]. Goode [2876, 2348, 2867]. goodeine [2517]. Goosander [1709]. gorbuscha [3384]. Gotland [622, 2507]. gouldii [2669]. gracilis [3158, 1810, 1986]. graded [219]. gradient [3191, 2086, 2692, 675, 577, 1974, 2047, 2299, 3112]. gradients [468, 1422, 242, 2390, 142, 1096, 3177]. grading [1710]. gradualism [3414]. graellsii [607]. grahami [3220, 3038]. Grammicolepis [1092]. Gran [642]. Grande [2650, 1495]. granular [2542]. granule [344]. granules [2542]. granulocytes [837]. grass [2394, 3031, 1254]. Graus [803]. gravel [1743]. gravel-bed [1743]. gravidity [3424]. gravity [1074]. Grayling [2077, 2261, 304, 2341, 98, 111, 323, 740, 1391, 141, 1571, 2732]. grazing [2569]. great [3325, 2511, 2468, 2224, 947, 2247, 2052, 3050, 1699, 599, 878, 65, 1011, 762]. greater [2241, 2391, 2451, 2895]. Greece [2784]. green [2796]. greenback [297, 298, 1392]. Greenland [34, 641, 1886, 2386]. greetings [2595]. gregory [1174]. grenadier [1410, 1079, 2943]. grenadiers [1065]. grey [1439, 500, 3383, 744, 2566, 1135]. grid [588, 3212]. grid-net [588]. groper [2669]. ground [346]. grounds [2007, 633, 240, 1099]. groundwater [687]. Group [1118, 3298, 346, 2693, 1241, 2484, 218, 1818, 2249, 1214, 85, 1036, 1740, 2216, 2598]. group-held [1740]. grouper [2643, 2876, 3039, 67, 3152, 1922, 2348, 2161, 359, 3108, 2578, 2225, 337, 2290, 762]. groupers [944]. groups [735, 783, 1360, 1596, 690, 3173]. growing [2684, 2716, 998, 3100, 68, 2587]. Growling [209]. grown [82]. grows [1270]. Growth [1885, 1991, 1415, 3077, 2214, 1995, 2667, 659, 1013, 1189, 692, 795, 372, 1029, 2096, 2384, 2062, 2998, 148, 1527, 2291, 1150, 2659, 68, 2690, 655, 1327, 1566, 3067, 1710, 1570, 577, 1103, 36, 2216, 441, 469, 72, 1657, 3109, 2819, 1695, 892, 1767, 1490, 2453, 350, 453, 3107, 1156, 1345, 413, 1859, 1604, 540, 2803, 346, 2907, 1672, 19, 1339, 188, 2055, 1577, 814, 392, 2941, 3076, 1891, 1453, 1977, 1023, 2215, 1240, 1418, 719, 735, 2283, 3240, 2611, 3093, 1086, 297, 2383, 3132, 2185, 1907, 1481, 979, 2780, 758, 3368, 1041, 3231, 3274, 2023, 101]. growth [532, 3460, 1249, 1660, 2610, 3371, 1998, 3158, 2157, 2392, 2616, 291, 387, 1766, 2213, 2836, 3097, 470, 1547, 2079, 2716, 2944, 3230, 2900, 2901, 1890, 3469, 2060, 864, 259, 1950, 194, 2484, 1637, 2931, 2633, 2849, 951, 98, 446, 1746, 202, 146, 658, 1968, 503, 1355, 1152, 2031, 1209, 723, 1151, 2037, 1211, 2845, 3196, 235, 473, 646, 2058, 2663, 340, 2032, 2810, 1113, 164, 25, 527, 2565, 3133, 3299, 2728, 3248, 396, 3467, 2398, 329, 518, 2634, 1017, 933, 2002, 3272, 2466, 2778, 2881, 2257, 1912, 393, 3155, 2587, 3392, 1756, 1109, 716, 1526, 1309, 2422, 147]. growth [1082, 195, 2119, 2035, 1789, 2428, 2427, 2869, 489, 3391, 2666, 1047, 53, 579, 1079, 1829, 2763, 1803, 2895, 363, 439, 970, 654, 48, 2260, 3258, 3300, 1361, 3089, 2335, 2187, 1089, 2589, 2505, 2605, 517, 3447, 2401, 1679, 2087, 3008, 2992, 388, 119, 288, 2429, 20, 878, 130, 1066, 2578, 1994, 232, 29, 976, 1764, 2182, 320, 3361, 1228, 3205, 1788, 3159, 1348, 3129, 3262, 150, 2285, 2354, 1623, 3091, 1325, 2937, 2574, 1274, 799, 1352, 69, 13, 2357, 2844, 1809, 2598, 1910, 1392, 1535, 2843, 1576, 3112, 2533, 1575, 3259, 1064, 2196, 237, 212]. growth [2820, 1307, 1640, 448, 1322, 3162, 2661, 2107, 318, 2629, 1270, 939, 365, 1768, 1099, 3251, 1881, 3352, 486, 3346, 377, 940]. Growth-rate [892]. growth-related [3299, 1109]. grunion [3267]. grunniens [1536]. grunter [360]. Guadalquivir [907]. Guadiana [2735, 1667]. guanine [1294]. guarding [2066]. Guatemala [3116]. guatemalensis [1353]. guatucupa [1911]. gudgeon [2092, 1605, 2757, 2760, 245, 2630, 2662, 2746, 2485, 3099]. Gudrun [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. guggenheim [1034]. Guiana [7]. Guidelines [1653]. guilds [3082, 3281, 921]. Guillaume [1405]. Guinea [1722, 3151, 1802]. Guinean [2646]. Güldenstädt [1707]. gulf [3379, 3392, 1487, 2526, 3073, 1168, 3386, 1147, 2610, 3233, 3152, 1504, 700, 2016, 1486, 2022, 3216, 1719, 2495, 1803, 2953, 36, 2919, 2871, 763, 799, 1352, 2268, 2389, 3470, 1846, 2290, 697, 3390, 3194]. gulfs [1922]. Gull [1939]. gulliveri [1453]. gulls [1336]. gulosus [1348]. gulper [2904]. Günther [2020, 2882, 1326, 2201, 2854]. guntheri [2780]. guppies [22, 2026, 1133, 3351, 1297, 665, 1517, 3456, 120, 61, 2051, 116, 1874, 2364, 3363, 739, 514]. guppy [1241, 3434, 2699, 3256, 1256, 3354, 683, 1267]. gurnard [209, 1439]. Gustatory [2028]. Gut [552, 1761, 210, 1353, 3157, 2783, 3328, 2284, 1136, 843, 2114, 1873, 726, 3362, 1480]. guttatus [1980, 1713]. guttulatus [2300, 2023, 2865]. Gwaii [2477]. Gymnocephalus [1545, 2933, 3061]. Gymnocypris [3438]. Gymnotidae [2528]. Gymnotiformes [3430, 2528]. gymnotrachelus [2731]. Gymnotus [2528]. Gymnura [3446, 3324]. gynogenesis [2740, 2983, 657]. gynogenetic [1995, 3293, 747, 3349]. Gyrodactylus [3056, 3197].

H [1654, 3410, 1775, 191]. h-changes [191]. habit [2602]. Habitat [1945, 1344, 2966, 1332, 2929, 2233, 2093, 1861, 2369, 1019, 1783, 111, 2961, 3134, 3480, 1313, 2375, 2962, 3194, 1672, 1777, 753, 1195, 3077, 2339, 2043, 1712, 890, 539, 2883, 1488, 2053, 816, 1243, 1244, 2447, 2902, 2781, 1750, 1982, 682, 125, 2367, 3075, 3190, 2199, 2420, 1683, 522, 2484, 871, 3102, 1191, 623, 2635, 942, 2660, 2222, 1209, 587, 391, 2772, 2412, 1311, 2730, 1550, 3273, 2967, 1669, 1759, 1275, 1953, 2193, 1220, 2495, 1784, 2286, 2606, 531, 732, 1954, 1291, 1955, 323, 740, 1391, 1105, 3138, 2418, 1897, 3101, 1944, 2091, 141, 824, 1548]. habitat [2239, 1958, 2669, 395, 1182, 1854, 948, 337, 2267, 2804, 609, 2560, 1233, 1235, 1268, 2615, 406, 2238, 2080, 569]. habitat- [1235]. Habitat-choice [2375]. habitat-specific [2730]. habitation [2613]. habitats [1704, 2084, 936, 3422, 1382, 1718, 1683, 2484, 1951, 1514, 397, 2690, 863, 1735, 485, 1957, 257]. habits [907, 2426, 148, 2896, 3208, 878, 2054, 2244, 754, 2088, 1059]. habituation [1967]. haddock [2773, 580, 196, 1951, 1756, 1708, 1448, 1334, 1450, 2264, 2090, 1389]. hadh2 [2859]. haematocrit [1699, 1509]. Haematological [442, 2581, 1524, 742]. Haematology [1643, 1762, 989, 772]. haematopoiesis [1832]. haematopoietic [721, 837]. haematopterus [2595]. haematoxylin [1345]. haemoglobin [3085, 1757]. haemoglobins [273]. haemolysis [771]. hagfish [2011]. Haida [2477]. hainanensis [2521]. hair [2385]. hake [115, 936, 3104, 3307, 541, 1464, 2258, 2259, 1686, 1543]. hakes [2537]. Halastatic [1138]. half [2161]. half-moon [2161]. halfbeak [2407]. halibut [864, 1280, 2899, 749, 1437, 25, 1366, 843, 1886, 879, 69, 807, 1024, 1424]. Halichoeres [2508]. Haliochoeres [1045]. halleri [2931, 2112]. Halobatrachus [2306, 2991]. Hals [3202]. hamatus [2467]. Hamilton [2892, 2135, 1519]. hamlet [2564]. hammerhead [3216]. hand [3085]. hand-held [3085]. Handbook [1652]. handle [3476]. handling [2773, 1320, 1804, 1276, 3325, 168, 481, 178, 2036]. hanging [230]. haplochromine [176]. Haplochromis [2020]. haploid [2722]. haplotype [751, 1470]. haplotypes [484, 1343]. haplotypic [2519]. happening [233]. harengus [3110, 2295, 2406, 1794, 2421, 2918]. harmless [2754]. harmlessness [2583]. Hart [1652]. harvesting [3070]. Hasler [618]. Hatch [3132, 1768, 3093, 2610, 3469, 1966]. Hatch-period-dependent [3132]. hatched [571, 1331]. hatcheries [1839]. hatchery [1172, 1852, 484, 1577, 583, 87, 478, 844, 1192, 1108, 2812, 3368, 2561, 784, 1634, 249, 2590, 1610, 864, 2327, 1562, 1563, 1208, 1587, 164, 1997, 1589, 3025, 1565, 11, 2923, 1569, 634, 2963, 1393, 1571, 1261, 2792, 870, 1535, 3455, 1747, 1806, 2958]. hatchery-reared [1172, 1577, 87, 844, 1192, 1108, 2812, 2561, 784, 1634, 864, 2327, 1997, 3025, 1565, 11, 2923, 1569, 634, 870, 1535, 3455, 1747]. Hatching [1625, 661, 196, 2408, 1482, 831, 3138, 3091, 2843]. Hatschekia [1979]. having [84]. Hawaiian [621, 521]. Hazard [2601]. Hazen [651]. head [668, 1681, 1714, 2318]. Healing [2935, 362, 1041]. health [3150, 2591, 3156]. Hearing [2360, 3465, 2221]. Heart [3127, 1884, 152, 1869, 3166, 543, 2698, 774, 447, 2318]. Heat [1915, 2773, 33, 105]. heater [2498]. heavily [382]. Heavy [272, 2396, 338, 3343, 2920, 1744, 2715]. hebvaicum [680]. Hedgehog [2010]. height [1797]. heightened [2366]. held [1821, 3085, 1740]. Helicolenus [1026, 3172, 2596]. Helminth [228, 163, 992]. Hemigaleus [2839]. Hemigymnus [168, 856, 1413]. Hemingway [1402]. hemiramphid [1758]. hemiramphids [2357]. Hemiscyllidae [2631]. Henicorhynchus [3366]. Henle [2800, 2631, 2408, 2837]. henlei [2953]. hepatic [2739, 1633, 2572, 474, 2559]. hepatocyte [3020]. Hepatocytes [2274, 1913, 301]. herbicide [1068]. herbivores [1488]. herbivorous [1758, 2275, 3357, 3396, 724, 81, 897, 65]. Herichthys [2085]. hermaphrodite [2661]. hermaphroditic [1664, 2656]. Hermaphroditism [640, 2649]. Herring [849, 3076, 1127, 3132, 1147, 260, 204, 3110, 387, 1766, 661, 2295, 2406, 1527, 673, 3019, 617, 1549, 2421, 48, 2087, 2068, 119, 26, 10, 603, 2918, 1986, 1161, 769, 709]. hertwigi [1677]. Herzegovina [2522]. herzi [3303]. heteroclitus [3331, 3148, 2647, 1817, 3349]. Heterodontus [2645, 2808, 2842]. Heterogeneity [2827, 217, 2367, 249, 3023, 2286, 1261, 948]. heterogeneous [723]. heterologous [1408]. Heteromorphic [2926]. heteroplasmy [739]. Heteropneustes [246, 1295]. Heterosigma [117]. Heterosis [683]. heterospecific [2165, 1718, 3398, 1154, 2583]. heterozygosity [950]. Hew [1404]. hexaploid [3294]. hexokinase [132]. hide [1734]. hierarchies [1260, 767]. High [855, 3075, 2223, 229, 3037, 2707, 1578, 3148, 2877, 2642, 659, 2712, 1517, 2684, 3381, 2242, 830, 444, 2793, 2147, 677, 983, 2299, 2399, 1536]. High-resolution [3075]. high-temperature [677]. highly [3295, 602, 2518]. Hildebrand [1405]. hill [1519]. hind [2961]. hinder [3453]. hippocampus [2300, 2185, 2300, 2023, 688, 2865, 2606, 695, 2377, 1078, 3374, 2474]. Hippoglossina [240]. hippoglossoides [1886]. Hippoglossus [2899, 2959, 1819]. hirame [3277]. histidine [891]. histochemical [1162, 2145, 1688, 2480, 1328, 1622, 3329]. histochemistry [2116, 351, 880, 1644]. histocompatibility [1858, 2918, 2519]. Histological [1162, 1058, 196, 1349, 3149, 309, 1864, 296, 2312, 1306, 347, 3356]. Histology [1688, 2116, 351, 315, 2471, 2268, 3322]. Histopathology [3053, 2997, 2665]. Historical [2616, 375, 3051, 3185, 2341, 2147, 599, 2794, 229]. histories [2956, 2110, 2067, 1374, 507]. History [1901, 2794, 1558, 115, 2081, 2866, 3036, 3151, 2723, 2192, 2670, 3000, 1489, 2023, 2021, 3429, 2179, 3075, 2638, 2550, 1429, 2745, 997, 1020, 656, 2753, 2218, 933, 2585, 2841, 2458, 887, 1075, 3256, 1756, 899, 154, 950, 1978, 53, 681, 579, 865, 2754, 322, 581, 1425, 634, 2919, 2087, 731, 1258, 3117, 3395, 3369, 1334, 29, 1571, 1970, 2544, 3324, 1597, 1574, 1575, 2957, 769, 212, 3449, 2444, 2537, 3342, 3412, 2431]. histrio [937]. hogfish [2656]. Hokkaido [2094, 2736, 3122, 3472, 2133]. Holacanthus [3263]. Holarctic [1070]. holbrooki [3077, 3021, 3013]. Holding [792, 313]. Holmberg [2698]. Holocephali [3359, 1835]. Home [604, 1639, 498, 2371, 1212, 3025, 2973, 3180]. home-range [2973]. Homing [525, 52, 973, 2042, 2764, 1214, 265]. homogenate [3282]. homoion [2662]. homologous [3293]. honest [1227, 1221]. honeycomb [2643]. honmasu [11]. honmoroko [515]. Honshu [630]. hook [1504, 3103, 1463]. hook-and-line [1504]. hooking [286]. Hoplias [2292, 2274, 2505, 573, 772, 2665, 3459, 3436]. Hoplosternum [388]. Hoplostethus [725]. Horabagrus [2201]. horizontal [411]. Hormonal [97, 761, 1387, 1472]. hormone [1604, 2214, 2145, 360, 3371, 1547, 2384, 503, 1063, 1113, 527, 937, 1017, 309, 649, 1566, 1082, 489, 2441, 2677, 3067, 439, 221, 319, 441, 1441, 1066, 879, 13, 2313]. hormone-induced [360, 503]. hormone-producing [2145]. hormones [187, 31, 2268]. horn [121]. hornyhead [3380]. horse [198, 385, 386, 2344]. Host [2395, 3052, 1239, 2675, 680, 2168, 2008, 479, 3054, 1259, 2943, 471, 1543, 1057, 3056]. host-parasite [1259]. hosts [3057, 3197]. hot [906, 1266]. Hotel [1824]. hound [2871]. houndshark [2105]. hours [3316]. houting [2070, 1833, 923]. Houttuyn [2806]. Hox [2916]. HPLC [2840]. Hsd17b10 [3217]. hsp30 [33]. Hsp70 [1837, 2773, 33]. hubbsi [3104]. Hubei [3294]. huchen [3344]. Hucho [1860, 3344]. hudsonius [2591]. HUFA [459]. human [615, 2341]. humans [3476]. humic [2437]. Hunger [2229, 1024, 2972]. Hureau [502]. Huso [3325, 2511]. huttoni [2204]. Hybopsis [3044]. hybrid [1872, 601, 817, 2237, 2782, 2115, 2733, 2691, 1400, 2216, 2795, 1707, 3214, 2313]. hybridization [2328, 2721, 1558, 3451, 610, 2052, 325, 3433, 3122, 1067, 3118, 1590, 2527, 2329, 598, 1839, 1601, 2315, 2133, 2313]. hybridogenetic [1667]. hybrids [234, 403, 1082, 3349, 460, 2206, 2861, 1103, 2476, 697, 145, 1625]. hydration [3427]. hydro [2689]. hydro-mechanical [2689]. Hydroacoustic [371, 857, 854, 3365]. hydrodynamics [2844]. hydroelectric [1958]. hydrography [728]. Hydrolagus [1835]. hydrologic [1059]. hydrology [2081, 2694]. hydrophila [2511]. hydrostatic [2073]. hydrosystem [2095]. hydroxylase [3354]. hyperbaric [2425]. Hyperbenthic [1132, 1467]. hypercapnia [379]. hypercarbia [3282]. Hyperoglyphe [2836]. hyperphagia [54]. Hyperplastic [3091]. Hypertension [1060]. hyperthermia [2157, 2392]. hypertrophic [3091]. hypertrophy [2245, 2219]. hyperventilation [1562]. hypo [30, 520]. hypo-osmoregulatory [30, 520]. Hypomesus [2084]. hypoosmoregulatory [2405]. Hypophthalmichthys [2457, 3221]. hypophthalmus [1036]. hypophysial [1379]. Hypoplectrus [2564]. Hyporhamphus [1758, 2407]. hypothalamic [398]. hypotheses [3097]. hypothesis [453, 1999]. hypoxanthine [404]. hypoxanthine-3-N-oxide [404]. Hypoxia [3395, 2747, 1097, 1475, 555, 2197, 260, 658, 2126, 1919, 3383, 1082, 2409, 590, 517, 999, 219, 1018, 3178, 1346, 2305, 231]. hypoxic [199, 2579, 2623, 3216, 3285]. Hypseleotris [2092]. hypselosomus [3257].

Iago [2105]. Ibbotson [792]. Iberian [3185, 607, 851, 582, 2735, 2745, 2850, 2896, 3026, 2708, 2148, 423, 2560]. ibitiensis [3435]. ice [230, 539, 1019, 328, 2962]. icefish [278, 1051, 1916, 476]. Iceland [34]. Icelandic [2062, 2998, 2342, 3284]. ICES [1824]. Ichthyodinium [175]. ichthyofauna [1137, 1153]. ichthyogeographic [2708]. Ichthyologie [1009]. Ichthyology [2255]. Ichthyophthirius [174]. ichthyoplankton [908, 1287]. ICP [1127]. ICP-MS [1127]. Ictaluridae [3440]. Ictalurus [3262]. Idaho [12]. ide [1583, 1607, 3180, 1119]. idella [1838]. identical [1117]. Identification [2546, 2145, 3039, 2861, 275, 2734, 2651, 2467, 3347, 2636, 1035, 957, 847, 1872, 1726, 582, 1552, 1793, 382, 2724, 243, 1471, 1464, 2898, 1835, 2206, 1284, 1905, 1367, 2971, 3064, 380, 1873, 2798, 2331, 422, 2537, 1787, 971, 3468]. identified [2725]. identify [873, 2795]. Identifying [2202, 2224, 2621, 2300]. identity [2300, 2135]. idus [1583, 1607, 3180]. if [3197]. IFREMER [1775]. IFREMER-IRD [1775]. IGF [1913]. IGF-I [1913]. ignorant [1219]. Iguacu [3043, 3439]. II [1858, 217, 355, 2417, 2918, 386, 2533, 2519, 1532]. IIB [3434]. III [1694]. Illinois [2759, 1000]. illumination [2995]. image [1002]. imagery [1787]. imaging [3110, 3433, 3033, 1713]. immature [54, 528, 466, 532, 489, 2476, 939, 220]. immediate [2198]. immigrant [244]. immigration [808]. Immnunohistochemical [785]. Immune [2511, 3033, 3196, 1688, 1289, 1659, 1838]. immunocompetence [3387]. Immunocytochemical [2542, 2145]. immunoglobulin [2715]. immunoglobulins [574]. Immunohistochemical [648, 1438, 2134, 1622]. immunological [2138]. immunoneutralization [1089]. immunoreactive [3328, 3354, 2887]. immunoreactivity [2330]. immunosorbent [348]. immunosuppressive [2138]. Impact [2445, 728, 1765, 1152, 1822, 1274, 2788, 78, 2215, 2558, 144, 2341, 1483, 1879, 1709, 1012]. impacted [2397]. Impacts [1312, 486, 2758, 1958]. impair [784]. impaired [3457]. impairment [1647, 1381]. imperilled [3440, 1559, 2414]. impingement [914]. implant [639, 1100, 2674, 2376, 2012]. implantation [1697, 830, 1822]. implanted [2374, 2283, 1041, 1071, 2810, 3111, 2696, 976]. implication [814, 3353]. Implications [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 3414, 1773, 3175, 1564, 176, 3185, 3366, 545, 882, 1946, 153, 3000, 758, 2336, 610, 3231, 958, 2550, 1429, 3171, 245, 3152, 698, 765, 1211, 1966, 1040, 704, 2730, 3288, 3221, 617, 2292, 2612, 2720, 3117, 2429, 1542, 1424, 1910, 2236, 3232, 1973, 2668, 2444, 1233, 2732]. imply [1816]. importance [2723, 1244, 3241, 3053, 3344, 1185, 1615, 2284, 967, 994, 72, 2114, 603, 2957, 1640, 835]. Important [722, 2557, 2357, 2750]. Impoundment [144, 3270]. improve [827, 1345, 1192, 3196]. improved [971]. Improving [1208]. Imsa [3133, 974]. in-depth [2755]. inadvertent [1375]. inanga [2204]. inbreeding [3460]. Incertae [3292]. Incidence [268, 580, 2157]. incision [1041]. including [457, 2193, 3219]. incompatibility [1625]. inconnu [520]. inconstans [1499]. incorporating [1029, 1803]. Incorporation [2645, 3268, 780]. increase [2739, 339, 923, 2681, 638, 1153]. Increased [3315, 126, 403, 503, 2682, 647, 3201, 1510, 1761]. increases [719, 752, 1566, 487, 2045, 2190]. Increasing [1455, 2681]. increment [977, 1383, 261, 1766, 3381, 3004, 3392, 2298, 1899, 2957]. increments [1345, 2900, 1679]. incubating [3267]. incubation [3193, 3093, 2408, 2056, 2223, 1310, 2651, 3135]. independent [587, 1118, 2136, 3262, 2148, 2566]. indeterminate [3257]. Index [1032, 1629, 1541, 1943, 1962, 1751, 1903, 2352, 2149, 2102, 2311, 2334, 2762, 2514, 2548, 2741, 3087, 3265, 3483, 2703, 1290, 1766, 700, 1478, 2488]. India [2201, 3345, 2135, 2377]. Indian [689, 3141, 2724, 2327, 2398, 734, 3375, 564, 2152, 3245, 1839, 2880, 3353]. indicate [3174, 2854]. indicated [2291, 513, 1110, 2210]. indicates [2717, 751, 1819]. Indications [1070]. indicative [1725]. indicator [961, 2770, 2061, 1255, 2169, 512, 1625, 1837]. indicators [2502, 2188, 3369, 2824, 185, 922]. indices [2387, 187, 1717, 3284, 781, 1101]. indicus [3207]. Indies [604, 605]. indigenous [862, 382, 2749, 2004, 1917, 2752]. indirect [205, 1981]. Individual [1384, 884, 1206, 1890, 634, 1693, 1236, 2991, 847, 2932, 1889, 1793, 1194, 725, 3469, 1100, 723, 865, 1117, 48, 1361, 2335, 2971, 1229, 1004, 2224]. Individual-based [1206, 1194, 723]. Individual-level [1890, 634]. individualized [97]. individuals [3247, 1143, 2795, 1711]. Indo [1428, 3207, 2041, 1115, 3030, 564, 1258, 762]. Indo-Malay [564]. Indonesia [2851, 2144, 2701, 3222, 2672, 2503]. indoor [1320]. induce [1935, 2363, 446, 474]. Induced [2271, 1387, 2980, 360, 544, 286, 1562, 2025, 503, 861, 1915, 2909, 3183, 1252, 1309, 2470, 1956, 319, 1362, 60, 3083, 1381, 1355]. inducement [1472]. inducer [2739]. induces [2245, 3201]. inducible [311]. inducing [678, 319]. Induction [258, 657, 118, 1153, 1408]. indus [903]. infected [1414, 3421, 219, 2219, 2573, 846, 206, 1416, 3055]. infecting [1999, 3056]. Infection [680, 175, 2622, 2558, 961, 575, 354, 2459, 2068, 9, 3063, 2506, 3128, 3277, 1364, 2029, 2982]. infections [3053, 3054]. inference [2343]. inferences [3437]. inferred [2988, 788, 2130, 2110, 3291, 2736, 1067, 3372, 2981, 1755, 854, 2455, 2919, 3227, 3037, 1470]. infestation [2225]. inflation [1834]. inflow [2496]. Influence [1888, 188, 2081, 2043, 2086, 1607, 1481, 3274, 170, 1163, 1342, 2550, 3230, 2294, 194, 723, 613, 3202, 2552, 2466, 489, 1829, 1420, 1620, 1310, 2156, 2559, 130, 1348, 1623, 337, 2231, 214, 2972, 1983, 890, 24, 2504, 555, 3190, 2446, 2373, 522, 652, 1335, 2694, 1350, 2682, 2164, 340, 2772, 2917, 1966, 1844, 2903, 2664, 681, 897, 2417, 2473, 1117, 142, 28, 2589, 1304, 647, 3098, 2432, 2614, 417, 1736, 2962, 1973, 3358, 3426]. influenced [1023, 382, 342, 2603, 1870, 1791]. Influences [1066, 495, 1171, 1732, 61, 1118, 2348, 2046, 1874, 1553, 3253, 143, 313, 451]. influencing [299, 3180, 1354, 1319, 822, 3312]. influx [687]. Information [1651, 2948, 1183, 3084, 956, 3181, 1665, 3272, 497, 1597, 1142]. informative [712]. infracommunity [1543]. infrared [307]. infrared-gates [307]. infrasound [2209]. infulatus [542]. ingesting [3381]. ingestion [1527]. ingredient [3149]. ingress [2236, 3232]. ingrowth [1847]. inhabitant [2851]. inhabiting [2956, 3452, 2692]. Inheritance [534, 2983, 1754, 2790, 739]. Inherited [668, 1292]. inhibition [962, 844, 279]. inhibitor [3282, 677]. Inhibitory [720, 2365, 738]. inhibits [448]. initial [217, 2758, 1660, 174, 1832, 3202, 3159, 69, 1834, 3352]. initiating [1854]. injection [344, 1838]. injury [268, 1362]. Inland [537, 2147, 548]. inlet [1357, 2209]. Innate [2442, 686, 3317]. inner [1696, 2385, 2113]. inoculation [3130]. inorganic [1528, 2393]. Insectivore [1353]. insertion [1879]. Inshore [1006, 2951, 1808]. insight [1473]. insights [2711, 3218, 160, 2150, 1370, 3089, 2609, 1929, 2706, 2044]. insignis [764]. inspection [718, 3424, 1234]. instability [2662]. instream [587, 1145]. Insulin [2728, 3274, 1547, 2533]. Insulin-like [2728, 3274, 1547, 2533]. intake [1242, 339, 1152, 473, 2767, 2051, 1357, 2200, 2968, 809, 690, 1761, 2598, 365]. Integrated [2279, 2309, 2381, 2436, 2465, 2492, 2411, 1653, 2513, 1989, 639, 872, 307, 1949, 1540, 3212, 3108]. Integrating [1824]. integration [121]. integumental [745]. Intensely [1395]. intensities [2584]. intensity [1238, 2611, 544, 2974, 580, 2295, 3102, 2834, 2607, 1740, 1256, 3385, 2965, 316, 9, 853, 1440]. intensive [882]. Inter [278, 992, 1361, 2167, 2269, 3232, 1269, 1889, 229, 1531]. Inter- [992, 2167, 2269, 1531]. Inter-annual [278]. Inter-individual [1361, 1889]. inter-oceanic [229]. Interaction [2902, 896, 2958, 1207, 1029, 1861, 1588, 1254, 1184, 1595]. Interactions [1901, 3050, 1563, 3387, 975, 569, 1241, 3302, 1245, 2756, 1249, 2332, 3021, 2868, 1593, 358, 2421, 1816, 1103, 2375, 1598, 3052, 1233, 1773]. interactive [686]. Interannual [2610, 3307, 2355, 2024, 624, 729, 2016, 70, 2217, 863, 2629, 3228]. Intercalibration [3365]. interception [2221]. Interdemic [1509, 1475]. interdrainage [375]. intergenic [2987]. Interhabitat [1145]. interhepatocytic [744]. intermittent [1958]. internal [685, 2045]. International [456, 477, 1706, 2077, 2788, 409, 436, 877, 1324, 1989, 2601, 2100]. interorgan [23]. Interpopulation [2587]. interpretation [2940]. interpreted [1597]. interpreting [1781, 1393]. interrelationships [1745, 916]. interrenal [983]. Intersex [1921, 1373]. intersexuality [296]. Interspawning [3254]. interspecies [800]. Interspecific [2328, 289, 1037, 3022, 2873, 2757, 295, 2059, 3122, 3123, 705, 2416, 2329, 1742, 1692, 750]. interstock [610]. intertidal [1488, 295, 2827, 255, 803, 81, 343, 1351, 2047, 3090, 2875, 2874]. intertidal-nesting [3090]. interval [3254, 1631, 1834]. intervals [2581]. Intestinal [1737, 13, 2203, 648, 1076, 2480, 344, 2586, 2559, 118, 1917]. intestinalis [3060, 531]. intestine [2195, 1612, 1027, 880]. Intra [1297, 2501, 991, 421, 992, 23, 705, 2269, 3232, 1692, 750, 1269]. intra- [23, 705, 1692, 750]. Intra-annual [421]. intra-estuary [3232]. intra-morph [992]. Intra-population [2501, 2269, 1269]. Intra-school [991]. Intra-sexual [1297]. intraannual [2355]. Intrabasin [2453]. intracellular [3282, 1033, 3319, 3080]. intraperitoneal [344, 1838]. Intrapopulational [2862]. intrarenal [1155]. Intraspecific [2065, 842, 480, 2226, 742, 2047, 524, 2328, 1807, 3123, 2167, 9, 1286, 1625, 1531]. Introduced [2506, 753, 1415, 2757, 2339, 855, 462, 2756, 2179, 3260, 2745, 621, 139, 2059, 202, 3021, 1067, 2868, 2143, 460, 2731, 896, 2501, 598, 2747, 1576, 1862, 2560, 1233, 3099, 163]. Introduction [2743, 3473, 732, 1627, 2758, 1242, 1512, 790, 2324, 2769]. introductions [614, 1170]. introgression [600, 1852, 2143, 612, 2794, 2148]. Introgressive [2133, 2527]. intron [2712]. intruder [3067]. intrusions [3398]. invasion [2751, 2743, 2755, 2759, 1233]. invasive [152, 1605, 1920, 2758, 1889, 1525, 2744, 3010, 1187, 2719, 3286, 2903, 1987, 1275, 1879, 2528, 2329, 2485, 3420, 2748, 3400]. invasively [1436]. inverse [2803]. inversion [1409]. invertebrate [682]. invertebrates [12]. investigate [303, 215, 294]. Investigating [1848, 2952, 587, 2841]. Investigation [3253, 1581, 1058, 3149, 105, 3346, 3356]. Investigational [3260]. investigations [2526, 1162, 1957]. investment [2504, 1547, 414, 164, 1188, 1218, 475]. iodine [879]. ion [554, 2330, 830, 425]. ion-poor [425]. Ionic [1446, 3160, 476, 301, 3384]. Ionoregulatory [3330, 284, 2454]. Ipnopidae [3326]. Iquitos [3449]. IRD [1775]. Ireland [1983, 302, 299, 449, 563, 1650, 2013, 2109, 3154]. iridovirus [3046]. Irish [2781, 616, 2634, 2264, 1747]. iron [1555, 2557, 2691, 173]. iron-binding [1555]. irradiance [1490]. irradiated [1914]. irritation [1293]. irwini [3107]. isaaci [3015]. ISBN [1403, 1496, 1497, 1406, 1654, 1402, 1772, 1651, 1404, 1773, 1774, 1493, 1775, 1405, 1652, 1495, 1494, 1653]. Ishikawa [2858]. Ishiyama [1682]. Isistius [733, 1396]. Island [1430, 1975, 3122, 3472, 787, 1649, 3372, 1336, 1909, 378, 2133, 468, 343, 2220, 642, 3416, 3050, 622, 2152, 1093, 2225, 799, 1010]. Islands [729, 2468, 1084, 664, 642, 2610, 2033, 1477, 2806, 3037]. Isle [2634]. Isles [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513, 1, 180, 3175, 1936, 1628]. isoform [338, 3343]. isoforms [987]. isogenic [437]. isolated [3451, 3437, 2590, 2200, 2524, 212, 3417]. Isolation [2705, 90, 471, 793, 1149, 2026, 2317, 3416, 751, 2314, 1755, 1260, 903]. isoloci [828]. Isopod [2225, 2395]. isopods [3050, 1699]. isotope [960, 764, 2771, 2756, 3095, 3075, 2303, 2482, 2281, 2493, 2284, 1539, 3445, 2833, 1144, 479, 775, 294, 2064, 3156]. isotopes [651, 2031, 3310]. Isotopic [997, 1758]. Isozyme [711]. issues [1992, 3156]. Isurus [2685]. Italian [751, 433]. Italy [2100, 1613, 607, 2139, 1457, 3341]. items [1176, 824, 1638]. Iteroparity [2765]. iteroparous [589, 324]. ITS1 [2130]. IV [1694, 1136]. ix [1404]. Izak [47].

J [3479, 1496, 1497, 1406, 1651, 1493, 1775, 3478, 3321, 1652, 1495]. jack [1518, 2568]. Jackson [497, 2808, 2842]. jamaicensis [1263]. januaria [2618]. Japan [3132, 630, 3136, 943, 2226, 2007, 2082, 2736, 3122, 2094, 3278, 3472, 1753, 579, 3309, 3455, 3277, 2322, 2133, 492, 3251]. Japanese [55, 743, 1189, 1249, 1660, 630, 2242, 2041, 2110, 519, 1151, 2007, 704, 2168, 3272, 926, 825, 2470, 336, 3385, 2992, 319, 3037, 1472, 2792, 591, 1445, 1865, 2285, 1279, 1294, 720, 628, 2377, 3455, 3277, 3074, 237, 1317, 2064, 347, 3251, 3228, 793]. japonica [1906, 3230, 3226, 3272, 1880, 1400, 2088, 3228]. japonicus [2041, 2291, 2007, 2858, 2806, 3079, 2285, 1723, 1767, 1768, 2568, 3251, 2318]. jar [2223]. jaw [2888]. jaws [2925]. Jean [1405, 2637]. Jean-Loup [1405]. Jenyns [2419]. Jersey [1772, 1405]. jetty [2776]. JFB [2977]. Joban [3277]. Johnston [2694]. Jordan [2114, 2768, 2661]. Jordanella [2867, 1482]. jordani [3309, 1812]. Journal [1991]. Jr [1495]. Júcar [156]. July [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2077, 2381, 2436, 2465, 2492, 2411, 2513]. jumping [1857]. June [1706, 1989, 1965]. junior [2132]. justification [1964]. Juvenile [1650, 2640, 2901, 522, 2249, 350, 413, 540, 1946, 1320, 2055, 1643, 2866, 392, 2941, 2424, 1038, 583, 1323, 932, 1023, 1240, 2070, 2086, 2642, 2283, 3302, 1580, 445, 1474, 2430, 466, 977, 2864, 2809, 27, 2185, 1607, 2188, 1041, 982, 1332, 293, 1430, 894, 3343, 3460, 2367, 1189, 1249, 3260, 868, 1207, 1584, 1646, 399, 1561, 3325, 2616, 387, 1766, 636, 1461, 2096, 3097, 2079, 619, 104, 2060, 444, 1950, 1197, 1459, 2124, 1191, 1861, 2025, 503, 126, 867, 1186, 223, 1152, 587, 1763, 1437, 1426, 235, 473, 646, 2058, 553, 2151, 1530, 25, 685, 2730]. juvenile [3298, 3137, 3131, 3181, 1668, 2687, 3278, 3004, 66, 58, 1154, 1000, 1591, 2466, 2881, 2284, 1219, 393, 1377, 3111, 1104, 1566, 2010, 2452, 2422, 487, 154, 526, 2298, 627, 53, 681, 1492, 2403, 3301, 3384, 485, 2507, 2416, 2417, 140, 113, 1117, 2320, 1190, 1125, 2801, 1717, 3376, 1570, 3300, 1361, 2838, 376, 2505, 517, 2418, 2567, 316, 2216, 119, 288, 1841, 1066, 3283, 29, 1764, 72, 1260, 83, 3128, 1865, 1623, 3091, 2819, 1466, 2556, 2937, 1338, 690, 1693, 2844, 2598, 2875, 826, 811, 2343, 815, 1294, 2476, 1499, 1161, 2957, 2584, 212, 1640]. juvenile [76, 1130, 2518, 1101, 2615, 3404, 493, 1368, 1767, 1556, 2958, 377, 841]. juveniles [412, 384, 953, 1726, 482, 233, 1606, 979, 1608, 901, 1611, 652, 519, 2007, 2083, 440, 701, 3183, 1467, 1864, 1565, 102, 3150, 2152, 2264, 3014, 1445, 2269, 1910, 3390, 1480, 1099, 3468, 1169].

Kaiser [3478]. kajose’ [488]. Kalandsvatnet [1469]. Kallmann [2782]. kammalensis [123]. Kao [3272]. Kao-Ping [3272]. kapalensis [2837]. Kapoor [3321, 3406, 3407, 3320]. Karaman [2000]. Kareius [2476]. Karlskrona [1253]. karst [1137, 1996]. karstic [2747]. karyological [3214]. karyotype [2516]. Karyotypic [3345, 3436]. kasmira [3065]. kasmiva [621]. Katsuwonus [2140, 2927]. Kattegat [1570, 1830]. Kaup [2611]. Kawamura [3478]. Kazakhstan [3219]. keeping [3275]. kelp [1336, 3108]. kelts [80]. Kenya [2758, 1998, 2324, 2527]. kept [309]. keratinization [2639, 1519]. keta [2686, 2990]. ketone [191]. ketotestosterone [1387, 172, 3311]. key [1069, 1835, 3245, 2651]. Keys [337]. kidney [1681, 3160, 2195, 2332, 562, 1148, 476, 2219, 2698, 2318]. Kii [3251]. killer [204]. killifish [2326, 2622, 2855, 3101, 1170]. killifishes [3125]. Kimberley [3119]. kin [3069, 2710, 3425]. kin-biased [2710]. kinase [3080]. kinds [3378]. kinematic [1857]. kinematics [338, 165, 1797, 1806]. kinetics [338, 1136]. King [392, 1380]. Kingdom [2743]. Kinship [3317]. kisutch [2812, 3195, 1499]. kite [2585]. Klunzinger [452]. klunzingeri [2092]. knifefish [2644]. Knipowitschia [1221, 3199]. knowledge [1989]. known [245]. Kobyliansky [2858]. Kodiak [799]. Koefoed [2711, 2985, 1729]. koi [1753]. Kokanee [1637, 1048, 2289]. kokopu [384, 953, 881]. Konkoure [1802]. Konosirus [2082]. Korean [1438, 96]. Kotthaus [2885]. kranjiensis [3468]. Krefft [2630]. krempfi [2613]. krill [1471]. Krøyer [2115]. Kryptolebias [2664]. kuda [2185]. kuhlii [3182, 3447, 3009]. Kuibyshev [1614]. Kullander [1495]. Kurtus [1453]. kurumeus [2168, 2008]. Kvarner [509, 1132]. Kyphosidae [3037]. Kyphosus [604, 724]. Kyushu [2054].

L [1654, 1402, 1772, 1651, 1404, 1657, 1536]. L. [214, 2773, 3290, 1672, 2739, 2131, 2424, 2227, 1840, 2771, 2339, 2043, 2717, 2693, 1240, 2086, 2388, 2121, 2558, 3069, 2198, 2897, 2599, 1923, 2387, 2319, 2481, 2809, 1606, 1583, 1607, 2358, 2902, 2781, 199, 2924, 2367, 2496, 2171, 2446, 2572, 604, 2893, 2888, 1650, 2400, 2684, 2330, 2878, 2302, 2617, 2821, 1833, 2460, 2555, 3344, 2195, 2438, 2647, 2115, 2920, 2355, 2013, 2484, 2341, 2062, 2899, 2406, 3328, 1797, 1737, 2509, 2845, 2733, 2917, 2126, 2342, 2471, 1969, 2815, 2170, 2632, 2767, 2865, 2822, 2834, 2903, 2143, 2634, 1591, 2691, 2841, 3372, 2778, 2245]. L. [2718, 3425, 2428, 2427, 2140, 1619, 1794, 2258, 2259, 2754, 2895, 2480, 2421, 2593, 2861, 1644, 2838, 2187, 2605, 2401, 2567, 2887, 2483, 2264, 2090, 2602, 2871, 472, 2824, 2539, 2748, 1265, 1597, 2918, 1372, 2343, 2843, 2533, 2288, 2048, 2322, 2681, 2668, 1707, 2123, 2519, 2874]. labelling [1049]. Labeo [3198]. lability [63]. Laboratory [957, 932, 1608, 402, 1238, 2424, 1281, 360, 55, 372, 519, 1966, 2428, 2399, 767, 1024, 1161, 3400, 1480]. laboratory-reared [360, 519, 1480]. Labrador [143, 35]. labrax [3153, 1923, 2387, 1608, 3457, 2195, 3183, 3391, 1829, 1710, 2824, 3362, 2533]. labrid [753, 1409, 856, 2508]. Labridae [1639, 1699, 2656]. labriform [3337]. Labrisomidae [343]. Labroides [856]. Lachnolaimus [2656]. Lack [2773, 2826, 916, 88, 3048, 286]. lacking [459]. lacrymatus [2189]. lactate [1276, 168, 1509, 2683]. lactoferrin [1681]. Lacustrine [1731, 1764, 1382, 148, 2817, 1939]. lacustris [3061]. laevifrons [2047]. laevis [642, 449]. lag [2415]. lagenar [1035]. Laggan [2317]. Lagocephalus [1727]. Lagodon [2614]. Lagoon [387, 546, 1953, 1671, 808, 1360, 2987, 254, 1287, 918, 167, 920]. lagoons [916, 2968, 2035, 1123, 669]. laid [431]. Lake [687, 2558, 2199, 46, 3045, 403, 3038, 3349, 1939, 207, 428, 3094, 2092, 2956, 1456, 312, 1534, 900, 2368, 3242, 1165, 189, 737, 2550, 2061, 2450, 2694, 2999, 1527, 5, 3357, 1102, 1311, 1969, 1150, 2936, 2868, 1000, 3372, 857, 2286, 732, 1080, 361, 634, 232, 2219, 8, 2399, 2209, 3281, 2844, 2843, 2933, 185, 257, 1322, 131, 1533, 2872, 1777, 2758, 3060, 1750, 1203, 2929, 1998, 1383, 15, 2324, 997, 2191, 2438, 696, 651, 334, 1098, 1746, 340, 1491, 2072, 236, 548, 3041, 3139, 2784, 1753, 3125, 1285, 2914, 2981, 2527, 550, 2473, 1677, 1144, 3089]. Lake [3098, 727, 2201, 2129, 1458, 2142, 2132, 2475, 2811, 488, 1638, 1707, 3162, 3342, 3222, 3012, 49, 2431, 2238, 163]. lake- [900]. lake-dwelling [232]. lake-type [1456]. lakes [162, 2905, 860, 659, 1243, 2021, 3241, 303, 2061, 529, 1527, 656, 1435, 1516, 1019, 776, 2625, 675, 1316, 2560]. lalandii [2800, 837]. lamellae [2302]. Lamna [2948, 1811]. Lamniformes [2891]. Lampetra [3437, 2013]. lamprey [3437, 42, 301, 2015, 2793, 6, 1318, 1670, 2696, 1503, 2887, 320, 1343, 2748]. lampreys [3002, 63, 1436, 2050, 1635, 1073, 2146]. Lampridiformes [2139]. Lamprologus [238]. lamta [1519]. land [890]. landlocked [42, 708, 2817, 1433, 105, 321]. Landscape [3431, 1945, 1499]. landscapes [1194]. langsdorfii [2525]. Languedoc [1360]. lantern [2995, 3000]. lanternfish [2911]. lapillar [2125]. Lapland [138]. Large [1031, 3011, 2461, 2362, 2445, 3003, 684, 1731, 901, 963, 1083, 686, 3249, 5, 972, 2764, 1388, 2005, 485, 2240, 2067, 323, 2612, 376, 2167, 1145, 229, 2239, 1873, 2844, 826, 822, 3449, 1322, 1329, 2982, 1153]. large-bodied [3003, 3249]. Large-scale [1031]. largely [3127]. largemouth [555, 610, 1041, 1378, 2354, 2212, 1973]. larger [838]. largescale [801]. largescaled [3030]. Lars [2103]. Larvae [1649, 1487, 1885, 1415, 2502, 3314, 1417, 2387, 233, 979, 360, 2725, 1430, 462, 55, 743, 1660, 1443, 2617, 2592, 3307, 2242, 2014, 1611, 661, 864, 1280, 2295, 519, 2450, 3317, 2406, 3328, 1076, 1460, 2823, 693, 710, 440, 3106, 518, 2634, 1734, 794, 1107, 655, 1864, 1464, 283, 1285, 1794, 1380, 3016, 3385, 1828, 2068, 469, 1686, 843, 2264, 1166, 3159, 591, 1445, 1736, 2651, 879, 768, 3362, 769, 3390, 3156, 1480, 1490, 3352]. Larval [3104, 1684, 1733, 2151, 2082, 2912, 147, 2898, 2594, 605, 1345, 1848, 2065, 2580, 719, 1127, 2443, 2611, 1054, 544, 1430, 2419, 580, 3157, 387, 1766, 2770, 2015, 715, 538, 898, 241, 342, 341, 2291, 426, 1063, 3275, 819, 1515, 155, 63, 2607, 1436, 1756, 2427, 1619, 1899, 745, 1829, 2050, 2416, 2417, 3406, 48, 955, 3300, 3089, 3008, 2992, 383, 3117, 521, 3227, 1401, 72, 1820, 1812, 1657, 2774, 2277, 2285, 1229, 1274, 1695, 1294, 2476, 2661, 2155, 1368, 1767, 3251, 3228, 486, 507, 835]. laser [1127]. last [3432, 2752, 2837, 2839]. lasting [1069]. Late [1169, 1391, 2559]. latency [2566]. lateolabracis [680]. Lateolabrax [2007, 2285, 2318]. later [2373, 2477]. Lateral [860, 2448, 928, 3209, 3418, 497, 3091, 3408]. lateral-plate [3418]. lateralis [2282, 869]. laterally [84, 1445]. Lates [3093, 2182, 2797]. Latham [2819]. latipinna [2006, 1475]. latitude [25, 2231]. Latitudinal [1422, 2390, 3028, 3331, 3140, 3258, 1796, 3419, 577, 1057]. Latris [1834]. Latrobe [2770]. latus [3040]. Laurentian [1534]. lavaretus [3206, 1311, 1969, 3089]. law [1603]. Lawrence [466, 1147, 3233, 700, 697]. layer [2362, 2445]. lead [392, 2917]. leafscale [2904]. leafy [816]. learn [1196, 884, 462, 1212, 3363]. learned [2744, 3059]. Learning [981, 478, 1198, 1108, 293, 1208, 2632, 1114, 102, 1599, 2874]. least [423]. leather [2910]. leave [1367, 3051]. Lebanese [935]. lectin [880, 1328, 1644]. left [718, 2959]. Leiostomus [1832]. lemon [872, 2079, 3283, 826]. Length [2713, 661, 2693, 1815, 64, 1966, 1754, 1040, 2249, 2008, 85, 2024, 1327, 536, 2050, 2455, 1703, 9, 213, 521, 2490, 512, 968]. lengths [2899, 572, 627]. lentic [327, 1346]. Lentipes [521]. leopardus [3315, 2052, 2727, 1979]. Lepeophtheirus [2115, 1479]. Lepidocephalichthys [226]. Lepidorhombus [2244]. Lepisosteidae [3036]. Lepomis [2339, 2179, 2745, 3317, 3337, 3205, 2456, 2560]. Leporinus [1896]. leptocephali [3229, 2398, 1912, 825, 590]. Leptocottus [2063]. Lernaeocera [2257]. Lessepsian [935, 852, 1700, 2372]. lesser [761, 984, 2888, 2385, 1155]. lesser-spotted [761, 984, 1155]. lessons [3422, 2749, 3475, 2696, 1955, 2795]. Lesueur [2440]. lethal [2579, 1364]. Lethenteron [971, 2146]. letnica [2000]. leucas [1803]. Leucaspius [2760, 1956, 2746, 2485]. Leuciscus [2693, 1276, 1583, 1607, 3180, 3387, 3112, 1707]. leucocyte [2332, 1762, 1499]. leucocytes [1555, 2332]. leucomaenis [1704, 3102, 3122, 2299, 2133, 2958]. Leucoraja [3386]. leucostictus [2527]. leukaemia [2980]. leukocytes [1720, 2117]. Leuresthes [3267]. level [2092, 2426, 3073, 1023, 2952, 1461, 1790, 1890, 1063, 634, 2229, 2559, 980, 437, 1623, 1024, 1099]. levels [1446, 1387, 1871, 589, 2214, 466, 2902, 2073, 2555, 2793, 168, 14, 2879, 937, 3203, 1879, 649, 489, 2441, 2677, 1399, 439, 358, 1177, 797, 1710, 316, 1066, 1498, 2573, 2707, 150, 870, 879, 3277, 2957, 3346]. lewini [3216]. LH [1056, 2677]. LH- [2677]. L’Histoire [1405]. LHRHa [1821]. Liangzi [428]. liberiensis [1802]. lice [575, 2682, 1691, 1479, 1012, 1364, 1416]. lice-infected [1416]. Liem [3388]. Life [1403, 2670, 3000, 2023, 1429, 2745, 656, 2218, 53, 1425, 731, 1597, 1574, 3449, 2444, 1901, 115, 2081, 3151, 2956, 478, 2497, 3316, 2021, 2179, 2496, 189, 2638, 692, 1463, 2684, 2550, 301, 997, 2373, 1020, 2826, 1051, 2552, 933, 2585, 2945, 2841, 887, 1075, 3256, 1756, 899, 154, 950, 1978, 681, 579, 2067, 322, 366, 581, 634, 2087, 1258, 3407, 3117, 521, 3395, 3101, 3369, 1334, 29, 2501, 1120, 1970, 1575, 769, 212, 2537, 131, 2431, 507]. Life-history [2745, 2218, 53, 1425, 731, 1597, 3449, 2444, 3151, 2179, 2638, 2550, 1020, 2585, 1978, 681, 581, 3117, 3395, 3369, 29, 1575, 2537]. life-stages [692]. life-style [2552]. lifespan [691]. lifestyle [549]. Lifetime [2693, 1567]. Light [1440, 217, 1238, 2611, 544, 2974, 901, 580, 2486, 1051, 1048, 2607, 1740, 1256, 416, 2965, 2393, 2229, 316, 3098, 853, 1166, 1024, 2584, 1381]. light-induced [1381]. Ligula [3060, 531]. Ligulosis [801]. Ligurian [1749, 1735]. like [2134, 3274, 1547, 2728, 1980, 1369, 1720, 2148, 2533]. likelihood [2710]. likely [1133, 3443, 2485, 2797]. Lim [1654]. Limanda [1848, 1952, 2634]. limbatus [2941, 3003]. limestone [442]. limestone-neutralized [442]. limit [3450]. limitation [2907, 1638, 3065, 1837]. limitations [2303]. limited [3371, 3452, 1508]. limiting [951]. limits [737, 1044, 1728, 1064]. limnetic [403]. limnological [2354]. Limnothrissa [2298]. line [2448, 1112, 928, 1504, 3209, 1369, 647, 3046]. line-fishing [647]. lineage [679, 3170, 3348, 1753]. lineages [707, 2796, 2718, 2732]. linear [3076]. lineata [1834]. lineatus [2493, 458, 3177]. lined [3374]. lines [1589, 2318]. lingcod [2229]. link [2657, 2802]. Linkage [2726]. linked [3240, 3237, 348]. Linking [1195, 2775, 2160, 3361, 1952, 1954, 3138]. Links [2605, 815]. Linnaeus [2730, 2527]. linoleic [214]. linolenic [214]. liolepis [193]. lion [2183]. lionfish [2719]. Lions [3073]. Liparidae [443]. lipase [342, 341, 990]. Lipid [459, 1633, 2419, 2242, 366, 3078, 1120, 2081, 3191, 1545, 3381, 342, 3278, 1783, 3020, 489, 2833, 681, 1620, 3370, 980, 1228, 1395]. lipids [170, 2555, 874]. Lipophrys [1786, 2874]. lipopolysaccharide [1350, 2828]. Lippe [2070]. lips [1476, 1519]. List [1, 1628]. Lithuania [2690]. Lithuanian [2210]. little [1726, 3386, 2225, 1939]. littoral [1534, 1243, 1020, 2999, 5, 775, 2844]. littorale [388]. Live [3033, 2967, 1192, 2611, 3433, 980, 472, 2425, 1973]. livebearing [1218]. lived [3158, 2097]. Liver [874, 2665, 3266, 3153, 191, 427, 2294, 1312, 562, 3078, 31, 744, 390, 3128, 2190, 2533, 2318]. living [3160, 1214, 1436]. Liza [2964, 3467, 1919, 3383]. Lizard [3050]. lizardfish [2451]. LLD [645]. loach [3294, 2717, 1680, 3428, 2471, 351, 1678, 1262]. Loaches [226, 601, 2733, 2722]. load [168, 2683]. loading [1536]. local [610, 3360, 3038, 1425, 3119]. localities [1284, 1346]. Localization [3354, 648, 2145, 2542, 2467]. localized [2609]. location [1108, 3292, 3103, 3227, 1024, 32]. locations [465, 2731]. Loch [2317, 3134, 668, 960, 2317, 2893, 1539]. loci [2517, 1698, 1940, 1938, 3237, 2792, 1932]. locomotion [2449]. locomotor [297, 84, 1919, 3203, 2038]. locus [3295, 2726, 2715, 1808, 2539, 1941]. Łodz [1706, 2601]. lolo [856]. Loma [1999, 472, 471]. lombavdoi [684]. Lomond [2893]. Long [1658, 1069, 2392, 1746, 2613, 1071, 1588, 2042, 432, 3108, 2109, 995, 628, 1983, 3141, 1513, 1869, 2481, 46, 2993, 2815, 732, 1916, 2097, 517, 772, 2665, 2277, 2533, 533, 1430]. long- [1513]. Long-distance [2042, 3141]. long-lived [2097]. Long-term [1658, 2392, 1746, 1071, 1588, 432, 3108, 2109, 995, 628, 1869, 2481, 46, 2993, 2815, 732, 1916, 517, 772, 2665, 2533]. Longer [2681]. longest [2827]. longfin [373, 558, 890]. longfinned [194, 1061, 1385, 2107]. longicephalus [1664]. longipinnis [300]. Longitudinal [2169, 2362, 2445, 3008, 1455, 3112]. loop [2140]. loosejaw [2960]. lophiid [2439]. Lophiiformes [2439]. Lophius [1502, 2071]. loricariid [897, 3015, 201]. Loricariidae [2600, 3429, 2832, 1996, 3015]. loricariids [349]. Loricariinae [3015]. Losing [2951]. Loss [2536, 2327, 3247, 872, 2437, 1361, 2720, 576]. lota [2365, 3098, 2365]. lotic [1346]. Lough [2013, 3154]. Louis [3168]. Loup [1405]. louse [635, 2115]. Low [2717, 3167, 1399, 1065, 2426, 2053, 2234, 189, 3381, 3230, 2555, 444, 951, 2222, 3106, 3155, 2437, 90, 2614, 983, 83, 472, 2425]. low-cost [2425]. low-level [2426]. low-lipid [3381]. low-order [2053]. low-salinity [2222]. low-virulence [472]. Lowe [3405, 1922, 2697, 2578, 2646]. lower [19, 913, 529, 2899, 2249, 3272, 44, 967, 53, 2884, 2861, 3459, 1395, 2732, 3279, 860, 2070, 577]. lowland [2777, 1581, 2061, 239, 3273, 2416, 2417, 2455, 624, 2378, 1119, 1329]. lozanoi [3063]. LR [2825]. Ltd [1774, 1652]. Lubbock [2458]. Lucia [604, 1855, 605]. lucioperca [2723, 2778, 2954, 3162]. lucius [2973, 2936, 2903, 1757, 2143, 1399, 3253, 2187]. Luiz [364]. Lujanidae [3375]. lumbriciformis [1898, 2154]. lumholtzi [1000, 548]. Lumping [2711]. lumpsuckers [2711]. lumpus [2171]. Lunar [3261, 3269, 221, 3331, 2479, 964]. lungfish [1998, 2324, 1016, 379, 2938]. lungfishes [3330]. Luoma [3408]. lupus [2062, 1737]. Lusitanian [2306, 2991, 678]. lusitanicus [1835]. Luteinizing [649]. luteoflamulatus [1843]. Lutjanidae [3005]. Lutjanus [621, 1785, 2022, 3005, 2144, 3065, 2930]. Lütken [1896]. Lycodes [251]. lymphoid [2316, 1720]. lymphoproliferative [2673]. Lyngbya [2098]. lysine [1885]. Lythrypnus [172, 2494].

M [3479, 1402, 3478, 3405, 1653, 1068]. M-4 [1068]. M. [2858]. M1 [3080]. M74 [429]. maccoyii [2616, 3381, 1762, 3353]. Maccullochella [2561, 3273, 2433]. Macintyre [2968]. Mackerel [1916, 1013, 1147, 3157, 278, 1797, 198, 1051, 513, 2428, 2427, 1110, 564, 591, 1445, 1865, 385, 386, 770, 2344, 2108, 2568]. mackerels [1518]. maclovinus [1716]. Macquaria [3463, 1662]. macquariensis [2561]. macro [541]. macro-structure [541]. macroalga [1275]. macroalgae [2484, 552]. macrocephalus [2848]. macrochirus [3317, 3337, 2456]. macroinvertebrates [103]. macrolepidotus [3331, 2243]. macroparasite [2866]. macroparasite- [2866]. macroparasites [2510]. Macrophage [173, 2665, 660]. macrophages [562, 744]. macrophyte [522, 1304]. macrophytes [1096]. macropomum [2707]. macrops [240]. Macroramphosus [2181]. macroscopic [1306, 945]. macrostigma [3153]. macrostructure [2057]. macrourid [101, 730]. macrozoobenthos [3241]. Macruronus [419]. maculatofasciatus [1828]. maculatus [453, 2204, 2419, 2052, 3299, 2039, 1160, 2220]. maculosa [2365]. Madang [1722]. made [1533]. Madeira [2314, 1909, 2806]. madtom [3440]. Magadi [49]. Magdalena [2562]. magellanicus [44, 419]. Magnetic [3209, 3110, 3319]. magnetosensitive [1880]. Magnhagen [3320]. magnuspinnatus [757]. mahseer [3345]. maiden [2240]. main [721, 529, 1550, 2707, 304]. Maine [3386, 2495, 1933, 190, 1719, 2268, 3470]. mainland [2009, 1397, 1835]. maintains [2815]. maintenance [679]. Maitland [3409]. Major [2500, 1858, 957, 2327, 587, 1688, 508, 536, 978, 2473, 1839, 1934, 2918, 2519]. majuscula [2098]. Makaira [1695]. make [3071, 1571]. making [496]. mako [2685]. Malaŵi [488, 1533, 2872, 997, 2199, 1491, 2072]. Malaŵian [494]. malabaricus [1552, 1688, 2292, 2274, 2505, 573, 772, 2665, 3459, 2144, 3436]. Malacoraja [3470]. Malacosteinae [2960]. Maladaptation [1586]. Malakichthys [1465]. Malay [564]. malayanus [3468]. Malaysia [123]. Male [2227, 631, 2248, 1335, 120, 3071, 1394, 252, 778, 1028, 669, 368, 1451, 2970, 2643, 2461, 679, 975, 200, 779, 821, 415, 2121, 268, 632, 2504, 888, 1423, 494, 2145, 684, 1297, 100, 665, 2460, 2120, 1250, 773, 1187, 3398, 2899, 1288, 2006, 2016, 389, 1976, 95, 584, 2867, 2051, 3311, 677, 1388, 3013, 986, 3421, 1254, 1507, 1109, 777, 2370, 3425, 2137, 2677, 499, 2470, 1042, 1741, 797, 1710, 1553, 258, 56, 3, 2219, 2573, 3420, 3027, 3090, 2695, 237, 1234, 3380, 220, 2379]. male- [1710]. male-male [779]. males [2044, 1533, 2241, 1501, 1133, 1248, 1517, 2571, 1091, 2508, 1505, 2817, 345, 358, 475, 2589, 238, 3363, 838, 1395, 813, 137, 1234]. malformations [2157, 1355]. Mallotus [1815, 2295]. malma [1728, 2133]. malpighian [362]. mammalian [1720]. mammalian-like [1720]. Man [2634]. managed [3171]. Management [537, 1324, 2601, 604, 1006, 605, 1653, 3366, 882, 1195, 1946, 3170, 911, 610, 1429, 3171, 3152, 3249, 3286, 611, 617, 3310, 1954, 2612, 2795, 1772, 1773, 3408]. managing [603]. mandarin [3352]. manganese [3268]. mangrove [908, 2388, 2827, 2482]. mangroves [2284]. manifested [387]. manipulated [2460]. manipulation [684, 1081, 2763]. Manipur [226]. Manual [1775]. Manuel [364]. mapping [3293, 2734]. maraena [2261]. Maranhão [364]. Marbled [1998, 2324]. Marcusenius [2243]. Margaritifera [206]. marginal [1766, 496]. marginatus [1922, 775, 2578, 1178]. Marguerite [1748, 1642, 1333]. Marine [1406, 1404, 1278, 364, 1777, 1704, 2084, 2866, 3367, 3427, 2812, 2336, 852, 606, 1750, 2023, 40, 604, 2893, 1832, 1683, 117, 2613, 3066, 613, 3202, 771, 3029, 3133, 2412, 2879, 1040, 2826, 1980, 51, 2959, 2553, 3474, 2987, 724, 1046, 2524, 2605, 2963, 1855, 3340, 755, 1421, 1724, 2423, 1093, 3361, 1308, 2338, 3250, 822, 605, 1236, 3156]. marinus [3002, 462, 2295, 2887, 2748]. marisalbi [251]. mark [966, 1998, 619, 3376, 3365, 2930]. mark-recapture [966, 1998]. mark-recaptured [619]. marked [2251, 1588, 3109]. marker [828, 612, 1136, 2918]. markers [2725, 1489, 2924, 2740, 382, 1754, 873, 1719, 2791, 2137, 1755, 3121, 275, 276, 380, 3347, 2537, 3438]. Marking [1900, 501, 639, 1100, 3310, 2376, 1331, 1723, 2969, 3156]. marks [639, 2224]. marlin [1157]. marmorata [3226, 1544, 2293, 1093, 3324, 3239]. marmoratus [234, 2664, 2646]. Marteinsdottir [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. martensi [1162, 2145]. Maryland [1651]. maskray [3182, 3447, 3009]. masou [2736, 3111, 2441, 3140, 2040]. Mass [512, 984, 2283, 2678, 1545, 178, 1703, 2505, 253, 841]. Mass-length [512]. Massachusetts [1824]. masses [1312]. massif [659]. mastacembelid [3342]. Mastacembelidae [1802]. Mastacembelus [1802]. masu [943, 2736, 3111, 2441, 3140, 2040, 1030, 220, 570]. Matano [3222]. match [1188]. matched [1388]. Mate [2699, 779, 192, 494, 100, 177, 718, 2554, 1251, 1335, 3423, 1180, 1795]. mate-choice [3423]. material [362, 1041]. Maternal [2237, 155, 1792, 979, 2157, 2392, 3358]. maternal-care [1792]. Maternally [2277]. mating [2991, 2104, 631, 1517, 1976, 120, 205, 1254, 2664, 116, 592, 777, 1222, 2765, 820, 1874, 252, 475, 2802, 1028, 782]. Matrilinear [3288]. matrix [1694]. matter [1388]. Matthes [2132]. Maturation [845, 799, 670, 962, 1387, 3076, 1891, 2346, 1633, 24, 335, 2120, 2565, 1589, 1188, 1109, 1618, 678, 865, 1365, 258, 319, 1093, 1352, 945, 313, 347]. maturation-inducing [678, 319]. mature [528, 268, 1288, 978, 3332]. maturing [466, 532, 249, 446]. maturity [188, 620, 3386, 2670, 64, 1637, 2062, 2998, 1290, 3028, 2002, 508, 865, 1291, 2882, 2156, 28, 2057, 2268, 3470, 348, 2180]. Mauritania [2657]. Mauritius [2152]. Maurolicus [1413, 2858, 2659, 267]. maxima [2821, 2920, 1657, 1265, 3319]. maximal [3337]. maximisation [1200]. maximize [1951]. maximizing [2104]. maximum [958, 64, 2623, 2814, 1736]. maximus [1995, 2656, 2507]. May [1428, 2601, 179, 1935, 894, 784, 1367]. Mayan [3178]. MCH [99]. meadow [2813]. meadows [824]. meagre [2287]. meal [2265, 3149, 305, 298, 385, 770, 3083, 2029]. meal-induced [3083]. meals [2563]. meaning [3430]. means [1502, 880]. measure [2814]. Measurement [2888, 2993, 910, 1525, 1529, 722, 371]. Measurements [2447, 3282, 3465, 2616, 931, 830, 227, 536, 2050, 3319]. measures [286, 1647, 3434, 213]. measuring [967, 3420, 1981, 1101]. mechanical [2689]. mechanism [962, 126, 17]. Mechanisms [3429, 1204, 2337, 1440, 1381]. Mechanistic [1893]. median [2644]. mediated [1645, 3072, 2407, 2755, 2341, 2321, 2063]. mediating [2412]. mediation [1210]. medical [3217]. medicinal [3196]. Mediterranean [2597, 1727, 1852, 753, 808, 3073, 78, 944, 1409, 2599, 500, 1923, 662, 1417, 1948, 1058, 101, 3207, 901, 2171, 2250, 2353, 3215, 465, 2647, 129, 688, 2041, 2885, 1810, 751, 2949, 3030, 1783, 1075, 916, 359, 1464, 633, 195, 2276, 1829, 1755, 2372, 254, 1287, 1816, 244, 450, 3237, 306, 2697, 1985, 775, 2578, 3, 2871, 1868, 1470, 50, 2344, 2646, 2244, 512, 754, 2180, 3356]. Mediterranean-type [2276]. mediterraneus [2344]. medium [1513, 111]. medium-term [1513]. Medullary [926]. Megachasma [2891]. Megachasmidae [2891]. Megalops [2683]. megamouth [2891]. meganuclease [3120]. meiogynogenetic [795]. meiotic [2740]. Mekong [3366]. Mekongina [3042]. Melanism [1749]. melano [2665]. melano-macrophage [2665]. Melanogrammus [2773, 2264, 2090]. melanoma [2782]. Melanophore [1793, 1645]. melanophores [99]. melanops [3268, 2875]. melanopterus [3308]. melanostictus [3008, 2992]. melanostomus [1501, 1826]. Melanotaenia [3247, 3119, 3403]. melanotheron [261]. melapterus [168, 856]. melas [902, 2903, 2896]. melastomus [549, 2543]. Melatonin [3275, 1814, 1044, 1529, 1617, 316]. members [1872, 356]. membership [2216]. membrane [3130, 769]. Memorial [3478]. menhaden [2770]. menhadens [2721, 2526]. mentella [3078, 697]. mercury [2929, 1129, 2155]. meridionalis [1790, 1798, 2944]. meristic [93, 3239]. Merlangius [2264, 3014, 2196]. merlangus [2264, 3014, 2196]. Merluccius [3104, 3307, 2258, 2259, 2537]. merra [2643]. mesembrinus [420]. mesocosm [170]. mesocosms [979]. mesodermal [3433]. mesopelagic [1120]. mesopotamicus [2698]. mesotrophic [675, 1548]. messieri [2912]. meta [606]. meta-analysis [606]. Metabolic [3128, 3129, 1500, 231, 1913, 3282, 1281, 2215, 317, 1524, 27, 894, 3381, 1798, 1126, 989, 2848, 1527, 969, 511, 2826, 3203, 717, 885, 154, 1082, 526, 140, 105, 2838, 1305, 1853, 1173, 493, 841]. Metabolism [897, 27, 1850, 2579, 191, 503, 198, 84, 68, 53, 681, 1708, 1498, 1386, 1033, 1536, 1368]. metabolite [437, 1346]. metabolites [3319]. metacercariae [3148, 741]. metal [272, 161, 3408]. metals [2396, 12]. metamorphic [2611, 1460, 1283]. metamorphosing [1277, 825, 1635]. metamorphosis [233, 1972, 372, 1151, 590, 2264, 879]. metapopulation [1210, 2345]. metapterygoid [2490]. Metazoan [419, 1284, 2943, 2478, 3062, 2806]. methaemoglobin [722]. methaemoglobinemia [818]. methanesulphonate [2198]. method [528, 3282, 3076, 3255, 1513, 1607, 2714, 1293, 1984, 64, 2014, 144, 898, 255, 2273, 2674, 1769, 1879, 2321, 3138, 1679, 2216, 980, 411, 2628, 2173, 3319, 1787]. Methods [158, 3394, 1534, 457, 42, 1366, 2414, 1905, 3365, 1900, 501, 2036, 257, 1101, 561]. methyltestosterone [220]. metric [1478]. Meunier [1405]. Mexican [2517, 2099]. mexicana [2153, 2648]. Mexico [1989, 883, 2526, 1168, 1504, 2016, 1486, 2022, 2914, 1803, 1671, 919, 763, 2389, 2290, 3116, 3390, 3194, 2910, 2953, 2562, 163]. Meyer [2808, 2842]. MHC [3434, 3423, 1117, 2323, 3051, 355, 312, 1180]. MHC-dependent [3423]. Michigan [1493, 1638]. Michoacán [163]. micreosatellite [382]. Micro [945, 541]. Micro- [945, 541]. microbiota [1917]. microcephalus [2386]. microchemistry [2110, 1356]. microcystin [2825]. microcystin-LR [2825]. Microgeographic [966]. Microgobius [1348]. Microhabitat [2757, 558, 2809, 295, 1849, 2905, 1948, 1807, 2755, 2069, 2664, 1786, 881, 1236]. microhardness [2001]. microincrement [1526, 1309]. microincrementation [352]. microincrements [388]. microlepi [2348]. microlepis [2876]. micronuclei [2714]. Micropogonias [1977, 3287, 1372, 920]. microps [254, 167]. Micropterus [2956, 290, 1806].

Microsatellite [847, 40, 249, 3293, 3223, 2143, 2791, 2927, 2541, 1819, 488, 1011, 2989, 966, 2924, 2740, 2517, 1698, 2983, 3047, 3295, 3121, 3237, 2792, 2539, 276, 2918, 1432, 1522, 3438, 2524]. Microsatellite-centromere [3293]. microsatellites [689, 2735, 3431, 2718, 90, 52]. microscopy [94, 1436, 543]. microsporidians [1999]. Microstomus [3136]. microstructural [3227]. microstructure [2621, 2291, 3004, 3019, 363, 469, 3079, 1768]. microtagging [2637]. microtidal [1153]. microtitre [1981]. Middleton [3416]. midland [1921]. midshipman [3090]. Midwestern [2093]. migrant [852, 1700, 1143, 2372]. migrants [53]. migrate [839]. migrating [1715, 340, 1892, 1905, 2338, 2668, 237]. Migration [2084, 1634, 529, 2172, 2825, 1945, 1847, 1121, 3229, 2070, 2086, 2388, 3243, 1171, 2481, 3233, 399, 1851, 859, 3045, 2657, 2613, 3066, 673, 2482, 1138, 1376, 1514, 2960, 1213, 2249, 1427, 239, 2094, 2217, 1744, 633, 1220, 1285, 516, 1433, 2688, 1730, 2567, 3098, 1503, 3227, 663, 1093, 1272, 3244, 10, 1124, 1319, 3225, 3074, 1169, 3279]. migrations [1704, 860, 1685, 2481, 796, 1106, 1486, 3236, 2154, 706]. Migratory [3389, 3133, 2210, 1119, 384, 953, 1840, 1087, 2812, 1489, 1484, 839, 2202, 468, 2015, 3344, 2110, 1811, 2218, 3299, 3472, 439, 1143, 516, 2421, 321, 2748, 1140, 2545, 3449]. Mikawa [1317]. mild [2392]. milk [2105]. milkfish [3271]. mill [1002]. Miller [1497, 2804]. Mills [1774]. milt [2429]. mimicry [3454]. mimics [1202]. minckleyi [2085, 946]. mind [1223]. mine [442]. Mineral [2090]. mineralization [2001]. miniature [1041, 1540]. minimize [1293]. minimizes [2683]. mining [189]. minnow [2021, 481, 3021, 732, 2754, 2221, 1307, 377]. minnows [404, 3072, 718, 2679, 3059, 531, 978, 2277, 2095, 939]. minor [536, 2998, 2058, 2186]. minuta [335, 691]. miodon [2298]. mirror [1002]. Misgurnus [3294, 2717, 1674]. mismatch [1586]. Missouri [510, 577, 2095]. mistaken [2300]. Mitchell [2429, 2433]. Mitchill [2340, 2347, 2836, 1968, 1759]. Mitigation [2601]. Mitochondrial [34, 3440, 3437, 2521, 2793, 2719, 2147, 190, 59, 2987, 96, 564, 2520, 586, 422, 378, 971, 2990, 689, 484, 2131, 2988, 2717, 2790, 2543, 2713, 91, 1489, 712, 956, 3174, 3431, 886, 1700, 2921, 2736, 622, 3221, 3167, 2140, 1755, 1399, 292, 2855, 2919, 1518, 127, 2794, 2325, 1343, 3037, 2148, 2938, 2540, 267, 1470, 2529]. mitogenomic [1818]. Mixed [705, 3173, 1894, 2632]. Mixed-species [705, 2632]. mixing [2634]. Miyako [3132]. MØ [660]. mobile [931, 3394]. mobility [327]. modal [2249]. modalities [2370]. mode [1216, 836, 1795, 2114]. model [2803, 545, 1894, 1893, 1977, 3229, 1856, 373, 2980, 2892, 1194, 2446, 1251, 1186, 723, 227, 3413, 771, 693, 2160, 2869, 2833, 1125, 1259, 1263, 1229, 1230, 1264, 1598, 2933, 386]. Modelling [1901, 2900, 673, 3272, 1893, 1213]. Models [2296, 1181, 1195, 957, 2404, 1199, 1202, 3157, 1206, 1771, 587, 1182, 2288]. modern [911]. modes [84, 496]. modification [1275, 1679]. modifications [1321, 2909, 1544]. modified [84]. modulates [1913]. modulation [56]. Modulatory [214]. Moenkhausia [3454]. Mogurnda [1684]. mohnikei [2377]. mojarras [2532]. mola [763]. Molecular [2535, 2988, 582, 3427, 3428, 1404, 1882, 3339, 2981, 3439, 3436, 2984, 50, 973, 1131, 132, 2711, 2985, 272, 2733, 598, 2737, 3324, 3214, 2706, 2315, 2146, 3468, 2977]. molecule [3275]. molecules [1821, 3218]. molitrix [3221]. mollies [495, 523, 3213, 2099, 2153, 2648, 1862]. molly [747]. moluccensis [1700]. monacanthid [2569]. Monica [2369]. monitoring [498, 2294, 3212, 1691, 405]. Monoamine [2698, 390]. monoaminergic [1386]. monoculus [2232]. monogamous [2595]. Monogenea [2662, 3277, 3197]. monogenean [856]. monogeneans [1308, 3057]. monroviae [1520]. Montana [12]. montane [1340]. Montenegro [1538]. Monthly [857, 2609]. months [2938]. moon [2161, 1642]. moorei [1714]. mordax [2340, 2347, 2592, 1759, 2429]. Moreton [2098]. morhua [2834, 1804, 2227, 2771, 1694, 2897, 2924, 2579, 2446, 2235, 3189, 3328, 2663, 2126, 2342, 1966, 1588, 3081, 3285, 2321, 2541, 3284, 3300, 2838, 2401, 2264, 3014, 2539, 2343, 2288, 1830, 1837]. Morizot [2782]. mormyrid [2243]. Mormyridae [1648, 2570, 3092]. Mormyrus [3298]. Morone [2270, 2607, 2216, 1736, 2313]. morph [3263, 992, 2564]. Morpho [369, 556]. Morpho-anatomy [369]. morpho-functional [556]. morphoedaphic [318]. Morphogenesis [707, 565]. Morphological [1052, 2526, 2622, 2339, 2340, 900, 2543, 2185, 234, 2353, 1020, 864, 2099, 3464, 1969, 996, 3238, 2304, 3219, 1458, 946, 2560, 1883, 3239, 347, 3468, 3099, 2711, 2985, 888, 3471, 743, 2363, 2392, 725, 196, 280, 1467, 3100, 3041, 3372, 3355, 1080, 369, 1022, 3245, 3172, 2269, 2230, 2129, 2117, 3214, 428, 1767, 2689]. morphologically [1842, 2520]. Morphology [3125, 1088, 774, 2044, 1292, 3304, 1035, 3388, 3206, 842, 2340, 172, 2487, 1484, 349, 1353, 1175, 1476, 264, 2472, 1318, 2451, 1966, 2471, 2909, 2069, 2967, 2552, 3378, 2005, 552, 1999, 2119, 776, 836, 1537, 1042, 149, 2662, 7, 3, 2842, 1401, 1762, 2264, 2544, 13, 2268, 201, 2475, 2054, 1167, 1317, 1129, 400]. Morphometric [1252, 748, 3090, 3324, 839, 1591, 2292, 1370, 430, 980, 3115, 2939]. morphometrics [2807, 2152, 47, 697, 2211]. Morphometry [1161, 893, 2529]. morphotypes [3305, 651, 1469]. morphs [480, 3045, 2486, 702, 992, 1252, 374, 2668, 1522, 2564]. Mortality [1721, 1687, 1848, 1415, 3229, 855, 286, 1647, 2251, 743, 3158, 2550, 619, 1280, 1419, 3463, 429, 117, 2025, 2577, 2412, 2946, 3467, 2687, 66, 58, 2223, 2778, 3111, 1336, 2046, 3103, 978, 1760, 647, 3008, 2808, 801, 3109, 353, 1264, 2880]. mortem [1046, 1492]. mortis [1492]. mosaic [747, 2722]. mosaics [1157]. mosquitofish [3077, 3013, 896]. mossambica [3227]. mossambicus [975, 1506, 3260, 3196, 3203, 155, 1498]. most [496, 2498, 2906]. mothers [3445]. motility [2396, 114, 1163, 893, 1076, 895, 1528, 671, 1450, 720, 2805, 2365]. motion [2456]. motivation [894, 177]. motor [926]. mound [1158]. mounds [3380]. mount [2313]. mountain [2368, 2319, 2756]. mounting [2716]. mouth [2020, 1476, 3398, 1040, 2082]. mouth-brooding [2020, 3398]. mouthbrooding [266, 345, 813]. Movement [1687, 2368, 2371, 328, 513, 1855, 2112, 1577, 3141, 2641, 816, 1243, 2023, 1206, 2147, 2184, 3025, 239, 3180, 1391, 760, 1340]. Movements [1662, 1840, 230, 539, 1332, 1504, 1811, 2042, 2337, 1110, 323, 1635, 1145, 2637, 405]. Moxostoma [886]. Mozambique [3260, 866, 3203]. mrakovcici [3199]. Mrf4 [3034]. mrigal [3198]. mrigala [3198]. mRNA [2739, 33, 14, 2923, 1441]. MS [1127]. MSH [99, 1820]. mtDNA [3290, 790, 1860, 3291, 641, 1753, 1070, 2718, 2527, 2321, 275, 1401, 3219, 2148, 52]. mucin [2480, 351]. mucosa [1749, 648, 427, 2480, 1644]. mucosal [825]. mucus [2407, 2509, 434, 1179, 1742, 1279, 2280]. mucus-mediated [2407]. mud [3366, 2284, 351]. mudskipper [2809, 2552, 757]. muelleri [1413, 2659]. Mugil [2709, 3291]. Mugilidae [2709, 3467, 1497]. mugilids [2964]. Muller [2631, 2800, 2408, 2659, 2837]. mullet [500, 2709, 1486, 3291, 3383, 229, 2566]. mullets [3073]. Mullidae [2039]. Mulloidichthys [2152]. Mullus [3073, 901, 1275]. Multi [3360, 3273, 3255, 1478, 2764, 1844, 1455]. multi-method [3255]. multi-metric [1478]. Multi-scale [3360, 3273]. multi-species [1844]. multi-stock [2764]. multi-year [1455]. multicolor [783]. multifiliis [174]. multilayer [2644]. Multiple [1874, 2563, 1704, 884, 2294, 3348, 3047, 1753, 820, 1101, 3257]. multiple-spawning [3257]. multiplex [2724]. multiradiatus [1251]. multispine [2649]. multispinis [2649]. multivariate [907, 468, 1746]. mummichog [2647, 3331, 1097]. Munro [2660]. muriei [1800]. Muros [228]. Murray [1306, 3286, 3273, 1662, 2433]. Muscle [1023, 110, 891, 1622, 1739, 1885, 1694, 3093, 1417, 2387, 338, 1725, 1126, 2899, 987, 1483, 2245, 3078, 130, 2906, 3091, 82, 121, 2533, 447, 810, 1624, 3080]. muscle-based [2387]. muscle-specific [3080]. muscles [2089]. musculature [3382, 2802]. music [2965]. muskellunge [27, 657]. mussel [1184, 3070, 1259]. mussels [2168, 2008, 2415, 3012]. Mustelus [211, 2953, 2871]. mutation [3045]. mutualism [1201]. myc [1149]. Mycobacterium [846]. Mycteroperca [2876, 2348, 2290]. myctophid [2840, 1899, 1166]. Myctophidae [2911, 1397]. Myctophiformes [2911, 1397]. Myctophum [2840]. Myf5 [3034]. mykiss [3130, 3302, 1907, 3368, 2557, 2908, 1610, 2942, 2623, 3024, 3327, 3139, 2945, 2881, 2116, 1462, 2765, 3441, 2312, 1841, 2442, 2038, 3393, 2989]. Myliobatoidei [3182]. myocardium [2582]. MyoD [752]. myofibrillar [987]. Myogenesis [3218]. Myogenic [3034, 752]. myology [1141, 458]. myonuclei [2245]. myosatellite [810]. myosin [338, 3343, 1744]. myotomal [2899, 2906]. Myotube [2899]. myristate [1555]. mystery [223]. mystus [2238]. myth [1216]. Myxobolus [354]. Myxus [2964, 3467].

N [1406, 3323, 3406, 3408, 214, 459, 1835, 404, 2555, 2783, 2164, 1144, 2586, 2489]. n-3 [214, 459]. n-6 [214]. N. [193]. naccarii [1321, 2409]. NADH [1860]. NADH-1 [1860]. Naivasha [2758]. naïve [553, 1330, 2452]. naked [3438]. namaycush [2550, 2286]. name [3168]. Names [1, 1628]. Namibian [981].

Nannoperca [2671]. Nantes [1901]. NAOI [1290]. Nardo [2353, 2855]. narrow [513, 1110, 2162]. narrow-barred [513, 1110]. nase [1606, 1615, 440, 3214]. Naso [848]. Nassau [2225]. nasus [1847, 1606, 1615, 3159, 3214]. Nations [1653]. Native [2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513, 1852, 1558, 1920, 2757, 2667, 2319, 3170, 2751, 462, 2756, 862, 2743, 2745, 425, 290, 2059, 2753, 2222, 2759, 3021, 3122, 1987, 2868, 2896, 1567, 2968, 2731, 2329, 1554, 598, 3178, 1846, 1862, 2750, 1233, 2095, 163]. nattereri [2638]. Natural [2021, 1067, 1336, 3294, 1578, 3193, 3206, 1241, 268, 3069, 2678, 817, 3429, 784, 310, 1976, 3212, 2471, 2700, 3397, 1857, 1591, 3423, 309, 2458, 1176, 2193, 1567, 2861, 1621, 1182, 2003, 767, 1274, 1535, 782, 1707, 3214, 2224]. naturally [2690]. nature [3302, 1580, 1473]. navigation [882, 1107]. ND4 [788]. ND4L [788]. Neagh [2013]. Near [223, 494, 2517, 1530, 2650, 1064, 2112]. near-extinct [2517]. Near-surface [223]. nearly [2148]. nearshore [397, 1099]. Neaxius [2851]. Nebish [46]. nebulosus [819, 2004, 2937]. necrosis [2982]. need [1196]. needs [3476]. Negaprion [2079, 3283]. Negative [396, 2558, 1393, 533]. Negro [1446, 425]. neighbouring [892]. nekton [1671]. Nemachilus [1262]. Nematoda [3063]. nematode [961, 563, 680, 1312]. Nematolebias [2619]. nemipterid [734]. Nemipteridae [2949]. Nemipterus [2041, 2949]. Nemo [2951]. neogaeus [1872]. Neogobius [1501, 1826, 2731]. Neoheterobothrium [3277]. Neolamprologus [1511, 3266, 1432, 3173]. Neomerinthe [889]. Neonatal [982, 683]. neonate [2941]. neonates [1027]. Neoparamoeba [2332, 2922]. Neoparamoebic [3054]. Neoteny [1379]. neotropical [315, 1085, 2081, 3270, 349, 772, 812, 2493, 2775, 2274, 3436]. Neotrygon [3182, 3447]. neozoic [3061]. Neptune [2468]. Neretva [2522]. nerka [2461, 1637, 3041, 2289, 3301, 3184, 3358]. Nerophis [1898, 2154]. nerves [370]. Nest [1985, 2066, 798, 2460, 391, 2867, 252, 2950, 3147, 2163, 2768, 3380]. Nesting [2883, 2219, 200, 2121, 1741, 1985, 3090]. nests [528, 3071]. net [954, 286, 588, 2831]. net- [2831]. Netherlands [923, 2278, 1119]. nets [3246, 2004, 2045]. Nettastomatidae [1912]. netting [1554, 2568]. network [1213, 1897]. networks [215, 1215, 2644]. neumayeri [1509]. neural [1213, 2644]. neurectodermal [3433]. neuroendocrine [2134]. neuromuscular [16]. neuronal [1076]. neurones [926]. neurons [3354, 2313]. neurotransmitters [1386]. neutral [1700, 3041]. neutralized [442]. newborn [2888]. Newfoundland [3367, 1332, 2447, 3045, 253, 1145, 35, 1939]. Newly [903, 1415, 607, 3368, 571, 2025, 1331]. newnesi [2467]. News [1773]. nezahualcoyotl [2782]. Nezumia [193, 2943]. niche [500, 3357, 3372, 2955, 2095]. niches [2954]. Nielsen [502]. niger [2394, 777, 2570, 778]. night [1748, 1044, 2273, 3391, 2231]. nigra [803]. nigri [2644]. nigricans [2713, 433, 1695]. nigrofasciatus [2248]. nigroviridis [873]. Nile [1485, 2534, 3274, 2555, 696, 2917, 2632, 677, 716, 1359, 1164, 330, 441, 1136]. niloticus [1885, 2534, 3274, 2555, 18, 215, 2917, 2632, 2691, 1526, 1309, 2997, 2527]. Nilsson [915]. nimbaria [1021]. nine [1139, 1845, 782, 2322]. nine-spined [1139, 1845, 782, 2322]. ninespine [2926]. Ninth [877]. nipponensis [2084]. nitidulum [2840]. nitrite [350, 438, 818]. Nitrogen [2359, 1536, 2684, 2942, 818, 754]. Nivelle [2083]. Nkhata [2199]. No [3069, 1700, 3222, 1653, 2044, 3097, 511]. nobilis [2457, 3221]. Nocomis [3380]. Nocturnal [1748, 57, 307, 2999, 1426, 1642, 1107, 881]. Nodularia [2801]. nodulosus [2558]. noise [2221]. Non [152, 2279, 2309, 2381, 2436, 2465, 2492, 500, 2753, 2411, 2764, 1770, 2513, 1777, 418, 1920, 2667, 466, 1889, 2751, 1750, 862, 1525, 2743, 2782, 3434, 2517, 2749, 2840, 2759, 2772, 3293, 2921, 3122, 3378, 2004, 1879, 1436, 2200, 1567, 1905, 3178, 3420, 2668, 2752, 1939, 1838, 3400, 2503, 2028, 3346, 3356, 2750]. non-additive [3434]. non-anadromous [1939]. non-coding [2921]. Non-competitive [500]. Non-destructive [1770, 1905]. non-diadromous [2772]. Non-direct [2764]. non-electrolyte [3378]. non-endangered [2517]. non-esterified [466]. non-feeding-stimulant [2028]. non-functional [418]. non-homologous [3293]. non-hybrid [2782]. non-indigenous [862, 2749, 2004, 2752]. Non-invasive [152, 1889, 1525, 1879, 3420, 3400]. non-invasively [1436]. non-isolated [2200]. non-marine [1777, 1750]. non-migrating [2668]. Non-Native [2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513, 2753, 1920, 2667, 2751, 2743, 2759, 3122, 1567, 3178, 2750]. non-paired [2840]. non-reproductive [3356]. non-specific [1838]. non-target [2503]. non-transgenic [3346]. norepinephrine [2073]. Normal [3433, 3206, 721, 1664, 2266, 1049, 1264]. Norman [2780]. normani [1468]. noronhanum [67]. North [346, 1901, 1830, 3228, 3185, 3234, 620, 3073, 3191, 2388, 3243, 1417, 2481, 1684, 3000, 730, 737, 2079, 1504, 2657, 3289, 2184, 375, 1486, 3446, 3291, 2217, 1502, 3019, 2784, 1783, 1719, 2428, 1835, 3375, 2135, 994, 2963, 3364, 1985, 1819, 2500, 1479, 364, 281, 228, 1808, 3064, 2183, 2539, 603, 464, 3186, 2288, 2244, 2685, 1396, 2106, 908, 2070, 2857, 179, 1733, 15, 2624, 2836, 1833, 2948, 902, 394, 1952, 223, 193, 1210, 1290, 2730, 2858, 190, 923, 3226, 68, 2181, 1931, 2791, 1918, 1220, 2337, 169, 371, 467, 2401, 1796, 1065, 469, 29, 809]. North [1343, 2529, 1932, 2088, 2345, 3419, 2365]. north-central [3064]. north-east [2481, 1684, 3000, 730, 2079, 1504, 2184, 1719, 2428, 1835, 2963, 3364, 364, 1808, 2183, 2539, 3186, 2288, 394, 223]. north-eastern [375, 3446, 1502, 3019, 2135, 464, 2244, 1396]. north-west [3234, 3073, 3243, 2657, 3291, 2784, 1819, 2500, 228, 2685]. north-western [3185, 620, 1417, 3289, 1486, 2217, 1783, 3375, 1985, 603]. Northern [1937, 302, 2013, 287, 1456, 3077, 1453, 2599, 2319, 179, 2270, 982, 3389, 3190, 2827, 2978, 943, 2016, 2037, 1922, 2022, 3446, 3029, 560, 641, 2547, 3278, 509, 2967, 3106, 2868, 863, 828, 2360, 3331, 633, 1110, 579, 1803, 178, 1284, 955, 2953, 376, 253, 663, 1367, 2144, 1808, 294, 2875, 1709, 2054, 2933, 138, 1064, 1286, 240, 1034, 2738, 492, 449, 3176]. northernmost [1398, 1288, 2736]. northward [2217]. northwest [508, 617, 597]. Norway [2753, 3202, 560, 3133, 2754, 29, 663, 203, 1709, 1469, 236]. Norwegian [3076, 737, 1634, 1690, 949, 1425, 2421, 1421, 1479, 1272, 3014, 26, 10]. notata [836]. notatus [345, 3090]. NOTCH1 [2980]. NOTCH1-induced [2980]. Note [1271, 2071, 2142]. Notes [2193, 1628, 1521, 2487, 251, 2372, 3334, 2386]. Nothobranchius [3101]. Notolabrus [2120, 1770]. notomelas [2862]. notothen [2780]. Notothenia [752, 274, 660]. nototheniid [1836]. Nototheniidae [2467]. nototheniids [1739]. notothenioid [1716, 110]. notothenioids [1666, 322]. Notropis [2591, 3169]. Noturus [3440]. nov [1680, 44, 3199, 889]. Nova [3349]. Novel [63, 2138, 1947, 1242, 1512, 894, 2924, 2582, 641, 3071, 1114, 2010, 2289, 1369, 395, 3319, 2636]. novelty [1799]. novemaculeata [3463]. November [1824]. novice [1041]. Nuclear [91, 3250, 2546, 3428, 1489, 3431, 1700, 1755, 2321, 2209, 3319]. nuclei [2207]. nucleic [469, 2540]. Nucleolar [2430]. nucleolus [1544, 2862]. nucleotide [788]. null [2404]. number [1604, 2899, 403, 3140, 1508, 2429, 638, 3306]. numbers [631, 901, 2137, 1179, 454]. numerical [673]. Nunavut [651, 3372]. nuptial [576]. nursery [346, 262, 2510, 2222, 2007, 1829, 2992, 807, 1099]. nursery-continental [2510]. nutrient [1672, 1418, 2557, 1304, 2873, 365]. nutrition [2090]. Nutritional [1277, 1813, 2502, 2430, 2188, 730, 1280, 519, 1570, 441].

O [1495, 346]. O-group [346]. Ob [3219]. obesus [3334]. Obituary [618, 645]. obligate [2675, 2655]. observation [859]. Observations [394, 3401, 253, 1906, 2424, 2948, 1411, 1502, 854, 149, 2230, 1432, 1266]. observed [3110, 479]. obstacles [161]. obtained [1578, 2419, 2617, 1710, 1855, 1472]. OC [645]. occipital [2975]. occupancy [3075]. occupying [2339]. occur [3080]. Occurrence [637, 2139, 2624, 789, 1922, 2650, 1677, 1519, 2377, 3417, 3294, 162, 1521, 642, 1729, 1606, 106, 688, 680, 3043, 23, 1520, 547, 2278, 3334, 1178, 1479, 1868, 3374, 3468, 3356]. occurrences [3334]. occurring [2964, 1520]. ocean [414, 2826, 985, 2963, 689, 3141, 2780, 2724, 2184, 2226, 2249, 1051, 2398, 3226, 1502, 2174, 2961, 3375, 2882, 2963, 244, 2152, 3364, 1065, 3245, 2183, 276, 2984, 3324, 2880, 1021, 3353, 3374]. ocean-ranched [985]. ocean-type [414]. Oceanic [3225, 3287, 229, 2220, 3228]. Oceanography [1010, 1229, 1406]. ocellatus [2168, 2008, 1866]. October [1824, 1965, 2100]. odd [653]. odd-broodline [653]. Odontesthes [2430, 2677]. Odontobutidae [1497]. Odontobutis [2921]. odorant [1212]. odour [965, 293, 2015, 1335]. odour-based [1335]. odours [1501, 293, 866, 1826, 1117]. oesophagus [1328]. oestradiol [1387, 466, 2413, 489, 530, 451, 1289]. oestradiol-17 [1387, 451]. oestrogen [150]. off [883, 2439, 642, 1397, 2079, 630, 2899, 2657, 2634, 3280, 1835, 1812, 2943, 799, 175, 2054, 240, 1034, 406, 421]. offshore [1808]. offspring [2157, 2392, 2248, 1081, 702, 1874, 2589, 1122, 3358, 1692]. often [3285]. Ogilby [2092, 2766]. oglinum [3019]. Ohrid [2129]. ohridanus [2129]. oil [699, 284, 2272, 2920, 255, 1371, 2593, 3370, 3078, 2190, 650]. okinawae [2402]. Okinawan [3005]. old [944, 2561, 708]. Olde [1955]. Older [3457]. oldest [1557]. olfaction [2030, 1077]. Olfactory [2020, 866, 2509, 2330, 3339, 3027, 1267, 1556]. Olifants [1374]. oligotrophic [764, 2450, 2868, 1065]. olivaceus [1249, 3385, 2792, 3455, 3277, 2064]. olorum [2725]. Oman [2105]. omanensis [2105]. Omitted [998]. omnivore [552]. omnivorous [315, 396]. Omono [782, 2322]. Ompok [1036]. Oncorhynchus [2461, 3130, 3302, 2812, 1907, 3368, 2557, 2908, 1610, 2901, 2942, 1637, 2623, 3024, 2736, 3327, 3041, 2945, 2881, 3111, 2116, 2686, 2137, 2441, 2289, 3140, 3301, 1462, 2765, 3384, 3441, 2040, 2320, 2312, 3195, 3184, 1841, 3361, 2442, 2038, 3393, 1499, 3358, 3204, 2615, 3404, 2990, 2989, 3279]. Oncorhynhus [3139]. one [684, 1376, 1993, 708, 678, 287, 3346]. one- [708]. one-sea-winter [1376, 287]. one-year [3346]. Oneida [1746]. ongoing [1233]. only [2804]. Onset [2057, 2825, 2900, 2390, 1179, 28]. onshore [1285]. Ontario [2080]. ontogenesis [744, 141]. Ontogenetic [814, 2832, 1550, 2069, 1673, 1757, 102, 2292, 2495, 81, 1956, 2243, 881, 3095, 1353, 139, 2373, 1718, 1949, 2837, 1786, 363, 2669, 3281, 2839, 1782, 2416]. Ontogenic [1703]. Ontogeny [2400, 1076, 3382, 990, 1257, 2316, 1307, 2313, 1052, 3304, 1506, 3072, 360, 665, 1832, 3033, 196, 3001, 3256, 1619, 2606, 1341, 2602, 123, 1806, 2280, 216]. Onyango [3407]. Oocyte [620, 151, 3315, 2472, 319, 1441, 150, 1392, 201, 3135, 2288]. oocytes [2467]. oogenesis [1544]. oophagous [2948]. Opeatogenys [1810]. open [901, 619, 3372, 1104, 2167]. open-water [1104]. operated [1502]. Operational [61, 2867, 1222]. operatively [2849, 1177, 1432]. Opercular [2165, 1173, 2456]. ophidiid [502]. Ophidiidae [1477]. Ophidiiformes [2620]. ophiocephalus [2394, 3031, 1254]. Ophiodon [2229]. Ophthalmolebias [2619]. Ophthalmotilapia [3398]. Opisthonema [3019]. Opportunistic [2776]. Opportunities [2601]. opportunity [61]. Opsanus [1832, 3392]. opsariichthines [2706]. opsin [793]. optimal [42]. Optimization [1870]. optimizing [1180]. optimum [814]. options [616]. Oral [924, 1959, 2349, 339]. orange [3039, 3152]. orange-spotted [3039, 3152]. order [2053]. Orectolobus [3017]. Oregon [1305, 1393]. Oreochromis [1885, 975, 2534, 1506, 3274, 3260, 2555, 71, 18, 215, 3196, 2917, 2632, 3203, 2691, 155, 1526, 1309, 2997, 2527, 1498, 2673]. Oreoglanis [542]. Orestias [3125]. Organ [571, 3092, 1657, 2243]. organic [1528]. Organised [2077]. organism [1949]. organism-environment [1949]. Organized [1495]. organizer [2862]. organoleptical [1594]. organs [196, 1688, 3016, 497, 988]. orientalis [2584]. orientation [2608, 1183, 798]. origin [607, 3231, 1610, 3300, 2159, 2136, 2795, 810, 3404]. Origins [3089, 1753]. Orkla [3202]. ornamental [1605]. ornamentation [3466]. ornate [3017, 3021]. ornatus [3021]. Oscillation [1290, 15]. Osmeridae [2222, 3289]. osmerids [271]. Osmeriformes [1818]. osmeroid [324]. Osmerus [2340, 2347, 2592, 1759, 2429]. osmolarity [1836]. osmolyte [2815]. Osmoregulation [122, 2265, 1612]. osmoregulatory [30, 470, 520, 2485]. osmotic [2365]. ossification [326, 556]. ossifications [3471]. Ostariophysi [1326, 1714, 2676, 3471, 3430]. Ostariophysii [2314]. Osteichthyes [2171, 2250, 3215, 1707]. Osteoglossomorpha [2570]. Osteologic [326]. Osteology [1141, 1537, 458]. OTC [3019]. other [2011, 40, 1141, 725, 3050, 1698, 1520, 3267, 90, 3334, 1661, 1600, 2386]. otic [369]. Otolith [1453, 1546, 1419, 1763, 3004, 3019, 2553, 3453, 2957, 2803, 1339, 2207, 392, 1977, 1407, 2125, 1972, 2261, 2202, 2616, 387, 1766, 2213, 303, 3230, 725, 2355, 2110, 2633, 1356, 2291, 1447, 2342, 352, 2659, 3461, 3372, 3392, 716, 2119, 3268, 1899, 3081, 363, 2401, 388, 2152, 469, 2057, 3227, 3079, 2210, 3074, 2344, 1167, 2820, 3065, 493, 1768, 1030, 2238]. Otolith-based [1419]. otolithic [769]. otoliths [827, 1345, 1035, 1127, 1116, 977, 233, 1481, 3075, 1383, 261, 2770, 1029, 830, 465, 898, 2727, 2031, 2482, 2730, 2959, 1526, 1309, 2298, 369, 3376, 1366, 3008, 824, 501, 2036, 1723, 780, 1161, 2196, 3353, 2969, 1767]. outbred [1573]. outbreeding [1625]. outfall [2112]. outflow [3166]. outlet [2843]. output [317]. ova [490, 1870]. Ovarian [2816, 1896, 760, 1317, 1058, 1163, 1069, 2294, 836]. ovaries [354, 2471, 2596, 1830]. Ovary [1886, 1821, 259, 2518]. ovata [3063]. over-compensation [446]. Over-summering [3186]. over-wintering [412, 652, 1459, 1190]. overestimation [3019]. overexploitation [1367]. overflow [1958]. overhead [1892]. overlap [500, 862, 3350, 2095]. overlaps [938]. overlooked [1810]. Overwinter [1545, 3029]. overwintering [2687, 53, 681, 3312]. oviparity [836]. Oviposition [353, 2487, 1230, 2695]. ovipositor [2008]. Övre [3139]. ovulated [220, 570]. owner [3067]. ownership [1603]. oxbow [1316]. Oxford [1402, 1773, 1774, 1652]. oxidase [390, 2698, 2331]. Oxidation [2190]. oxidative [771, 3404]. oxide [404, 1136]. Oxleyan [2671]. oxleyana [2671]. Oxygen [511, 3131, 1017, 938, 45, 350, 1384, 555, 2282, 783, 1907, 77, 931, 2684, 2944, 1765, 2031, 310, 3106, 985, 717, 2817, 517, 999, 83, 748, 1033, 2228, 2628, 1973, 1173, 2683]. oxygenation [2902]. oxygeneios [1398]. oxyrinchus [1643, 1833, 2685]. oxytetracycline [2645, 1900, 3109]. oyster [128, 16].

P [1497, 1654, 648, 1775, 3321, 1652, 1494, 3407, 3480, 3409, 3319]. Pacific [1428, 3226, 1065, 883, 2866, 1805, 3132, 3207, 730, 931, 2582, 630, 3307, 2948, 2041, 2848, 2369, 3289, 2184, 1763, 1437, 2291, 2226, 3291, 2082, 1115, 2858, 2249, 2217, 787, 1649, 1045, 2002, 3030, 102, 147, 994, 1366, 1258, 2562, 1401, 2943, 640, 2183, 3324, 3172, 807, 1024, 1424, 2584, 2529, 597, 2088, 2683, 1451, 2990, 3228, 762]. pacificus [1815, 2670, 2386]. pacu [390]. paddlefish [114, 161, 2316, 1720]. Padogobius [2713, 1162, 2145, 433]. paedomorphic [335, 2886, 787, 1661]. Pagellus [3149, 1618, 1619, 1622, 1623, 3091, 1597]. Pagothenia [1060, 273]. Pagrus [1031, 1492]. pair [403, 3447]. paired [2840]. pairs [1388]. pale [396, 744]. pale-grey [744]. palearctic [275]. Palk [2377]. Pallas [2608, 2610, 1860, 2056, 2422]. pallasi [1499]. pallasii [3132]. pallid [2828, 1787]. Palmyra [3308]. palpation [1769]. paludica [654]. paludinosus [2929, 1799]. Pamvotis [2784]. Pan [1808, 2539]. Panama [3347]. pandora [1619, 1597]. panel [2568]. Panfili [1775]. pangasiid [127]. Pangasius [2613]. panizzae [1221]. panmixia [2707]. pans [1613]. Pantophysin [1808, 2539]. paper [1002]. Papers [924, 1959, 2349, 569, 479, 3473]. papilla [172]. papillomata [530]. Papua [1722, 3151]. Parablennius [2535, 1250]. Paracheirodon [2712, 2437]. Parachromis [1702]. paradigms [508]. Paradiplozoon [2662]. paradisefish [534]. paradox [3240, 3388]. paradoxus [2854, 2537]. Paragobiodon [2675]. Paralabrax [1975, 1828]. Paralichthys [1249, 2730, 3385, 2792, 3455, 3277, 2064]. Paraliparis [443]. Parallel [2524]. parallelus [3048]. parameters [1583, 114, 972, 887, 742, 3447, 2614, 995]. Paraná [2528, 1326, 2862, 3177, 201]. Paranapanema [3459]. Paranotothenia [1869]. parapatric [2817]. Parapercis [2661]. Parascylliidae [2487]. Parascyllium [2487]. parasite [1239, 2622, 1920, 3053, 2332, 856, 2909, 2775, 2257, 1719, 1677, 479, 1259, 228, 3128, 3244, 3052, 1862, 3065, 2108]. parasite-induced [2909]. Parasites [779, 2934, 1255, 1543, 1308, 1057, 1911, 2870, 2807, 2262, 3062, 756, 1690, 3049, 513, 1110, 419, 1284, 2806, 2372, 2943, 2066, 1854, 2591, 3176, 2478, 3276]. Parasitic [1699, 1254, 2395, 9, 3303]. parasitism [1131, 3398, 3059, 1005, 3387, 167]. paraspermatozoa [1411]. Paratrygon [3160]. Parcel [364]. parentage [2115, 3173]. Parental [3024, 417, 1181, 2443, 1003, 1752, 819, 3071, 1227, 1482, 3301, 2470, 3300, 1795, 1122, 2873]. parents [983]. Parin [2858]. Park [364]. parmifera [3124]. Parodontidae [3435]. parr [57, 2907, 528, 1195, 188, 1891, 1748, 268, 1200, 532, 1382, 148, 1288, 1426, 1642, 329, 627, 11, 2923, 976, 1145, 1510, 870, 2962, 1231, 2835, 2190, 1500, 220]. parr-smolt [2190]. parrotfish [2822, 3448, 2158]. part [3307]. Parthenogenesis [3003]. parthenogenetic [2784]. Partial [132, 1433, 1053, 3037, 748, 348]. partially [212]. participation [843]. particle [3229]. particles [1765, 3209]. particular [1937, 373, 3443, 3373]. partitioning [2932, 2263, 2551, 3452, 2846, 623, 2635, 2194, 440, 1759, 1953, 2837, 1333, 1365, 2167, 3281, 2839]. partitus [2125]. partner [1188]. parts [1497]. parva [1605, 2757, 851, 1538, 2760, 789, 2278, 2221, 2746, 2485, 3099]. parvalbumin [987]. pass [431, 1079]. passage [2204, 210, 3184, 1503, 44]. passive [639, 872, 307, 1540, 3212, 3108, 3313, 405]. past [3051]. Patagonia [211, 3452, 1922]. Patagonian [1084, 2207, 2368, 2912, 2033, 2868, 776, 557, 380]. patagonicus [44]. Patagonotothen [2780]. patch [1515]. patches [1176]. patchy [3452]. Paternal [1334, 1030, 2241, 2237, 345, 2163, 2768]. paternity [3047]. pathology [206]. pathway [1967]. pathways [3427]. Patos [546]. Pattern [54, 335, 638, 1945, 1852, 1418, 1645, 2940, 297, 2893, 2942, 156, 673, 401, 1565, 3331, 3081, 2206, 1166, 2003, 813, 802]. pattern-related [1645]. Patterns [816, 862, 425, 1950, 414, 2087, 1392, 3279, 1578, 1131, 1848, 2516, 3076, 1748, 523, 2388, 659, 2185, 2641, 2261, 1546, 2023, 92, 1422, 3360, 864, 2679, 3431, 1746, 1186, 2660, 1612, 1811, 1426, 2226, 1530, 756, 3025, 1550, 992, 3288, 1688, 3041, 2587, 2397, 1898, 1938, 1257, 3258, 1795, 2605, 479, 2992, 1855, 119, 2894, 1503, 812, 2483, 1656, 2210, 395, 3057, 2880, 1432, 2236, 1374, 1057, 2378, 2112, 798, 1268, 2629, 1329, 3200, 207, 365, 1782, 3468, 1531, 1532]. Pátzcuaro [163]. pauciradiatus [1649]. pause [126]. pause-travel [126]. pavement [1321]. pavo [2554, 3147, 3027]. PCB [1968]. PCB-contaminated [1968]. PCNA [752]. PCR [2739, 2131, 2724, 641, 2140]. PCR-based [2724]. PCR-RFLP [2131]. peacock [2554, 694, 773, 1091, 358, 3147, 3027, 1442]. peak [2944]. peaking [2999]. Pearl [237]. pearlfish [1283]. pearlside [2659]. pebble [3380]. pectinata [2819]. pectinirostris [2809]. pectoral [2676, 1141, 89, 2089, 869, 458, 2456]. peduncle [534]. peelii [3273, 2433, 3273, 2433]. pejerrey [2430, 2677]. Pelagic [2580, 1724, 131, 2065, 1732, 289, 2624, 1020, 5, 1971, 2960, 2005, 2160, 2421, 3385, 169, 755, 1367, 3064, 1873, 1274, 2054, 257, 3390]. pelagios [2891]. pelamis [2140, 2927]. pelleted [1136]. pelvic [487]. Pelvicachromis [3318]. Peninsula [851, 582, 2850, 2896, 2708]. peninsular [2721, 123]. People [3405]. pepsinogen [3039, 325]. peptide [2134, 648]. Perca [1633, 2558, 1545, 2438, 3004, 2159, 1551, 2123]. percentage [2441]. perception [2644]. perch [957, 569, 2972, 2368, 1685, 2362, 2445, 1014, 2251, 2363, 303, 2438, 1100, 696, 2450, 1350, 2828, 446, 2469, 146, 998, 5, 2671, 3004, 1438, 1114, 1327, 1594, 543, 186, 1553, 1440, 72, 2489, 2375, 353, 690, 1693, 2598, 1064, 2378, 1638, 1454, 2123, 2318]. Percichthyidae [2561]. Percichthys [2368]. Percidae [1614, 2778, 2794]. percids [675]. Perciformes [2532, 1398, 2988, 2822, 1649, 2949, 51, 2847, 1520, 3375, 2329, 3037, 2142, 2132, 2811, 1465]. perciformis [2836]. Percilia [3107]. percnurus [2056]. percomorph [2532]. Percophis [2934]. percula [2497]. Performance [1461, 405, 1523, 1895, 2262, 3305, 2283, 610, 2447, 1203, 1850, 2579, 931, 2818, 291, 3097, 1126, 444, 2848, 2623, 1151, 481, 3137, 432, 701, 3183, 985, 1919, 1452, 3111, 1567, 1082, 626, 3301, 3384, 1160, 1620, 2186, 1296, 2954, 976, 320, 1623, 1596, 2228, 1542, 846, 1576, 121, 181, 3358, 206, 1012, 1364, 2305, 3306, 2689, 367, 835]. performances [1842]. perfusion [761]. peri [1460]. peri-metamorphic [1460]. period [311, 3132, 2122, 125, 15, 301, 1637, 1785, 2408, 693, 1452, 2427, 1568, 2455, 140]. periodicity [1383, 2630, 2022, 3261, 2455, 169, 3447, 1389]. periods [1908, 935, 2368, 1829, 1466, 2476, 2231]. Periophthalmus [2552, 757]. peripheral [2714]. peristalsis [17]. permanent [1153]. peroxidation [3020]. Peroxide [771]. Peroxide-triggered [771]. Persian [1846]. persist [2026, 2277]. persistence [1987, 559]. Persistent [2663, 3301, 3011]. perspective [410, 3034, 1801]. perspectives [2985, 2753, 2164, 63, 1600, 762]. perturbations [1309]. Peru [3449]. Peruvian [2034]. pest [2754]. petenensis [2099]. Peter [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. petersii [1648, 1390]. petimba [1521]. Petitgas [1901]. Petrocephalus [3092]. Petromyzon [3002, 2887, 2748]. Petromyzontidae [3437, 2793]. Petroscirtes [2163]. Petrovich [2382]. Pfeffer [1537]. pH [1555, 3282, 2356, 1015, 189, 927, 2437, 1033, 2399, 3319, 1536, 3080, 2365]. pH-resistant [1555]. PHABSIM [791]. phagocytic [2312, 2673]. Phagocytosis [660]. pharyngeal [1799, 213]. pharynx [3362]. phase [2015, 529, 1642, 2412, 1880, 1813, 2518, 131, 1059]. Phases [516, 2358]. Phenology [2401]. phenomena [1983, 767]. phenotype [2494]. phenotypes [3351, 389]. Phenotypic [153, 861, 1370, 137, 1935, 2373, 2469, 2909, 3183, 3391, 254, 1586]. phenoxyethanol [2198]. pheromone [510, 3011, 2748, 368]. pheromones [384, 953, 2890, 941, 676]. Philip [3408]. Philippines [695]. Philometra [3063]. Philometva [680]. pholis [1786, 2874]. phorbol [1555]. phorbol-12-myristate-13-acetate [1555]. phosphate [3319]. phosphoglucomutase [1941]. phosphorus [2359, 1150]. photo [142]. photo- [142]. photographic [2230]. photography [1247, 3157]. Photoperiod [1859, 1609, 2265, 1485, 3235, 187, 3274, 2663, 1066]. photoperiodic [3419]. Photophilic [2648]. photophores [1413]. Photopigment [2072]. photoreceptor [549]. photorepair [1381]. Photoresponse [2823]. Phoxinus [1872, 3190, 2754]. phthalate [785]. Phyllopteryx [1788]. Phylogenetic [788, 2857, 2130, 840, 886, 292, 1358, 762, 3185, 410, 3437, 1141, 2619, 3430]. Phylogenetics [160, 378]. Phylogeny [2853, 1035, 2988, 3428, 2361, 3221, 2150, 2981, 166, 127, 2520, 586, 2146]. Phylogeographic [1470, 3437, 956, 3432, 2325]. Phylogeographical [3289]. phylogeographically [712]. Phylogeography [3290, 2326, 2131, 3044, 3291, 3341, 1286, 2521, 3288, 2708, 3436]. Physical [1403, 1901, 299, 3480, 268, 1732, 2550, 3102, 1861, 2769, 1736, 2856]. Physical-biological [1901]. physicochemical [2396]. physics [835]. Physiological [1485, 578, 610, 1610, 2801, 3184, 43, 767, 1124, 1012, 438, 1511, 2940, 442, 1889, 2996, 286, 372, 1077, 2126, 3111, 3103, 781, 2047, 3156]. physiologically [1169]. physiology [2812, 2466, 1377, 439, 1741, 2965, 3330, 3054, 1305, 835, 3406]. phytoestrogens [2138]. phytoplankton [170]. Phytotechnology [1653]. Piaractus [2698]. picking [2039]. piecewise [3076]. Pierre [1901]. Pigment [1812]. Pigmentation [87, 1852, 3035, 1856, 1189, 1063, 2050]. pigments [1157, 871, 2840]. Pike [2363, 1339, 1607, 189, 1443, 1419, 1191, 2973, 2936, 2903, 2143, 1070, 178, 3253, 2490, 501, 948, 1313, 2375, 1331, 1454]. pikeminnow [2097, 376]. Pikeperch [2954, 1897, 2723, 1607, 1071, 1150, 1516, 239, 305, 432, 328, 655, 1454, 3162]. pilchard [1417]. pilchardus [2705]. pilings [2776]. Pimelodidae [1489].

Pimephales [788, 3072, 3059, 1499]. Pimozide [118]. pinetorum [3136]. pinfish [427, 2614]. Ping [3272]. Pinguipes [2870, 3176]. pink [871, 653, 2403, 3384, 1764]. pioneers [1931]. pipefish [2044, 2970, 3151, 915, 986, 2154, 2301, 2595]. pipefishes [888, 1216]. piranha [2638, 3401]. piscatorius [1502, 2071]. Pisces [3036, 2439, 443, 2145, 1684, 2725, 1203, 1810, 2451, 2912, 44, 1898, 2154, 2329, 452, 2697, 2285, 2301, 1714, 2141]. piscine [1638]. piscivore [1407, 2954, 1865]. piscivores [1203]. piscivorous [2563, 3270, 2032, 3397, 1445]. Piscivory [963, 3241, 1102, 3260, 2160, 2260, 591]. pit [497, 988, 639, 307, 3108, 1635, 2172, 1878]. PIT-tagged [2172]. PIT-tagging [1878]. pitfalls [1563]. pituitary [2145]. placement [2532]. placentatrophic [3445]. Places [3405, 1107]. plagiosum [2864, 1853]. plague [3021]. plaice [346, 445, 1474, 796, 104, 126, 867, 1186, 1210, 2767, 1220, 959, 48]. plainfin [3090]. plains [3167, 878]. planktivore [2954]. planktivorous [2054]. Plankton [2039, 538]. Plankton-picking [2039]. planktonic [1380]. plant [3196, 1357, 2593, 2209]. planting [484]. plants [1463]. Plasma [821, 466, 1044, 1836, 437, 2214, 1276, 1547, 168, 1128, 649, 1507, 2413, 489, 2677, 874, 797, 1710, 316, 221, 1066, 995, 1791, 451]. plasticity [2469, 861, 3183, 3396, 254, 2208]. Plata [3177, 3287]. platana [2882]. plate [3418, 1981]. Plateau [2650, 2520]. platei [776]. platessa [2767]. Platichthys [1672, 2608, 2821, 2484, 2476, 1099]. platy [562]. platycephala [2921]. platycephalus [1358, 2689]. Platysalmo [1358]. Plecoglossus [1754, 1914, 2531]. Plectroglyphidodon [2189]. plectropomi [1979]. plectropomid [1052]. Plectropomus [453, 3315, 2052, 2727, 1979]. Plectvopomus [762]. Pleistocene [3044]. plerocercoid [2558]. plerocercoids [2506, 138]. pleurocidin [2542]. Pleuronectes [2086, 1460, 2767, 2507]. Pleuronectida [3136]. Pleuronectidae [2608]. Pleuronectiformes [2516, 2846]. Pleuroscopus [2650]. ploidy [1667]. Plotosidae [458]. Plotosus [458]. plunderfish [210]. pluripotent [1112]. plutodus [1396]. pocket [967]. pockets [73]. Poecilia [1241, 3351, 3213, 3434, 3456, 2692, 2099, 2006, 2153, 3256, 1123, 1256, 3354, 2648, 3363, 1475, 1862, 2804, 1267]. poeciliid [2678, 1027]. Poeciliidae [3213, 2692, 2648, 2804]. Poey [2664]. poison [1463]. poissons [1405]. Poland [1706, 2601, 1606, 1615, 1616, 1316]. polar [284]. polarization [1696, 2385]. Poll [2132]. Pollachius [1658]. pollack [1620]. pollan [616]. pollock [729, 2610, 2687, 3278, 2422, 854, 90, 2, 492]. pollutants [1528]. polluted [768, 3052]. pollution [272, 3177, 2591]. polyacanthus [1281, 719]. polychlorinated [2873]. polychromatic [3222]. Polyculture [1580]. polygynous [2402]. polymerase [32]. polymorpha [3012]. polymorphic [668, 2085, 3454]. Polymorphism [1614, 333, 2323, 960, 2807, 1969, 1754, 1539, 2321, 3101, 2862, 1941]. polymorphisms [107]. Polynesia [3065]. polyommata [2766]. polyploid [2735, 2915]. Polyprion [1398]. Polyprionidae [1398]. polyunsaturated [263]. pomacentrid [2594]. Pomacentridae [2796, 794, 2166, 2189]. Pomacentrus [2046, 2594]. Pomatomus [1407]. Pomatoschistus [808, 3402, 254, 167]. pomfret [2088]. Pomphorhynchus [449]. pond [2084, 1606, 3252]. ponds [862, 2004, 948, 1313]. Pontificia [1495]. Pontual [1775]. pool [3472, 252, 2835]. pooled [545]. Pools [77, 299, 396]. poor [425, 41]. pop [3194]. pop-up [3194]. Population [3366, 1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 3151, 1456, 3422, 101, 653, 3286, 1592, 1079, 3117, 1600, 1929, 2531, 3161, 966, 2925, 1292, 1945, 2800, 1946, 2207, 1577, 3036, 1891, 1281, 855, 1983, 268, 3437, 1559, 2319, 2481, 3220, 2270, 1489, 2447, 3379, 2367, 1206, 1517, 2554, 3360, 3171, 2079, 2647, 575, 3045, 2013, 1637, 751, 2749, 2978, 202, 1867, 520, 3066, 613, 1614, 1976, 2348, 1754, 756, 3248, 3467, 2223, 622, 236, 1252, 1591, 3167, 68, 1075, 41, 3280, 2337, 612, 460, 2286, 732, 90, 1433, 2335, 731, 1065, 2429, 3365, 1819]. population [2578, 2500, 2501, 1367, 35, 2144, 1543, 2299, 1170, 2210, 276, 2269, 1374, 488, 2545, 2854, 2322, 2529, 1011, 1939, 1078, 2738, 3239, 2969, 670, 1269, 1270, 3040, 2990, 3417, 3438]. populations [1722, 1937, 2693, 900, 2678, 1579, 3084, 759, 2710, 2760, 1923, 3170, 2125, 1933, 606, 2419, 862, 2536, 956, 796, 3053, 2353, 2745, 2523, 2034, 174, 2115, 2692, 2793, 913, 304, 914, 2978, 98, 3399, 1288, 1845, 646, 2951, 3223, 2736, 59, 2659, 3462, 3028, 1759, 3041, 2004, 3421, 2699, 2817, 1719, 599, 617, 1940, 1938, 1755, 1080, 1123, 1399, 3121, 1370, 2329, 1425, 1710, 1595, 2187, 2068, 1796, 1122, 1996, 1691, 29, 1305, 3459, 2943, 1808, 2442, 3129, 2530, 2792, 2343, 1575, 1057, 1862, 1161, 212, 1602, 2738, 52, 903, 49, 2989]. Porcupine [667, 241]. Porichthys [3090]. porphyropsin [2840]. Port [497, 2808, 2842]. portable [405]. portal [2572]. Porter [3349]. Porto [1495]. Portugal [1314, 262, 295, 2096, 902, 1667, 1835, 918, 3063, 175, 421]. Portuguese [2535, 2339, 1397, 2510, 2904, 1092]. portusjacksoni [2645, 2808, 2842]. position [1200, 3428, 2929, 3241, 2159, 1358, 2874]. positional [991]. positions [588, 1260, 991]. positively [2944]. possessing [484]. possible [549, 3271, 3471, 1647, 3429, 2330, 429, 803, 401, 1753, 2837, 2415, 2293, 17, 2221, 769]. Post [3299, 3139, 1046, 11, 948, 3455, 424, 484, 1840, 3044, 2502, 3093, 2387, 1825, 2974, 3389, 189, 1561, 1634, 1890, 3463, 196, 223, 2682, 3202, 1063, 710, 2412, 2946, 3025, 1690, 1669, 985, 828, 3134, 1380, 1492, 416, 2229, 1421, 1820, 287, 879, 1124, 2683]. post-emergent [1825]. post-exercise [985, 2683]. post-exogenous [3093]. post-flexion [2387]. post-glacial [3044, 828]. post-hatching [196]. post-juvenile [223]. post-larvae [710, 1380]. post-larval [1063, 1820]. post-mining [189]. post-mortem [1492]. post-planting [484]. Post-release [3139, 11, 3455, 1561, 3463, 3025]. Post-settlement [3299, 2412, 2229]. post-smolt [2682, 1669, 3134, 416, 287]. post-smolts [424, 2974, 3389, 1634, 3202, 1690, 1421, 1124]. post-spawning [1840]. Post-stocking [948, 2946]. post-weaned [2502]. Poster [285]. posterior [1027]. Posters [925, 1960, 2350]. postglacial [2547]. Posthodiplostomum [741]. postlarval [2086, 3275]. postovulatory [2617].

Postprandial [3381, 1798, 2944]. postsmolts [829]. postspawning [2225]. posture [2010]. pot [2474]. pot-bellied [2474]. potassium [2365]. potency [1056, 3027]. Potential [1847, 1697, 2825, 2991, 1920, 855, 3149, 963, 2273, 2037, 1427, 3421, 828, 2664, 2895, 141, 2183, 2628, 1709, 2288, 1364, 1713, 3257]. potentially [1605]. pouch [888]. pouches [888]. pout [2826]. power [914, 1357, 820, 2209, 3405]. pp [1403, 1496, 1497, 1406, 1654, 1402, 1772, 1651, 1404, 1773, 1774, 1493, 1775, 1405, 1652, 1495, 1494, 1653]. practical [3149]. Practices [1901]. praematura [787, 1661]. pragmatic [3444]. prawn [2766]. Pre [3191, 1143, 3314, 839, 301, 1890, 3139, 2844, 3362]. pre- [1890]. pre-alpine [2844]. pre-larvae [3362]. Pre-migratory [1143, 839]. pre-settlement [3314]. pre-spawning [301, 3139]. Pre-winter [3191]. precedence [2379]. precision [827, 457]. precursors [150]. Predation [829, 2450, 548, 978, 1389, 3351, 3242, 1607, 3190, 1517, 2550, 67, 2295, 3105, 1118, 704, 2896, 1327, 2046, 2776, 531, 2808, 2442, 1884, 1709, 3455, 1130, 3012]. Predator [1172, 1241, 3424, 1562, 553, 776, 2448, 758, 532, 718, 1014, 1205, 1443, 1197, 3317, 126, 1208, 1102, 1330, 1734, 938, 2421, 3308, 1816, 3453, 2418, 1018, 2442, 2183, 566, 385, 1234, 2831]. predator-caught [2831]. predator-dominated [3308]. Predator-induced [1562]. Predator-inspection [3424]. predators [2952, 1203, 2363, 290, 2375]. predatory [545, 1894, 1893, 2426, 549, 462, 2903, 938]. predetermined [2383]. predict [1856, 373, 494, 3368, 2973, 967, 2628]. Predicting [2263]. Prediction [1122, 2517]. predictions [227, 1568]. predictive [587]. predictor [317, 1173]. predictors [1222]. predicts [2248]. predisposes [2661]. preface [1405]. prefer [3363, 838]. preference [2970, 2098, 684, 587, 791, 3247, 1673, 2881, 1786, 2395, 1265, 991, 2768]. preferences [3318, 2053, 1297, 1517, 111, 1391, 3476, 2267, 807, 1849, 1235, 1267, 2569, 2080]. preferenda [910]. preferential [2583]. pregnant [2948]. pregnen [678]. prejuvenile [2151]. Preliminary [3404, 453, 1323, 1070, 105, 1622]. Preparing [1679, 898]. prescription [3376]. Presence [883, 944, 2646, 2535, 373, 2330, 3296]. present [368]. presentation [79]. preservation [2014, 1356, 2281, 1046, 2050]. preserved [1116]. Prespawning [2157, 2368]. Press [1403, 1772, 1493, 1405, 1494]. pressure [2751, 2073, 1159, 2442, 2628, 1173]. pressures [911, 748]. prevalence [2834]. prevent [3020]. prevents [2437]. Previously [2240, 2617, 1282, 2067]. Prey [3304, 2504, 2032, 2452, 2046, 3283, 2563, 424, 2011, 216, 2876, 2608, 1241, 1192, 1108, 3270, 3314, 758, 1147, 682, 1014, 1205, 1185, 2295, 867, 1186, 1952, 342, 1971, 1102, 2700, 1516, 3397, 518, 2607, 1176, 2160, 2422, 2403, 938, 178, 2421, 1816, 2954, 2505, 3453, 1193, 213, 1440, 1018, 1865, 1736, 2354, 948, 2375, 386, 770, 3009, 1638, 1454, 1806, 2831, 2417]. prey-selection [424]. preying [376]. Pribilof [729]. Price [1403, 1496, 1497, 1406, 1654, 1402, 1772, 1651, 1404, 1773, 1774, 1493, 1775, 1405, 1652, 1495, 1494]. prickleback [1488]. primarily [2030]. Primary [2312, 263, 1513, 2508, 1505, 2274].

Primed [1049]. primers [641]. priming [676, 368]. primitive [1716]. Prince [1811, 2118]. Princeton [1403]. Principles [1954]. Printer [567]. Prionace [2304]. prior [1192]. Pristiformes [1745]. Pristis [2819]. probability [2045]. Probable [2547, 2910, 3263, 678]. Probing [3116]. problems [1847, 1697, 1582]. proboscirostris [3298]. procedure [3420, 423]. procedures [2489]. process [1715, 3025, 2720]. Processes [2601, 2337, 11, 2236]. prochilodontid [1286]. Prochilodontidae [2493]. Prochilodus [2493, 1917, 3177]. produced [672, 3293]. producing [2145, 2242, 2802]. production [1888, 975, 1439, 2306, 263, 1947, 1580, 912, 1546, 665, 1044, 2899, 1062, 2259, 1508, 2189, 383, 2500, 2873, 2221, 3420]. productivity [2061, 2788]. products [2798]. productus [3307]. profile [2362, 2445, 2677]. profiles [466, 2083, 1128, 309, 3384, 221, 437]. progenetic [1818]. progenitor [752]. progeny [983]. Programme [1653]. programmes [3231, 2294, 1601]. progress [1063, 922]. progressive [260, 3383]. project [1613]. proliferation [2358, 2245, 2593, 60, 810]. prolonged [2026, 1842, 1617, 3018, 463, 231]. promelas [3072, 3059, 1499]. Promiscuous [1986]. promotes [3155]. proof [3399]. propagation [1563]. Propagule [2751]. Propanil [1068]. propensity [1711]. properties [2655]. propogation [1611]. proposal [1835]. prospects [1582, 923]. prostaglandin [263]. prostaglandins [941, 676]. protandry [2508]. proteases [1128]. protected [1770, 1308, 2338]. protection [606, 3071]. protective [2917]. protein [1023, 3149, 2555, 2942, 2728, 3131, 1483, 53, 745, 469, 843, 1791, 1623, 1929]. protein-1 [2728]. protein-concentrate [3149]. proteins [987, 2593, 150, 1404]. Proteolytic [1128]. protistan [175]. protocols [377, 940]. protogynous [453, 1409, 1045, 1505, 2161, 2661]. protogyny [418, 1722, 947, 2508]. proton [325]. Protopterus [2324, 3330, 2938]. Prototroctes [2261]. prove [1219]. provide [1183]. provided [753]. Province [3294]. Proximate [2097, 1245, 1382, 3068, 1151, 3041, 1717]. Prussian [2915, 1572]. Przewalski [3438]. przewalskii [3438]. Psammobatis [786, 1468]. Psetta [2821, 2920, 1657, 1265, 3319]. Pseudochondrostoma [3185]. Pseudocrenilabrus [783]. pseudodorsalis [2650]. Pseudogobius [2725]. pseudoharengus [3075]. Pseudophyllidea [138]. Pseudoplatystoma [1489]. Pseudopleuronectes [2443, 1609, 1561, 2869, 3200]. Pseudorasbora [1067, 1605, 2757, 851, 1538, 2760, 789, 2278, 2221, 2746, 2485, 3099]. Pseudosciaena [1864, 2982]. Pseudotropheus [2872]. Pseudotvopheus [684]. Pseudupeneus [2039, 2220]. Pterocryptis [491]. Pteroplatytrygon [2624]. Pterygoplichthys [2600, 3010]. Puberty [1507]. public [3203]. Published [1403, 1496, 1497, 1406, 1654, 1402, 1772, 1651, 1404, 1773, 1774, 1493, 1775, 1405, 1652, 1495, 1494, 1653, 924, 925, 1959, 1960, 2349, 2350]. Publishing [1654, 1404]. puffadder [1336]. puffer [2646]. pulchella [2943]. pulcher [1511, 3266, 1432, 3173]. pulmonary [379]. pulse [1436]. pulse-echo [1436]. pump [325]. pumping [2456]. pumpkinseed [2339, 900, 2179, 2745, 139, 1020, 1154, 2560]. pumpkinseeds [1718]. punctata [710, 1915, 3037]. punctatum [2631]. punctatus [2082, 3262]. Pungitius [2926, 2322]. Pungtungia [3303]. Puntius [74]. pupfish [3348]. pure [1794]. Purification [1279, 348]. pusillus [1141]. putative [941, 3203, 797, 2498, 3454]. puye [2419]. pygmy [3451, 3440, 2671]. Pygocentrus [2638]. pyloric [2909, 17]. pyrenaicus [2148]. pyrenaicus-like [2148].

Qinghai [2520]. quality [1579, 457, 1423, 2222, 2369, 1062, 1594, 1227, 2422, 1221, 1105, 3376, 258, 313]. quantification [2739]. Quantitative [2997, 2568, 3002, 2792, 972]. Quattro [3479]. Quebec [641, 1333, 2637, 1748, 1642]. queen [3263]. Queensland [2098, 1684, 2147, 3182, 3447, 3276]. queketti [2250]. quelen [821, 2356]. question [335]. quiescent [105].

R [3479, 1406, 3410, 3406, 1495]. R$ [1495]. rabbitfish [2098, 221, 964]. rabbitfishes [434]. racer [2731]. races [2326]. Rachycentron [2400, 2684]. radiata [51, 2010, 2268]. radiation [3168, 2446, 378]. Radio [185, 2374, 1003, 1041, 1293, 1071, 1514, 432, 3111, 981, 2172, 2556]. radio-tagged [1514, 2172]. radio-tagging [432]. radio-telemetry [2556]. Radio-transmitted [185]. radioactive [1331]. radiocarbon [1116, 1366]. radiotelemetry [498]. Radulinopsis [1792]. Rafinesque [2896, 2685]. Rahel [1651]. rain [1353]. Rainbow [338, 438, 152, 672, 2203, 3130, 554, 1641, 2215, 2214, 1418, 317, 3302, 1349, 1555, 1681, 884, 2557, 2347, 33, 805, 1525, 1814, 2818, 2743, 191, 249, 2908, 2592, 635, 339, 2165, 2942, 2623, 3431, 3024, 1529, 2164, 1118, 1002, 527, 972, 2700, 1668, 370, 3327, 3139, 2945, 1483, 1999, 2881, 2266, 2116, 1109, 526, 88, 489, 1420, 1740, 2625, 1049, 3441, 2312, 1369, 474, 2001, 1089, 56, 2429, 1373, 1441, 2559, 130, 2794, 271, 1066, 2501, 463, 598, 804, 472, 1386, 2038, 1362, 3393, 1466, 60, 454, 1004, 1981, 810, 368, 448, 2028, 570]. rainbow [2989, 3408]. rainbowfish [174, 375, 3021, 3247]. rainfed [1945]. rainforest [2856]. raised [640]. Raja [2331]. Rajidae [2408, 2766, 2585, 3280, 2882, 2649, 2331]. raker [2054, 638]. ramps [2204]. ranched [1588, 985, 1103]. Randall [3375]. randalli [2949]. Randallichthys [3375]. range [1398, 3044, 604, 245, 3456, 1683, 2371, 1639, 2973, 3180, 1391, 2541, 2372, 471, 1064, 3239, 3099, 599]. ranges [3185, 498, 3025, 2266, 1258]. ranging [3000, 588, 1182]. rank [1242]. Rankin [294]. ranks [1512]. Rannoch [668]. Ranzania [642]. RAPD [582, 1692]. rapeseed [3370]. Rapid [3416, 2413, 3277, 2661, 2211, 2874, 2717, 2558, 2321, 2835, 378]. Rare [3118, 3334, 2960, 157]. rarefaction [144]. Rastrineobola [3060]. Rate [2687, 152, 954, 761, 19, 1384, 2346, 2215, 210, 317, 1869, 3127, 2667, 2487, 3368, 3274, 894, 1147, 692, 387, 1798, 2944, 2242, 1126, 2165, 2848, 969, 1211, 2845, 340, 3248, 2700, 2659, 3378, 3203, 3155, 717, 154, 1082, 2067, 178, 140, 48, 1570, 577, 36, 130, 1791, 3129, 1884, 3112, 892, 493, 1480, 841]. rates [392, 2876, 3389, 1998, 2616, 2684, 301, 2235, 194, 2406, 1527, 2037, 2032, 1514, 1357, 885, 2046, 3005, 3138, 1361, 1708, 2187, 288, 2429, 1444, 2182, 2819, 2880, 237]. rather [894]. ratio [42, 2942, 2058, 61, 2867, 627, 1222, 2833, 80, 647, 1870, 2618, 1385]. ration [2216, 330, 2559, 1865, 1338, 3306]. rations [545, 1894, 194, 698, 178]. ratios [3315, 795, 2096, 519, 646, 2320, 3308, 2210, 294, 3074, 754, 2064]. Rauchenberger [2782]. ray [3446, 2150]. ray-finned [2150]. rays [2641, 2503]. razor [1409]. razorback [2946]. rDNA [2130, 3296, 1049]. Re [149, 1578, 2660, 2965, 2665, 1958]. re-circulating [1578, 2965]. re-established [2660]. Re-examination [149]. re-feeding [2665]. re-watering [1958]. reach [3272]. reaction [32]. Reactions [2209, 2374, 1734]. read [1219]. reading [898]. readings [541]. ready [839]. ready-to-migrate [839]. real [2739]. real-time [2739]. realized [2037]. Reared [1594, 1172, 1577, 583, 87, 478, 844, 1192, 1108, 1559, 2387, 2812, 1606, 979, 360, 2561, 784, 372, 93, 1634, 1610, 1611, 864, 519, 2327, 1208, 2845, 646, 1997, 1966, 1434, 3025, 66, 58, 3247, 1565, 716, 2428, 11, 2923, 1569, 634, 1570, 870, 1596, 1265, 1535, 1161, 3455, 1747, 1480]. Rearing [652, 1607, 3368, 3231, 715, 1350, 1587, 1564, 1829, 2559, 2651, 1884]. rearrangement [2921]. rearrangements [3185]. reasons [1572]. recapture [966, 1998, 3365, 1721, 2685, 2930]. recaptured [1847, 619, 2045]. recaptures [58]. Received [38, 134, 408, 876, 905, 1095, 1302, 1776, 1876, 1927, 2018, 2075, 2176, 2253, 2308, 2435, 2576, 2653, 2786]. recently [710, 2524, 212]. receptor [339, 1212, 1744]. receptors [2979, 1441]. reciprocal [2476, 1340]. recognition [2991, 3344, 1251, 510, 3059, 1252, 1464, 3425, 2442, 2651]. recognize [2460]. recolonization [923]. Recommended [1901]. Reconstructing [627, 2238]. reconstruction [3075, 213, 2544]. record [3294, 2597, 3161, 1727, 1777, 1398, 2009, 2911, 2247, 851, 2599, 662, 1397, 1750, 2171, 2886, 2250, 180, 449, 3215, 1674, 2850, 2647, 129, 2525, 2822, 787, 2914, 3375, 266, 452, 306, 3063, 203, 1092, 3214, 1396]. recorded [2287, 1768]. recorder [769]. Recording [1869, 152, 2457]. recordings [1556]. Records [2851, 1628, 1521, 664, 67, 2620, 902, 2482, 464]. Recovery [1320, 1804, 2474, 2913, 1559, 1524, 613, 1363, 102, 1973, 1500]. recreational [3444, 1663]. recrudescence [214, 1507, 451]. recruit [64]. recruited [2690]. Recruitment [3236, 3234, 882, 1417, 3233, 3360, 46, 74, 3399, 1135, 70, 3211, 3238, 1898, 1055, 1553, 3253, 3079, 2574, 1274, 3232, 3313]. recruits [848, 3237]. rectal [761, 984]. red [3073, 947, 1112, 535, 338, 2638, 1867, 2022, 499, 3309, 2418, 2144, 2906, 1900, 3109, 294, 3319, 1011, 2886, 2885, 3373, 889]. red-bellied [2638]. red-throat [947]. red-transparent [1867]. redband [1305]. redbelly [3190]. redbreast [3205]. redd [2080]. redds [2499, 73, 310]. Redescription [1802]. redfin [2204, 1374]. redfish [518, 3078, 697]. redfishes [859]. Rediscovery [2458]. Redistribution [3185, 2491]. redness [961]. redside [3335]. redspotted [402]. reduce [1699, 2951, 1668, 1357, 3350, 1602]. Reduced [2759, 1002, 3100, 846, 631, 470, 2195, 2767, 2257, 1377, 2573, 991]. reduces [2197]. reduction [1320, 1700, 3418]. Reef [2052, 1494, 762, 699, 222, 418, 2065, 3314, 758, 2272, 1194, 1175, 2275, 92, 1422, 3360, 621, 168, 3399, 1214, 1107, 1520, 2397, 265, 2656, 2046, 2395, 3005, 3308, 3022, 1258, 3117, 364, 2852, 1229, 1274, 2701, 1431, 2231, 2629, 561, 2930, 947, 3050, 1785, 1699, 65, 1011]. reef-fish [1107]. reefs [1712, 1947, 2233, 3452, 1515]. refeeding [31, 2533]. Referees [1303, 1626, 1655, 1942, 1961, 1988, 2351, 2380, 2761, 2787, 3113]. reference [171, 607, 1937, 373, 3443, 1753, 2581, 971]. refinement [3420]. reflect [295, 1193]. reflectivity [1491]. reflects [1786, 1739]. refractive [2488]. Refuge [2234, 85, 412, 2932, 652, 2229]. Refuge-use [2234]. refuges [2835]. refugia [77, 3432, 176]. regalis [3312]. regarding [3476]. regeneration [2051]. regime [217, 24, 297, 784, 636, 3230, 3249, 1743, 3418, 878]. regimes [153, 1485]. region [607, 2713, 2975, 712, 1141, 2610, 956, 382, 1860, 2724, 465, 3174, 751, 3291, 2921, 3167, 2493, 2140, 2495, 458, 369, 3037, 3014, 3449]. Regional [729, 2217, 361, 2001, 1932, 1146, 3119]. regions [1115, 3106, 2862, 2544]. regius [2287]. regression [3076, 627, 1896]. regressive [1781]. Regular [569, 479]. regularis [1758, 2407]. regulated [2777, 2496, 3249, 2767, 3211, 1730]. regulating [3055]. Regulation [1063, 97, 535, 3160, 301, 425, 843, 476]. regulator [2293]. regulatory [3034, 1941]. regurgitation [566]. rehabilitating [2319]. Reinhardtius [1886]. reintroduced [2536]. reintroductions [1198]. Reis [1495]. related [315, 2328, 162, 746, 1533, 1821, 3073, 1842, 687, 1645, 642, 1488, 729, 1948, 3333, 610, 100, 2638, 2394, 2944, 2832, 1419, 2459, 2314, 1818, 164, 25, 3299, 1590, 1783, 1109, 899, 526, 1898, 1938, 574, 1088, 775, 2562, 1820, 2955, 2442, 45, 1676, 318]. relating [2778]. relation [821, 1038, 1633, 539, 1423, 187, 2447, 3198, 682, 3429, 1189, 1646, 661, 3179, 3249, 2361, 2371, 969, 242, 1971, 1150, 2008, 518, 2634, 1218, 655, 3013, 3026, 2370, 154, 3349, 1042, 1257, 351, 2354, 1313, 2162, 1028, 1266, 1883, 940]. relations [429, 141]. Relationship [302, 349, 103, 2845, 1743, 2441, 2429, 1450, 2955, 201, 19, 1339, 1848, 2675, 549, 1242, 1546, 387, 290, 1683, 1209, 1460, 1081, 3221, 508, 1075, 717, 2677, 1717, 2662, 1703, 3008, 9, 2415, 870, 2628, 3135, 2820, 512, 493, 3346]. Relationships [2532, 893, 1126, 1663, 363, 1096, 804, 788, 2857, 2130, 840, 289, 1141, 1014, 1205, 805, 2658, 64, 1949, 2633, 886, 3068, 2007, 1040, 1067, 3125, 734, 3150, 1784, 2416, 2417, 2731, 292, 479, 1022, 1470, 1289, 428]. Relative [1056, 1748, 2723, 2192, 99, 2974, 177, 2745, 2294, 5, 2659, 154, 675, 1316, 2880, 1307]. Release [1273, 1814, 1561, 2908, 1056, 3463, 3025, 3139, 649, 11, 1315, 1780, 1973, 1576, 3455]. released [1501, 1647, 784, 1525, 66, 58, 3103, 1570, 1596, 1264, 1535, 2064]. releaser [570]. releases [1602]. releasing [1559, 649, 2313]. relevance [1146]. relevant [791]. reliable [1766]. reloaded [3425]. remains [1816]. remarkable [1980]. remarks [2639]. Remediation [3480]. Remodelling [1283, 3393]. remote [498, 2971]. remotely [1502]. removal [1164, 1834]. removing [1697]. Renal [850, 1148, 2211]. renewal [1737]. Renibacterium [88]. renin [1155]. renin-angiotensin [1155]. reoligotrophication [2558]. repaid [2683]. repeat [1452]. repeatability [154]. Repeated [1967, 1710, 2794]. repeats [1604]. Repetitive [1296]. replacement [2593]. reply [792]. report [733]. represent [2449]. reproducing [2156]. Reproduction [3010, 359, 3182, 920, 2836, 2948, 146, 915, 3446, 2766, 2033, 1309, 1789, 1047, 1079, 2470, 345, 654, 574, 1088, 47, 2386, 1881]. Reproductive [2800, 935, 415, 786, 1716, 1168, 1409, 1412, 2317, 1682, 1975, 2904, 1585, 630, 1091, 1187, 3152, 3398, 2694, 1139, 2105, 1670, 2964, 2366, 2671, 1218, 986, 1468, 3421, 592, 3280, 2154, 1055, 2258, 1042, 2612, 2953, 1351, 1337, 1272, 2871, 1788, 461, 3079, 2301, 1021, 2389, 1469, 2461, 459, 97, 1414, 821, 2227, 1501, 632, 952, 888, 1583, 3017, 3060, 100, 1689, 1982, 1248, 1731, 125, 3457, 1547, 2191, 694, 2657, 2314, 481, 2297, 1252, 1394, 703, 2150, 2370, 2656, 633, 676, 2666, 2289, 705, 1177, 1620, 3309, 1741, 254, 1365, 778, 3, 2501]. reproductive [3147, 167, 1026, 3090, 1325, 2268, 201, 3332, 669, 1078, 903, 2672, 2503, 3306, 3200, 3303, 3257, 3356]. Republic [2861, 1572, 2556, 3214]. requiem [3003]. require [2241]. rerio [3120, 2980, 2892, 3218, 3033, 3217, 3378, 2193, 1887, 2208, 2155, 1269]. Research [2077, 1964, 906, 911, 3339, 3310, 1929, 3217, 3323]. Reserve [2079, 606, 1855, 669, 605]. reserves [604, 1783, 1640]. reservoir [3270, 2445, 686, 1614, 981, 2894, 878, 1548, 3204, 802]. reservoirs [2339, 2362, 1043, 2560]. resettlement [395]. Residence [2641, 2567]. residency [2192, 3279]. resident [3130, 2667, 2383, 1689, 1715, 575, 865, 1462, 1143, 2040, 321, 2454]. Resistance [186, 199, 174, 1113, 2257]. resistant [1555]. Resolution [1518, 3075]. resolved [3428, 3138]. resonance [3110, 3319]. Resource [686, 2194, 3364, 2925, 2932, 911, 2263, 2846, 324, 1969, 2837, 1222, 2159, 3281]. resources [2504, 3371, 522, 989, 2469, 705, 2167]. respect [480, 2068]. Respiration [301, 927, 351, 772]. Respiratory [783, 176, 930]. respirometer [931, 84, 2111]. respirometer-confined [2111]. resplendens [3451]. respond [1159]. responding [2593]. Response [2387, 1718, 438, 3094, 761, 1418, 3235, 2430, 2234, 1798, 538, 1350, 680, 1113, 2126, 2767, 3131, 1330, 1275, 655, 2200, 2817, 3103, 489, 2403, 2111, 56, 3054, 288, 3283, 983, 2442, 3178, 1346, 2673, 1024, 1389, 2584, 2982, 1556, 2568, 3419, 940]. Responses [3325, 1367, 3480, 1913, 2773, 1281, 1501, 1485, 404, 3302, 1645, 3422, 3072, 1658, 284, 442, 2020, 1244, 3198, 189, 77, 636, 1826, 3456, 3398, 3317, 2828, 1208, 1587, 1363, 1077, 2511, 370, 941, 1836, 2409, 499, 1369, 577, 2838, 1089, 2614, 43, 2938, 1884, 1574, 1981, 2957, 1853, 1130, 2155, 2028, 377]. responsibility [615]. Responsible [1801, 2760]. responsiveness [99, 2038]. rest [152, 1643]. resting [1798, 2944, 493]. restocking [2319, 1613]. Restoration [251, 1144]. restored [1415]. Restoring [3399]. restricted [498, 126, 1148]. restrictions [716]. result [2555, 503, 2260, 2187, 1393]. resulting [2899]. results [1194, 2582, 202, 1594, 2855]. resumption [743]. retain [3457]. retardation [1355]. Retention [2012, 1418, 1071, 1354, 3156]. reticulata [1241, 3351, 3434, 3456, 3256, 1256, 3354, 3363, 1267]. reticulatus [1722, 3215]. retinae [549]. retinal [1157, 871]. retinomotor [2584]. Retracted [2916]. Retrograde [17]. Retropinna [2957]. Retrospective [2841, 1144]. Return [1376, 516, 2067]. returning [3186]. Reunion [2152, 1093]. reveal [91, 2147, 3142, 2806, 3439]. revealed [3290, 1131, 2807, 2527, 1644, 3008, 3079, 2539, 1834, 3074]. reveals [960, 2138, 2582, 2719, 2143, 2984, 3266, 488, 697, 2797, 903]. reversal [3235, 309, 2320, 647]. Reversals [3136]. reversed [2554, 773, 2154]. Review [3479, 1403, 1010, 1496, 1406, 1654, 2928, 3410, 1555, 1402, 1772, 3323, 1651, 1404, 3322, 1773, 1774, 1493, 1775, 3478, 3405, 3321, 3311, 236, 1665, 1652, 3406, 1495, 1494, 3407, 3480, 3408, 1653, 3320, 3411, 3409, 457, 3443, 2763, 491, 1036, 3057, 2748, 2752, 922, 1465, 1406]. Reviews [2017, 21, 37, 108, 133, 183, 224, 269, 331, 407, 435, 455, 504, 643, 713, 766, 875, 904, 1094, 1497, 2175, 2252, 2307, 2434, 2575, 2652, 2785, 2976, 3032, 3143, 3163, 1875, 1301, 1924, 1298, 1134, 1926, 1925, 1299, 1405, 1300, 2074]. Revillagigedo [2910]. revised [1772, 334]. Revision [1039, 314]. revisited [3428, 627]. Reykjanes [3078]. Reynolds [1652]. RFLP [2131, 2724, 2140, 3347]. RFLPs [2718]. Rhône [637]. Rhabdoblennius [2768]. Rhadinocentrus [3021]. Rhamdia [821, 2356]. rheotactic [2600, 965]. rheotaxis [497]. Rhincodon [3141]. Rhine [3243, 1121, 860, 2070, 529, 310]. Rhinogobius [1976, 3074]. Rhinoptera [2641]. Rhizoprionodon [2800, 3465, 2105, 2016, 837]. Rhodes [852]. Rhodeus [2168, 2008, 3211]. rhodopsin [2840]. rhombeus [1831]. rhomboides [2614]. rhythm [1617]. Rhythmic [1418, 964]. rhythmicity [2479]. rhythms [2391, 3331, 3269]. Ria [918]. ribbonfish [2139, 1749]. ribosomal [292, 32]. Rican [1353]. rice [1945, 3149]. rich [1672, 2611, 2701]. Richardson [2670, 2669, 2982]. richness [601, 144, 1683]. Ricker [645]. Ridge [3078, 2547]. rifting [728]. right [718, 2959]. rigor [1492]. Rim [2990]. Rineloricaria [3015]. ringens [2478]. rings [2031]. Rio [1495, 3177, 1446, 425, 2650, 1326, 3454, 3287]. Riparian [890, 373, 558, 2551]. Risk [3391, 3190, 2743, 3059, 1118, 1218, 1327, 531, 1367, 1884]. Risk-taking [3391]. risks [2751]. Risso [2391, 2554]. River [3234, 1748, 1137, 2122, 234, 12, 1667, 2472, 2420, 3401, 2314, 1642, 2759, 1135, 3236, 3043, 2094, 3211, 2528, 2968, 1462, 1316, 2169, 3184, 1996, 2501, 1572, 1846, 201, 2322, 237, 3214, 1802, 2095, 1050, 670, 2913, 2080, 1195, 1320, 1085, 1908, 2956, 1415, 2777, 1456, 3255, 890, 1244, 2638, 1860, 2015, 2093, 1663, 3249, 2588, 156, 1540, 1670, 2627, 1743, 2082, 2764, 1550, 239, 3261, 2297, 929, 2397, 111, 579, 2232, 1080, 485, 2240, 2067, 2986, 323, 361, 376, 7, 1635, 1272, 1393, 3205, 2506, 3262, 2285, 1596, 3112, 2378, 2444, 1061, 52, 1329, 367]. river [3417, 3366, 600, 19, 637, 1983, 3243, 3170, 1086, 466, 2812, 1606, 302, 1106, 2647, 144, 529, 156, 310, 560, 1997, 974, 3133, 1434, 1376, 1514, 2069, 1000, 3272, 1333, 1537, 80, 2884, 1662, 516, 2688, 2861, 577, 1096, 3439, 2522, 3015, 3459, 3219, 2862, 2707, 2860, 2285, 2556, 1306, 1709, 1374, 2637, 1849, 1602, 2095, 1059, 2238, 3279]. river-impacted [2397]. river-type [1456]. Riverine [2812, 2693, 539, 2093, 3029, 1514, 740, 2473, 1257]. Rivers [1083, 3480, 57, 2907, 1456, 764, 1323, 1937, 1581, 1933, 2561, 2635, 2147, 1540, 1530, 190, 3273, 2547, 1921, 2276, 2416, 2417, 2455, 624, 1343, 3205, 3186, 768, 1286, 1119, 3202]. rivularis [2431]. rivulatus [2161]. Rivulidae [1843, 2619]. RNA [2096, 519, 646, 2869, 32]. roach [569, 1323, 2536, 2363, 2191, 698, 951, 969, 5, 553, 1435, 685, 2390, 1273, 296, 2239, 991, 1289, 1640, 3214, 2315]. roacl [723]. robust [1793]. robustum [1348]. robustus [290]. Rocas [2079]. rock [900, 1003, 2037, 1491, 252, 799, 1458]. rock-dwelling [1491]. rock-pool [252]. Rockall [667]. rockfish [858, 3268, 366]. rockfishes [1858]. rockling [1047]. rockmover [1158]. rockpool [255]. Rocks [1916, 2780]. rockskipper [3210]. rocky [1805, 1488, 295, 2233, 2199, 863, 3396, 1735, 81, 917, 2875, 921]. Rodhe [3479]. Rodlet [1027, 1817, 1473, 3123, 1068, 2195, 23, 547]. rohita [3198]. rohu [3198]. roissali [1985]. Role [2412, 3055, 1172, 2092, 1632, 1920, 1198, 3235, 2890, 3351, 3427, 30, 2554, 89, 773, 942, 1077, 1616, 3327, 3423, 1075, 2154, 676, 178, 988, 2066, 738, 2163, 1097]. roles [293, 3416, 3045]. Rondelet [1405]. root [1894, 1893]. rosa [1680]. Ross [443, 283]. rostrata [3231, 3233, 2236, 3232]. rosy [2168, 2008]. rosylip [1351]. rosyside [524]. rotenone [699]. rotifers [3328]. roughhead [1079]. Roule [509]. round [2931, 2087, 2112, 3197]. rousseauxii [3449]. Roussillon [1360]. Route [3243, 2557]. routes [828]. routine [27, 503, 885, 1361, 1368]. roving [2275]. rows [1799]. Roy [3323]. RT [2739]. RT-PCR [2739]. ruber [2535]. Rudava [2069]. rudd [2315]. rudis [1468]. ruffe [125, 2933, 3061]. Rufiji [1537]. Rule [865]. rume [3298]. run [3002, 746, 575, 2249, 3472, 2040, 2109, 232]. Running [1159, 1935]. runs [2347]. Russia [1614]. Russian [1931, 282, 3453, 281]. ruthenus [1252, 1591]. rutilus [2315, 2536, 2999, 2845, 3288, 3214]. Rutlius [2315]. Ruzzante [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127].

S [1403, 1654, 1493, 1495, 3408, 1270, 3409, 3326]. S. [262]. S72 [2712]. sAAT [828]. sAAT-1 [828]. sablefish [286, 1647, 1338]. sac [1023, 2832, 429, 1081, 2427, 1295]. saddle [3458]. saddleback [2822]. Sado [1314]. sagax [3381, 3106]. sagittae [1679]. sagittal [2125, 2959, 1167, 1767]. Sahajärvi [3162]. sailfin [2099]. Saimaa [708]. Sainte [1748, 1642, 1333]. Sainte-Marguerite [1748, 1642, 1333]. saira [2217]. saithe [954, 179, 371]. Sakhalin [73, 150]. salangids [1818]. Salanginae [3289]. Salanx [3289]. salar [2907, 3193, 2043, 2642, 3069, 3450, 2974, 3389, 2367, 2572, 1207, 1584, 2157, 2392, 1851, 3469, 2115, 1612, 3066, 3202, 3443, 3131, 3181, 3261, 1590, 2895, 3121, 2986, 2923, 2593, 2605, 2963, 2567, 3129, 1598, 3186, 2962, 2545, 3083, 2922, 2519]. Salaria [2554, 2297, 1080, 3147, 3027, 798, 2444]. salaris [3056, 3197]. Sale [1494]. salina [2611]. saline [578, 2746]. salinities [470, 2195, 646, 1813, 2485, 2211]. Salinity [3240, 3178, 1632, 2086, 1856, 1608, 3231, 1163, 2993, 2222, 819, 1377, 3103, 3349, 938, 2156, 1348, 2488, 1536, 2615, 1368]. Salinity-linked [3240]. Salmo [2907, 1852, 3153, 3193, 2131, 1840, 2043, 1240, 2642, 3069, 2319, 3170, 3450, 2756, 2974, 3389, 234, 2367, 2572, 1207, 1584, 2157, 2392, 2893, 1851, 3097, 3469, 2115, 3179, 3174, 2371, 1612, 3066, 3202, 3029, 3133, 3443, 3248, 3131, 3181, 3122, 1278, 3261, 1590, 2718, 3134, 3067, 2895, 3121, 2986, 2923, 2593, 3258, 2605, 2963, 2567, 2000, 1260, 3129, 2544, 1598, 3186, 2962, 1884, 1358, 2129, 3332, 2843, 2545, 3083, 1624, 2922, 2519, 3336]. salmoides [1806, 290]. Salmon [1008, 1207, 148, 2461, 746, 57, 2907, 424, 412, 528, 1195, 1946, 188, 2055, 637, 2265, 2866, 3193, 589, 1891, 847, 1456, 1748, 192, 3191, 932, 1023, 2043, 1937, 1983, 2642, 3069, 844, 1192, 1108, 2283, 1513, 1171, 1559, 1200, 1524, 2812, 1481, 1935, 1889, 1933, 3450, 1560, 293, 2122, 1484, 302, 894, 2974, 3389, 177, 1248, 1382, 829, 2551, 2367, 1793, 1984, 692, 2572, 1584, 1646, 2157, 2392, 399, 2640, 931, 1851, 1634, 619, 893, 2948, 2901, 1126, 1890, 3469, 2060, 2115, 1765, 652, 1459, 429, 117, 871, 653, 1562, 943, 414, 1861, 623, 2635]. salmon [1288, 503, 1209, 587, 389, 1612, 3066, 2682, 3202, 1426, 1642, 1113, 560, 1530, 1997, 974, 1290, 2627, 2565, 3443, 1212, 1936, 1589, 1138, 1434, 1376, 2083, 1514, 2042, 2764, 1892, 3131, 708, 3181, 190, 1062, 2366, 2736, 641, 1690, 611, 3261, 1669, 1857, 1590, 985, 1017, 933, 929, 3041, 1219, 393, 1931, 3111, 1744, 1566, 2452, 885, 154, 950, 1567, 2686, 967, 627, 2441, 1930, 1333, 1940, 1938, 53, 681, 3140, 2403, 11, 3301, 3384, 2895, 439, 994, 80, 485, 2240, 2067, 3121, 2986, 1568, 1179, 140, 2923, 2040, 516, 1508, 2320, 2593, 416, 2688, 1310]. salmon [1569, 1365, 344, 105, 2586, 2605, 2963, 2567, 3195, 3184, 316, 1905, 219, 930, 2109, 1691, 1421, 2172, 976, 1145, 1510, 1764, 463, 83, 3361, 43, 287, 2860, 3129, 870, 2489, 1934, 1596, 1573, 2269, 1124, 1598, 13, 1761, 3186, 2962, 1231, 2835, 1910, 2190, 811, 1709, 815, 181, 3358, 3204, 2545, 206, 212, 3083, 1601, 1602, 1130, 1941, 1939, 1929, 1932, 2922, 1747, 1012, 1416, 1500, 1625, 597, 2615, 3404, 709, 2831, 650, 493, 2519, 1030, 220, 570, 1237, 2379, 2990, 367, 1586, 1774]. salmonae [1999, 472, 471]. salmonid [2447, 73, 2332, 1587, 1211, 941, 3138, 1122, 1364, 1340, 3480]. Salmonidae [2000]. salmonids [800, 2192, 2188, 1984, 2743, 1950, 2124, 2059, 1563, 1208, 1769, 1190, 96, 376, 3054, 1600, 2750]. Salmoniformes [1833, 2136]. Salmoninae [166]. salmoninarum [88]. salmonis [2115, 1479]. salmopunctatus [2458]. salpa [1822]. Salt [199, 342, 341, 1613, 2734]. salt-dependent [342, 341]. Saltatory [2448]. saltatrix [1407]. Salvelinus [3161, 3290, 1324, 1704, 2905, 2932, 2198, 2756, 3460, 2550, 3045, 3102, 3029, 3122, 2909, 3372, 2726, 2286, 1728, 3370, 2299, 3128, 1854, 3058, 3056, 3197, 2133, 2958]. same [1610, 1711]. sample [144, 2833]. sampled [901, 2610, 3404]. samples [107, 2831]. Sampling [257, 699, 1534, 2198, 1631, 255, 1554, 2414, 1678, 2609]. Samuel [1405, 3408]. San [1494, 2222]. sand [1672, 936, 779, 842, 2484, 2284, 3103, 2779, 1395, 669, 754, 2689]. sand-goby [2779]. sandbank [1085]. sandbar [982]. sanddab [2369]. sandeel [2295, 1389]. Sander [1871, 2723, 3343, 3171, 2778, 3155, 2954, 1870, 3162]. sandperch [2870, 3176]. sandwich [3138]. Santa [1975, 2369, 343, 2501]. São [1137, 2472, 3401, 1996, 1896]. sapidissima [52]. saprolegniosis [2590]. Sarangi [3323]. sarasinorum [2947]. sarba [2211]. Sarcocheilichthys [2853]. Sarda [2180]. Sardina [2705]. Sardine [3106, 2769, 2992, 469, 2562, 175, 638, 421]. Sardinella [2657]. sardines [3381, 2705]. Sardinops [3381, 3106, 3008, 2992]. Sarotherodon [261]. Sarpa [1822]. Satellite [1073, 3292, 2854, 3194]. satiation [1185]. Saurenchelys [1912]. Saurida [2451]. saury [2217]. Savannah [2420, 670]. sawfish [2819]. sawfishes [1745]. sawtail [2543]. saxatilis [2270, 2607, 2216]. scad [3207, 564]. Scaglia [1843]. Scale [2106, 2451, 1031, 1983, 1183, 3360, 3066, 2303, 2845, 3273, 2634, 3448, 2159, 35, 1957, 2475, 2322]. scale-eating [2475]. scales [107, 1697, 2336, 3190, 2303, 2603, 641, 2847, 536, 3258, 1144, 3119, 2489, 2609, 2856]. Scaling [984, 48, 993, 1703, 2689]. Scalloped [3216, 2139, 1749]. Scanning [543, 94]. Scardinius [2315]. scarecrow [1291]. Scaridae [2822]. Scarring [2880]. Scartella [3396, 1042]. Scartichthys [81]. Scartomyzon [886]. sceleratus [1727]. Schädigung [3411]. scherzeri [3352]. Schilbeidae [1282]. Schindleria [2886, 787, 1661]. Schindleriidae [2886]. Schinz [2883]. Schistocephalus [1414, 2262, 3421, 2219, 2573, 2506, 2029]. Schistocephalus-infected [2219]. Schizothoracinae [2520]. schmitti [211]. Schneider [2692, 2826, 2144]. schoenleinii [3025]. school [991]. Schooling [2655, 2584, 260, 204, 426, 991]. schools [2632]. schreyeni [2132]. Schroeder [1405]. Schultz [2437]. schwanenfeldii [2850]. sciadicus [3167]. Sciaenidae [1156, 1977, 1480]. Sciaenops [1866]. Science [1496, 1402, 1774, 1652, 3408, 2977]. Sciences [1493, 1405]. Scientific [1, 1628, 1404]. scissortail [1053]. sclateri [1105]. Sclerochronology [1775]. Scomber [3157, 1797, 2428, 2427, 2806]. Scomberomorus [2108]. scombrus [3157, 1797, 2428, 2427]. scope [3189, 717]. Scophthalmus [1995, 2507]. scoping [1243]. Scorpaena [836]. Scorpaenidae [3172, 2596]. scorpaeniform [1696]. Scorpaeniformes [44]. Scotia [3349]. Scotian [1686, 1658, 1719]. Scotland [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 668, 960, 2481, 2317, 2893, 913, 1539, 3186, 2345, 3409]. scotti [1777, 1750]. Scottish [3134]. Screening [107]. scrutiny [718]. sculpin [1792, 973, 942, 887, 2470, 2063, 1351, 121]. sculpins [806, 705]. Scyliorhinus [549, 2888, 2330, 2878, 2993, 2385, 1090, 2483, 1721]. SE [1253]. Sea [2139, 1121, 443, 2007, 2565, 3472, 283, 2396, 3002, 746, 263, 97, 2271, 932, 311, 664, 2502, 642, 2387, 1413, 1524, 1112, 42, 128, 910, 1608, 2781, 286, 1015, 928, 829, 730, 1205, 15, 868, 2640, 326, 291, 795, 3457, 1851, 1056, 619, 2523, 2373, 444, 575, 2115, 2920, 2840, 658, 1318, 2682, 3202, 310, 1613, 3029, 3275, 2170, 1376, 2249, 483, 701, 3183, 622, 329, 496, 309, 63, 1436, 1565, 649, 173, 885, 374, 3134, 3391, 3150, 1055, 579, 2986, 2040, 416, 28, 1103, 3078, 574, 517, 999, 1905, 31, 3364, 1503, 1635, 2887]. sea [1691, 232, 1479, 1724, 2423, 320, 1343, 476, 983, 2602, 417, 287, 748, 2183, 2824, 2748, 3362, 586, 2533, 1981, 2205, 2244, 240, 1012, 1364, 1416, 2797, 2211, 3251, 2318, 3257, 793, 2597, 1727, 346, 3073, 2877, 2070, 1749, 1409, 2952, 729, 179, 982, 2781, 3207, 2171, 1682, 2886, 2624, 2250, 2640, 387, 3215, 1833, 2647, 2295, 129, 2041, 2885, 3174, 1952, 3105, 1318, 1210, 2291, 2226, 1025, 1997, 1434, 2865, 611, 509, 1132, 2967, 923, 3472, 2634, 2949, 68, 3030, 1912, 1735, 1220, 2337, 1549, 3285, 3373, 2986, 1074, 2507, 2372, 157, 452, 955, 169]. Sea [371, 467, 1816, 2401, 1193, 450, 306, 2697, 469, 1985, 2578, 809, 2483, 1721, 2264, 2871, 1543, 3244, 2162, 1868, 1352, 203, 2870, 3176, 2048, 2646, 1602, 968, 1830, 709, 2345, 1782, 3257, 2180, 889, 3356]. sea-cage [2205]. sea-cages [309]. sea-ranched [1103]. Sea-run [3472, 746, 2249, 2040, 232]. sea-trout [868]. seabass [1578]. seabream [459, 2546, 1608, 3149, 1618, 2965, 1622, 1623, 3091, 3040]. seadragons [816]. seagrass [2851, 2284, 1380, 824, 2569]. seagrass-associated [1380]. seahorse [350, 2185, 2606, 695, 2377, 3374, 2474]. seahorses [888, 1429, 2865, 1216, 1337]. Seal [2931, 1336]. seals [1336]. seamount [1978]. seaperch [2285]. seaperches [3172]. Seaport [1824]. Sear [3480]. search [2448, 965, 2554, 126, 867, 1186, 873]. season [1084, 2093, 2879, 929, 1620, 416, 8, 313, 1289, 1640, 754]. season-dependent [1289]. Seasonal [2306, 3153, 1417, 30, 187, 989, 2016, 5, 1811, 2700, 2008, 23, 2834, 863, 1789, 2259, 854, 485, 323, 467, 878, 1479, 3027, 2163, 870, 1910, 813, 121, 3176, 1064, 1322, 2454, 3200, 207, 1782, 3251, 1131, 2081, 2994, 860, 500, 27, 730, 2900, 1611, 2294, 139, 2588, 3401, 16, 1311, 708, 2973, 2936, 3238, 1953, 151, 654, 56, 1065, 2047, 624, 2556, 3228, 1532]. seasonality [2656]. seasonally [2167, 2399]. seasons [1111, 1162, 291, 249]. seatrout [2937, 1480, 1368]. Seaward [1892, 2825, 399]. seawater [2265, 1321, 885, 1898, 1296, 463, 206, 1169]. seawater-acclimated [1321]. Sebastes [1858, 3268, 1284, 3078, 1812, 2875, 697]. Sebert [3407]. Second [306, 1573, 2797, 2379]. second-generation [1573]. Secondary [3195, 2732, 2326, 2060, 2718, 1380]. secretion [761, 3031, 671, 1742]. secretions [3027]. sectatrix [604]. sections [2027]. sedentary [2944]. Sediment [104, 807, 953, 768]. sedis [3292]. seen [1999, 3426]. segregation [960, 495, 523, 1204, 1083, 686, 295, 2755, 2420, 3357, 1311, 2468, 2955]. segregations [2483]. select [1951]. selected [2106, 3305, 2400, 2449, 2895, 983]. selection [1181, 424, 3304, 1344, 2966, 736, 555, 445, 3270, 3314, 2809, 2902, 817, 3416, 104, 1375, 1861, 623, 61, 1102, 3273, 2705, 2452, 3391, 3005, 3373, 2417, 2954, 3283, 2323, 1182, 1854, 353, 300, 1638, 145, 76, 1233, 3051, 881, 2080]. Selective [1935, 344, 3158, 1419, 460, 2114]. selectivity [216, 3246, 1048]. self [2391, 2664, 1740]. self-feeding [2391, 1740]. self-fertilizing [2664]. Semaprochilodus [764]. semelparity [335]. Semelparous [324, 589]. Semen [672, 3327]. semi [3193, 3367, 2724, 1997, 3423, 1105, 2479]. semi-arid [1105]. semi-enclosed [3367]. semi-lunar [2479]. semi-multiplex [2724]. semi-natural [3193, 3423]. semi-wild [1997]. semilaevis [3293]. seminal [1128, 671]. semoni [2957]. Semotilus [1968]. Senegal [372, 3269]. senegalensis [262, 2611, 2096, 3269, 2720, 1820]. Senegalese [2720]. sensitive [1443, 1837]. sensitivity [1506, 866, 2124, 2509, 927, 2366, 2153, 983]. Sensory [576, 2466, 3016, 497]. senta [3470]. sentinel [781]. separate [2543, 267]. separated [3170, 91, 804]. September [2255, 2424]. Sequence [312, 3290, 34, 2790, 2713, 2130, 712, 2924, 790, 1860, 18, 886, 1753, 96, 1518, 3037, 3219, 422, 556, 971]. sequences [788, 840, 956, 3174, 3431, 2981, 1049, 292, 1401, 2540, 267, 2529]. sequencing [2789, 2706]. sequential [1834]. Serbia [1538, 2544]. sergeant [1053]. serial [3157]. Series [1653, 263, 3267, 676, 2343]. Seriola [662, 2391, 244]. serotonergic [746]. serranid [1975, 2666, 2612]. Serranidae [2052, 2564]. Serranochromis [290]. Serranus [1167]. Serrapinnus [2862]. Serrasalmus [3401, 2314]. Serum [1387, 2120, 150, 2897, 1930, 574, 2922, 3346]. set [2303, 3121]. Seton [3184]. settled [710]. Settlement [2781, 1191, 3314, 2772, 2412, 3299, 794, 2046, 2229, 521, 3232]. seven [2272, 3345]. several [327, 3295, 1666]. severe [2622, 2299]. Severn [19]. Sex [1131, 284, 2440, 2849, 2089, 1211, 2508, 205, 2246, 574, 2618, 2820, 1268, 1713, 1052, 2643, 153, 814, 3435, 2271, 311, 1409, 172, 2554, 795, 1547, 515, 1250, 773, 243, 749, 61, 2879, 2867, 1980, 937, 3311, 873, 309, 1273, 2161, 1436, 2402, 3026, 1222, 2677, 2726, 2154, 1829, 80, 3296, 2926, 361, 2320, 1710, 3308, 647, 1905, 2494, 1887, 1262, 1791, 2163, 1307, 2661, 1385, 514]. Sex-biased [1211, 514]. sex-changing [3311]. sex-determining [2726]. Sex-related [574]. sex-role [773]. Sex-specific [1131, 2820]. sexes [1667]. Sexing [1769]. Sexual [2676, 584, 887, 2656, 2137, 430, 2364, 1061, 962, 1181, 188, 200, 2346, 24, 3386, 1297, 177, 3213, 2120, 61, 2565, 401, 2002, 63, 2699, 649, 2097, 3195, 1373, 2468, 1093, 2494, 2483, 2618, 238, 1854, 2268, 3470, 348, 2142, 1862, 2290, 2661, 2180]. sexually [2449, 978]. sGnRH [118]. sGnRH-A [118]. shad [162, 3095, 1043, 2823, 3028, 2405, 2841, 1169]. Shadow [868]. Shag [1916, 2780]. Shallow [3432, 3172, 3242, 862, 546, 1683, 1435, 2004, 1953, 920, 921]. shallow-water [546]. shanny [2874]. Shape [1996, 3351, 725, 2832, 2355, 215, 1040, 2961, 1123, 2602, 2955, 2477, 2344, 697, 2939, 3065, 1747]. shaped [2362, 2445, 2894]. shaping [3220]. shared [1343]. Shark [2925, 2800, 3141, 3465, 3003, 2864, 2995, 667, 3000, 3017, 2904, 2948, 394, 2105, 2016, 2184, 3216, 2174, 2304, 3445, 2891, 497, 988, 2953, 2808, 2842, 2871, 3064, 2183, 2389, 2839, 1853, 2386]. shark-fin [3064]. sharks [2941, 982, 872, 89, 2079, 264, 1504, 1159, 1471, 2005, 3308, 3364, 3283, 2468, 1182, 2880, 826, 2672, 2386]. sharksucker [1142]. sharpnose [2800, 3465, 1504, 2016]. sharptail [763]. sharptooth [2969]. Shedding [1051]. shelf [2780, 3379, 2510, 277, 2640, 1783, 1034, 3200, 1658, 1719, 1686]. shell [1458]. shell-bed [1458]. shelter [3452, 1459, 1190, 1569, 2874]. shelter-limited [3452]. shi [1328]. shift [3170, 338, 139, 1718, 2679, 698, 2999, 3066, 2292, 1786, 319, 881]. Shifts [141, 2484, 2837, 2416, 1956, 2109, 775, 2669, 1322]. Shiisorapuchi [2094]. shiner [1233]. shiners [1342, 2591]. shoal [523, 718, 1887]. shoal-mate [718]. Shoaling [2486, 124, 1342, 859, 248, 3247, 3383, 1256, 1887, 1269]. shoals [588, 1844]. shock [2773, 33, 3007, 1363]. shocks [208]. shore [1805, 1247, 1735]. shore-bases [1247]. shoreline [485, 2625]. shores [3396, 917]. Short [691, 1504, 3210, 490, 3259, 1416, 54, 392, 2877, 2387, 3158, 2126, 1571, 121, 1130]. short-horn [121]. short-lived [3158]. Short-term [1504, 3210, 490, 3259, 1416, 54, 392, 2877, 2387, 2126, 1130]. shortbelly [366]. shortens [3094]. shorter [2899]. shortest [3101]. shortfin [839, 2685]. shortfinned [194, 2530]. shorthorn [887]. shortnose [2440, 3137]. shovelnose [114, 1787]. show [2571]. shown [938, 1306]. shows [836]. shrimp [2851, 240]. shrimps [468]. Shrinkage [1815, 2821, 1046, 1492, 1283]. Shumagin [2610]. shy [1004]. shysharks [1336]. sialic [804]. siamensis [3366]. Siamese [3366, 391, 584]. Siberia [3290]. Siberian [2740, 1860, 715, 942, 511, 565]. Sibling [216, 217, 3179]. Sibling-size [3179]. siblings [1117]. Sicily [1982]. Sicyopterus [3079]. sidearm [1531, 1532]. sidearm-system [1531, 1532]. Sierra [1843]. sieve [71]. siganid [935]. Siganus [2098, 1980, 1304]. signal [1459, 1221]. signals [2872, 1246, 1154, 1227, 3096, 185]. signature [2493, 2323]. signatures [2730, 2281, 1144]. significance [888, 2619, 2181, 1262, 1431, 3217]. Significant [3012, 3040, 608]. significantly [2215, 1455]. Sillaginodes [710]. Sillago [3103]. Siluranodon [1282]. silurid [1036]. Siluridae [491]. siluriform [2134]. Siluriformes [2658, 3440, 3429, 1141, 2832, 3382, 2370, 542, 458, 1996]. Silurus [1790, 1798, 2944, 59]. silver [2825, 936, 1641, 2356, 3243, 839, 170, 1715, 2983, 572, 3221, 1420, 1686, 2376, 3244, 2378, 2211]. silvering [1715]. silversides [548]. similar [2213, 3464, 2260, 377]. similarity [2523, 2858, 2416, 3096]. Simojoki [1997, 1434]. simple [64, 1718, 1769, 2971]. simplex [2068]. Simpsonichthys [2619]. simulated [1320, 445, 2623, 3139, 878, 3252, 2]. Simulation [1957, 1608, 1186, 1229]. Simulation-based [1957]. simulations [2795]. Simultaneous [2722]. sinensis [3046]. Singapore [1404]. single [790, 2030, 1665, 3383, 2720]. Siniperca [3352]. Sinnamary [7]. Sinocyclocheilus [3220]. Sinos [2169]. Sinusoidal [802]. sirenian [1142]. Sisoridae [542]. sister [3096]. Site [2771, 2936, 265, 2932, 1474, 1832, 1985, 353, 1531]. Site- [1531]. sites [3296]. situ [325, 3433, 227, 794, 1049, 727, 726, 2313, 32]. Sivas [1707]. six [2516, 3428, 349, 1698, 1371, 1639, 2167, 1742, 3466, 1600]. Size [3246, 3386, 1948, 1807, 1014, 1249, 3469, 2692, 242, 2408, 993, 1290, 66, 58, 2882, 775, 2354, 815, 2768, 1888, 960, 3107, 954, 19, 1084, 216, 217, 3073, 569, 2227, 1038, 2192, 1241, 1983, 523, 735, 890, 555, 3270, 729, 3333, 302, 1546, 958, 177, 1165, 1189, 604, 2640, 110, 387, 303, 3360, 3171, 2248, 2900, 2294, 2832, 1185, 2438, 144, 538, 3179, 522, 1419, 1637, 1771, 2469, 845, 803, 198, 942, 248, 969, 3011, 1460, 1971, 2058, 1174, 1081, 2164, 584, 340, 1118, 1589, 2348, 1434, 1040, 1516, 2973, 2936, 305, 1188, 1752, 1048]. size [85, 2659, 3180, 1216, 75, 3028, 1515, 1388, 933, 655, 3013, 2553, 111, 1053, 3026, 2666, 865, 3140, 2403, 3301, 1803, 897, 1123, 1291, 3373, 1160, 1916, 831, 1874, 252, 1190, 2097, 1760, 2063, 2156, 1710, 3308, 3300, 1001, 1708, 2187, 376, 2229, 36, 2216, 1065, 298, 330, 2109, 521, 3365, 1887, 8, 3159, 238, 1865, 1466, 948, 26, 2, 1338, 3266, 1264, 566, 69, 1761, 807, 1424, 2598, 2268, 3470, 385, 770, 1140, 1638, 2566, 237, 2604, 1676, 754, 2661, 1385, 2503, 2029, 1368, 1030, 2531, 2689, 486]. Size-assortative [2768]. Size-dependent [1807, 1014, 3469, 58]. size-frequency [1760]. size-matched [1388]. Size-related [1948, 3073, 729, 3333, 1676]. Size-scaling [993]. size-selection [1638]. size-selective [1419]. size-selectivity [1048]. Size-specific [2354, 807]. sizes [3314, 3045, 178]. Skagerrak [1425, 2343]. Skate [3105, 842, 1726, 3386, 3124, 2766, 2002, 2585, 3280, 2649, 2984, 2268, 3470]. skates [1682, 1468, 2984, 2331, 464]. Skeletal [93, 1885, 1506]. skeleton [326, 1295, 1714]. skills [478]. Skin [1005, 745, 1820, 1742, 362, 2300, 2935, 3210, 2509, 1077, 434, 757, 1914, 82, 1279, 3329]. Skin-type [745]. skipjack [2140, 2927]. skipped [253]. Skjoldal [3410]. skull [2778, 3078, 556]. skunk [3333]. sleeper [2184, 2183, 2386]. Sleeping [1158]. slender [1899, 2938]. slender-type [1899]. slimy [806]. slope [445, 101, 1682, 2904, 3326, 44, 2118, 3390]. slopes [667, 2547]. Slovenia [234]. slow [2245]. Small [2634, 1728, 3448, 3234, 2461, 907, 2757, 3255, 1183, 1948, 912, 1204, 2929, 2251, 2330, 2878, 103, 1100, 1799, 1563, 2273, 2674, 2603, 2627, 2027, 2936, 66, 1388, 2137, 2776, 321, 2159, 1018, 1636, 1272, 768, 2856, 2804, 892, 3222, 207, 1782, 1340, 3403, 841]. Small-scale [2634, 3448, 1183, 2159]. small-spotted [2330]. smallmouth [2956, 327, 46, 2111]. smalltooth [2819]. smell [1212]. smelt [2084, 2340, 2670, 2347, 2592, 2164, 1677, 2429, 271, 846, 669, 2957, 1322, 1676, 754, 1630]. smelts [2222]. Smith [253]. smithi [2885]. Smitt [2033]. smoked [2798]. smokey [1233]. smolt [1171, 2122, 2682, 2565, 1434, 708, 1669, 1744, 627, 3134, 2923, 416, 663, 287, 1573, 2190]. smoltification [1179, 1910]. Smolting [708, 885, 870]. smolts [424, 2265, 2812, 910, 2974, 3389, 1634, 1610, 943, 3202, 1997, 1690, 439, 2923, 2688, 1421, 1124, 1709]. smooth [2871, 3470]. smooth-hound [2871]. smoothhound [2953]. snailfish [443]. Snailfishes [2230]. snakehead [1945, 2847]. snapper [107, 621, 2022, 3005, 3375, 2418, 3065]. snappers [2144]. sneaker [1091]. sneaking [3426]. Sneaky [2947, 116]. snipefish [2181]. snook [3048]. snout [3329]. SNP [2321]. SNPs [2546]. snyderi [2661]. sobaity [309]. Soca [234]. Social [478, 844, 1108, 3302, 2632, 1215, 393, 1114, 2494, 962, 3094, 158, 583, 1198, 1511, 1249, 2849, 1208, 3256, 358, 1103, 3201, 3266, 811, 1232]. socially [60]. Society [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2381, 2436, 2465, 2492, 1651, 2411, 2513]. Socioeconomic [2601, 909]. sockeye [2461, 847, 1456, 2283, 1889, 1484, 2640, 389, 1857, 3041, 2452, 3301, 1365, 3184, 1764, 83, 3358, 650]. sodium [2437, 720]. sofie [1949]. soft [917, 921]. sole [346, 1277, 1314, 2902, 372, 751, 2037, 3269, 2720, 1352, 556].

Solea [262, 2611, 2902, 2510, 2096, 3269, 2720, 1820, 50]. soles [2096, 799]. solidus [1414, 2262, 3421, 2573, 2029]. solidus-infected [3421]. solute [850]. solutions [3378]. somatic [392, 2534, 387, 2213, 3463, 2845, 747, 716, 1309, 2119, 363, 3387, 2820]. Somatolactin [2923]. Somatostatin [448, 489, 1089]. somatostatin-14 [1089]. Some [839, 3285, 271, 1446, 171, 373, 2359, 3260, 1476, 2030, 1980, 1528, 1107, 1783, 3070, 1599, 1600, 922]. something [233]. Somniosus [2386, 2386]. Songsjøen [236]. sonic [16]. sonicate [2332]. sophore [74]. sordidus [2369]. Soufrière [604, 605]. Sound [975, 1439, 847, 1811, 2189, 253, 2221, 2306, 1546, 1357, 3297, 2802, 1430, 1830]. sounding [411]. sounds [1533, 3458, 3333, 2287, 1107, 3402]. source [957, 2120, 2473]. sources [2642, 3297]. South [3479, 787, 1649, 2174, 1336, 1495, 557, 7, 2468, 1306, 2201, 3107, 2800, 1398, 3191, 2934, 1347, 2098, 2439, 2641, 2756, 1982, 839, 1650, 2275, 2620, 2724, 465, 2630, 2846, 623, 1139, 1845, 2660, 2964, 2002, 2181, 1520, 703, 1468, 1719, 2708, 1784, 1938, 1123, 1662, 831, 2882, 2320, 3182, 3447, 2152, 3245, 1178, 1305, 2871, 2943, 2618, 1656, 461, 1352, 2357, 1358, 1458, 2377, 1057, 1385, 2107, 2141, 2911, 2368, 912, 1412, 2658, 2420, 3348, 3082, 2858, 1394, 2650, 1410, 1036, 322, 1916, 2806, 557, 379, 1896, 2737, 3172, 1374, 1911, 1286, 2750, 512, 378, 2243]. South [1050, 3042]. south-central [3107, 1139, 1845, 2871, 1358]. south-east [1347, 2098, 2181, 3182, 3447, 1036]. south-eastern [2800, 2439, 839, 2630, 2846, 2660, 2964, 2002, 1123, 1662, 1305, 2943, 2618, 1352, 2357, 2377, 1385, 2107]. south-west [1398, 2934, 2641, 2756, 465, 623, 703, 1468, 1719, 2708, 1784, 2882, 2320, 2152, 1178, 1656, 1057, 512]. south-western [2275, 2620, 2724, 1520, 831, 3245]. southeastern [598]. southern [78, 1168, 1409, 211, 1606, 3450, 2781, 2735, 1147, 1360, 2616, 2658, 3381, 3307, 3152, 700, 1952, 2369, 1025, 2912, 2585, 3272, 3372, 3238, 2174, 2304, 359, 1728, 1399, 675, 419, 3309, 1570, 1808, 2747, 669, 65, 1638, 2290, 920, 2537, 3353, 889, 346, 2780, 299, 1051, 276, 2984]. Southernmost [106, 1178, 1940]. southwestern [92]. Sox [356]. Soybean [3083, 2265]. sp [1680, 1699, 44, 3199, 1835, 2395, 2313, 889, 488]. sp. [2919]. space [604, 1566]. spacer [2987]. Spacing [204]. spacings [3066]. Spain [1937, 600, 614, 156, 1784, 228]. span [2497, 3101, 2501]. Spanish [513, 1110, 896, 1343, 1170, 591, 1445, 1865, 2108]. sparid [1347, 916]. Sparidae [586]. Sparisoma [2822, 3448, 2158]. Sparus [1604, 1608, 2781, 1650, 3275, 3223, 2170, 3150, 2965, 2602, 2211]. spathaceum [3058]. Spatial [1946, 2336, 2420, 622, 1671, 824, 1656, 2937, 968, 2629, 1277, 2055, 1937, 1183, 3084, 3190, 1422, 3452, 2633, 2037, 2603, 2168, 2008, 112, 1665, 2160, 3284, 3258, 521, 2856, 3176, 2378, 3065]. Spatially [817, 2421, 3170, 2422, 3138, 2091]. Spatio [2496, 3353]. Spatio-temporal [2496, 3353]. spawned [387, 2240, 2067]. spawner [613, 1351]. spawners [367]. Spawning [1792, 3379, 1106, 2459, 1785, 2287, 1226, 1636, 3480, 1854, 1306, 2910, 3002, 975, 528, 935, 3076, 2227, 1840, 192, 2877, 3458, 687, 24, 2504, 2368, 1423, 2347, 1731, 291, 1851, 2294, 301, 2438, 2420, 3398, 998, 1486, 2022, 1138, 1831, 819, 2217, 3226, 1216, 941, 2390, 496, 3139, 3267, 3331, 2686, 633, 959, 2258, 874, 938, 2625, 2455, 740, 3269, 2421, 1620, 797, 2612, 2479, 1503, 3365, 2057, 3227, 253, 1450, 2971, 2136, 26, 10, 2162, 799, 1319, 300, 1986, 3225, 2290, 1939, 1713, 3306, 3257, 1881, 2080]. Spawning-related [2459]. spawning-run [3002]. spawnings [360]. special [171, 2480, 3407, 2425]. specialization [2085, 2909, 3372, 2395, 2839]. Specialized [3388, 2039, 836]. Speciation [2919, 2852, 3416, 3415, 333, 3340, 1073, 1854, 2853]. Species [2872, 384, 601, 2796, 2537, 2503, 570, 1052, 1533, 2081, 226, 1035, 1521, 2516, 3451, 1387, 415, 3073, 2988, 607, 2192, 2134, 2857, 94, 582, 1111, 2543, 1413, 443, 2130, 1684, 3017, 2272, 3207, 1203, 234, 2929, 289, 1733, 1360, 277, 256, 295, 2233, 2213, 307, 1833, 2724, 2294, 2061, 725, 3344, 144, 2472, 1810, 2110, 3136, 1044, 1187, 354, 1683, 2846, 2633, 2517, 2719, 1098, 2709, 2124, 2273, 1698, 3357, 2314, 2147, 1664, 2759, 2733, 1471, 1639, 1491, 3299, 1102, 3292, 2632, 1040, 2826, 1067, 403, 1027, 1844, 2508, 1987, 1734, 2896, 608, 3295, 1783, 559, 2587]. species [3345, 2360, 2528, 916, 1482, 592, 2791, 3445, 3280, 2968, 734, 1918, 1770, 2495, 1046, 3092, 1666, 1835, 705, 2754, 711, 897, 182, 542, 491, 1036, 2135, 2206, 2455, 2884, 149, 2156, 2189, 3330, 695, 2169, 3096, 3439, 2167, 275, 279, 1258, 3201, 3365, 624, 1326, 2500, 2873, 1018, 3015, 176, 1742, 1073, 2774, 2485, 2399, 2498, 2906, 3057, 2651, 1839, 603, 2984, 2655, 727, 1231, 267, 1374, 1458, 2811, 422, 2054, 3135, 50, 2701, 2231, 2529, 3342, 1883, 2797, 378, 1236, 2831, 1465, 971, 3042, 2141, 3454, 2913, 762, 3115]. species-rich [2701]. Species-specific [384, 2537, 570]. Specific [1074, 1339, 1131, 384, 3035, 2642, 293, 399, 1790, 791, 2730, 1734, 3038, 288, 2354, 807, 2820, 2537, 1838, 3080, 570, 1149]. specifically [474]. specificity [1308]. specimen [1799, 502]. specimens [1321]. Spectral [2153, 2124, 2072]. spectrins [2467]. spectroscopic [2840]. spectrum [1238, 580, 1516]. speed [2215, 2818, 3337, 3026, 869, 2814, 371, 826, 3332, 185]. speeds [3243, 1844, 1675, 2681]. spent [494]. Sperm [1423, 3031, 3335, 2604, 2123, 2805, 551, 2396, 2044, 1199, 1087, 42, 1163, 1069, 665, 2571, 2915, 893, 264, 389, 927, 3293, 895, 1408, 1216, 1528, 2784, 116, 1315, 336, 258, 1795, 2364, 3195, 1448, 1450, 1870, 1780, 1809, 720, 628, 760, 1129, 1237, 2379, 2365]. sperm-dependent [2784]. Sperm-duct [3031]. Spermatic [3430]. spermatocrit [1450]. spermatogenesis [1507]. spermatogenetic [2358]. spermatogenic [573]. spermatozoa [672, 114, 431, 934, 149, 105, 573, 420, 2596, 2805, 629]. spermatozol [149]. spermatozoon [671]. spermiation [258, 118]. spermiogenesis [420]. Spermonde [2851]. Sphoeroides [1821, 1148, 2646]. Sphyraena [452, 2224]. Sphyraenidae [452]. Sphyrna [3216]. Spiggin [2573]. spike [534]. spike-tailed [534]. spinax [2995, 3000]. spine [3359, 487]. spined [54, 413, 1131, 1414, 2241, 1842, 2262, 601, 2121, 3254, 3084, 3422, 100, 1163, 2460, 3424, 3416, 243, 1335, 1139, 1845, 2089, 3413, 2226, 2733, 2471, 756, 2249, 403, 3472, 3423, 3071, 3421, 2413, 487, 3425, 3418, 2524, 2801, 1741, 731, 781, 2432, 2219, 2950, 2323, 576, 3420, 1346, 2477, 782, 2322, 2668, 3426, 1234, 1235, 3400, 3306, 1270, 2029, 939, 3419, 3417, 551, 377, 940]. spines [1770, 1723]. spinifer [2141]. spinosus [2711]. spiny [210, 2815, 3047, 1802]. spiny-eel [1802]. spirlin [2069]. spleen [1312, 562, 3128, 2318]. splendens [2138]. splice [2636]. Spontaneous [330, 503, 3383]. spot [684, 3025]. spotlighting [2273]. spottail [2591]. spotted [761, 984, 2888, 3039, 2330, 2878, 470, 3152, 2998, 3102, 2058, 895, 3122, 3247, 2385, 1047, 536, 579, 865, 1740, 3182, 3447, 1985, 2937, 124, 1155, 1480, 1368, 2133, 2958, 3352]. spp [1156, 2872, 2721, 2526, 808, 1842, 2262, 2600, 179, 3017, 1204, 2332, 3105, 3236, 3341, 2181, 2791, 785, 1899, 1795, 2544, 2673, 846, 3074, 1167, 2922]. sprat [2424, 3377, 1549, 1074, 119, 2057]. Spratelloides [3158, 1986]. Sprattus [2424, 3377, 1703, 2057]. spread [563]. spreading [1264]. Spree [1106]. Spring [3010, 492, 808, 3076, 2640, 387, 1025, 2217, 2421, 3387, 1897, 2136, 26, 10, 3404]. spring-spawned [387]. spring-spawning [3076, 2421, 2136, 26, 10]. spumigena [2801]. spumigena-exposed [2801]. spurt [453]. Squalius [679, 2735, 1667, 2148]. squaloid [264]. Squaloidea [2995]. Squalus [2815, 3047]. squamosus [667, 2904, 264]. square [1894, 1893]. squares [423]. Squatina [1034]. Sr [2202, 830, 2210, 1331, 3074]. S(R) [3403]. SSCP [355]. St [466, 1147, 3233, 604, 700, 2637, 697, 605, 1855]. stability [2092, 958, 2115, 2355, 1146, 1209, 140, 2187, 3259, 3176]. Stable [960, 2756, 3095, 2303, 2482, 3445, 754, 2771, 651, 2031, 3025, 2281, 2284, 1539, 2833, 3310, 479, 294, 2064, 3156]. Stable-isotope [3445, 2833, 3156]. stage [1856, 3315, 1825, 1715, 1463, 1746, 1966, 2056, 2637, 945]. stage-explicit [1746]. stage-structured [1856]. stages [1023, 2611, 839, 165, 2496, 189, 692, 2684, 2373, 656, 3024, 1460, 2826, 2366, 2945, 3038, 885, 366, 3089, 2559, 2491, 1120, 2537, 2180]. staghorn [2063]. Staging [1294, 3267, 1306]. Staiger [502]. stain [1345]. stained [3376]. staining [827, 898, 1688, 344, 1720, 2173]. Stam(R) [1068]. stanauli [3440]. standard [2215, 830, 198, 885, 154, 140]. standardizing [2288]. Starks [2768, 2661]. starry [2608, 2476, 1173]. start [2262, 2296, 1375, 2155]. start-feeding [1375]. starts [121]. Starvation [2555, 743, 2815, 1017, 2427]. starved [519, 1813]. Starving [2424, 301]. State [1616, 364, 10, 1437, 1435, 508, 3332]. state-dependence [1437]. State-dependent [10]. States [2439]. Static [623, 2614]. station [792, 914]. stationary [405]. Statistical [1977, 820, 9]. Statistics [3323]. Status [609, 625, 158, 1633, 1511, 3235, 2711, 2985, 2502, 3034, 2188, 470, 3424, 2849, 658, 393, 3111, 41, 633, 1310, 1570, 1518, 2494, 3324, 45, 3266, 1432, 811, 2752, 512]. steady [1895]. steelhead [583, 3368, 1610, 927, 2765, 1575]. Stegastes [2125, 2166]. Steindachner [2691, 2344]. stelgidolepis [193]. stellatus [2608, 2476]. Steller [2183]. stenohaline [540]. stenolepis [2959]. stenothermal [280, 774]. stenotopy [2473]. stepfather [2460]. Stephanoberycidae [1918]. Stepping [1347]. stercusmuscarum [1684, 1684, 599]. Stereolepis [2523]. Stereological [633]. sterlet [1252, 1591]. Steroid [937, 309, 821, 3011, 1062, 649, 151, 2677, 678, 221, 1791, 3420, 2268, 1392]. steroidogenesis [1045]. steroidogenic [319]. steroids [2020, 1547, 1273, 2413, 797, 1710, 1791]. Stichaeidae [1488]. stickleback [1414, 2241, 3414, 355, 2121, 3254, 1163, 1069, 3416, 243, 1335, 1139, 1845, 2089, 2226, 2249, 3472, 1253, 3421, 2413, 3418, 2926, 2524, 731, 781, 2432, 2219, 2323, 576, 3420, 1346, 2477, 782, 2322, 2668, 3426, 1235, 3412, 3306, 2029, 3419, 551, 377, 940]. sticklebacks [54, 413, 1131, 200, 1842, 2262, 3084, 3422, 100, 2460, 3424, 3415, 3413, 756, 403, 85, 3423, 3071, 487, 3425, 1180, 2801, 1741, 2573, 2950, 145, 1234, 3400, 1270, 939]. sticky [71]. still [2950]. stillwater [1576]. stimulant [2028]. stimulates [2437]. stimulation [1002, 489]. stimuli [761, 2165]. stimulus [2965]. stingray [3160, 2624, 2931, 850, 1263, 2112]. Stock [2443, 1582, 2342, 1110, 288, 2108, 3276, 689, 19, 2934, 2558, 3242, 1430, 302, 796, 696, 2341, 2764, 704, 513, 1592, 1464, 857, 1284, 143, 2090, 3219, 1543, 294, 1542, 1424, 1576, 1723, 1801, 1911, 2870, 3176, 2344, 2478, 670]. stock-enhanced [670]. Stock-specific [288]. stocked [1687, 600, 3255, 2070, 1581, 2561, 382, 1419, 2690, 612, 460, 1568]. Stocking [1577, 78, 1938, 1331, 2972, 1606, 3231, 805, 1419, 202, 2946, 2986, 2210, 948, 1586]. Stocking-related [1938]. stocks [1852, 751, 1589, 2143, 3287, 1621, 2338]. stoichiometry [2361]. Stolephorus [123]. Stomach [429, 545, 1894, 2563, 2866, 79, 1631, 427, 325, 1952, 508, 557, 775, 809, 566, 1354]. stomach-content [2866]. stomachless [552]. stomachs [1987, 559, 1816, 1676]. Stomiidae [2960]. Stomiiformes [2960, 2858]. stone [1347, 1678, 1262, 2476]. stoneroller [1698]. stoplight [3448]. storage [2081, 3282, 42, 722, 490, 1994, 2045, 1870, 1228, 2036]. stored [2596]. stores [188, 3191, 685, 2]. storm [1159]. straight [2929]. strain [1588, 2726, 2895]. strains [473, 3355, 1829, 2895, 437, 2038, 1934]. Strait [633]. strata [2553]. strategies [888, 2234, 326, 53, 322, 29, 1182, 1231, 1469]. strategy [1716, 1559, 2319, 1731, 125, 2904, 1952, 950, 681, 2258, 2891, 3369, 3079, 201, 669]. straying [974]. streaked [209]. Stream [1010, 1200, 3192, 1687, 1320, 525, 583, 2757, 900, 3069, 2319, 1948, 3450, 2756, 2234, 234, 15, 2658, 636, 103, 1950, 290, 3179, 2630, 2059, 414, 2635, 2194, 1037, 1210, 1976, 3248, 2094, 933, 1075, 111, 823, 1955, 654, 585, 581, 3138, 321, 1519, 524, 2091, 1636, 2299, 870, 2012, 1535, 2856, 512, 881, 2958]. stream- [870]. stream-and [414]. stream-dwelling [525, 900, 1950, 2059, 1037, 3248, 2094, 933, 585]. Stream/Oceanography [1010]. streams [1852, 373, 558, 2053, 1684, 2447, 2551, 1353, 256, 468, 1563, 623, 1968, 2273, 3212, 2027, 496, 949, 2843, 892, 1340]. streber [250]. strength [2446, 1064]. Stress [286, 454, 3094, 1859, 1320, 1804, 1821, 1281, 3198, 199, 77, 1525, 2157, 3024, 1363, 1113, 1915, 1879, 2769, 2200, 2817, 1526, 102, 487, 1359, 2593, 2838, 56, 3201, 43, 2038, 2824, 60, 1981, 3404, 2474]. stress-coping [3024]. stress-induced [1915]. stress-related [1821]. stressful [2728]. stressor [208]. stressors [3325, 3203, 781]. striata [1945]. striatus [3210]. strike [2608]. striped [2282, 544, 2270, 1157, 1486, 2607, 2216, 229, 2301, 1354, 670, 2313]. Strong [591, 2719, 3167]. strontium [1481, 830, 3372, 501, 1331, 2238]. Structural [3367, 1413, 2114]. Structure [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 2847, 1164, 2118, 757, 917, 2113, 966, 3107, 689, 2207, 908, 3130, 1456, 569, 2404, 2043, 94, 1534, 2710, 1805, 3271, 3220, 2735, 1430, 101, 796, 73, 2619, 2275, 1422, 2079, 728, 538, 541, 522, 2341, 3102, 653, 3249, 1968, 242, 3401, 1210, 1614, 1025, 2342, 2348, 3467, 2921, 2736, 1980, 622, 401, 2705, 513, 3167, 2397, 671, 41, 1079, 3121, 675, 2855, 1287, 564, 3119, 1974, 1065, 3117, 1819, 35, 1742, 2144, 1739, 1543, 276, 294, 1661, 774, 2322, 2344, 2529, 1011, 1929, 968, 3239, 1713]. structure [2108, 49, 3040, 2531, 2990, 3228, 3276]. structured [1856, 3248, 2421]. structures [415, 1701, 2645, 572, 3461, 1519, 507]. structuring [1632, 1982, 751, 3432, 2337, 2539, 2345]. studied [1503, 1272]. studies [2925, 1181, 2771, 932, 1199, 307, 2052, 1588, 51, 2337, 1930, 1144, 411, 2114, 3219, 2939, 3400, 3412]. Study [348, 3318, 1847, 1323, 2134, 888, 2710, 2481, 2980, 2145, 1041, 2751, 1204, 2446, 1375, 1251, 2371, 84, 1822, 771, 1118, 3248, 236, 1438, 2644, 1070, 1864, 612, 2480, 1160, 369, 142, 1328, 999, 1635, 980, 1958, 1914, 1596, 2673, 1455, 2117, 3056, 1270, 3329, 3257, 2859, 3438]. studying [2866, 1263, 395]. stuhlmannii [1537]. Stunted [3162]. sturgeon [3094, 1387, 114, 3325, 2440, 2740, 1611, 715, 2828, 2511, 3137, 2409, 565, 1787, 3046]. sturgeons [1582]. stutter [90]. style [2552]. styles [3024]. sub [2907, 1247, 2579, 110, 2635, 2118, 3334, 1364]. sub-Antarctic [110, 2118]. sub-arctic [2635, 2907]. sub-lethal [2579, 1364]. sub-tropical [3334]. subalpine [732]. subarctic [57, 2032, 1102, 1311, 1969, 1019, 80, 485, 2240, 2067, 323, 2688, 2088]. subdivision [3366]. subfamily [2520]. subgenera [2619]. subisoform [3080]. subject [1517]. subjected [2773, 783, 2617, 636, 2437]. sublethal [189, 3456, 1369]. Sublittoral [1863]. subordinate [1260]. subordinates [2849]. subsequent [2215, 784, 1310, 2567, 3259]. subsidy [3452]. substance [648]. substance-P [648]. substances [2437]. Substantial [2720]. substitute [1647]. substitution [2148]. substrate [1240, 1265]. substratum [3005]. substructure [2328]. substructuring [488]. subterranean [1137]. subtidal [2024, 863]. subtle [52]. subtropical [2994, 2684, 397, 3396, 2169, 822]. subunit [2331, 2636]. subunits [2677]. subyearling [1892, 3204]. Success [2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513, 484, 1581, 42, 631, 100, 1248, 1069, 2745, 694, 1187, 2749, 2727, 1976, 3211, 1482, 1567, 252, 1365, 2429, 1985, 1448, 1870, 2172, 2802, 167, 1308, 1325, 1693, 1331, 811, 1028, 1454, 1078]. Successful [2625, 3184]. sucker [539, 14, 2946]. suckermouth [2832]. suckers [801]. sudden [2957]. suffering [746, 3444]. suffering-centred [3444]. suggest [1458]. suggesting [1999]. suggestions [601, 922]. suggests [2995, 3348, 3432]. Suitability [1116, 740, 213, 1982, 1136]. suitable [672, 3197]. Sul [1495]. Sulawesi [2851, 2947]. sulcus [1156]. sulphidic [2804]. Sulphur [2639]. sulphuraria [2804]. Sumatra [182]. summary [2594]. Summer [2635, 1784, 2088, 2383, 77, 1530, 2730, 1062, 151, 2276, 1391, 2054]. summering [3186]. sunbleak [1920, 2760, 1187, 1341, 1956, 2746, 2485]. Sundadanio [2676]. Sundasalangidae [1818]. Sundasalanx [1818]. sunfish [139, 3205, 406]. sunfishes [2747]. supernatants [1555]. Supplement [924, 925, 1626, 1959, 1960, 1942, 1961, 2349, 2350, 2351, 2761, 1943, 1962, 2352, 2762]. supplementary [1170]. supplementation [719, 259, 658]. supplemented [1885, 2203, 891]. Supplements [2333, 2549, 2742, 2149]. supply [1348, 605]. support [2907, 3097, 511]. supports [790]. suppresses [677, 447]. suppression [1530]. surface [3002, 1893, 94, 2619, 223, 3401, 2153, 1090, 2648, 1996, 812, 1097, 1834]. surface- [2153, 2648]. surface-dependent [1893]. surfperch [3168, 2282]. surgeon [1041]. surgical [1697]. surgically [2374, 2283, 1071, 2810, 3111, 2696, 976]. surmuletus [3073, 1275]. surplus [845]. surprising [2582]. survey [2927, 3159, 556]. surveys [3394, 3064]. Survival [189, 1378, 440, 1669, 329, 2843, 1131, 3130, 3255, 719, 2283, 1580, 1111, 3132, 1607, 1041, 2023, 3389, 1660, 1984, 2499, 2157, 2392, 1561, 2242, 1419, 1504, 98, 1211, 2682, 3202, 3196, 2810, 1997, 3133, 1434, 819, 3378, 518, 1218, 949, 2769, 967, 3301, 1568, 585, 2320, 3385, 2589, 2605, 2963, 320, 83, 3159, 683, 2285, 948, 1573, 3056, 2685, 1490, 3352, 486]. survive [3018]. survivors [2804]. survivorship [1889, 2491]. susceptibility [575, 3046]. suspended [953]. suspension [71]. suspension-feeding [71]. Sustainability [2601]. sustained [649]. sustained-release [649]. sustaining [3089]. suttkusi [306]. suture [1041, 2810]. suvmuletus [901]. Sweden [622, 3139, 1253]. Swedish [282, 917, 1636, 921]. Swim [1734, 931, 1126, 347]. swimbladder [961, 563, 3110, 1312, 3244, 1834]. swimbladders [227]. Swimming [1578, 3305, 3235, 1850, 985, 2409, 3384, 2668, 367, 152, 625, 1523, 1842, 1895, 2215, 2296, 230, 2283, 2282, 1560, 338, 2974, 260, 2579, 1584, 931, 2818, 2848, 1797, 84, 1363, 710, 432, 701, 3183, 1844, 2644, 3100, 1452, 3111, 3026, 3301, 1160, 113, 416, 371, 1296, 976, 141, 1466, 2228, 2628, 826, 1675, 846, 121, 181, 185, 2681, 1449, 1012, 1364, 2305, 2689, 802, 835]. Swiss [1921]. switched [2899]. switched-off [2899]. Switzerland [3055]. swordfish [620, 91, 2358, 3062, 1763]. swordtails [499]. sydneyanus [724]. symbionts [3373]. symbioses [1246]. symbiosis [1259]. Sympatric [3226, 3290, 3073, 262, 1684, 1689, 1204, 2551, 2317, 2394, 3045, 623, 2635, 202, 3464, 2314, 1311, 1969, 702, 992, 1759, 1468, 2528, 2837, 3350, 1462, 1190, 2423, 1764, 2955, 2129, 3466, 1469, 3115, 697, 145, 3222, 1522, 903]. sympatry [1073]. Symphodus [1985]. Symphysodon [2325]. Symposium [537, 1324, 1902, 1963, 1706, 1824, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2381, 2436, 2465, 2492, 2788, 2411, 2513, 409, 436, 877]. Sympterygia [2408, 831]. Symptevygia [703]. Synaphobranchidae [1912]. synapomorphies [1141]. synapomorphy [3471]. Synbranchiformes [1802]. synchronization [221]. synchronize [2456]. synchronized [3080]. Synchrony [1239, 2688]. syndrome [2822]. Syngnathidae [2865, 1898, 2154, 2301]. Syngnathoides [3151]. Syngnathus [2970, 759, 2873, 1970, 2301]. Synodontidae [2451]. synonym [2132]. synonymy [334, 2858]. synopsis [1733]. synthesis [3120, 1681, 1593, 1483]. system [3153, 263, 2956, 2134, 549, 1171, 2980, 928, 563, 1634, 2100, 2840, 2482, 1822, 16, 2764, 3106, 1394, 1176, 1357, 205, 1864, 2664, 592, 2961, 3134, 1222, 3296, 2965, 2293, 3219, 1374, 2322, 1624, 3400, 1155, 1531, 1532, 3092]. systematic [957]. systematics [1035, 1745, 50, 2706]. Systems [1651, 2510, 307, 1251, 3372, 649, 2765, 1105, 919, 3, 1657, 405].

T [1654, 3478, 2980]. T-cell [2980]. tabira [3309]. Tableand [375]. tactic [2289, 1140]. tactics [2461, 1248, 2554, 1186, 324, 120, 1254, 777, 1042, 475, 778, 3303]. Tadeusz [3006]. taenia [2733, 2471]. taeniatus [3318, 1896]. taeniolatus [1788]. taeniops [2597, 2961]. tag [872, 2810, 1879, 1721, 2685]. tag-recapture [1721, 2685]. tagged [1811, 1514, 58, 2172, 2172]. tagging [1474, 872, 2674, 432, 981, 3108, 2045, 3400, 3276, 1878]. tags [2374, 2934, 639, 179, 2924, 2251, 18, 1540, 2810, 3394, 3142, 419, 1994, 2045, 1900, 2376, 2012, 533, 3194]. Tagus [262, 2096]. Taihu [2238]. tail [2606, 991, 3046]. tail-fin [3046]. tailed [534]. taimen [73, 150, 1860]. Taiwan [1428, 765, 3272, 3115]. take [2877]. Takeuchi [3478]. Taking [3475, 3391]. tale [2460]. tambaqui [2197, 2707]. Tambre [2860]. tamoxifen [530]. Tana [1203, 560, 1376, 516, 1709]. tandem [1604]. Tanganyika [1383, 334, 1098, 550, 1458, 2132, 2475, 2811, 3342]. Tanganyikan [2142]. tank [1675]. tanks [1578, 583, 309, 1596]. tannery [82]. Tanzania [2981]. Tapajós [3454]. taranetzi [1792]. target [1633, 2503]. tarpon [2582, 2683]. Tarquinia [1613]. task [1183]. Taste [1476, 370, 674]. tau [1832]. taxa [2011, 2960, 2858, 267]. taxifolia [753, 1275]. taxon [251]. taxonomic [2711, 2985, 3221, 41, 1974, 1518, 3324, 2529, 2732]. Taxonomy [2146, 2709, 3288, 2806]. Tay [960]. Taylor [1406]. TCDD [1355]. TCDD-induced [1355]. Teaching [1261]. technical [909]. technique [2138, 3193, 1889, 2572, 2582, 588, 1665, 1909, 2971, 971]. Techniques [2425, 827, 639, 327, 2996, 1203, 3453, 1448]. technological [1594]. Technology [1824]. Tecocomulco [2914]. teeth [2925, 1799]. Tejo [2735]. Telemetering [1247]. telemetry [2771, 1889, 2371, 2810, 2946, 1855, 1503, 1635, 1272, 2556, 2637, 181, 533]. teleost [410, 2816, 311, 3035, 326, 1832, 3001, 1980, 2847, 273, 2715, 1896, 1661, 3051]. Teleostei [2516, 415, 2009, 3435, 3044, 2600, 601, 2711, 2985, 2713, 1488, 2532, 3220, 3428, 1729, 1141, 1733, 2619, 2658, 2353, 3213, 3430, 2620, 3326, 2885, 2692, 3136, 334, 1098, 2709, 3292, 1831, 2865, 2851, 2547, 2778, 552, 3125, 2370, 1918, 2987, 2779, 3199, 3092, 2135, 2731, 458, 2166, 2855, 3022, 3048, 2520, 573, 2794, 2325, 1326, 2000, 2737, 1458, 1661, 2804, 2596, 3303, 2141]. teleosts [1035, 1446, 2272, 2030, 280, 1128, 1528, 2005, 3208, 1473, 812, 2467, 2117, 1499, 1981]. Telestes [2913]. Telmatherina [2947, 3222]. Telmatochromis [334, 1098, 1458]. Temaiken [2408]. Temperate [1447, 2580, 2876, 1716, 3240, 2359, 2496, 1975, 303, 3360, 2588, 1371, 280, 2007, 2412, 3467, 863, 724, 1046, 2669, 3135, 920, 2969, 1329]. temperate-zone [3240, 1371]. Temperature [1506, 3158, 2235, 2848, 1371, 1437, 2422, 3081, 316, 1033, 2453, 214, 2970, 954, 392, 1891, 3377, 1023, 2642, 1856, 3093, 1481, 187, 33, 286, 2251, 1189, 1660, 1609, 387, 2213, 1631, 1461, 3230, 2373, 2191, 2438, 2093, 803, 1152, 969, 1460, 3202, 2058, 2663, 1447, 1530, 3443, 1966, 1831, 305, 701, 677, 1673, 1987, 2390, 3337, 1452, 1377, 717, 1756, 1789, 3268, 1829, 1420, 1310, 48, 28, 1708, 2505, 288, 2559, 3369, 320, 463, 2, 1338, 726, 1973, 3135, 385, 770, 1853, 1266, 313, 1640, 76, 3055, 406, 493, 3080, 1368, 3251, 1881, 835]. Temperature-dependent [2235, 1033, 1530]. temperatures [625, 3198, 2179, 2640, 3230, 194, 951, 2056, 440, 1062, 2366, 626, 1898, 2614, 3018, 2906]. Temporal [2055, 1871, 2932, 1171, 2347, 1732, 3171, 3456, 2115, 14, 2037, 154, 1898, 2067, 1717, 2187, 2992, 119, 521, 2423, 2875, 1676, 1277, 908, 1937, 2336, 2496, 3075, 2202, 1547, 2420, 1146, 112, 622, 2160, 1671, 2937, 451, 1830, 3353, 365]. temporalis [334, 1098]. tench [2302, 2617, 989, 490, 1659]. tending [2867, 2817]. tengana [2135]. Teno [80]. tension [1033]. tentacles [3329]. tentacularis [1250]. tentative [2532]. tenuis [3267]. Tephrosia [1463]. terapon [3030, 3030]. teres [2087]. term [54, 392, 2877, 1983, 1513, 1869, 1658, 2387, 2481, 2392, 1029, 46, 2993, 1504, 1746, 3210, 1071, 2126, 1588, 2815, 432, 490, 732, 1916, 3108, 517, 2109, 772, 2665, 995, 628, 2533, 533, 3259, 1130, 1416]. Terminos [1671]. terraenovae [3465, 2016]. Terrebone [2941]. terrestrially [3267]. Territorial [95, 2166, 1224, 2043, 736, 482, 3102, 1002, 3067]. territoriality [803]. territories [2966, 1508]. territory [1174, 2583]. Terwinghe [3322]. Tesch [1496]. test [453, 3076, 2404, 2932, 1771, 1209, 1002, 1184, 145, 3197]. testacea [2837]. testes [1216]. testicular [221, 964]. Testing [1568, 2930, 2263, 1182, 423]. Testis [264, 263, 2599, 1162, 2293, 2518]. testosterone [1387, 2413, 530, 451, 1289, 1659]. tests [2818]. testudineus [2972, 3020, 1148]. Tethering [3201]. tetra [2712]. Tetracapsuloides [3055]. tetradactylum [3276]. tetranucleotide [90]. Tetraodon [873]. tetras [2237, 2437]. tetricus [2120, 1770]. texanus [2946]. Texas [1342]. Texoma [548]. textbooks [1219]. texture [3002]. Thai [1883]. thalassinid [2851]. Thalassoma [3458, 67, 1274]. Thaleichthys [1815, 2670]. Thames [911, 119]. their [2825, 459, 1035, 171, 860, 1581, 2532, 1202, 958, 829, 1014, 2571, 1056, 3430, 2460, 301, 1589, 1138, 1754, 1040, 1067, 3221, 3100, 1473, 460, 2240, 2206, 366, 2186, 369, 1103, 2589, 3146, 2943, 1791, 2485, 591, 870, 3057, 2476, 1057, 2668, 697, 2444, 3303, 1149]. theories [1182]. theory [2550, 511, 1213, 3272, 1219, 1233, 2930]. Theragra [1815, 2985, 1729, 2610, 3278, 2422]. theraps [3030]. there [2044, 686, 465, 1217, 1220, 2860]. Thermal [2270, 3097, 2427, 113, 970, 1305, 451, 153, 814, 932, 910, 3075, 77, 737, 2658, 3213, 914, 14, 3131, 2826, 1836, 4, 2838, 878, 2614, 2208, 2112, 1837]. Thermally [3183]. thermo [142]. thermo-gradients [142]. Thermolabile [515, 814]. thermoregulation [3204]. these [8]. Thiaminase [709, 2831]. thiamine [746, 1871]. thicklip [1699]. thiols [2639]. thorny [2268]. Thorpe [1496]. thread [3019]. threadfin [162, 2041, 401, 2949, 102, 3276]. Threat [1443, 602, 1130]. Threat-sensitive [1443]. threatened [2261, 2694, 3365, 1374]. Three [3413, 2399, 1052, 54, 413, 1533, 545, 1131, 2081, 1521, 1414, 2241, 1387, 2777, 1842, 2262, 2134, 94, 2121, 3254, 3084, 3422, 888, 482, 1948, 100, 2924, 1175, 295, 2460, 3424, 3416, 2472, 243, 1335, 2124, 2089, 646, 2226, 1491, 1311, 1969, 756, 2249, 403, 3472, 941, 3423, 51, 3071, 949, 3421, 2528, 2413, 487, 2791, 2968, 3425, 612, 3418, 897, 1315, 2455, 149, 2524, 2801, 1741, 2187, 731, 781, 2432, 2500, 176, 2219, 2950, 1120, 3219, 2774, 2323, 576, 3420, 3129, 1780, 1346, 2477, 2357, 1374, 2467, 2054, 2668, 3426, 1234, 1235, 3400, 2095, 3306]. three [1270, 2029, 939, 207, 3419, 551, 377, 940]. Three-spined [3413, 54, 413, 1131, 1414, 2241, 1842, 2262, 2121, 3254, 3084, 3422, 100, 2460, 3424, 3416, 243, 1335, 2089, 2226, 756, 2249, 403, 3472, 3423, 3071, 3421, 2413, 487, 3425, 3418, 2524, 2801, 1741, 731, 781, 2432, 2219, 2950, 2323, 576, 3420, 1346, 2477, 2668, 3426, 1234, 1235, 3400, 3306, 1270, 2029, 939, 3419, 551, 377, 940]. threespine [355]. thresher [2174]. threshold [2623]. throat [947, 1011]. throated [2120]. thrombocytes [2117]. throughout [997, 196, 2541, 258, 2602]. Thryssa [123]. thumb [2943]. Thunnus [629, 1895, 2599, 2130, 2616, 3381, 3334, 1762, 2113, 422, 2584, 3353, 3194, 3356]. Thymallus [1850, 2341, 2732]. thynnus [629, 2599, 2113, 3356]. thyroid [187, 1063, 489, 2441, 31, 1441, 1841, 2887, 1066, 879]. thyroxine [1617, 2887]. Tibetan [2520]. Tidal [2388, 1210, 911, 915, 3472, 2634, 2562, 1153]. tidal-mixing [2634]. tidepool [973, 1805]. tidepools [2875]. tides [1669]. tiger [2606, 2290]. Tigris [1846, 2290]. tilapia [153, 814, 1485, 2534, 1579, 3274, 3260, 866, 2555, 3196, 2917, 2632, 677, 3203, 2691, 716, 1082, 1359, 1164, 330, 441, 2734, 1136, 2673, 49, 354, 785, 2206]. tilapiine [696, 1059]. Tilesius [2848]. Timbalier [2941]. time [2739, 1748, 1418, 210, 24, 494, 661, 2679, 1419, 2450, 2645, 3275, 1104, 2427, 3391, 1160, 178, 1316, 2172, 26, 1693, 2343, 3135, 1625, 1881]. time-keeping [3275]. time-to-fatigue [1160]. times [884, 1243, 1987, 1107, 559]. Timing [2346, 2070, 2908, 2438, 2031, 3133, 1589, 3261, 959, 2470, 80, 2063, 2109, 2843, 2518, 3352]. Tinca [2302, 2617, 1791]. tissue [1885, 3282, 1847, 2374, 132, 170, 18, 1612, 2281, 1841, 2887, 983, 1149]. tissues [1633, 837, 2316, 1720]. titanium [1136]. Titicaca [3125]. toad [462]. toadfish [2306, 2991, 128, 16, 3392, 678, 1291, 306]. Tocantins [1137, 1996]. Tödürge [1707]. tolerance [2270, 3213, 155, 1082, 590, 2734, 1018, 1305, 2208, 2747, 2277, 1097]. tolerant [2399]. tomato [3433]. tool [3318, 1165, 963, 3310, 2870]. tools [2996, 2733, 1427, 1284, 2478]. toothcarps [896]. toothfish [1084, 2207, 2033, 557, 276, 380]. Toothless [1203, 1282]. topminnow [3167]. topmouth [1605, 2757, 2760, 2221, 2746, 2485, 3099]. Torpedo [1544, 2293]. Tortugas [3263]. total [589, 1150, 980]. tournaments [1973]. towing [286]. toxic [462, 2801, 2774]. toxicity [1463]. toxins [1005]. toxostoma [1949]. TR [2559]. Trace [465, 824]. tracer [2064]. tracers [2482]. Trachidermus [2110]. Trachurus [2988, 3207, 2344, 1767, 1768, 2568]. tracked [816]. tracking [528, 3229, 1822, 828, 3394]. tract [315, 648, 1015, 1175, 2400, 427, 3166, 196, 1619, 2158, 351, 3362]. Tractable [1182]. trade [2034, 3064, 406]. trade-off [406]. trader [3064]. trahira [2274, 2505]. trahiras [2472]. trailing [1847, 991]. training [478, 2245]. trait [2638, 120, 2449]. Traits [682, 1423, 2179, 2745, 1250, 1335, 1590, 2585, 3013, 1109, 2292, 2763, 581, 634, 254, 2364, 3195, 3117, 778, 3395, 1334, 1656, 2792, 1597, 2444]. trammel [1554]. trans [3229]. trans-Atlantic [3229]. transboundary [1456]. transcription [1604, 2593, 2715]. transect [3394]. transfer [1320, 2341, 155, 116, 885, 2527, 574]. transferred [2484]. transferrin [3169]. transferrins [1930]. transfers [128, 2993]. transformation [2923, 2190]. transformed [2025]. Transgene [2534]. Transgenerational [3310, 3156]. transgenesis [2534, 1082]. transgenic [3120, 3371, 2716, 2384, 1113, 985, 1017, 3100, 441, 13, 1761, 3346]. transition [1052, 2282, 2412, 869, 2814]. transitional [453, 2088]. transitions [363, 1575]. Translocations [2199]. Transmission [3061]. transmitted [1690, 185]. transmitter [1697, 1822, 826, 533]. transmitters [2283, 1003, 1041, 1293, 1071, 3111, 2696, 976, 2228, 181]. transparency [948, 1313, 2375]. transparent [952, 1867]. transplants [1474]. transponder [639, 872, 307, 1540, 3212, 3108, 405]. transport [554, 3337, 1380, 999, 3404, 2474, 3228]. transthyretin [31]. trapped [2954]. travel [126, 3332]. Traverse [1493]. trawl [2766, 566]. treated [1681, 82, 220]. Treatment [677, 1964, 1387, 1555, 649, 2833, 258, 150]. treatments [3130]. Trematochromis [2132]. Trematoda [2262]. trematode [3059]. Trematomus [1777, 1750, 2467]. Trench [3014]. trend [1073]. Trends [913, 1983, 1244, 2553, 3284, 2752]. Trent [2884]. Trewavas [2527]. tri [2609]. tri-monthly [2609]. triacylglycerol [1794]. Triaenophorus [3161, 2558]. trials [1984, 441, 650]. tributaries [19, 310, 3279]. tributyltin [2040]. tricaine [2198]. Trichiurus [2529]. Trichomycteridae [2658]. Trichomycterus [2053, 2658, 1088]. tridentatus [2793]. triggered [307, 2030, 771]. triggerfishes [482]. Trihydroxy [678]. triiodothyronine [1441]. trim [89]. trimaculatus [2508, 1045]. Trimma [2402]. Trinidadian [22, 1241, 116]. Triphoturus [1401]. triplefin [2233]. triploid [625, 1560, 291, 3457, 2302, 2901, 1867, 626, 43, 246, 1295, 2722]. triploids [2237]. triploidy [1113]. Triplophysa [1680]. Tripterygiidae [2233]. triradiatus [3329, 2832]. Trisopterus [41]. tRNAs [2921]. Troadec [1775]. Trophic [2807, 2846, 2868, 1034, 1292, 3073, 500, 2952, 2756, 2929, 289, 3241, 1422, 651, 2879, 1040, 2909, 1759, 916, 1539, 3396, 675, 479, 1974, 2159, 2955, 2609]. trophic-level [2952]. Trophically [1690, 2085]. Tropical [1007, 699, 753, 1632, 2994, 3036, 2065, 1281, 2404, 906, 1805, 1684, 2272, 2359, 2263, 468, 2014, 3062, 1683, 2846, 2194, 2613, 2603, 2851, 592, 1918, 2606, 1046, 3334, 2182, 2618, 2220, 1986, 2231, 1853, 2683, 2280, 1159]. tropicus [3036]. Trough [667]. Trout [1008, 2544, 1687, 453, 438, 600, 152, 625, 1172, 1558, 1320, 672, 1577, 2203, 2131, 3130, 554, 1840, 78, 1641, 1121, 2215, 2214, 736, 1240, 1418, 855, 687, 268, 317, 539, 3302, 1457, 1111, 2667, 2710, 1349, 2319, 3170, 1555, 1681, 1581, 884, 910, 1243, 535, 2557, 712, 338, 2756, 33, 1689, 532, 1825, 682, 829, 2551, 956, 2317, 2499, 2561, 15, 77, 868, 805, 1525, 1814, 299, 2893, 12, 2818, 2743, 191, 1851, 1634, 249, 2908, 2550, 636, 1461, 3097, 2590, 382, 635, 339, 103, 444, 575, 1375, 2942, 3179, 1197, 3174, 2623]. trout [2727, 623, 2635, 656, 942, 2273, 3024, 1209, 2371, 1037, 1363, 1971, 310, 927, 1529, 2032, 3029, 2810, 164, 1118, 1002, 527, 2218, 3133, 3443, 3248, 972, 2083, 2700, 3122, 1668, 483, 370, 1921, 3327, 622, 329, 236, 496, 3139, 2945, 1483, 1999, 2881, 949, 559, 1452, 828, 2266, 2116, 2718, 1109, 526, 88, 626, 489, 612, 460, 3134, 1333, 676, 2286, 978, 3067, 1420, 1143, 2986, 1740, 2625, 1049, 3441, 113, 970, 585, 581, 1433, 2312, 1369, 634, 474, 3258, 2001, 1296, 1089, 1103, 2589, 321, 56, 1905, 1373, 1441, 2559, 130, 2522]. trout [1066, 744, 1479, 1724, 2423, 2501, 1305, 463, 1260, 598, 1636, 1272, 804, 472, 1386, 2038, 1362, 3393, 1466, 2003, 60, 454, 1004, 1231, 2835, 1455, 2012, 1535, 1846, 1319, 3332, 2843, 185, 1981, 810, 1624, 76, 368, 448, 892, 2028, 570, 3336, 2989, 2080, 3279]. trouts [2052]. trucha [2368]. trumpeter [544]. trunk [562]. truss [215]. Trust [2077]. trutta [3153, 1624, 1852, 3153, 2131, 1840, 1240, 2319, 3170, 2756, 2893, 1851, 3097, 3179, 3174, 2371, 3029, 3133, 3443, 3248, 3122, 1278, 2718, 3134, 3067, 2986, 3258, 1260, 1884, 2129, 3332, 2843, 1624]. Trygonoptera [2837]. Trypsin [2502]. tryptic [843]. tshawytscha [2901, 2137, 2320, 3361, 3204, 2615, 3404]. Tsukamoto [3478]. tubular [1841]. tubules [1148]. tucunaré [2232]. tumbil [2451]. tumour [2982]. tumour-necrosis [2982]. tuna [689, 2599, 1058, 2616, 3381, 633, 2119, 2140, 2927, 3334, 640, 2498, 2906, 2113, 348, 422, 2584, 3353, 3194, 3356]. tunas [3334]. Tundra [282]. tuner [2728]. tunnel [931]. turbid [3357]. turbidities [2393]. Turbidity [2006, 544, 2603, 2432, 1736, 1551, 1455]. Turbinenbedingte [3411]. turbot [24, 784, 2821, 2920, 342, 341, 235, 473, 646, 717, 2507, 1570, 517, 999, 995, 1265, 1264, 1470, 3319]. turbulence [1160, 3385]. turbulent [1207, 1584, 1646, 241]. Turkey [1707, 1358]. Turkish [2131, 2988]. turn [2608]. turning [2608, 2296, 126]. turnover [53]. tusk [647]. tuskfish [3025]. Tuusulanjärvi [1677]. twentieth [918]. Twenty [3316, 2924, 90]. Twenty-four [3316]. Twenty-three [2924]. Two [443, 2782, 2469, 2804, 315, 2328, 625, 2803, 2526, 1722, 3451, 957, 808, 935, 415, 3073, 2262, 549, 1133, 1534, 3170, 1413, 2362, 3017, 1204, 3241, 2638, 2510, 2561, 77, 604, 480, 2213, 2394, 67, 465, 264, 3136, 1044, 2633, 2719, 3432, 2059, 2635, 1845, 3464, 803, 3357, 2147, 646, 2759, 2007, 3299, 2964, 3467, 708, 702, 1019, 3273, 1027, 3462, 401, 1252, 1836, 1468, 3421, 1075, 1104, 916, 2837, 592, 626, 2403, 705, 1829, 363, 491, 2135, 2206, 2625, 2884, 2473, 1284, 475, 3284, 2189, 3330, 3287, 3008, 2414, 279, 2520, 271, 919, 624, 1224]. two [29, 1479, 896, 2873, 1401, 437, 2955, 1739, 123, 2038, 2747, 2136, 2236, 267, 124, 2843, 1458, 2475, 1469, 257, 1161, 1638, 3115, 2529, 212, 2820, 2732, 1522, 750, 2569, 971, 2386, 2379, 3468, 1340, 3279]. two-banded [465]. two-dimension [2803]. two-second [2379]. two-spotted [124]. two-year-old [708]. type [1456, 1694, 3270, 3368, 2061, 414, 2810, 1114, 1899, 2276, 745, 3252, 770, 782, 76, 2560, 793, 84]. type-indicator [2061]. Types [433, 1694, 3314, 110, 2403, 2040, 1622, 1742, 386, 782, 2322]. typhle [2970]. typical [2148]. typological [3082]. typus [3141]. tyrannus [2770]. tyrosine [3354]. Tyrrhenian [2139].

U.K [1654, 1901, 2279, 2309, 2381, 2436, 2465, 2492, 2411, 2513]. U.K. [2077, 2457, 1187, 3421, 1553, 3253]. U.S. [2236]. U.S.A [1901]. U.S.A. [2941, 3386, 1975, 12, 3010, 2093, 2420, 2759, 548, 1462, 1348, 2747, 337, 3192, 3012, 670]. UK [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127]. ukliva [2913]. Ukrainian [6]. ultrasonic [1855, 1713, 1787]. ultrasonically [816]. Ultrastructural [1411, 1999, 1914, 888, 1321, 1413, 1162, 51]. Ultrastructure [629, 573, 427, 547, 149, 420]. Ultraviolet [1491, 1381, 2280, 2446]. umbla [3161]. Umbrina [880, 1644]. Ume [1602]. unaffectedby [330]. underestimate [561]. underestimated [533]. undergoing [2558, 254]. underneath [289]. understand [1889]. Understanding [2749, 1944, 3305, 758, 2337, 3232]. underwater [3023]. undulatory [2644]. undulatus [1372]. unequal [3045]. Unexpected [901]. unexploited [2925]. unicornis [848]. uniporum [2794]. Unique [2164]. unit [126, 2143]. United [1653, 2439, 2743]. uniting [3471]. units [91, 3286, 3043, 611, 608, 2343]. Universidade [1495]. University [1403, 1902, 1963, 1706, 1990, 2019, 2049, 2078, 2101, 2127, 2279, 2309, 2077, 2381, 2436, 2465, 2492, 2411, 2513]. Unpredictability [70]. unreliable [2300]. unsolved [223]. unsuccessful [981]. until [879]. untreated [442]. Unusual [2921, 895, 1467, 479, 1661, 2023, 84, 3296]. update [2733]. Upeneus [1700]. upland [1581, 2561, 2219]. upon [1525, 2120]. upper [2813, 101, 737, 3439, 1996, 1326, 2285, 1572, 1137, 2558, 2314, 3043, 2528, 3092, 1974, 1059]. upriver [1484, 1851]. Upstream [2204, 1514, 2426, 2688, 1905]. upsurge [696]. uptake [1641, 2070, 2557, 931, 2684, 1668, 1017, 562, 2437, 88, 118, 1791, 3083]. Uranoscopidae [2650]. urban [1415, 2776, 2680]. urinary [2356]. Urine [220, 368, 570]. Urobatis [2931, 1263, 2112]. urolophid [2837]. Urolophus [2837]. urophthalmus [3178]. Uruguay [3287, 3015, 1034]. US$48.00 [1495]. USA [1824]. Use [936, 1183, 2261, 382, 2588, 1408, 1719, 388, 1345, 1945, 1964, 412, 171, 2757, 2905, 558, 890, 539, 457, 1552, 3085, 1948, 816, 1332, 1807, 2234, 718, 2929, 3190, 2579, 686, 604, 2233, 307, 1422, 2093, 255, 652, 1191, 2660, 3293, 1588, 3131, 1550, 1427, 2069, 85, 1436, 1566, 2664, 3134, 1786, 2286, 531, 732, 323, 1190, 2229, 1897, 1635, 3101, 176, 141, 2943, 1548, 2239, 2490, 1313, 2962, 348, 1057, 2344, 609, 1849, 1236, 2238]. used [171, 2294, 213, 3313, 1586]. useful [179, 2206]. Using [2980, 651, 215, 1144, 2795, 2268, 2064, 453, 545, 689, 2739, 672, 2404, 3255, 3229, 94, 1947, 1127, 1116, 1869, 3315, 2125, 2725, 2996, 840, 498, 3157, 2572, 931, 2818, 3110, 2740, 898, 1771, 2031, 1209, 3293, 1002, 641, 1665, 2644, 2004, 2718, 734, 2140, 1770, 3384, 1568, 585, 3008, 275, 1518, 3245, 1678, 1721, 2114, 472, 1597, 2331, 1354, 1158, 3319, 2288, 3115, 3400, 3194, 2969, 1787, 2224]. Utah [1447]. utility [1754]. Utilization [1000, 299, 2942, 2168, 2008, 3150, 3364, 20]. utilizes [3216]. Utsjoki [2688, 2506]. UV [1914, 1490]. UV-B [1914, 1490].

V [1694, 1404, 3320, 2347, 2893, 1594, 3340, 2172, 1808, 2231, 1653]. vacca [869]. vaccination [472]. Valenciennea [300]. Valenciennes [2597, 2535, 2430, 2161, 2961, 2677]. valid [1447]. Validated [3447, 2933]. validates [2543]. Validation [977, 1383, 261, 2572, 2298, 1899, 2969, 3282, 457, 1116, 1889, 352, 3467, 3004, 3019, 2898, 1125, 1366, 2357, 1101]. validity [2535]. Valley [3192]. Vanneaugobius [2967, 157]. Varden [2094, 2133]. Variability [2991, 1908, 979, 1667, 2493, 3112, 2107, 3234, 908, 2952, 2712, 1246, 2536, 1732, 2619, 2096, 2294, 725, 1250, 538, 2355, 1700, 613, 3223, 2069, 2634, 2024, 1075, 3123, 3331, 857, 1803, 1554, 2047, 624, 2423, 1812, 824, 2707, 2937, 3232, 815, 3176, 946, 2750, 3251, 3099, 3228]. Variable [1604, 2182, 1977, 2693, 1526, 2091, 1274, 1536, 2518]. variables [2396, 1859, 2800, 3151, 2496, 1486, 3138, 2468, 2354, 1542, 2268, 2454]. variance [2227]. variant [472, 471, 2636]. Variation [1084, 764, 3270, 1111, 27, 1248, 3249, 473, 2784, 2817, 2370, 959, 2470, 1477, 9, 1018, 690, 1385, 3107, 2306, 1946, 34, 2055, 1384, 2131, 3036, 2777, 2065, 1038, 3206, 1937, 842, 900, 2790, 3351, 1171, 3437, 2430, 2713, 729, 631, 3333, 1889, 1933, 712, 1843, 1147, 40, 234, 2893, 278, 2353, 139, 2324, 2692, 1251, 3179, 751, 2633, 2327, 1098, 117, 2016, 2037, 1971, 235, 25, 1936, 2348, 70, 3467, 708, 2736, 2547, 112, 59, 2143, 933, 3167, 2585, 3372, 887, 1070, 863, 1953, 1509, 154, 2968, 2961, 374, 2140, 2927, 2495, 3391]. variation [2726, 1940, 1938, 865, 3140, 2259, 711, 1399, 485, 2067, 2541, 1370, 2524, 361, 581, 1425, 3448, 3284, 96, 955, 2001, 848, 731, 2720, 1796, 3022, 1621, 878, 521, 2522, 1996, 2501, 1228, 1170, 2539, 2331, 1695, 2875, 2540, 1470, 422, 1475, 1286, 137, 1011, 1929, 2560, 892, 1747, 3353, 1442, 2564, 1269, 1782, 3419, 2990, 421]. Variations [1315, 2455, 29, 2453, 814, 2994, 2387, 1948, 1015, 2638, 2217, 3028, 1312, 3100, 899, 254, 467, 574, 780, 1064, 2854, 1780]. varies [2932, 3332]. varieties [2673]. variety [3174]. variolatum [2487]. various [2816, 1813, 2288]. varius [3458]. vary [8]. varying [404, 817, 1881]. vascular [761, 3393]. vasoactive [648]. vasotocin [1386]. vaterite [1161].

Vaz [1843]. Vaz-Ferreira [1843]. Vecht [1119]. vegetable [2190]. vegetated [1672]. vegetation [373, 558, 2551, 1551]. vehicle [1502]. vein [2572]. velifera [2099]. velocities [3137]. velocity [1825, 587, 1466, 2604]. velvet [2995]. Vembanad [2201]. vendace [569, 2450]. venenosa [2290]. Venezuela [1083, 2709]. Venice [1953]. ventilation [152, 3127, 379, 45]. ventralis [3398]. ventricular [447]. venus [647]. Verde [2961]. Verhandlungen [1009]. Verification [2740, 1718, 994, 1846]. verified [3468]. verifies [2528]. Verlag [1497]. Vermont [3192]. vertebrae [1138, 2001]. vertebral [1310, 1283]. vertebrate [3101]. vertebrates [2296, 3146, 1263]. Vertical [241, 1685, 859, 2960, 1074, 416, 3098]. vesicle [671]. vevio [738]. VI [1694]. via [2485]. viability [671]. viarius [1738]. vicinity [2118]. Victoria [2981, 3060, 2324, 696, 2473]. victoriae [783]. video [1247, 1822, 2230]. VIE [639]. Vietnam [542, 491]. view [46]. vii [1406]. viii [1496, 1773]. villosus [1815, 2295]. Vimba [3432]. Vinciguerria [1021]. Vindelälven [1602]. violacea [2624]. virens [1658]. virginalis [488]. viride [3448]. viridis [2966, 2796, 81]. virulence [472]. viscerocranial [326]. viscoelasticity [2888]. visible [639, 1100, 2674, 2376, 2012]. Vision [1442, 2488]. visit [3285]. visiting [2932]. Vistula [387]. Visual [1260, 400, 561, 1349, 1157, 773, 2840, 3181, 2466, 3023, 2160, 1909, 1166, 1442]. visual-census [3023]. visually [2063]. vitamin [658, 1310]. vitelline [150]. vitellogenesis [1062, 2366]. vitellogenin [1392, 348]. vitreus [1871, 3343, 3171, 3155, 1870]. vitro [2203, 99, 1645, 1555, 1681, 33, 2617, 2332, 1045, 626, 279, 810, 2379]. vivipara [2692, 1123]. viviparous [858]. viviparus [2914]. vivo [3120, 33, 2617]. vivoliana [662]. Voer [1849]. Vogel [1403]. vogeliana [1463]. vogelii [1463]. Vol [1497, 1406, 1404, 1652]. Volarja [234]. volitional [3332]. Volta [3092]. voltages [171]. Volume [1032, 1541, 2310, 2333, 2549, 2742, 2515, 2704, 3086, 3264, 535, 3482, 1751, 1903, 2149, 2102, 2311, 2334, 2514, 2548, 2741, 3087, 3265, 3483, 2703, 2993, 1081, 3393]. Volusia [3010]. vulgaris [378]. Vulnerability [1978, 609]. vulnerable [3021]. vulpes [2996]. vulpinus [2174].

W [3479, 1496, 1654, 1651, 3407]. W.I. [1855]. Waikato [3236]. Walbaum [2443, 2812, 2901, 2942, 2623, 2705, 2881, 2686, 2869, 2765, 2320, 2615]. Wales [796]. Walla [1462, 1462]. walleye [729, 3343, 2610, 3171, 1746, 845, 1527, 2687, 3278, 3155, 2422, 90, 2260, 878, 2489, 2, 492]. Wallingford [1654]. walvisensis [2858]. wardi [2851]. warm [2876, 2640, 546]. warm-temperate [2876]. warming [642]. Warta [144, 1316]. was [330]. wastage [1740]. Water [2601, 2356, 803, 1529, 540, 2092, 932, 906, 1183, 2642, 2558, 2439, 1524, 128, 187, 1072, 1825, 286, 189, 1525, 1069, 2302, 1461, 3230, 546, 2438, 2093, 264, 2920, 194, 2633, 3249, 587, 861, 1614, 850, 1447, 1530, 3443, 3216, 1150, 3137, 2826, 2366, 2687, 329, 1502, 1273, 1357, 1953, 1104, 885, 3268, 1835, 2232, 3285, 2965, 1296, 476, 8, 1791, 82, 948, 1313, 2375, 2, 2209, 1097, 2844, 2036, 2488, 2231, 1834, 782, 1266, 3055, 367].

Water-based [1529]. water-management [3249]. water-surface [1834]. waterbodies [74, 2416, 2417]. waterborne [1641, 3456, 2917, 2413]. watercourse [2032]. waterfalls [1857]. watering [1958]. waters [2800, 2424, 2139, 664, 578, 1430, 839, 901, 1412, 619, 3062, 1810, 2062, 2998, 2105, 2025, 241, 3029, 1464, 823, 994, 3334, 2950, 2746, 2485, 2210, 1886, 1021, 1372, 2529, 609, 1092, 2244, 2088, 3390, 889]. Watershed [1653, 2841, 2473, 1974, 1145, 2637]. waterway [2680]. waterways [2776]. wave [936]. waveforms [171]. way [2608]. Weak [2738, 1154, 488, 2978]. weaken [1587]. weakfish [3347, 3312]. weakly [1648, 3298]. weaned [2502]. weasel [2839]. weather [2717]. web [2181]. Weberian [2676]. weight [385]. weighted [423]. weir [2426]. weirs [2777]. Welfare [3478, 3474, 3444, 805, 3477, 1992, 3475, 3473]. well [368]. Welsh [2122]. West [604, 354, 605, 3234, 1398, 3073, 2934, 3243, 2641, 2756, 796, 465, 623, 193, 2657, 3291, 2634, 2784, 703, 1468, 1719, 599, 2708, 3199, 1784, 2882, 2320, 917, 2152, 1819, 1178, 2500, 228, 1656, 2220, 1886, 1469, 1057, 921, 2685, 2345, 641, 2779, 2156, 1802, 512]. Western [3119, 2106, 3185, 2526, 620, 2092, 2424, 3141, 1417, 2532, 3386, 1058, 1733, 2275, 106, 2836, 2620, 2724, 2719, 3289, 1486, 2858, 2217, 3226, 1783, 1520, 2650, 3375, 831, 1985, 775, 2578, 3245, 2264, 3361, 2669, 2544, 2819, 603, 2268, 3470, 1374, 2054, 2529, 3419, 3228, 2535, 1686, 294, 2231]. Westin [2103]. westslope [598]. Westward [2372]. wet [2231]. wetland [176]. wetlands [2588, 2747, 3178]. whale [3141, 1471, 2880]. whaler [2672]. whales [204]. Wheeler [1965, 1628]. Which [2608]. while [204, 1675]. whiskered [1314]. white [2247, 539, 2504, 1054, 3160, 1725, 14, 3102, 3401, 987, 1486, 3122, 1055, 536, 579, 865, 1740, 2393, 130, 2468, 318, 2133, 2958, 2839]. white-spotted [3122, 536, 579, 865, 1740, 2133, 2958]. whitefin [245]. whitefish [312, 3206, 202, 1102, 1969, 992, 1144, 3089, 138, 2604, 1490]. whitei [1078]. whitemouth [1345, 3287]. whitespotted [2864, 1853]. whiting [954, 1344, 392, 538, 241, 3103, 1380, 169, 809, 2264, 3014, 566, 1389, 2196]. Whitley [2407, 2851, 2671]. who [1270]. whole [1276, 729, 3311, 2437, 1001, 781, 2313]. whole-body [3311, 781]. whole-mount [2313]. wide [3000]. wide-ranging [3000]. widely [2793, 1600]. Widespread [1839, 3118]. width [1766, 2957]. Wiesbaden [1497]. Wild [2895, 600, 1172, 424, 412, 188, 1320, 1577, 1821, 1891, 1633, 1937, 3255, 855, 2642, 2457, 3254, 2897, 2599, 1559, 1581, 2812, 1608, 3389, 204, 1793, 868, 1584, 399, 93, 1634, 1585, 1610, 1611, 2034, 2115, 519, 1562, 1563, 2682, 391, 584, 2810, 1997, 1588, 1589, 1434, 3223, 2083, 1690, 2822, 58, 1617, 2143, 1857, 1590, 1593, 1594, 1565, 1566, 116, 3474, 1618, 3067, 1143, 1569, 1570, 1595, 1103, 2963, 1450, 980, 976, 1762, 2090, 1393, 1839, 1934, 879, 2269, 1265, 2574, 1124, 1598, 1910, 826, 2205, 1601, 1289, 1078, 1747, 3404]. wild [2969, 1806, 2958]. wild-capture [3474]. wild-caught [1608]. wild-reared [1610]. William [1811, 1405, 645]. Windermere [1553, 3253]. windowpane [363]. Winter [539, 3278, 2884, 3191, 1726, 230, 99, 1645, 2443, 1332, 1430, 1609, 325, 1191, 2577, 1376, 1019, 1278, 1719, 2869, 745, 990, 2542, 1391, 1569, 1145, 1510, 83, 287, 2, 2962, 3200, 507]. wintering [412, 652, 1459, 1190]. winterkill [2021]. wire [2251, 1900]. wireless [152]. Wisconsin [46]. Wislok [1606]. Within [2286, 581, 788, 3367, 262, 1111, 2641, 1933, 1203, 2496, 2658, 2517, 3348, 2099, 588, 1550, 2508, 3372, 2397, 2726, 731, 1897, 8, 603, 690, 1379, 1470, 422, 2322, 2596, 1385, 922, 2519, 2379]. Within-population [2286]. within-species [603]. Within-stream [581]. without [1907, 1527, 1566, 1046, 1466]. wobbegong [3017]. wolf [68]. wolffish [571, 470, 2062, 2998, 2058, 895, 313, 451]. wolffishes [2791]. Woo [1654]. Workshop [2100, 1901]. World [1404, 3478, 2827, 1403]. worldwide [1073]. worm [986]. worming [3161]. wound [362]. wounds [2935]. wrasse [3458, 67, 2120, 1045, 430, 1985, 1158, 137]. wrasses [1639]. Wright [1902, 1963, 1990, 2019, 2049, 2078, 2101, 2127, 1775].

x [1497, 1654, 1060]. X-cell [1060]. xanthocephalus [2898]. xanthurus [1832]. Xenomystus [2644]. Xenurolebias [2619]. xi [1774]. XII [2255, 1403, 1497, 2255]. xiii [1652]. Xingu [3454]. Xiphias [2358, 1763, 1816]. Xiphophorus [2782]. xiv [1494]. xv [1493]. xx [1402, 1772]. XXVII [456, 477]. Xyrauchen [2946].

Yangtze [2238]. year [3234, 1672, 3191, 2905, 944, 2693, 3069, 729, 1243, 1013, 2122, 1850, 1793, 963, 2610, 2561, 2446, 1419, 2484, 3189, 2469, 1785, 2037, 1426, 756, 70, 708, 236, 949, 1114, 1327, 2119, 2869, 2416, 2417, 2455, 2884, 741, 3253, 1361, 1708, 1677, 2567, 1679, 3098, 853, 624, 1440, 948, 1313, 1551, 807, 1455, 1064, 145, 3197, 1556, 492, 3346]. year-group [2484]. year-old [944, 2561]. year-round [3197]. year-to-year [3234]. years [2640, 3253, 2477, 1929]. Yearsley [2837]. Yellow [1221, 2637, 2739, 957, 2780, 839, 1715, 303, 1350, 2828, 446, 146, 3461, 1440, 1263, 2489, 1638, 2982, 1782, 3257, 2291, 1025, 1782]. yellow-fin [2780]. Yellow-stage [2637]. yellowfin [2290, 1233, 3194, 3040]. yellownose [2002]. Yellowstone [577]. yellowtail [858, 426, 1507, 891, 319]. yield [64]. yolk [1023, 2832, 429, 1081, 693, 2056, 2427, 2491]. yolk-feeding [2056]. yolk-sac [2832, 2427]. yolk-sac-fry [429]. York [2077, 3200]. young [2739, 3191, 3069, 1243, 1013, 963, 3075, 2363, 303, 623, 1426, 70, 1734, 2024, 1114, 1327, 2119, 2869, 475, 741, 142, 1361, 1708, 1677, 1679, 7, 388, 3098, 853, 624, 123, 807, 145]. young-of-the-year [3069, 1013, 963, 1426, 70, 1114, 1327, 2119, 2869, 741, 1361, 1708, 1677, 1679, 3098, 853, 624, 807, 145]. young-of-year [3191, 1243]. younger [1599]. Yurii [2382]. YY [677].

Zalewski [1653]. Zambezi [1059]. zanandreai [3437]. zander [578, 648]. Zanzibar [2284]. Zealand [107, 373, 558, 890, 2233, 340, 1135, 2772, 3236, 1743, 608, 823, 2137, 2530, 1827, 1630, 378, 881]. zebra [1780, 3012]. zebrafish [3120, 1384, 2980, 2892, 3218, 1463, 93, 3033, 3217, 1355, 1076, 1698, 1081, 3378, 996, 2193, 1315, 2491, 2208, 2267, 2574, 768, 2518, 2155, 2305, 1269, 2859].

Zeidae [2697]. Zenopsis [2697]. Zimbabwe [290]. zinc [3456, 2691, 173, 344]. Zingel [1086, 3316, 250]. Zoarces [2826]. zona [51]. zone [1908, 1534, 3240, 1243, 817, 803, 2222, 1371, 3216, 2024, 1104, 81, 343, 2844]. zones [2339, 2999, 5]. zoogeographic [92]. zooplanktivore [2114]. zooplanktivores [2181]. zooplanktivorous [2263]. zooplankton [170, 3328, 993, 655, 2403, 2421, 2801, 1313, 879]. Zosterisessor [2394, 3031, 1254]. Zuger [2223]. ZW [3435, 3296]. ZZ [3435, 3296]. ZZ/ZW [3435, 3296].

References

Anonymous:1992:LCS

[1] Anonymous. A list of the common and scientific names of fishes of the British Isles. Journal of Fish Biology, 41(sB):17–26, September 1992. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See updates [1628].

Sogard:2000:ESW

[2] S. M. Sogard and B. L. Olla. Endurance of simulated winter conditions by age-0 walleye pollock: effects of body size, water temperature and energy stores. Journal of Fish Biology, 56(1):1–21, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richtarski:2000:CMM

[3] U. Richtarski and R. A. Patzner. Comparative morphology of male reproductive systems in Mediterranean blennies (Blenniidae). Journal of Fish Biology, 56(1):22–36, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luksiene:2000:ETE

[4] D. Lukšienė, O. Sandström, L. Lounasheimo, and J. Andersson. The effects of thermal effluent exposure on the gametogenesis of female fish. Journal of Fish Biology, 56(1):37–50, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horppila:2000:SCD

[5] J. Horppila, J. Ruuhijärvi, M. Rask, C. Karppinen, K. Nyberg, and M. Olin. Seasonal changes in the diets and relative abundances of perch and roach in the littoral and pelagic zones of a large lake. Journal of Fish Biology, 56(1):51–72, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holcik:2000:NDU

[6] J. Holčík and A. Delić. New discovery of the Ukrainian brook lamprey in Croatia. Journal of Fish Biology, 56(1):73–86, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ponton:2000:CMD

[7] D. Ponton and S. Mérigoux. Comparative morphology and diet of young cichlids in the dammed Sinnamary river, French Guiana, South America. Journal of Fish Biology, 56(1):87–102, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sarre:2000:DDC

[8] G. A. Sarre, M. E. Platell, and I. C. Potter. Do the dietary compositions of Acanthopagrus butcheri in four estuaries and a coastal lake vary with body size and season and within and amongst these water bodies? Journal of Fish Biology, 56(1):103–122, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulin:2000:VIR

[9] R. Poulin. Variation in the intraspecific relationship between fish length and intensity of parasitic infection: biological and statistical causes. Journal of Fish Biology, 56(1):123–137, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slotte:2000:SDS

[10] A. Slotte and Ø. Fiksen. State-dependent spawning migration in Norwegian spring-spawning herring. Journal of Fish Biology, 56(1):138–162, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munakata:2000:PRA

[11] A. Munakata, B. Th. Björnsson, E. Jönsson, M. Amano, K. Ikuta, S. Kitamura, T. Kurokawa, and K. Aida. Post-release adaptation processes of hatchery-reared honmasu salmon parr. Journal of Fish Biology, 56 (1):163–172, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farag:2000:DMD

[12] A. M. Farag, M. J. Suedkamp, J. S. Meyer, R. Barrows, and D. F. Woodward. Distribution of metals during digestion by cutthroat trout fed benthic invertebrates contaminated in the Clark Fork River, Montana and the Coeur d’Alene River, Idaho, U.S.A., and fed artificially contaminated Artemia. Journal of Fish Biology, 56(1):173–190, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2000:IMG

[13] E. D. Stevens and R. H. Devlin. Intestinal morphology in growth hormone transgenic Coho salmon. Journal of Fish Biology, 56(1):191–195, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardewig:2000:TCE

[14] I. Hardewig, P. L. M. Van Dijk, S. C. Leary, and C. D. Moyes. Temporal changes in enzyme activity and mRNA levels during thermal challenge in white sucker. Journal of Fish Biology, 56(1):196–207, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2000:EPS

[15] J. M. Elliott, M. A. Hurley, and S. C. Maberly. The emergence period of sea trout fry in a Lake District stream correlates with the North Atlantic Oscillation. Journal of Fish Biology, 56(1):208–210, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2000:ASC

[16] M. S. Johnson, T. D. Waybright, D. W. Matt, J. J. Feher, and M. L. Fine. Absence of a seasonal cycle in the sonic neuromuscular system of the oyster toadfish. Journal of Fish Biology, 56(1):211–215, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ronnestad:2000:RPP

[17] I. Rønnestad, C. R. Rojas-Garcia, and J. Skadal. Retrograde peristalsis; a possible mechanism for filling the pyloric caeca? Journal of Fish Biology, 56(1):216–218, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamilton:2000:EST

[18] L. C. Hamilton, G. R. Macpherson, and J. M. Wright. Expressed sequence tags derived from brain tissue of Oreochromis niloticus. Journal of Fish Biology, 56(1):219–222, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aprahamian:2000:GRE

[19] M. W. Aprahamian. The growth rate of eel in tributaries of the lower River Severn, England, and its relationship with stock size. Journal of Fish Biology, 56(1):223–227, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qian:2000:CGF

[20] X. Qian, Y. Cui, B. Xiong, and Y. Yang. Compensatory growth, feed utilization and activity in gibel carp, following feed deprivation. Journal of Fish Biology, 56(1):228–232, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRa

[21] Anonymous. Book reviews. Journal of Fish Biology, 56(1):233–236, January 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Becher:2000:GFT

[22] S. A. Becher and A. E. Magurran. Gene flow in Trinidadian guppies. Journal of Fish Biology, 56(2):241–249, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koponen:2000:SCI

[23] K. Koponen and M. S. Myers. Seasonal changes in intra- and interorgan occurrence of rodlet cells in freshwater bream. Journal of Fish Biology, 56 (2):250–263, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bromley:2000:IFR

[24] P. J. Bromley, C. Ravier, and P. R. Witthames. The influence of feeding regime on sexual maturation, fecundity and atresia in first-time spawning turbot. Journal of Fish Biology, 56(2):264–278, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonassen:2000:GVG

[25] T. M. Jonassen, A. K. Imsland, R. Fitzgerald, S. W. Bonga, E. V. Ham, G. Nævdal, M. O. Stefánsson, and S. O. Stefansson. Geographic variation in growth and food conversion efficiency of juvenile Atlantic halibut related to latitude. Journal of Fish Biology, 56(2):279–294, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slotte:2000:EFS

[26] A. Slotte, A. Johannessen, and O. S. Kjesbu. Effects of fish size on spawning time in Norwegian spring-spawning herring. Journal of Fish Biology, 56(2):295–310, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chipps:2000:VRM

[27] S. R. Chipps, D. F. Clapp, and D. H. Wahl. Variation in routine metabolism of juvenile muskellunge: evidence for seasonal metabolic compensation in fishes. Journal of Fish Biology, 56(2):311–318, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pawson:2000:ITO

[28] M. G. Pawson, G. D. Pickett, and P. R. Witthames. The influence of temperature on the onset of first maturity in sea bass. Journal of Fish Biology, 56(2):319–327, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2000:VJG

[29] A. H. Rikardsen and J. M. Elliott. Variations in juvenile growth, energy allocation and life-history strategies of two populations of Arctic charr in North Norway. Journal of Fish Biology, 56(2):328–346, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claireaux:2000:SCH

[30] G. Claireaux and C. Audet. Seasonal changes in the hypo-osmoregulatory ability of brook charr: the role of environmental factors. Journal of Fish Biology, 56(2):347–373, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Power:2000:EFD

[31] D. M. Power, J. Melo, and C. R. A. Santos. The effect of food deprivation and refeeding on the liver, thyroid hormones and transthyretin in sea bream. Journal of Fish Biology, 56(2):374–387, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2000:CLG

[32] Q. Zhang, R. K. Cooper, and T. R. Tiersch. Chromosomal location of the 28S ribosomal RNA gene of channel catfish by in situ polymerase chain reaction. Journal of Fish Biology, 56(2):388–397, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Currie:2000:EHS

[33] S. Currie, C. D. Moyes, and B. L. Tufts. The effects of heat shock and acclimation temperature on hsp70 and hsp30 mRNA expression in rainbow trout: in vivo and in vitro comparisons. Journal of Fish Biology, 56(2):398–408, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnason:2000:MCD

[34] E. Árnason, P. H. Petersen, K. Kristinsson, H. Sigurgíslason, and S. Pálsson. Mitochondrial cytochrome b DNA sequence variation of Atlantic cod from Iceland and Greenland. Journal of Fish Biology, 56(2): 409–430, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruzzante:2000:BSP

[35] D. E. Ruzzante, J. S. Wroblewski, C. T. Taggart, R. K. Smedbol, D. Cook, and S. V. Goddaard. Bay-scale population structure in coastal Atlantic cod in Labrador and Newfoundland, Canada. Journal of Fish Biology, 56(2):431–447, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pfeiler:2000:GRA

[36] E. Pfeiler, D. Padrón, and R. E. Crabtree. Growth rate, age and size of bonefish from the Gulf of California. Journal of Fish Biology, 56(2): 448–453, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRb

[37] Anonymous. Book reviews. Journal of Fish Biology, 56(2):454–457, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRc

[38] Anonymous. Book received. Journal of Fish Biology, 56(2):457, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Aa

[39] Anonymous. Announcements. Journal of Fish Biology, 56(2):458–459, February 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeWoody:2000:MVM

[40] J. A. DeWoody and J. C. Avise. Microsatellite variation in marine, freshwater and anadromous fishes compared with other animals. Journal of Fish Biology, 56(3):461–473, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mattiangelli:2000:AAG

[41] V. Mattiangelli, E. A. Bourke, A. W. Ryan, J. Mork, and T. F. Cross. Allozyme analyses of the genus Trisopterus: taxonomic status and population structure of the poor cod. Journal of Fish Biology, 56(3):474–494, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciereszko:2000:FLS

[42] A. Ciereszko, J. Glogowski, and K. Dabrowski. Fertilization in landlocked sea lamprey: storage of gametes, optimal sperm: egg ratio, and methods of assessing fertilization success. Journal of Fish Biology, 56(3):495– 505, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadler:2000:PSR

[43] J. Sadler, N. W. Pankhurst, P. M. Pankhurst, and H. King. Physiological stress responses to confinement in diploid and triploid Atlantic salmon. Journal of Fish Biology, 56(3):506–518, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matallanas:2000:DCP

[44] J. Matallanas and G. Pequeño. Description of Careproctus patagonicus sp. nov. and C. magellanicus sp. nov. (Pisces: Scorpaeniformes) from the lower slope of Drake Passage. Journal of Fish Biology, 56(3):519– 527, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soncini:2000:OAB

[45] R. Soncini and M. L. Glass. Oxygen and acid-base status related drives to gill ventilation in carp. Journal of Fish Biology, 56(3):528–541, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillooly:2000:LTV

[46] J. F. Gillooly, T. C. O’Keefe, S. P. Newman, and J. R. Baylis. A longterm view of density-dependent recruitment in smallmouth bass from Nebish Lake, Wisconsin. Journal of Fish Biology, 56(3):542–551, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richardson:2000:ADM

[47] A. J. Richardson, G. Maharaj, L. J. V. Compagno, R. W. Leslie, D. A. Ebert, and M. J. Gibbons. Abundance, distribution, morphometrics, reproduction and diet of the Izak catshark. Journal of Fish Biology, 56 (3):552–576, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Overnell:2000:SEA

[48] J. Overnell and R. S. Batty. Scaling of enzyme activity in larval herring and plaice: effects of temperature and individual growth rate on aerobic and anaerobic capacity. Journal of Fish Biology, 56(3):577–589, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2000:GSL

[49] P. J. Wilson, C. M. Wood, J. N. Maina, and B. N. White. Genetic structure of Lake Magadi tilapia populations. Journal of Fish Biology, 56(3):590–603, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tinti:2000:MSA

[50] F. Tinti and C. Piccinetti. Molecular systematics of the Atlanto– Mediterranean Solea species. Journal of Fish Biology, 56(3):604–614, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2000:CUS

[51] Y. H. Li, C. C. Wu, and J. S. Yang. Comparative ultrastructural studies of the zona radiata of marine fish eggs in three genera in Perciformes. Journal of Fish Biology, 56(3):615–621, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waters:2000:HBF

[52] J. M. Waters, J. M. Epifanio, T. Gunter, and B. L. Brown. Homing behaviour facilitates subtle genetic differentiation among river populations of Alosa sapidissima: microsatellites and mtDNA. Journal of Fish Biology, 56(3):622–636, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2000:LHS

[53] I. J. Morgan, I. D. McCarthy, and N. B. Metcalfe. Life-history strategies and protein metabolism in overwintering juvenile Atlantic salmon: growth is enhanced in early migrants through lower protein turnover. Journal of Fish Biology, 56(3):637–647, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ali:2000:PHI

[54] M. Ali and R. J. Wootton. Pattern of hyperphagia in immature threespined sticklebacks after short-term food deprivation. Journal of Fish Biology, 56(3):648–653, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dou:2000:FBJ

[55] S. Dou, T. Seikai, and K. Tsukamoto. Feeding behaviour of Japanese flounder larvae under laboratory conditions. Journal of Fish Biology, 56 (3):654–666, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:2000:CSM

[56] T. G. Pottinger and T. R. Carrucj. Contrasting seasonal modulation of the stress response in male and female rainbow trout. Journal of Fish Biology, 56(3):667–675, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amundsen:2000:FCA

[57] P.-A. Amundsen, H.-M. Gabler, T. Herfindal, and L. S. Riise. Feeding chronology of Atlantic salmon parr in subarctic rivers: consistency of nocturnal feeding. Journal of Fish Biology, 56(3):676–686, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kristiansen:2000:SDM

[58] T. S. Kristiansen, H. Otterå, and T. Svåsand. Size-dependent mortality of juvenile Atlantic cod, estimated from recaptures of released reared cod and tagged wild cod. Journal of Fish Biology, 56(3):687–712, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krieg:2000:MDV

[59] F. Krieg, A. Triantafyllidis, and R. Guyomard. Mitochondrial DNA variation in European populations of Silurus glanis. Journal of Fish Biology, 56(3):713–724, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloman:2000:DSI

[60] K. A. Sloman, K. M. Gilmour, N. B. Metcalfe, and A. C. Taylor. Does socially induced stress in rainbow trout cause chloride cell proliferation? Journal of Fish Biology, 56(3):725–738, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jirotkul:2000:OSR

[61] M. Jirotkul. Operational sex ratio influences the opportunity for sexual selection in guppies. Journal of Fish Biology, 56(3):739–741, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ab

[62] Anonymous. Announcements. Journal of Fish Biology, 56(3):742, March 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowartz:2000:NPS

[63] S. M. Lowartz and F. W. H. Beamish. Novel perspectives in sexual lability through gonadal biopsy in larval sea lampreys. Journal of Fish Biology, 56(4):743–757, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Froese:2000:ERE

[64] R. Froese and C. Binohlan. Empirical relationships to estimate asymptotic length, length at first maturity and length at maximum yield per recruit in fishes, with a simple method to evaluate length frequency data. Journal of Fish Biology, 56(4):758–773, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townsend:2000:BDH

[65] K. A. Townsend and I. R. Tibbetts. Biomass and distribution of herbivorous blennies in the southern Great Barrier Reef. Journal of Fish Biology, 56(4):774–791, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kristianen:2000:SDM

[66] T. S. Kristianen, H. Otterå, and T. Svåsand. Size–dependent mortality of juvenile reared Atlantic cod released in a small fjord. Journal of Fish Biology, 56(4):792–801, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Francini–Filho:2000:CWT

[67] R. B. Francini-Filho, R. L. Moura, and I. Sazima. Cleaning by the wrasse Thalassoma noronhanum, with two records of predation by its grouper client Cephalopholis fulva. Journal of Fish Biology, 56(4):802–809, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liao:2000:GDM

[68] Y.-Y. Liao and M. C. Lucas. Growth, diet and metabolism of common wolf–fish in the North Sea, a fast–growing population. Journal of Fish Biology, 56(4):810–825, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stefansson:2000:EDI

[69] M. Ö. Stefánsson, A. K. Imsland, M. D. Jenssen, T. M. Jonassen, S. O. Stefansson, and R. FitzGerald. The effect of different initial size distributions on the growth of Atlantic halibut. Journal of Fish Biology, 56(4):826–836, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karjalainen:2000:UFR

[70] J. Karjalainen, H. Auvinen, H. Helminen, T. J. Arjomäki, T. Niva, J. Sarvala, and M. Vilhanen. Unpredictability of fish recruitment: interannual variation in young-of-the-year abundance. Journal of Fish Biology, 56 (4):837–857, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goodrich:2000:BAS

[71] J. S. Goodrich, S. L. Sanderson, I. E. Batjakas, and L. S. Kaufman. Branchial arches of suspension-feeding Oreochromis esculentus: sieve or sticky filter? Journal of Fish Biology, 56(4):858–875, April 2000. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romare:2000:GLJ

[72] P. Romare. Growth of larval and juvenile perch: the importance of diet and fish density. Journal of Fish Biology, 56(4):876–889, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edo:2000:SDR

[73] K. Edo, H. Kawamula, and S. Higashi. The structure and dimensions of redds and egg pockets of the endangered salmonid, Sakhalin taimen. Journal of Fish Biology, 56(4):890–904, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halls:2000:DDR

[74] A. S. Halls, K. Debnath, G. P. Kirkwood, and A. I. Payne. Densitydependent recruitment of Puntius sophore in floodplain waterbodies in Bangladesh. Journal of Fish Biology, 56(4):905–914, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahti:2000:ESA

[75] K. Lahti and N. Lower. Effects of size asymmetry on aggression and food acquisition in Arctic charr. Journal of Fish Biology, 56(4):915–922, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vehanen:2000:EFF

[76] T. Vehanen, P. L. Bjerke, J. Heggenes, A. Huusko, and A. Mäki-Petäys. Effect of fluctuating flow and temperature on cover type selection and behaviour by juvenile brown trout in artificial flumes. Journal of Fish Biology, 56(4):923–937, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2000:PRB

[77] J. M. Elliott. Pools as refugia for brown trout during two summer droughts: trout responses to thermal and oxygen stress. Journal of Fish Biology, 56(4):938–948, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berrebi:2000:SIA

[78] P. Berrebi, C. Poteaux, M. Fissier, and G. Cattaneo-Berrebi. Stocking impact and allozyme diversity in brown trout from Mediterranean southern France. Journal of Fish Biology, 56(4):949–960, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeCrespindeBilly:2000:BPD

[79] V. De Crespin de Billy, S. Dolegec, and D. Chessel. Biplot presentation of diet composition data: an alternative for fish stomach contents analysis. Journal of Fish Biology, 56(4):961–973, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niemela:2000:ASR

[80] E. Niemelä, T. S. Makinen, K. Moen, E. Hassinen, J. Erkinaro, M. Lansman, and M. Julkunen. Age, sex ratio and timing of the catch of kelts and ascending Atlantic salmon in the subarctic River Teno. Journal of Fish Biology, 56(4):974–985, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz:2000:OCD

[81] A. A. Muñoz and F. P. Ojeda. Ontogenetic changes in the diet of the herbivorous Scartichthys viridis in a rocky intertidal zone in central Chile. Journal of Fish Biology, 56(4):986–998, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sivakumar:2000:CCS

[82] P. Sivakumar, R. Arichandran, L. Suguna, M. Mariappan, and G. Chandrakasan. The composition and characteristics of skin and muscle collagens from a freshwater catfish grown in biologically treated tannery effuent water. Journal of Fish Biology, 56(4):999–1012, April 2000. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruggerone:2000:DSJ

[83] G. T. Ruggerone. Differential survival of juvenile sockeye and Coho salmon exposed to low dissolved oxygen during winter. Journal of Fish Biology, 56(4):1013–1016, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hove:2000:MBT

[84] J. R. Hove, M. S. Gordon, P. W. Webb, and D. Weihs. A modified Bläzkatype respirometer for the study of swimming metabolism in fishes having deep, laterally compressed bodies or unusual locomotor modes. Journal of Fish Biology, 56(4):1017–1022, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krause:2000:RUS

[85] J. Krause, P. Longworth, and G. D. Ruxton. Refuge use in sticklebacks as a function of body length and group size. Journal of Fish Biology, 56 (4):1023–1027, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ac

[86] Anonymous. Announcements. Journal of Fish Biology, 56(4):1028, April 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolker:2000:PDH

[87] J. A. Bolker and C. R. Hill. Pigmentation development in hatcheryreared flatfishes. Journal of Fish Biology, 56(5):1029–1052, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mcintosh:2000:LUR

[88] D. Mcintosh, B. Austin, E. Fla, A. Villena, J. A. Matinez-Pereda, and J. Tarazona. Lack of uptake of Renibacterium salmoninarum by gill epithelia of rainbow trout. Journal of Fish Biology, 56(5):1053–1061, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fish:2000:RPF

[89] F. E. Fish and L. D. Shannahan. The role of the pectoral fins in body trim of sharks. Journal of Fish Biology, 56(5):1062–1073, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OReilly:2000:ITL

[90] P. T. O’Reilly, M. F. Canino, K. M. Bailey, and P. Bentzen. Isolation of twenty low stutter di- and tetranucleotide microsatellites for population analyses of walleye pollock and other gadoids. Journal of Fish Biology, 56(5):1074–1086, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chow:2000:NMD

[91] S. Chow and H. Takeyama. Nuclear and mitochondrial DNA analyses reveal four genetically separated breeding units of the swordfish. Journal of Fish Biology, 56(5):1087–1098, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Floeter:2000:SAR

[92] S. R. Floeter and J. L. Gasparini. The southwestern Atlantic reef fish fauna: composition and zoogeographic patterns. Journal of Fish Biology, 56(5):1099–1114, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreri:2000:SCW

[93] F. Ferreri, C. Nicolais, C. Boglione, and B. Bmertoline. Skeletal characterization of wild and reared zebrafish: anomalies and meristic characters. Journal of Fish Biology, 56(5):1115–1128, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breining:2000:ESS

[94] T. Breining and R. Britz. Egg surface structure of three clingfish species, using scanning electron microscopy. Journal of Fish Biology, 56(5):1129– 1137, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Itzkowitz:2000:TBM

[95] M. Itzkowitz, A. Ludlow, and M. Haley. Territorial boundaries of the male beaugregory damselfish. Journal of Fish Biology, 56(5):1138–1144, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Park:2000:MCS

[96] J. Y. Park, H. J. Lee, W. J. Kim, J. H. Lee, and K. S. Min. Mitochondrial cytochrome b sequence variation in Korean salmonids. Journal of Fish Biology, 56(5):1145–1154, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Asturiano:2000:HRE

[97] J. F. Asturiano, L. A. Sorbera, J. Ramos, D. E. Kime, M. Carrilo, and S. Zanuy. Hormonal regulation of the European sea bass reproductive cycle: an individualized female approach. Journal of Fish Biology, 56 (5):1155–1172, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haugen:2000:ESG

[98] T. O. Haugen. Early survival and growth in populations of grayling with recent common ancestors — field experiments. Journal of Fish Biology, 56(5):1173–1191, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:2000:RVR

[99] D. Burton and J. E. Vokey. The relative in vitro responsiveness of melanophores of winter flounder to α -MSH and MCH. Journal of Fish Biology, 56(5):1192–1200, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cubillos:2000:AFR

[100] E. R. Cubillos and H. E. Guderley. Analysis of the factors related with mate choice and reproductive success in male three-spined sticklebacks. Journal of Fish Biology, 56(5):1201–1216, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DOnghia:2000:PSA

[101] G. D’Onghia, M. Basanisi, and A. Tursi. Population structure, age and growth of macrourid fish from the upper slope of the Eastern-Central Mediterranean. Journal of Fish Biology, 56(5):1217–1238, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Masuda:2000:OCL

[102] R. Masuda and D. A. Ziemann. Ontogenetic changes of learning capability and stress recovery in Pacific threadfin juveniles. Journal of Fish Biology, 56(5):1239–1247, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giroux:2000:RBD

[103] F. Giroux, M. Ovidio, J.-C. Phillippart, and E. Baras. Relationship between the drift of macroinvertebrates and the activity of brown trout in a small stream. Journal of Fish Biology, 56(5):1248–1257, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:2000:SSJ

[104] R. N. Gibson and L. Robb. Sediment selection in juvenile plaice and its behavioural basis. Journal of Fish Biology, 56(5):1258–1275, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penttinen:2000:PIC

[105] O-P. Penttinen, H. Hyväurinen, and J. Piironen. A preliminary investigation to compare metabolic heat dissipation by quiescent and activated spermatozoa of landlocked salmon. Journal of Fish Biology, 56(5):1276– 1279, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Figurroa:2000:SOA

[106] D. E. Figurroa, J. M. Diaz De Astarloa, and M. B. Ccousseau. Southernmost occurrence of the aguavina on the western Atlantic coast of Argentina. Journal of Fish Biology, 56(5):1280–1282, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adcodk:2000:SDP

[107] G. J. Adcodk, J. H. Bernal Ramázrez, L. Hauser, P. Smith, and G. R. Carvalho. Screening of DNA polymorphisms in samples of archived scales from New Zealand snapper. Journal of Fish Biology, 56(5):1283–1287, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRd

[108] Anonymous. Book reviews. Journal of Fish Biology, 56(5):1288–1291, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ad

[109] Anonymous. Announcements. Journal of Fish Biology, 56(5):1292–1293, May 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernandez:2000:MFT

[110] D. A. Fernandez, J. Calvo, C. E. Franklin, and I. A. Johnston. Muscle fibre types and size distribution in sub-Antarctic notothenioid fishes. Journal of Fish Biology, 56(6):1295–1311, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mallet:2000:HPE

[111] J. P. Mallet, N. Lamouroux, P. Sagnes, and H. Persat. Habitat preferences of European grayling in a medium size stream, the Ain river, France. Journal of Fish Biology, 56(6):1312–1326, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kraus:2000:FBC

[112] G. Kraus, A. Müller, K. Trella, and F. W. Köuster. Fecundity of Baltic cod: temporal and spatial variation. Journal of Fish Biology, 56(6):1327– 1341, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ojanguren:2000:TDS

[113] A. F. Ojanguren and F. Brañta. Thermal dependence of swimming endurance in juvenile brown trout. Journal of Fish Biology, 56(6):1342– 1347, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cosson:2000:AMP

[114] J. Cosson, O. Linhart, S. D. Mims, W. L. Shelton, and M. Rodina. Analysis of motility parameters from paddlefish and shovelnose sturgeon spermatozoa. Journal of Fish Biology, 56(6):1348–1367, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arneri:2000:AEL

[115] E. Arneri and B. Morales-Nin. Aspects of the early life history of European hake from the central Adriatic. Journal of Fish Biology, 56(6): 1368–1380, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matthews:2000:EST

[116] I. M. Matthews and A. E. Magurran. Evidence for sperm transfer during sneaky mating in wild Trinidadian guppies. Journal of Fish Biology, 56 (6):1381–1386, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hard:2000:GVM

[117] J. J. Hard, L. Connell, W. K. Hershberger, and L. W. Harrell. Genetic variation in mortality of Chinook salmon during a bloom of the marine alga Heterosigma akashiwo. Journal of Fish Biology, 56(6):1387–1397, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roelants:2000:ISC

[118] I. Roelants, T. Mikolajczyk, P. Epler, F. Ollevier, J. Chyb, and B. Breton. Induction of spermiation in common carp after enhanced intestinal uptake of sGnRH-A and pimozide. Journal of Fish Biology, 56(6):1398– 1407, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Power:2000:TAP

[119] M. Power, M. J. Attrill, and R. M. Thomas. Temporal abundance patterns and growth of juvenile herring and sprat from the Thames estuary 1977–1992. Journal of Fish Biology, 56(6):1408–1426, June 2000. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jirotkul:2000:MTD

[120] M. Jirotkul. Male trait distribution determined alternative mating tactics in guppies. Journal of Fish Biology, 56(6):1427–1434, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temple:2000:SCF

[121] G. K. Temple, J. M. Wakeling, and I. A. Johnston. Seasonal changes in fast-starts in the short-horn sculpin: integration of swimming behaviour and muscle performance. Journal of Fish Biology, 56(6):1435–1449, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birrell:2000:ODD

[122] L. Birrell, G. Cramb, and N. Hazon. Osmoregulation during the development of glass eels and elvers. Journal of Fish Biology, 56(6):1450–1459, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sarpeadonti:2000:DOY

[123] V. Sarpéadonti, D. Ponton, and Chong Ving Ching. Description and ontogeny of young Stolephorus baganensis and Thryssa kammalensis, two Engraulididae from Peninsular Malaysia. Journal of Fish Biology, 56(6):1460–1476, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svensson:2000:SBT

[124] P. A. Svensson, I. Barber, and E. Forsgren. Shoaling behaviour of the two-spotted goby. Journal of Fish Biology, 56(6):1477–1487, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Devine:2000:CRS

[125] J. A. Devine, C. E. Adams, and P. S. Maitland. Changes in reproductive strategy in the ruffe during a period of establishment in a new habitat. Journal of Fish Biology, 56(6):1488–1496, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hill:2000:ITP

[126] S. Hill, M. T. Burrows, and R. N. Hughes. Increased turning per unit distance as an area-restricted search mechanism in a pause-travel predator, juvenile plaice, foraging for buried bivalves. Journal of Fish Biology, 56(6):1497–1508, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pouyaud:2000:CPP

[127] L. Pouyaud, G. G. Teugels, R. Gustiano, and M. Legendre. Contribution to the phylogeny of pangasiid catfishes based on allozymes and mitochondrial DNA. Journal of Fish Biology, 56(6):1509–1538, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claiborne:2000:BAB

[128] J. B. Claiborne, D. Compton-Mccullough, and J. S. Walton. Branchial acid-base transfers in the euryhaline oyster toadfish during exposure to dilute sea water. Journal of Fish Biology, 56(6):1539–1544, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golani:2000:FRB

[129] D. Golani. First record of the bluespotted cornetfish from the Mediterranean Sea. Journal of Fish Biology, 56(6):1545–1547, June 2000. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rasmussen:2000:IGR

[130] R. S. Rasmussen and T. H. Ostenfeld. Influence of growth rate on white muscle dynamics in rainbow trout and brook trout. Journal of Fish Biology, 56(6):1548–1552, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wanzenbouck:2000:PEL

[131] J. Wanzenböuck, B. Lahnsteiner, and K. Maier. Pelagic early life phase of the bullhead in a freshwater lake. Journal of Fish Biology, 56(6):1553– 1557, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blin:2000:PMC

[132] C. Blin, S. Panerat, S. Kaushik, and R. Krishnamoorthy. Partial molecular cloning and tissue distribution of hexokinase I cDNA in common carp. Journal of Fish Biology, 56(6):1558–1561, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRe

[133] Anonymous. Book reviews. Journal of Fish Biology, 56(6):1562–1565, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRf

[134] Anonymous. Book received. Journal of Fish Biology, 56(6):1565, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ae

[135] Anonymous. Announcements. Journal of Fish Biology, 56(6):1566–1567, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:E

[136] Anonymous. Erratum. Journal of Fish Biology, 56(6):1568, June 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uglem:2000:PVB

[137] I. Uglem, G. Rosenqvist, and H. S. Wasslavik. Phenotypic variation between dimorphic males in corkwing wrasse. Journal of Fish Biology, 57(1):1–14, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tolonen:2000:ADD

[138] A. Tolonen, H. Rita, and H. Peltonen. Abundance and distribution of Diphyllobothrium ditremum Creplin (Cestoda: Pseudophyllidea) plerocercoids in benthic whitefish, in northern Finnish Lapland. Journal of Fish Biology, 57(1):15–28, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Berthou:2000:FIP

[139] E. García-Berthou and R. Moreno-Amich. Food of introduced pumpkinseed sunfish: ontogenetic diet shift and seasonal variation. Journal of Fish Biology, 57(1):29–40, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OConnor:2000:SSM

[140] K. I. O’Connor, A. C. Taylor, and N. B. Metcalfe. The stability of standard metabolic rate during a period of food deprivation in juvenile Atlantic salmon. Journal of Fish Biology, 57(1):41–51, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sagnes:2000:SDS

[141] P. Sagnes, J-Y. Champagne, and R. Morel. Shifts in drag and swimming potential during grayling ontogenesis: relations with habitat use. Journal of Fish Biology, 57(1):52–68, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pavlov:2000:ESY

[142] D. S. Pavlov, R. V. Sadkovskii, V. V. Kostin, and A. I. Lupandin. Experimental study of young fish distribution and behaviour under combined influence of baro-, photo- and thermo-gradients. Journal of Fish Biology, 57(1):69–81, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Power:2000:EIE

[143] M. Power, J. B. Dempson, G. Power, and J. D. Reist. Environmental influences on an exploited anadromous Arctic charr stock in Labrador. Journal of Fish Biology, 57(1):82–98, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glowacki:2000:IIF

[144] Ł. Głowacki and T. Penczak. Impoundment impact on fish in the Warta River: species richness and sample size in the rarefaction method. Journal of Fish Biology, 57(1):99–108, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vamosi:2000:TES

[145] S. M. Vamosi, T. Hatfield, and D. Schluter. A test of ecological selection against young-of-the-year hybrids of sympatric sticklebacks. Journal of Fish Biology, 57(1):109–121, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2000:ACE

[146] B. A. Henderson, T. Trivedi, and N. Collins. Annual cycle of energy allocation to growth and reproduction of yellow perch. Journal of Fish Biology, 57(1):122–133, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McFarlane:2000:LGD

[147] M. B. McFarlane, D. J. Cripe, and S. H. Thompson. Larval growth and development of cultured Pacific bonito. Journal of Fish Biology, 57(1): 134–144, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Halvorsen:2000:GAS

[148] M. Halvorsen and M.-A. Svenning. Growth of Atlantic salmon parr in fluvial and lacustrine habits. Journal of Fish Biology, 57(1):145–160, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okamura:2000:RES

[149] A. Okamura, H. Zhang, S. Tanaka, N. Hore, N. Mikawa, T. Utoh, and H. P. Oka. Re-examination of the spermatozol ultrastructure of eels: observations of the external morphology of spermatozoa in three species. Journal of Fish Biology, 57(1):161–169, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shimizu:2000:SLP

[150] M. Shimizu, H. Fukada, T. Fujita, N. Hiramatsu, and A. Hara. Serum levels of precursors to vitelline envelope proteins (choriogenins) in Sakhalin taimen after treatment with oestrogen and during oocyte growth. Journal of Fish Biology, 57(1):170–181, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merson:2000:ODS

[151] R. R. Merson, C. S. Casey, C. Martinez, B. Soffientino, M. Chandlee, and J. L. Specker. Oocyte development in summer flounder: seasonal changes and steroid correlates. Journal of Fish Biology, 57(1):182–196, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Altimiras:2000:NIR

[152] J. Altimiras and E. Larsen. Non-invasive recording of heart rate and ventilation rate in rainbow trout during rest and swimming. Fish go wireless! Journal of Fish Biology, 57(1):197–209, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baras:2000:PSD

[153] E. Baras, C. Prignon, G. Gohoungo, and C. Méalard. Phenotypic sex differentiation of blue tilapia under constant and fluctuating thermal regimes and its adaptive and evolutionary implications. Journal of Fish Biology, 57(1):210–223, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:2000:TRR

[154] I. D. McCarthy. Temporal repeatability of relative standard metabolic rate in juvenile Atlantic salmon and its relation to life history variation. Journal of Fish Biology, 57(1):224–238, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2000:MTC

[155] H. C. Lin, S. C. Hsu, and P. P. Hwang. Maternal transfer of cadmium tolerance in larval Oreochromis mossambicus. Journal of Fish Biology, 57(1):239–248, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hernandez:2000:DPR

[156] R. Hernández, R. T. Lacomba, Y. N. Uviñtas, and R. Oltra. Distribution pattern of river blennies in the Júcar River basin (eastern Spain). Journal of Fish Biology, 57(1):250–254, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pallaoro:2000:VDB

[157] A. Pallaoro and M. Kovačića. Vanneaugobius dollfusi Brownell, 1978 a rare fish new to the Adriatic Sea. Journal of Fish Biology, 57(1):255– 257, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bailey:2000:MAS

[158] J. Bailey, A. Alanäuräu, and E. Bräunnäus. Methods for assessing social status in Arctic charr. Journal of Fish Biology, 57(1):258–261, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Af

[159] Anonymous. Announcements. Journal of Fish Biology, 57(1):262–263, July 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haunfling:2000:PEC

[160] B. Häunfling and R. Brandl. Phylogenetics of European cyprinids: insights from allozymes. Journal of Fish Biology, 57(2):265–276, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gurgens:2000:EAM

[161] C. Gurgens, D. F. Russell, and L. A. Wlikens. Electrosensory avoidance of metal obstacles by the paddlefish. Journal of Fish Biology, 57(2):277– 290, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:2000:FRG

[162] M. S. Allen, M. V. Hoyer, and D. E. Canfield, Jr. Factors related to gizzard shad and the threadfin shad occurrence and abundance in Florida lakes. Journal of Fish Biology, 57(2):291–302, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deLeoan:2000:HCN

[163] G. Péarez-Ponce de Leóan, L. Garcíaa-Prieto, V. Leóan-Réggagnon, and A. Choudhury. Helminth communities of native and introduced fishes in Lake Pátzcuaro, Michoacán, México. Journal of Fish Biology, 57(2):303– 325, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsson:2000:FRE

[164] J. I. Johnsson, E. Jöusson, E. Petersson, T. Jäurvi, and B. Th. Bjöurnsson. Fitness-related effects of growth investment in brown trout under field and hatchery conditions. Journal of Fish Biology, 57(2):326– 336, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dowling:2000:FKB

[165] N. A. Dowling, S. J. Hall, and J. G. Mitchell. Foraging kinematics of barramundi during early stages of development. Journal of Fish Biology, 57(2):337–353, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osinov:2000:GDP

[166] A. G. Osinov and V. S. Lebedev. Genetc divergence and phylogeny of the Salmoninae based on allozyme data. Journal of Fish Biology, 57 (2):354–381, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sasal:2000:ARS

[167] P. Sasal and C. Pampoulie. Asymmetry, reproductive success and parasitism of Pomatoschistus microps in a French lagoon. Journal of Fish Biology, 57(2):382–390, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grutter:2000:ECH

[168] A. S. Grutter and N. W. Pankhurst. The effects of capture, handling, confinement and ectoparasite load on plasma levels of cortisol, glucose and lactate in the coral reef fish Hemigymnus melapterus. Journal of Fish Biology, 57(2):391–401, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedersen:2000:FCD

[169] J. Pedersen. Food consumption and daily feeding periodicity: comparison between pelagic and demersal whiting in the North Sea. Journal of Fish Biology, 57(2):402–416, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domaizon:2000:IZP

[170] I. Domaizon, C. Desvilettes, D. Debroas, and G. Bourdier. Influence of zooplankton and phytoplankton on the fatty acid composition of digesta and tissue lipids of silver carp: mesocosm experiment. Journal of Fish Biology, 57(2):417–432, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beaumoint:2000:ESE

[171] W. R. C. Beaumoint, M. J. Lle, and M. A. Rouen. An evaluation of some electrical waveforms and voltages used for electric fishing; with special reference to their use in backpack electric fishing gear. Journal of Fish Biology, 57(2):433–444, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carisle:2000:EKG

[172] S. L. Carisle, S. K. Marxer-Miller, A. V. M. Canario, R. F. Oliveira, L. Carneiro, and M. S. Grober. Effects of 11-ketotestosterone on genital papilla morphology in the sex changing fish Lythrypnus dalli. Journal of Fish Biology, 57(2):445–456, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manera:2000:MAG

[173] M. Manera, R. Sera, G. Isani, and E. Carpenéa. Macrophage aggregates in gilthead sea bream fed copper, iron and zinc enriched diets. Journal of Fish Biology, 57(2):457–465, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gleeson:2000:DIA

[174] D. J. Gleeson, H. I. McCalum, and I. P. F. Owens. Differences in initial and acquired resistance to Ichthyophthirius multifiliis between populations of rainbowfish. Journal of Fish Biology, 57(2):466–475, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stratoudakis:2000:ISE

[175] Y. Stratoudakis, A. Barbosa, and I. Meneses. Infection of sardine eggs by the protistan endoparasite Ichthyodinium chabelardi off Portugal. Journal of Fish Biology, 57(2):476–482, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rosenberger:2000:RCT

[176] A. E. Rosenberger and L. J. Chapman. Respiratory characters of three species of haplochromine cichlids: Implications for use of wetland refugia. Journal of Fish Biology, 57(2):483–501, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeGaudemar:2000:ECR

[177] B. De Gaudemar, J. M. Bonzom, and E. Beall. Effects of courtship and relative mate size on sexual motivation in Atlantic salmon. Journal of Fish Biology, 57(2):502–515, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nilisson:2000:RGE

[178] P. A. Nilisson and C. Bröunamark. The role of gastric evacuation rate in handling time of equal-mass rations of different prey sizes in northern pike. Journal of Fish Biology, 57(2):516–524, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Comiskey:2000:CSM

[179] P. Comiskey and K. MacKenzie. Corynosoma spp. may be useful biological tags for saithe in the northern North Sea. Journal of Fish Biology, 57(2):525–528, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2000:FRA

[180] D. W. Evans and M. A. Matthews. First record of Argulus foliaceus on the European eel in the British Isles. Journal of Fish Biology, 57(2):529– 530, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorstad:2000:ETT

[181] E. B. Thorstad, F. Økland, and B. Finstad. Effects of telemetry transmitters on swimming performance of adult Atlantic salmon. Journal of Fish Biology, 57(2):531–535, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2000:NSE

[182] H. H. Ng and H. H. Tan. A new species of Encheloclarias from Sumatra. Journal of Fish Biology, 57(2):536–539, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRg

[183] Anonymous. Book reviews. Journal of Fish Biology, 57(2):540–543, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ag

[184] Anonymous. Announcements. Journal of Fish Biology, 57(2):544–545, August 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorstad:2000:RTE

[185] E. B. Thorstad, F. Økland, A. Koed, and R. S. McKinley. Radiotransmitted electromyogram signals as indicators of swimming speed in lake trout and brown trout. Journal of Fish Biology, 57(3):547–561, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paxton:2000:RPE

[186] C. G. M. Paxton and L. G. Willoughby. Resistance of perch eggs to attack by aquatic fungi. Journal of Fish Biology, 57(3):562–570, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Comeau:2000:SCT

[187] L. A. Comeau, S. E. Campana, J. M. Hanson, and G. A. Chouinard. Seasonal changes of thyroid hormones in field-collected Atlantic cod in relation to condition indices, water temperature and photoperiod. Journal of Fish Biology, 57(3):571–588, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arndt:2000:ISM

[188] S. K. A. Arndt. Influence of sexual maturity on feeding, growth and energy stores of wild Atlantic salmon parr. Journal of Fish Biology, 57 (3):589–596, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duis:2000:SSR

[189] K. Duis and A. Oberemm. Survival and sublethal responses of early life stages of pike exposed to low pH in artificial post-mining lake water. Journal of Fish Biology, 57(3):597–613, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:2000:MDD

[190] T. L. King, A. P. Spidle, M. S. Eackles, B. A. Lubinski, and W. B. Schill. Mitochondrial DNA diversity in North American and European Atlantic salmon with emphasis on the Downeast rivers of Maine. Journal of Fish Biology, 57(3):614–630, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Figueroa:2000:EFD

[191] R. I. Figueroa, R. Rodríguez-Sabarís, M. Aldegunde, and J. L. Soengas. Effects of food deprivation on 24 h-changes in brain and liver carbohydrate and ketone body metabolism of rainbow trout. Journal of Fish Biology, 57(3):631–646, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berejikian:2000:FMC

[192] B. A. Berejikian, E. P. Tezak, and A. L. LaRae. Female mate choice and spawning behaviour of Chinook salmon under experimental conditions. Journal of Fish Biology, 57(3):647–661, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoff:2000:BEN

[193] G. R. Hoff, T. W. Buckley, J. C. Drazen, and K. M. Duncan. Biology and ecology of Nezumia liolepis and N. stelgidolepis from the west coast of North America. Journal of Fish Biology, 57(3):662–680, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Graynoth:2000:IDR

[194] E. Graynoth and M. J. Taylor. Influence of different rations and water temperatures on the growth rates of shortfinned eels and longfinned eels. Journal of Fish Biology, 57(3):681–699, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Megalofonou:2000:AGM

[195] P. Megalofonou. Age and growth of Mediterranean albacore. Journal of Fish Biology, 57(3):700–715, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamlin:2000:HME

[196] H. J. Hamlin, I. Hunt Von Herbing, and L. J. Kling. Histological and morphological evaluations of the digestive tract and associated organs of haddock throughout post-hatching ontogeny. Journal of Fish Biology, 57 (3):716–732, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreau:2000:BAA

[197] R. Moreau and K. Dabrowski. Biosynthesis of ascorbic acid by extant actinopterygians. Journal of Fish Biology, 57(3):733–745, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herrmann:2000:EBS

[198] J-P. Herrmann and E. C. Enders. Effect of body size on the standard metabolism of horse mackerel. Journal of Fish Biology, 57(3):746–760, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeBoeck:2000:SSR

[199] G. De Boeck, A. Vlaeminck, A. Van der Linden, and R. Blust. Salt stress and resistance to hypoxic challenges in the common carp (Cyprinus carpio L.). Journal of Fish Biology, 57(3):761–776, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:2000:CBS

[200] I. Barber, S. A. Arnott, V. A. Braithwaite, J. Andrew, W. Mullen, and F. A. Huntingford. Carotenoid-based sexual coloration and body condition in nesting male sticklebacks. Journal of Fish Biology, 57(3):777–790, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suzuki:2000:RBO

[201] H. I. Suzuki, A. A. Agostinho, and K. O. Winemiller. Relationship between oocyte morphology and reproductive strategy in loricariid catfishes of the Paraná River, Brazil. Journal of Fish Biology, 57(3):791–807, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heikineheimo:2000:DGC

[202] O. Heikineheimo, M. Miinalainen, and H. Peltonen. Diet, growth and competitive abilities of sympatric whitefish forms in a dense introduced population: results of a stocking experiment. Journal of Fish Biology, 57 (3):808–827, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stefanni:2000:FRN

[203] S. Stefanni. First record of the Norway goby in the Adriatic Sea. Journal of Fish Biology, 57(3):828–830, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domenici:2000:SWS

[204] P. Domenici, R. S. Batty, and T. Similä. Spacing of wild schooling herring while encircled by killer whales. Journal of Fish Biology, 57(3):831–836, September 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magurran:2000:SDB

[205] A. E. Magurran and C. Maciás Garcia. Sex differences in behaviour as an indirect consequence of mating system. Journal of Fish Biology, 57 (4):839–857, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Treasure:2000:PSP

[206] J. W. Treasure and T. Turnbull. The pathology and seawater performance of farmed Atlantic salmon infected with glochidia of Margaritifera margaritifera. Journal of Fish Biology, 57(4):858–866, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2000:SPF

[207] S. Xie, Y. Cui, T. Zhang, and Z. Li. Seasonal patterns in feeding ecology of three small fishes in the Biandantang Lake, China. Journal of Fish Biology, 57(4):867–880, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tanck:2000:CSS

[208] M. W. T. Tanck, G. H. R. Booms, E. H. Eding, S. E. Wendellar Bonga, and J. Komen. Cold shocks: a stressor for common carp. Journal of Fish Biology, 57(4):881–894, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2000:GFA

[209] M. C. P. Amorim and A. D. Hawkins. Growling for food: acoustic emissions during competitive feeding of the streaked gurnard. Journal of Fish Biology, 57(4):895–907, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boyce:2000:DRG

[210] S. J. Boyce, A. W. A. Murray, and L. S. Peck. Digestion rate, gut passage time and absorption efficiency in the Antarctic spiny plunderfish. Journal of Fish Biology, 57(4):908–929, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chiaramonte:2000:BMS

[211] G. E. Chiaramonte and A. D. Pettovello. The biology of Mustelus schmitti in southern Patagonia, Argentina. Journal of Fish Biology, 57 (4):930–942, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Unwin:2000:DJG

[212] M. J. Unwin, T. P. Quinn, M. T. Kinnison, and N. C. Boustead. Divergence in juvenile growth and life history in two recently colonized and partially isolated Chinook salmon populations. Journal of Fish Biology, 57(4):943–960, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Radke:2000:SPB

[213] R. J. Radke, T. Petzold, and C. Wolter. Suitability of pharyngeal bone measures commonly used for reconstruction of prey fish length. Journal of Fish Biology, 57(4):961–967, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Acharia:2000:MET

[214] K. Acharia, B. Lal, and T. P. Singh. Modulatory effect of temperature on the influence of dietary linolenic (18 : 3 n-3) and linoleic (18 : 2 n-6) acids on the gonadal recrudescence in Clarias batrachus (L.). Journal of Fish Biology, 57(4):968–980, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hockaday:2000:UTN

[215] S. Hockaday, T. A. Beddow, M. Stone, P. Hancock, and L. G. Ross. Using truss networks to estimate the biomass of Oreochromis niloticus, and to investigate shape characteristics. Journal of Fish Biology, 57(4): 981–1000, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baras:2000:SCDa

[216] E. Baras, M. Ndao, M. Y. J. Maxi, D. Jeandrain, J. P. Thomé, P. Vanadewalle, and C. Mélard. Sibling cannibalism in dorada under experimental conditions. I. Ontogeny, dynamics, bioenergetics of cannibalism and prey size selectivity. Journal of Fish Biology, 57(4):1001–1020, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baras:2000:SCDb

[217] E. Baras, M. Y. J. Maxi, M. Ndao, and C. Mélard. Sibling cannibalism in dorada under experimental conditions. II. Effect of initial size heterogeneity, diet and light regime on early cannibalism. Journal of Fish Biology, 57(4):1021–1036, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartl:2000:BGG

[218] M. G. J. Hartl, S. Hutchinson, L. E. Hawkins, and M. Eledjam. The biometry of gills of 0-group European flounder. Journal of Fish Biology, 57(4):1037–1046, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powell:2000:EGH

[219] M. D. Powell, D. Fisk, and B. F. Nowak. Effects of graded hypoxia on Atlantic salmon infected with amoebic gill disease. Journal of Fish Biology, 57(4):1047–1057, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yambe:2000:UOF

[220] H. Yambe and F. Yamazaki. Urine of ovulated female masu salmon attracts immature male parr treated with methyltestosterone. Journal of Fish Biology, 57(4):1058–1064, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahman:2000:LST

[221] M. S. Rahman, A. Takemura, and K. Takano. Lunar synchronization of testicular development and plasma steroid hormone profiles in the golden rabbitfish. Journal of Fish Biology, 57(4):1065–1074, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnal:2000:DBC

[222] C. Arnal and I. M. Côté. Diet of broadstripe cleaning gobies on a Barbadian reef. Journal of Fish Biology, 57(4):1075–1082, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hislop:2000:NSC

[223] J. R. G. Hislop, J. C. Holst, and D. Skagen. Near-surface captures of post-juvenile anglerfish in the North-east Atlantic — an unsolved mystery. Journal of Fish Biology, 57(4):1083–1087, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRh

[224] Anonymous. Book reviews. Journal of Fish Biology, 57(4):1088–1090, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ah

[225] Anonymous. Announcements. Journal of Fish Biology, 57(4):1091–1092, October 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arunkumar:2000:LGL

[226] L. Arunkumar. Loaches of the genus Lepidocephalichthys, from Manipur, with description of a new species. Journal of Fish Biology, 57(5):1093– 1104, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horne:2000:CAM

[227] J. K. Horne, P. D. Walline, and J. M. Jech. Comparing acoustic model predictions to in situ backscatter measurements of fish with dualchambered swimbladders. Journal of Fish Biology, 57(5):1105–1121, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanmartin:2000:HPC

[228] M. L. Sanmartín, M. F. Alvarez, D. Peris, R. Iglesias, and J. Leiro. Helminth parasite communities of the conger eel in the estuaries of Arousa and Muros (Galicia, north-west Spain). Journal of Fish Biology, 57(5):1122–1133, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rocha-Olivares:2000:HGD

[229] A. Rocha-Olivares, N. M. Garber, and K. C. Stuck. High genetic diversity, large inter-oceanic divergence and historical demography of the striped mullet. Journal of Fish Biology, 57(5):1134–1149, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2000:EHI

[230] R. S. Brown, G. Power, S. Beltaos, and T. A. Beddow. Effects of hanging ice dams on winter movements and swimming activity of fish. Journal of Fish Biology, 57(5):1150–1159, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhou:2000:MAC

[231] B. S. Zhou, R. S. S. Wu, D. J. Randall, P. K. S. Lam, Y. K. Ip, and S. F. Chew. Metabolic adjustments in the common carp during prolonged hypoxia. Journal of Fish Biology, 57(5):1160–1171, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2000:CGD

[232] A. H. Rikardsen, P.-A. Amundsen, P. A. Bjørn, and M. Johansen. Comparison of growth, diet and food consumption of sea-run and lakedwelling Arctic charr. Journal of Fish Biology, 57(5):1172–1188, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cieri:2000:DBO

[233] M. D. Cieri and J. D. McCleave. Discrepancies between otoliths of larvae and juveniles of the American eel: is something fishy happening at metamorphosis? Journal of Fish Biology, 57(5):1189–1198, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delling:2000:MVH

[234] B. Delling, A. J. Crivelli, J-F. Rubin, and P. Berrebi. Morphological variation in hybrids between Salmo marmoratus, and alien Salmo species in the Volarja stream, Soca River basin, Slovenia. Journal of Fish Biology, 57(5):1199–1212, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsland:2000:CVG

[235] A. K. Imsland, A. Foss, G. Névdal, T. Cross, S. W. Bonga, E. A. Ham, and S. O. Stefansson. Countergradient variation in growth and food conversion efficiency of juvenile turbot. Journal of Fish Biology, 57(5): 1213–1226, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Langeland:2000:BRY

[236] A. Langeland and T. Pedersen. Book review: A 27-year study of brown trout population dynamics and exploitation in Lake Songsjøen, central Norway. Journal of Fish Biology, 57(5):1227–1244, November 2000. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tzeng:2000:DSG

[237] W. N. Tzeng, H. R. Lin, C. H. Wang, and S. N. Xu. Differences in size and growth rates of male and female migrating Japanese eels in Pearl River, China. Journal of Fish Biology, 57(5):1245–1253, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schutz:2000:GMD

[238] D. Schütz and M. Taborsky. Giant males or dwarf females: what determines the extreme sexual size dimorphism in Lamprologus callipterus? Journal of Fish Biology, 57(5):1254–1265, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koed:2000:AMM

[239] A. Koed, P. Mejlhede, K. Balleby, and K. Aarestrup. Annual movement and migration of adult pikeperch in a lowland river. Journal of Fish Biology, 57(5):1266–1279, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Villarroel:2000:CDS

[240] J. C. Villarroel and E. Acuña. Consumption of deep-sea shrimp by bigeye flounder Hippoglossina macrops in fishing grounds off northern Chile. Journal of Fish Biology, 57(5):1280–1289, November 2000. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hillgruber:2000:VDF

[241] N. Hillgruber and M. Kloppmann. Vertical distribution and feeding of larval blue whiting in turbulent waters above Porcupine Bank. Journal of Fish Biology, 57(5):1290–1311, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmgren:2000:SSB

[242] K. Holmgren and M. Appelberg. Size structure of benthic freshwater fish communities in relation to environmental gradients. Journal of Fish Biology, 57(5):1312–1330, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2000:DSI

[243] R. Griffiths, K. L. Orr, A. Adam, and I. Barber. DNA sex identification in the three-spined stickleback. Journal of Fish Biology, 57(5):1331–1334, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pizzicori:2000:SCN

[244] P. Pizzicori, L. Castriota, G. Marino, and F. Andaloro. Seriola carpenteri: a new immigrant in the Mediterranean from the Atlantic Ocean. Journal of Fish Biology, 57(5):1335–1338, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freyhof:2000:EKR

[245] J. Freyhof, M. Scholten, A. Bischoff, J. Wanzenböck, S. Staas, and C. Wolter. Extensions to the known range of the whitefin gudgeon to Europe and biogeographical implications. Journal of Fish Biology, 57(5): 1339–1342, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiwary:2000:GDT

[246] B. K. Tiwary, R. Kirubagaran, and A. K. Ray. Gonadal development in triploid Heteropneustes fossilis. Journal of Fish Biology, 57(5):1343– 1348, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Ai

[247] Anonymous. Announcements. Journal of Fish Biology, 57(5):1349, November 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoare:2000:BSS

[248] D. J. Hoare, J. Krause, N. Peuhkuri, and J.-G. J. Godin. Body size and shoaling in fish. Journal of Fish Biology, 57(6):1351–1366, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fishback:2000:MAH

[249] A. G. Fishback, R. G. Danzmann, and M. M. Ferguson. Microsatellite allelic heterogeneity among hatchery rainbow trout maturing in different seasons. Journal of Fish Biology, 57(6):1367–1380, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovac:2000:EDZ

[250] V. Kováč. Early development of Zingel streber. Journal of Fish Biology, 57(6):1381–1403, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moller:2000:RTL

[251] P. R. Møller. Restoration of the taxon Lycodes marisalbi, with notes on its disjunct Arctic distribution. Journal of Fish Biology, 57(6):1404–1415, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2000:MMS

[252] R. F. Oliveira, J. A. Miranda, N. Carvalho, E. J. Gonçalves, M. S. Grober, and R. S. Santos. Male mating success in the Azorean rock-pool blenny: the effects of body size, male behaviour and nest characteristics. Journal of Fish Biology, 57(6):1416–1428, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2000:OMA

[253] R. M. Rideout, M. P. M. Burton, and G. A. Rose. Observations on mass atresia and skipped spawning in northern Atlantic cod, from Smith Sound, Newfoundland. Journal of Fish Biology, 57(6):1429–1440, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pampoulie:2000:AVR

[254] C. Pampoulie, J. L. Bouchereau, E. Rosecchi, G. Poizat, and A. J. Crivelli. Annual variations in the reproductive traits of Pomatoschistus microps in a Mediterranean lagoon undergoing environmental changes: evidence of phenotypic plasticity. Journal of Fish Biology, 57(6):1441–1452, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2000:UCO

[255] S. P. Griffiths. The use of clove oil as an anaesthetic and method for sampling intertidal rockpool fishes. Journal of Fish Biology, 57(6):1453– 1464, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faber:2000:CGF

[256] J. E. Faber and M. M. White. Comparison of gene flow estimates between species of darters in different streams. Journal of Fish Biology, 57(6): 1465–1473, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tischler:2000:SCT

[257] G. Tischler, H. Gassner, and J. Wanzenböck. Sampling characteristics of two methods for capturing age-0 fish in pelagic lake habitats. Journal of Fish Biology, 57(6):1474–1487, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez:2000:IMS

[258] L. Pérez, J. F. Aturiano, A. Tomás, S. Zegrari, R. Barrera, F. J. Espinós, J. C. Navarro, and M. Jover. Induction of maturation and spermiation in the male European eel: assessment of sperm quality throughout treatment. Journal of Fish Biology, 57(6):1488–1504, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grandi:2000:EDS

[259] G. Grandi, F. Poerio, G. Colombo, and M. Chicca. Effects of diet supplementation with carp ovary on gonad differentiation and growth of the European eel. Journal of Fish Biology, 57(6):1505–1525, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Domenici:2000:EPH

[260] P. Domenici, J. F. Steffensen, and R. S. Batty. The effect of progressive hypoxia on swimming activity and schooling in Atlantic herring. Journal of Fish Biology, 57(6):1526–1538, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ekau:2000:VDI

[261] W. Ekau and J. Blay. Validation of daily increment deposition and early development in the otoliths of Sarotherodon melanotheron. Journal of Fish Biology, 57(6):1539–1549, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cabral:2000:CFE

[262] H. N. Cabral. Comparative feeding ecology of sympatric Solea solea and S. senegalensis, within the nursery areas of the Tagus estuary, Portugal. Journal of Fish Biology, 57(6):1550–1562, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Asturiano:2000:EPF

[263] J. F. Asturiano, L. A. Sorbera, S. Zanuy, and M. Carrillo. Effects of polyunsaturated fatty acids and gonadotropin on prostaglandin series E production in a primary testis cell culture system for the European sea bass. Journal of Fish Biology, 57(6):1563–1574, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Girard:2000:TSM

[264] M. Girard, P. Rivalan, and G. Sinquin. Testis and sperm morphology in two deep-water squaloid sharks, Centroscymnus coelolepis and Centrophorus squamosus. Journal of Fish Biology, 57(6):1575–1589, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marnane:2000:SFH

[265] M. J. Marnane. Site fidelity and homing behaviour in coral reef cardinalfishes. Journal of Fish Biology, 57(6):1590–1600, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ochi:2000:FRB

[266] H. Ochi, A. Rossiter, and Y. Yanagisawa. The first record of a biparental mouthbrooding catfish. Journal of Fish Biology, 57(6):1601–1604, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suneetha:2000:AMD

[267] K. B. Suneetha, G. Dahle, and G. Nævdal. Analysis of mitochondrial DNA sequences from two Maurolicus taxa: evidence for separate species? Journal of Fish Biology, 57(6):1605–1609, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broberg:2000:IPI

[268] M. M. Broberg, E. E. Nielsen, and C. Dieperink. Incidence of physical injury of mature male parr in a natural population of brown trout. Journal of Fish Biology, 57(6):1610–1612, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:BRi

[269] Anonymous. Book reviews. Journal of Fish Biology, 57(6):1613–1615, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:Aj

[270] Anonymous. Announcements. Journal of Fish Biology, 57(6):1616, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raymond:2000:SCF

[271] J. A. Raymond and A. Hassel. Some characteristics of freezing avoidance in two osmerids, rainbow smelt and capelin. Journal of Fish Biology, 57 (sA):1–7, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2000:HMP

[272] C. W. Evans, J. M. Hills, and J. M. J. Dickson. Heavy metal pollution in Antarctica: a molecular ecotoxicological approach to exposure assessment. Journal of Fish Biology, 57(sA):8–19, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riccio:2000:FDH

[273] A. Riccio, M. Tamburrini, V. Carratore, and G. di Prisco. Functionally distinct haemoglobins of the cryopelagic Antarctic teleost Pagothenia borchgrevinki. Journal of Fish Biology, 57(sA):20–32, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cocca:2000:CFG

[274] E. Cocca, H. W. Detrich III, S. K. Parker, and G. di Prisco. A cluster of four globin genes from the Antarctic fish Notothenia coriiceps. Journal of Fish Biology, 57(sA):33–50, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Politov:2000:IPC

[275] D. V. Politov, N. Yu. Gordon, K. I. Afanasiev, Yu. P. Altukhov, and J. W. Bickham. Identification of palearctic coregonid fish species using mtDNA and allozyme genetic markers. Journal of Fish Biology, 57(sA): 51–71, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2000:AMD

[276] P. Smith and M. McVeagh. Allozyme and microsatellite DNA markers of toothfish population structure in the Southern Ocean. Journal of Fish Biology, 57(sA):72–83, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eastman:2000:FAC

[277] J. T. Eastman and A. R. McCune. Fishes on the Antarctic continental shelf: evolution of amarine species flock? Journal of Fish Biology, 57 (sA):84–102, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Everson:2000:IAV

[278] I. Everson, K.-H. Kock, and J. Ellison. Inter-annual variation in the gonad cycle of the mackerel icefish. Journal of Fish Biology, 57(sA): 103–111, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powell:2000:VCD

[279] M. D. Powell, M. E. Forster, and W. Davision. In vitro carbon dioxide excretion from erythrocytes of two species of Antarctic fishes and its inhibition by catecholamines. Journal of Fish Biology, 57(sA):112–120, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Icardo:2000:BAS

[280] J. M. Icardo, E. Colvee, M. C. Cerra, and B. Tota. The bulbus arteriosus of stenothermal and temperate teleosts: a morphological approach. Journal of Fish Biology, 57(sA):121–135, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salmenkova:2000:GDC

[281] E. A. Salmenkova, V. T. Omelchenko, A. A. Kolesnikov, and T. V. Malinina. Genetic differentiation of charrs in the Russian north and far east. Journal of Fish Biology, 57(sA):136–157, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Novikov:2000:FEF

[282] G. G. Novikov, D. V. Politov, A. A. Makhorov, T. V. Malinina, K. I. Afanasiev, and B. Fernholm. Freshwater and estuarine fishes of the Russian Arctic coast (the Swedish–Russian expedition ‘tundra ecology — 94 ‘). Journal of Fish Biology, 57(sA):158–162, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mesa:2000:FDL

[283] M. La Mesa, A. Granata, L. Guglielmo, and M. Vacchi. First description of larvae of Acanthodraco dewitti in the Ross Sea, Antarctica. Journal of Fish Biology, 57(sA):163–166, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christiansen:2000:SDI

[284] J. S. Christiansen. Sex differences in ionoregulatory responses to dietary oil exposure in polar cod. Journal of Fish Biology, 57(sA):167–170, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2000:PA

[285] Anonymous. Poster abstract. Journal of Fish Biology, 57(sA):171–179, December 2000. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:2001:SIH

[286] M. W. Davis, B. L. Olla, and C. B. Schreck. Stress induced by hooking, net towing, elevated sea water temperature and air in sablefish: lack of concordance between mortality and physiological measures of stress. Journal of Fish Biology, 58(1):1–15, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salminen:2001:DPS

[287] M. Salminen, E. Erkamo, and J. Salmi. Diet of post-smolt and one-sea-winter Atlantic salmon in the Bothnian Sea, Northern Baltic. Journal of Fish Biology, 58(1):16–35, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purchase:2001:SSC

[288] C. F. Purchase and J. A. Brown. Stock-specific changes in growth rates, food conversion efficiencies, and energy allocation in response to temperature change in juvenile Atlantic cod. Journal of Fish Biology, 58 (1):36–52, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deudero:2001:ITR

[289] S. Deudero. Interspecific trophic relationships among pelagic fish species underneath FADs. Journal of Fish Biology, 58(1):53–67, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gratwicke:2001:RBE

[290] B. Gratwicke and B. E. Marshall. The relationship between the exotic predators Micropterus salmoides and Serranochromis robustus and native stream fishes in Zimbabwe. Journal of Fish Biology, 58(1):68–75, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Felip:2001:CGP

[291] A. Felip, F. Piferrer, S. Zanuy, and M. Carrillo. Comparative growth performance of diploid and triploid European sea bass over the first four spawning seasons. Journal of Fish Biology, 58(1):76–88, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okazaki:2001:PRB

[292] M. Okazaki, K. Naruse, A. Shima, and R. Arai. Phylogenetic relationships of bitterlings based on mitochondrial 12S ribosomal DNA sequences. Journal of Fish Biology, 58(1):89–106, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Courtenay:2001:DFS

[293] S. C. Courtenay, T. P. Quinn, H. M. C. Dupuis, C. Groot, and P. A. Larkin. Discrimination of family-specific odours by juvenile Coho salmon: roles of learning and odour concentration. Journal of Fish Biology, 58 (1):107–125, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stephenson:2001:ASI

[294] P. C. Stephenson, J. S. Edmons, M. J. Moran, and N. Caputi. Analysis of stable isotope ratios to investigate stock structure of red emperor and Rankin cod in northern Western Australia. Journal of Fish Biology, 58 (1):126–144, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faria:2001:MST

[295] C. Faria and V. Almada. Microhabitat segregation in three rocky intertidal fish species in Portugal: does it reflect interspecific competition?

Journal of Fish Biology, 58(1):145–159, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nolan:2001:HDI

[296] M. Nolan, S. Jobling, G. Brighty, J. P. Sumpter, and C. R. Tyler. A histological description of intersexuality in the roach. Journal of Fish Biology, 58(1):160–176, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2001:EFR

[297] W-M. Chen and G. J. Purser. The effect of feeding regime on growth, locomotor activity pattern and the development of food anticipatory activity in greenback flounder. Journal of Fish Biology, 58(1):177–187, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purser:2001:EMS

[298] G. J. Purser and W-M. Chen. The effect of meal size and meal duration on food anticipatory activity in greenback flounder. Journal of Fish Biology, 58(1):188–200, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elso:2001:PCI

[299] J. I. Elso and P. S. Giller. Physical characteristics influencing the utilization of pools by brown trout in an afforested catchment in Southern Ireland. Journal of Fish Biology, 58(1):201–221, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takegaki:2001:EFA

[300] T. Takegaki. Environmental factors affecting the spawning burrow selection by the gobiid Valenciennea longipinnis. Journal of Fish Biology, 58 (1):222–229, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gamper:2001:RRA

[301] N. L. Gamper, M. V. Savina, I. V. Brailovskaya, and A. A. Vereninov. Respiration rates, ATP content and ionic regulation in hepatocytes of starving lamprey during the pre-spawning period of their life cycle. Journal of Fish Biology, 58(1):230–239, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crozier:2001:RBF

[302] W. W. Crozier and G. J. A. Kennedy. Relationship between freshwater angling catch of Atlantic salmon and stock size in the River Bush, Northern Ireland. Journal of Fish Biology, 58(1):240–247, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fitzgerald:2001:AOA

[303] D. G. Fitzgerald, A. R. Dale, M. V. Thomas, and P. F. Sale. Application of otolith analyses to investigate broad size distributions of young yellow perch in temperate lakes. Journal of Fish Biology, 58(1):248–263, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gross:2001:GDE

[304] R. Gross, R. Kühn, M. Baars, W. Schröder, H. Stein, and O. Rottmann. Genetic differentiation of European grayling populations across the Main, Danube and Elbe drainages in Bavaria. Journal of Fish Biology, 58(1): 264–280, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koed:2001:EMS

[305] A. Koed. The effects of meal size, body size and temperature on gastric evacuation in pikeperch. Journal of Fish Biology, 58(1):281–290, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ragonese:2001:SRT

[306] S. Ragonese, G. B. Giusto, and J. H. Caruso. Second record of the toadfish Chaunax suttkusi Caruso, 1989 in the Mediterranean Sea. Journal of Fish Biology, 58(1):291–294, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fischer:2001:UPI

[307] P. Fischer, H. Kautz, H. Weber, and W. Obergfell. The use of passive integrated transponder systems (PIT) triggered by infrared-gates for behavioural studies in nocturnal, bottom-dwelling fish species. Journal of Fish Biology, 58(1):295–298, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ea

[308] Anonymous. Erratum. Journal of Fish Biology, 58(1):299, January 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lone:2001:SHP

[309] K. P. Lone, S. Ablani, and A. Al-Yaqout. Steroid hormone profiles and correlative gonadal histological changes during natural sex reversal of sobaity kept in tanks and sea-cages. Journal of Fish Biology, 58(2):305– 324, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingendahl:2001:DOC

[310] D. Ingendahl. Dissolved oxygen concentration and emergence of sea trout fry from natural redds in tributaries of the River Rhine. Journal of Fish Biology, 58(2):325–341, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blazquez:2001:CPA

[311] M. Blázquez, A. Felip, S. Zanuy, M. Carrillo, and F. Piferrer. Critical period of androgen-inducible sex differentiation in a teleost fish, the European sea bass. Journal of Fish Biology, 58(2):342–358, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Binz:2001:SDM

[312] T. Binz, C. Largiader, R. Müller, and C. Wedekind. Sequence diversity of Mhc genes in lake whitefish. Journal of Fish Biology, 58(2):359–373, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tveiten:2001:HTD

[313] H. Tveiten, S. E. Solevåg, and H. K. Johnsen. Holding temperature during the breeding season influences final maturation and egg quality in common wolffish. Journal of Fish Biology, 58(2):374–385, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2001:RAC

[314] H. H. Ng and P. K. L. Ng. A revision of the akysid catfish genus Acrochordonichthys Bleeker. Journal of Fish Biology, 58(2):386–418, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albrecht:2001:AFH

[315] M. P. Albrecht, M. F. N. Ferreira, and E. P. Caramaschi. Anatomical features and histology of the digestive tract of two related neotropical omnivorous fishes (Characiformes; Anostomidae). Journal of Fish Biology, 58(2):419–430, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Porter:2001:TLI

[316] M. J. R. Porter, N. Duncan, S. O. Handeland, S. O. Stafansson, and N. R. Bromage. Temperature, light intensity and plasma melatonin levels in juvenile Atlantic salmon. Journal of Fish Biology, 58(2):431–438, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brodeur:2001:ACO

[317] J. C. Brodeur, D. G. Dixon, and R. S. McKinly. Assessment of cardiac output as a predictor of metabolic rate in rainbow trout. Journal of Fish Biology, 58(2):439–452, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilde:2001:CRM

[318] G. R. Wilde and M. I. Muoneke. Climate-related and morphoedaphic correlates of growth in white bass. Journal of Fish Biology, 58(2):453– 461, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahman:2001:GIS

[319] M. A. Rahman, K. Ohta, H. Chuda, S. Nakano, K. Maruyama, and M. Matsuyama. Gonadotropin-induced steroidogenic shift towards maturation-inducing hormone in Japanese yellowtail during final oocyte maturation. Journal of Fish Biology, 58(2):462–474, February 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Muncoz:2001:ETD

[320] R. Rodríguez-Muñcoz, A. G. Nicieza, and F. Braña. Effects of temperature on developmental performance, survival and growth of sea lamprey embryos. Journal of Fish Biology, 58(2):475–486, February 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pettersson:2001:DDL

[321] J. C. E. Pettersson, M. M. Hansen, and T. Bohlin. Does dispersal from landlocked trout explain the coexistence of resident and migratory trout females in a small stream? Journal of Fish Biology, 58(2):487–495, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

North:2001:ELH

[322] A. W. North. Early life history strategies of notothenioids at South Georgia. Journal of Fish Biology, 58(2):496–505, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nykanen:2001:SCH

[323] M. Nykänen, A. Huusko, and A. Mäki-Petäys. Seasonal changes in the habitat use and movements of adult European grayling in a large subarctic river. Journal of Fish Biology, 58(2):506–519, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Iguchi:2001:SIR

[324] K. Iguchi and Y. Tsukamoto. Semelparous or iteroparous: resource allocation tactics in the ayu, an osmeroid fish. Journal of Fish Biology, 58 (2):520–528, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gawlicka:2001:CEP

[325] A. Gawlicka, C. T. Leggiadro, J. W. Gallant, and S. E. Douglas. Cellular expression of the pepsinogen and gastric proton pump genes in the stomach of winter flounder as determined by in situ hybridization. Journal of Fish Biology, 58(2):529–536, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faustino:2001:ODV

[326] M. Faustino and D. M. Power. Osteologic development of the viscerocranial skeleton in sea bream: alternative ossification strategies in teleost fish. Journal of Fish Biology, 58(2):537–572, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2001:CST

[327] S. J. Cooke, C. M. Bunt, J. F. Schreer, and D. H. Wahl. Comparison of several techniques for mobility and activity estimates of smallmouth bass in lentic environments. Journal of Fish Biology, 58(2):573–587, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lappalainen:2001:MAP

[328] J. Lappalainen and M. Vinni. Movement of age-1 pikeperch under the ice cover. Journal of Fish Biology, 58(2):588–590, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landergren:2001:SGS

[329] P. Landergren. Survival and growth of sea trout parr in fresh and brackish water. Journal of Fish Biology, 58(2):591–593, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qian:2001:SAW

[330] X. Qian, Y. Cui, B. Xiong, S. Xie, X. Zhu, and Y. Yang. Spontaneous activity was unaffectedby ration size in Nile tilapia and gibel carp. Journal of Fish Biology, 58(2):594–598, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:BRa

[331] Anonymous. Book reviews. Journal of Fish Biology, 58(2):599–602, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Aa

[332] Anonymous. Announcements. Journal of Fish Biology, 58(2):603, February 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2001:PSA

[333] B. Jonsson and N. Jonsson. Polymorphism and speciation in Arctic charr. Journal of Fish Biology, 58(3):605–638, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanssens:2001:RST

[334] M. Hanssens and J. Snoeks. A revised synonymy of Telmatochromis temporalis (Teleostei, Cichlidae) from Lake Tanganyika (East Africa). Journal of Fish Biology, 58(3):639–655, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caputo:2001:PGM

[335] V. Caputo, G. Candi, M. La Mesa, and E. Arneri. Pattern of gonad maturation and the question of semelparity in the paedomorphic goby Aphia minuta. Journal of Fish Biology, 58(3):656–669, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohta:2001:CSJ

[336] H. Ohta, K. Kawamura, T. Unuma, and Y. Takegoshi. Cryopreservation of the sperm of the Japanese bitterling. Journal of Fish Biology, 58 (3):670–681, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sluka:2001:IHG

[337] R. D. Sluka, M. Chiappone, and K. M. Sullivan Sealey. Influence of habitat on grouper abundance in the Florida Keys, U.S.A. Journal of Fish Biology, 58(3):682–700, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coughlin:2001:RTD

[338] D. J. Coughlin, J. Burdick, K. A. Stauffer, and F. E. Weaver. Rainbow trout display a developmental shift in red muscle kinetics, swimming kinematics and myosin heavy chain isoform. Journal of Fish Biology, 58(3):701–715, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gelineau:2001:OAC

[339] A. Gélineau and T. Boujard. Oral administration of cholecystokinin receptor antagonists increase feed intake in rainbow trout. Journal of Fish Biology, 58(3):716–724, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2001:IGR

[340] D. J. Jellyman. The influence of growth rate on the size of migrating female eels in Lake Ellesmere, New Zealand. Journal of Fish Biology, 58(3):725–736, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoehne-Reitan:2001:DBS

[341] K. Hoehne-Reitan, E. Kjørsvik, and D. R. Gjellesvik. Development of bile salt-dependent lipase in larval turbot. Journal of Fish Biology, 58 (3):737–745, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoehne-Reitan:2001:BSD

[342] K. Hoehne-Reitan, E. Kjørsvik, and K. I. Reitan. Bile salt-dependent lipase in larval turbot, as influenced by density and lipid content of fed prey. Journal of Fish Biology, 58(3):746–754, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nieder:2001:ABF

[343] J. Nieder. Amphibious behaviour and feeding ecology of the four-eyed blenny (Dialommus fuscus, Labrisomidae) in the intertidal zone of the island of Santa Cruz (Galapagos, Ecuador). Journal of Fish Biology, 58 (3):755–767, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paulsen:2001:SSD

[344] S. M. Paulsen, B. Sveinbjørnsson, and B. Robertsen. Selective staining and disintegration of intestinal eosinophilic granule cells in Atlantic salmon after intraperitoneal injection of the zinc chelator dithizone. Journal of Fish Biology, 58(3):768–775, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okuda:2001:CRM

[345] N. Okuda. The costs of reproduction to males and females of a paternal mouthbrooding cardinalfish Apogon notatus. Journal of Fish Biology, 58(3):776–787, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amara:2001:FEG

[346] R. Amara, P. Laffargue, J. M. Dewarumez, C. Maryniak, F. Lagardére, and C. Luzac. Feeding ecology and growth of O-group flatfish (sole, dab and plaice) on a nursery ground (Southern Bight of the North Sea). Journal of Fish Biology, 58(3):788–803, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamada:2001:MHC

[347] Y. Yamada, H. Zhang, A. Okamura, S. Tanaka, N. Horie, N. Mikawa, T. Utoh, and H. P. Oka. Morphological and histological changes in the swim bladder during maturation of the Japanese eel. Journal of Fish Biology, 58(3):804–814, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Susca:2001:SSM

[348] V. Susca, A. Corriero, C. R. Bridges, and G. De Metrio. Study of the sexual maturity of female bluefin tuna: purification and partial characterization of vitellogenin and its use in an enzyme-linked immunosorbent assay. Journal of Fish Biology, 58(3):815–831, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delariva:2001:RBM

[349] R. L. Delariva and A. A. Agostinho. Relationship between morphology and diets of six neotropical loricariids. Journal of Fish Biology, 58(3): 832–847, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2001:EAC

[350] M. B. Adams, M. D. Powell, and G. J. Purser. Effect of acute and chronic ammonia and nitrite exposure on oxygen consumption and growth of juvenile big bellied seahorse. Journal of Fish Biology, 58(3):848–860, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Park:2001:HMH

[351] J. Y. Park and I. S. Kim. Histology and mucin histochemistry of the gastrointestinal tract of the mud loach, in relation to respiration. Journal of Fish Biology, 58(3):861–872, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joyeux:2001:AVO

[352] J-C. Joyeux, C. Aliaume, and A. Zerbi. An alternative to validation of otolith microincrementation. Journal of Fish Biology, 58(3):873–879, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2001:OSS

[353] C. Smith, A. Douglas, and P. Jurajda. Oviposition site selection and embryo mortality in perch. Journal of Fish Biology, 58(3):880–882, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grankoto:2001:MDI

[354] A. Grankoto, C. Pampoulie, A. Marques, and G. N. Sakiti. Myxobolus dahomeyensis infection in ovaries of Tilapia species from Benin (West Africa). Journal of Fish Biology, 58(3):883–886, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Binz:2001:SAM

[355] T. Binz, T. B. H. Reusch, C. Wedekind, and M. Milinski. SSCP analysis of Mhc class II B genes in the threespine stickleback. Journal of Fish Biology, 58(3):887–890, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhou:2001:FMG

[356] R. Zhou, Q. Zhang, T. R. Tiersch, and R. K. Cooper. Four members of the Sox gene family in channel catfish. Journal of Fish Biology, 58 (3):891–894, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ab

[357] Anonymous. Announcements. Journal of Fish Biology, 58(3):895, March 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2001:ALS

[358] R. F. Oliveira, V. C. Almada, E. J. Gonçalves, E. Forsgren, and A. V. M. Canario. Androgen levels and social interactions in breeding males of the peacock blenny. Journal of Fish Biology, 58(4):897–908, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marino:2001:RDG

[359] G. Marino, E. Azzurro, A. Massari, M. G. Finoia, and A. Mandich. Reproduction in the dusky grouper from the southern Mediterranean. Journal of Fish Biology, 58(4):909–927, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Close:2001:EOC

[360] P. G. Close, C. G. Barlow, and L. J. Rodgers. Early ontogeny of coal grunter from hormone-induced spawnings and laboratory-reared embryos and larvae. Journal of Fish Biology, 58(4):928–942, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2001:RVE

[361] K. Oliveira, J. D. McCleave, and G. S. Wippelhauser. Regional variation and the effect of lake: river area on sex distribution of American eels. Journal of Fish Biology, 58(4):943–952, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aasbakk:2001:EFM

[362] K. Åsbakk. Elimination of foreign material by epidermal malpighian cells during wound healing in fish skin. Journal of Fish Biology, 58(4):953– 966, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neuman:2001:RBO

[363] M. J. Neuman, D. A. Witting, and K. W. Able. Relationships between otolith microstructure, otolith growth, somatic growth and ontogenetic transitions in two cohorts of windowpane. Journal of Fish Biology, 58(4): 967–984, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rocha:2001:BAR

[364] L. A. Rocha and I. L. Rosa. Baseline assessment of reef fish assemblages of Parcel Manuel Luiz Marine State Park, Maranhão, north-east Brazil. Journal of Fish Biology, 58(4):985–998, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2001:CGG

[365] S. Xie, X. Zhu, Y. Cui, R. J. Wootton, W. Lei, and Y. Yang. Compensatory growth in the gibel carp following feed deprivation: temporal patterns in growth, nutrient deposition, feed intake and body composition. Journal of Fish Biology, 58(4):999–1009, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Norton:2001:LCD

[366] E. C. Norton, R. B. MacFarlane, and M. S. Mohr. Lipid class dynamics during development in early life stages of shortbelly rockfish and their application to condition assessment. Journal of Fish Biology, 58(4):1010– 1024, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ytrestoyl:2001:SPB

[367] T. Ytrestøyl, B. Finstad, and R. S. McKinley. Swimming performance and blood chemistry in Atlantic salmon spawners exposed to acid river water with elevated aluminium concentrations. Journal of Fish Biology, 58(4):1025–1038, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vermeirssen:2001:MPP

[368] E. L. M. Vermeirssen and A. P. Scott. Male priming pheromone is present in bile, as well as urine, of female rainbow trout. Journal of Fish Biology, 58(4):1039–1045, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parmentier:2001:MAO

[369] E. Parmentier, P. Vandewalle, and F. Lagardère. Morpho-anatomy of the otic region in carapid fishes: eco-morphological study of their otoliths. Journal of Fish Biology, 58(4):1046–1061, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kohbara:2001:TRF

[370] J. Kohbara and J. Caprio. Taste responses of the facial and glossopharyngeal nerves to amino acids in the rainbow trout. Journal of Fish Biology, 58(4):1062–1072, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedersen:2001:HMS

[371] J. Pedersen. Hydroacoustic measurement of swimming speed of North Sea saithe in the field. Journal of Fish Biology, 58(4):1073–1085, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernandez-Diaz:2001:GPC

[372] C. Fernández-Díaz, M. Yýfera, J. P. Cañavate, F. J. Moyano, F. J. Alarcón, and M. Díaz. Growth and physiological changes during metamorphosis of Senegal sole reared in the laboratory. Journal of Fish Biology, 58(4):1086–1097, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broad:2001:MPP

[373] T. L. Broad, C. R. Townsend, C. J. Arbuckle, and D. J. Jellyman. A model to predict the presence of longfin eels in some New Zealand streams, with particular reference to riparian vegetation and elevation. Journal of Fish Biology, 58(4):1098–1112, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medioni:2001:GDV

[374] E. Medioni, R. Lecomte Finiger, N. Louveiro, and S. Planes. Genetic and demographic variation among colour morphs of cabrilla sea bass. Journal of Fish Biology, 58(4):1113–1124, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hurwood:2001:HID

[375] D. A. Hurwood and J. M. Hughies. Historical interdrainage dispersal of eastern rainbowfish from the atherton tableand, north-eastern Australia. Journal of Fish Biology, 58(4):1125–1136, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petersen:2001:DAB

[376] J. H. Petersen. Density, aggregation, and body size of northern pikeminnow preying on juvenile salmonids in a large river. Journal of Fish Biology, 58(4):1137–1148, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhu:2001:CCG

[377] X. Zhu, Y. Cui, M. Ali, and R. J. Wootton. Comparison of compensatory growth responses of juvenile three-spined stickleback and minnow following similar food deprivation protocols. Journal of Fish Biology, 58 (4):1149–1165, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waters:2001:MDP

[378] J. M. Waters and G. P. Wallis. Mitochondrial DNA phylogenetics of the Galaxias vulgaris complex from South Island, New Zealand: rapid radiation of a species flock. Journal of Fish Biology, 58(4):1166–1180, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez:2001:EEH

[379] A. P. Sanchez and M. L. Glass. Effects of environmental hypercapnia on pulmonary ventilation of the South American lungfish. Journal of Fish Biology, 58(4):1181–1189, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2001:GMI

[380] P. J. Smith, P. M. Gaffney, and M. Purves. Genetic markers for identification of Patagonian toothfish and Antarctic toothfish. Journal of Fish Biology, 58(4):1190–1194, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ac

[381] Anonymous. Announcements. Journal of Fish Biology, 58(4):1195, April 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fritzner:2001:UMM

[382] N. G. Fritzner, M. M. Hansen, S. S. Madsen, and K. Kristiansen. Use of micreosatellite markers for identification of indigenous brown trout in a geographical region heavily influenced by stocked domensticated trout. Journal of Fish Biology, 58(5):1197–1210, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Privitera:2001:CEL

[383] L. A. Privitera. Characteristics of egg and larval production in captive bluespotted gobies. Journal of Fish Biology, 58(5):1211–1220, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baker:2001:SSA

[384] C. F. Baker and J. C. Montgomery. Species-specific attraction of migratory banded kokopu juveniles to adult pheromones. Journal of Fish Biology, 58(5):1221–1229, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temming:2001:GEHa

[385] A. Temming and J.-P. Herrmann. Gastric evacuation in horse mackerel. I. The effects of meal size, temperature and predator weight. Journal of Fish Biology, 58(5):1230–1245, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temming:2001:GEHb

[386] A. Temming and J.-P. Herrmann. Gastric evacuation of horse mackerel. II. The effects of different prey types on the evacuation model. Journal of Fish Biology, 58(5):1246–1256, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fey:2001:DTC

[387] D. P. Fey. Differences in temperature conditions and somatic growth rate of larval and early juvenile spring-spawned herring from the vistula lagoon, Baltic Sea manifested in the otolith to fish size relationship. Journal of Fish Biology, 58(5):1257–1273, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ponton:2001:UOM

[388] D. Ponton, J. H. Mol, and J. Panfili. Use of otolith microincrements for estimating the age and growth of young armoured catfish Hoplosternum littorale. Journal of Fish Biology, 58(5):1274–1285, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoysak:2001:FDS

[389] D. J. Hoysak and N. R. Liley. Fertilization dynamics in sockeye salmon and a comparison of sperm from alternative male phenotypes. Journal of Fish Biology, 58(5):1286–1300, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salles:2001:BCL

[390] C. Salles, R. H. S. Souza, J. B. Salles, V. L. F. Cunha Bastos, R. A. Dias, and J. Cunha Bastos. Biochemical characteristics of liver and brain monoamine oxidase from pacu. Journal of Fish Biology, 58(5):1301– 1310, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaroensutasinee:2001:BNH

[391] M. Jaroensutasinee and K. Jaroensutansinee. Bubble nest habitat characteristics of wild Siamese fighting fish. Journal of Fish Biology, 58(5):

1311–1319, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:2001:DEF

[392] M. C. Barber and G. P. Jenkins. Differential effects of food and temperature lead to decoupling of short-term otolith and somatic growth rates in juvenile King George whiting. Journal of Fish Biology, 58(5):1320– 1330, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maclean:2001:SSA

[393] A. Maclean and N. B. Metcalfe. Social status, access to food, and compensatory growth in juvenile Atlantic salmon. Journal of Fish Biology, 58(5):1331–1346, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2001:OBE

[394] A. C. Henderson, K. Flannery, and J. Dunne. Observations on the biology and ecology of the blue shark in the North-east Atlantic. Journal of Fish Biology, 58(5):1347–1358, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silberschneider:2001:NAH

[395] V. Silberschneider, B. C. Pease, and D. J. Booth. A novel artificial habitat collection device for studying resettlement patterns in anguillid glass eels. Journal of Fish Biology, 58(5):1359–1370, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katano:2001:NEA

[396] O. Katano and Y. Aonuma. Negative effect of ayu on the growth of omnivorous pale chub in experimental pools. Journal of Fish Biology, 58(5):1371–1382, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kupschus:2001:ABF

[397] S. Kupschus and D. Tremain. Associations between fish assemblages and environmental factors in nearshore habitats of a subtropical estuary. Journal of Fish Biology, 58(5):1383–1403, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dePedro:2001:FHC

[398] N. de Pedro, M. J. Delgado, and M. Alonso-Bedate. Fasting and hypothalamic catecholamines in goldfish. Journal of Fish Biology, 58(5):

1404–1413, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ewing:2001:CGK

[399] R. D. Ewing, G. S. Ewing, and T. D. Satterthwaite. Changes in gill Na^{+}, K^{+}-ATPase specific activity during seaward migration of wild juvenile Chinook salmon. Journal of Fish Biology, 58(5):1414–1426, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welch:2001:VMF

[400] D. W. Welch and P. M. Pankhurst. Visual morphology and feeding behaviour of the daggertooth. Journal of Fish Biology, 58(5):1427–1437, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lau:2001:GSS

[401] P. P. F. Lau and Y. Sadovy. Gonad structure and sexual pattern in two threadfin breams and possible function of the dorsal accessory duct. Journal of Fish Biology, 58(5):1438–1453, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aedo:2001:GER

[402] G. Aedo and H. Arancibia. Gastric evacuation of the redspotted catshark under laboratory conditions. Journal of Fish Biology, 58(5):1454–1457, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kraak:2001:INH

[403] S. B. M. Kraak, B. Mundwiler, and P. J. B. Hart. Increased number of hybrids between benthic and limnetic three-spined sticklebacks in Enos Lake, Canada; the collapse of a species pair? Journal of Fish Biology, 58(5):1458–1464, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2001:BRF

[404] G. E. Brown, J. C. Adrian, Jr., and M. L. Shih. Behavioural responses of fathead minnows to hypoxanthine-3-N-oxide at varying concentrations. Journal of Fish Biology, 58(5):1465–1470, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zydlewski:2001:PSP

[405] G. Barbin Zydlewski, A. Haro, K. G. Whalen, and S. D. McCormick. Performance of stationary and portable passive transponder detection systems for monitoring of fish movements. Journal of Fish Biology, 58 (5):1471–1475, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wildhaber:2001:TBF

[406] M. L. Wildhaber. The trade-off between food and temperature in the habitat choice of bluegill sunfish. Journal of Fish Biology, 58(5):1476– 1478, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:BRb

[407] Anonymous. Book reviews. Journal of Fish Biology, 58(5):1479–1485, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:BRc

[408] Anonymous. Books received. Journal of Fish Biology, 58(5):1486, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:EISa

[409] Anonymous. Eighth international symposium on the ecology of fluvial fishes. Journal of Fish Biology, 58(5):1487, May 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alves-Gomes:2001:EEB

[410] J. A. Alves-Gomes. The evolution of electroreception and bioelectrogenesis in teleost fish: a phylogenetic perspective. Journal of Fish Biology, 58(6):1489–1511, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romare:2001:EHE

[411] P. Romare. An evaluation of horizontal echo sounding as a method for behavioural studies of 0+ fish in field experiments. Journal of Fish Biology, 58(6):1512–1523, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2001:DDR

[412] J. D. Armstrong and S. W. Griffiths. Density-dependent refuge use among over-wintering wild Atlantic salmon juveniles. Journal of Fish Biology, 58(6):1524–1530, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ali:2001:CGC

[413] M. Ali and R. J. Wootton. Capacity for growth compensation in juvenile three-spined sticklebacks experiencing cycles of food deprivation. Journal of Fish Biology, 58(6):1531–1544, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Healey:2001:PGI

[414] M. C. Healey. Patterns of gametic investment by female stream-and ocean-type Chinook salmon. Journal of Fish Biology, 58(6):1545–1556, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barni:2001:RAM

[415] A. Barni, C. Mazzoldi, and M. B. Rasotto. Reproductive apparatus and male accessory structures in two batrachoid species (Teleostei, Batrachoididae). Journal of Fish Biology, 58(6):1557–1569, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oppedal:2001:ALS

[416] F. Oppedal, J-E. Juell, G. L. Tarranger, and T. Hansen. Artificial light and season affects vertical distribution and swimming behaviour of postsmolt Atlantic salmon in sea cages. Journal of Fish Biology, 58(6):1570– 1584, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saillant:2001:PIE

[417] E. Saillant, B. Chatain, A. Fostier, C. Przybyla, and C. Fauvel. Parental influence on early development in the European sea bass. Journal of Fish Biology, 58(6):1585–1600, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Asoh:2001:GDN

[418] K. Asoh, T. Yoshikawa, and M. Kasuya. Gonadal development and nonfunctional protogyny in a coral-reef damselfish, Dascyllus albisella Gill. Journal of Fish Biology, 58(6):1601–1616, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva:2001:MPM

[419] M. E. Oliva. Metazoan parasites of Macruronus magellanicus from southern Chile as biological tags. Journal of Fish Biology, 58(6):1617–1622, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quagio-Grassiotto:2001:USS

[420] I. Quagio-Grassiotto, C. Oliveria, and A. E. Gosztonyi. The ultrastructure of spermiogenesis and spermatozoa in Diplomystes mesembrinus. Journal of Fish Biology, 58(6):1623–1632, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zwolinski:2001:IAV

[421] J. Zwolinski, Y. Stratoudakis, and E. Sares. Intra-annual variation in the batch fecundity of sardine off Portugal. Journal of Fish Biology, 58 (6):1633–1645, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takeyama:2001:MDS

[422] H. Takeyama, S. Chow, H. Tsuzuki, and T. Matunaga. Mitochondrial DNA sequence variation within and between tuna Thunnus species and its application to species identification. Journal of Fish Biology, 58(6): 1646–1657, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vila-Gispert:2001:FCA

[423] A. Vila-Gispert and R. Moreno-Amich. Fish condition analysis by a weighted least squares procedure: testing geographical differences of an endangered Iberian cyprinodontid. Journal of Fish Biology, 58(6):1658– 1666, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andreassen:2001:FPS

[424] P. M. R. Andreassen, M. B. Martinussen, N. A. Hvidsten, and S. O. Stefansson. Feeding and prey-selection of wild Atlantic salmon postsmolts. Journal of Fish Biology, 58(6):1667–1679, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez:2001:PIR

[425] R. J. Gonzalez and R. W. Wilson. Patterns of ion regulation in acidophilic fish native to the ion-poor, acidic Rio Negro. Journal of Fish Biology, 58(6):1680–1690, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ishizaki:2001:EDD

[426] Y. Ishizaki, R. Masuda, K. Uematsu, K. Shimizu, M. Arimoto, and T. Takeuchi. The effect of dietary docosahexaenoic acid on schooling behaviour and brain development in larval yellowtail. Journal of Fish Biology, 58(6):1691–1703, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gallagher:2001:CUG

[427] M. L. Gallagher, J. J. Luczkovich, and E. J. Stellwag. Characterization of the ultrastructure of the gastrointestinal tract mucosa, stomach contents and liver enzyme activity of the pinfish during development. Journal of Fish Biology, 58(6):1704–1713, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2001:DMR

[428] S. Xie, Y. Cui, and Z. Li. Dietary-morphological relationships of fishes in Liangzi Lake, China. Journal of Fish Biology, 58(6):1714–1729, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansson:2001:SAB

[429] S. Hansson, L. Karlsson, E. Ikonen, O. Christensen, A. Mitans, D. Uzars, E. Petersson, and B. Ragnarsson. Stomach analyses of Baltic salmon from 1959–1962 and 1994–1997: possible relations between diet and yolk-sac-fry mortality (M74). Journal of Fish Biology, 58(6):1730–1745, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Park:2001:SDM

[430] I-S. Park, C. I. Zhang, and Y-D. Lee. Sexual dimorphism in morphometric characteristics of cocktail wrasse. Journal of Fish Biology, 58 (6):1746–1749, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giulianini:2001:CGS

[431] P. G. Giulianini, D. Ota, M. Marchesan, and E. A. Ferrero. Can goby spermatozoa pass through the filament adhesion apparatus of laid eggs? Journal of Fish Biology, 58(6):1750–1752, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koed:2001:LTE

[432] A. Koed and E. B. Thorstad. Long-term effect of radio-tagging on the swimming performance of pikeperch. Journal of Fish Biology, 58(6): 1753–1756, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nocita:2001:TIF

[433] A. Nocita and S. Vanni. Types of the Italian freshwater goby Padogobius nigricans (Canestrini, 1867). Journal of Fish Biology, 58(6):1757–1760, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nagashima:2001:AFS

[434] Y. Nagashima, A. Sendo, K. Shimakura, K. Shiomi, T. Kobayashi, B. Kimura, and T. Fujii. Antibacterial factors in skin mucus of rabbitfishes. Journal of Fish Biology, 58(6):1761–1765, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:BRd

[435] Anonymous. Book reviews. Journal of Fish Biology, 58(6):1766–1767, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:EISb

[436] Anonymous. Eighth international symposium on the ecology of fluvial fishes. Journal of Fish Biology, 58(6):1770, June 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruane:2001:PCM

[437] N. M. Ruane, E. A. Huisman, and J. Komen. Plasma cortisol and metabolite level profiles in two isogenic strains of common carp during confinement. Journal of Fish Biology, 59(1):1–12, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aggergaard:2001:CCP

[438] S. Aggergaard and F. B. Jensen. Cardiovascular changes and physiological response during nitrite exposure in rainbow trout. Journal of Fish Biology, 59(1):13–27, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nielsen:2001:GDP

[439] C. Nielsen, G. Holdensgaard, H. C. Petersen, B. Th. Björnsson, and S. S. Madsen. Genetic differences in physiology, growth hormone levels and migratory behaviour of Atlantic salmon smolts. Journal of Fish Biology, 59(1):28–44, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keckeis:2001:SDF

[440] H. Keckeis, E. Kamler, E. Bauer-Nemeschkal, and K. Schneeweiss. Survival, development and food energy partitioning of nase larvae and early juveniles at different temperatures. Journal of Fish Biology, 59(1):45– 61, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahman:2001:GNT

[441] M. A. Rahman, A. Ronyai, B. Z. Engidaw, K. Jauncey, G-L. Hwang, A. Smith, E. Roderick, D. Penman, L. Varadi, and N. Maclean. Growth and nutritional trials on transgenic Nile tilapia containing an exogenous fish growth hormone gene. Journal of Fish Biology, 59(1):62–78, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cole:2001:HPR

[442] M. B. Cole, D. E. Arnole, B. J. Watten, and W. F. Krise. Haematological and physiological responses of brook charr, to untreated and limestoneneutralized acid mine drainage. Journal of Fish Biology, 59(1):79–91, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chernova:2001:TNS

[443] N. V. Chernova and J. T. Eastman. Two new species of snailfish genus Paraliparis (Pisces: Liparidae) from the Ross Sea, Antarctica. Journal of Fish Biology, 59(1):92–104, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2001:CPJ

[444] K. A. Glover, J. B. Taggart, Ø. Skalla, and A. J. Teale. Comparative performance of juvenile sea trout families in high and low feeding environments. Journal of Fish Biology, 59(1):105–115, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burrows:2001:DSB

[445] M. T. Burrows. Depth selection behaviour during activity cycles of juvenile plaice on a simulated beach slope. Journal of Fish Biology, 59(1): 116–125, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayward:2001:FIC

[446] R. S. Hayward and N. Wang. Failure to induce over-compensation of growth in maturing yellow perch. Journal of Fish Biology, 59(1):126– 140, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tittu:2001:CAS

[447] V. Tittu and M. Vornanen. Cold adaptation suppresses the contractility of both atrial and ventricular muscle of the crucian carp heart. Journal of Fish Biology, 59(1):141–156, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Very:2001:SIG

[448] N. M. Very, D. Knutson, J. D. Kittilson, and M. A. Sheridan. Somatostatin inhibits growth of rainbow trout. Journal of Fish Biology, 59(1): 157–165, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2001:FRP

[449] D. W. Evans, M. A. Matthews, and C. A. McClintock. First record of Pomphorhynchus laevis (Acanthocephala) in fishes from Northern Ireland. Journal of Fish Biology, 59(1):166–168, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pujolar:2001:ADB

[450] J. M. Pujolar, M. I. Roldán, and C. Pla. Allozyme differentiation of bonito in the Mediterranean Sea. Journal of Fish Biology, 59(1):169– 174, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tveiten:2001:TIT

[451] H. Tveiten and H. K. Johnsen. Thermal influences on temporal changes in plasma testosterone and oestradiol-17β concentrations during gonadal recrudescence in female common wolffish. Journal of Fish Biology, 59 (1):175–178, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pallaoro:2001:FRS

[452] A. Pallaoro and J. Dulčić. First record of the Sphyraena chrysotaenia (Klunzinger, 1884) (Pisces, Sphyraenidae) from the Adriatic Sea. Journal of Fish Biology, 59(1):179–182, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2001:PTT

[453] S. Adams and A. J. Williams. A preliminary test of the transitional growth spurt hypothesis using the protogynous coral trout Plectropomus maculatus. Journal of Fish Biology, 59(1):183–185, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloman:2001:SAE

[454] K. A. Sloman, A. C. Taylor, N. B. Metcalfe, and K. M. Gilmour. Stress from air emersion fails to alter chloride cell numbers in the gills of rainbow trout. Journal of Fish Biology, 59(1):186–190, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:BRe

[455] Anonymous. Book reviews. Journal of Fish Biology, 59(1):191–194, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:XICa

[456] Anonymous. XXVII International Council of Ethologists — Ethologische Gesellschaft E.v. Journal of Fish Biology, 59(1):195, July 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campana:2001:APQ

[457] S. E. Campana. Accuracy, precision and quality control in age determination, including a review of the use and abuse of age validation methods. Journal of Fish Biology, 59(2):197–242, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveria:2001:OMC

[458] C. Oliveria, R. Diogo, P. Vandewalle, and M. Chardon. Osteology and myology of the cephalic region and pectoral girdle of Plotosus lineatus, with comments on Plotosidae (Teleostei: Siluriformes) autapomorphies. Journal of Fish Biology, 59(2):243–266, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almansa:2001:LFA

[459] E. Almansa, M. V. Martían, J. R. Cejas, P. Badí, S. Jerez, and A. Lorenzo. Lipid and fatty acid composition of female gilthead seabream during their reproductive cycle: effects of a diet lacking n-3 HUFA. Journal of Fish Biology, 59(2):267–286, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mezzera:2001:ESA

[460] M. Mezzera and C. R. Largiadèr. Evidence for selective angling of introduced trout and their hybrids in a stocked brown trout population. Journal of Fish Biology, 59(2):287–301, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sbaihi:2001:RBC

[461] M. Sbaihi, M. Fouchereau-Peron, F. Meunier, P. Elie, I. Mayer, E. Burzawa-Gerard, B. Vidal, and S. Dufour. Reproductive biology of the conger eel from the south coast of Brittany, France and comparison with the Europe eel. Journal of Fish Biology, 59(2):302–318, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crossland:2001:APN

[462] M. R. Crossland. Ability of predatory native Australian fishes to learn to avoid toxic larvae of the introduced toad Bufo marinus. Journal of Fish Biology, 59(2):319–329, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rorvik:2001:BTS

[463] K-A. Rørvik, P. O. Skjervold, S. O. Fjæra, T. Mørkøre, and S. H. Steien. Body temperature and seawater adaptation in farmed Atlantic salmon and rainbow trout during prolonged chilling. Journal of Fish Biology, 59(2):330–337, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stehmann:2001:FRA

[464] M. F. W. Stehmann and N. R. Merrett. First records of advanced embryos and egg capsules of Bathyraja skates from the deep north-eastern Atlantic. Journal of Fish Biology, 59(2):338–349, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillanders:2001:TEO

[465] B. M. Gillanders, P. Sanchez-Jerez, J. Bayle-Sempere, and A. Ramos-Espla. Trace elements in otoliths of the two-banded bream from a coastal region in the south-west Mediterranean: are there differences among locations? Journal of Fish Biology, 59(2):350–363, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ccottrill:2001:PNE

[466] R. A. Ccottrill, R. S. McKinley, G. Van Der Kaak, J-D. Dutil, K. B. Reid, and K. J. McGrath. Plasma non-esterified fatty acid profiles and 17βoestradiol levels of juvenile immature and maturing adult American eels in the st Lawrence River. Journal of Fish Biology, 59(2):364–379, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedersen:2001:SVE

[467] J. Pedersen and J. R. G. Hislop. Seasonal variations in the energy density of fishes in the North Sea. Journal of Fish Biology, 59(2):380–389, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fievet:2001:DMF

[468] E. Fièvet, S. Dolédec, and P. Lim. Distribution of migratory fishes and shrimps along multivariate gradients in tropical island streams. Journal of Fish Biology, 59(2):390–402, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramirez:2001:GNA

[469] T. Ramírez, D. Cortés, and A. Garcíaa. Growth of North Alboran Sea sardine larvae estimated by otolith microstructure, nucleic acids and protein content. Journal of Fish Biology, 59(2):403–415, August 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Foss:2001:ERS

[470] A. Foss, T. H. Evensen, A. K. Imsland, and V. Øiestad. Effects of reduced salinities on growth, food conversion efficiency and osmoregulatory status in the spotted wolffish. Journal of Fish Biology, 59(2):416–426, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez:2001:ILS

[471] J. G. Sánchez, D. J. Speare, R. J. F. Markham, and S. R. M. Jones. Isolation of a Loma salmonae variant: biological characteristics and host range. Journal of Fish Biology, 59(2):427–441, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez:2001:EVR

[472] J. G. Sánchez, D. J. Speare, R. J. F. Markham, and S. R. M. Jones. Experimental vaccination of rainbow trout against Loma salmonae using a live low-virulence variant of L. salmonae. Journal of Fish Biology, 59 (2):442–448, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsland:2001:VFI

[473] A. K. Imsland, A. Foss, and S. O. Stefansson. Variation in food intake, food conversion efficiency and growth of juvenile turbot from different geographic strains. Journal of Fish Biology, 59(2):449–454, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Panserat:2001:DFD

[474] S. Panserat, E. Capilla, J. Gutierrez, C. Vachot, E. Plagnes-Juan, P. Aguirre, J. Brèque, and S. Kaushik. Dietary fructose does not specifically induce hepatic glucokinase expression in rainbow trout. Journal of Fish Biology, 59(2):455–458, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2001:GIY

[475] R. F. Oliveira, E. J. Gonçalves, and R. S. Santos. Gonadal investment of young males in two blenniid fishes with alternative mating tactics. Journal of Fish Biology, 59(2):459–462, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romao:2001:IRK

[476] S. Romão, C. A. Freire, and E. Fanta. Ionic regulation and Na^{+}, K^{+}- ATPase activity in gills and kidney of the Antarctic aglomerular cod icefish exposed to dilute sea water. Journal of Fish Biology, 59(2):463– 468, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:XICb

[477] Anonymous. XXVII International Council of Ethologists — Ethologische Gesellschaft E.v. Journal of Fish Biology, 59(2):469–470, August 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2001:SLL

[478] C. Brown and K. Laland. Social learning and life skills training for hatchery reared fish. Journal of Fish Biology, 59(3):471–493, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinnegar:2001:RPU

[479] J. K. Pinnegar, N. Campbell, and N. V. C. Polunin. Regular papers: Unusual stable isotope fractionation patterns observed for fish host — parasite trophic relationships. Journal of Fish Biology, 59(3):494–503, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Espmark:2001:IBA

[480] Y. Espmark and T. Knudsen. Intraspecific brood adoption in the convict cichlid with respect to fry of two colour morphs. Journal of Fish Biology, 59(3):504–514, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kahl:2001:EHE

[481] M. D. Kahl, K. M. Jensen, J. J. Korte, and G. T. Ankley. Effects of handling on endocrinology and reproductive performance of the fathead minnow. Journal of Fish Biology, 59(3):515–523, September 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2001:FTB

[482] T-C. Chen, R. F. G. Ormond, and H-K. Mok. Feeding and territorial behaviour in juveniles of three co-existing triggerfishes. Journal of Fish Biology, 59(3):524–532, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knutsen:2001:FAB

[483] J. A. Knutsen, H. Knutsen, J. Gjøsæter, and B. Jonsson. Food of anadromous brown trout at sea. Journal of Fish Biology, 59(3):533–543, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ashford:2001:EPP

[484] B. D. Ashford and R. G. Danzmann. An examination of the post-planting success of hatchery brook charr possessing different mitochondrial DNA haplotypes. Journal of Fish Biology, 59(3):544–554, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niemela:2001:SVD

[485] E. Niemelä, M. Julkunen, J. Erkinaro, and T. S. Mäukinen. Seasonal variation in density of juvenile Atlantic salmon in fluvial shoreline habitats of a large subarctic river. Journal of Fish Biology, 59(3):555–568, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2001:IEL

[486] Y. Zhao, Y. Chen, and J. A. Brown. Impacts of egg and larval size on survival and growth of Atlantic cod under different feeding conditions. Journal of Fish Biology, 59(3):569–581, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzi:2001:ASI

[487] D. Mazzi and T. C. M. Bakker. Acid stress increases pelvic spine asymmetry in juvenile three-spined sticklebacks. Journal of Fish Biology, 59 (3):582–592, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2001:MDR

[488] M. I. Taylor and E. Verheyen. Microsatellite data reveals weak population substructuring in Copadichromis sp. ‘virginalis kajose’, a demersal cichlid from Lake Malaŵi, Africa. Journal of Fish Biology, 59(3):593–604, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mercure:2001:IPL

[489] F. Mercure, A. C. Holloway, D. R. Tocher, M. A. Sheridan, G. Van der Kraak, and J. F. Leatherland. Influence of plasma lipid changes in response to 17β-oestradiol stimulation on plasma growth hormone, somatostatin, and thyroid hormone levels in immature rainbow trout. Journal of Fish Biology, 59(3):605–615, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linhart:2001:STS

[490] O. Linhart, D. Gela, M. Rodina, and M. Rodriguez Gutierrez. Shortterm storage of ova of common carp and tench in extenders. Journal of Fish Biology, 59(3):616–623, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2001:RCG

[491] H. H. Ng and J. Freyhof. A review of the catfish genus Pterocryptis (Siluridae) in Vietnam, with the description of two new species. Journal of Fish Biology, 59(3):624–644, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamura:2001:SCY

[492] O. Yamamura, K. Yabuki, O. Shida, K. Watanabe, and S. Honda. Spring cannibalism on 1 year walleye pollock in the Doto area, northern Japan: is it density dependent? Journal of Fish Biology, 59(3):645–656, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2001:RBO

[493] P. J. Wright, P. Fallon-Cousins, and J. D. Armstrong. The relationship between otolith accretion and resting metabolic rate in juvenile Atlantic salmon during a change in temperature. Journal of Fish Biology, 59(3): 657–666, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ccouldridge:2001:DTS

[494] V. C. K. Ccouldridge and G. J. Alexander. Does the time spent near a male predict female mate choice in a malaŵian cichlid? Journal of Fish Biology, 59(3):667–672, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradner:2001:BCI

[495] J. Bradner and S. P. McRobert. Background colouration influences body colour segregation in mollies. Journal of Fish Biology, 59(3):673–681, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Limburg:2001:FMA

[496] K. E. Limburg, P. Landergren, L. Westin, M. Elfman, and P. Kristiansson. Flexible modes of anadromy in Baltic sea trout: making the most of marginal spawning streams. Journal of Fish Biology, 59(3):682–695, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peach:2001:DLP

[497] M. B. Peach. The dorso-lateral pit organs of the Port Jackson shark contribute sensory information for rheotaxis. Journal of Fish Biology, 59 (3):696–704, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

David:2001:CRM

[498] B. O. David and G. P. Closs. Continuous remote monitoring of fish activity with restricted home ranges using radiotelemetry. Journal of Fish Biology, 59(3):705–715, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mirza:2001:DRM

[499] R. S. Mirza, J. J. Scott, and D. P. Chivers. Differential responses of male and female red swordtails to chemical alarm cues. Journal of Fish Biology, 59(3):716–728, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cardona:2001:NCC

[500] L. Cardona. Non-competitive coexistence between Mediterranean grey mullet: evidence from seasonal changes in food availability, niche breadth and trophic overlap. Journal of Fish Biology, 59(3):729–744, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skov:2001:MPF

[501] C. Skov, P. Grønkjær, and C. Nielsen. Marking pike fry otoliths with alizarin complexone and strontium: an evaluation of methods. Journal of Fish Biology, 59(3):745–750, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:2001:FSD

[502] N. R. Merrett and J. G. Nielsen. Fourth specimen of the deep benthopelagic ophidiid fish, Apagesoma delosommatus (Hureau, Staiger & Nielsen, 1979). Journal of Fish Biology, 59(3):751–753, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herbert:2001:EGH

[503] N. A. Herbert, J. D. Armstrong, and B. Th. Björnsson. Evidence that growth hormone-induced elevation in routine metabolism of juvenile Atlantic salmon is a result of increased spontaneous activity. Journal of Fish Biology, 59(3):754–757, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:BRf

[504] Anonymous. Book reviews. Journal of Fish Biology, 59(3):758–762, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Eb

[505] Anonymous. Erratum. Journal of Fish Biology, 59(3):763, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ad

[506] Anonymous. Announcements. Journal of Fish Biology, 59(3):764–765, September 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vonHerbing:2001:DFS

[507] I. Hunt von Herbing. Development of feeding structures in larval fish with different life histories: winter flounder and Atlantic cod. Journal of Fish Biology, 59(4):767–782, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Link:2001:RBS

[508] J. S. Link and J. Burnett. The relationship between stomach contents and maturity state for major northwest Atlantic fishes: new paradigms?

Journal of Fish Biology, 59(4):783–794, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

KovacIc:2001:BRG

[509] M. KovačIć. The biology of Roule’s goby in the Kvarner area, northern Adriatic Sea. Journal of Fish Biology, 59(4):795–809, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haney:2001:APR

[510] D. C. Haney, J. C. Vokoun, and D. B. Noltie. Alarm pheromone recognition in a Missouri darter assemblage. Journal of Fish Biology, 59(4): 810–817, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:2001:OCE

[511] A. Drud Jordan, M. Jungersen, and J. F. Steffensen. Oxygen consumption of East Siberian cod: no support for the metabolic cold adaptation theory. Journal of Fish Biology, 59(4):818–823, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vila-Gispert:2001:MLR

[512] A. Vila-Gispert and R. Moreno-Amich. Mass-length relationship of Mediterranean barbel as an indicator of environmental status in Southwest European stream ecosystems. Journal of Fish Biology, 59(4):824– 832, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lester:2001:MSS

[513] R. J. G. Lester, C. Thompson, H. Moss, and S. C. Barker. Movement and stock structure of narrow-barred Spanish mackerel as indicated by parasites. Journal of Fish Biology, 59(4):833–842, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watt:2001:FCG

[514] P. J. Watt, A. J. Shohet, and K. Renshaw. Female choice for good genes and sex-biased broods in guppies. Journal of Fish Biology, 59(4):843– 850, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fujioka:2001:TSD

[515] Y. Fujioka. Thermolabile sex determination in honmoroko. Journal of Fish Biology, 59(4):851–861, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okland:2001:RMA

[516] F. Økland, J. Erkinaro, K. Moen, E. Niemelä, P. Fiske, R. S. McKinley, and E. B. Thorstab. Return migration of Atlantic salmon in the River Tana: Phases of migratory behaviour. Journal of Fish Biology, 59(4): 862–874, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pichavant:2001:CEL

[517] K. Pichavant, J. Person-Le-Ruyet, N. Le Bayon, A. Severe, A. Le Roux, and G. Boeuf. Comparative effects of long-term hypoxia on growth, feeding and oxygen consumption in juvenile turbot and European sea bass. Journal of Fish Biology, 59(4):875–883, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurel:2001:BGS

[518] B. J. Laurel, J. A. Brown, and R. Anderson. Behaviour, growth and survival of redfish larvae in relation to prey availability. Journal of Fish Biology, 59(4):884–901, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gwak:2001:DCR

[519] W. S. Gwak and M. Tanaka. Developmental change in RNA: DNA ratios of fed and starved laboratory-reared Japanese flounder larvae and juveniles, and its application to assessment of nutritional condition for wild fish. Journal of Fish Biology, 59(4):902–915, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Howland:2001:CHO

[520] K. L. Howland, W. M. Tonn, and G. Goss. Contrasts in the hypoosmoregulatory abilities of a freshwater and an anadromous population of inconnu. Journal of Fish Biology, 59(4):916–927, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Radtke:2001:TSV

[521] R. L. Radtke, R. A. Kinzie III, and D. J. Shafer. Temporal and spatial variation in length of larval life and size at settlement of the Hawaiian amphidromous goby Lentipes concolor. Journal of Fish Biology, 59(4): 928–938, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grenouillet:2001:JFM

[522] G. Grenouillet and D. Pont. Juvenile fishes in macrophyte beds: influence of food resources, habitat structure and body size. Journal of Fish Biology, 59(4):939–959, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradner:2001:ESS

[523] J. Bradner and S. P. McRobert. The effect of shoal size on patterns of body colour segregation in mollies. Journal of Fish Biology, 59(4):960– 967, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rincon:2001:IAR

[524] P. A. Rincón and G. D. Grossman. Intraspecific aggression in rosyside dace, a drift-feeding stream cyprinid. Journal of Fish Biology, 59(4):968– 986, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belangerz:2001:HBS

[525] G. Bélangerz and M. A. Rodríguez. Homing behaviour of stream-dwelling brook charr following experimental displacement. Journal of Fish Biology, 59(4):987–1001, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mccarthy:2001:CAR

[526] I. D. Mccarthy. Competitive ability is related to metabolic asymmetry in juvenile rainbow trout. Journal of Fish Biology, 59(4):1002–1014, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2001:CDG

[527] I. Jones, P. Kille, and G. Sweeney. Cadmium delays growth hormone expression during rainbow trout development. Journal of Fish Biology, 59 (4):1015–1022, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2001:MTB

[528] J. D. Armstrong, S. Einum, I. A. Fleming, and P. Rycroft. A method for tracking the behaviour of mature and immature salmon parr around nests during spawning. Journal of Fish Biology, 59(4):1023–1032, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grift:2001:MBB

[529] R. E. Grift, A. D. Buljse, J. G. P. Klein Breteler, W. L. T. van Densen, M. A. M. Machiels, and J. J. M. Backx. Migration of bream between the main channel and floodplain lakes along the lower River Rhine during the connection phase. Journal of Fish Biology, 59(4):1033–1055, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Premdas:2001:EOT

[530] P. D. Premdas, C. D. Metcalfe, and S. Brown. The effects of 17βoestradiol, testosterone and tamoxifen on the development of papillomata in Catostomus commersoni. Journal of Fish Biology, 59(4):1056–1069, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Museth:2001:ELI

[531] J. Museth. Effects of Ligula intestinalis on habitat use, predation risk and catchability in European minnows. Journal of Fish Biology, 59(4): 1070–1080, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dannewit:2001:ABG

[532] J. Dannewit and E. Petersson. Association between growth, body condition and anti-predator behaviour in maturing and immature brown trout parr. Journal of Fish Biology, 59(4):1081–1091, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorstad:2001:LTN

[533] E. B. Thorstad, F. Økland, and T. G. Heggberget. Are long term negative effects from external tags underestimated? Fouling of an externally attached telemetry transmitter. Journal of Fish Biology, 59(4):1092–1094, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frankel:2001:ICP

[534] J. S. Frankel. Inheritance of caudal peduncle banding in the spike-tailed paradisefish. Journal of Fish Biology, 59(4):1095–1097, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooper:2001:VRR

[535] A. Cooper, E. W. Taylor, and T. Wang. Volume regulation by red blood cells from brown trout. Journal of Fish Biology, 59(4):1098–1103, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2001:BCF

[536] K. Morita. Back-calculation of fork length of white-spotted charr from scales: a comparison between major and minor axes measurements. Journal of Fish Biology, 59(4):1104–1107, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:EIF

[537] Anonymous. European Inland Fisheries Advisory Commission (EIFAC) Symposium on Inland Fisheries Management and the Aquatic Environment. Journal of Fish Biology, 59(4):1108–1109, October 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Quiroas:2001:DBV

[538] R. González-Quiróas and R. Anadóan. Diet breadth variability in larval blue whiting as a response to plankton size structure. Journal of Fish Biology, 59(5):1111–1125, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2001:WMH

[539] R. S. Brown, G. Power, and S. Beltaoa. Winter movements and habitat use of riverine brown trout, white sucker and common carp in relation to flooding and ice break-up. Journal of Fish Biology, 59(5):1126–1141, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Altinokand:2001:EBW

[540] I. Altinokand and J. M. Grizzle. Effects of brackish water on growth, feed conversion and energy absorption efficiency by juvenile euryhaline and freshwater stenohaline fishes. Journal of Fish Biology, 59(5):1142–1152, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gordoa:2001:CBM

[541] A. Gordoa, N. Raventóas, and F. Dealie. Comparison between micro- and macro-structure readings in the age estimations of Cape hake. Journal of Fish Biology, 59(5):1153–1163, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2001:OIN

[542] H. H. Ng and J. Freyhof. Oreoglanis infulatus, a new species of glyptosternine catfish (Siluriformes: Sisoridae) from central Vietnam. Journal of Fish Biology, 59(5):1164–1169, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munshi:2001:SEM

[543] J. S. D. Munshi, K. R. Olson, P. K. Roy, and U. Ghosh. Scanning electron microscopy of the heart of the climbing perch. Journal of Fish Biology, 59(5):1170–1180, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cobcroft:2001:ELI

[544] J. M. Cobcroft, P. M. Pankhurst, P. R. Hart, and S. C. Battalene. The effects of light intensity and algae-induced turbidity on feeding behaviour of larval striped trumpeter. Journal of Fish Biology, 59(5):1181–1197, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2001:GEM

[545] N. G. Andersen. A gastric evacuation model for three predatory gadoids and implications of using pooled field data of stomach contents to estimate food rations. Journal of Fish Biology, 59(5):1198–1217, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2001:DSW

[546] A. M. Garcia, J. P. Vieira, and K. O. Winemiller. Dynamics of the shallow-water fish assemblage of the Patos Lagoon estuary (Brazil) during cold and warm ENSO episodes. Journal of Fish Biology, 59(5):1218– 1238, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manera:2001:EDO

[547] M. Manera, E. Simoni, and B. S. Dezfuli. The effect of dexamethasone on the occurrence and ultrastructure of rodlet cells in goldfish. Journal of Fish Biology, 59(5):1239–1248, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lienesch:2001:PIS

[548] P. W. Lienesch and M. Gophen. Predation by inland silversides on an exotic cladoceran, Daphnia lumholtzi, in Lake Texoma, U.S.A. Journal of Fish Biology, 59(5):1249–1257, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bozzanao:2001:PSR

[549] A. Bozzanao, R. Murgia, S. Vallerga, J. Hirano, and S. Archer. The photoreceptor system in the retinae of two dogfishes, Scyliorhinus canicula and Galeus melastomus: possible relationship with depth distribution and predatory lifestyle. Journal of Fish Biology, 59(5):1258–1278, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ochi:2001:ACB

[550] H. Ochi, T. Onchiand, and Y. Yanagisawa. Alloparental care between catfishes in Lake Tanganyika. Journal of Fish Biology, 59(5):1279–1286, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zbinden:2001:SAT

[551] M. Zbinden, C. R. Largiadèr, and T. C. M. Bakker. Sperm allocation in the three-spined stickleback. Journal of Fish Biology, 59(5):1287–1297, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Logothetis:2001:GMF

[552] E. A. Logothetis, M. H. Horn, and K. A. Dickson. Gut morphology and function in Atherinops affinis (Teleostei: Atherinopsidae), a stomachless omnivore feeding on macroalgae. Journal of Fish Biology, 59(5):1298– 1312, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jachner:2001:APB

[553] A. Jachner. Anti-predator behaviour of naïve compared experienced juvenile roach. Journal of Fish Biology, 59(5):1313–1322, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barrett:2001:CIC

[554] K. J. Barrett, D. G. McDonald, and M. J. O.’Donnell. Changes in ion content and transport during development of embryonic rainbow trout. Journal of Fish Biology, 59(5):1323–1335, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burleson:2001:IFS

[555] M. L. Burleson, D. R. Wilhelm, and N. J. Smatresk. The influence of fish size size on the avoidance of hypoxia and oxygen selection by largemouth bass. Journal of Fish Biology, 59(5):1336–1349, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagemans:2001:DBS

[556] F. Wagemans and P. Vandewalle. Development of the bony skull in common sole: brief survey of morpho-functional aspects of ossification sequence. Journal of Fish Biology, 59(5):1350–1369, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pilling:2001:SCP

[557] G. M. Pilling, M. G. Purves, T. M. Daw, D. A. Agnew, and J. C. Xavier. The stomach contents of Patagonian toothfish around South Georgia (South Atlantic). Journal of Fish Biology, 59(5):1370–1384, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broad:2001:MUL

[558] T. L. Broad, C. R. Townsend, G. P. Closs, and D. J. Jellyman. Microhabitat use by longfin eels in New Zealand streams with contrasting riparian vegetation. Journal of Fish Biology, 59(5):1385–1400, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacNeil:2001:BCP

[559] C. MacNeil, R. W. Elwood, and J. T. A. Dick. Brief communications: Persistence times of four amphipod species in the stomachs of brown trout. Journal of Fish Biology, 59(5):1401–1404, November 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johansen:2001:EFF

[560] M. Johansen. Evidence of freshwater feeding by adult salmon in the Tana River, northern Norway. Journal of Fish Biology, 59(5):1405–1407, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Willis:2001:VCM

[561] T. J. Willis. Visual census methods underestimate density and diversity of cryptic reef fishes. Journal of Fish Biology, 59(5):1408–1411, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leknes:2001:UFF

[562] I. L. Leknes. The uptake of foreign ferritin by macrophages in the spleen, trunk kidney and liver of platy. Journal of Fish Biology, 59(5):1412–1415, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2001:SES

[563] D. W. Evans, M. A. Matthews, and C. A. McClintock. The spread of the eel swimbladder nematode Anguillicola crassus through the Erne system, Ireland. Journal of Fish Biology, 59(5):1416–1420, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perrin:2001:MDA

[564] C. Perrin and P. Borsa. Mitochondrial DNA analysis of the geographic structure of Indian scad mackerel in the Indo-Malay archipelago. Journal of Fish Biology, 59(5):1421–1426, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez:2001:MES

[565] A. Rodríguez and E. Gisbert. Morphogenesis of the eye of Siberian sturgeon. Journal of Fish Biology, 59(5):1427–1429, November 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staniland:2001:RSC

[566] I. J. Staniland, P. J. B. Hart, and P. J. Bromley. The regurgitation of stomach contents in trawl caught whiting, evidence of a predator size effect. Journal of Fish Biology, 59(5):1430–1432, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:PE

[567] Anonymous. Printer’s erratum. Journal of Fish Biology, 59(5):1433, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ae

[568] Anonymous. Announcements. Journal of Fish Biology, 59(5):1434–1435, November 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beier:2001:RPH

[569] U. Beier. Regular papers: Habitat distribution and size structure in freshwater fish communities: effects of vendace on interactions between perch and roach. Journal of Fish Biology, 59(6):1437–1454, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yambe:2001:SSR

[570] H. Yambe and F. Yamazaki. Species-specific releaser effect of urine from ovulated female masu salmon and rainbow trout. Journal of Fish Biology, 59(6):1455–1464, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Falk-Petersen:2001:ODN

[571] I. B. Falk-Petersen and T. K. Hansen. Organ differentiation in newly hatched common wolffish. Journal of Fish Biology, 59(6):1465–1482, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johdal:2001:CBC

[572] M. S. Johdal, H. R. Esmaeili, and K. K. Tandon. A comparison of backcalculated lengths of silver carp derived from bony structures. Journal of Fish Biology, 59(6):1483–1493, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quagio-Grassiotto:2001:USC

[573] I. Quagio-Grassiotto, J. N. C. Negrão, E. D. Carvalho, and F. Foresti. Ultrastructure of spermatogenic cells and spermatozoa in Hoplias malabaricus (Teleostei, Characiformes, Erythrinidae). Journal of Fish Biology, 59(6):1494–1502, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Picchitti:2001:SRV

[574] S. Picchitti, G. Scapigliati, M. Fanelli, F. Barbato, S. Canese, L. Mastrolla, M. Mazzini, and L. Abelli. Sex-related variations of serum immunoglobulins during reproduction in gilthead sea bream and evidence for a transfer from the female to the eggs. Journal of Fish Biology, 59(6): 1503–1511, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2001:DSS

[575] K. A. Glover, F. Nilsen, Ø. Skaala, J. B. Taggart, and A. J. Teale. Differences in susceptibility to sea lice infection between a sea run and a freshwater resident population of brown trout. Journal of Fish Biology, 59(6):1512–1519, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2001:SDN

[576] R. J. Scott. Sensory drive and nuptial colour loss in the three-spined stickleback. Journal of Fish Biology, 59(6):1520–1528, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pegg:2001:GRR

[577] M. A. Pegg and C. L. Pierce. Growth rate responses of Missouri and Lower Yellowstone River fishes to a latitudinal gradient. Journal of Fish Biology, 59(6):1529–1543, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2001:PES

[578] J. A. Brown, W. M. Moore, and E. S. Quabius. Physiological effects of saline waters on zander. Journal of Fish Biology, 59(6):1544–1555, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2001:GHA

[579] K. Morita. The growth history of anadromous white-spotted charr in northern Japan: a comparison between river and sea life. Journal of Fish Biology, 59(6):1556–1565, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Downing:2001:ELI

[580] G. Downing and M. K. Litvak. The effect of light intensity and spectrum on the incidence of first feeding by larval haddock. Journal of Fish Biology, 59(6):1566–1578, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2001:WSV

[581] E. M. Olsen and L. A. Vøllestad. Within-stream variation in early lifehistory traits in brown trout. Journal of Fish Biology, 59(6):1579–1588, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Callejas:2001:MIR

[582] C. Callejas and M. D. Ochando. Molecular identification (RAPD) of the eight species of the genus Barbus (Cyprinidae) in the Iberian Peninsula. Journal of Fish Biology, 59(6):1589–1599, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berejikian:2001:CAS

[583] B. A. Berejikian, E. P. Tezak, S. C. Riley, and A. L. LaRae. Competitive ability and social behaviour of juvenile steelhead reared in enriched and conventional hatchery tanks and a stream environment. Journal of Fish Biology, 59(6):1600–1613, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaroensutasinee:2001:SSD

[584] M. Jaroensutasinee and K. Jaroensutasinee. Sexual size dimorphism and male contest in wild Siamese fighting fish. Journal of Fish Biology, 59(6): 1614–1621, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2001:ESS

[585] E. M. Olsen and L. A. Vollestad. Estimates of survival of stream-dwelling brown trout using. Journal of Fish Biology, 59(6):1622–1637, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Summerer:2001:MPB

[586] M. Summerer, R. Hanel, and C. Sturmbauer. Mitochondrial phylogeny and biogeographic affinities of sea breams of the genus Diplodus (Sparidae). Journal of Fish Biology, 59(6):1638–1652, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holm:2001:IMA

[587] C. F. Holm, J. D. Armstrong, and D. J. Gilvear. Investigating a major assumption of predictive instream habitat models: is water velocity preference of juvenile Atlantic salmon independent of discharge? Journal of Fish Biology, 59(6):1653–1666, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoare:2001:GNT

[588] D. J. Hoare, A. J. W. Ward, I. D. Couzin, D. P. Croft, and J. Krause. A grid-net technique for the analysis of fish positions within free-ranging shoals. Journal of Fish Biology, 59(6):1667–1672, December 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barry:2001:FTC

[589] T. P. Barry, M. J. Unwin, J. A. Malison, and T. P. Quinn. Free and total cortisol levels in semelparous and iteroparous Chinook salmon. Journal of Fish Biology, 59(6):1673–1676, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pfeiler:2001:CHT

[590] E. Pfeiler. Changes in hypoxia tolerance during metamorphosis of bonefish leptocephali. Journal of Fish Biology, 59(6):1677–1681, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shoji:2001:SPJ

[591] J. Shoji and M. Tanka. Strong piscivory of Japanese Spanish mackerel larvae from their first feeding. Journal of Fish Biology, 59(6):1682–1685, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzoldi:2001:RAM

[592] C. Mazzoldi. Reproductive apparatus and mating system in two tropical goby species. Journal of Fish Biology, 59(6):1686–1691, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ec

[593] Anonymous. Erratum. Journal of Fish Biology, 59(6):1692, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ed

[594] Anonymous. Erratum. Journal of Fish Biology, 59(6):1693, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Ee

[595] Anonymous. Erratum. Journal of Fish Biology, 59(6):1694, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2001:Af

[596] Anonymous. Announcements. Journal of Fish Biology, 59(6):1695–1697, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waples:2001:CDS

[597] R. S. Waples, R. G. Gustafson, L. A. Weitkamp, J. M. Myers, O. W. Jjohnson, P. J. Busby, J. J. Hard, G. J. Bryant, F. W. Waknitz, K. Nelly, D. Teel, W. S. Grant, G. A. Winans, S. Phelps, A. Marshall, and B. M. Baker. Characterizing diversity in salmon from the Pacific Northwest. Journal of Fish Biology, 59(sA):1–41, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rubidge:2001:MAH

[598] E. Rubidge, P. Corbett, and E. B. Taylor. A molecular analysis of hybridization between native westslope cutthroat trout and introduced rainbow trout in southeastern British Columbia, Canada. Journal of Fish Biology, 59(sA):42–54, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McGlashan:2001:GEH

[599] D. J. McGlashan and J. M. Hughes. Genetic evidence for historical continuity between populations of the Australian freshwater fish Craterocephalus stercusmuscarum (Atherinidae) east and west of the Great Dividing Range. Journal of Fish Biology, 59(sA):55–67, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almodovar:2001:GIB

[600] A. Almodóvar, J. Suárez, G. G. Nicola, and M. Nuevo. Genetic introgression between wild and stocked brown trout in the Douro River basin, Spain. Journal of Fish Biology, 59(sA):68–74, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bohlen:2001:SHR

[601] J. Bohlen and P. Ráb. Species and hybrid richness in spined loaches of the genus Cobitis (Teleostei: Cobitidae), with a checklist of European forms and suggestions for conservation. Journal of Fish Biology, 59(sA):75–89, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadovy:2001:TFH

[602] Y. Sadovy. The threat of fishing to highly fecund fishes. Journal of Fish Biology, 59(sA):90–108, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smedbol:2001:IMW

[603] R. K. Smedbol and R. Stephenson. The importance of managing withinspecies diversity in cod and herring fisheries of the north-western Atlantic. Journal of Fish Biology, 59(sA):109–128, December 2001. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eristhee:2001:HRS

[604] N. Eristhee and H. A. Oxenford. Home range size and use of space by Bermuda chub Kyphosus sectatrix (L.) in two marine reserves in the Soufrière Marine Management Area, St Lucia, West Indies. Journal of Fish Biology, 59(sA):129–151, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valles:2001:LSM

[605] H. Valles, S. Sponaugle, and H. A. Oxenford. Larval supply to a marine reserve and adjacent fished area in the Soufrière Marine Management Area, St Lucia, West Indies. Journal of Fish Biology, 59(sA):152–177, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2001:EMR

[606] I. M. Côté, I. Mosqueira, and J. D. Reynolds. Effects of marine reserve characteristics on the protection of fish populations: a meta-analysis. Journal of Fish Biology, 59(sA):178–189, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bianco:2001:ACF

[607] P. G. Bianco and V. Ketmaier. Anthropogenic changes in the freshwater fish fauna of Italy, with reference to the central region and Barbus graellsii, a newly established alien species of Iberian origin. Journal of Fish Biology, 59(sA):190–208, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ling:2001:CDS

[608] N. Ling, D. M. Gleeson, K. J. Willis, and S. U. Binzegger. Creating and destroying species: the ‘new’ biodiversity and evolutionarily significant units among New Zealand’s galaxiid fishes. Journal of Fish Biology, 59(sA):209–222, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uiblein:2001:SHU

[609] F. Uiblein, A. Jagsch, W. Honsig-Erlenburg, and S. Weiss. Status, habitat use, and vulnerability of the European ggrayling in Austrian waters. Journal of Fish Biology, 59(sA):223–247, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2001:PPL

[610] S. J. Cooke, T. W. Kassler, and D. P. Philipp. Physiological performance of largemouth bass related to local adaptation and interstock hybridization: implications for conservation and management. Journal of Fish Biology, 59(sA):248–268, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koljonen:2001:CGF

[611] M-L. Koljonen. Conservation goals and fisheries management units for Atlantic salmon in the Baltic Sea area. Journal of Fish Biology, 59(sA): 269–288, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mezzera:2001:CAI

[612] M. Mezzera and C. R. Largiadèr. Comparative analysis of introgression at three marker classes: a case study in a stocked population of brown trout. Journal of Fish Biology, 59(sA):289–305, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutchings:2001:IPD

[613] J. A. Hutchings. Influence of population decline, fishing, and spawner variability on the recovery of marine fishes. Journal of Fish Biology, 59(sA):306–322, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elvira:2001:FFI

[614] B. Elvira and A. Almodóvar. Freshwater fish introductions in Spain: facts and figures at the beginning of the 21st century. Journal of Fish Biology, 59(sA):323–331, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fricke:2001:CHR

[615] H. Fricke. Coelacanths: a human responsibility. Journal of Fish Biology, 59(sA):332–338, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrod:2001:IPC

[616] C. Harrod, D. Griffiths, T. K. McCarthy, and R. Rosell. The Irish pollan, Coregonus autumnalis: options for its conservation. Journal of Fish Biology, 59(sA):339–355, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McPherson:2001:GDC

[617] A. A. McPherson, R. L. Stephenson, P. T. O’Reilly, M. W. Jones, and C. T. Taggart. Genetic diversity of coastal Northwest Atlantic herring populations: implications for management. Journal of Fish Biology, 59(sA):356–370, December 2001. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:OAH

[618] Anonymous. Obituary: Arthur Hasler, 1908–2001. Journal of Fish Biology, 60(1):1–2, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fukuwaka:2002:ESM

[619] M. Fukuwaka and T. Suzuki. Early sea mortality of mark-recaptured juvenile chum salmon in open coastal waters. Journal of Fish Biology, 60(1):3–12, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arocha:2002:ODM

[620] F. Arocha. Oocyte development and maturity classification of swordfish from the north-western Atlantic. Journal of Fish Biology, 60(1):13–27, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Friedlander:2002:EIS

[621] M. Friedlander, J. D. Parrish, and R. C. DeFelice. Ecology of the introduced snapper Lutjanus kasmiva (Forsskål) in the reef fish assemblage of a Hawaiian bay. Journal of Fish Biology, 60(1):28–48, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laikre:2002:STP

[622] L. Laikre, T. Järvi, L. Johansson, S. Palm, J-F. Rubin, C. E. Glimsäter, P. Landergren, and N. Ryman. Spatial and temporal population structure of sea trout at the Island of Gotland, Sweden, delineated from mitochondrial DNA. Journal of Fish Biology, 60(1):49–71, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggenes:2002:SHP

[623] J. Heggenes, S. J. Saltveit, D. Bird, and R. Grew. Static habitat partitioning and dynamic selection by sympatric young Atlantic salmon and brown trout in south-west England streams. Journal of Fish Biology, 60(1):72–86, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reichard:2002:IVS

[624] M. Reichard, P. Jurajda, and M. Ondračková. Interannual variability in seasonal dynamics and species composition of drifting young-of-the-year fishes in two European lowland rivers. Journal of Fish Biology, 60(1):87– 101, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Altimiras:2002:CST

[625] J. Altimiras, M. Axelsoon, G. Claireaux, C. Lefrançois, C. Mercier, and A. P. Farrell. Cardiorespiratory status of triploid brown trout during swimming at two acclimation temperatures. Journal of Fish Biology, 60 (1):102–116, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mercier:2002:VCP

[626] C. Mercier, M. Axelsson, N. Imbert, G. Claireaux, C. Lefrançois, J. Altimiras, and A. P. Farrell. In vitro cardiac performance in triploid brown trout at two acclimation temperatures. Journal of Fish Biology, 60(1): 117–133, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Middlemas:2002:RLJ

[627] S. J. Middlemas and J. D. Armstrong. Reconstructing the lengths of juvenile Atlantic salmon from atlas bones: estimating the parr to smolt ratio from regression analyses revisited. Journal of Fish Biology, 60(1): 134–138, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tanaka:2002:LTC

[628] S. Tanaka, H. Zhang, N. Horie, Y. Yamada, A. Okamura, T. Utoh, N. Mikawa, H. P. Oka, and H. Kurokura. Long-term cryopreservation of sperm of Japanese eel. Journal of Fish Biology, 60(1):139–146, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abascal:2002:UTT

[629] F. J. Abascal, A. Medina, C. Megina, and A. Calzada. Ultrastructure of Thunnus thynnus and Euthynnus allettevatus spermatozoa. Journal of Fish Biology, 60(1):147–153, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Funamoto:2002:REJ

[630] T. Funamoto and I. Aoki. Reproductive ecology of Japanese anchovy off the Pacific coast of eastern Honshu, Japan. Journal of Fish Biology, 60 (1):154–169, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cleveland:2002:MVM

[631] L. Cleveland, M. Itzkowitz, and M. Haley. Male variation in mating success after female numbers are reduced. Journal of Fish Biology, 60 (1):170–177, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown-Peterson:2002:ACG

[632] N. J. Brown-Peterson, H. J. Grier, and R. M. Overstreet. Annual changes in germinal epithelium determine male reproductive classes of the cobia. Journal of Fish Biology, 60(1):178–202, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medina:2002:SAR

[633] A. Medina, F. J. Abascal, C. Megina, and A. García. Stereological assessment of the reproductive status of female Atlantic northern bluefin tuna during migration to Mediterranean spawning grounds through the Strait of Gibraltar. Journal of Fish Biology, 60(1):203–217, January 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pakkasmaa:2002:ILA

[634] S. Pakkasmaa and M. Jones. Individual-level analysis of early life history traits in hatchery-reared lake trout. Journal of Fish Biology, 60(1):218– 225, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gault:2002:BCF

[635] N. F. S. Gault, D. J. Kllpatrick, and M. T. Stewart. Biological control of the fish louse in a rainbow trout fishery. Journal of Fish Biology, 60 (1):226–237, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flodmark:2002:CGR

[636] L. E. W. Flodmark, H. A. Urke, J. H. Halleraker, J. V. Arnekleiv, L. A. Vøllestad, and A. B. S. Poléo. Cortisol and glucose responses in juvenile brown trout subjected to a fluctuating flow regime in an artificial stream. Journal of Fish Biology, 60(1):238–248, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagliniere:2002:OAA

[637] J. L. Baglinière, K. Tabet Aoul, and J. Y. Menella. Occurrence of an adult Atlantic salmon in the River Rhône, France. Journal of Fish Biology, 60(1):249–255, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Villalobos:2002:PIG

[638] H. Villalobos and R. Rodriguez-Sánchez. Pattern of increase in gill raker number of the California sardine. Journal of Fish Biology, 60(1):256– 259, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruyndoncx:2002:EPI

[639] L. Bruyndoncx, G. Knaepkens, W. Meeus, L. Bervoets, and M. Eens. The evaluation of passive integrated transponder (PIT) tags and visible implant elastomer (VIE) marks as new marking techniques for the bullhead. Journal of Fish Biology, 60(1):260–262, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sawada:2002:HCR

[640] Y. Sawada, M. Seoka, T. Okada, S. Miyashita, O. Murata, and H. Kumai. Hermaphroditism in a captive-raised Pacific bluefin tuna. Journal of Fish Biology, 60(1):263–265, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knox:2002:GAA

[641] D. Knox, K. Lehmann, D. G. Reddin, and E. Verspoor. Genotyping of archival Atlantic salmon scales from northern Quebec and West Greenland using novel PCR primers for degraded mtDNA. Journal of Fish Biology, 60(1):266–270, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castro:2002:ORL

[642] J. J. Castro and A. G. Ramos. The occurrence of Ranzania laevis off the Island of Gran Canaria, the Canary Islands, related to sea warming. Journal of Fish Biology, 60(1):271–273, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRa

[643] Anonymous. Book reviews. Journal of Fish Biology, 60(1):274–276, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:Aa

[644] Anonymous. Announcements. Journal of Fish Biology, 60(1):277–278, January 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beamish:2002:OWE

[645] R. J. Beamish. Obituary: William Edwin Ricker OC, FRSC, LLD, DSc. Journal of Fish Biology, 60(2):285–287, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsland:2002:CGR

[646] A. K. Imsland, A. Foss, S. W. Bonga, E. van Ham, and S. O. Stefansson. Comparison of growth and RNA: DNA ratios in three populations of juvenile turbot reared at two salinities. Journal of Fish Biology, 60(2): 288–300, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Platten:2002:IIL

[647] J. R. Platten, I. R. Tibbetts, and M. J. Sheaves. The influence of increased line-fishing mortality on the sex ratio and age of sex reversal of the venus tusk fish. Journal of Fish Biology, 60(2):301–318, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

CInar:2002:ILG

[648] K. Çinar and A. Dller. Immunohistochemical localization of glucagon, substance-p and vasoactive intestinal peptide in gastrointestinal tract mucosa of zander. Journal of Fish Biology, 60(2):319–327, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mananos:2002:LHS

[649] E. Mañanós, M. Carrillo, L. A. Sorbera, C. C. Mylonas, J. F. Asturiano, M. J. Bayarri, Y. Zohar, and S. Zanuy. Luteinizing hormone and sexual steroid plasma levels after treatment of European sea bass with sustained-release delivery systems for gonadotropin-releasing hormone analogue. Journal of Fish Biology, 60(2):328–339, February 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woody:2002:COA

[650] C. A. Woody, J. Nelson, and K. Ramstad. Clove oil as an anaesthetic for adult sockeye salmon: field trials. Journal of Fish Biology, 60(2): 340–347, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guiguer:2002:USI

[651] K. R. R. A. Guiguer, J. D. Reist, M. Power, and J. A. Babaluk. Using stable isotopes to confirm the trophic ecology of Arctic charr morphotypes from Lake Hazen, Nunavut, Canada. Journal of Fish Biology, 60 (2):348–362, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2002:RCI

[652] S. W. Griffiths and J. D. Armstrong. Rearing conditions influence refuge use among over-wintering Atlantic salmon juveniles. Journal of Fish Biology, 60(2):363–369, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hawkins:2002:PSO

[653] S. L. Hawkins, N. V. Varnavskaya, E. A. Matzak, V. V. Efremov, C. M. Guthrie III, R. L. Wilmot, H. Mayama, F. Yamazaki, and A. J. Gharrett. Population structure of odd-broodline Asian pink salmon and its contrast to the even-broodline structure. Journal of Fish Biology, 60(2):370–388, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva-Paterna:2002:AGR

[654] F. J. Oliva-Paterna, M. M. Torralva, and C. Fernández-Delgado. Age, growth and reproduction of Cobitis paludica in a seasonal stream. Journal of Fish Biology, 60(2):389–404, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ljunggren:2002:GRP

[655] L. Ljunggren. Growth response of pikeperch larvae in relation to body size and zooplankton abundance. Journal of Fish Biology, 60(2):405–414, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesthagen:2002:LHC

[656] T. Hesthagen and B. Jonsson. Life history characteristics of brown trout in lakes at different stages of acidification. Journal of Fish Biology, 60 (2):415–426, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rlnchard:2002:IGG

[657] J. Rlnchard, M. A. Garcia-Abiado, K. Dabrowski, J. Ottobre, and D. Schmidt. Induction of gynogenesis and gonad development in the muskellunge. Journal of Fish Biology, 60(2):427–441, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henrique:2002:EDA

[658] M. M. F. Henrique, M. F. Gouillou-Coustans, and E. Gomes. Effect of dietary ascorbic acid supplementation and chronic hypoxia on sea bream growth and vitamin C status. Journal of Fish Biology, 60(2):442–452, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cavalli:2002:GPA

[659] L. Cavalli, N. Pech, L. Tron, and R. Chappaz. Growth patterns of Arctic charr in five high altitude lakes in a French Alpine massif. Journal of Fish Biology, 60(2):453–465, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2002:PGC

[660] J. R. M. C. Silva, N. A. Staines, F. J. Hernandez-Blazquez, L. R. Porto-Neto, and J. C. S. Borges. Phagocytosis and giant cell formation at 0^{o}C by macrophage (MØ) of Notothenia coriiceps. Journal of Fish Biology, 60 (2):466–478, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffen:2002:LHL

[661] A. J. Geffen. Length of herring larvae in relation to age and time of hatching. Journal of Fish Biology, 60(2):479–485, February 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castriota:2002:FRS

[662] L. Castriota, S. Greco, G. Marino, and F. Andaloro. First record of Seriola vivoliana Cuvier, 1833 in the Mediterranean. Journal of Fish Biology, 60(2):486–488, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2002:FBC

[663] H. Rikardsen, P.-A. Amundsen, and P. J. Bodin. Foraging behaviour changes of Arctic charr during smolt migration in northern Norway. Journal of Fish Biology, 60(2):489–491, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brickle:2002:NRD

[664] P. Brickle and V. Laptikhovsky. New records of deep-sea fishes from the waters around the Falkland Islands. Journal of Fish Biology, 60 (2):492–494, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2002:OCC

[665] J. P. Evans, T. E. Pitcher, and A. E. Magurran. The ontogeny of courtship, colour and sperm production in male guppies. Journal of Fish Biology, 60(2):495–498, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:Ea

[666] Anonymous. Erratum. Journal of Fish Biology, 60(2):499, February 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:2002:AEE

[667] M. W. Clarke, P. L. Connolly, and J. J. Bracken. Age estimation of the exploited deepwater shark Centrophorus squamosus from the continental slopes of the Rockall Trough and Porcupine Bank. Journal of Fish Biology, 60(3):501–514, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2002:IDH

[668] C. E. Adams and F. A. Huntingford. Inherited differences in head allometry in polymorphic Arctic charr from Loch Rannoch, Scotland. Journal of Fish Biology, 60(3):515–520, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomasini:2002:MRS

[669] J. A. Tomasini and T. Laugier. Male reproductive strategy and reserve allocation in sand smelt from brackish lagoons of southern France. Journal of Fish Biology, 60(3):521–531, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Will:2002:MFS

[670] T. A. Will, T. R. Reinert, and C. A. Jennings. Maturation and fecundity of a stock-enhanced population of striped bass in the Savannah River Estuary, U.S.A. Journal of Fish Biology, 60(3):532–544, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mansour:2002:SAC

[671] N. Mansour, F. Lahnsteiner, and R. A. Patzner. The spermatozoon of the African catfish: fine structure, motility, viability and its behaviour in seminal vesicle secretion. Journal of Fish Biology, 60(3):545–560, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babiak:2002:SRT

[672] I. Babiak, J. Glogowski, S. Dobosz, H. Kuzminki, and K. Goryczko. Semen from rainbow trout produced using cryopreserved spermatozoa is more suitable for cryopreservation. Journal of Fish Biology, 60(3):561– 570, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huse:2002:MCM

[673] G. Huse, S. Railsback, and A. Feronö. Modelling changes in migration pattern of herring: collective behaviour and numerical domination. Journal of Fish Biology, 60(3):571–582, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harvey:2002:CTB

[674] R. Harvey and R. S. Batty. Cutaneous taste buds in gadoid fishes. Journal of Fish Biology, 60(3):583–592, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olin:2002:FCS

[675] M. Olin, M. Rask, J. Ruuhljärvi, M. Kurkilahti, P. Ala-Opas, and O. Ylönen. Fish community structure in mesotrophic and eutrophic lakes of southern Finland: the relative abundances of percids and cyprinids along a trophic gradient. Journal of Fish Biology, 60(3):593–612, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2002:RFS

[676] A. Moore, K. H. Olsén, N. Lower, and H. Kindahl. The role of F-series prostaglandins as reproductive priming pheromones in the brown trout. Journal of Fish Biology, 60(3):613–624, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kwon:2002:TAI

[677] J. Y. Kwon, B. J. McAndrew, and D. J. Penman. Treatment with an aromatase inhibitor suppresses high-temperature feminization of genetic male (YY) Nile tilapia. Journal of Fish Biology, 60(3):625–636, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Modesto:2002:TPO

[678] T. Modesto and A. V. M. Canário. 17α, 20β, 21-trihydroxy-4-pregnen-3-one: the probable maturation-inducing steroid of the Lusitanian toadfish. Journal of Fish Biology, 60(3):637–648, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alves:2002:GMA

[679] M. J. Alves, M. J. Collares-Pereira, T. E. Dowling, and M. M. Coelho. The genetics of maintenance of an all-male lineage in the Squalius albuvnoides complex. Journal of Fish Biology, 60(3):649–662, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesp:2002:IGG

[680] S. A. Hesp, R. P. Hobbs, and I. C. Potter. Infection of the gonads of Glaucosoma hebvaicum by the nematode Philometva lateolabracis: occurrence and host response. Journal of Fish Biology, 60(3):663–673, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2002:ILH

[681] I. J. Morgan, I. D. McCarthy, and N. B. Metcalfe. The influence of lifehistory strategy on lipid metabolism in overwintering juvenile Atlantic salmon. Journal of Fish Biology, 60(3):674–686, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeCrespinDeBilly:2002:TBT

[682] V. De Crespin De Billy and P. Usseglio-Polatera. Traits of brown trout prey in relation to habitat characteristics and benthic invertebrate communities. Journal of Fish Biology, 60(3):687–714, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shikano:2002:HNS

[683] T. Shikano and N. Taniguchi. Heterosis for neonatal survival in the guppy. Journal of Fish Biology, 60(3):715–725, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Couldridge:2002:EMM

[684] V. C. K. Couldridge. Experimental manipulation of male eggspots demonstrates female preference for one large spot in Pseudotvopheus lombavdoi. Journal of Fish Biology, 60(3):726–730, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kamjunke:2002:ADC

[685] N. Kamjunke, R. Mendonca, I. Hardewig, and T. Mehner. Assimilation of different cyanobacteria as food and the consequences for internal energy stores of juvenile roach. Journal of Fish Biology, 60(3):731–738, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edds:2002:RUL

[686] D. R. Edds, W. J. Matthews, and F. P. Gelwick. Resource use by large catfishes in a reservoir: is there evidence for interactive segregation and innate differences? Journal of Fish Biology, 60(3):739–750, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brabrand:2002:LSB

[687] Å. Brabrand, A. G. Koestler, and R. Borgstrøm. Lake spawning of brown trout related to groundwater influx. Journal of Fish Biology, 60(3):751– 763, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golani:2002:OHF

[688] D. Golani and M. Fine. On the occurrence of Hippocampus fuscus in the eastern Mediterranean. Journal of Fish Biology, 60(3):764–766, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Appleyard:2002:GSS

[689] S. A. Appleyard, R. D. Ward, and P. M. Grewe. Genetic stock structure of bigeye tuna in the Indian Ocean using mitochondrial DNA and microsatellites. Journal of Fish Biology, 60(3):767–770, March 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staffan:2002:VFI

[690] F. Staffan, C. Magnhagen, and A. Alanärä. Variation in food intake within groups of juvenile perch. Journal of Fish Biology, 60(3):771– 774, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caputo:2002:SLA

[691] V. Caputo, G. Candi, E. Arneri, M. La Mesa, C. Cinti, M. Provinciali, P. Nisi Cerioni, and A. Gregorini. Short lifespan and apoptosis in Aphia minuta. Journal of Fish Biology, 60(3):775–779, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Einum:2002:GRC

[692] S. Einum, E. B. Thorstad, and T. F. Næsje. Growth rate correlations across life-stages in female Atlantic salmon. Journal of Fish Biology, 60 (3):780–784, March 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jaworski:2002:DBM

[693] A. Jaworski and E. Kamler. Development of a bioenergetics model for fish embryos and larvae during the yolk feeding period. Journal of Fish Biology, 60(4):785–809, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2002:FAR

[694] D. M. Gonçalves, P. C. Simões, A. C. Chumbinho, M. J. Correia, T. Fagundes, and R. F. Oliveira. Fluctuating asymmetries and reproductive success in the peacock blenny. Journal of Fish Biology, 60(4):810–820, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perante:2002:BSS

[695] N. C. Perante, M. G. Pajaro, J. J. Meeuwig, and A. C. J. Vincent. Biology of a seahorse species, Hippocampus comes in the central Philippines. Journal of Fish Biology, 60(4):821–837, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goudswaard:2002:TFS

[696] P. C. Goudswaard, F. Witte, and E. F. B. Katunzi. The tilapiine fish stock of Lake Victoria before and after the Nile perch upsurge. Journal of Fish Biology, 60(4):838–856, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valentin:2002:GMR

[697] A. Valentin, J.-M. Sévigny, and J.-P. Chanut. Geometric morphometrics reveals body shape differences between sympatric redfish Sebastes mentella, Sebastes fasdatus and their hybrids in the Gulf of St Lawrence. Journal of Fish Biology, 60(4):857–875, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haertel:2002:DDS

[698] S. S. Haertel and R. Eckmann. Diel diet shift of roach and its implications for the estimation of daily rations. Journal of Fish Biology, 60(4):876– 892, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ackerman:2002:CEC

[699] J. L. Ackerman and D. R. Bellwood. Comparative efficiency of clove oil and rotenone for sampling tropical reef fish assemblages. Journal of Fish Biology, 60(4):893–901, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanson:2002:DAC

[700] J. M. Hanson and G. A. Chouinard. Diet of Atlantic cod in the southern Gulf of St Lawrence as an index of ecosystem change, 1959–2000. Journal of Fish Biology, 60(4):902–922, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koumoundouros:2002:ETS

[701] G. Koumoundouros, D. G. Sfakianakis, P. Divanach, and M. Kentouri. Effect of temperature on swimming performance of sea bass juveniles. Journal of Fish Biology, 60(4):923–932, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klemetsen:2002:EGD

[702] A. Klemetsen, J. M. Elliott, R. Knudsen, and P. Sørensen. Evidence for genetic differences in the offspring of two sympatric morphs of Arctic charr. Journal of Fish Biology, 60(4):933–950, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mabragana:2002:REA

[703] E. Mabragaña, L. O. Lucifora, and A. M. Massa. The reproductive ecology and abundance of Symptevygia bonapavtii endemic to the south-west Atlantic. Journal of Fish Biology, 60(4):951–967, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kellison:2002:DDP

[704] G. T. Kellison, D. B. Eggleston, and M. Tanaka. Density-dependent predation and implications for stock enhancement with Japanese flounder. Journal of Fish Biology, 60(4):968–980, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munehara:2002:MSB

[705] H. Munehara, V. G. Sideleva, and A. Goto. Mixed-species brooding between two Baikal sculpins: field evidence for intra- and interspecific competition for reproductive resources. Journal of Fish Biology, 60(4): 981–988, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slavik:2002:FAM

[706] O. Slavík and L. Bartoš. Factors affecting migrations of burbot. Journal of Fish Biology, 60(4):989–998, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fijan:2002:MBC

[707] N. Fijan. Morphogenesis of blood cell lineages in channel catfish. Journal of Fish Biology, 60(4):999–1014, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kiiskinen:2002:SSV

[708] P. Kiiskinen, H. Hyvärinen, and J. Piironen. Smolting and seasonal variation in the smolt characteristics of one- and two-year-old Saimaa landlocked salmon under fish farm conditions. Journal of Fish Biology, 60(4):1015–1030, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wistbacka:2002:TAG

[709] S. Wistbacka, A. Heinonen, and G. Bylund. Thiaminase activity of gastrointestinal contents of salmon and herring from the Baltic Sea. Journal of Fish Biology, 60(4):1031–1042, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jenkins:2002:SAR

[710] G. P. Jenkins and D. C. Welsford. The swimming abilities of recently settled post-larvae of Sillaginodes punctata. Journal of Fish Biology, 60 (4):1043–1050, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Na-Nakorn:2002:IVA

[711] U. Na-Nakorn, P. Sodsuk, P. Wongrat, S. Janekitkarn, and D. M. Bartley. Isozyme variation among four species of the catfish genus Clarias. Journal of Fish Biology, 60(4):1051–1057, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cortey:2002:EPI

[712] M. Cortey and J-L. García-Marín. Evidence for phylogeographically informative sequence variation in the mitochondrial control region of Atlantic brown trout. Journal of Fish Biology, 60(4):1058–1063, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRb

[713] Anonymous. Book reviews. Journal of Fish Biology, 60(4):1064–1068, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:Ab

[714] Anonymous. Announcements. Journal of Fish Biology, 60(4):1069–1070, April 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gisbert:2002:ALR

[715] E. Gisbert and P. Williot. Advances in the larval rearing of Siberian sturgeon. Journal of Fish Biology, 60(5):1071–1092, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Massou:2002:EDF

[716] A. M. Massou, J. Panfili, R. Laë, J. F. Baroiller, O. Mikolasek, G. Fontenelle, and P.-Y. Le Bail. Effects of different food restrictions on somatic and otolith growth in Nile tilapia reared under controlled conditions. Journal of Fish Biology, 60(5):1093–1104, May 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mallekh:2002:ETD

[717] R. Mallekh and J. P. Lagardère. Effect of temperature and dissolved oxygen concentration on the metabolic rate of the turbot and the relationship between metabolic scope and feeding demand. Journal of Fish Biology, 60(5):1105–1115, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSanti:2002:CLR

[718] A. De Santi, A. Bisazza, and G. Vallortigara. Complementary left and right eye use during predator inspection and shoal-mate scrutiny in minnows. Journal of Fish Biology, 60(5):1116–1125, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Booth:2002:FSI

[719] D. Booth and R. Alquezar. Food supplementation increases larval growth, condition and survival of Acanthochromis polyacanthus. Journal of Fish Biology, 60(5):1126–1133, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tanaka:2002:IES

[720] S. Tanaka, H. Zhang, Y. Yamada, A. Okamura, N. Horie, T. Utoh, N. Mikawa, H. P. Oka, and H. Kurokura. Inhibitory effect of sodium bicarbonate on the motility of sperm of Japanese eel. Journal of Fish Biology, 60(5):1134–1141, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fijan:2002:CMH

[721] N. Fijan. Composition of main haematopoietic compartments in normal and bled channel catfish. Journal of Fish Biology, 60(5):1142–1154, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacey:2002:ICM

[722] J. A. Lacey and K. J. Rodnick. Important considerations for methaemoglobin measurement in fish blood: assay choice and storage conditions. Journal of Fish Biology, 60(5):1155–1169, May 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holker:2002:IAH

[723] F. Hölker and B. Breckling. Influence of activity in a heterogeneous environment on the dynamics of fish growth: an individual-based model of roacl. Journal of Fish Biology, 60(5):1170–1189, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2002:DEC

[724] D. Moran and K. D. Clements. Diet and endogenous carbohydrases in the temperate marine herbivorous fish Kyphosus sydneyanus. Journal of Fish Biology, 60(5):1190–1203, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gauldie:2002:EME

[725] R. W. Gauldie and J. S. Crampton. An eco-morphological explanation of individual variability in the shape of the fish otolith: comparison of the otolith of Hoplostethus atlanticus with other species by depth. Journal of Fish Biology, 60(5):1204–1221, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Specziar:2002:SEG

[726] A. Specziàr. In situ estimates of gut evacuation and its dependence on temperature in five cyprinids. Journal of Fish Biology, 60(5):1222– 1236, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Specziar:2002:SEF

[727] A. Specziár. An in situ estimate of food consumption of five cyprinid species in Lake Balaton. Journal of Fish Biology, 60(5):1237–1251, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giddelo:2002:IRH

[728] C. S. Giddelo, A. D. Arndt, and F. A. M. Volckaert. Impact of rifting and hydrography on the genetic structure of Clarias gariepinus in eastern Africa. Journal of Fish Biology, 60(5):1252–1266, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciannelli:2002:RIS

[729] L. Ciannelli, A. J. Paul, and R. D. Brodeur. Regional, interannual and size-related variation of age 0 year walleye pollock whole body energy content around the Pribilof Islands, Bering Sea. Journal of Fish Biology, 60(5):1267–1279, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drazen:2002:SAN

[730] J. C. Drazen. A seasonal analysis of the nutritional condition of deep-sea macrourid fishes in the north-east Pacific. Journal of Fish Biology, 60 (5):1280–1295, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poizat:2002:LHV

[731] G. Poizat, E. Rosecchi, and A. J. Crivelli. Life-history variation within a three-spined stickleback population in the Camargue. Journal of Fish Biology, 60(5):1296–1307, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Museth:2002:IEM

[732] J. Museth, R. Borgstrøm, J. E. Brittain, I. Herberg, and C. Naalsund. Introduction of the European minnow into a subalpine lake: habitat use and long-term changes in population dynamics. Journal of Fish Biology, 60(5):1308–1321, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gadig:2002:FRE

[733] O. B. F. Gadig and U. L. Gomes. First report on embryos of Isistius brasiliensis. Journal of Fish Biology, 60(5):1322–1325, May 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menezes:2002:GRA

[734] M. R. Menezes, U. Arolkar, and V. Rathod. Genetic relationships among five nemipterid fish species from the Indian coast using allozyme analysis. Journal of Fish Biology, 60(5):1326–1330, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brannas:2002:AGE

[735] E. Brännäs, J. Linnér, and L.-O. Eriksson. Aggression and growth as an effect of size composition in groups of Arctic charr. Journal of Fish Biology, 60(5):1331–1334, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bohlin:2002:STA

[736] T. Bohlin, J. C. E. Pettersson, and J. I. Johnsson. Is selection for territorial aggression in brown trout density-dependent? Journal of Fish Biology, 60(5):1335–1337, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2002:UCT

[737] J. M. Elliott and A. Klemetsen. The upper critical thermal limits for alevins of Arctic charr from a Norwegian lake north of the Arctic circle. Journal of Fish Biology, 60(5):1338–1341, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Serra:2002:RCI

[738] É. L. Serra, C. C. Medalha, R. O. Bueno, and R. Mattioli. The role of chlorpheniramine in inhibitory avoidance in Danio vevio. Journal of Fish Biology, 60(5):1342–1345, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2002:IHG

[739] J. S. Taylor and F. Breden. The inheritance of heteroplasmy in guppies. Journal of Fish Biology, 60(5):1346–1350, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nykanen:2002:SCS

[740] M. Nykänen and A. Huusko. Suitability criteria for spawning habitat of riverine European grayling. Journal of Fish Biology, 60(5):1351–1354, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ondrackova:2002:DPC

[741] M. Ondračková, P. Jurajda, and M. Gelnar. The distribution of Posthodiplostomum cuticola metacercariae in young-of-the-year cyprinid fishes. Journal of Fish Biology, 60(5):1355–1357, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nespolo:2002:IAH

[742] R. F. Nespolo and M. Rosenmann. Intraspecific allometry of haematological parameters in Basilichthys australis. Journal of Fish Biology, 60 (5):1358–1362, May 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dou:2002:FRM

[743] S. Dou, R. Masuda, M. Tanaka, and K. Tsukamoto. Feeding resumption, morphological changes and mortality during starvation in Japanese flounder larvae. Journal of Fish Biology, 60(6):1363–1380, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Resende:2002:IMP

[744] A. D. Resende, E. Rocha, R. A. F. Monteiro, and P. Rodrigues. The interhepatocytic macrophages and pale-grey cells in brown trout liver ontogenesis. Journal of Fish Biology, 60(6):1381–1390, June 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:2002:STA

[745] H. M. Murray, C. L. Hew, and G. L. Fletcher. Skin-type antifreeze protein expression in integumental cells of larval winter flounder. Journal of Fish Biology, 60(6):1391–1406, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amcoff:2002:ADS

[746] P. Amcoff, U. O. E. Elofsson, H. Börjeson, L. Norrgren, and G. E. Nilsson. Alterations of dopaminergic and serotonergic activity in the brain of sea-run Baltic salmon suffering a thiamine deficiency-related disorder. Journal of Fish Biology, 60(6):1407–1416, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lamatsch:2002:SMG

[747] D. K. Lamatsch, M. Schmid, and M. Schartl. A somatic mosaic of the gynogenetic Amazon molly. Journal of Fish Biology, 60(6):1417–1422, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saroglia:2002:MAS

[748] M. Saroglia, G. Terova, A. De Stradis, and A. Caputo. Morphometric adaptations of sea bass gills to different dissolved oxygen partial pressures. Journal of Fish Biology, 60(6):1423–1430, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hendry:2002:GSD

[749] C. I. Hendry, D. J. Martin-Robichaud, and T. J. Benfey. Gonadal sex differentiation in Atlantic halibut. Journal of Fish Biology, 60(6):1431– 1442, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whiteman:2002:CAT

[750] E. A. Whiteman and I. M. Côté. Cleaning activity of two Caribbean cleaning gobies: intra- and interspecific comparisons. Journal of Fish Biology, 60(6):1443–1458, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guarnieo:2002:CRH

[751] I. Guarnieo, S. Franzellitti, N. Ungaro, S. Tommasini, C. Piccinetti, and F. Tinti. Control region haplotype variation in the central Mediterranean common sole indicates geographical isolation and population structuring in Italian stocks. Journal of Fish Biology, 60(6):1459–1474, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brodeur:2002:FIM

[752] J. C. Brodeur, L. S. Peck, and I. A. Johnston. Feeding increases MyoD and PCNA expression in myogenic progenitor cells of Notothenia coviiceps. Journal of Fish Biology, 60(6):1475–1485, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arigoni:2002:ACM

[753] S. Arigoni, P. Francour, M. Harmelin-Vivien, and L. Zaninetti. Adaptive colouration of Mediterranean labrid fishes to the new habitat provided by the introduced tropical alga Caulerpa taxifolia. Journal of Fish Biology, 60(6):1486–1497, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vizzini:2002:SCN

[754] S. Vizzini and A. Mazzola. Stable carbon and nitrogen ratios in the sand smelt from a Mediterranean coastal area: feeding habits and effect of season and size. Journal of Fish Biology, 60(6):1498–1510, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riis-Vestergaard:2002:EDM

[755] J. Riis-Vestergaard. Energy density of marine pelagic fish eggs. Journal of Fish Biology, 60(6):1511–1528, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalbe:2002:DPP

[756] M. Kalbe, K. M. Wegner, and T. B. H. Reusch. Dispersion patterns of parasites in 0+ year three-spined sticklebacks: a cross population comparison. Journal of Fish Biology, 60(6):1529–1542, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Park:2002:SSA

[757] J. Y. Park. Structure of the skin of an air-breathing mudskipper, Periophthalmus magnuspinnatus. Journal of Fish Biology, 60(6):1543–1550, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Connell:2002:EPP

[758] S. D. Connell. Effects of a predator and prey on a foraging reef fish: implications for understanding density-dependent growth. Journal of Fish Biology, 60(6):1551–1561, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cakic:2002:BAS

[759] P. Cakić, M. Lenhardt, D. Mićković, N. Sekulić, and L. J. Budakov. Biometric analysis of Syngnathus abaster populations. Journal of Fish Biology, 60(6):1562–1569, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:2002:OFE

[760] E. Turner and R. Montgomerie. Ovarian fluid enhances sperm movement in Arctic charr. Journal of Fish Biology, 60(6):1570–1579, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2002:CCS

[761] W. G. Anderson, J. P. Good, and N. Hazon. Changes in chloride secretion rate and vascular perfusion in the rectal gland of the European lesser-spotted dogfish in response to environmental and hormonal stimuli. Journal of Fish Biology, 60(6):1580–1590, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vanHerwerden:2002:PEP

[762] L. van Herwerden, C. R. Davies, and J. H. Choat. Phylogenetic and evolutionary perspectives of the Indo–Pacific grouper Plectvopomus species on the Great Barrier Reef, Australia. Journal of Fish Biology, 60(6): 1591–1596, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seitz:2002:BSM

[763] A. C. Seitz, K. C. Weng, A. M. Boustany, and B. A. Block. Behaviour of a sharptail mola in the Gulf of Mexico. Journal of Fish Biology, 60 (6):1597–1602, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benedito-Cecilio:2002:VCI

[764] E. Benedito-Cecilio and C. A. R. M. Araujo-Lima. Variation in the carbon isotope composition of Semaprochilodus insignis, a detritivorous fish associated with oligotrophic and eutrophic Amazonian rivers. Journal of Fish Biology, 60(6):1603–1607, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Han:2002:EAE

[765] Y.-S. Han, C.-H. Yu, H.-T. Yu, C.-W. Chang, I-C. Liao, and W.-N. Tzeng. The exotic American eel in Taiwan: ecological implications. Journal of Fish Biology, 60(6):1608–1612, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRc

[766] Anonymous. Book reviews. Journal of Fish Biology, 60(6):1613–1617, June 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sloman:2002:PED

[767] K. A. Sloman and J. D. Armstrong. Physiological effects of dominance hierarchies: laboratory artefacts or natural phenomena? Journal of Fish Biology, 61(1):1–23, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strmac:2002:ESE

[768] M. Strmac, A. Oberemm, and T. Braunbeck. Effects of sediment eluates and extracts from differently polluted small rivers on zebrafish embryos and larvae. Journal of Fish Biology, 61(1):24–38, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tytler:2002:GOM

[769] P. Tytler, C. J. Fox, and A. Folkvord. Glycoconjugates in the otolithic membrane of herring larvae: a possible framework for encoding the life history recorder in fishes. Journal of Fish Biology, 61(1):39–49, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temming:2002:GEM

[770] A. Temming, B. Bøhle, D. W. Skagen, and F. R. Knudsen. Gastric evacuation in mackerel: the effects of meal size, prey type and temperature. Journal of Fish Biology, 61(1):50–70, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janssens:2002:PTE

[771] B. J. Janssens, R. Le Gall, and J.-F. Rees. Peroxide-triggered erythrocytes haemolysis as a model for the study of oxidative damage in marine fishes. Journal of Fish Biology, 61(1):71–84, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rios:2002:ELT

[772] F. S. Rios, A. L. Kalinin, and F. T. Rantin. The effects of long-term food deprivation on respiration and haematology of the neotropical fish Hoplias malabaricus. Journal of Fish Biology, 61(1):85–95, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2002:RMV

[773] D. M. Gonçalves, E. N. Barata, R. F. Oliveira, and A. V. M. Canário. The role of male visual and chemical cues on the activation of female courtship behaviour in the sex-role reversed peacock blenny. Journal of Fish Biology, 61(1):96–105, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiitu:2002:MFS

[774] V. Tiitu and M. Vornamen. Morphology and fine structure of the heart of the burbot, a cold stenothermal fish. Journal of Fish Biology, 61(1):106– 121, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Renones:2002:SRD

[775] O. Reñones, N. V. C. Polunin, and R. Goni. Size related dietary shifts of Epinephelus marginatus in a western Mediterranean littoral ecosystem: an isotope and stomach content analysis. Journal of Fish Biology, 61(1):

122–137, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milano:2002:PAM

[776] D. Milano, V. E. Cussac, P. J. Macchi, D. E. Ruzzante, M. F. Alonso, P. H. Vigliano, and M. A. Denegri. Predator associated morphology in Galaxias platei in Patagonian lakes. Journal of Fish Biology, 61(1):138– 156, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzoldi:2002:AMM

[777] C. Mazzoldi and M. B. Rasotto. Alternative male mating tactics in Gobius niger. Journal of Fish Biology, 61(1):157–172, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rasotto:2002:MTA

[778] M. B. Rasotto and C. Mazzoldi. Male traits associated with alternative reproductive tactics in Gobius niger. Journal of Fish Biology, 61(1): 173–184, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:2002:PMM

[779] I. Barber. Parasites, male-male competition and female mate choice in the sand goby. Journal of Fish Biology, 61(1):185–198, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tohse:2002:DVC

[780] H. Tohse and Y. Mugiya. Diel variations in carbonate incorporation into otoliths in goldfish. Journal of Fish Biology, 61(1):199–206, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:2002:TSS

[781] T. G. Pottinger, T. R. Carrick, and W. E. Yeomans. The three-spined stickleback as an environmental sentinel: effects of stressors on wholebody physiological indices. Journal of Fish Biology, 61(1):207–229, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsuruta:2002:ETA

[782] T. Tsuruta, H. Takahashi, and A. Goto. Evidence for type assortative mating between the freshwater and Omono types of nine-spined stickleback in natural fresh water. Journal of Fish Biology, 61(1):230–241, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chapman:2002:ROC

[783] L. J. Chapman, F. G. Nordlie, and A. Seifert. Respiratory oxygen consumption among groups of Pseudocrenilabrus multicolor victoriae subjected to different oxygen concentrations during development. Journal of Fish Biology, 61(1):242–251, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2002:ADR

[784] T. Ellis, R. N. Hughes, and B. R. Howell. Artificial dietary regime may impair subsequent foraging behaviour of hatchery-reared turbot released into the natural environment. Journal of Fish Biology, 61(1):252–264, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menghi:2002:IDP

[785] G. Menghi, M. G. Sabbieti, M. Menghi, A. Roda, S. Materazzi, and L. Marchetti. Immnunohistochemical detection of phthalate esters in the alimentary canal of Tilapia spp. Journal of Fish Biology, 61(1):265– 271, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braccini:2002:RBP

[786] J. M. Braccini and G. E. Chiaramonte. Reproductive biology of Psammobatis extenta. Journal of Fish Biology, 61(1):272–288, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landaeta:2002:FRP

[787] M. F. Landaeta, R. Veas, and L. R. Castro. First record of the paedomorphic goby Schindleria praematura, Easter Island, South Pacific. Journal of Fish Biology, 61(1):289–292, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bielawski:2002:PRW

[788] J. P. Bielawski, A. Brault, and J. R. Gold. Phylogenetic relationships within the genus Pimephales as inferred from ND4 and ND4L nucleotide sequences. Journal of Fish Biology, 61(1):293–297, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gozlan:2002:OAC

[789] R. E. Gozlan, A. C. Pinder, and J. Shelley. Occurrence of the Asiatic cyprinid Pseudorasbora parva in England. Journal of Fish Biology, 61 (1):298–300, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Froufe:2002:MSD

[790] E. Froufe, I. Magyary, I. Lehoczky, and S. Weiss. mtDNA sequence data supports an Asian ancestry and single introduction of the common carp into the Danube Basin. Journal of Fish Biology, 61(1):301–304, July 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibbotson:2002:DBD

[791] A. Ibbotson and M. Dunbar. Are differences between discharge-specific preference functions relevant for the application of PHABSIM? Journal of Fish Biology, 61(2):305–307, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See reply [792].

Armstrong:2002:HSA

[792] J. D. Armstrong, C. F. Holm, and D. J. Gilvear. Holding station against the flow: a reply to Ibbotson & Dunbar. Journal of Fish Biology, 61 (2):308–312, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [791].

Zhang:2002:IFD

[793] H. Zhang, K. Futami, Y. Yamada, N. Horie, A. Okamura, T. Utoh, N. Mikawa, S. Tanaka, N. Okamoto, and H. P. Oka. Isolation of freshwater and deep-sea type opsin genes from the common Japanese conger. Journal of Fish Biology, 61(2):313–324, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leis:2002:SSB

[794] J. M. Leis and B. M. Carson-Ewart. In situ settlement behaviour of damselfish (Pomacentridae) larvae. Journal of Fish Biology, 61(2):325– 346, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Felip:2002:GGD

[795] A. Felip, F. Piferrer, M. Carrillo, and S. Zanuy. Growth, gonadal development and sex ratios of meiogynogenetic diploid sea bass. Journal of Fish Biology, 61(2):347–359, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunn:2002:SSM

[796] M. R. Dunn and M. G. Pawson. The stock structure and migrations of plaice populations on the west coast of England and Wales. Journal of Fish Biology, 61(2):360–393, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pankhurst:2002:CPL

[797] N. W. Pankhurst and R. E. Peter. Changes in plasma levels of gonadal steroids and putative gonadotropin in association with spawning and brooding behaviour of male demoiselles. Journal of Fish Biology, 61 (2):394–404, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinyoles:2002:NOP

[798] D. Vinyoles, I. M. Côté, and A. De Sostoa. Nest orientation patterns in Salaria fluviatilis. Journal of Fish Biology, 61(2):405–416, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stark:2002:MSG

[799] J. W. Stark and D. A. Somerton. Maturation, spawning and growth of rock soles off Kodiak Island in the Gulf of Alaska. Journal of Fish Biology, 61(2):417–431, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babiak:2002:FIA

[800] I. Babiak, S. Dobosz, H. Kuzminski, K. Goryczko, S. Ciesielski, P. Brzuzan, B. Urbányi, Á. Horváth, F. Lahnsteiner, and J. Piironen. Failure of interspecies androgenesis in salmonids. Journal of Fish Biology, 61(2):432–447, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shields:2002:LAM

[801] B. A. Shields, K. L. Groves, C. Rombaugh, and R. Bellmore. Ligulosis associated with mortality in largescale suckers. Journal of Fish Biology, 61(2):448–455, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

cEch:2002:SCS

[802] M. čEch and J. Kubečka. Sinusoidal cycling swimming pattern of reservoir fishes. Journal of Fish Biology, 61(2):456–471, August 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hernandez:2002:WTF

[803] C. E. Hernández, P. E. Neill, J. M. Pulgar, F. P. Ojeda, and F. Bozinovic. Water temperature fluctuations and territoriality in the intertidal zone: two possible explanations for the elevational distribution of body size in Graus nigra. Journal of Fish Biology, 61(2):472–488, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salvolini:2002:RBK

[804] E. Salvolini, L. Nanetti, N. Moretti, A. Vignini, D. Martarelli, A. Frontini, A. M. Santroni, G. Falcioni, and L. Mazzanti. Relationships between Na^{+}, K^{+}-ATPase, sialic acid content and fluidity in rainbow trout density-separated erythrocytes. Journal of Fish Biology, 61(2):489–491, August 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2002:RBS

[805] T. Ellis, B. North, A. P. Scott, N. R. Bromage, M. Porter, and D. Gadd. The relationships between stocking density and welfare in farmed rainbow trout. Journal of Fish Biology, 61(3):493–531, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mirza:2002:ASS

[806] R. S. Mirza and P. Chiers. Attraction of slimy sculpins to chemical cues of brook charr eggs. Journal of Fish Biology, 61(3):532–539, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoner:2002:SPS

[807] A. W. Stoner and A. A. Abookire. Sediment preferences and sizespecific distribution of young-of-the-year Pacific halibut in an Alaska nursery. Journal of Fish Biology, 61(3):540–559, September 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bardin:2002:EFC

[808] O. Bardin and D. Pont. Environmental factors controlling the spring immigration of two estuarine fishes Atherina boyeri and Pomatoschistus spp. into a Mediterranean lagoon. Journal of Fish Biology, 61(3):560– 578, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rindorf:2002:ESF

[809] A. Rindorf. The effect of stomach fullness on food intake of whiting in the North Sea. Journal of Fish Biology, 61(3):579–593, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valente:2002:EGO

[810] L. M. P. Valente, G. Paboeuf, and B. Fauconneau. Effect of genetic origin of the fish on in vitro proliferation of muscle myosatellite cells of rainbow trout. Journal of Fish Biology, 61(3):594–605, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suter:2002:ECJ

[811] H. C. Suter and F. A. Huntingford. Eye colour in juvenile Atlantic salmon: effects of social status, aggression and foraging success. Journal of Fish Biology, 61(3):606–614, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rizzo:2002:ASP

[812] E. Rizzo, Y. Sato, B. P. Barreto, and H. P. Godinho. Adhesiveness and surface patterns of eggs in neotropical freshwater teleosts. Journal of Fish Biology, 61(3):615–632, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takeyama:2002:SPF

[813] T. Takeyama, N. Okuda, and Y. Yanagisawa. Seasonal pattern of filial cannibalism by Apogon doederleini mouthbrooding males. Journal of Fish Biology, 61(3):633–644, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baras:2002:OVT

[814] E. Baras, A. Mpo’n’tcha, H. Driouch, Ch. Prignon, and C. Mélard. Ontogenetic variations of thermal optimum for growth, and its implication on thermolabile sex determination in blue tilapia. Journal of Fish Biology, 61(3):645–660, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swansburg:2002:SVJ

[815] E. Swansburg, G. Chaput, D. Moore, D. Caissie, and N. El-Jabi. Size variability of juvenile Atlantic salmon: links to environmental conditions. Journal of Fish Biology, 61(3):661–683, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Connolly:2002:PMH

[816] R. M. Connolly, A. J. Melville, and K. M. Preston. Patterns of movement and habitat use by leafy seadragons tracked ultrasonically. Journal of Fish Biology, 61(3):684–695, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crespin:2002:SVN

[817] L. Crespin, P. Berrebi, and J-D. Lebreton. Spatially varying natural selection in a fish hybrid zone. Journal of Fish Biology, 61(3):696–711, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huang:2002:EAB

[818] C.-Y. Huang and J.-C. Chen. Effects on acid-base balance, methaemoglobinemia and nitrogen excretion of European eel after exposure to elevated ambient nitrite. Journal of Fish Biology, 61(3):712–725, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kucera:2002:EPA

[819] C. J. Kucera, C. K. Faulk, and G. J. Holt. The effect of parental acclimation to spawning salinity on the survival of larval Cynoscion nebulosus. Journal of Fish Biology, 61(3):726–738, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neff:2002:ASP

[820] B. D. Neff and T. E. Pitcher. Assessing the statistical power of genetic analyses to detect multiple mating in fishes. Journal of Fish Biology, 61 (3):739–750, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barcellos:2002:PSC

[821] L. J. G. Barcellos, G. F. Wassermann, A. P. Scott, V. M. Woehl, R. M. Quevedo, I. Ittzés, M. H. Krieger, and F. Lulhier. Plasma steroid concentrations in relation to the reproductive cycle of cultured male Rhamdia quelen. Journal of Fish Biology, 61(3):751–763, September 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Travers:2002:FIC

[822] M. J. Travers and I. C. Potter. Factors influencing the characteristics of fish assemblages in a large subtropical marine embayment. Journal of Fish Biology, 61(3):764–784, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCleave:2002:DNZ

[823] J. D. McCleave and D. J. Jellyman. Discrimination of New Zealand stream waters by glass eels of Anguilla australis and Anguilla dieffenbachii. Journal of Fish Biology, 61(3):785–800, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez-Jerez:2002:SVT

[824] P. Sanchez-Jerez, B. M. Gillanders, and M. J. Kingsford. Spatial variability of trace elements in fish otoliths: comparison with dietary items and habitat constituents in seagrass meadows. Journal of Fish Biology, 61 (3):801–821, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakamura:2002:DEM

[825] O. Nakamura, M. Saeki, H. Kamiya, K. Muramoto, and T. Watanabe. Development of epidermal and mucosal galectin containing cells in metamorphosing leptocephali of Japanese conger. Journal of Fish Biology, 61 (3):822–833, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundstrom:2002:ECT

[826] L. F. Sundström and S. H. Gruber. Effects of capture and transmitter attachments on the swimming speed of large juvenile lemon sharks in the wild. Journal of Fish Biology, 61(3):834–838, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alagador:2002:CST

[827] D. Alagador and A. G. Murta. A comparison of staining techniques to improve precision of age estimation from fish otoliths. Journal of Fish Biology, 61(3):839–841, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makhrov:2002:ASI

[828] A. A. Makhrov, O. Skaala, and Yu.P. Altukhov. Alleles of sAAT-1,2* isoloci in brown trout: potential diagnostic marker for tracking routes of post-glacial colonization in northern Europe. Journal of Fish Biology, 61 (3):842–846, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dieperink:2002:PAS

[829] C. Dieperink, B. D. Bak, L.-F. Pedersen, M. I. Pedersen, and S. Pedersen. Predation on Atlantic salmon and sea trout during their first days as postsmolts. Journal of Fish Biology, 61(3):848–852, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gauldie:2002:DNC

[830] R. W. Gauldie and A. Markwitz. Demonstration of a new calibration standard for strontium measurements in otoliths based on high energy Sr ion implantation. Journal of Fish Biology, 61(3):853–857, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oddone:2002:ECS

[831] M. C. Oddone and C. M. Vooren. Egg-cases and size at hatching of Sympterygia acuta in the south-western Atlantic. Journal of Fish Biology, 61(3):858–861, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:Eb

[832] Anonymous. Erratum. Journal of Fish Biology, 61(3):862, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:Ec

[833] Anonymous. Erratum. Journal of Fish Biology, 61(3):863, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:Ac

[834] Anonymous. Announcement. Journal of Fish Biology, 61(3):864, September 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vonHerbing:2002:ETL

[835] I. Hunt von Herbing. Effects of temperature on larval fish swimming performance: the importance of physics to physiology. Journal of Fish Biology, 61(4):865–876, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz:2002:OMS

[836] M. Muñoz, M. Casadevall, and S. Bonet. The ovarian morphology of Scorpaena notata shows a specialized mode of oviparity. Journal of Fish Biology, 61(4):877–887, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pacheco:2002:DGH

[837] F. J. Pacheco, S. O. S. Pacheco, H. R. C. Segreto, R. A. Segreto, M. R. R. Silva, and M. I. Egami. Development of granulocytes in haematopoietic tissues of Rhizoprionodon lalandii. Journal of Fish Biology, 61(4):888– 898, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suk:2002:FPM

[838] H. Y. Suk and J. C. Choe. Females prefer males with larger first dorsal fins in the common freshwater goby. Journal of Fish Biology, 61(4):899– 914, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeSilva:2002:SMB

[839] S. S. De Silva, R. M. Gunasekera, and R. O. Collins. Some morphometric and biochemical features of ready-to-migrate silver and pre-migratory yellow stages of the shortfin eel of south-eastern Australian waters. Journal of Fish Biology, 61(4):915–928, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cunha:2002:PRE

[840] C. Cunha, N. Mesquita, T. E. Dowling, A. Gilles, and M. M. Coelho. Phylogenetic relationships of Eurasian and American cyprinids using cytochrome b sequences. Journal of Fish Biology, 61(4):929–944, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

vonHerbing:2002:EBM

[841] I. Hunt von Herbing and L. White. The effects of body mass and feeding on metabolic rate in small juvenile Atlantic cod. Journal of Fish Biology, 61(4):945–958, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braccini:2002:IVE

[842] J. M. Braccini and G. E. Chiaramonte. Intraspecific variation in the external morphology of the sand skate. Journal of Fish Biology, 61(4): 959–972, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rojas-Garcia:2002:CTA

[843] C. R. Rojas-García and I. Rønnestad. Cholecystokinin and tryptic activity in the gut and body of developing Atlantic halibut larvae: evidence for participation in the regulation of protein digestion. Journal of Fish Biology, 61(4):973–986, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2002:SES

[844] C. Brown and K. Laland. Social enhancement and social inhibition of foraging behaviour in hatchery-reared Atlantic salmon. Journal of Fish Biology, 61(4):987–998, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2002:MWA

[845] B. A. Henderson and G. E. Morgan. Maturation of walleye by age, size and surplus energy. Journal of Fish Biology, 61(4):999–1011, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swanson:2002:RSP

[846] C. Swanson, D. V. Baxa, P. S. Young, J. J. Cech, Jr., and R. P. Hedrick. Reduced swimming performance in delta smelt infected with Mycobacterium spp. Journal of Fish Biology, 61(4):1012–1020, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beacham:2002:MII

[847] T. D. Beacham, B. McIntosh, and C. MacConnachie. Microsatellite identification of individual sockeye salmon in Barkley Sound, British Columbia. Journal of Fish Biology, 61(4):1021–1032, October 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Planes:2002:GDV

[848] S. Planes, G. Lecaillon, P. Lenfant, and M. Meekan. Genetic and demographic variation in new recruits of Naso unicornis. Journal of Fish Biology, 61(4):1033–1049, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skaret:2002:HC

[849] G. Skaret, B. E. Axelsen, L. Nøttestad, A. Fernö, and A. Johannessen. Herring as cannibals. Journal of Fish Biology, 61(4):1050–1052, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janech:2002:RWS

[850] M. G. Janech and P. M. Piermarini. Renal water and solute excretion in the Atlantic stingray in fresh water. Journal of Fish Biology, 61(4): 1053–1057, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caiola:2002:FRA

[851] N. Caiola and A. de Sostoa. First record of the Asiatic cyprinid Pseudorasbora parva in the Iberian Peninsula. Journal of Fish Biology, 61(4): 1058–1060, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corsini:2002:LMF

[852] M. Corsini, G. Kondilatos, and P. S. Economidis. Lessepsian migrant Fistularia commersonii from the Rhodes marine area. Journal of Fish Biology, 61(4):1061–1062, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reichard:2002:ELI

[853] M. Reichard, P. Jurajda, and M. Ondračková. The effect of light intensity on the drift of young-of-the-year cyprinid fishes. Journal of Fish Biology, 61(4):1063–1066, October 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neilson:2002:SAC

[854] J. D. Neilson, L. Annis, P. Perley, A. Clay, C. Croft, and M. O’Connor. Seasonal aggregations of Canadian east coast pollock as inferred from the commercial fishery and hydroacoustic observations. Journal of Fish Biology, 61(5):1067–1084, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borgstrosm:2002:HMI

[855] R. Borgstrosm, Ø. Skaala, and A. H. Aastveit. High mortality in introduced brown trout depressed potential gene flow to a wild population. Journal of Fish Biology, 61(5):1085–1097, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grutter:2002:ECF

[856] A. S. Grutter, M. R. Deveney, I. D. Whittington, and R. J. G. Lester. The effect of the cleaner fish Labroides dimidiatus on the capsalid monogenean Benedenia lolo parasite of the labrid fish Hemigymnus melapterus. Journal of Fish Biology, 61(5):1098–1108, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mehner:2002:MVH

[857] T. Mehner and M. Schulz. Monthly variability of hydroacoustic fish stock estimates in a deep lake and its correlation to gillnet catches. Journal of Fish Biology, 61(5):1109–1121, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eldridge:2002:EED

[858] M. B. Eldridge, E. C. Norton, B. M. Jarvis, and R. B. Macfarlane. Energetics of early development in the viviparous yellowtail rockfish. Journal of Fish Biology, 61(5):1122–1134, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gauthier:2002:AOD

[859] S. Gauthier and G. A. Rose. Acoustic observation of diel vertical migration and shoaling behaviour in Atlantic redfishes. Journal of Fish Biology, 61(5):1135–1153, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borcherding:2002:LMF

[860] J. Borcherding, M. Bauerfeld, D. Hintzen, and D. Neumann. Lateral migrations of fishes between floodplain lakes and their drainage channels at the Lower Rhine: diel and seasonal aspects. Journal of Fish Biology, 61(5):1154–1170, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imre:2002:PPB

[861] I. Imre, R. L. McLaughlin, and D. L. G. Noakes. Phenotypic plasticity in brook charr: changes in caudal fin induced by water flow. Journal of Fish Biology, 61(5):1171–1181, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Declerck:2002:PDO

[862] S. Declerck, G. Louette, T. De Bie, and L. De Meester. Patterns of diet overlap between populations of non-indigenous and native fishes in shallow ponds. Journal of Fish Biology, 61(5):1182–1197, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magill:2002:SIV

[863] S. H. Magill and M. D. J. Sayer. Seasonal and interannual variation in fish assemblages of northern temperate rocky subtidal habitats. Journal of Fish Biology, 61(5):1198–1216, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gisbert:2002:MDA

[864] E. Gisbert, G. Merino, J. B. Muguet, D. Bush, R. H. Piedrahita, and D. E. Conklin. Morphological development and allometric growth patterns in hatchery-reared California halibut larvae. Journal of Fish Biology, 61(5):1217–1229, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2002:RAS

[865] K. Morita and S. H. Morita. Rule of age and size at maturity: individual variation in the maturation history of resident white-spotted charr. Journal of Fish Biology, 61(5):1230–1238, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frade:2002:OSM

[866] P. Frade, P. C. Hubbard, E. N. Barata, and A. V. M. Canario. Olfactory sensitivity of the Mozambique tilapia to conspecific odours. Journal of Fish Biology, 61(5):1239–1254, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hill:2002:ASJ

[867] S. Hill, M. T. Burrows, and R. N. Hughes. Adaptive search in juvenile plaice foraging for aggregated and dispersed prey. Journal of Fish Biology, 61(5):1255–1267, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2002:SCW

[868] J. M. Elliott. Shadow competition in wild juvenile sea-trout. Journal of Fish Biology, 61(5):1268–1281, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mussi:2002:GTS

[869] M. Mussi, A. P. Summers, and P. Domenici. Gait transition speed, pectoral fin-beat frequency and amplitude in Cymatogaster aggregata, Embiotoca lateralis and Damalichthys vacca. Journal of Fish Biology, 61(5):1282–1293, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shrimpton:2002:SCA

[870] J. M. Shrimpton and S. D. McCormick. Seasonal changes in androgen levels in stream- and hatchery-reared Atlantic salmon parr and their relationship to smolting. Journal of Fish Biology, 61(5):1294–1304, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2002:CCR

[871] E. I. Hasegawa, T. Saito, and J. Seki. Composition changes in retinal pigments according to habitat of chum and pink salmon. Journal of Fish Biology, 61(5):1305–1308, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feldheim:2002:GTD

[872] K. A. Feldheim, S. H. Gruber, J. R. C. de Marignac, and M. V. Ashley. Genetic tagging to determine passive integrated transponder tag loss in lemon sharks. Journal of Fish Biology, 61(5):1309–1313, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2002:ESD

[873] Y. Li, J. A. Hill, G. H. Yue, F. Chen, and L. Orban. Extensive search does not identify genomic sex markers in Tetraodon nigroviridis. Journal of Fish Biology, 61(5):1314–1317, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mustonen:2002:LPL

[874] A-M. Mustonen, P. Nieminen, and H. Hyvaurinen. Liver and plasma lipids of spawning burbot. Journal of Fish Biology, 61(5):1318–1322, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRd

[875] Anonymous. Book reviews. Journal of Fish Biology, 61(5):1323–1325, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRe

[876] Anonymous. Book received. Journal of Fish Biology, 61(5):1326, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:NIS

[877] Anonymous. Ninth international symposium on the ecology of fluvial fishes. Journal of Fish Biology, 61(5):1327, November 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quist:2002:SVC

[878] M. C. Quist, C. S. Guy, R. J. Bernot, and J. L. Stephen. Seasonal variation in condition, growth and food habits of walleye in a Great Plains reservoir and simulated effects of an altered thermal regime. Journal of Fish Biology, 61(6):1329–1344, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solbakken:2002:DIT

[879] J. S. Solbakken, M. H. G. Berntessen, B. Norberg, K. Pittman, and K. Hamre. Different iodine and thyroid hormone levels between Atlantic halibut larvae fed wild zooplankton or Artemia from first exogenous feeding until post metamorphosis. Journal of Fish Biology, 61(6):1345–1362, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedini:2002:GCI

[880] V. Pedini, P. Scocco, A. M. Gargiulo, P. Ceccarelli, and S. Lorvik. Glycoconjugate characterization in the intestine of Umbrina cirrosa by means of lectin histochemistry. Journal of Fish Biology, 61(6):1363–1372, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitehead:2002:OSN

[881] A. L. Whitehead, B. O. David, and G. P. Closs. Ontogenetic shift in nocturnal microhabitat selection by giant kokopu in a New Zealand stream. Journal of Fish Biology, 61(6):1373–1385, December 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arlinghaus:2002:FRC

[882] R. Arlinghaus, C. Engelhardt, A. Sukhodolov, and C. Wolter. Fish recruitment in a canal with intensive navigation: implications for ecosystem management. Journal of Fish Biology, 61(6):1386–1402, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguirre:2002:PEC

[883] H. Aguirre, V. J. Madrid, and J. A. Virgen. Presence of Echinorhinus cookei off central Pacific Mexico. Journal of Fish Biology, 61(6):1403– 1409, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2002:IRT

[884] W.-M. Chen and M. Tabata. Individual rainbow trout can learn and anticipate multiple daily feeding times. Journal of Fish Biology, 61(6): 1410–1422, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maxime:2002:ETS

[885] V. Maxime. Effects of transfer to sea water on standard and routine metabolic rates in smolting Atlantic salmon at different stages of seawater adaptability. Journal of Fish Biology, 61(6):1423–1432, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harris:2002:PRM

[886] P. M. Harris, R. L. Mayden, H. S. Espinosa Pérez, and F. García de Leon. Phylogenetic relationships of Moxostoma and Scartomyzon (Catostomidae) based on mitochondrial cytochrome b sequence data. Journal of Fish Biology, 61(6):1433–1452, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luksenburg:2002:SGV

[887] J. A. Luksenburg and T. Pedersen. Sexual and geographical variation in life history parameters of the shorthorn sculpin. Journal of Fish Biology, 61(6):1453–1464, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carcupino:2002:FSM

[888] M. Carcupino, A. Baldacci, M. Mazzini, and P. Franzoi. Functional significance of the male brood pouch in the reproductive strategies of pipefishes and seahorses: a morphological and ultrastructural comparative study on three anatomically different pouches. Journal of Fish Biology, 61(6):1465–1480, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zajonz:2002:NBS

[889] U. Zajonz and W. Klausewitz. Neomerinthe bathypevimensis sp. nov. from deep waters of the southern Red Sea. Journal of Fish Biology, 61(6):1481–1488, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Broad:2002:RLU

[890] T. L. Broad, C. R. Townsend, G. P. Closs, and D. J. Jellyman. Riparian land use and accessibility to fishers influence size class composition and habitat use by longfin eels in a New Zealand river. Journal of Fish Biology, 61(6):1489–1503, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ogata:2002:MBC

[891] H. Y. Ogata. Muscle buffering capacity of yellowtail fed diets supplemented with crystalline histidine. Journal of Fish Biology, 61(6):1504– 1512, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vollestad:2002:GRV

[892] L. A. Vøllestad, E. M. Olsen, and T. Forseth. Growth-rate variation in brown trout in small neighbouring streams: evidence for densitydependence? Journal of Fish Biology, 61(6):1513–1527, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gage:2002:RBS

[893] M. J. G. Gage, C. Macfarlane, S. Yeates, R. Shackleton, and G. A. Parker. Relationships between sperm morphometry and sperm motility in the Atlantic salmon. Journal of Fish Biology, 61(6):1528–1539, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cutts:2002:FMF

[894] C. J. Cutts, N. B. Metcalfe, and A. C. Taylor. Fish may fight rather than feed in a novel environment: metabolic rate and feeding motivation in juvenile Atlantic salmon. Journal of Fish Biology, 61(6):1540–1548, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kime:2002:UMC

[895] D. E. Kime and H. Tveiten. Unusual motility characteristics of sperm of the spotted wolffish. Journal of Fish Biology, 61(6):1549–1559, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rincon:2002:IBI

[896] P. A. Rincón, A. M. Correas, F. Morcillo, P. Risueño, and J. Lobón-Cerviá. Interaction between the introduced eastern mosquitofish and two autochthonous Spanish toothcarps. Journal of Fish Biology, 61(6): 1560–1585, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nelson:2002:MTS

[897] J. A. Nelson. Metabolism of three species of herbivorous loricariid catfishes: influence of size and diet. Journal of Fish Biology, 61(6):1586– 1599, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Green:2002:CEM

[898] B. S. Green, S. M. Reilly, and M. I. McCormick. cost-effective method of preparing larval fish otoliths for reading using enzyme digestion and staining. Journal of Fish Biology, 61(6):1600–1605, December 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzoni:2002:ERL

[899] R. Mazzoni and R. Iglesias-Rios. Environmentally related life history variations in Geophagus brasiliensis. Journal of Fish Biology, 61(6):1606– 1618, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brinsmead:2002:MVB

[900] J. Brinsmead and M. G. Fox. Morphological variation between lakeand stream-dwelling rock bass and pumpkinseed populations. Journal of Fish Biology, 61(6):1619–1638, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deudero:2002:ULN

[901] S. Deudero. Unexpected large numbers of Mullus suvmuletus juveniles in open waters of the Mediterranean sampled with light attraction devices. Journal of Fish Biology, 61(6):1639–1642, December 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gante:2002:FRN

[902] H. F. Gante and C. D. Santos. First records of the North American catfish Ameiurus melas in Portugal. Journal of Fish Biology, 61(6):1643–1646, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Westin:2002:NDR

[903] L. Westin and K. E. Limburg. Newly discovered reproductive isolation reveals sympatric populations of Esox indus in the Baltic. Journal of Fish Biology, 61(6):1647–1652, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRf

[904] Anonymous. Book reviews. Journal of Fish Biology, 61(6):1653–1655, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:BRg

[905] Anonymous. Book received. Journal of Fish Biology, 61(6):1656, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blaber:2002:FHW

[906] S. J. M. Blaber. ‘fish in hot water”: the challenges facing fish and fisheries research in tropical estuaries. Journal of Fish Biology, 61(sA):1–20, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baldoa:2002:MAF

[907] F. Baldoa and P. Drake. A multivariate approach to the feeding habits of small fishes in the Guadalquivir Estuary. Journal of Fish Biology, 61 (sA):21–32, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barletta-Bergan:2002:CST

[908] A. Barletta-Bergan, M. Barletta, and U. Saint-Paul. Community structure and temporal variability of ichthyoplankton in North Brazilian mangrove creeks. Journal of Fish Biology, 61(sA):33–51, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chaves:2002:BTS

[909] P. Chaves, H. Pichler, and M. Robert. Biological, technical and socioeconomic aspects of the fishing activity in a Brazilian estuary. Journal of Fish Biology, 61(sA):52–59, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clough:2002:EMT

[910] S. C. Clough, A. W. H. Turnpenny, and S. D. J. Holden. Experimental measurement of thermal preferenda in sea trout smolts. Journal of Fish Biology, 61(sA):60–63, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colclough:2002:FFT

[911] S. R. Colclough, G. Gray, A. Bark, and B. Knights. Fish and fisheries of the tidal Thames: management of the modern resource, research aims and future pressures. Journal of Fish Biology, 61(sA):64–73, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowley:2002:BPE

[912] P. D. Cowley and A. K. Whitfield. Biomass and production estimates of a fish community in a small South African estuary. Journal of Fish Biology, 61(sA):74–89, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greenwood:2002:TAB

[913] M. F. D. Greenwood, A. S. Hill, and D. S. McLusky. Trends in abundance of benthic and demersal fish populations of the lower Forth Estuary, East Scotland, from 1982–2001. Journal of Fish Biology, 61(sA):90–104, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hadderingh:2002:CFI

[914] R. H. Hadderingh and Z. Jager. Comparison of fish impingement by a thermal power station with fish populations in the Ems Estuary. Journal of Fish Biology, 61(sA):105–124, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hiddink:2002:ARN

[915] J. G. Hiddink and Z. Jager. Abundance and reproduction of Nilsson’s pipefish on tidal flats. Journal of Fish Biology, 61(sA):125–137, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marian:2002:LCB

[916] S. Marian, A. Maccaroni, F. Massa, M. Rampacci, and L. Tancioni. Lack of consistency between the trophic interrelationships of five sparid species in two adjacent central Mediterranean coastal lagoons. Journal of Fish Biology, 61(sA):138–147, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pihl:2002:SDF

[917] L. Pihl and H. Wennhage. Structure and diversity of fish assemblages on rocky and soft bottom shores on the Swedish west coast. Journal of Fish Biology, 61(sA):148–166, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pombo:2002:CFF

[918] L. Pombo, M. Elliott, and J. E. Rebelo. Changes in the fish fauna of the Ria de Aveiro estuarine lagoon (Portugal) during the twentieth century. Journal of Fish Biology, 61(sA):167–181, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raz-Guzman:2002:FCT

[919] A. Raz-Guzman and L. Huidobro. Fish communities in two environmentally different estuarine systems of Mexico. Journal of Fish Biology, 61(sA):182–195, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vizziano:2002:RMF

[920] D. Vizziano, F. Forni, G. Saona, and W. Norbis. Reproduction of Micropogonias funieri in a shallow temperate coastal lagoon in the southern Atlantic. Journal of Fish Biology, 61(sA):196–206, December 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wennhage:2002:FFG

[921] H. Wennhage and L. Pihl. Fish feeding guilds in shallow rocky and soft bottom areas on the Swedish west coast. Journal of Fish Biology, 61(sA): 207–228, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2002:FIE

[922] A. K. Whitfield and M. Elliott. Fishes as indicators of environmental and ecological changes within estuaries: a review of progress and some suggestions for the future. Journal of Fish Biology, 61(sA):229–250, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kranenbarg:2002:RIN

[923] J. Kranenbarg, H. V. Winter, and J. J. G. M. Backx. Recent increase of North Sea houting and prospects for recolonization in The Netherlands. Journal of Fish Biology, 61(sA):251–253, December 2002. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:AOP

[924] Anonymous. Abstracts of oral papers not published in the supplement. Journal of Fish Biology, 61(sA):254–272, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2002:APP

[925] Anonymous. Abstracts of posters not published in the supplement. Journal of Fish Biology, 61(sA):273–281, December 2002. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mukuda:2003:MMN

[926] T. Mukuda and M. Ando. Medullary motor neurones associated with drinking behaviour of Japanese eels. Journal of Fish Biology, 62(1):1– 12, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingermann:2003:RST

[927] R. L. Ingermann, M. L. Robinson, and J. G. Cloud. Respiration of steelhead trout sperm: sensitivity to pH and carbon dioxide. Journal of Fish Biology, 62(1):13–23, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diaz:2003:DLL

[928] J. P. Diaz, M. Prié-Granié, M. Kentouri, S. Varsamos, and R. Connes. Development of the lateral line system in the sea bass. Journal of Fish Biology, 62(1):24–40, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lilja:2003:ESR

[929] J. Lilja and A. Romakkaniemi. Early-season river entry of adult Atlantic salmon: its dependency on environmental factors. Journal of Fish Biology, 62(1):41–50, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powell:2003:ABR

[930] M. D. Powell and B. F. Nowak. Acid–base and respiratory effects of confinement in Atlantic salmon affected with amoebic gill disease. Journal of Fish Biology, 62(1):51–63, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2003:FBM

[931] A. P. Farrell, C. G. Lee, K. Tierney, A. Hodaly, S. Clutterham, M. Healey, S. Hinch, and A. Lotto. Field-based measurements of oxygen uptake and swimming performance with adult Pacific salmon using a mobile respirometer swim tunnel. Journal of Fish Biology, 62(1):64–84, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Birtwell:2003:LSE

[932] I. K. Birtwell, J. S. Korstrom, and A. E. Brotherston. Laboratory studies on the effects of thermal change on the behaviour and distribution of juvenile chum salmon in sea water. Journal of Fish Biology, 62(1):85– 96, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Letcher:2003:ELH

[933] B. H. Letcher and G. Gries. Effects of life history variation on size and growth in stream-dwelling Atlantic salmon. Journal of Fish Biology, 62 (1):97–114, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2003:SEC

[934] F. Lahnsteiner. The spermatozoa and eggs of the cardinal fish. Journal of Fish Biology, 62(1):115–128, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bariche:2003:RCS

[935] M. Bariche, M. Harmelin-Vivien, and J.-P. Quignard. Reproductive cycles and spawning periods of two Lessepsian siganid fishes on the Lebanese coast. Journal of Fish Biology, 62(1):129–142, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Auster:2003:USW

[936] P. J. Auster, J. Lindholm, S. Schaub, G. Funnell, L. S. Kaufman, and P. C. Valentine. Use of sand wave habitats by silver hake. Journal of Fish Biology, 62(1):143–152, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kroon:2003:SHL

[937] F. J. Kroon, P. L. Munday, and N. W. Pankhurst. Steroid hormone levels and bi-directional sex change in Gobiodon histrio. Journal of Fish Biology, 62(1):153–167, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neuenfeldt:2003:OSC

[938] S. Neuenfeldt and J. E. Beyer. Oxygen and salinity characteristics of predator–prey distributional overlaps shown by predatory Baltic cod during spawning. Journal of Fish Biology, 62(1):168–183, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2003:ECF

[939] L. Wu, S. Xie, Y. Cui, and R. J. Wootton. Effect of cycles of feed deprivation on growth and food consumption of immature three-spined sticklebacks and European minnows. Journal of Fish Biology, 62(1):184– 194, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhu:2003:CGR

[940] X. Zhu, L. Wu, Y. Cui, Y. Yang, and R. J. Wootton. Compensatory growth response in three-spined stickleback in relation to feeddeprivation protocols. Journal of Fish Biology, 62(1):195–205, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laberge:2003:BER

[941] F. Laberge and T. J. Hara. Behavioural and electrophysiological responses to F-prostaglandins, putative spawning pheromones, in three salmonid fishes. Journal of Fish Biology, 62(1):206–221, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesthagen:2003:CHD

[942] T. Hesthagen and J. Heggenes. Competitive habitat displacement of brown trout by Siberian sculpin: the role of size and density. Journal of Fish Biology, 62(1):222–236, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayano:2003:ACM

[943] H. Hayano, Y. Miyakoshi, M. Nagata, K. Sugiwaka, and J. R. Irvine. Age composition of masu salmon smolts in northern Japan. Journal of Fish Biology, 62(1):237–241, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bodilis:2003:PYO

[944] P. Bodilis, A. Ganteaume, and P. Francour. Presence of 1 year-old dusky groupers along the French Mediterranean coast. Journal of Fish Biology, 62(1):242–246, January 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomkiewicz:2003:MMC

[945] J. Tomkiewicz, L. Tybjerg, and Å. Jespersen. Micro- and macroscopic characteristics to stage gonadal maturation of female Baltic cod. Journal of Fish Biology, 62(2):253–275, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trapani:2003:MVC

[946] J. Trapani. Morphological variability in the Cuatro Cienegas cichlid, Cichlasoma minckleyi. Journal of Fish Biology, 62(2):276–298, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bean:2003:GDE

[947] K. Bean, B. D. Mapstone, C. R. Davies, C. D. Murchie, and A. J. Williams. Gonad development and evidence of protogyny in the redthroat emperor on the Great Barrier Reef. Journal of Fish Biology, 62 (2):299–310, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skov:2003:PSS

[948] C. Skov, L. Jacobsen, and S. Berg. Post-stocking survival of 0+ year pike in ponds as a function of water transparency, habitat complexity, prey availability and size heterogeneity. Journal of Fish Biology, 62(2): 311–322, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lund:2003:FYS

[949] E. Lund, E. M. Olsen, and L. A. Vøllestad. First-year survival of brown trout in three Norwegian streams. Journal of Fish Biology, 62(2):323– 340, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:2003:AHD

[950] I. D. McCarthy, J. A. Sánchez, and G. Blanco. Allozyme heterozygosity, date of first feeding and life history strategy in Atlantic salmon. Journal of Fish Biology, 62(2):341–357, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardewig:2003:DCL

[951] I. Hardewig and P. L. M. Van Dijk. Is digestive capacity limiting growth at low temperatures in roach? Journal of Fish Biology, 62(2):358–374, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caputo:2003:RBC

[952] V. Caputo, M. L. Mesa, G. Candi, and P. N. Cerioni. The reproductive biology of the crystal goby with a comparison to that of the transparent goby. Journal of Fish Biology, 62(2):375–385, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baker:2003:EAP

[953] C. F. Baker. Effect of adult pheromones on the avoidance of suspended sediment by migratory banded kokopu juveniles. Journal of Fish Biology, 62(2):386–394, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2003:EFC

[954] N. G. Andersen and J. Riis-Vestergaard. The effects of food consumption rate, body size and temperature on net food conversion efficiency in saithe and whiting. Journal of Fish Biology, 62(2):395–412, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parmanne:2003:AVG

[955] R. Parmanne and K. Lindström. Annual variation in gobiid larval density in the northern Baltic Sea. Journal of Fish Biology, 62(2):413–426, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duftner:2003:EPI

[956] N. Duftner, S. Weiss, N. Medgyesy, and C. Sturmbauer. Enhanced phylogeographic information about Austrian brown trout populations derived from complete mitochondrial control region sequences. Journal of Fish Biology, 62(2):427–435, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bajer:2003:LET

[957] P. G. Bajer, G. W. Whitledge, R. S. Hayward, and R. D. Zweifel. Laboratory evaluation of two bioenergetics models applied to yellow perch: identification of a major source of systematic error. Journal of Fish Biology, 62(2):436–454, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davenport:2003:ACS

[958] J. Davenport. Allometric constraints on stability and maximum size in flying fishes: implications for their evolution. Journal of Fish Biology, 62 (2):455–463, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2003:VAT

[959] M. J. Morgan. Variation with age in the timing and duration of spawning in American plaice. Journal of Fish Biology, 62(2):464–473, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2003:SIA

[960] C. Adams, D. Fraser, I. McCarthy, S. Shields, S. Waldron, and G. Alexander. Stable isotope analysis reveals ecological segregation in a bimodal size polymorphism in Arctic charr from Loch Tay, Scotland. Journal of Fish Biology, 62(2):474–481, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crean:2003:ARE

[961] S. R. Crean, J. T. A. Dick, D. W. Evans, R. W. Elwood, and R. S. Rosell. Anal redness in European eels as an indicator of infection by the swimbladder nematode, Anguillicola crassus. Journal of Fish Biology, 62 (2):482–485, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aday:2003:MSI

[962] D. D. Aday, D. H. Wahl, and D. P. Philipp. A mechanism for social inhibition of sexual maturation in bluegill. Journal of Fish Biology, 62 (2):486–490, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorner:2003:PLE

[963] H. Dörner and J. Benndorf. Piscivory by large eels on young-of-the-year fishes: its potential as a biomanipulation tool. Journal of Fish Biology, 62 (2):491–494, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahman:2003:RCT

[964] M. S. Rahman, A. Takemura, S. Nakamura, and K. Takano. Rhythmic changes in testicular activity with lunar cycle in the forktail rabbitfish. Journal of Fish Biology, 62(2):495–499, February 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carton:2003:ERC

[965] A. G. Carton and J. C. Montgomery. Evidence of a rheotactic component in the odour search behaviour of freshwater eels. Journal of Fish Biology, 62(3):501–516, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2003:MPS

[966] B. K. Adams and J. A. Hutchings. Microgeographic population structure of brook charr: a comparison of microsatellite and mark-recapture data. Journal of Fish Biology, 62(3):517–533, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meyer:2003:IMB

[967] C. B. Meyer. The importance of measuring biotic and abiotic factors in the lower egg pocket to predict Coho salmon egg survival. Journal of Fish Biology, 62(3):534–548, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinje:2003:SSL

[968] F. Vinje, M. Heino, U. Dieckmann, O. R. GodØ, and J. Mork. Spatial structure in length at age of cod in the Barents Sea. Journal of Fish Biology, 62(3):549–564, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holker:2003:MRR

[969] F. Hölker. The metabolic rate of roach in relation to body size and temperature. Journal of Fish Biology, 62(3):565–579, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ojanguren:2003:TDE

[970] A. F. Ojanguren and F. Braña. Thermal dependence of embryonic growth and development in brown trout. Journal of Fish Biology, 62(3):580– 590, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamazaki:2003:MDS

[971] Y. Yamazaki, A. Goto, and M. Nishida. Mitochondrial DNA sequence divergence between two cryptic species of Lethenteron, with reference to an improved identification technique. Journal of Fish Biology, 62(3):591– 609, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kause:2003:BBQ

[972] A. Kause, O. Ritola, T. Paananen, U. Eskelinen, and E. Mäntysaari. Big and beautiful? Quantitative genetic parameters for appearance of large rainbow trout. Journal of Fish Biology, 62(3):610–622, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Altman:2003:MAH

[973] A. D. Altman and E. B. Taylor. A molecular assessment of homing in the tidepool sculpin. Journal of Fish Biology, 62(3):623–640, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2003:ASS

[974] B. Jonsson, N. Jonsson, and L. P. Hansen. Atlantic salmon straying from the River Imsa. Journal of Fish Biology, 62(3):641–657, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2003:SPD

[975] M. C. P. Amorim, P. J. Fonseca, and V. C. Almada. Sound production during courtship and spawning of Oreochromis mossambicus: male– female and male–male interactions. Journal of Fish Biology, 62(3):658– 672, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robertson:2003:ESI

[976] M. J. Robertson, D. A. Scruton, and J. A. Brown. Effects of surgically implanted transmitters on swimming performance, food consumption and growth of wild Atlantic salmon parr. Journal of Fish Biology, 62(3):673– 678, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cermeno:2003:VDI

[977] P. Cermeño, A. Uriarte, A. M. De Murguía, and B. Morales-Nin. Validation of daily increment formation in otoliths of juvenile and adult European anchovy. Journal of Fish Biology, 62(3):679–691, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Museth:2003:PBT

[978] J. Museth, R. BorgstrØm, T. Hame, and L. Å. Holen. Predation by brown trout: a major mortality factor for sexually mature European minnows. Journal of Fish Biology, 62(3):692–705, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clemmesen:2003:VCG

[979] C. Clemmesen, V. Bühler, G. Carvalho, R. Case, G. Evans, L. Hauser, W. F. Hutchinson, O. S. Kjesbu, H. Mempel, E. Moksness, H. Otteraa, H. Paulsen, A. Thorsen, and T. Svaasand. Variability in condition and growth of Atlantic cod larvae and juveniles reared in mesocosms: environmental and maternal effects. Journal of Fish Biology, 62(3):706–723, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2003:MME

[980] A. H. Rikardsen and M. Johansen. A morphometric method for estimation of total lipid level in live Arctic charr: a case study of its application on wild fish. Journal of Fish Biology, 62(3):724–734, March 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okland:2003:LUR

[981] F. Økland, C. J. Hay, T. F. Næsje, N. Nickandor, and E. B. Thorstad. Learning from unsuccessful radio tagging of common carp in a Namibian reservoir. Journal of Fish Biology, 62(3):735–739, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costantini:2003:NJS

[982] M. Costantini and M. Affronte. Neonatal and juvenile sandbar sharks in the northern Adriatic Sea. Journal of Fish Biology, 62(3):740–743, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rotllant:2003:DIT

[983] J. Rotllant, D. Montero, M. J. Caballero, L. Robaina, M. S. Izquierdo, and L. Tort. Differences in interrenal tissue, biosynthetic capacity and ACTH sensitivity in progeny of sea bream from parents selected for high or low cortisol response. Journal of Fish Biology, 62(3):744–748, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2003:SRG

[984] W. G. Anderson, J. P. Good, C. E. Franklin, and N. Hazon. Scaling of rectal gland mass in the European lesser-spotted dogfish. Journal of Fish Biology, 62(3):749–752, March 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2003:SPO

[985] C. G. Lee, R. H. Devlin, and A. P. Farrell. Swimming performance, oxygen consumption and excess post-exercise oxygen consumption in adult transgenic and ocean-ranched Coho salmon. Journal of Fish Biology, 62(4):753–766, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lyons:2003:RCM

[986] D. O. Lyons and J. J. Dunne. Reproductive costs to male and female worm pipefish. Journal of Fish Biology, 62(4):767–773, April 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huriaux:2003:EMP

[987] F. Huriaux, E. Baras, P. Vandewalle, and B. Focant. Expression of myofibrillar proteins and parvalbumin isoforms in white muscle of dorada during development. Journal of Fish Biology, 62(4):774–792, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peach:2003:BRP

[988] M. B. Peach. The behavioural role of pit organs in the epaulette shark. Journal of Fish Biology, 62(4):793–802, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guijarro:2003:SCH

[989] A. I. Guijarro, M. A. Lopez-Patiño, M. L. Pinillos, E. Isorna, N. De pedro, A. L. Alonso-Gómez, M. Alonso-Bedate, and M. J. Delgado. Seasonal changes in haematology and metabolic resources in the tench. Journal of Fish Biology, 62(4):803–815, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:2003:OLE

[990] H. M. Murray, J. W. Gallant, J. C. Perez-Casanova, S. C. Johnson, and S. E. Douglas. Ontogeny of lipase expression in winter flounder. Journal of Fish Biology, 62(4):816–833, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2003:ISP

[991] J. C. Svendsen, J. Skov, M. Bildsoe, and J. F. Steffensen. Intra-school positional preference and reduced tail beat frequency in trailing positions in schooling roach under experimental conditions. Journal of Fish Biology, 62(4):834–846, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knudsen:2003:IIM

[992] R. Knudsen, P.-A. Amundsen, and A. Klemetsen. Inter- and intra-morph patterns in helminth communities of sympatric whitefish morphs. Journal of Fish Biology, 62(4):847–859, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jansen:2003:SSZ

[993] P. A. Jansen, A. G. Finstad, and A. Langeland. Size-scaling of zooplankton foraging in Arctic charr. Journal of Fish Biology, 62(4):860– 870, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nielsen:2003:IGV

[994] J. L. Nielsen, I. Williams, G. K. Sage, and C. E. Zimmerman. The importance of genetic verification for determination of Atlantic salmon in north Pacific waters. Journal of Fish Biology, 62(4):871–878, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruyet:2003:LTA

[995] J. Person le Ruyet, A. Lamers, A. le Roux, A. Sévère, G. Boeuf, and N. Mayer-Gostan. Long-term ammonia exposure of turbot: effects on plasma parameters. Journal of Fish Biology, 62(4):879–894, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maack:2003:MDG

[996] G. Maack and H. Segner. Morphological development of the gonads in zebrafish. Journal of Fish Biology, 62(4):895–906, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Genner:2003:ICT

[997] M. J. Genner, S. J. Hawkins, and G. F. Turner. Isotopic change throughout the life history of a Lake Malaŵi cichlid fish. Journal of Fish Biology, 62(4):907–917, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmgren:2003:OSC

[998] K. Holmgren. Omitted spawning in compensatory-growing perch. Journal of Fish Biology, 62(4):918–927, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pichavant:2003:CSB

[999] K. Pichavant, V. Maxime, P. Soulier, G. Boeuf, and G. Nonnotte. A comparative study of blood oxygen transport in turbot and sea bass: effect of chronic hypoxia. Journal of Fish Biology, 62(4):928–937, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lemke:2003:UEC

[1000] A. M. Lemke, J. A. Stoeckel, and M. A. Pegg. Utilization of the exotic cladoceran Daphnia lumholtzi by juvenile fishes in an Illinois River floodplain lake. Journal of Fish Biology, 62(4):938–954, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payne:2003:ECS

[1001] A. G. Payne, C. Smith, and A. C. Campbell. The effect of clutch size on whole clutch cannibalism in the beaugregory damselfish. Journal of Fish Biology, 62(4):955–958, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsson:2003:RTD

[1002] J. I. Johnsson, J. Parkkonen, and L. Förlin. Reduced territorial defence in rainbow trout fry exposed to a paper mill effluent: using the mirror image stimulation test as a behavioural bioassay. Journal of Fish Biology, 62(4):959–964, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2003:EAR

[1003] S. J. Cooke. Externally attached radio transmitters do not affect the parental care behaviour of rock bass. Journal of Fish Biology, 62(4): 965–970, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sneddon:2003:BSI

[1004] L. U. Sneddon. The bold and the shy: individual differences in rainbow trout. Journal of Fish Biology, 62(4):971–975, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munday:2003:STE

[1005] P. L. Munday, M. Schubert, J. A. Baggio, G. P. Jones, M. J. Caley, and A. S. Grutter. Skin toxins and external parasitism of coral-dwelling gobies. Journal of Fish Biology, 62(4):976–981, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pawson:2003:IFM

[1006] Mike Pawson. Inshore fisheries management. Journal of Fish Biology, 62 (4):982, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2003:TEF

[1007] Mike Elliott. Tropical estuarine fishes. Journal of Fish Biology, 62(4): 983–984, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2003:TS

[1008] J. D. Armstrong. Trout and salmon. Journal of Fish Biology, 62(4):984– 985, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2003:VGI

[1009] Anonymous. Verhandlungen der Gesellschaft für Ichthyologie. Journal of Fish Biology, 62(4):986, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2003:BRI

[1010] Anonymous. Book review: Island in the Stream/Oceanography and Fisheries of The Charleston Bump. Journal of Fish Biology, 62(4):986, April 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanHerwerden:2003:MVP

[1011] L. Van Herwerden, J. Benzie, and C. Davies. Microsatellite variation and population genetic structure of the red throat emperor on the Great Barrier Reef. Journal of Fish Biology, 62(5):987–999, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2003:PIS

[1012] G. N. Wagner, R. S. McKinley, P. A. Bjørn, and B. Finstad. Physiological impact of sea lice on swimming performance of Atlantic salmon. Journal of Fish Biology, 62(5):1000–1009, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cotano:2003:GYY

[1013] U. Cotano and P. Álvarez. Growth of young-of-the-year mackerel in the Bay of Biscay. Journal of Fish Biology, 62(5):1010–1020, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorner:2003:SDP

[1014] H. Dörner and A. Wagner. Size-dependent predator–prey relationships between perch and their fish prey. Journal of Fish Biology, 62(5):1021– 1032, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deguara:2003:EAP

[1015] S. Deguara, K. Jauncey, and C. Agius. Enzyme activities and pH variations in the digestive tract of gilthead sea bream. Journal of Fish Biology, 62(5):1033–1043, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2003:AEG

[1016] F. B. Jensen, J. Brahm, P. Koldkjær, T. Wang, D. J. McKenzie, and E. W. Taylor. Anion exchange in the giant erythrocytes of African lungfish. Journal of Fish Biology, 62(5):1044–1052, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leggatt:2003:OUG

[1017] R. A. Leggatt, R. H. Devlin, A. P. Farrell, and D. J. Randall. Oxygen uptake of growth hormone transgenic Coho salmon during starvation and feeding. Journal of Fish Biology, 62(5):1053–1066, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robb:2003:VTH

[1018] T. Robb and M. V. Abrahams. Variation in tolerance to hypoxia in a predator and prey species: an ecological advantage of being small? Journal of Fish Biology, 62(5):1067–1081, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klemetsen:2003:HDF

[1019] A. Klemetsen, R. Knudsen, F. J. Staldvik, and P.-A. Amundsen. Habitat, diet and food assimilation of Arctic charr under the winter ice in two subarctic lakes. Journal of Fish Biology, 62(5):1082–1098, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillespie:2003:MLH

[1020] G. J. Gillespie and M. G. Fox. Morphological and life-history differentiation between littoral and pelagic forms of pumpkinseed. Journal of Fish Biology, 62(5):1099–1115, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stequert:2003:RBV

[1021] B. Stequert, F. Menard, and E. Marchal. Reproductive biology of Vinciguerria nimbaria in the equatorial waters of the eastern Atlantic Ocean. Journal of Fish Biology, 62(5):1116–1136, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pouilly:2003:DMR

[1022] M. Pouilly, F. Lino, J.-G. Bretenoux, and C. Rosales. Dietary– morphological relationships in a fish assemblage of the Bolivian Amazonian floodplain. Journal of Fish Biology, 62(5):1137–1158, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bjornevik:2003:MGJ

[1023] M. Bjørnevik, C. Beattie, T. Hansen, and A. Kiessling. Muscle growth in juvenile Atlantic salmon as influenced by temperature in the egg and yolk sac stages and diet protein level. Journal of Fish Biology, 62(5):1159– 1175, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoner:2003:HLL

[1024] A. W. Stoner. Hunger and light level alter response to bait by Pacific halibut: laboratory analysis of detection, location and attack. Journal of Fish Biology, 62(5):1176–1193, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jin:2003:FAS

[1025] X. Jin, B. Xu, and Q. Tang. Fish assemblage structure in the East China Sea and southern Yellow Sea during autumn and spring. Journal of Fish Biology, 62(5):1194–1205, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sequeira:2003:NAR

[1026] V. Sequeira, I. Figueiredo, M. Muñoz, and L. S. Gordo. New approach to the reproductive biology of Helicolenus dactylopterus. Journal of Fish Biology, 62(5):1206–1210, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kramer:2003:RCP

[1027] C. R. Kramer and H. Potter. Rodlet cells in the posterior intestine of embryos and neonates of two poeciliid species. Journal of Fish Biology, 62(5):1211–1216, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svensson:2003:MMS

[1028] O. Svensson and E. Forsgren. Male mating success in relation to food availability in the common goby. Journal of Fish Biology, 62(5):1217– 1221, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finstad:2003:GBB

[1029] A. G. Finstad. Growth backcalculations based on otoliths incorporating an age effect: adding an interaction term. Journal of Fish Biology, 62 (5):1222–1225, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2003:PEO

[1030] T. Yamamoto and U. G. Reinhardt. Paternal effect on otolith size at emergence in masu salmon. Journal of Fish Biology, 62(5):1226–1231, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ball:2003:LSG

[1031] A. O. Ball, G. R. Sedberry, J. H. Wessel III, and R. W. Chapman. Largescale genetic differentiation of Pagrus pagrus in the Atlantic. Journal of Fish Biology, 62(5):1232–1237, May 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2003:IV

[1032] Anonymous. Index, volume 62. Journal of Fish Biology, 62(6):9–30, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sartoris:2003:TDC

[1033] F. J. Sartoris, C. Bock, I. Serendero, G. Lannig, and H. O. Pörtner. Temperature-dependent changes in energy metabolism, intracellular pH and blood oxygen tension in the Atlantic cod. Journal of Fish Biology, 62(6):1239–1253, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vogler:2003:TES

[1034] R. Vögler, A. C. Milessi, and R. A. Quiñones. Trophic ecology of Squatina guggenheim on the continental shelf off Uruguay and northern Argentina. Journal of Fish Biology, 62(6):1254–1267, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Assis:2003:LOT

[1035] C. A. Assis. The lagenar otoliths of teleosts: their morphology and its application in species identification, phylogeny and systematics. Journal of Fish Biology, 62(6):1268–1295, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2003:ROH

[1036] H. H. Ng. A review of the Ompok hypophthalmus group of silurid catfishes with the description of a new species from South-East Asia. Journal of Fish Biology, 62(6):1296–1311, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmen:2003:ICB

[1037] J. Holmen, E. M. Olsen, and L. A. Vøllestad. Interspecific competition between stream-dwelling brown trout and Alpine bullhead. Journal of Fish Biology, 62(6):1312–1325, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benhaim:2003:BVJ

[1038] D. Benhaïm, S. Skúlason, and B. R. Hansen. Behavioural variation in juvenile Arctic charr in relation to body size. Journal of Fish Biology, 62(6):1326–1338, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:2003:RAC

[1039] A. D. Jordan, P. R. Møller, and J. G. Nielsen. Revision of the Arctic cod genus Arctogadus. Journal of Fish Biology, 62(6):1339–1352, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karpouzi:2003:RBM

[1040] V. S. Karpouzi and K. I. Stergiou. The relationships between mouth size and shape and body length for 18 species of marine fishes and their trophic implications. Journal of Fish Biology, 62(6):1353–1365, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2003:ESM

[1041] S. J. Cooke, B. D. S. Graeb, C. D. Suski, and K. G. Ostrand. Effects of suture material on incision healing, growth and survival of juvenile largemouth bass implanted with miniature radio transmitters: case study of a novice and experienced fish surgeon. Journal of Fish Biology, 62(6): 1366–1380, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2003:RMR

[1042] F. C. Neat, L. Locatello, and M. B. Rasotto. Reproductive morphology in relation to alternative male reproductive tactics in Scartella cristata. Journal of Fish Biology, 62(6):1381–1391, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gido:2003:EGS

[1043] K. B. Gido. Effects of gizzard shad on benthic communities in reservoirs. Journal of Fish Biology, 62(6):1392–1404, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gozdowska:2003:PCC

[1044] M. Gozdowska, E. Sokołowska, and E. Kulczykowska. Plasma Ca^{2+} concentration limits melatonin night production in two fish species. Journal of Fish Biology, 62(6):1405–1413, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leatherland:2003:VSG

[1045] J. F. Leatherland, K. Ogasawara, M. S. Rahman, R. Renaud, H. Yamashiro, and A. Takemura. In vitro steroidogenesis of the gonads of the protogynous Pacific wrasse, Haliochoeres trimaculatus. Journal of Fish Biology, 62(6):1414–1434, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morison:2003:PMS

[1046] A. K. Morison, I. W. Brown, and G. K. Jones. Post mortem shrinkage of four species of temperate and tropical marine fishes, without freezing or preservation. Journal of Fish Biology, 62(6):1435–1449, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morato:2003:RAG

[1047] T. Morato, P. Afonso, R. S. Santos, H. M. Krug, and R. D. M. Nash. The reproduction, age and growth of the spotted rockling. Journal of Fish Biology, 62(6):1450–1455, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koski:2003:ELS

[1048] M. L. Koski, B. M. Johnson, and T. M. Stecklein. Effects of light on size-selectivity of Kokanee feeding on Daphnia. Journal of Fish Biology, 62(6):1456–1461, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ocalewicz:2003:PSL

[1049] K. Ocalewicz and I. Babiak. Primed in situ labelling detection of 5S rDNA sequences in normal and androgenetic rainbow trout. Journal of Fish Biology, 62(6):1462–1466, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whitfield:2003:RFF

[1050] A. K. Whitfield and T. D. Harrison. River flow and fish abundance in a South African estuary. Journal of Fish Biology, 62(6):1467–1472, June 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kock:2003:SNL

[1051] K.-H. Kock and I. Everson. Shedding new light on the life cycle of mackerel icefish in the Southern Ocean. Journal of Fish Biology, 63(1): 1–21, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2003:MOG

[1052] S. Adams. Morphological ontogeny of the gonad of three plectropomid species through sex differentiation and transition. Journal of Fish Biology, 63(1):22–36, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manica:2003:EBS

[1053] A. Manica. The effect of brood size and age on partial filial cannibalism in the scissortail sergeant. Journal of Fish Biology, 63(1):37–47, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cara:2003:ADE

[1054] J. B. Cara, F. J. Moyano, S. Cárdenas, C. Fernández-Díaz, and M. Yúfera. Assessment of digestive enzyme activities during larval development of white bream. Journal of Fish Biology, 63(1):48–58, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morato:2003:RBR

[1055] T. Morato, P. Afonso, P. Lourinho, R. D. M. Nash, and R. S. Santos. Reproductive biology and recruitment of the white sea bream in the Azores. Journal of Fish Biology, 63(1):59–72, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fornies:2003:RPF

[1056] M. A. Forniés, M. Carrillo, E. Mañanós, L. A. Sorbera, Y. Zohar, and S. Zanuy. Relative potency of the forms of GnRH and their analogs on LH release in sea bass. Journal of Fish Biology, 63(1):73–89, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Timi:2003:PAA

[1057] J. T. Timi. Parasites of Argentine anchovy in the south-west Atlantic: latitudinal patterns and their use for discrimination of host populations. Journal of Fish Biology, 63(1):90–107, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corriero:2003:HIO

[1058] A. Corriero, S. Desantis, M. Deflorio, F. Acone, C. R. Bridges, J. M. De la Serna, P. Megalofonou, and G. De Metrio. Histological investigation on the ovarian cycle of the bluefin tuna in the western and central Mediterranean. Journal of Fish Biology, 63(1):108–119, July 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winemiller:2003:FHT

[1059] K. O. Winemiller and L. C. Kelso-Winemiller. Food habits of tilapiine cichlids of the Upper Zambezi River and floodplain during the descending phase of the hydrologic cycle. Journal of Fish Biology, 63(1):120–128, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davison:2003:HPB

[1060] W. Davison and C. E. Franklin. Hypertension in Pagothenia borchgrevinki caused by X-cell disease. Journal of Fish Biology, 63(1):129– 136, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:2003:SDG

[1061] C. T. Walsh, B. C. Pease, and D. J. Booth. Sexual dimorphism and gonadal development of the Australian longfinned river eel. Journal of Fish Biology, 63(1):137–152, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:2003:EES

[1062] H. R. King, N. W. Pankhurst, M. Watts, and P. M. Pankhurst. Effect of elevated summer temperatures on gonadal steroid production, vitellogenesis and egg quality in female Atlantic salmon. Journal of Fish Biology, 63(1):153–167, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jegstrup:2003:RPL

[1063] I. M. Jegstrup and P. Rosenkilde. Regulation of post-larval development in the European eel: thyroid hormone level, progress of pigmentation and changes in behaviour. Journal of Fish Biology, 63(1):168–175, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tolonen:2003:SGY

[1064] A. Tolonen, J. Lappalainen, and E. Pulliainen. Seasonal growth and year class strength variations of perch near the northern limits of its distribution range. Journal of Fish Biology, 63(1):176–186, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priede:2003:LAS

[1065] I. G. Priede, A. R. Deary, D. M. Bailey, and K. L. Smith, Jr. Low activity and seasonal change in population size structure of grenadiers in the oligotrophic abyssal central North Pacific Ocean. Journal of Fish Biology, 63(1):187–196, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reddy:2003:IPA

[1066] P. K. Reddy and J. F. Leatherland. Influences of photoperiod and alternate days of feeding on plasma growth hormone and thyroid hormone levels in juvenile rainbow trout. Journal of Fish Biology, 63(1):197–212, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konishi:2003:NHB

[1067] M. Konishi, K. Hosoya, and K. Takata. Natural hybridization between endangered and introduced species of pseudorasbora, with their genetic relationships and characteristics inferred from allozyme analyses. Journal of Fish Biology, 63(1):213–231, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dezfuli:2003:ARC

[1068] B. S. Dezfuli, L. Giari, E. Simoni, D. Palazzi, and M. Manera. Alteration of rodlet cells in chub caused by the herbicide Stam(R) M-4 (propanil). Journal of Fish Biology, 63(1):232–239, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elofsson:2003:LLS

[1069] H. Elofsson, B. G. Mcallister, D. E. Kime, I. Mayer, and B. Borg. Long lasting stickleback sperm; is ovarian fluid a key to success in fresh water? Journal of Fish Biology, 63(1):240–253, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maes:2003:IRH

[1070] G. E. Maes, J. K. J. Van Houdt, D. De Charleroy, and F. A. M. Volckaert. Indications for a recent Holarctic expansion of pike based on a preliminary study of mtDNA variation. Journal of Fish Biology, 63(1): 254–259, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jepsen:2003:LTR

[1071] N. Jepsen. Long-term retention of surgically implanted radio transmitters in pikeperch. Journal of Fish Biology, 63(1):260–262, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Copp:2003:FCC

[1072] G. H. Copp. Is fish condition correlated with water conductivity? Journal of Fish Biology, 63(1):263–266, July 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salewski:2003:SSL

[1073] V. Salewski. Satellite species in lampreys: a worldwide trend for ecological speciation in sympatry? Journal of Fish Biology, 63(2):267–279, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nissling:2003:SGV

[1074] A. Nissling, A. Müller, and H.-H. Hinrichsen. Specific gravity and vertical distribution of sprat eggs in the Baltic Sea. Journal of Fish Biology, 63 (2):280–299, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magalhaes:2003:RLH

[1075] M. F. Magalhães, I. J. Schlosser, and M. J. Collares-Pereira. The role of life history in the relationship between population dynamics and environmental variability in two Mediterranean stream fishes. Journal of Fish Biology, 63(2):300–317, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmberg:2003:OIM

[1076] A. Holmberg, T. Schwerte, R. Fritsche, B. Pelster, and S. Holmgren. Ontogeny of intestinal motility in correlation to neuronal development in zebrafish embryos and larvae. Journal of Fish Biology, 63(2):318– 331, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ide:2003:ROB

[1077] L. M. Ide, E. C. Urbinati, and A. Hoffmann. The role of olfaction in the behavioural and physiological responses to conspecific skin extract in Brycon cephalus. Journal of Fish Biology, 63(2):332–343, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vincent:2003:CRS

[1078] A. C. J. Vincent and B. G. Giles. Correlates of reproductive success in a wild population of Hippocampus whitei. Journal of Fish Biology, 63 (2):344–355, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murua:2003:PSG

[1079] H. Murua. Population structure, growth and reproduction of roughhead grenadier on the Flemish Cap and Flemish Pass. Journal of Fish Biology, 63(2):356–373, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2003:BMD

[1080] F. C. Neat, W. Lengkeek, E. P. Westerbeek, B. Laarhoven, and J. J. Videler. Behavioural and morphological differences between lake and river populations of Salaria fluviatilis. Journal of Fish Biology, 63(2): 374–387, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jardine:2003:DYS

[1081] D. Jardine and M. K. Litvak. Direct yolk sac volume manipulation of zebrafish embryos and the relationship between offspring size and yolk sac volume. Journal of Fish Biology, 63(2):388–397, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2003:EGH

[1082] D. J. McKenzie, R. Martínez, A. Morales, J. Acosta, R. Morales, E. W. Taylor, J. F. Steffensen, and M. P. Estrada. Effects of growth hormone transgenesis on metabolic rate, exercise performance and hypoxia tolerance in tilapia hybrids. Journal of Fish Biology, 63(2):398–409, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duque:2003:DSA

[1083] A. Barbarino Duque and K. O. Winemiller. Dietary segregation among large catfishes of the Apure and Arauca Rivers, Venezuela. Journal of Fish Biology, 63(2):410–427, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arkhipkin:2003:VDP

[1084] A. Arkhipkin, P. Brickle, and V. Laptikhovsky. Variation in the diet of the Patagonian toothfish with size, depth and season around the Falkland Islands. Journal of Fish Biology, 63(2):428–441, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arrington:2003:DCS

[1085] D. A. Arrington and K. O. Winemiller. Diel changeover in sandbank fish assemblages in a neotropical floodplain river. Journal of Fish Biology, 63(2):442–459, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cavalli:2003:DGE

[1086] L. Cavalli, N. Pech, and R. Chappaz. Diet and growth of the endangered Zingel asper in the Durance River. Journal of Fish Biology, 63(2):460– 471, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carolsfeld:2003:CSB

[1087] J. Carolsfeld, H. P. Godinho, E. Zaniboni Filho, and B. J. Harvey. Cryopreservation of sperm in Brazilian migratory fish conservation. Journal of Fish Biology, 63(2):472–489, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pouilly:2003:MRC

[1088] M. Pouilly and G. Miranda. Morphology and reproduction of the cavefish Trichomycterus chaberti and the related epigean Trichomycterus cf. barbouri. Journal of Fish Biology, 63(2):490–505, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peterson:2003:EAS

[1089] B. C. Peterson, P. R. Simpson, K. D. Cain, R. H. Hardy, G. T. Schelling, and T. L. Ott. Effects of administration of somatostatin-14 and immunoneutralization of somatostatin on endocrine and growth responses in rainbow trout. Journal of Fish Biology, 63(2):506–522, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nokhbatolfoghahai:2003:CCS

[1090] M. Nokhbatolfoghahai and J. R. Downie. Ciliated cells on the surface of embryos of Scyliorhinus canicula. Journal of Fish Biology, 63(2):523– 527, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2003:RBS

[1091] D. Gonçalves, T. Fagundes, and R. Oliveira. Reproductive behaviour of sneaker males of the peacock blenny. Journal of Fish Biology, 63(2):528– 532, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vasconcelos:2003:FRG

[1092] P. Vasconcelos, M. N. Santos, and M. B. Gaspar. First record of Grammicolepis brachiusculus in Portuguese waters. Journal of Fish Biology, 63(2):533–537, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinet:2003:ASM

[1093] T. Robinet, M. Sbaihi, S. Guyet, B. Mounaix, S. Dufour, and E. Feunteun. Advanced sexual maturation before marine migration of Anguilla bicolor bicolor and Anguilla marmorata at Réunion Island. Journal of Fish Biology, 63(2):538–542, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2003:BRa

[1094] Anonymous. Book reviews. Journal of Fish Biology, 63(2):543–545, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2003:BRb

[1095] Anonymous. Books received. Journal of Fish Biology, 63(2):546, August 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petry:2003:RBF

[1096] P. Petry, P. B. Bayley, and D. F. Markle. Relationships between fish assemblages, macrophytes and environmental gradients in the Amazon River floodplain. Journal of Fish Biology, 63(3):547–579, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stierhoff:2003:HTM

[1097] K. L. Stierhoff, T. E. Targett, and P. A. Grecay. Hypoxia tolerance of the mummichog: the role of access to the water surface. Journal of Fish Biology, 63(3):580–592, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanssens:2003:NSG

[1098] M. Hanssens and J. Snoeks. A new species and geographical variation in the Telmatochromis temporalis complex (Teleostei, Cichlidae) from Lake Tanganyika. Journal of Fish Biology, 63(3):593–616, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamashita:2003:CGF

[1099] Y. Yamashita, O. Tominaga, H. Takami, and H. Yamada. Comparison of growth, feeding and cortisol level in Platichthys bicoloratus juveniles between estuarine and nearshore nursery grounds. Journal of Fish Biology, 63(3):617–630, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldsmith:2003:EVI

[1100] R. J. Goldsmith, G. P. Closs, and H. Steen. Evaluation of visible implant elastomer for individual marking of small perch and common bully. Journal of Fish Biology, 63(3):631–636, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:2003:DVM

[1101] L. P. Weber, P. S. Higgins, R. I. Carlson, and D. M. Janz. Development and validation of methods for measuring multiple biochemical indices of condition in juvenile fishes. Journal of Fish Biology, 63(3):637–658, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kahilainen:2003:PPS

[1102] K. Kahilainen and H. Lehtonen. Piscivory and prey selection of four predator species in a whitefish dominated subarctic lake. Journal of Fish Biology, 63(3):659–672, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petersson:2003:GSI

[1103] E. Petersson and T. Järvi. Growth and social interactions of wild and sea-ranched brown trout and their hybrids. Journal of Fish Biology, 63 (3):673–686, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marchand:2003:DTB

[1104] F. Marchand, P. Magnan, and D. Boisclair. Differential time budgets of two forms of juvenile brook charr in the open-water zone. Journal of Fish Biology, 63(3):687–698, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva-Paterna:2003:CBS

[1105] F. J. Oliva-Paterna, A. Vila-Gispert, and M. Torralva. Condition of Barbus sclateri from semi-arid aquatic systems: effects of habitat quality disturbances. Journal of Fish Biology, 63(3):699–709, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fredrich:2003:SMC

[1106] F. Fredrich, S. Ohmann, B. Curio, and F. Kirschbaum. Spawning migrations of the chub in the River Spree, Germany. Journal of Fish Biology, 63 (3):710–723, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leis:2003:CRS

[1107] J. M. Leis, B. M. Carson-Ewart, A. C. Hay, and D. H. Cato. Coralreef sounds enable nocturnal navigation by some reef-fish larvae in some places and at some times. Journal of Fish Biology, 63(3):724–737, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2003:SLP

[1108] C. Brown, A. Markula, and K. Laland. Social learning of prey location in hatchery-reared Atlantic salmon. Journal of Fish Biology, 63(3):738–745, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martyniuk:2003:GAC

[1109] C. J. Martyniuk, G. M. L. Perry, H. K. Mogahadam, M. M. Ferguson, and R. G. Danzmann. The genetic architecture of correlations among growthrelated traits and male age at maturation in rainbow trout. Journal of Fish Biology, 63(3):746–764, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2003:SDM

[1110] B. R. Moore, R. C. Buckworth, H. Moss, and R. J. G. Lester. Stock discrimination and movements of narrow-barred Spanish mackerel across northern Australia as indicated by parasites. Journal of Fish Biology, 63 (3):765–779, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlson:2003:VBB

[1111] S. M. Carlson and B. H. Letcher. Variation in brook and brown trout survival within and among seasons, species, and age classes. Journal of Fish Biology, 63(3):780–794, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2003:DPE

[1112] S.-L. Chen, H.-Q. Ye, Z.-X. Sha, and Y. Hong. Derivation of a pluripotent embryonic cell line from red sea bream blastulas. Journal of Fish Biology, 63(3):795–805, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jhingan:2003:DRS

[1113] E. Jhingan, R. H. Devlin, and G. K. Iwama. Disease resistance, stress response and effects of triploidy in growth hormone transgenic Coho salmon. Journal of Fish Biology, 63(3):806–823, September 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnhagen:2003:SLY

[1114] C. Magnhagen and F. Staffan. Social learning in young-of-the-year perch encountering a novel food type. Journal of Fish Biology, 63(3):824–829, September 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keith:2003:BEA

[1115] P. Keith. Biology and ecology of amphidromous Gobiidae of the Indo– Pacific and the Caribbean regions. Journal of Fish Biology, 63(4):831– 847, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campana:2003:SGP

[1116] S. E. Campana, R. D. Stanley, and S. Wischniowski. Suitability of glycerin-preserved otoliths for age validation using bomb radiocarbon. Journal of Fish Biology, 63(4):848–854, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2003:IDD

[1117] K. H. Olsén, M. Grahn, and J. Lohm. The influence of dominance and diet on individual odours in MHC identical juvenile Arctic charr siblings. Journal of Fish Biology, 63(4):855–862, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnsson:2003:GSI

[1118] J. I. Johnsson. Group size influences foraging effort independent of predation risk: an experimental study on rainbow trout. Journal of Fish Biology, 63(4):863–870, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winter:2003:MBI

[1119] H. V. Winter and F. Fredrich. Migratory behaviour of ide: a comparison between the lowland rivers Elbe, Germany, and Vecht, The Netherlands. Journal of Fish Biology, 63(4):871–880, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sabates:2003:LCE

[1120] A. Sabatés, S. Rossi, and E. Reyes. Lipid content in the early life stages of three mesopelagic fishes. Journal of Fish Biology, 63(4):881–891, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

BijdeVaate:2003:STM

[1121] A. Bij de Vaate, A. W. Breukelaar, T. Vriese, G. De Laak, and C. Dijkers. Sea trout migration in the Rhine delta. Journal of Fish Biology, 63(4): 892–908, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Primmer:2003:POF

[1122] C. R. Primmer, P.-A. Landry, E. Ranta, J. Merilä, J. Piironen, K. Tiira, N. Peuhkuri, S. Pakkasmaa, and P. Eskelinen. Prediction of offspring fitness based on parental genetic diversity in endangered salmonid populations. Journal of Fish Biology, 63(4):909–927, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neves:2003:BSS

[1123] F. M. Neves and L. R. Monteiro. Body shape and size divergence among populations of Poecilia vivipara in coastal lagoons of south-eastern Brazil. Journal of Fish Biology, 63(4):928–941, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stefansson:2003:PCW

[1124] S. O. Stefansson, B. Th. Björnsson, K. Sundell, G. Nyhammer, and S. D. McCormick. Physiological characteristics of wild Atlantic salmon postsmolts during estuarine and coastal migration. Journal of Fish Biology, 63(4):942–955, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paakkonen:2003:DVB

[1125] J.-P. J. Pääkkönen, O. Tikkanen, and J. Karjalainen. Development and validation of a bioenergetics model for juvenile and adult burbot. Journal of Fish Biology, 63(4):956–969, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geist:2003:RBM

[1126] D. R. Geist, R. S. Brown, V. I. Cullinan, M. G. Mesa, S. P. VanderKooi, and C. A. McKinstry. Relationships between metabolic rate, muscle electromyograms and swim performance of adult Chinook salmon. Journal of Fish Biology, 63(4):970–989, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brophy:2003:DEL

[1127] D. Brophy, B. S. Danilowicz, and T. E. Jeffries. The detection of elements in larval otoliths from Atlantic herring using laser ablation ICP-MS. Journal of Fish Biology, 63(4):990–1007, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kowalski:2003:PAE

[1128] R. Kowalski, J. Glogowski, D. Kucharczyk, K. Goryczko, S. Dobosz, and A. Ciereszko. Proteolytic activity and electrophoretic profiles of proteases from seminal plasma of teleosts. Journal of Fish Biology, 63 (4):1008–1019, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanLook:2003:ASM

[1129] K. J. W. Van Look and D. E. Kime. Automated sperm morphology analysis in fishes: the effect of mercury on goldfish sperm. Journal of Fish Biology, 63(4):1020–1033, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vehanen:2003:AFB

[1130] T. Vehanen. Adaptive flexibility in the behaviour of juvenile Atlantic salmon: short-term responses to food availability and threat from predation. Journal of Fish Biology, 63(4):1034–1045, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnold:2003:SSS

[1131] K. E. Arnold, A. Adam, K. J. Orr, R. Griffiths, and I. Barber. Sexspecific survival and parasitism in three-spined sticklebacks: seasonal patterns revealed by molecular analysis. Journal of Fish Biology, 63(4): 1046–1050, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovacic:2003:HGK

[1132] M. Kovačić. Hyperbenthic gobies in the Kvarner area, Adriatic Sea. Journal of Fish Biology, 63(4):1051–1055, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brewster:2003:FGM

[1133] J. Brewster and A. Houde. Are female guppies more likely to flee when approached by two males? Journal of Fish Biology, 63(4):1056–1059, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrod:2003:BR

[1134] Chris Harrod, Melanie L. J. Stiassny, Peter Miller, and Ben Wilson. Book reviews. Journal of Fish Biology, 63(4):1060–1065, October 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2003:FAR

[1135] D. J. Jellyman and P. W. Lambert. Factors affecting recruitment of glass eels into the Grey River, New Zealand. Journal of Fish Biology, 63(5): 1067–1079, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richter:2003:EPF

[1136] H. Richter, C. Lückstädt, U. Focken, and K. Becker. Evacuation of pelleted feed and the suitability of titanium(IV) oxide as a feed marker for gut kinetics in Nile tilapia. Journal of Fish Biology, 63(5):1080–1099, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bichuette:2003:ESI

[1137] M. E. Bichuette and E. Trajano. Epigean and subterranean ichthyofauna from the São Domingos karst area, Upper Tocantins River basin, Central Brazil. Journal of Fish Biology, 63(5):1100–1121, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kacem:2003:HDV

[1138] A. Kacem and F. J. Meunier. Halastatic demineralization in the vertebrae of Atlantic salmon, during their spawning migration. Journal of Fish Biology, 63(5):1122–1130, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heins:2003:REN

[1139] D. C. Heins, J. M. Johnson, and J. A. Baker. Reproductive ecology of the nine-spined stickleback from south-central Alaska. Journal of Fish Biology, 63(5):1131–1143, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Theriault:2003:BSA

[1140] V. Thériault and J. J. Dodson. Body size and the adoption of a migratory tactic in brook charr. Journal of Fish Biology, 63(5):1144–1159, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Diogo:2003:OMC

[1141] R. Diogo, M. Chardon, and P. Vandewalle. Osteology and myology of the cephalic region and pectoral girdle of Erethistes pusillus, comparison with other erethistids, and comments on the synapomorphies and phylogenetic relationships of the Erethistidae (Teleostei: Siluriformes). Journal of Fish Biology, 63(5):1160–1175, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2003:ESA

[1142] E. H. Williams, Jr., A. A. Mignucci-Giannoni, L. Bunkley-Williams, R. K. Bonde, C. Self-Sullivan, A. Preen, and V. G. Cockcroft. Echeneid– sirenian associations, with information on sharksucker diet. Journal of Fish Biology, 63(5):1176–1183, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nielsen:2003:PMD

[1143] C. Nielsen, K. Aarestrup, U. Nørum, and bS. S. Madsen. Pre-migratory differentiation of wild brown trout into migrant and resident individuals. Journal of Fish Biology, 63(5):1184–1196, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perga:2003:UWS

[1144] M. E. Perga and D. Gerdeaux. Using the δ ^{13}C and δ ^{15}N of whitefish scales for retrospective ecological studies: changes in isotope signatures during the restoration of Lake Geneva, 1980–2001. Journal of Fish Biology, 63(5):1197–1207, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robertson:2003:IIM

[1145] M. J. Robertson, K. D. Clarke, D. A. Scruton, and J. A. Brown. Interhabitat and instream movements of large Atlantic salmon parr in a Newfoundland watershed in winter. Journal of Fish Biology, 63(5):1208–1218, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hinz:2003:ERT

[1146] H. Hinz, M. J. Kaiser, M. Bergmann, S. I. Rogers, and M. J. Armstrong. Ecological relevance of temporal stability in regional fish catches. Journal of Fish Biology, 63(5):1219–1234, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Darbyson:2003:DVF

[1147] E. Darbyson, D. P. Swain, D. Chabot, and M. Castonguay. Diel variation in feeding rate and prey composition of herring and mackerel in the southern Gulf of St Lawrence. Journal of Fish Biology, 63(5):1235–1257, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prodocimo:2003:GRT

[1148] V. Prodocimo and C. A. Freire. Glomeruli and renal tubules are restricted to the cranial kidney of the adult estuarine Sphoeroides testudineus. Journal of Fish Biology, 63(5):1258–1265, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2003:IGC

[1149] H. Zhang, K. Futami, and N. Okamoto. Isolation of c-myc genes from goldfish and their tissue specific expression. Journal of Fish Biology, 63(5):1266–1273, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keskinen:2003:GPR

[1150] T. Keskinen and T. J. Marjomäki. Growth of pikeperch in relation to lake characteristics: total phosphorus, water colour, lake area and depth. Journal of Fish Biology, 63(5):1274–1282, November 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hossain:2003:CGP

[1151] M. A. R. Hossain, M. Tagawa, R. Masuda, and M. Tanaka. Changes in growth performance and proximate composition in Japanese flounder during metamorphosis. Journal of Fish Biology, 63(5):1283–1294, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofmann:2003:ITF

[1152] N. Hofmann and P. Fischer. Impact of temperature on food intake and growth in juvenile burbot. Journal of Fish Biology, 63(5):1295–1305, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:2003:IAC

[1153] G. C. Young and I. C. Potter. Induction of annual cyclical changes in the ichthyofauna of a large microtidal estuary following an artificial and permanent increase in tidal flow. Journal of Fish Biology, 63(5):

1306–1330, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leduc:2003:DCH

[1154] A. O. H. C. Leduc, M. K. Noseworthy, J. C. Adrian, Jr., and G. E. Brown. Detection of conspecific and heterospecific alarm signals by juvenile pumpkinseed under weak acidic conditions. Journal of Fish Biology, 63(5):1331–1336, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wells:2003:EIR

[1155] A. Wells, W. G. Anderson, and N. Hazon. Evidence for an intrarenal renin-angiotensin system in the European lesser-spotted dogfish. Journal of Fish Biology, 63(5):1337–1340, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguirre:2003:AGS

[1156] W. E. Aguirre. Allometric growth of the sulcus in Cynoscion spp. (Sciaenidae). Journal of Fish Biology, 63(5):1341–1346, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fritsches:2003:CVP

[1157] K. A. Fritsches, L. Litherland, N. Thomas, and J. Shand. Cone visual pigments and retinal mosaics in the striped marlin. Journal of Fish Biology, 63(5):1347–1351, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takayanagi:2003:SMC

[1158] S. Takayanagi, Y. Sakai, H. Hashimoto, and K. Gushima. Sleeping mound construction using coral fragments by the rockmover wrasse. Journal of Fish Biology, 63(5):1352–1356, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heupel:2003:RBS

[1159] M. R. Heupel, C. A. Simpfendorfer, and R. E. Hueter. Running before the storm: blacktip sharks respond to falling barometric pressure associated with Tropical Storm Gabrielle. Journal of Fish Biology, 63(5):1357–1363, November 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nikora:2003:EFS

[1160] V. I. Nikora, J. Aberle, B. J. F. Biggs, I. G. Jowett, and J. R. E. Sykes. Effects of fish size, time-to-fatigue and turbulence on swimming performance: a case study of Galaxias maculatus. Journal of Fish Biology, 63(6):1365–1382, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomas:2003:MCA

[1161] J. Tomás and A. J. Geffen. Morphometry and composition of aragonite and vaterite otoliths of deformed laboratory reared juvenile herring from two populations. Journal of Fish Biology, 63(6):1383–1401, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cinquetti:2003:HHE

[1162] R. Cinquetti and L. Dramis. Histological, histochemical, enzyme histochemical and ultrastructural investigations of the testis of Padogobius martensi between annual breeding seasons. Journal of Fish Biology, 63(6):1402–1428, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elofsson:2003:ISO

[1163] H. Elofsson, K. Van Look, B. Borg, and I. Mayer. Influence of salinity and ovarian fluid on sperm motility in the fifteen-spined stickleback. Journal of Fish Biology, 63(6):1429–1438, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morrison:2003:SER

[1164] C. M. Morrison, B. Pohajdak, M. Henry, and J. R. Wright, Jr. Structure and enzymatic removal of the chorion of embryos of the Nile tilapia. Journal of Fish Biology, 63(6):1439–1453, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeLeeuw:2003:BSD

[1165] J. J. De Leeuw, L. A. J. Nagelkerke, W. L. T. Van Densen, K. Holmgren, P. A. Jansen, and J. Vijverberg. Biomass size distributions as a tool for characterizing lake fish communities. Journal of Fish Biology, 63(6): 1454–1475, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sabates:2003:FPV

[1166] A. Sabatés, A. Bozzano, and I. Vallvey. Feeding pattern and the visual light environment in myctophid fish larvae. Journal of Fish Biology, 63(6):1476–1490, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuset:2003:CMS

[1167] V. M. Tuset, A. Lombarte, J. A. González, J. F. Pertusa, and MaJ. Lorente. Comparative morphology of the sagittal otolith in Serranus spp. Journal of Fish Biology, 63(6):1491–1504, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brule:2003:RBG

[1168] T. Brulé, C. Déniel, T. Colás-Marrufo, and X. Renán. Reproductive biology of gag in the southern Gulf of Mexico. Journal of Fish Biology, 63(6):1505–1520, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zydlewski:2003:LMS

[1169] J. Zydlewski, S. D. McCormick, and J. G. Kunkel. Late migration and seawater entry is physiologically disadvantageous for American shad juveniles. Journal of Fish Biology, 63(6):1521–1537, December 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schonhuth:2003:EFE

[1170] S. Schönhuth, G. Luikart, and I. Doadrio. Effects of a founder event and supplementary introductions on genetic variation in a captive breeding population of the endangered Spanish killifish. Journal of Fish Biology, 63(6):1538–1551, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrne:2003:TEI

[1171] C. J. Byrne, R. Poole, G. Rogan, M. Dillane, and K. F. Whelan. Temporal and environmental influences on the variation in Atlantic salmon smolt migration in the Burrishoole system 1970–2000. Journal of Fish Biology, 63(6):1552–1564, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alvarez:2003:PAB

[1172] D. Álvarez and A. G. Nicieza. Predator avoidance behaviour in wild and hatchery-reared brown trout: the role of experience and domestication. Journal of Fish Biology, 63(6):1565–1577, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valle:2003:ODP

[1173] A. Z. Dalla Valle, R. Rivas-Diaz, and G. Claireaux. Opercular differential pressure as a predictor of metabolic oxygen demand in the starry flounder. Journal of Fish Biology, 63(6):1578–1588, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jan:2003:DTS

[1174] R.-Q. Jan, C.-T. Ho, and F.-K. Shiah. Determinants of territory size of the dusky gregory. Journal of Fish Biology, 63(6):1589–1597, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elliott:2003:ATM

[1175] J. P. Elliott and D. R. Bellwood. Alimentary tract morphology and diet in three coral reef fish families. Journal of Fish Biology, 63(6):1598–1609, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maclean:2003:CAF

[1176] A. Maclean, M. S. Miles, and J. D. Armstrong. A computer-activated food delivery system for dispensing natural prey items among complex arrays of patches. Journal of Fish Biology, 63(6):1610–1614, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2003:ALR

[1177] R. F. Oliveira, K. Hirschenhauser, A. V. M. Canário, and M. Taborsky. Androgen levels of reproductive competitors in a co-operatively breeding cichlid. Journal of Fish Biology, 63(6):1615–1620, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rico:2003:SOE

[1178] M. R. Rico and E. M. Acha. Southernmost occurrence of Epinephelus marginatus in the south-west Atlantic. Journal of Fish Biology, 63(6): 1621–1624, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OByrne-Ring:2003:CMC

[1179] N. O’Byrne-Ring, K. Dowling, D. Cotter, K. Whelan, and U. MacEvilly. Changes in mucus cell numbers in the epidermis of the Atlantic salmon at the onset of smoltification. Journal of Fish Biology, 63(6):1625–1630, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milinski:2003:FMC

[1180] M. Milinski. The function of mate choice in sticklebacks: optimizing Mhc genetics. Journal of Fish Biology, 63(s1):1–16, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amundsen:2003:FMS

[1181] T. Amundsen. Fishes as models in studies of sexual selection and parental care. Journal of Fish Biology, 63(s1):17–52, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sims:2003:TMT

[1182] D. W. Sims. Tractable models for testing theories about natural strategies: foraging behaviour and habitat selection of free-ranging sharks. Journal of Fish Biology, 63(s1):53–73, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braithwaite:2003:UWF

[1183] V. A. Braithwaite and J. R. Girvan. Use of water flow direction to provide spatial information in a small-scale orientation task. Journal of Fish Biology, 63(s1):74–83, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:2003:BMI

[1184] S. C. Mills and J. D. Reynolds. The bitterling–mussel interaction as a test case for co-evolution. Journal of Fish Biology, 63(s1):84–104, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gill:2003:DPC

[1185] A. B. Gill. The dynamics of prey choice in fish: the importance of prey size and satiation. Journal of Fish Biology, 63(s1):105–116, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hill:2003:EAS

[1186] S. L. Hill, M. T. Burrows, and R. N. Hughes. The efficiency of adaptive search tactics for different prey distribution patterns: a simulation model based on the behaviour of juvenile plaice. Journal of Fish Biology, 63 (s1):117–130, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gozlan:2003:RSM

[1187] R. E. Gozlan, C. J. Flower, and A. C. Pinder. Reproductive success in male sunbleak, a recent invasive fish species in the U.K. Journal of Fish Biology, 63(s1):131–143, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolm:2003:DIB

[1188] N. Kolm and J. Olsson. Differential investment in the Banggai cardinalfish: can females adjust egg size close to egg maturation to match the attractiveness of a new partner? Journal of Fish Biology, 63(s1): 144–151, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dou:2003:GPA

[1189] S. Dou, M. J. Miller, and K. Tsukamoto. Growth, pigmentation and activity of juvenile Japanese eels in relation to temperature and fish size. Journal of Fish Biology, 63(s1):152–165, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orpwood:2003:EBS

[1190] J. E. Orpwood, S. W. Griffiths, and J. D. Armstrong. Effects of body size on sympatric shelter use in over-wintering juvenile salmonids. Journal of Fish Biology, 63(s1):166–173, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hawkins:2003:SHU

[1191] L. A. Hawkins, J. D. Armstrong, and A. E. Magurran. Settlement and habitat use by juvenile pike in early winter. Journal of Fish Biology, 63 (s1):174–186, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2003:EEP

[1192] C. Brown, T. Davidson, and K. Laland. Environmental enrichment and prior experience of live prey improve foraging behaviour in hatcheryreared Atlantic salmon. Journal of Fish Biology, 63(s1):187–196, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinnegar:2003:DDC

[1193] J. K. Pinnegar, V. M. Trenkel, A. N. Tidd, W. A. Dawson, and M. H. Du buit. Does diet in Celtic Sea fishes reflect prey availability? Journal of Fish Biology, 63(s1):197–212, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dytham:2003:HLA

[1194] C. Dytham. How landscapes affect the evolution of dispersal behaviour in reef fishes: results from an individual-based model. Journal of Fish Biology, 63(s1):213–225, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2003:LMA

[1195] J. D. Armstrong, C. F. Holm, P. S. Kemp, and D. J. Gilvear. Linking models of animal behaviour and habitat management: Atlantic salmon parr and river discharge. Journal of Fish Biology, 63(s1):226, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braithwaite:2003:FCW

[1196] V. Braithwaite. Fish cognition: what can we learn, and what do they need to learn? Journal of Fish Biology, 63(s1):226–227, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2003:FCA

[1197] S. Griffiths, J. Höjesjö, and J. Johnsson. Familiarity confers anti-predator and foraging advantages on juvenile brown trout. Journal of Fish Biology, 63(s1):226, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2003:RSL

[1198] C. Brown and K. Laland. The role of social learning in conservation and fisheries reintroductions. Journal of Fish Biology, 63(s1):227, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Candolin:2003:BMS

[1199] U. Candolin and J. D. Reynolds. Bitterling as models for studies of sperm competition. Journal of Fish Biology, 63(s1):227, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carter:2003:SPC

[1200] T. Carter. Stream position choice in salmon parr: maximisation of foraging efficiency. Journal of Fish Biology, 63(s1):228, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cheney:2003:CBC

[1201] K. L. Cheney and I. M. Côté. Costs and benefits of a cleaning mutualism. Journal of Fish Biology, 63(s1):228, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2003:CCM

[1202] I. M. Côté. The cost of cleanerfish mimics to their models. Journal of Fish Biology, 63(s1):228–229, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeGraaf:2003:TPF

[1203] M. De Graaf, F. A. Sibbing, and J. W. M. Osse. Toothless predators: feeding performance and techniques among the piscivores within Lake Tana’s endemic barbus species flock (Pisces: Cyprinidae). Journal of Fish Biology, 63(s1):229, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dejen:2003:DSB

[1204] E. Dejen and F. A. Sibbing. Diet segregation between two sympatric ‘small’ Barbus spp: an experimental study of mechanisms. Journal of Fish Biology, 63(s1):229–230, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dulvy:2003:DPP

[1205] N. K. Dulvy and N. V. C. Polunin. Detecting predator–prey relationships in the sea. Journal of Fish Biology, 63(s1):230, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Edwards:2003:IBM

[1206] H. J. Edwards. Individual-based models of cod movement and population dynamics. Journal of Fish Biology, 63(s1):230, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Enders:2003:ITF

[1207] E. C. Enders, T. Buffin-Bélanger, D. Boisclair, and A. G. Roy. On the interaction of turbulent flow and feeding behaviour of juvenile Atlantic salmon (Salmo salar). Journal of Fish Biology, 63(s1):231, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirvonen:2003:IAP

[1208] H. Hirvonen, S. Vilhunen, C. Brown, V. Lintunen, and K. N. Laland. Improving anti-predator responses of hatchery reared salmonids by social learning. Journal of Fish Biology, 63(s1):232, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hojesjo:2003:RBD

[1209] J. Höjesjö, L. Stradmeyer, J. Armstrong, and S. Griffiths. The relationship between dominance behaviour and growth as a function of habitat stability and community complexity: a test using salmon and trout. Journal of Fish Biology, 63(s1):232, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hunter:2003:TSM

[1210] E. Hunter, J. D. Metcalfe, G. P. Arnold, and J. D. Reynolds. Tidal stream mediation of metapopulation structure in North Sea plaice. Journal of Fish Biology, 63(s1):233, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutchings:2003:SBD

[1211] J. A. Hutchings. Sex-biased dispersal in a salmonid fish: implications for growth rate and survival. Journal of Fish Biology, 63(s1):233, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:2003:ASC

[1212] W. C. Jordan, J. P. Dukes, R. Deaville, M. W. Bruford, and A. F. Youngson. Atlantic salmon change the expression of odorant receptor genes to learn the smell of home. Journal of Fish Biology, 63(s1):233–234, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2003:NMC

[1213] Y-H. Kim. New modelling of complex fish migration by application of chaos theory and neural network. Journal of Fish Biology, 63(s1):234, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolm:2003:CHB

[1214] N. Kolm. Causes of homing behaviour in a group living reef fish. Journal of Fish Biology, 63(s1):234–235, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krause:2003:SNF

[1215] J. Krause, D. Croft, and R. James. Social networks in fish. Journal of Fish Biology, 63(s1):235, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kvarnemo:2003:MSC

[1216] C. Kvarnemo, L. W. Simmons, M. Ah-King, and H. Elofsson. The myth of the sperm cloud — or why is testes size not affected by spawning mode in seahorses and pipefishes. Journal of Fish Biology, 63(s1):235, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laland:2003:WTB

[1217] K. N. Laland. Why there is better evidence for culture in fish than chimpanzees. Journal of Fish Biology, 63(s1):236, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindstrom:2003:RIR

[1218] K. Lindström, C. St Mary, B. Gunnels, R. Hale, C. Osenberg, S. Stevens, J. Vonesh, and J. Wilson. Reproductive investment in relation to survival risk in a livebearing fish. Journal of Fish Biology, 63(s1):236, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacLean:2003:FDR

[1219] A. MacLean, F. A. Huntingford, J. D. Armstrong, and G. D. Ruxton. Fish don’t read textbooks: juvenile salmon prove ignorant of foraging theory. Journal of Fish Biology, 63(s1):236–237, December 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2003:BTG

[1220] J. Metcalfe and E. Hunter. Being there and getting there: balancing the costs of migration against the benefits of changing habitat for North Sea plaice. Journal of Fish Biology, 63(s1):237–238, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzoldi:2003:YBH

[1221] C. Mazzoldi, M. Massironi, and M. B. Rasotto. Yellow belly as honest signal of female quality in Knipowitschia panizzae (Gobiidae). Journal of Fish Biology, 63(s1):237, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mills:2003:OSR

[1222] S. C. Mills and J. D. Reynolds. Operational sex ratio and resource defence as predictors of the mating system in European bitterling. Journal of Fish Biology, 63(s1):238, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nash:2003:GBM

[1223] J. Nash. Gender bent but not in the mind of fish. Journal of Fish Biology, 63(s1):238–239, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Righton:2003:TFB

[1224] D. Righton. Territorial and foraging behaviour of two corallivorous butterflyfishes. Journal of Fish Biology, 63(s1):239–240, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Payne:2003:AFC

[1225] A. Payne. Adaptive filial cannibalism in the beaugregory damselfish. Journal of Fish Biology, 63(s1):240, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowe:2003:SBA

[1226] S. Rowe and J. A. Hutchings. Spawning behaviour and acoustic communication in Atlantic cod. Journal of Fish Biology, 63(s1):240, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mary:2003:CCT

[1227] C. M. St Mary and K. Lindström. Components of care: are they honest signals of parental care quality? Journal of Fish Biology, 63(s1):240–241, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salvanes:2003:CVB

[1228] A. G. V. Salvanes, J. E. Skjæraasen, and T. Nilsen. Countergradient variation in behaviour, growth and lipid storage in Atlantic cod; environmental and genetic effects. Journal of Fish Biology, 63(s1):241, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simpson:2003:DBE

[1229] S. D. Simpson, C. Dytham, and M. G. Meekan. Does behaviour explain the larval dispersal of coral reef fishes better than oceanography? — An individual based model simulation approach. Journal of Fish Biology, 63(s1):241, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2003:EBM

[1230] C. Smith. The European bitterling: a model for oviposition decision. Journal of Fish Biology, 63(s1):241–242, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stradmeyer:2003:BSS

[1231] L. Stradmeyer, J. Höjesjö, J. Armstrong, and S. Griffiths. Behavioural strategies for species co-existence during environmental extremes: competition between trout and salmon parr during dewatering. Journal of Fish Biology, 63(s1):242, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taborsky:2003:ECA

[1232] M. Taborsky. The evolution of cooperation and advanced social behaviour. Journal of Fish Biology, 63(s1):242, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2003:BIB

[1233] C. M. Wagner and G. D. Grossman. Behavioural interactions between native smokey dace and introduced yellowfin shiner: implications from habitat selection theory for an ongoing invasion. Journal of Fish Biology, 63(s1):243, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walling:2003:DCM

[1234] C. A. Walling, N. Dawnay, A. J. N. Kazem, R. Hickling, and J. Wright. Do competing males cooperate? Familiarity and its effect on cooperation during predator inspection in male three-spined sticklebacks (Gasterosteus aculeatus). Journal of Fish Biology, 63(s1):243–244, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:2003:EHD

[1235] A. J. W. Ward, J. Krause, and P. J. B. Hart. The effects of habitat- and diet-based cues on association preferences in the three-spined stickleback (Gasterosteus aculeatus). Journal of Fish Biology, 63(s1):244, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whiteman:2003:IDM

[1236] E. A. Whiteman and I. M. Côté. Individual differences in microhabitat use in a Caribbean cleaning goby: a buffer effect in a marine species? Journal of Fish Biology, 63(s1):244–245, December 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeates:2003:FSC

[1237] S. Yeates and M. Gage. Fertilisation and sperm competition dynamics in salmon. Journal of Fish Biology, 63(s1):245, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:2003:ELI

[1238] I. Barber, J. Young, E. Morrison-Smith, and E. Matthews. Effects of light intensity and spectrum on fish behaviour in laboratory aquaria. Journal of Fish Biology, 63(s1):246, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barber:2003:SBP

[1239] I. Barber and P. A. Svensson. Synchrony between parasite development and host behaviour change. Journal of Fish Biology, 63(s1):246, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolliet:2003:ESE

[1240] V. Bolliet, A. Bardonnet, and J.-C. Vignes. Effects of substrate embeddedness on juvenile brown trout (Salmo trutta L.) growth and behaviour. Journal of Fish Biology, 63(s1):247, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Botham:2003:PPI

[1241] M. Botham and J. Krause. Predator prey interactions: predator choice for prey group size and composition in Crenicichla alta, a natural predator of the Trinidadian guppy, Poecilia reticulata. Journal of Fish Biology, 63(s1):247, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carter:2003:RBD

[1242] C. G. Carter, A-W. A. Abdel-Warith, and S. J. Davies. The relationship between dominance rank and feed intake following the introduction of novel feeds to African catfish. Journal of Fish Biology, 63(s1):248–247, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coombs:2003:DLZ

[1243] M. F. Coombs and M. S. Ridgway. Diffusion in the littoral zone: scoping emergence times and movement to essential habitat for young-of-year brook trout in lakes. Journal of Fish Biology, 63(s1):248, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Copp:2003:FBR

[1244] G. H. Copp, M. G. Carter, and H. Faulkner. Fish behavioural responses to river discharge trends and the importance of habitat connectivity. Journal of Fish Biology, 63(s1):248–249, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2003:ETP

[1245] I. M. Côté and P. P. Molloy. Ectoparasites: are they the proximate cause of cleaning interactions? Journal of Fish Biology, 63(s1):249, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2003:EVC

[1246] I. M. Côté and E. A. Whiteman. Up or down? Explaining the variability of communication signals in cleaning symbioses. Journal of Fish Biology, 63(s1):249–250, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davies:2003:TSA

[1247] R. Davies and S. Humphries. Telemetering sub aqua video photography to boat- and shore-bases: a new approach. Journal of Fish Biology, 63 (s1):250, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeGaudemar:2003:VRB

[1248] B. De Gaudemar and E. Beall. Variation of reproductive behaviour and success of males adopting different tactics in Atlantic salmon. Journal of Fish Biology, 63(s1):250, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dou:2003:SDA

[1249] S. Dou, R. Msuda, M. Tanaka, and K. Tsukamoto. Size dominance affecting the social interactions and growth of juvenile Japanese flounder, Paralichthys olivaceus. Journal of Fish Biology, 63(s1):251, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giacomello:2003:FVM

[1250] E. Giacomello. Function and variability of male sex traits in a blenniid fish, Parablennius tentacularis. Journal of Fish Biology, 63(s1):251, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Zuarth:2003:AGM

[1251] C. Gonzalez-Zuarth, E. Avila, and C. Macías Garcia. The Amarillo (Girardinichthys multiradiatus) a fish model to study geographic variation in mate recognition systems. Journal of Fish Biology, 63(s1):251–252, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lenhardt:2003:MRT

[1252] M. Lenhardt, P. Cakic, J. Kolarevic, and Z. Gacic. Morphometric recognition of two morphs in sterlet (Acipenser ruthenus) population induced by different reproductive behaviour. Journal of Fish Biology, 63(s1):252, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Linden:2003:ASG

[1253] M. Lindén, C. Påhlson, and J. Nilsson. The aberrant stickleback Gasterosteus aculeatus in the Karlskrona archipelago, SE Sweden. Journal of Fish Biology, 63(s1):252–253, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malavasi:2003:AMT

[1254] S. Malavasi, M. Lugli, P. Torricelli, and D. Mainardi. Alternative mating tactics and acoustic behaviour in the grass goby Zosterisessor ophiocephalus: a behavioural analysis of the interaction between the parasitic and the bourgeois male. Journal of Fish Biology, 63(s1):253, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malek:2003:PIF

[1255] M. Malek. Parasites as indicator of fish behaviour. Journal of Fish Biology, 63(s1):253–254, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OConnor:2003:ELI

[1256] E. O’Connor and J. Krause. Effect of light intensity on the shoaling behaviour of the guppy (Poecilia reticulata). Journal of Fish Biology, 63(s1):254, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ondrackova:2003:ODA

[1257] M. Ondračková, E. Mazurová, M. Reichard, and P. Jurajda. Ontogeny of the diel activity patterns in relation to downstream dispersal in riverine cyprinid fishes. Journal of Fish Biology, 63(s1):254, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Priest:2003:DEL

[1258] M. A. Priest, S. D. Simpson, and C. Dytham. Do early life history characteristics determine species ranges and dispersal of coral reef fishes in the Indo–Pacific? Journal of Fish Biology, 63(s1):254–255, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reichard:2003:BMS

[1259] M. Reichard. The bitterling–mussel symbiosis: a model for host-parasite adaptation. Journal of Fish Biology, 63(s1):255, December 2003. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roussel:2003:VIF

[1260] J-M. Roussel and B. Agnès. Visual isolation furthers access to driftfeeding positions for subordinate juvenile brown trout (Salmo trutta) in dominance hierarchies. Journal of Fish Biology, 63(s1):255–256, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salvanes:2003:THC

[1261] A. G. V. Salvanes and V. Braithwaite. Teaching hatchery cod about environmental heterogeneity. Journal of Fish Biology, 63(s1):256, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sari:2003:ASS

[1262] A. Sari. Adaptive significance of sex change in the stone loach (Nemachilus angorae). Journal of Fish Biology, 63(s1):256, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sherman:2003:YSU

[1263] R. L. Sherman, J. W. Barnes, J. P. Huston, and R. E. Spieler. The yellow stingray, Urobatis jamaicensis, as a model for studying cerebellar function in vertebrates. Journal of Fish Biology, 63(s1):256–257, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sparrevohn:2003:ESD

[1264] C. R. Sparrevohn and A. Nielsen. Estimating size dependent diet, advection, diffusion, catchability and mortality applying a normal distribution model to spreading of released turbot. Journal of Fish Biology, 63(s1): 257, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sparrevohn:2003:BSP

[1265] C. R. Sparrevohn and J. G. Støttrup. Bottom substrate preference in wild and reared turbot Psetta maxima L. Journal of Fish Biology, 63 (s1):257, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:2003:CHW

[1266] K. Turner and D. Righton. Cod in hot water: observations of behaviour in relation to temperature. Journal of Fish Biology, 63(s1):257–258, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watt:2003:FPB

[1267] P. J. Watt and A. J. Shohet. Female preferences based on olfactory cues in the guppy (Poecilia reticulata). Journal of Fish Biology, 63(s1):258, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wearmouth:2003:SDD

[1268] V. J. Wearmouth, D. W. Sims, J. C. Partridge, and I. C. Cuthill. Sex differences in dogfish behaviour: alternative activity patterns and habitat choice. Journal of Fish Biology, 63(s1):258, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2003:IIP

[1269] D. Wright, L. B. Rimmer, V. L. Pritchard, R. K. Butlin, and J. Krause. Inter and intra-population variation in shoaling and boldness in the zebrafish (Danio rerio). Journal of Fish Biology, 63(s1):258–259, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2003:WDG

[1270] H. A. Wright, R. J. Wootton, and I. Barber. (S)he who dares, grows! A cross-population study of growth and behaviour in three-spined sticklebacks. Journal of Fish Biology, 63(s1):259, December 2003. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2004:EN

[1271] John Craig. Editorial note. Journal of Fish Biology, 64(1):1, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rustadbakken:2004:RMB

[1272] A. Rustadbakken, J. H. L’Abée-Lund, J. V. Arnekleiv, and M. Kraabøl. Reproductive migration of brown trout in a small Norwegian river studied by telemetry. Journal of Fish Biology, 64(1):2–15, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lower:2004:RSS

[1273] N. Lower, A. P. Scott, and A. Moore. Release of sex steroids into the water by roach. Journal of Fish Biology, 64(1):16–33, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sponaugle:2004:IVP

[1274] S. Sponaugle and D. R. Pinkard. Impact of variable pelagic environments on natural larval growth and recruitment of the reef fish Thalassoma bifasciatum. Journal of Fish Biology, 64(1):34–54, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Levi:2004:BRM

[1275] F. Levi and P. Francour. Behavioural response of Mullus surmuletus to habitat modification by the invasive macroalga Caulerpa taxifolia. Journal of Fish Biology, 64(1):55–64, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bracewell:2004:EHE

[1276] P. Bracewell, I. G. Cowx, and R. F. Uglow. Effects of handling and electrofishing on plasma glucose and whole blood lactate of Leuciscus cephalus. Journal of Fish Biology, 64(1):65–71, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amara:2004:NCM

[1277] R. Amara and R. Galois. Nutritional condition of metamorphosing sole: spatial and temporal analyses. Journal of Fish Biology, 64(1):72–88, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knutsen:2004:MFA

[1278] J. A. Knutsen, H. Knutsen, E. M. Olsen, and B. Jonsson. Marine feeding of anadromous Salmo trutta during winter. Journal of Fish Biology, 64 (1):89–99, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sumi:2004:PFC

[1279] T. Sumi, Y. Hama, H. Nakagawa, D. Maruyama, and M. Asakawa. Purification and further characterization of a glycoprotein from the skin mucus of Japanese eels. Journal of Fish Biology, 64(1):100–115, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gisbert:2004:EDF

[1280] E. Gisbert, D. B. Conklin, and R. H. Piedrahita. Effects of delayed first feeding on the nutritional condition and mortality of California halibut larvae. Journal of Fish Biology, 64(1):116–132, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Begg:2004:EMR

[1281] K. Begg and N. W. Pankhurst. Endocrine and metabolic responses to stress in a laboratory population of the tropical damselfish Acanthochromis polyacanthus. Journal of Fish Biology, 64(1):133–145, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golubtsov:2004:DAC

[1282] A. S. Golubtsov, K. A. Moots, and K. F. Dzerjinskii. Dentition in the African catfishes Andersonia (Amphiliidae) and Siluranodon (Schilbeidae) previously considered toothless. Journal of Fish Biology, 64(1):146– 158, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parmentier:2004:RVA

[1283] E. Parmentier, D. Lecchini, and P. Vandewalle. Remodelling of the vertebral axis during metamorphic shrinkage in the pearlfish. Journal of Fish Biology, 64(1):159–169, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva:2004:MPS

[1284] M. E. Oliva and M. T. Gonzalez. Metazoan parasites of Sebastes capensis from two localities in northern Chile as tools for stock identification. Journal of Fish Biology, 64(1):170–175, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miler:2004:DOM

[1285] O. Miler and P. Fischer. Distribution and onshore migration behaviour of burbot larvae in Lake Constance, Germany. Journal of Fish Biology, 64(1):176–185, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:2004:PIG

[1286] T. F. Turner, M. V. McPhee, P. Campbell, and K. O. Winemiller. Phylogeography and intraspecific genetic variation of prochilodontid fishes endemic to rivers of northern South America. Journal of Fish Biology, 64(1):186–201, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez-Ruzafa:2004:CSD

[1287] A. Pérez-Ruzafa, J. I. Quispe-Becerra, J. A. García-Charton, and C. Marcos. Composition, structure and distribution of the ichthyoplankton in a Mediterranean coastal lagoon. Journal of Fish Biology, 64(1): 202–218, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heinimaa:2004:CMM

[1288] S. Heinimaa and J. Erkinaro. Characteristics of mature male parr in the northernmost Atlantic salmon populations. Journal of Fish Biology, 64 (1):219–226, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vainikka:2004:GSD

[1289] A. Vainikka, E. I. Jokinen, R. Kortet, and J. Taskinen. Gender- and season-dependent relationships between testosterone, oestradiol and immune functions in wild roach. Journal of Fish Biology, 64(1):227–240, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2004:SAM

[1290] N. Jonsson and B. Jonsson. Size and age of maturity of Atlantic salmon correlate with the North Atlantic Oscillation Index (NAOI). Journal of Fish Biology, 64(1):241–247, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newman:2004:STH

[1291] S. P. Newman, S. H. Gruber, and R. D. Handy. The scarecrow toadfish: habitat, abundance and size at maturity at bimini, Bahamas. Journal of Fish Biology, 64(1):248–252, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alexander:2004:ECE

[1292] G. D. Alexander and C. E. Adams. Exposure to a common environment erodes inherited between-population trophic morphology differences in Arctic charr. Journal of Fish Biology, 64(1):253–257, January 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crook:2004:MEA

[1293] D. A. Crook. A method for externally attaching radio transmitters to minimize dermal irritation. Journal of Fish Biology, 64(1):258–261, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2004:SLE

[1294] M. Takahashi and Y. Watanabe. Staging larval and early juvenile Japanese anchovy based on the degree of guanine deposition. Journal of Fish Biology, 64(1):262–267, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiwary:2004:AAS

[1295] B. K. Tiwary and A. K. Ray. Alterations in air-sac and skeleton of triploid Heteropneustes fossilis. Journal of Fish Biology, 64(1):268–272, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedersen:2004:RAS

[1296] L. F. Pedersen and H. Malte. Repetitive acceleration swimming performance of brown trout in fresh water and after acute seawater exposure. Journal of Fish Biology, 64(1):273–278, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Croft:2004:ISP

[1297] D. P. Croft, B. J. Arrowsmith, M. Webster, and J. Krause. Intra-sexual preferences for familiar fish in male guppies. Journal of Fish Biology, 64 (1):279–283, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferguson:2004:BR

[1298] A. Ferguson. Book reviews. Journal of Fish Biology, 64(1):284–285, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2004:BR

[1299] Peter J. Miller. Book reviews. Journal of Fish Biology, 64(1):285–286, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:2004:BR

[1300] Gordon Mcgregor Reid. Book reviews. Journal of Fish Biology, 64(1): 286–288, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2004:BR

[1301] I-Shiung Chen. Book reviews. Journal of Fish Biology, 64(1):288–289, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:BR

[1302] Anonymous. Books received. Journal of Fish Biology, 64(1):290, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:R

[1303] Anonymous. Referees 2003. Journal of Fish Biology, 64(1):291–296, January 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pillans:2004:FCS

[1304] R. D. Pillans, C. E. Franklin, and I. R. Tibbetts. Food choice in Siganus fuscescens: influence of macrophyte nutrient content and availability. Journal of Fish Biology, 64(2):297–309, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodnick:2004:TTM

[1305] K. J. Rodnick, A. K. Gamperl, K. R. Lizars, M. T. Bennett, R. N. Rausch, and E. R. Keeley. Thermal tolerance and metabolic physiology among redband trout populations in south-eastern Oregon. Journal of Fish Biology, 64(2):310–335, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2004:SDC

[1306] B. B. Smith and K. F. Walker. Spawning dynamics of common carp in the River Murray, South Australia, shown by macroscopic and histological staging of gonads. Journal of Fish Biology, 64(2):336–354, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanAerle:2004:OGS

[1307] R. Van Aerle, T. J. Runnalls, and C. R. Tyler. Ontogeny of gonadal sex development relative to growth in fathead minnow. Journal of Fish Biology, 64(2):355–369, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sasal:2004:PMP

[1308] P. Sasal, Y. Desdevises, E. Durieux, P. Lenfant, and P. Romans. Parasites in marine protected areas: success and specificity of monogeneans. Journal of Fish Biology, 64(2):370–379, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Massou:2004:EPI

[1309] A. M. Massou, J. Panfili, P.-Y. Le Bail, R. Laë, O. Mikolasek, G. Fontenelle, and J. F. Baroiller. Evidence of perturbations induced by reproduction on somatic growth and microincrement deposition in Oreochromis niloticus otoliths. Journal of Fish Biology, 64(2):380–398, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ornsrud:2004:IEV

[1310] R. Ørnsrud, A. Wargelius, Ø. Sæle, K. Pittman, and R. Waagbø. Influence of egg vitamin a status and egg incubation temperature on subsequent development of the early vertebral column in Atlantic salmon fry. Journal of Fish Biology, 64(2):399–417, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kahilainen:2004:DSH

[1311] K. Kahilainen, T. Malinen, A. Tuomaala, and H. Lehtonen. Diel and seasonal habitat and food segregation of three sympatric Coregonus lavaretus forms in a subarctic lake. Journal of Fish Biology, 64(2):418–434, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefebvre:2004:ISN

[1312] F. Lefebvre, B. Mounaix, G. Poizat, and A. J. Crivelli. Impacts of the swimbladder nematode Anguillicola crassus on Anguilla anguilla: variations in liver and spleen masses. Journal of Fish Biology, 64(2):435–447, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skov:2004:HUY

[1313] C. Skov and A. Koed. Habitat use of 0+ year pike in experimental ponds in relation to cannibalism, zooplankton, water transparency and habitat complexity. Journal of Fish Biology, 64(2):448–459, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amaral:2004:EWS

[1314] V. Amaral and H. N. Cabral. Ecology of the whiskered sole in the Sado Estuary, Portugal. Journal of Fish Biology, 64(2):460–474, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Njiwa:2004:VSR

[1315] J. R. Kemadjou Njiwa, P. Müller, and R. Klein. Variations of sperm release in three batches of zebrafish. Journal of Fish Biology, 64(2): 475–482, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Penczak:2004:FAC

[1316] T. Penczak, W. Galicka, Ł. Głowacki, H. Koszaliński, A. Kruk, G. Ziȩba, J. Kostrzewa, and L. Marszał. Fish assemblage changes relative to environmental factors and time in the Warta River, Poland, and its oxbow lakes. Journal of Fish Biology, 64(2):483–501, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Utoh:2004:OMJ

[1317] T. Utoh, N. Mikawa, A. Okamura, Y. Yamada, S. Tanaka, N. Horie, A. Akazawa, and H. P. Oka. Ovarian morphology of the Japanese eel in Mikawa Bay. Journal of Fish Biology, 64(2):502–513, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holcik:2004:DMS

[1318] J. Holčík, A. Delić, M. Kučinić, V. Bukvić, and M. Vater. Distribution and morphology of the sea lamprey from the Balkan coast of the Adriatic Sea. Journal of Fish Biology, 64(2):514–527, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2004:FIS

[1319] J. C. Svendsen, A. Koed, and K. Aarestrup. Factors influencing the spawning migration of female anadromous brown trout. Journal of Fish Biology, 64(2):528–540, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnekleiv:2004:RWJ

[1320] J. V. Arnekleiv, H. A. Urke, T. Kristensen, J. H. Halleraker, and L. E. W. Flodmark. Recovery of wild, juvenile brown trout from stress of flow reduction, electrofishing, handling and transfer from river to an indoor simulated stream channel. Journal of Fish Biology, 64(2):541–552, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carmona:2004:CCP

[1321] R. Carmona, M. García-Gallego, A. Sanz, A. Domezaín, and M. V. Ostos-Garrido. Chloride cells and pavement cells in gill epithelia of Acipenser naccarii: ultrastructural modifications in seawater-acclimated specimens. Journal of Fish Biology, 64(2):553–566, February 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinni:2004:SBD

[1322] M. Vinni, J. Lappalainen, T. Malinen, and H. Peltonen. Seasonal bottlenecks in diet shifts and growth of smelt in a large eutrophic lake. Journal of Fish Biology, 64(2):567–579, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beresford:2004:EDJ

[1323] N. Beresford, S. Jobling, R. Williams, and J. P. Sumpter. Endocrine disruption in juvenile roach from English rivers: a preliminary study. Journal of Fish Biology, 64(2):580–586, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:ISC

[1324] Anonymous. International Symposium on the Conservation and Management of Arctic Charr Salvelinus alpinus. Journal of Fish Biology, 64 (2):587–588, February 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Small:2004:EDC

[1325] B. C. Small. Effect of dietary cortisol administration on growth and reproductive success of channel catfish. Journal of Fish Biology, 64(3): 589–596, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ribeiro:2004:NSC

[1326] A. C. Ribeiro, R. C. Benine, and C. A. Figueiredo. A new species of creagrutus Günther (Teleostei: Ostariophysi: Characiformes), from the upper Rio Paraná basin, central Brazil. Journal of Fish Biology, 64(3):

597–611, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnhagen:2004:GLB

[1327] C. Magnhagen and E. Heibo. Growth in length and in body depth in young-of-the-year perch with different predation risk. Journal of Fish Biology, 64(3):612–624, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedini:2004:LHS

[1328] V. Pedini, C. Dall’Aglio, F. Parillo, and P. Scocco. A lectin histochemical study of the oesophagus of shi drum. Journal of Fish Biology, 64(3):625– 631, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wolter:2004:DDP

[1329] C. Wolter and J. Freyhof. Diel distribution patterns of fishes in a temperate large lowland river. Journal of Fish Biology, 64(3):632–642, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kristensen:2004:APR

[1330] E. A. Kristensen and G. P. Closs. Anti-predator response of naïve and experienced common bully to chemical alarm cues. Journal of Fish Biology, 64(3):643–652, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sutela:2004:SSN

[1331] T. Sutela, P. Korhonen, and K. Nyberg. Stocking success of newly hatched pike evaluated by radioactive strontium (^{85}Sr) marking. Journal of Fish Biology, 64(3):653–664, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2004:HUE

[1332] D. Cote, S. Moulton, P. C. B. Frampton, D. A. Scruton, and R. S. McKinley. Habitat use and early winter movements by juvenile Atlantic cod in a coastal area of Newfoundland. Journal of Fish Biology, 64(3): 665–679, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mookerji:2004:FPB

[1333] N. Mookerji, Z. Weng, and A. Mazumder. Food partitioning between coexisting Atlantic salmon and brook trout in the Sainte-Marguerite River ecosystem, Quebec. Journal of Fish Biology, 64(3):680–694, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2004:PEH

[1334] R. M. Rideout, E. A. Trippel, and M. K. Litvak. Paternal effects on haddock early life history traits. Journal of Fish Biology, 64(3):695– 701, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haberli:2004:MTI

[1335] M. A. Häberli and P. B. Aeschlimann. Male traits influence odour-based mate choice in the three-spined stickleback. Journal of Fish Biology, 64 (3):702–710, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2004:NMP

[1336] R. A. Martin. Natural mortality of puffadder shysharks due to Cape fur seals and black-backed kelp gulls at Seal Island, South Africa. Journal of Fish Biology, 64(3):711–716, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poortenaar:2004:RBF

[1337] C. W. Poortenaar, C. M. C. Woods, P. J. James, F. M. Giambartolomei, and P. M. Lokman. Reproductive biology of female big-bellied seahorses. Journal of Fish Biology, 64(3):717–725, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sogard:2004:EAJ

[1338] S. M. Sogard and M. L. Spencer. Energy allocation in juvenile sablefish: effects of temperature, ration and body size. Journal of Fish Biology, 64(3):726–738, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2004:RBS

[1339] J. D. Armstrong, P. S. Fallon-Cousins, and P. J. Wright. The relationship between specific dynamic action and otolith growth in pike. Journal of Fish Biology, 64(3):739–749, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:2004:ESM

[1340] M. K. Young and D. A. Schmetterling. Electrofishing and salmonid movement: reciprocal effects in two small montane streams. Journal of Fish Biology, 64(3):750–761, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinder:2004:EOS

[1341] A. C. Pinder and R. E. Gozlan. Early ontogeny of sunbleak. Journal of Fish Biology, 64(3):762–775, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farmer:2004:IFS

[1342] N. A. Farmer, D. O. Ribble, and D. G. Miller iii. Influence of familiarity on shoaling behaviour in Texas and blacktail shiners. Journal of Fish Biology, 64(3):776–782, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Munoz:2004:ASM

[1343] R. Rodríguez-Muñoz, J. R. Waldman, C. Grunwald, N. K. Roy, and I. Wirgin. Absence of shared mitochondrial DNA haplotypes between sea lamprey from North American and Spanish rivers. Journal of Fish Biology, 64(3):783–787, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Atkinson:2004:HSW

[1344] C. J. L. Atkinson, M. Bergmann, and M. J. Kaiser. Habitat selection in whiting. Journal of Fish Biology, 64(3):788–793, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albuquerque:2004:UHS

[1345] C. Q. Albuquerque and J. H. Muelbert. The use of haematoxylin stain to improve contrast of growth increments in otoliths of larval whitemouth croaker. Journal of Fish Biology, 64(3):794–798, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sneddon:2004:DRH

[1346] L. U. Sneddon and J. Yerbury. Differences in response to hypoxia in the three-spined stickleback from lotic and lentic localities: dominance and an anaerobic metabolite. Journal of Fish Biology, 64(3):799–804, March 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burridge:2004:SSG

[1347] C. P. Burridge, A. C. Hurt, L. W. Farrington, P. C. Coutin, and C. M. Austin. Stepping stone gene flow in an estuarine-dwelling sparid from south-east Australia. Journal of Fish Biology, 64(4):805–819, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schofield:2004:ISC

[1348] P. J. Schofield. Influence of salinity, competition and food supply on the growth of Gobiosoma robustum and Microgobius gulosus from Florida Bay, U.S.A. Journal of Fish Biology, 64(4):820–832, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvalho:2004:BHB

[1349] P. S. M. Carvalho, D. B. Noltie, and D. E. Tillitt. Biochemical, histological and behavioural aspects of visual function during early development of rainbow trout. Journal of Fish Biology, 64(4):833–850, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haukenes:2004:CCR

[1350] A. H. Haukenes and B. A. Barton. Characterization of the cortisol response following an acute challenge with lipopolysaccharide in yellow perch and the influence of rearing density. Journal of Fish Biology, 64 (4):851–862, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petersen:2004:RBR

[1351] C. W. Petersen, K. A. Zarrella, C. A. Ruben, and C. Mazzoldi. Reproductive biology of the rosylip sculpin, an intertidal spawner. Journal of Fish Biology, 64(4):863–875, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stark:2004:CMG

[1352] J. W. Stark. A comparison of the maturation and growth of female flathead sole in the central Gulf of Alaska and south-eastern Bering Sea. Journal of Fish Biology, 64(4):876–889, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drewe:2004:IFO

[1353] K. E. Drewe, M. H. Horn, K. A. Dickson, and A. Gawlicka. Insectivore to frugivore: ontogenetic changes in gut morphology and digestive enzyme activity in the characid fish Brycon guatemalensis from Costa Rican rain forest streams. Journal of Fish Biology, 64(4):890–902, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sutton:2004:DFI

[1354] T. M. Sutton, M. J. Cyterski, J. J. Ney, and M. C. Duval. Determination of factors influencing stomach content retention by striped bass captured using gillnets. Journal of Fish Biology, 64(4):903–910, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hill:2004:CTI

[1355] A. Hill, V. Howard, and A. Cossins. Characterization of TCDD-induced craniofacial malformations and retardation of zebrafish growth. Journal of Fish Biology, 64(4):911–922, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedges:2004:EEP

[1356] K. J. Hedges, S. A. Ludsin, and B. J. Fryer. Effects of ethanol preservation on otolith microchemistry. Journal of Fish Biology, 64(4):923–937, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maes:2004:FES

[1357] J. Maes, A. W. H. Turnpenny, D. R. Lambert, J. R. Nedwell, A. Parmentier, and F. Ollevier. Field evaluation of a sound system to reduce estuarine fish intake rates at a power plant cooling water inlet. Journal of Fish Biology, 64(4):938–946, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Susnik:2004:PPS

[1358] S. Sušnik, J. Schöffmann, and A. Snoj. Phylogenetic position of Salmo (Platysalmo) platycephalus Behnke 1968 from south-central Turkey, evidenced by genetic data. Journal of Fish Biology, 64(4):947–960, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreira:2004:CSN

[1359] P. S. A. Moreira and G. L. Volpato. Conditioning of stress in Nile tilapia. Journal of Fish Biology, 64(4):961–969, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dumay:2004:FGL

[1360] O. Dumay, P. S. Tari, J. A. Tomasini, and D. Mouillot. Functional groups of lagoon fish species in languedoc roussillon, southern France. Journal of Fish Biology, 64(4):970–983, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peck:2004:IID

[1361] M. A. Peck, L. J. Buckley, and D. A. Bengtson. Inter-individual differences in rates of routine energy loss and growth in young-of-the-year juvenile Atlantic cod. Journal of Fish Biology, 64(4):984–995, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schreer:2004:EIC

[1362] J. F. Schreer, S. J. Cooke, and K. B. Connors. Electrofishing-induced cardiac disturbance and injury in rainbow trout. Journal of Fish Biology, 64(4):996–1014, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyvarinen:2004:EAC

[1363] P. Hyvärinen, S. Heinimaa, and H. Rita. Effects of abrupt cold shock on stress responses and recovery in brown trout exhausted by swimming. Journal of Fish Biology, 64(4):1015–1026, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2004:ASS

[1364] G. N. Wagner and R. S. Mckinley. Anaemia and salmonid swimming performance: the potential effects of sub-lethal sea lice infection. Journal of Fish Biology, 64(4):1027–1038, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Patterson:2004:EEC

[1365] D. A. Patterson, J. S. Macdonald, S. G. Hinch, M. C. Healey, and A. P. Farrell. The effect of exercise and captivity on energy partitioning, reproductive maturation and fertilization success in adult sockeye salmon. Journal of Fish Biology, 64(4):1039–1059, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piner:2004:PHC

[1366] K. R. Piner and S. G. Wischniowski. Pacific halibut chronology of bomb radiocarbon in otoliths from 1944 to 1981 and a validation of ageing methods. Journal of Fish Biology, 64(4):1060–1071, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:2004:RNA

[1367] C. J. Robinson. Responses of the northern anchovy to the dynamics of the pelagic environment: identification of fish behaviours that may leave the population under risk of overexploitation. Journal of Fish Biology, 64(4):1072–1087, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wuenschel:2004:EBS

[1368] M. J. Wuenschel, R. G. Werner, and D. E. Hoss. Effect of body size, temperature, and salinity on the routine metabolism of larval and juvenile spotted seatrout. Journal of Fish Biology, 64(4):1088–1102, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ossum:2004:CNF

[1369] C. G. Ossum, E. K. Hoffmann, M. M. Vijayan, S. E. Holt, and N. C. Bols. Characterization of a novel fibroblast-like cell line from rainbow trout and responses to sublethal anoxia. Journal of Fish Biology, 64 (4):1103–1116, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OReilly:2004:PVA

[1370] K. M. O’Reilly and M. H. Horn. Phenotypic variation among populations of Atherinops affinis (Atherinopsidae) with insights from a geometric morphometric analysis. Journal of Fish Biology, 64(4):1117–1135, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoskonen:2004:TEA

[1371] P. Hoskonen and J. Pirhonen. Temperature effects on anaesthesia with clove oil in six temperate-zone fishes. Journal of Fish Biology, 64(4): 1136–1142, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2004:MUA

[1372] M. Stevens, G. Rappé, J. Maes, B. Van Asten, and F. Ollevier. Micropogonias undulatus (L.), another exotic arrival in European waters. Journal of Fish Biology, 64(4):1143–1146, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quillet:2004:ASD

[1373] E. Quillet, L. Labbe, and I. Queau. Asymmetry in sexual development of gonads in intersex rainbow trout. Journal of Fish Biology, 64(4):1147– 1151, April 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swartz:2004:CGP

[1374] E. R. Swartz, A. F. Flemming, and P. le F. N. Mouton. Contrasting genetic patterns and population histories in three threatened redfin species (Cyprinidae) from the Olifants River system, western South Africa. Journal of Fish Biology, 64(5):1153–1167, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2004:SID

[1375] K. A. Glover, J. B. Taggart, Ø. Skaala, and A. J. Teale. A study of inadvertent domestication selection during start-feeding of brown trout families. Journal of Fish Biology, 64(5):1168–1178, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karppinen:2004:RMO

[1376] P. Karppinen, J. Erkinaro, E. Niemelä, K. Moen, and F. Økland. Return migration of one-sea-winter Atlantic salmon in the River Tana. Journal of Fish Biology, 64(5):1179–1192, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magill:2004:ERT

[1377] S. H. Magill and M. D. J. Sayer. The effect of reduced temperature and salinity on the blood physiology of juvenile Atlantic cod. Journal of Fish Biology, 64(5):1193–1205, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henry:2004:SLB

[1378] T. B. Henry and J. M. Grizzle. Survival of largemouth bass, bluegill and channel catfish embryos after electroshocking. Journal of Fish Biology, 64(5):1206–1216, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suter:2004:NDH

[1379] H. C. Suter, M. Kennedy, and G. Young. Neoteny or not? The distribution of the hypophysial duct within the galaxioidea. Journal of Fish Biology, 64(5):1217–1225, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2004:ESP

[1380] S. M. Moran, G. P. Jenkins, M. J. Keough, and J. S. Hindell. Evidence for secondary planktonic transport of post-larvae of seagrass-associated King George whiting. Journal of Fish Biology, 64(5):1226–1241, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wiegand:2004:ULI

[1381] M. D. Wiegand, D. L. W. Young, B. M. Gajda, D. J. M. Thuen, D. A. H. Rittberg, J. D. Huebner, and N. L. Loadman. Ultraviolet light-induced impairment of goldfish embryo development and evidence for photorepair mechanisms. Journal of Fish Biology, 64(5):1242–1256, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dempson:2004:CPB

[1382] J. B. Dempson, C. J. Schwarz, M. Shears, and G. Furey. Comparative proximate body composition of Atlantic salmon with emphasis on parr from fluvial and lacustrine habitats. Journal of Fish Biology, 64(5):1257– 1271, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egger:2004:VPI

[1383] B. Egger, M. Meekan, W. Salzburger, L. Mwape, L. Makasa, R. Shapola, and C. Sturmbauer. Validation of the periodicity of increment formation in the otoliths of a cichlid fish from Lake Tanganyika, East Africa. Journal of Fish Biology, 64(5):1272–1284, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bang:2004:IVR

[1384] A. Bang, P. Grønkjær, and H. Malte. Individual variation in the rate of oxygen consumption by zebrafish embryos. Journal of Fish Biology, 64(5):1285–1296, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:2004:VSR

[1385] C. T. Walsh, B. C. Pease, and D. J. Booth. Variation in the sex ratio, size and age of longfinned eels within and among coastal catchments of south-eastern Australia. Journal of Fish Biology, 64(5):1297–1312, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sangiao-Alvarellos:2004:ECA

[1386] S. Sangiao-Alvarellos, M. Lapido, J. M. Miguez, and J. L. Soengas. Effects of central administration of arginine vasotocin on monoaminergic neurotransmitters and energy metabolism of rainbow trout brain. Journal of Fish Biology, 64(5):1313–1329, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barannikova:2004:SLT

[1387] I. A. Barannikova, L. V. Bayunova, and T. B. Semenkova. Serum levels of testosterone, 11-ketotestosterone and oestradiol-17β in three species of sturgeon during gonadal development and final maturation induced by hormonal treatment. Journal of Fish Biology, 64(5):1330–1338, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leiser:2004:DSM

[1388] J. K. Leiser, J. L. Gagliardi, and M. Itzkowitz. Does size matter? Assessment and fighting in small and large size-matched pairs of adult male convict cichlids. Journal of Fish Biology, 64(5):1339–1350, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temming:2004:PWH

[1389] A. Temming, S. Götz, N. Mergardt, and S. Ehrich. Predation of whiting and haddock on sandeel: aggregative response, competition and diel periodicity. Journal of Fish Biology, 64(5):1351–1372, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Terleph:2004:FAD

[1390] T. A. Terleph. The function of agonistic display behaviours in Gnathonemus petersii. Journal of Fish Biology, 64(5):1373–1385, May 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nykanen:2004:CMR

[1391] M. Nykänen, A. Huusko, and M. Lahti. Changes in movement, range and habitat preferences of adult grayling from late summer to early winter. Journal of Fish Biology, 64(5):1386–1398, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2004:POG

[1392] B. Sun and N. W. Pankhurst. Patterns of oocyte growth, vitellogenin and gonadal steroid concentrations in greenback flounder. Journal of Fish Biology, 64(5):1399–1412, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ryer:2004:CGB

[1393] C. Ryer. Comparison of genetically based antipredator behaviour in hatchery and wild Coho fry from a coastal Oregon river: interpreting a negative result. Journal of Fish Biology, 64(5):1413–1418, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopes:2004:MRS

[1394] D. C. J. R. Lopes, N. Bazzoli, M. F. G. Brito, and T. A. Maria. Male reproductive system in the South American catfish Conorhynchus conirostris. Journal of Fish Biology, 64(5):1419–1424, May 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svensson:2004:CCD

[1395] O. Svensson, A. Nyman, and C. Kvarnemo. Costly courtship or dishonest display? Intensely displaying sand goby males have lower lipid content. Journal of Fish Biology, 64(5):1425–1429, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zidowitz:2004:FRI

[1396] H. Zidowitz, H. O. Fock, C. Pusch, and H. Von Westernhagen. A first record of Isistius plutodus in the north-eastern Atlantic. Journal of Fish Biology, 64(5):1430–1434, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2004:FRD

[1397] P. M. Costa. First record of Diaphus dumerilii (Myctophiformes: Myctophidae) off the Portuguese mainland coast. Journal of Fish Biology, 64(5):1435–1438, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barreiros:2004:FRP

[1398] J. P. Barreiros, L. Machado, M. Hostim-Silva, I. Sazima, and P. C. Heemstra. First record of Polyprion oxygeneios (Perciformes: Polyprionidae) for the south-west Atlantic and a northernmost range extension. Journal of Fish Biology, 64(5):1439–1441, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nicod:2004:LLM

[1399] J.-C. Nicod, Y. Z. Wang, L. Excoffier, and C. R. Largiadèr. Low levels of mitochondrial DNA variation among central and southern European Esox lucius populations. Journal of Fish Biology, 64(5):1442–1449, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okamura:2004:AHB

[1400] A. Okamura, H. Zhang, T. Utoh, A. Akazawa, Y. Yamada, N. Horie, N. Mikawa, S. Tanaka, and H. P. Oka. Artificial hybrid between Anguilla anguilla and A. japonica. Journal of Fish Biology, 64(5):1450–1454, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Grana:2004:DTF

[1401] L. Rodríguez-Graña, G. Herrera, L. Herrera, and L. R. Castro. Divergence of two forms of Triphoturus in the eastern Pacific based on mtDNA cytochrome b gene sequences and larval morphology. Journal of Fish Biology, 64(5):1455–1461, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claridge:2004:BRF

[1402] Peter N. Claridge. Book review: Fishes in Estuaries. Edited by M. Elliott and K. L. Hemingway. xx 636 pp. Published by Blackwell Science, Oxford, 2002. Price £89.50. ISBN 0-632-05733-5. Journal of Fish Biology, 64(5):1462–1463, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alexander:2004:BRC

[1403] R. Mcneill Alexander. Book review: Comparative Biomechanics: Life’s Physical World. Edited by S. Vogel. xii 582 pp. Published by Princeton University Press, 2003. Price £39.95. ISBN 0-691-11297-5. Journal of Fish Biology, 64(5):1463–1464, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Egginton:2004:BRF

[1404] Stuart Egginton. Book review: Fish antifreeze proteins. Edited by K. V. Ewart and C. L. Hew. ix 236 pp. Published by World Scientific Publishing Co. (Molecular Aspects of Fish and Marine Biology, Vol. 1), Singapore, 2002. Price £46.00. ISBN 981-02-4899-7. Journal of Fish Biology, 64 (5):1464–1465, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2004:BRL

[1405] Peter J. Miller. Book reviews: L’Histoire entière des poissons. By Guillaume Rondelet (preface by F. Meunier and Jean-Loup D’Hondt). 418 + 181 pp. Published by CTHS (Collection CTHS Sciences 2), 2002. Price 40 euros. ISBN 2-7355-0501-4 and FISHES OF CHESAPEAKE BAY. Edited by Samuel F. Hildebrand and William C. Schroeder. 388 pp. Published by The Blackburn Press, Caldwell, New Jersey, 2002. Price $69.95. ISBN 1-930665-74-1. Journal of Fish Biology, 64(5):1465–1466, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:BRO

[1406] Anonymous. Book review: Oceanography and Marine Biology/An Annual Review. Vol. 41. Edited by R. N. Gibson and R. J. A. Atkinson. vii 435 pp. Published by Taylor and Francis, 2003. Price £107.00. ISBN 0-415-25463-9. Journal of Fish Biology, 64(5):1467, May 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buckel:2004:EDO

[1407] J. A. Buckel, B. L. Sharack, and V. S. Zdanowicz. Effect of diet on otolith composition in Pomatomus saltatrix, an estuarine piscivore. Journal of Fish Biology, 64(6):1469–1484, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kirankumar:2004:UHS

[1408] S. Kirankumar and T. J. Pandian. Use of heterologous sperm for the dispermic induction of androgenesis in barbs. Journal of Fish Biology, 64(6):1485–1497, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Candi:2004:RCS

[1409] G. Candi, L. Castriota, F. Andaloro, M. G. Finoia, and G. Marino. Reproductive cycle and sex inversion in razor fish, a protogynous labrid in the southern Mediterranean Sea. Journal of Fish Biology, 64(6):1498– 1513, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morley:2004:BBG

[1410] S. A. Morley, T. Mulvey, J. Dickson, and M. Belchier. The biology of the bigeye grenadier at South Georgia. Journal of Fish Biology, 64(6):1514– 1529, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayakawa:2004:UOE

[1411] Y. Hayakawa and H. Munehara. Ultrastructural observations of euspermatozoa and paraspermatozoa in a copulatory cottoid fish Blepsias cirrhosus. Journal of Fish Biology, 64(6):1530–1539, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Driggers:2004:RBC

[1412] W. B. Driggers III, D. A. Oakley, G. Ulrich, J. K. Carlson, B. J. Cullum, and J. M. Dean. Reproductive biology of Carcharhinus acronotus in the coastal waters of South Carolina. Journal of Fish Biology, 64(6):1540– 1551, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cavallaro:2004:SUC

[1413] M. Cavallaro, C. L. Mammola, and R. Verdiglione. Structural and ultrastructural comparison of photophores of two species of deep-sea fishes:

Argyropelecus hemigymnus and Maurolicus muelleri. Journal of Fish Biology, 64(6):1552–1567, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bagamian:2004:BCR

[1414] K. H. Bagamian, D. C. Heins, and J. A. Baker. Body condition and reproductive capacity of three-spined stickleback infected with the cestode Schistocephalus solidus. Journal of Fish Biology, 64(6):1568–1576, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bartl:2004:GMI

[1415] E. Bartl and H. Keckeis. Growth and mortality of introduced fish larvae in a newly restored urban river. Journal of Fish Biology, 64(6):1577– 1592, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wagner:2004:STF

[1416] G. N. Wagner, R. S. Mckinley, P. A. Bjørn, and B. Finstad. Short-term freshwater exposure benefits sea lice-infected Atlantic salmon. Journal of Fish Biology, 64(6):1593–1604, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catalan:2004:SDM

[1417] I. A. Catalán, I. A. Johnston, and M. P. Olivar. Seasonal differences in muscle fibre recruitment of pilchard larvae in the north-western Mediterranean. Journal of Fish Biology, 64(6):1605–1615, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolliet:2004:RPG

[1418] V. Bolliet, M. Jarry, and T. Boujard. Rhythmic pattern of growth and nutrient retention in response to feeding time in the rainbow trout. Journal of Fish Biology, 64(6):1616–1624, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gronkjaer:2004:OBA

[1419] P. Grønkjær, C. Skov, and S. Berg. Otolith-based analysis of survival and size-selective mortality of stocked 0+ year pike related to time of stocking. Journal of Fish Biology, 64(6):1625–1637, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nichols:2004:ITS

[1420] J. W. Nichols and R. C. Playle. Influence of temperature on silver accumulation and depuration in rainbow trout. Journal of Fish Biology, 64(6):1638–1654, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2004:GDM

[1421] A. H. Rikardsen, M. Haugland, P. A. Bjørn, B. Finstad, R. Knudsen, J. B. Dempson, J. C. Holst, N. A. Hvidsten, and M. Holm. Geographical differences in marine feeding of Atlantic salmon post-smolts in Norwegian fjords. Journal of Fish Biology, 64(6):1655–1679, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Floeter:2004:LGA

[1422] S. R. Floeter, C. E. L. Ferreira, A. Dominici-Arosemena, and I. R. Zalmon. Latitudinal gradients in Atlantic reef fish communities: trophic structure and spatial use patterns. Journal of Fish Biology, 64(6):1680– 1699, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casselman:2004:STR

[1423] S. J. Casselman and R. Montgomerie. Sperm traits in relation to male quality in colonial spawning bluegill. Journal of Fish Biology, 64(6):1700– 1711, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoner:2004:FDS

[1424] A. W. Stoner and M. L. Ottmar. Fish density and size alter Pacific halibut feeding: implications for stock assessment. Journal of Fish Biology, 64(6):1712–1724, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2004:LHV

[1425] E. M. Olsen, H. Knutsen, J. Gjøsæter, P. E. Jorde, J. A. Knutsen, and N. C. Stenseth. Life-history variation among local populations of Atlantic cod from the Norwegian Skagerrak coast. Journal of Fish Biology, 64(6): 1725–1730, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imre:2004:AED

[1426] I. Imre and D. Boisclair. Age effects on diel activity patterns of juvenile Atlantic salmon: parr are more nocturnal than young-of-the-year. Journal of Fish Biology, 64(6):1731–1736, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knaepkens:2004:UGT

[1427] G. Knaepkens, E. Verheyen, P. Galbusera, and M. Eens. The use of genetic tools for the evaluation of a potential migration barrier for the bullhead. Journal of Fish Biology, 64(6):1737–1744, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:IPF

[1428] Anonymous. 7^{th} Indo–Pacific Fish Conference, May 2005, Taiwan. Journal of Fish Biology, 64(6):1745–1746, June 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Foster:2004:LHE

[1429] S. J. Foster and A. C. J. Vincent. Life history and ecology of seahorses: implications for conservation and management. Journal of Fish Biology, 65(1):1–61, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crivello:2004:GSS

[1430] J. F. Crivello, D. J. Danila, E. Lorda, M. Keser, and E. F. Roseman. The genetic stock structure of larval and juvenile winter flounder larvae in Connecticut waters of eastern Long Island Sound and estimations of larval entrainment. Journal of Fish Biology, 65(1):62–76, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Townsend:2004:ESC

[1431] K. A. Townsend and I. R. Tibbetts. The ecological significance of the combtoothed blenny in a coral reef ecosystem. Journal of Fish Biology, 65(1):77–90, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stiver:2004:DPS

[1432] K. A. Stiver, P. Dierkes, M. Taborsky, and S. Balshine. Dispersal patterns and status change in a co-operatively breeding cichlid Neolamprologus pulcher: evidence from microsatellite analyses and behavioural observations. Journal of Fish Biology, 65(1):91–105, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsson:2004:PML

[1433] I. C. Olsson and L. A. Greenberg. Partial migration in a landlocked brown trout population. Journal of Fish Biology, 65(1):106–121, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kallio-nyberg:2004:ABE

[1434] I. Kallio-nyberg, E. Jutila, I. Saloniemi, and E. Jokikokko. Association between environmental factors, smolt size and the survival of wild and reared Atlantic salmon from the Simojoki River in the Baltic Sea. Journal of Fish Biology, 65(1):122–134, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2004:DRB

[1435] L. Jacobsen, S. Berg, N. Jepsen, and C. Skov. Does roach behaviour differ between shallow lakes of different environmental state? Journal of Fish Biology, 65(1):135–147, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maeva:2004:UPE

[1436] E. Maeva, I. Bruno, B. S. Zielinski, M. F. Docker, F. M. Severin, and R. G. Maev. The use of pulse-echo acoustic microscopy to non-invasively determine the sex of living larval sea lampreys. Journal of Fish Biology, 65(1):148–156, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hurst:2004:TSD

[1437] T. P. Hurst. Temperature and state-dependence of feeding and gastric evacuation in juvenile Pacific halibut. Journal of Fish Biology, 65(1): 157–169, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2004:ISG

[1438] J. H. Lee, S. K. Ku, K. D. Park, and H. S. Lee. Immunohistochemical study of the gastrointestinal endocrine cells in the Korean aucha perch. Journal of Fish Biology, 65(1):170–181, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2004:SPD

[1439] M. C. P. Amorim, Y. Stratoudakis, and A. D. Hawkins. Sound production during competitive feeding in the grey gurnard. Journal of Fish Biology, 65(1):182–194, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richmond:2004:LIP

[1440] H. E. Richmond, T. R. Hrabik, and A. F. Mensinger. Light intensity, prey detection and foraging mechanisms of age 0 year yellow perch. Journal of Fish Biology, 65(1):195–205, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raine:2004:EEO

[1441] J. C. Raine, C. Cameron, M. M. Vijayan, J. Lamarre, and J. F. Leatherland. The effect of elevated oocyte triiodothyronine content on development of rainbow trout embryos and expression of mRNA encoding for thyroid hormone receptors. Journal of Fish Biology, 65(1):206–226, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2004:VVV

[1442] E. M. White, D. M. GonÇalves, J. C. Partridge, and R. F. Oliveira. Vision and visual variation in the peacock blenny. Journal of Fish Biology, 65(1):227–250, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Engstrom-Ost:2004:TSP

[1443] J. Engström-Öst and M. Lehtiniemi. Threat-sensitive predator avoidance by pike larvae. Journal of Fish Biology, 65(1):251–261, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rindorf:2004:EER

[1444] A. Rindorf. Estimation of evacuation rates in the field. Journal of Fish Biology, 65(1):262–281, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shoji:2004:FPJ

[1445] J. Shoji and M. Tanaka. Feeding of piscivorous Japanese Spanish mackerel juveniles on larvae with an elongate and laterally compressed body. Journal of Fish Biology, 65(1):282–286, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baldisserotto:2004:ILG

[1446] B. Baldisserotto, K. López-Vásquez, L. V. F. Silva, J. I. Golombieski, and A. L. Val. Ionic levels of the gallbladder bile of some teleosts from the Rio Negro, Amazon. Journal of Fish Biology, 65(1):287–292, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnson:2004:TUC

[1447] J. B. Johnson and M. C. Belk. Temperate Utah chub form valid otolith annuli in the absence of fluctuating water temperature. Journal of Fish Biology, 65(1):293–298, July 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2004:DHA

[1448] R. M. Rideout, E. A. Trippel, and M. K. Litvak. The development of haddock and Atlantic cod sperm cryopreservation techniques and the effect of sperm age on cryopreservation success. Journal of Fish Biology, 65(2):299–311, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanDenThillart:2004:ESE

[1449] G. Van Den Thillart, V. Van Ginneken, F. Körner, R. Heijmans, R. Van Der Linden, and A. Gluvers. Endurance swimming of European eel. Journal of Fish Biology, 65(2):312–318, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2004:RBS

[1450] R. M. Rideout, E. A. Trippel, and M. K. Litvak. Relationship between sperm density, spermatocrit, sperm motility and spawning date in wild and cultured haddock. Journal of Fish Biology, 65(2):319–332, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wong:2004:MCD

[1451] B. B. M. Wong. Male competition is disruptive to courtship in the Pacific blue-eye. Journal of Fish Biology, 65(2):333–341, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacNutt:2004:ETA

[1452] M. J. MacNutt, S. G. Hinch, A. P. Farrell, and S. Topp. The effect of temperature and acclimation period on repeat swimming performance in cutthroat trout. Journal of Fish Biology, 65(2):342–353, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berra:2004:ODA

[1453] T. M. Berra and D. D. Aday. Otolith description and age-and-growth of Kurtus gulliveri from northern Australia. Journal of Fish Biology, 65 (2):354–362, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turesson:2004:FBC

[1454] H. Turesson and C. Brönmark. Foraging behaviour and capture success in perch, pikeperch and pike and the effects of prey density. Journal of Fish Biology, 65(2):363–375, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stuart-Smith:2004:ITS

[1455] R. D. Stuart-Smith, A. M. M. Richardson, and R. W. G. White. Increasing turbidity significantly alters the diet of brown trout: a multi-year longitudinal study. Journal of Fish Biology, 65(2):376–388, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beacham:2004:PSL

[1456] T. D. Beacham, B. McIntosh, and C. MacConnachie. Population structure of lake-type and river-type sockeye salmon in transboundary rivers of northern British Columbia. Journal of Fish Biology, 65(2):389–402, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caputo:2004:GDB

[1457] V. Caputo, M. Giovannotti, P. Nisi Cerioni, M. L. Caniglia, and A. Splendiani. Genetic diversity of brown trout in central Italy. Journal of Fish Biology, 65(2):403–418, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2004:MGD

[1458] T. Takahashi. Morphological and genetic distinctness of rock and shellbed dwelling Telmatochromis (Teleostei, Cichlidae) in the south of Lake Tanganyika suggest the existence of two species. Journal of Fish Biology, 65(2):419–435, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2004:CSA

[1459] S. W. Griffiths, P. Collen, and J. D. Armstrong. Competition for shelter among over-wintering signal crayfish and juvenile Atlantic salmon. Journal of Fish Biology, 65(2):436–447, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutchinson:2004:RBT

[1460] S. Hutchinson and L. E. Hawkins. The relationship between temperature and the size and age of larvae and peri-metamorphic stages of Pleuronectes flesus. Journal of Fish Biology, 65(2):448–459, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flodmark:2004:PJB

[1461] L. E. W. Flodmark, L. A. Vøllestad, and T. Forseth. Performance of juvenile brown trout exposed to fluctuating water level and temperature.

Journal of Fish Biology, 65(2):460–470, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Narum:2004:GDS

[1462] S. R. Narum, C. Contor, A. Talbot, and M. S. Powell. Genetic divergence of sympatric resident and anadromous forms of Oncorhynchus mykiss in the Walla Walla River, U.S.A. Journal of Fish Biology, 65(2):471– 488, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ekanem:2004:ELS

[1463] A. P. Ekanem, T. Meinelt, W. Kloas, and K. Knopf. Early life stage toxicity of extracts from the African fish poison plants Tephrosia vogelii Hook. f. and Asystasia vogeliana Benth. on zebrafish embryos. Journal of Fish Biology, 65(2):489–494, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mattiucci:2004:GIA

[1464] S. Mattiucci, P. Abaunza, L. Ramadori, and G. Nascetti. Genetic identification of Anisakis larvae in European hake from Atlantic and Mediterranean waters for stock recognition. Journal of Fish Biology, 65(2):495– 510, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamanoue:2004:RGM

[1465] Y. Yamanoue and K. Matsuura. A review of the genus Malakichthys Döderlein (Perciformes: Acropomatidae) with the description of a new species. Journal of Fish Biology, 65(2):511–529, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simpkins:2004:CBS

[1466] D. G. Simpkins, W. A. Hubert, C. Martinez Del Rio, and D. C. Rule. Constraints of body size and swimming velocity on the ability of juvenile rainbow trout to endure periods without food. Journal of Fish Biology, 65(2):530–544, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovacic:2004:UME

[1467] M. Kovačić. Unusual morphological and ecological characteristics of hyperbenthic juveniles of Gobius cruentatus. Journal of Fish Biology, 65 (2):545–558, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mabragana:2004:RBT

[1468] E. Mabragaña and M. B. Cousseau. Reproductive biology of two sympatric skates in the south-west Atlantic: Psammobatis rudis and Psammobatis normani. Journal of Fish Biology, 65(2):559–573, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Telnes:2004:RST

[1469] T. Telnes and H. Sægrov. Reproductive strategies in two sympatric morphotypes of Arctic charr in Kalandsvatnet, west Norway. Journal of Fish Biology, 65(2):574–579, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suzuki:2004:PRW

[1470] N. Suzuki, M. Nishida, K. Yoseda, C. Üstündağ, T. Şahïn, and K. Amaoka. Phylogeographic relationships within the Mediterranean turbot inferred by mitochondrial DNA haplotype variation. Journal of Fish Biology, 65(2):580–585, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jarman:2004:DBS

[1471] S. N. Jarman and S. G. Wilson. DNA-based species identification of krill consumed by whale sharks. Journal of Fish Biology, 65(2):586– 591, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seoka:2004:FAA

[1472] M. Seoka, S. Yamada, and H. Kumai. Free amino acids in Japanese eel eggs obtained by hormonal inducement. Journal of Fish Biology, 65 (2):592–596, August 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manera:2004:RCT

[1473] M. Manera and B. S. Dezfuli. Rodlet cells in teleosts: a new insight into their nature and functions. Journal of Fish Biology, 65(3):597–619, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burrows:2004:ADS

[1474] M. T. Burrows, R. N. Gibson, L. Robb, and A. Maclean. Alongshore dispersal and site fidelity of juvenile plaice from tagging and transplants. Journal of Fish Biology, 65(3):620–634, September 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Timmerman:2004:HIV

[1475] C. M. Timmerman and L. J. Chapman. Hypoxia and interdemic variation in Poecilia latipinna. Journal of Fish Biology, 65(3):635–650, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fishelson:2004:TBL

[1476] L. Fishelson and Y. Delarea. Taste buds on the lips and mouth of some blenniid and gobiid fishes: comparative distribution and morphology. Journal of Fish Biology, 65(3):651–665, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyegaard:2004:VDG

[1477] M. Nyegaard, A. Arkhipkin, and P. Brickle. Variation in the diet of Genypterus blacodes (Ophidiidae) around the Falkland Islands. Journal of Fish Biology, 65(3):666–682, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2004:MMF

[1478] T. D. Harrison and A. K. Whitfield. A multi-metric fish index to assess the environmental condition of estuaries. Journal of Fish Biology, 65(3): 683–710, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2004:SOS

[1479] A. H. Rikardsen. Seasonal occurrence of sea lice Lepeophtheirus salmonis on sea trout in two north Norwegian fjords. Journal of Fish Biology, 65 (3):711–722, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wuenschel:2004:CGE

[1480] M. J. Wuenschel and R. G. Werner. Consumption and gut evacuation rate of laboratory-reared spotted seatrout (Sciaenidae) larvae and juveniles. Journal of Fish Biology, 65(3):723–743, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:2004:IGT

[1481] L. M. Clarke and K. D. Friedland. Influence of growth and temperature on strontium deposition in the otoliths of Atlantic salmon. Journal of Fish Biology, 65(3):744–759, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mary:2004:EEE

[1482] C. M. St Mary, E. Gordon, and R. E. Hale. Environmental effects on egg development and hatching success in Jordanella floridae, a species with parental care. Journal of Fish Biology, 65(3):760–768, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lortie:2004:IFA

[1483] M. B. Lortie, T. Arnason, S. G. Dugan, J. G. Nickerson, and T. W. Moon. The impact of feeding β_{2}-adrenergic agonists on rainbow trout muscle β_{2}-adrenoceptors and protein synthesis. Journal of Fish Biology, 65(3): 769–787, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crossin:2004:EMS

[1484] G. T. Crossin, S. G. Hinch, A. P. Farrell, D. A. Higgs, A. G. Lotto, J. D. Oakes, and M. C. Healey. Energetics and morphology of sockeye salmon: effects of upriver migratory distance and elevation. Journal of Fish Biology, 65(3):788–810, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biswas:2004:PRN

[1485] A. K. Biswas, M. Maita, G. Yoshizaki, and T. Takeuchi. Physiological responses in Nile tilapia exposed to different photoperiod regimes. Journal of Fish Biology, 65(3):811–821, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibanez:2004:CVS

[1486] A. L. Ibáñez and O. Gutiérrez Benítez. Climate variables and spawning migrations of the striped mullet and white mullet in the north-western area of the Gulf of Mexico. Journal of Fish Biology, 65(3):822–831, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aceves-Medina:2004:FLA

[1487] G. Aceves-Medina, S. P. A. Jiménez-Rosenberg, A. Hinojosa-Medina, R. Funes-Rodríguez, R. J. Saldierna-Martínez, and P. E. Smith. Fish larvae assemblages in the Gulf of California. Journal of Fish Biology, 65 (3):832–847, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chan:2004:DEA

[1488] A. S. Chan, M. H. Horn, K. A. Dickson, and A. Gawlicka. Digestive enzyme activities in carnivores and herbivores: comparisons among four closely related prickleback fishes (Teleostei: Stichaeidae) from a California rocky intertidal habitat. Journal of Fish Biology, 65(3):848–858, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coronel:2004:DPH

[1489] J. S. Coronel, G. E. Maes, S. Claus, P. A. Van Damme, and F. A. M. Volckaert. Differential population history in the migratory catfishes Brachyplatystoma flavicans and Pseudoplatystoma fasciatum (Pimelodidae) from the Bolivian Amazon assessed with nuclear and mitochondrial DNA markers. Journal of Fish Biology, 65(3):859–868, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ylonen:2004:GSE

[1490] O. Ylönen and J. Karjalainen. Growth and survival of European whitefish larvae under enhanced UV-B irradiance. Journal of Fish Biology, 65(3): 869–875, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:2004:URT

[1491] R. Jordan, K. Kellogg, F. Juanes, D. Howe, J. Staufferals, Jr., E. Loew, and G. Losey. Ultraviolet reflectivity in three species of Lake Malaŵi rock-dwelling cichlids. Journal of Fish Biology, 65(3):876–882, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morison:2004:RMC

[1492] A. K. Morison. Is rigor mortis the cause of post-mortem shrinkage in juvenile Pagrus auratus? Journal of Fish Biology, 65(3):883–888, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2004:BRB

[1493] Paul J. B. Hart. Book review: Biology and Ecology of Fishes. By J. S. Diana. xv 498 pp. 2nd ed. Published by Biological Sciences Press, Traverse City, Michigan, 2004. Price $80.00. ISBN 1-884125-98-0. Journal of Fish Biology, 65(3):889–890, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Polunin:2004:BRC

[1494] Nicholas Polunin. Book review: Coral Reef Fishes/Dynamics and Diversity in a Complex Ecosystem. Edited by P. F. Sale. xiv 549 pp. Published by Academic Press, San Diego, 2002. Price £69.95. ISBN 0-12-615185-7. Journal of Fish Biology, 65(3):890–891, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parenti:2004:BRF

[1495] Lynne R. Parenti. Book review: The Freshwater Fishes of South and Central America. Organized by R. E. Reis, S. O. Kullander and C. J. Ferraris, Jr. 742 pp. Published by EDIPUCRS (Editora Pontificia Universidade Católica do Rio Grande do Sul), Porto Alegro, Brazil, 2003. Price R$ 146.00 or US$48.00. ISBN 85-7430-361-5. Journal of Fish Biology, 65(3):891–892, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:BRE

[1496] Anonymous. Book review: The Eel. By F.-W. Tesch (edited by J. E. Thorpe). viii 408 pp. Published by Blackwell Science, 2003. Price £89.50. ISBN 0-632-06389-0. Journal of Fish Biology, 65(3):893, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:BRM

[1497] Anonymous. Book reviews: Mugilidae, Atherinidae, Atherinopsidae, Blenniidae, Odontobutidae and Gobiidae 1. xii + 404 pp., 2003. Price EUR 109.00. ISBN 3-89104-668-5 and Gobiidae 2. x + 478 pp., 2004. Price EUR 129.00. ISBN 3-89104-669-3. Edited by P. J. Miller. Published by AULA-Verlag, Wiesbaden (The Freshwater Fishes of Europe, Vol. 8, parts 1 and 2). Journal of Fish Biology, 65(3):893–894, September 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ros:2004:ALE

[1498] A. F. H. Ros, K. Becker, A. V. M. Canário, and R. F. Oliveira. Androgen levels and energy metabolism in Oreochromis mossambicus. Journal of Fish Biology, 65(4):895–905, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tierney:2004:DLL

[1499] K. B. Tierney, A. P. Farrell, and C. J. Kennedy. The differential leucocyte landscape of four teleosts: juvenile Oncorhynchus kisutch, Clupea pallasi, Culaea inconstans and Pimephales promelas. Journal of Fish Biology, 65(4):906–919, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wakefield:2004:MRA

[1500] A. M. Wakefield, R. A. Cunjak, and J. D. Kieffer. Metabolic recovery in Atlantic salmon fry and parr following forced activity. Journal of Fish Biology, 65(4):920–932, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belanger:2004:BER

[1501] A. J. Belanger, W. J. Arbuckle, L. D. Corkum, D. B. Gammon, W. Li, A. P. Scott, and B. S. Zielinski. Behavioural and electrophysiological responses by reproductive female Neogobius melanostomus to odours released by conspecific males. Journal of Fish Biology, 65(4):933–946, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurenson:2004:DWO

[1502] C. H. Laurenson, I. R. Hudson, D. O. B. Jones, and I. G. Priede. Deep water observations of Lophius piscatorius in the north-eastern Atlantic Ocean by means of a remotely operated vehicle. Journal of Fish Biology, 65(4):947–960, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quintella:2004:BPS

[1503] B. R. Quintella, N. O. Andrade, A. Koed, and P. R. Almeida. Behavioural patterns of sea lampreys’ spawning migration through difficult passage areas, studied by electromyogram telemetry. Journal of Fish Biology, 65 (4):961–972, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gurshin:2004:STS

[1504] C. W. D. Gurshin and S. T. Szedlmayer. Short-term survival and movements of Atlantic sharpnose sharks captured by hook-and-line in the north-east Gulf of Mexico. Journal of Fish Biology, 65(4):973–986, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2004:EGD

[1505] M. Liu and Y. Sadovy. Early gonadal development and primary males in the protogynous epinepheline, Cephalopholis boenak. Journal of Fish Biology, 65(4):987–1002, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campinho:2004:TSS

[1506] M. A. Campinho, K. A. Moutou, and D. M. Power. Temperature sensitivity of skeletal ontogeny in Oreochromis mossambicus. Journal of Fish Biology, 65(4):1003–1025, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manning:2004:PRC

[1507] A. J. Manning, M. P. M. Burton, and L. W. Crim. Puberty and recrudescence in cultured male yellowtail flounder: plasma androgens and spermatogenesis. Journal of Fish Biology, 65(4):1026–1046, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Okland:2004:ASP

[1508] F. Økland, E. B. Thorstad, and T. F. Næsje. Is Atlantic salmon production limited by number of territories? Journal of Fish Biology, 65 (4):1047–1055, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinez:2004:IVH

[1509] M. L. Martinez, L. J. Chapman, J. M. Grady, and B. B. Rees. Interdemic variation in haematocrit and lactate dehydrogenase in the African cyprinid Barbus neumayeri. Journal of Fish Biology, 65(4):1056–1069, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robertson:2004:EIF

[1510] M. J. Robertson, C. J. Pennell, D. A. Scruton, G. J. Robertson, and J. A. Brown. Effect of increased flow on the behaviour of Atlantic salmon parr in winter. Journal of Fish Biology, 65(4):1070–1079, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buchner:2004:PES

[1511] A. S. Buchner, K. A. Sloman, and S. Balshine. The physiological effects of social status in the cooperatively breeding cichlid Neolamprologus pulcher. Journal of Fish Biology, 65(4):1080–1095, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carter:2004:CFD

[1512] C. G. Carter and S. J. Davies. Changes to feeding and dominance ranks following the introduction of novel feeds to African catfish. Journal of Fish Biology, 65(4):1096–1107, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Butler:2004:DEM

[1513] R. Butler and B. F. Nowak. A dual enzyme method for the establishment of long- and medium-term primary cultures of epithelial and fibroblastic cells from Atlantic salmon gills. Journal of Fish Biology, 65(4):1108– 1125, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keefer:2004:UMR

[1514] M. L. Keefer, C. A. Peery, M. A. Jepson, and L. C. Stuehrenberg. Upstream migration rates of radio-tagged adult Chinook salmon in riverine habitats of the Columbia River basin. Journal of Fish Biology, 65(4): 1126–1141, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lecchini:2004:EPS

[1515] D. Lecchini, V. Dufour, J. Carleton, S. Strand, and R. Galzin. Estimating the patch size of larval fishes during colonization on coral reefs. Journal of Fish Biology, 65(4):1142–1146, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keskinen:2004:DPS

[1516] T. Keskinen and T. J. Marjomäki. Diet and prey size spectrum of pikeperch in lakes in central Finland. Journal of Fish Biology, 65(4): 1147–1153, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2004:FMP

[1517] J. P. Evans, A. Bisazza, and A. Pilastro. Female mating preferences for colourful males in a population of guppies subject to high predation. Journal of Fish Biology, 65(4):1154–1159, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulin:2004:RTS

[1518] E. Poulin, L. Cárdenas, C. E. Hernández, I. Kornfield, and F. P. Ojeda. Resolution of the taxonomic status of Chilean and Californian jack mackerels using mitochondrial DNA sequence. Journal of Fish Biology, 65(4): 1160–1164, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinky:2004:OKS

[1519] Pinky, S. Mittal, M. Yashpal, J. Ojha, and A. K. Mittal. Occurrence of keratinization in the structures associated with lips of a hill stream fish Garra lamta (Hamilton) (Cyprinidae, Cypriniformes). Journal of Fish Biology, 65(4):1165–1172, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luiz-Junior:2004:OAM

[1520] O. J. Luiz-Júnior, S. R. Floeter, J. L. Gasparini, C. E. L. Ferreira, and P. Wirtz. The occurrence of Acanthurus monroviae (Perciformes: Acanthuridae) in the south-western Atlantic, with comments on other eastern Atlantic reef fishes occurring in Brazil. Journal of Fish Biology, 65(4): 1173–1179, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2004:FRF

[1521] J. M. N. Azevedo, P. M. Raposeiro, and L. Rodrigues. First records of Fistularia petimba and Diodon eydouxii for the Azores, with notes on the occurrence of three additional species. Journal of Fish Biology, 65 (4):1180–1184, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Westgaard:2004:GDB

[1522] J. I. Westgaard, A. Klemetsen, and R. Knudsen. Genetic differences between two sympatric morphs of Arctic charr confirmed by microsatellite DNA. Journal of Fish Biology, 65(4):1185–1191, October 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2004:FFD

[1523] R. W. Blake. Fish functional design and swimming performance. Journal of Fish Biology, 65(5):1193–1222, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cech:2004:CSH

[1524] J. J. Cech, Jr., M. McEnroe, and D. J. Randall. Coho salmon haematological, metabolic and acid-base changes during exercise and recovery in sea water. Journal of Fish Biology, 65(5):1223–1232, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2004:NIS

[1525] T. Ellis, J. D. James, C. Stewart, and A. P. Scott. A non-invasive stress assay based upon measurement of free cortisol released into the water by rainbow trout. Journal of Fish Biology, 65(5):1233–1252, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Massou:2004:ECS

[1526] A. M. Massou, P.-Y. Le Bail, J. Panfili, R. Laë, J. F. Baroiller, O. Mikolasek, G. Fontenelle, and B. Auperin. Effects of confinement stress of variable duration on the growth and microincrement deposition in the otoliths of Oreochromis niloticus (Cichlidae). Journal of Fish Biology, 65(5):1253–1269, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2004:GIR

[1527] B. A. Henderson, G. E. Morgan, and A. Vaillancourt. Growth, ingestion rates and metabolic activity of walleye in lakes with and without lake herring. Journal of Fish Biology, 65(5):1270–1282, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2004:EIO

[1528] F. Lahnsteiner, N. Mansour, and B. Berger. The effect of inorganic and organic pollutants on sperm motility of some freshwater teleosts. Journal of Fish Biology, 65(5):1283–1297, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

James:2004:WBM

[1529] J. D. James, T. Ellis, and A. P. Scott. Water-based measurement of rainbow trout melatonin. Journal of Fish Biology, 65(5):1298–1304, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:2004:DAP

[1530] P. Johnston, N. E. Bergeron, and J. J. Dodson. Diel activity patterns of juvenile Atlantic salmon in rivers with summer water temperature near the temperature-dependent suppression of diurnal activity. Journal of Fish Biology, 65(5):1305–1318, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zitek:2004:FDDa

[1531] A. Zitek, S. Schmutz, G. Unfer, and A. Ploner. Fish drift in a Danube sidearm-system: I. Site-, inter- and intraspecific patterns. Journal of Fish Biology, 65(5):1319–1338, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zitek:2004:FDDb

[1532] A. Zitek, S. Schmutz, and A. Ploner. Fish drift in a Danube sidearmsystem: II. Seasonal and diurnal patterns. Journal of Fish Biology, 65(5):

1339–1357, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2004:DSM

[1533] M. C. P. Amorim, M. E. Knight, Y. Stratoudakis, and G. F. Turner. Differences in sounds made by courting males of three closely related Lake Malaŵi cichlid species. Journal of Fish Biology, 65(5):1358–1371, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

BrindAmour:2004:CBT

[1534] A. Brind’Amour and D. Boisclair. Comparison between two sampling methods to evaluate the structure of fish communities in the littoral zone of a Laurentian lake. Journal of Fish Biology, 65(5):1372–1384, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundstrom:2004:DDG

[1535] L. F. Sundström, T. Bohlin, and J. I. Johnsson. Density-dependent growth in hatchery-reared brown trout released into a natural stream. Journal of Fish Biology, 65(5):1385–1391, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:2004:NME

[1536] P. J. Walsh, Z. Wei, C. M. Wood, A. M. Loong, K. C. Hiong, S. M. L. Lee, W. P. Wong, S. F. Chew, and Y. K. Ip. Nitrogen metabolism and excretion in Allenbatrachus grunniens (L): effects of variable salinity, confinement, high pH and ammonia loading. Journal of Fish Biology, 65(5):1392–1411, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:2004:OMC

[1537] A. M. Murray. Osteology and morphology of the characiform fish Alestes stuhlmannii Pfeffer, 1896 (Alestidae) from the Rufiji River basin, East Africa. Journal of Fish Biology, 65(5):1412–1430, November 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cakic:2004:DAC

[1538] P. Cakic, M. Lenhardt, J. Kolarevic, B. Mickovic, and A. Hegedis. Distribution of the Asiatic cyprinid Pseudorasbora parva in Serbia and Montenegro. Journal of Fish Biology, 65(5):1431–1434, November 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCarthy:2004:SIA

[1539] I. D. McCarthy, D. Fraser, S. Waldron, and C. E. Adams. A stable isotope analysis of trophic polymorphism among Arctic charr from Loch Ericht, Scotland. Journal of Fish Biology, 65(5):1435–1440, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibbotson:2004:CRA

[1540] A. T. Ibbotson, W. R. C. Beaumont, D. Collinson, A. Wilkinson, and A. C. Pinder. A cross-river antenna array for the detection of miniature passive integrated transponder tags in deep, fast flowing rivers. Journal of Fish Biology, 65(5):1441–1443, November 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:IV

[1541] Anonymous. Index, volume 65. Journal of Fish Biology, 65(6):i–xxiv, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoner:2004:EEV

[1542] A. W. Stoner. Effects of environmental variables on fish feeding ecology: implications for the performance of baited fishing gear and stock assessment. Journal of Fish Biology, 65(6):1445–1471, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sardella:2004:PAH

[1543] N. H. Sardella and J. T. Timi. Parasites of Argentine hake in the Argentine Sea: population and infracommunity structure as evidence for host stock discrimination. Journal of Fish Biology, 65(6):1472–1488, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prisco:2004:NMD

[1544] M. Prisco, L. Ricchiari, R. Montella, A. Liguoro, and P. Andreuccetti. The nucleolus and its modifications during oogenesis of Torpedo marmorata. Journal of Fish Biology, 65(6):1489–1497, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eckmann:2004:OCM

[1545] R. Eckmann. Overwinter changes in mass and lipid content of Perca fluviatilis and Gymnocephalus cernuus. Journal of Fish Biology, 65(6): 1498–1511, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cruz:2004:OSR

[1546] A. Cruz and A. Lombarte. Otolith size and its relationship with colour patterns and sound production. Journal of Fish Biology, 65(6):1512– 1525, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frantzen:2004:EFT

[1547] M. Frantzen, B. Damsgård, H. Tveiten, S. Moriyama, M. Iwata, and H. K. Johnsen. Effects of fasting on temporal changes in plasma concentrations of sex steroids, growth hormone and insulin-like growth factor i, and reproductive investment in Arctic charr. Journal of Fish Biology, 65(6):1526–1542, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schulze:2004:CAH

[1548] T. Schulze, U. Kahl, R. J. Radke, and J. Benndorf. Consumption, abundance and habitat use of Anguilla anguilla in a mesotrophic reservoir. Journal of Fish Biology, 65(6):1543–1562, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mollmann:2004:FEC

[1549] C. Möllmann, G. Kornilovs, M. Fetter, and F. W. Köster. Feeding ecology of central Baltic Sea herring and sprat. Journal of Fish Biology, 65(6): 1563–1581, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:2004:OPH

[1550] A. J. King. Ontogenetic patterns of habitat use by fishes within the main channel of an Australian floodplain river. Journal of Fish Biology, 65(6): 1582–1603, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Snickars:2004:ABY

[1551] M. Snickars, A. Sandström, and J. Mattila. Antipredator behaviour of 0+ year Perca fluviatilis: effect of vegetation density and turbidity. Journal of Fish Biology, 65(6):1604–1613, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chiou:2004:UEC

[1552] L. M. Chiou, M. C. Chung, P. G. Tung, T. Hsu, and J. S. Yang. The use of egg chorion glycoprotein of Epinephelus malabaricus for egg identification. Journal of Fish Biology, 65(6):1614–1621, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paxton:2004:BAI

[1553] C. G. M. Paxton, I. J. Winfield, J. M. Fletcher, D. G. George, and D. P. Hewitt. Biotic and abiotic influences on the recruitment of male perch in Windermere, U.K. Journal of Fish Biology, 65(6):1622–1642, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paukert:2004:CET

[1554] C. P. Paukert. Comparison of electrofishing and trammel netting variability for sampling native fishes. Journal of Fish Biology, 65(6):1643–1652, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cecchini:2004:BRE

[1555] S. Cecchini, A. R. Caputo, and F. Mecca. Book review: Enhancement of pH-resistant iron-binding activity in supernatants of rainbow trout blood leucocytes by in vitro treatment with phorbol-12-myristate-13-acetate. Journal of Fish Biology, 65(6):1653–1656, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yacoob:2004:ERO

[1556] S. Y. Yacoob, H. I. Browman, and P. A. Jensen. Electroencephalogram recordings from the olfactory bulb of juvenile (0 year) Atlantic cod in response to amino acids. Journal of Fish Biology, 65(6):1657–1664, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balon:2004:AOD

[1557] E. K. Balon. About the oldest domesticates among fishes. Journal of Fish Biology, 65(s1):1–27, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Araguas:2004:BBT

[1558] R. M. Araguas, N. Sanz, C. Pla, and J. L. García-Marín. Breakdown of the brown trout evolutionary history due to hybridization between native and cultivated fish. Journal of Fish Biology, 65(s1):28–37, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carr:2004:ERC

[1559] J. W. Carr, F. Whoriskey, and P. O’Reilly. Efficacy of releasing captive reared broodstock into an imperilled wild Atlantic salmon population as a recovery strategy. Journal of Fish Biology, 65(s1):38–54, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cotterell:2004:ESD

[1560] S. P. Cotterell and C. S. Wardle. Endurance swimming of diploid and triploid Atlantic salmon. Journal of Fish Biology, 65(s1):55–68, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fairchild:2004:FAP

[1561] E. A. Fairchild and W. H. Howell. Factors affecting the post-release survival of cultured juvenile Pseudopleuronectes americanus. Journal of Fish Biology, 65(s1):69–87, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hawkins:2004:PIH

[1562] L. A. Hawkins, J. D. Armstrong, and A. E. Magurran. Predator-induced hyperventilation in wild and hatchery Atlantic salmon fry. Journal of Fish Biology, 65(s1):88–100, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hayes:2004:IBE

[1563] S. A. Hayes, M. H. Bond, C. V. Hanson, and R. B. MacFarlane. Interactions between endangered wild and hatchery salmonids: can the pitfalls of artificial propagation be avoided in small coastal streams? Journal of Fish Biology, 65(s1):101–121, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2004:IDR

[1564] F. A. Huntingford. Implications of domestication and rearing conditions for the behaviour of cultivated fishes. Journal of Fish Biology, 65(s1): 122–142, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malavasi:2004:DPA

[1565] S. Malavasi, V. Georgalas, M. Lugli, P. Torricelli, and D. Mainardi. Differences in the pattern of antipredator behaviour between hatcheryreared and wild European sea bass juveniles. Journal of Fish Biology, 65 (s1):143–155, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin-Smith:2004:GHI

[1566] K. M. Martin-Smith, J. D. Armstrong, J. I. Johnsson, and B. Th. Björnsson. Growth hormone increases growth and dominance of wild juvenile Atlantic salmon without affecting space use. Journal of Fish Biology, 65(s1):156–172, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McGinnity:2004:DLS

[1567] P. McGinnity, P. Prodöhl, N. Ó Maoiléidigh, R. Hynes, D. Cotter, N. Baker, B. O’Hea, and A. Ferguson. Differential lifetime success and performance of native and non-native Atlantic salmon examined under communal natural conditions. Journal of Fish Biology, 65(s1):173–187, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nislow:2004:TPC

[1568] K. H. Nislow, S. Einum, and C. L. Folt. Testing predictions of the critical period for survival concept using experiments with stocked Atlantic salmon. Journal of Fish Biology, 65(s1):188–200, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orpwood:2004:EDC

[1569] J. E. Orpwood, S. W. Griffiths, and J. D. Armstrong. Effect of density on competition between wild and hatchery-reared Atlantic salmon for shelter in winter. Journal of Fish Biology, 65(s1):201–209, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paulsen:2004:GRN

[1570] H. Paulsen and J. G. Støttrup. Growth rate and nutritional status of wild and released reared juvenile turbot in southern Kattegat, Denmark. Journal of Fish Biology, 65(s1):210–230, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salonen:2004:SHH

[1571] A. Salonen and N. Peuhkuri. A short hatchery history: does it make a difference to aggressiveness in European grayling? Journal of Fish Biology, 65(s1):231–239, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slavik:2004:WRP

[1572] O. Slavík and L. Bartoš. What are the reasons for the Prussian carp expansion in the upper Elbe River, Czech Republic? Journal of Fish Biology, 65(s1):240–253, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smoker:2004:ESS

[1573] W. W. Smoker, I. A. Wang, A. J. Gharrett, and J. J. Hard. Embryo survival and smolt to adult survival in second-generation outbred Coho salmon. Journal of Fish Biology, 65(s1):254–262, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorpe:2004:LHR

[1574] J. E. Thorpe. Life history responses of fishes to culture. Journal of Fish Biology, 65(s1):263–285, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thrower:2004:GAG

[1575] F. P. Thrower, J. J. Hard, and J. E. Joyce. Genetic architecture of growth and early life-history transitions in anadromous and derived freshwater populations of steelhead. Journal of Fish Biology, 65(s1):286–307, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2004:DSD

[1576] A. A. L. Taylor, J. R. Britton, and I. G. Cowx. Does the stock density of stillwater catch and release fisheries affect the growth performance of introduced cultured barbel? Journal of Fish Biology, 65(s1):308–313, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baer:2004:SHR

[1577] J. Baer. Stocking hatchery-reared brown trout in different densities into a wild population — a comparison of growth and movement. Journal of Fish Biology, 65(s1):314, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anras:2004:SAS

[1578] M.-L. Bégout Anras, J. P. Lagardère, and D. Covès. Swimming activity of seabass: comparing patterns obtained in natural environment and in re-circulating tanks under high density. Journal of Fish Biology, 65(s1): 314–315, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brummett:2004:GQD

[1579] R. E. Brummett and R. Ponzoni. Genetic quality of domesticated African tilapia populations. Journal of Fish Biology, 65(s1):315, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burke:2004:PPJ

[1580] J. S. Burke, J. Morris, and J. Kenworthy. Polyculture production of juvenile fishes for survival in nature. Journal of Fish Biology, 65(s1):315, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chatterji:2004:EIS

[1581] R. K. Chatterji, D. N. Stubbing, A. W. Bark, and W. P. Williams. Experimental investigation of the success of stocked, domestic brown trout and their effects on wild brown trout in upland and lowland rivers. Journal of Fish Biology, 65(s1):315–316, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chebanov:2004:SEC

[1582] M. S. Chebanov, E. V. Galich, and Y. N. Chmir. Stock enhancement and conservation culture of sturgeons: problems and prospects. Journal of Fish Biology, 65(s1):316, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciesla:2004:EDI

[1583] M. Ciesla and R. Wojda. Effect of domestication on ide Leuciscus idus L. reproductive parameters. Journal of Fish Biology, 65(s1):316, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Enders:2004:DEC

[1584] E. C. Enders, D. Boisclair, and A. G. Roy. Differences in the energetic cost of swimming in turbulent flow between wild, farmed and domesticated juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 65(s1):317, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fleming:2004:REC

[1585] I. A. Fleming. Reproductive ecology of cultured fish in the wild. Journal of Fish Biology, 65(s1):317, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

deLeaniz:2004:MPM

[1586] C. García de Leániz, S. Consuegra, and A. Serdio. Maladaptation and phenotypic mismatch in cultured Atlantic salmon used for stocking. Journal of Fish Biology, 65(s1):317–318, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hirvonen:2004:WDS

[1587] H. Hirvonen and M. V. M. Laakkonen. Why do salmonid antipredator responses weaken in hatchery rearing? Journal of Fish Biology, 65(s1): 318, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jorstad:2004:LTS

[1588] K. E. Jørstad, G. Nævdal, Ø. Karlsen, S. Torkildsen, O. I. Paulsen, and H. Otterå. Long-term studies on genetic interaction between wild and ranched cod Gadus morhua by use of a genetic marked strain. Journal of Fish Biology, 65(s1):318–319, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Joyce:2004:CMT

[1589] J. E. Joyce, A. C. Wertheimer, A. Gray, and F. Thrower. Comparison of maturation timing, egg size and fecundity between hatchery lines of Chinook salmon and their wild donor stocks. Journal of Fish Biology, 65(s1):319, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lawlor:2004:CFR

[1590] J. L. Lawlor and J. A. Hutchings. Consequences to fitness-related traits of hybridization between farmed and wild Atlantic salmon, Salmo salar. Journal of Fish Biology, 65(s1):319, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lenhardt:2004:CAM

[1591] M. Lenhardt, M. Prokes, I. Z. Jaric, V. Barus, J. Kolarevic, I. Krupka, G. Cvijanovic, P. Cakic, and Z. Gacic. Comparative analysis of morphometric characters of juvenile sterlet Acipenser ruthenus L. from natural population and aquaculture. Journal of Fish Biology, 65(s1):320, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorenzen:2004:PDF

[1592] K. Lorenzen. Population dynamics of fisheries stock enhancement. Journal of Fish Biology, 65(s1):320, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorenzen:2004:DCB

[1593] K. Lorenzen, M. Beveridge, and M. Mangel. Domestication, comparative biology and interactions of wild and cultured fish: convenor’s synthesis. Journal of Fish Biology, 65(s1):320–321, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mairesse:2004:RPV

[1594] G. Mairesse, M. Thomas, J.-N. Gardeur, and J. Brun-Bellut. Reared perch v. wild perch — first results about organoleptical and technological quality. Journal of Fish Biology, 65(s1):321, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peeler:2004:DIB

[1595] E. J. Peeler and A. G. Murray. Disease interaction between farmed and wild fish populations. Journal of Fish Biology, 65(s1):321–322, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skilbrei:2004:CSP

[1596] O. T. Skilbrei and V. Wennevik. A comparative study of the performance of family groups of Atlantic salmon reared in tanks and released in a river. Journal of Fish Biology, 65(s1):322, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spedicato:2004:LHT

[1597] M. T. Spedicato, P. Carbonara, and G. Lembo. Life-history traits of the common pandora Pagellus erythrinus L., interpreted using information from aquaculture experiments. Journal of Fish Biology, 65(s1):322, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stephens:2004:EMI

[1598] A. Stephens and A. B. Cooper. Ecological model of interactions between escaped and wild Atlantic salmon Salmo salar. Journal of Fish Biology, 65(s1):323, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suquet:2004:WDS

[1599] M. Suquet, M.-J. Rochet, and J. L. Gaignon. Why do some fish do it younger than others? learning from experiments. Journal of Fish Biology, 65(s1):323, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Utter:2004:PGC

[1600] F. Utter. Population genetics, conservation and evolution in salmonids and other widely cultured fishes: some perspectives over six decades. Journal of Fish Biology, 65(s1):323–324, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaha:2004:EAF

[1601] J.-P. K. Vähä and C. R. Primmer. Efficiency of allele frequency-based Bayesian programmes for detecting hybridization between farmed and wild salmon. Journal of Fish Biology, 65(s1):324, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vasemagi:2004:CRR

[1602] A. Vasemägi. Compensatory releases reduce genetic differentiation among Atlantic salmon populations in the Baltic Sea: evidence from the River Ume–Vindelälven. Journal of Fish Biology, 65(s1):324–325, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walrut:2004:DAO

[1603] B. P. Walrut. Domestication and absolute ownership of fish in the English common law. Journal of Fish Biology, 65(s1):325, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almuly:2004:VNT

[1604] R. Almuly, Y. Poleg-Danin, Y. Kashi, S. Gorshkov, G. Gorshkova, A. Dyman, B. Cavari, B. Rapoport, M. Soller, and B. Funkenstein. Variable number of tandem repeats in the growth hormone gene of Sparus aurata: association with growth and effect on gene transcription. Journal of Fish Biology, 65(s1):326, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beyer:2004:EPI

[1605] K. Beyer. Escapees of potentially invasive fishes from an ornamental aquaculture facility: the case of topmouth gudgeon Pseudorasbora parva. Journal of Fish Biology, 65(s1):326–327, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciesla:2004:CNC

[1606] M. Ciesla, J. Sliwinski, P. Konieczny, and R. Litynski. Changes of nase Chondrostoma nasus L. occurrence in the Wislok River, southern Poland, after stocking with pond reared juveniles. Journal of Fish Biology, 65 (s1):327, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciesla:2004:IRM

[1607] M. Ciesla and Z. Kaczkowski. Influence of rearing method on ide Leuciscus idus L. juvenile survival under pike and pikeperch predation. Journal of Fish Biology, 65(s1):327, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conides:2004:LSE

[1608] A. J. Conides, B. Glamuzina, and C. Papaconstantinou. Laboratory simulation of the effects of environmental salinity on wild-caught juveniles of European sea bass Dicentrarchus labrax and gilthead seabream Sparus aurata. Journal of Fish Biology, 65(s1):327–328, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duchemin:2004:PTE

[1609] M. B. Duchemin, C. Audet, and Y. Lambert. Photoperiod and temperature effects on gametogenesis in winter flounder Pseudopleuronectes americanus. Journal of Fish Biology, 65(s1):328, December 2004. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gale:2004:PBD

[1610] W. L. Gale, M. S. Hill, and G. B. Zydlewski. Physiological and behavioural differences of hatchery and wild-reared steelhead Oncorhynchus mykiss smolts of the same genetic origin. Journal of Fish Biology, 65(s1): 328–329, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galich:2004:CES

[1611] E. V. Galich and M. S. Chebanov. Comparative evaluation of sturgeon larvae and juveniles reared under control of seasonal propogation of the wild and domestic breeders. Journal of Fish Biology, 65(s1):329, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubert:2004:GEP

[1612] S. Hubert, A. M. O’Keeffe, D. Cotter, N. P. Wilkins, and M. T. Cairns. Gene expression patterns in Atlantic salmon Salmo salar: gene expression during osmoregulation in intestine tissue. Journal of Fish Biology, 65 (s1):329–330, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ingle:2004:FPA

[1613] E. Ingle and L. Venzi. Feasibility of a project for active restocking of the sea adjacent the salt pans of Tarquinia (Italy). Journal of Fish Biology, 65(s1):330, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Islam:2004:DPP

[1614] A. Islam, O. V. Gorshkov, V. M. Chernov, and V. A. Kuznetsov. DNA polymorphism of Percidae population structure in the Kuibyshev water reservoir of Russia. Journal of Fish Biology, 65(s1):330, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaczkowski:2004:IAN

[1615] Z. Kaczkowski, M. Ciesla, and M. Zalewski. The importance of aquaculture to nase Chondrostoma nasus conservation in Poland. Journal of Fish Biology, 65(s1):331, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaczkowski:2004:SRA

[1616] Z. Kaczkowski and C. Miroslaw. State and role of active conservation of barbel Barbus barbus in Poland. Journal of Fish Biology, 65(s1):331, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kulczykowska:2004:DMR

[1617] E. Kulczykowska, E. Sokolowska, M. Gozdowska, and H. Kalamarz. Disruption of the melatonin rhythm in wild and farmed fish: a consequence of prolonged thyroxine administration and calcium depletion. Journal of Fish Biology, 65(s1):331–332, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Micale:2004:GMB

[1618] V. Micale, G. Maricchiolo, and L. Genovese. Gonadal maturation in the blackspot seabream Pagellus bogaraveo: a comparison between a farmed and a wild broodstock. Journal of Fish Biology, 65(s1):332, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Micale:2004:OAT

[1619] V. Micale, M. Garaffo, L. Genovese, M. T. Spedicato, and U. Muglia. The ontogeny of the alimentary tract of larval pandora, Pagellus erythrinus L. Journal of Fish Biology, 65(s1):332–333, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Omnes:2004:IDD

[1620] M. H. Omnes, S. Recek, H. Barone, H. Le Delliou, A. Schmitz, A. Mutelet, M. Suquet, and J. H. Robin. Influence of dry diets on reproductive performance and egg lipid composition during the first spawning season of captive pollack. Journal of Fish Biology, 65(s1):333, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pujolar:2004:DGV

[1621] J. M. Pujolar, G. E. Maes, and F. A. M. Volckaert. Distribution of genetic variation in farmed and natural stocks of European eel. Journal of Fish Biology, 65(s1):333–334, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowlerson:2004:MDC

[1622] A. M. Rowlerson, P. Silva, E. Rocha, M. Olmedo, and L. M. P. Valente. Muscle development in cultured blackspot seabream Pagellus bogaraveo: preliminary histochemical and immunohistochemical data on the fibre types. Journal of Fish Biology, 65(s1):334, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2004:IDP

[1623] P. Silva, C. A. P. Andrade, M. F. A. Timóteo Viriato, E. Rocha, and L. M. P. Valente. Influence of dietary protein level on growth performance and body composition of juvenile blackspot seabream, Pagellus bogaraveo (Brunnich, 1768). Journal of Fish Biology, 65(s1):334–335, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Varricchio:2004:CCS

[1624] E. Varricchio, T. Rubino, S. Paino, T. Di Lascio, O. Paciello, and M. Langella. The calpastatin/calpain system in trout Salmo trutta trutta muscle. Journal of Fish Biology, 65(s1):335, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2004:HTI

[1625] I. A. Wang, W. W. Smoker, S. E. Gilk, D. S. Oxman, and A. J. Gharrett. Hatching time as an indicator of environmental incompatibility and outbreeding depression in intraspecific salmon hybrids. Journal of Fish Biology, 65(s1):335–336, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:RS

[1626] Anonymous. Referees for this supplement. Journal of Fish Biology, 65 (s1):337, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wheeler:2004:I

[1627] A. C. Wheeler, N. R. Merrett, and D. T. G. Quigley. Introduction. Journal of Fish Biology, 65(s2):i–iii, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2004:ARN

[1628] Anonymous. Additional records and notes for Wheeler’s (1992) List of the Common and Scientific Names of Fishes of the British Isles. Journal of Fish Biology, 65(s2):1–34, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [1].

Anonymous:2004:I

[1629] Anonymous. Index. Journal of Fish Biology, 65(s2):35–40, December 2004. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:2005:NZC

[1630] F. J. Ward, T. G. Northcote, and J. A. T. Boubée. The New Zealand common smelt: biology and ecology. Journal of Fish Biology, 66(1):1– 32, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finstad:2005:ESI

[1631] A. G. Finstad. Effect of sampling interval and temperature on the accuracy of food consumption estimates from stomach contents. Journal of Fish Biology, 66(1):33–44, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barletta:2005:RSS

[1632] M. Barletta, A. Barletta-Bergan, U. Saint-Paul, and G. Hubold. The role of salinity in structuring the fish assemblages in a tropical estuary.

Journal of Fish Biology, 66(1):45–72, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blanchard:2005:LCF

[1633] G. Blanchard, X. Druart, and P. Kestemont. Lipid content and fatty acid composition of target tissues in wild Perca fluviatilis females in relation to hepatic status and gonad maturation. Journal of Fish Biology, 66 (1):73–85, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finstad:2005:MHR

[1634] B. Finstad, F. Økland, E. B. Thorstad, P. A. BjØrn, and R. S. McKinley. Migration of hatchery-reared Atlantic salmon and wild anadromous brown trout post-smolts in a Norwegian fjord system. Journal of Fish Biology, 66(1):86–96, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quintella:2005:UPT

[1635] B. R. Quintella, N. O. Andrade, R. Espanhol, and P. R. Almeida. The use of PIT telemetry to study movements of ammocoetes and metamorphosing sea lampreys in river beds. Journal of Fish Biology, 66(1):97–106, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rubin:2005:SCA

[1636] J.-F. Rubin, C. Glimsäter, and T. Jarvi. Spawning characteristics of the anadromous brown trout in a small Swedish stream. Journal of Fish Biology, 66(1):107–121, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grover:2005:CSA

[1637] M. C. Grover. Changes in size and age at maturity in a population of Kokanee Oncorhynchus nerka during a period of declining growth conditions. Journal of Fish Biology, 66(1):122–134, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Truemper:2005:GLP

[1638] H. A. Truemper and T. E. Lauer. Gape limitation and piscine prey sizeselection by yellow perch in the extreme southern area of Lake Michigan, with emphasis on two exotic prey items. Journal of Fish Biology, 66 (1):135–149, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2005:HRA

[1639] K. M. M. Jones. Home range areas and activity centres in six species of Caribbean wrasses (Labridae). Journal of Fish Biology, 66(1):150–166, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanDijk:2005:ERD

[1640] P. L. M. Van Dijk, I. Hardewig, and F. Hölker. Energy reserves during food deprivation and compensatory growth in juvenile roach: the importance of season and temperature. Journal of Fish Biology, 66(1): 167–181, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bertram:2005:EWC

[1641] B. O. B. Bertram and R. C. Playle. Effects of waterborne complexing agents on silver uptake and depuration in rainbow trout. Journal of Fish Biology, 66(1):182–197, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imre:2005:MPN

[1642] I. Imre and D. Boisclair. Moon phase and nocturnal density of Atlantic salmon parr in the Sainte-Marguerite River, Québec. Journal of Fish Biology, 66(1):198–207, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baker:2005:HJA

[1643] D. W. Baker, A. M. Wood, M. K. Litvak, and J. D. Kieffer. Haematology of juvenile Acipenser oxyrinchus and Acipenser brevirostrum at rest and following forced activity. Journal of Fish Biology, 66(1):208–221, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pedini:2005:GDG

[1644] V. Pedini, C. Dall’Aglio, F. Parillo, and P. Scocco. Glycoconjugate distribution in gastric fundic mucosa of Umbrina cirrosa L. revealed by lectin histochemistry. Journal of Fish Biology, 66(1):222–229, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:2005:VPR

[1645] D. Burton. In vitro pattern-related effects of forskolin on melanophore adrenoceptor-mediated responses of winter flounder. Journal of Fish Biology, 66(1):230–241, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Enders:2005:FBJ

[1646] E. C. Enders, T. Buffin-Bélanger, D. Boisclair, and A. G. Roy. The feeding behaviour of juvenile Atlantic salmon in relation to turbulent flow. Journal of Fish Biology, 66(1):242–253, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davis:2005:BIC

[1647] M. W. Davis. Behaviour impairment in captured and released sablefish: ecological consequences and possible substitute measures for delayed discard mortality. Journal of Fish Biology, 66(1):254–265, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greisman:2005:AFC

[1648] L. Greisman and P. Moller. The anal fin complex in a weakly discharging electric fish, Gnathonemus petersii (Mormyridae). Journal of Fish Biology, 66(1):266–275, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landaeta:2005:LCP

[1649] M. F. Landaeta, G. Herrera, and L. R. Castro. Larvae of Crystallodytes pauciradiatus (Perciformes: Creediidae) from Easter Island, South Pacific. Journal of Fish Biology, 66(1):276–282, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahy:2005:JSA

[1650] E. Fahy, P. Green, and D. T. G. Quigley. Juvenile Sparus aurata L. on the south coast of Ireland. Journal of Fish Biology, 66(1):283–289, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eastwood:2005:BRG

[1651] Paul Eastwood. Book review: Geographic Information Systems in Fisheries. Edited by W. L. Fisher and F. J. Rahel. 275 pp. Published by American Fisheries Society, Bethesda, Maryland, 2004. Price $69.00. ISBN 1-888569-57-3. Journal of Fish Biology, 66(1):290–291, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2005:BRH

[1652] Peter J. Miller. Book review: Handbook of Fish Biology and Fisheries, Vol. 1, FISH BIOLOGY. Edited by P. J. B. Hart and J. D. Reynolds. xiii 413 pp. Published by Blackwell Science Ltd, Oxford, 2002. Price £70.00. ISBN 0-632-05412-3. Journal of Fish Biology, 66(1):291–292, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welcomme:2005:BRG

[1653] Robin Welcomme. Book review: Guidelines for the Integrated Management of the Watershed — Phytotechnology and Ecohydrology. Edited By M. Zalewski. v 188 pp. Published By United Nations Environment Programme (Freshwater Management Series No 5), 2002. ISBN 92-807-2059-7. Journal of Fish Biology, 66(1):292–293, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:BRD

[1654] Anonymous. Book review: Diseases and Disorders of Finfish in Cage Culture. Edited by P. T. K. Woo, D. W. Bruno and L. H. S. Lim. x + 354 pp. Published by CABI Publishing, Wallingford, U.K., 2002. Price £85.00. ISBN 0-85199-443-1. Journal of Fish Biology, 66(1):294, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:R

[1655] Anonymous. Referees 2004. Journal of Fish Biology, 66(1):295–299, January 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santoul:2005:SPB

[1656] F. Santoul, J. Cayrou, S. Mastrorillo, and R. Céréghino. Spatial patterns of the biological traits of freshwater fish communities in south-west France. Journal of Fish Biology, 66(2):301–314, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sala:2005:GOS

[1657] R. Sala, C. A. Santamaría, and S. Crespo. Growth of organ systems of Dentex dentex (L) and Psetta maxima (L) during larval development. Journal of Fish Biology, 66(2):315–326, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carruthers:2005:LTC

[1658] E. H. Carruthers, J. D. Neilson, C. Waters, and P. Perley. Long-term changes in the feeding of Pollachius virens on the Scotian Shelf: responses to a dynamic ecosystem. Journal of Fish Biology, 66(2):327–347, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vainikka:2005:ETG

[1659] A. Vainikka, E. I. Jokinen, R. Kortet, S. Paukku, J. Pirhonen, M. J. Rantala, and J. Taskinen. Effects of testosterone and β-glucan on immune functions in tench. Journal of Fish Biology, 66(2):348–361, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dou:2005:ETD

[1660] S. Z. Dou, R. Masuda, M. Tanaka, and K. Tsukamoto. Effects of temperature and delayed initial feeding on the survival and growth of Japanese flounder larvae. Journal of Fish Biology, 66(2):362–377, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thacker:2005:UGS

[1661] C. Thacker and H. Grier. Unusual gonad structure in the paedomorphic teleost Schindleria praematura (Teleostei: Gobioidei): a comparison with other gobioid fishes. Journal of Fish Biology, 66(2):378–391, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OConnor:2005:MMA

[1662] J. P. O’Connor, D. J. O’Mahony, and J. M. O’Mahony. Movements of Macquaria ambigua, in the Murray River, south-eastern Australia. Journal of Fish Biology, 66(2):392–403, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Growns:2005:RBR

[1663] I. Growns and M. James. Relationships between river flows and recreational catches of Australian bass. Journal of Fish Biology, 66(2):404– 416, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Iglesias:2005:DNH

[1664] S. P. Iglésias, D. Y. Sellos, and K. Nakaya. Discovery of a normal hermaphroditic chondrichthyan species: Apristurus longicephalus. Journal of Fish Biology, 66(2):417–428, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurel:2005:BRN

[1665] B. J. Laurel, C. J. Laurel, J. A. Brown, and R. S. Gregory. Book review: A new technique to gather 3-D spatial information using a single camera. Journal of Fish Biology, 66(2):429–441, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Motta:2005:CCA

[1666] C. M. Motta, S. Tammaro, P. Simoniello, M. Prisco, L. Ricchiari, P. Andreuccetti, and S. Filosa. Characterization of cortical alveoli content in several species of Antarctic notothenioids. Journal of Fish Biology, 66 (2):442–453, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes-Ferreira:2005:VDF

[1667] A. Gomes-Ferreira, F. Ribeiro, L. Moreira da Costa, I. G. Cowx, and M. J. Collares-Pereira. Variability in diet and foraging behaviour between sexes and ploidy forms of the hybridogenetic Squalius alburnoides complex (Cyprinidae) in the Guadiana River basin, Portugal. Journal of Fish Biology, 66(2):454–467, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kjoss:2005:DDR

[1668] V. A. Kjoss, C. N. Kamunde, S. Niyogi, M. Grosell, and C. M. Wood. Dietary Na does not reduce dietary Cu uptake by juvenile rainbow trout. Journal of Fish Biology, 66(2):468–484, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lacroix:2005:SBP

[1669] G. L. Lacroix, D. Knox, and M. J. W. Stokesbury. Survival and behaviour of post-smolt Atlantic salmon in coastal habitat with extreme tides. Journal of Fish Biology, 66(2):485–498, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jang:2005:RER

[1670] M.-H. Jang and M. C. Lucas. Reproductive ecology of the river lamprey. Journal of Fish Biology, 66(2):499–512, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ramos-Miranda:2005:STC

[1671] J. Ramos-Miranda, L. Quiniou, D. Flores-Hernandez, T. Do-Chi, L. Ayala-Perez, and A. Sosa-Lopez. Spatial and temporal changes in the nekton of the Terminos Lagoon, Campeche, Mexico. Journal of Fish Biology, 66(2):513–530, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2005:FEG

[1672] B. S. Andersen, J. D. Carl, P. Grønkjær, and J. G. Støttrup. Feeding ecology and growth of age 0 year Platichthys flesus (L.) in a vegetated and a bare sand habitat in a nutrient rich fjord. Journal of Fish Biology, 66(2):531–552, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lafrance:2005:OCT

[1673] P. Lafrance, M. Castonguay, D. Chabot, and C. Audet. Ontogenetic changes in temperature preference of Atlantic cod. Journal of Fish Biology, 66(2):553–567, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freyhof:2005:FRM

[1674] J. Freyhof and E. Korte. The first record of Misgurnus anguillicaudatus in Germany. Journal of Fish Biology, 66(2):568–571, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2005:AAB

[1675] J. C. Svendsen, A. Koed, and M. C. Lucas. The angle of attack of the body of common bream while swimming at different speeds in a flume tank. Journal of Fish Biology, 66(2):572–577, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinni:2005:TSR

[1676] M. Vinni, J. Lappalainen, and J. Horppila. Temporal and size-related changes in the frequency of empty stomachs in smelt. Journal of Fish Biology, 66(2):578–582, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pekcan-Hekim:2005:OPG

[1677] Z. Pekcan-Hekim, R. Rahkonen, and J. Horppila. Occurrence of the parasite Glugea hertwigi in young-of-the-year smelt in Lake Tuusulanjärvi. Journal of Fish Biology, 66(2):583–588, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reyjol:2005:ESB

[1678] Y. Reyjol, G. Loot, and S. Lek. Estimating sampling bias when using electrofishing to catch stone loach. Journal of Fish Biology, 66(2):589– 591, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaza:2005:PSE

[1679] G. Plaza, H. Honda, H. Sakaji, and K. Nashida. Preparing sagittae for examination of daily growth increments of young-of-the-year fishes: a modification of the embed method. Journal of Fish Biology, 66(2): 592–597, February 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2005:TRS

[1680] X. Y. Chen and J. X. Yang. Triplophysa rosa sp. nov.: a new blind loach from China. Journal of Fish Biology, 66(3):599–608, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cecchini:2005:EVE

[1681] S. Cecchini, A. R. Caputo, and F. Mecca. Evaluation of in vitro endocytosis and antibody synthesis by rainbow trout head kidney cells treated with bovine lactoferrin. Journal of Fish Biology, 66(3):609–617, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebert:2005:RBS

[1682] D. A. Ebert. Reproductive biology of skates, Bathyraja (Ishiyama), along the eastern Bering Sea continental slope. Journal of Fish Biology, 66(3): 618–649, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gratwicke:2005:RBF

[1683] B. Gratwicke and M. R. Speight. The relationship between fish species richness, abundance and habitat complexity in a range of shallow tropical marine habitats. Journal of Fish Biology, 66(3):650–667, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Close:2005:LDS

[1684] P. G. Close, B. J. Pusey, and A. H. Arthington. Larval description of the sympatric species, Craterocephalus stercusmuscarum stercusmuscarum (Pisces: Atherinidae) and Mogurnda adspersa (Pisces: Eleotridae) from tropical streams of north-east Queensland, Australia. Journal of Fish Biology, 66(3):668–684, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cech:2005:DVM

[1685] M. Čech, M. Kratochvíl, J. Kubečka, V. Draštík, and J. Matěna. Diel vertical migrations of bathypelagic perch fry. Journal of Fish Biology, 66(3):685–702, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reiss:2005:FES

[1686] C. Reiss, I. McLaren, P. Avendaño, and C. Taggart. Feeding ecology of silver hake larvae on the Western Bank, Scotian Shelf, and comparison with Atlantic cod. Journal of Fish Biology, 66(3):703–720, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aarestrup:2005:MMS

[1687] K. Aarestrup, N. Jepsen, A. Koed, and S. Pedersen. Movement and mortality of stocked brown trout in a stream. Journal of Fish Biology, 66(3):721–728, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2005:HHE

[1688] H.-T. Lin, H.-Y. Lin, and H.-L. Yang. Histology and histochemical enzyme-staining patterns of major immune organs in Epinephelus malabaricus. Journal of Fish Biology, 66(3):729–740, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curry:2005:ARC

[1689] R. A. Curry. Assessing the reproductive contributions of sympatric anadromous and freshwater-resident brook trout. Journal of Fish Biology, 66(3):741–757, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knudsen:2005:TTP

[1690] R. Knudsen, A. H. Rikardsen, J. B. Dempson, P. A. BjØrn, B. Finstad, M. Holm, and P.-A. Amundsen. Trophically transmitted parasites in wild Atlantic salmon post-smolts from Norwegian fjords. Journal of Fish Biology, 66(3):758–772, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Revie:2005:EEC

[1691] C. W. Revie, G. Gettinby, J. W. Treasurer, and C. Wallace. Evaluating the effect of clustering when monitoring the abundance of sea lice populations on farmed Atlantic salmon. Journal of Fish Biology, 66(3): 773–783, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Toth:2005:GAO

[1692] B. Tóth, E. Várkonyi, A. Hidas, E. Edviné Meleg, and L. Váradi. Genetic analysis of offspring from intra- and interspecific crosses of Carassius auratus gibelio by chromosome and RAPD analysis. Journal of Fish Biology, 66(3):784–797, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staffan:2005:IFS

[1693] F. Staffan, C. Magnhagen, and A. Alanärä. Individual feeding success of juvenile perch is consistent over time in aquaria and under farming conditions. Journal of Fish Biology, 66(3):798–809, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bruggemann:2005:EMG

[1694] D. A. Brüggemann and M. A. Lawson. The extracellular matrix of Gadus morhua muscle contains types III, v, VI and IV collagens in addition to type I. Journal of Fish Biology, 66(3):810–821, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sponaugle:2005:GVL

[1695] S. Sponaugle, K. L. Denit, S. A. Luthy, J. E. Serafy, and R. K. Cowen. Growth variation in larval Makaira nigricans. Journal of Fish Biology, 66(3):822–835, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lovell:2005:PIE

[1696] J. M. Lovell, M. M. Findlay, R. M. Moate, and D. A. Pilgrim. The polarization of inner ear ciliary bundles from a scorpaeniform fish. Journal of Fish Biology, 66(3):836–846, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bauer:2005:PPR

[1697] C. Bauer. Potential problems with removing scales before surgical transmitter implantation. Journal of Fish Biology, 66(3):847–850, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Holmen:2005:CSA

[1698] J. Holmen, L. A. Vøllestad, K. S. Jakobsen, and C. R. Primmer. Crossspecies amplification of zebrafish and central stoneroller microsatellite loci in six other cyprinids. Journal of Fish Biology, 66(3):851–859, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2005:PIG

[1699] C. M. Jones and A. S. Grutter. Parasitic isopods (Gnathia sp.) reduce haematocrit in captive blackeye thicklip (Labridae) on the Great Barrier Reef. Journal of Fish Biology, 66(3):860–864, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hassan:2005:NRN

[1700] M. Hassan and F. Bonhomme. No reduction in neutral variability of mitochondrial and nuclear genes for a Lessepsian migrant, Upeneus moluccensis. Journal of Fish Biology, 66(3):865–870, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huysentruyt:2005:ASE

[1701] F. Huysentruyt and D. Adriaens. Adhesive structures in the eggs of Corydoras aeneus (Gill, 1858; Callichthyidae). Journal of Fish Biology, 66(3):871–876, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tobler:2005:FDC

[1702] M. Tobler. Feigning death in the Central American cichlid Parachromis friedrichsthalii. Journal of Fish Biology, 66(3):877–881, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peck:2005:OCA

[1703] M. A. Peck, C. Clemmesen, and J.-P. Herrmann. Ontogenic changes in the allometric scaling of the mass and length relationship in Sprattus sprattus. Journal of Fish Biology, 66(3):882–887, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arai:2005:EMM

[1704] T. Arai and K. Morita. Evidence of multiple migrations between freshwater and marine habitats of Salvelinus leucomaenis. Journal of Fish Biology, 66(3):888–895, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:Ea

[1705] Anonymous. Erratum. Journal of Fish Biology, 66(3):896, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:ISE

[1706] Anonymous. 10^{th} International Symposium on the Ecology of Fluvial Fishes, University of Łodz, Poland 26–29 June 2006. Journal of Fish Biology, 66(3):897, March 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Unver:2005:NHL

[1707] B. Ünver and F. Erk’akan. A natural hybrid of Leuciscus cephalus (L.) and Chalcalburnus chalcoides (Güldenstädt) (Osteichthyes–Cyprinidae) from Lake Tödürge (Sivas, Turkey). Journal of Fish Biology, 66(4):899– 910, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peck:2005:ETB

[1708] M. A. Peck, L. J. Buckley, and D. A. Bengtson. Effects of temperature, body size and feeding on rates of metabolism in young-of-the-year haddock. Journal of Fish Biology, 66(4):911–923, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svenning:2005:GPP

[1709] M.-A. Svenning, S. E. Fagermo, R. T. Barrett, R. Borgstrøm, W. Vader, T. Pedersen, and S. Sandring. Goosander predation and its potential impact on Atlantic salmon smolts in the River Tana estuary, northern Norway. Journal of Fish Biology, 66(4):924–937, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papadaki:2005:GSD

[1710] M. Papadaki, F. Piferrer, S. Zanuy, E. Maingot, P. Divanach, and C. C. Mylonas. Growth, sex differentiation and gonad and plasma levels of sex steroids in male- and female-dominant populations of Dicentrarchus labrax obtained through repeated size grading. Journal of Fish Biology, 66(4):938–956, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svenning:2005:CAC

[1711] M.-A. Svenning and R. Borgstrøm. Cannibalism in Arctic charr: do all individuals have the same propensity to be cannibals? Journal of Fish Biology, 66(4):957–965, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bozec:2005:AFH

[1712] Y.-M. Bozec, S. Dolédec, and M. Kulbicki. An analysis of fish-habitat associations on disturbed coral reefs: chaetodontid fishes in New Caledonia. Journal of Fish Biology, 66(4):966–982, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whiteman:2005:SSP

[1713] E. A. Whiteman, C. A. Jennings, and R. S. Nemeth. Sex structure and potential female fecundity in a Epinephelus guttatus spawning aggregation: applying ultrasonic imaging. Journal of Fish Biology, 66(4):983– 995, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vandewalle:2005:EDH

[1714] P. Vandewalle, G. Germeau, P. Besancenet, E. Parmentier, and E. Baras. Early development of the head skeleton in Brycon moorei (Pisces, Ostariophysi, Characidae). Journal of Fish Biology, 66(4):996–1024, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Durif:2005:SPA

[1715] C. Durif, S. Dufour, and P. Elie. The silvering process of Anguilla anguilla: a new classification from the yellow resident to the silver migrating stage. Journal of Fish Biology, 66(4):1025–1043, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brickle:2005:RSP

[1716] P. Brickle, V. Laptikhovsky, and A. Arkhipkin. Reproductive strategy of a primitive temperate notothenioid Eleginops maclovinus. Journal of Fish Biology, 66(4):1044–1059, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pangle:2005:TCR

[1717] K. L. Pangle and T. M. Sutton. Temporal changes in the relationship between condition indices and proximate composition of juvenile Coregonus artedi. Journal of Fish Biology, 66(4):1060–1072, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golub:2005:RCH

[1718] J. L. Golub, V. Vermette, and G. E. Brown. Response to conspecific and heterospecific alarm cues by pumpkinseeds in simple and complex habitats: field verification of an ontogenetic shift. Journal of Fish Biology, 66(4):1073–1081, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McClelland:2005:UPG

[1719] G. McClelland, J. Melendy, J. Osborne, D. Reid, and S. Douglas. Use of parasite and genetic markers in delineating populations of winter flounder from the central and south-west Scotian Shelf and north-east Gulf of Maine. Journal of Fish Biology, 66(4):1082–1100, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petrie-Hanson:2005:APL

[1720] L. Petrie-Hanson and A. E. Peterman. American paddlefish leukocytes demonstrate mammalian-like cytochemical staining characteristics in lymphoid tissues. Journal of Fish Biology, 66(4):1101–1115, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-cabello:2005:MES

[1721] C. Rodríguez-cabello and F. Sánchez. Mortality estimates of Scyliorhinus canicula in the Cantabrian Sea using tag-recapture data. Journal of Fish Biology, 66(4):1116–1126, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Asoh:2005:GDD

[1722] K. Asoh. Gonadal development and diandric protogyny in two populations of Dascyllus reticulatus from Madang, Papua New Guinea. Journal of Fish Biology, 66(4):1127–1148, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2005:BMO

[1723] M. D. Taylor, D. S. Fielder, and I. M. Suthers. Batch marking of otoliths and fin spines to assess the stock enhancement of Argyrosomus japonicus. Journal of Fish Biology, 66(4):1149–1162, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2005:PMF

[1724] A. H. Rikardsen and P.-A. Amundsen. Pelagic marine feeding of Arctic charr and sea trout. Journal of Fish Biology, 66(4):1163–1166, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Francois:2005:WMA

[1725] N. R. Le François, S. G. Lamarre, and P. U. Blier. Is white muscle anaerobic glycolysis capacity indicative of competitive ability in Arctic charr? Journal of Fish Biology, 66(4):1167–1176, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bremer:2005:GIC

[1726] J. R. Alvarado Bremer, M. G. Frisk, T. J. Miller, J. Turner, J. Viñas, and K. Kwil. Genetic identification of cryptic juveniles of little skate and winter skate. Journal of Fish Biology, 66(4):1177–1182, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Akyol:2005:FCR

[1727] O. Akyol, V. Ünal, T. Ceyhan, and M. Bilecenoglu. First confirmed record of Lagocephalus sceleratus (Gmelin, 1789) in the Mediterranean Sea. Journal of Fish Biology, 66(4):1183–1186, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2005:SAS

[1728] K. Morita, T. Arai, D. Kishi, and J. Tsuboi. Small anadromous Salvelinus malma at the southern limits of its distribution. Journal of Fish Biology, 66(4):1187–1192, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Christiansen:2005:OTF

[1729] J. S. Christiansen, S-E. Fevolden, and I. Byrkjedal. The occurrence of Theragra finnmarchica Koefoed, 1956 (Teleostei, Gadidae), 1932–2004. Journal of Fish Biology, 66(4):1193–1197, April 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paragamian:2005:ERD

[1730] V. L. Paragamian, R. Hardy, and B. Gunderman. Effects of regulated discharge on burbot migration. Journal of Fish Biology, 66(5):1199– 1213, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeGraaf:2005:LSN

[1731] M. De Graaf, E. D. Nentwich, J. W. M. Osse, and F. A. Sibbing. Lacustrine spawning: is this a new reproductive strategy among ‘large’ African cyprinid fishes? Journal of Fish Biology, 66(5):1214–1236, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dempster:2005:TVP

[1732] T. Dempster. Temporal variability of pelagic fish assemblages around fish aggregation devices: biological and physical influences. Journal of Fish Biology, 66(5):1237–1260, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ditty:2005:LDF

[1733] J. G. Ditty, R. F. Shaw, and L. A. Fuiman. Larval development of five species of blenny (Teleostei: Blenniidae) from the western central North Atlantic, with a synopsis of blennioid family characters. Journal of Fish Biology, 66(5):1261–1284, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lehtiniemi:2005:SHP

[1734] M. Lehtiniemi. Swim or hide: predator cues cause species specific reactions in young fish larvae. Journal of Fish Biology, 66(5):1285–1299, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mesa:2005:ABA

[1735] G. La Mesa and M. Vacchi. Analysis of the blennioid assemblages associated with different rocky shore habitats in the Ligurian Sea. Journal of Fish Biology, 66(5):1300–1327, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shoji:2005:FEM

[1736] J. Shoji, E. W. North, and E. D. Houde. The feeding ecology of Morone americana larvae in the Chesapeake Bay estuarine turbidity maximum: the influence of physical conditions and prey concentrations. Journal of Fish Biology, 66(5):1328–1341, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hellberg:2005:IEA

[1737] H. Hellberg and I. Bjerkås. Intestinal epithelium in Anarhichas lupus L., with emphasis on cell renewal. Journal of Fish Biology, 66(5):1342– 1356, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arezo:2005:EDA

[1738] M. J. Arezo, L. Pereiro, and N. Berois. Early development in the annual fish Cynolebias viarius. Journal of Fish Biology, 66(5):1357–1370, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanger:2005:MFS

[1739] A. M. Sänger, W. Davison, and S. Egginton. Muscle fine structure reflects ecotype in two nototheniids. Journal of Fish Biology, 66(5):1371–1386, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Noble:2005:DLI

[1740] C. Noble, K. Mizusawa, and M. Tabata. Does light intensity affect selffeeding and food wastage in group-held rainbow trout and white-spotted charr? Journal of Fish Biology, 66(5):1387–1399, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pall:2005:CRP

[1741] M. K. Páll, A. Hellqvist, M. Schmitz, P.-E. Olsson, I. Mayer, and B. Borg. Changes in reproductive physiology and behaviour over the nesting cycle in male three-spined sticklebacks. Journal of Fish Biology, 66(5):1400– 1410, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sadovy:2005:SSS

[1742] Y. Sadovy, J. E. Randall, and M. B. Rasotto. Skin structure in six dragonet species (Gobiesociformes; Callionymidae): interspecific differences in glandular cell types and mucus secretion. Journal of Fish Biology, 66 (5):1411–1418, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jowett:2005:RBF

[1743] I. G. Jowett, J. Richardson, and M. L. Bonnett. Relationship between flow regime and fish abundances in a gravel-bed river, New Zealand. Journal of Fish Biology, 66(5):1419–1436, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manttari:2005:EDM

[1744] S. Mänttäri, K. Anttila, and M. Järvilehto. Effects of downstream migration on myosin heavy chain expression and dihydropyridine receptor density in farmed smolt of Atlantic salmon. Journal of Fish Biology, 66 (5):1437–1446, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deynat:2005:NDS

[1745] P. P. Deynat. New data on the systematics and interrelationships of sawfishes (Elasmobranchii, Batoidea, Pristiformes). Journal of Fish Biology, 66(5):1447–1458, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2005:LTP

[1746] J. X. He, L. G. Rudstam, J. L. Forney, A. J. VanDeValk, and D. J. Stewart. Long-term patterns in growth of Oneida Lake walleye: a multivariate and stage-explicit approach for applying the von Bertalanffy growth function. Journal of Fish Biology, 66(5):1459–1470, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VonCramon-Taubadel:2005:DBS

[1747] N. Von Cramon-Taubadel, E. N. Ling, D. Cotter, and N. P. Wilkins. Determination of body shape variation in Irish hatchery-reared and wild Atlantic salmon. Journal of Fish Biology, 66(5):1471–1482, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bedard:2005:NDP

[1748] M.-E. Bédard, I. Imre, and D. Boisclair. Nocturnal density patterns of Atlantic salmon parr in the Sainte-Marguerite River, Québec, relative to the time of night. Journal of Fish Biology, 66(5):1483–1488, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bottaro:2005:MGM

[1749] M. Bottaro, S. Ferrando, T. Ferrando, P. N. Psomadakis, and M. Vacchi. Melanism in the gastric mucosa of the scalloped ribbonfish from the Ligurian Sea. Journal of Fish Biology, 66(5):1489–1492, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cromer:2005:TSB

[1750] L. Cromer, R. Williams, and J. A. E. Gibson. Trematomus scotti in Beaver Lake: the first record of a fish from a non-marine Antarctic habitat. Journal of Fish Biology, 66(5):1493–1497, May 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:IVa

[1751] Anonymous. Index, volume 66. Journal of Fish Biology, 66(6):i–xxvi, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolm:2005:DES

[1752] N. Kolm and I. Ahnesjö. Do egg size and parental care coevolve in fishes? Journal of Fish Biology, 66(6):1499–1515, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mabuchi:2005:DAL

[1753] K. Mabuchi, H. Senou, T. Suzuki, and M. Nishida. Discovery of an ancient lineage of Cyprinus carpio from Lake Biwa, central Japan, based on mtDNA sequence data, with reference to possible multiple origins of koi. Journal of Fish Biology, 66(6):1516–1528, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kakehi:2005:IAF

[1754] Y. Kakehi, K. Nakayama, K. Watanabe, and M. Nishida. Inheritance of amplified fragment length polymorphism markers and their utility in population genetic analysis of Plecoglossus altivelis. Journal of Fish Biology, 66(6):1529–1544, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakadate:2005:GIB

[1755] M. Nakadate, J. Viñas, A. Corriero, S. Clarke, N. Suzuki, and S. Chow. Genetic isolation between Atlantic and Mediterranean albacore populations inferred from mitochondrial and nuclear DNA markers. Journal of Fish Biology, 66(6):1545–1557, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martell:2005:ETD

[1756] D. J. Martell, J. D. Kieffer, and E. A. Trippel. Effects of temperature during early life history on embryonic and larval development and growth in haddock. Journal of Fish Biology, 66(6):1558–1575, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lancioni:2005:OCC

[1757] H. Lancioni, L. Lucentini, and F. Panara. Ontogenetic changes of circulating erythroid cells and haemoglobin components in Esox lucius. Journal of Fish Biology, 66(6):1576–1588, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carseldine:2005:DAH

[1758] L. Carseldine and I. R. Tibbetts. Dietary analysis of the herbivorous hemiramphid Hyporhamphus regularis ardelio: an isotopic approach. Journal of Fish Biology, 66(6):1589–1600, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lecomte:2005:DTH

[1759] F. Lecomte and J. J. Dodson. Distinguishing trophic and habitat partitioning among sympatric populations of the estuarine fish Osmerus mordax Mitchill. Journal of Fish Biology, 66(6):1601–1623, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ostrovsky:2005:AMC

[1760] I. Ostrovsky. Assessing mortality changes from size-frequency curves. Journal of Fish Biology, 66(6):1624–1632, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stevens:2005:GSG

[1761] E. D. Stevens and R. H. Devlin. Gut size in GH-transgenic Coho salmon is enhanced by both the GHtransgene and increased food intake. Journal of Fish Biology, 66(6):1633–1648, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rough:2005:HLM

[1762] K. M. Rough, B. F. Nowak, and R. E. Reuter. Haematology and leucocyte morphology of wild caught Thunnus maccoyii. Journal of Fish Biology, 66(6):1649–1659, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Humphreys:2005:OEF

[1763] R. L. Humphreys, Jr., S. E. Campana, and E. E. DeMartini. Otolith elemental fingerprints of juvenile Pacific swordfish Xiphias gladius. Journal of Fish Biology, 66(6):1660–1670, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robins:2005:LGJ

[1764] J. B. Robins, C. A. Abrey, T. P. Quinn, and D. E. Rogers. Lacustrine growth of juvenile pink salmon and a comparison with sympatric sockeye salmon. Journal of Fish Biology, 66(6):1671–1680, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Greig:2005:ICP

[1765] S. M. Greig, D. A. Sear, D. Smallman, and P. A. Carling. Impact of clay particles on the cutaneous exchange of oxygen across the chorion of Atlantic salmon eggs. Journal of Fish Biology, 66(6):1681–1691, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fey:2005:MOI

[1766] D. P. Fey. Is the marginal otolith increment width a reliable recent growth index for larval and juvenile herring? Journal of Fish Biology, 66(6):1692–1703, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2005:GMD

[1767] S. Xie, Y. Watanabe, T. Saruwatari, R. Masuda, Y. Yamashita, C. Sassa, and Y. Konishi. Growth and morphological development of sagittal otoliths of larval and early juvenile Trachurus japonicus. Journal of Fish Biology, 66(6):1704–1719, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2005:HDD

[1768] S. Xie and Y. Watanabe. Hatch date-dependent differences in early growth and development recorded in the otolith microstructure of Trachurus japonicus. Journal of Fish Biology, 66(6):1720–1734, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kano:2005:SFP

[1769] Y. Kano. Sexing fish by palpation: a simple method for gonadal assessment of fluvial salmonids. Journal of Fish Biology, 66(6):1735–1739, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalf:2005:NDA

[1770] S. J. Metcalf and S. E. Swearer. Non-destructive ageing in Notolabrus tetricus using dorsal spines with an emphasis on the benefits for protected, endangered and fished species. Journal of Fish Biology, 66(6): 1740–1747, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grutter:2005:DCS

[1771] A. S. Grutter, S. Glover, and R. Bshary. Does client size affect cleaner fish choice of client? An empirical test using client fish models. Journal of Fish Biology, 66(6):1748–1752, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:2005:BRE

[1772] Malcolm Clarke. Book review: The Economics of Fisheries Management. By L. G. Anderson. 2nd revised ed. xx 296 pp. Published by The Blackburn Press, Caldwell, New Jersey, 2004. Price $54.95. ISBN 1-930665-98-9. Journal of Fish Biology, 66(6):1753–1754, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Furness:2005:BRI

[1773] Robert W. Furness. Book review: Interactions between Fish and Birds: Implications for Management. Edited by I. G. Cowx. viii 376 pp. Published by Fishing News Books, Oxford, 2003. Price £89.50. ISBN 0-632-06385-8. Journal of Fish Biology, 66(6):1754–1755, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goldspink:2005:BRS

[1774] Christopher R. Goldspink. Book review: Salmon at the Edge. Edited by D. Mills. xi 307 pp. Published by Blackwell Science Ltd, Oxford, 2003. Price £85.00: ISBN 0-632-06457-9. Journal of Fish Biology, 66(6):1755– 1756, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2005:BRM

[1775] Paul J. B. Hart. Book review: Manual of Fish Sclerochronology. Edited by J. Panfili, H. de Pontual, H. Troadec and P. J. Wright. 463 pp and DVD. Published by IFREMER-IRD, Brest, France, 2002. Price EUR 120. ISBN 2-7099-1490-5. Journal of Fish Biology, 66(6):1756–1757, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:BRa

[1776] Anonymous. Books received. Journal of Fish Biology, 66(6):1758, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:TSB

[1777] Anonymous. Trematomus scotti in Beaver Lake: the first record of a fish from a non-marine Antarctic habitat. Journal of Fish Biology, 66(6): 1759, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:Eb

[1778] Anonymous. Erratum. Journal of Fish Biology, 66(6):1759, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:Ec

[1779] Anonymous. Erratum. Journal of Fish Biology, 66(6):1759, June 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skinner:2005:CVS

[1780] A. M. J. Skinner. Comments on ‘Variations of sperm release in three batches of zebra fish’. Journal of Fish Biology, 67(1):1–2, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Romero:2005:ERE

[1781] A. Romero and S. M. Green. The end of regressive evolution: examining and interpreting the evidence from cave fishes. Journal of Fish Biology, 67(1):3–32, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xue:2005:SDO

[1782] Y. Xue, X. Jin, B. Zhang, and Z. Liang. Seasonal, diel and ontogenetic variation in feeding patterns of small yellow croaker in the central Yellow Sea. Journal of Fish Biology, 67(1):33–50, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lloret:2005:HRD

[1783] J. Lloret, R. Galzin, L. Gil de Sola, A. Souplet, and M. Demestre. Habitat related differences in lipid reserves of some exploited fish species in the north-western Mediterranean continental shelf. Journal of Fish Biology, 67(1):51–65, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran-Lopez:2005:SHR

[1784] R. Morán-López, J. L. Pérez-Bote, E. Da Silva Rubio, and C. Corbacho Amado. Summer habitat relationships of barbels in south-west Spain. Journal of Fish Biology, 67(1):66–82, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heyman:2005:SAL

[1785] W. D. Heyman, B. Kjerfve, R. T. Graham, K. L. Rhodes, and L. Garbutt. Spawning aggregations of Lutjanus cyanopterus (Cuvier) on the Belize Barrier Reef over a 6 year period. Journal of Fish Biology, 67(1):83– 101, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monteiro:2005:DPR

[1786] N. M. Monteiro, S. M. Quinteira, K. Silva, M. N. Vieira, and V. C. Almada. Diet preference reflects the ontogenetic shift in microhabitat use in Lipophrys pholis. Journal of Fish Biology, 67(1):102–113, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wildhaber:2005:GIS

[1787] M. L. Wildhaber, D. M. Papoulias, A. J. DeLonay, D. E. Tillitt, J. L. Bryan, M. L. Annis, and J. A. Allert. Gender identification of shovelnose sturgeon using ultrasonic and endoscopic imagery and the application of the method to the pallid sturgeon. Journal of Fish Biology, 67(1):114– 132, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sanchez-Camara:2005:RCG

[1788] J. Sanchez-Camara, D. J. Booth, and X. Turon. Reproductive cycle and growth of Phyllopteryx taeniolatus. Journal of Fish Biology, 67(1):133– 148, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mello:2005:SGA

[1789] L. G. S. Mello and G. A. Rose. Seasonal growth of Atlantic cod: effects of temperature, feeding and reproduction. Journal of Fish Biology, 67 (1):149–170, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2005:EFL

[1790] S. J. Fu, X. J. Xie, and Z. D. Cao. Effect of feeding level and feeding frequency on specific dynamic action in Silurus meridionalis. Journal of Fish Biology, 67(1):171–181, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2005:RUS

[1791] A. P. Scott, M. L. Pinillos, and M. Huertas. The rate of uptake of sex steroids from water by Tinca tinca is influenced by their affinity for sex steroid binding protein in plasma. Journal of Fish Biology, 67(1):182– 200, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abe:2005:SMC

[1792] T. Abe and H. Munehara. Spawning and maternal-care behaviours of a copulating sculpin, Radulinopsis taranetzi. Journal of Fish Biology, 67(1):201–212, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Donaghy:2005:MCA

[1793] M. J. Donaghy, A. F. Youngson, and P. J. Bacon. Melanophore constellations allow robust individual identification of wild 0+ year Atlantic salmon. Journal of Fish Biology, 67(1):213–222, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morais:2005:DAP

[1794] S. Morais, C. R. Rojas-Garcia, L. E. C. Conceição, and I. Rønnestad. Digestion and absorption of a pure triacylglycerol and a free fatty acid by Clupea harengus L. larvae. Journal of Fish Biology, 67(1):223–238, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petersen:2005:FMS

[1795] C. W. Petersen, C. Mazzoldi, K. A. Zarrella, and R. E. Hale. Fertilization mode, sperm characteristics, mate choice and parental care patterns in Artedius spp. (Cottidae). Journal of Fish Biology, 67(1):239–254, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Power:2005:LVF

[1796] M. Power, J. B. Dempson, J. D. Reist, C. J. Schwarz, and G. Power. Latitudinal variation in fecundity among Arctic charr populations in eastern North America. Journal of Fish Biology, 67(1):255–273, July 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2005:ECF

[1797] P. He and C. S. Wardle. Effect of caudal fin height on swimming kinematics in the mackerel Scomber scombrus L. Journal of Fish Biology, 67(1):274–278, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2005:EFR

[1798] S. J. Fu, X. J. Xie, and Z. D. Cao. Effect of fasting on resting metabolic rate and postprandial metabolic response in Silurus meridionalis. Journal of Fish Biology, 67(1):279–285, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golubtsov:2005:FRP

[1799] A. S. Golubtsov, K. F. Dzerjinskii, and A. M. Prokofiev. Four rows of pharyngeal teeth in an aberrant specimen of the small African barb Barbus paludinosus (Cyprinidae): novelty or atavistic alteration? Journal of Fish Biology, 67(1):286–291, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Randle:2005:ABB

[1800] A. M. Randle and L. J. Chapman. Air-breathing behaviour of the African anabantoid fish Ctenopoma muriei. Journal of Fish Biology, 67(1):292– 298, July 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2005:REF

[1801] M. D. Taylor, P. J. Palmer, D. S. Fielder, and I. M. Suthers. Responsible estuarine finfish stock enhancement: an Australian perspective. Journal of Fish Biology, 67(2):299–331, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vreven:2005:RML

[1802] E. J. Vreven and G. G. Teugels. Redescription of Mastacembelus liberiensis Boulenger, 1898 and description of a new West African spiny-eel (Synbranchiformes: Mastacembelidae) from the Konkoure River basin, Guinea. Journal of Fish Biology, 67(2):332–369, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neer:2005:AGC

[1803] J. A. Neer, B. A. Thompson, and J. K. Carlson. Age and growth of Carcharhinus leucas in the northern Gulf of Mexico: incorporating variability in size at birth. Journal of Fish Biology, 67(2):370–383, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Artigas:2005:RHS

[1804] M. L. Artigas, J. E. Skjæraasen, A. C. Utne-Palm, and T. Nilsen. Recovery from handling stress in Gadus morhua. Journal of Fish Biology, 67(2):384–391, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castellanos-Galindo:2005:CSA

[1805] G. A. Castellanos-Galindo, A. Giraldo, and E. A. Rubio. Community structure of an assemblage of tidepool fishes on a tropical eastern Pacific rocky shore, Colombia. Journal of Fish Biology, 67(2):392–408, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wintzer:2005:CPC

[1806] A. P. Wintzer and P. J. Motta. A comparison of prey capture kinematics in hatchery and wild Micropterus salmoides floridanus: effects of ontogeny and experience. Journal of Fish Biology, 67(2):409–427, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davey:2005:SDM

[1807] A. J. H. Davey, S. J. Hawkins, G. F. Turner, and C. P. Doncaster. Sizedependent microhabitat use and intraspecific competition in Cottus gobio. Journal of Fish Biology, 67(2):428–443, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sarvas:2005:PPL

[1808] T. H. Sarvas and S. E. Fevolden. Pantophysin (Pan i) locus divergence between inshore v. offshore and northern v. southern populations of Atlantic cod in the north-east Atlantic. Journal of Fish Biology, 67(2): 444–469, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoltz:2005:AGS

[1809] J. A. Stoltz, B. D. Neff, and J. D. Olden. Allometric growth and sperm competition in fishes. Journal of Fish Biology, 67(2):470–480, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goncalves:2005:OGP

[1810] E. J. Gonçalves, R. Beldade, and M. Henriques. Opeatogenys gracilis (Pisces: Gobiesocidae): an overlooked species or another ‘Mediterranean endemism’ found in Atlantic waters? Journal of Fish Biology, 67(2):481– 489, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulbert:2005:SFM

[1811] L. B. Hulbert, A. M. Aires da Silva, V. F. Gallucci, and J. S. Rice. Seasonal foraging movements and migratory patterns of female Lamna ditropis tagged in Prince William Sound, Alaska. Journal of Fish Biology, 67(2):490–509, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakuma:2005:PVL

[1812] K. M. Sakuma, C. A. Taylor, and W. Watson. Pigment variability in larval Sebastes jordani off central California. Journal of Fish Biology, 67 (2):510–520, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez:2005:NCA

[1813] A. Rodríguez, E. Gisbert, and F. Castelló-Orvay. Nutritional condition of Anguilla anguilla starved at various salinities during the elver phase. Journal of Fish Biology, 67(2):521–534, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2005:BRF

[1814] T. Ellis, J. D. James, and A. P. Scott. Branchial release of free cortisol and melatonin by rainbow trout. Journal of Fish Biology, 67(2):535– 540, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buchheister:2005:SCL

[1815] A. Buchheister and M. T. Wilson. Shrinkage correction and length conversion equations for Theragra chalcogramma, Mallotus villosus and Thaleichthys pacificus. Journal of Fish Biology, 67(2):541–548, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peristeraki:2005:WCR

[1816] P. Peristeraki, G. Tserpes, and E. Lefkaditou. What cephalopod remains from Xiphias gladius stomachs can imply about predator–prey interactions in the Mediterranean Sea? Journal of Fish Biology, 67(2):549–554, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kramer:2005:RCD

[1817] C. R. Kramer, A. J. Kramer, and A. Konovalov. Rodlet cell distribution in the gall bladder epithelium of Fundulus heteroclitus. Journal of Fish Biology, 67(2):555–560, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ishiguro:2005:SSP

[1818] N. B. Ishiguro, M. Miya, J. G. Inoue, and M. Nishida. Sundasalanx (Sundasalangidae) is a progenetic clupeiform, not a closely-related group of salangids (Osmeriformes): mitogenomic evidence. Journal of Fish Biology, 67(2):561–569, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reid:2005:MAI

[1819] D. P. Reid, S. Pongsomboon, T. Jackson, C. McGowan, C. Murphy, D. Martin-Robichaud, and M. Reith. Microsatellite analysis indicates an absence of population structure among Hippoglossus hippoglossus in the north-west Atlantic. Journal of Fish Biology, 67(2):570–576, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruane:2005:SDR

[1820] N. M. Ruane, P. Makridis, P. H. M. Balm, and M. T. Dinis. Skin darkness is related to cortisol, but not MSH, content in post-larval Solea senegalensis. Journal of Fish Biology, 67(2):577–581, August 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barron-Vivanco:2005:ELE

[1821] B. Barrón-Vivanco, N. J. Duncan, N. García-Aguilar, J. Gutiérrez, and A. García-Gasca. Effect of LHRHa on the expression of stress-related molecules in the ovary of wild caught Sphoeroides annulatus held in captivity. Journal of Fish Biology, 67(2):582–588, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jadot:2005:IIT

[1822] C. Jadot, A. Donnay, M. Ylieff, and P. Poncin. Impact implantation of a transmitter on Sarpa salpa behaviour: study with a computerized video tracking system. Journal of Fish Biology, 67(2):589–595, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:Ed

[1823] Anonymous. Erratum. Journal of Fish Biology, 67(2):596, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:ISF

[1824] Anonymous. ICES Symposium on Fishing Technology in the 21^{st} century Integrating Fishing and Ecosystem Conservation 30 October–3 November 2006 The Seaport Hotel, Boston, Massachusetts, USA. Journal of Fish Biology, 67(2):597, August 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Daufresne:2005:DDP

[1825] M. Daufresne, H. Capra, and P. Gaudin. Downstream displacement of post-emergent brown trout: effects of development stage and water velocity. Journal of Fish Biology, 67(3):599–614, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gammon:2005:BRF

[1826] D. B. Gammon, W. Li, A. P. Scott, B. S. Zielinski, and L. D. Corkum. Behavioural responses of female Neogobius melanostomus to odours of conspecifics. Journal of Fish Biology, 67(3):615–626, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2005:EGD

[1827] P. J. Smith, S. M. McVeagh, and R. Allibone. Extensive genetic differentiation in Gobiomorphus breviceps from New Zealand. Journal of Fish Biology, 67(3):627–639, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pena:2005:EDF

[1828] R. Peña and S. Dumas. Effect of delayed first feeding on development and feeding ability of Paralabrax maculatofasciatus larvae. Journal of Fish Biology, 67(3):640–651, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mylonas:2005:IRT

[1829] C. C. Mylonas, L. Anezaki, P. Divanach, S. Zanuy, F. Piferrer, B. Ron, A. Peduel, I. Ben Atia, S. Gorshkov, and A. Tandler. Influence of rearing temperature during the larval and nursery periods on growth and sex differentiation in two Mediterranean strains of Dicentrarchus labrax. Journal of Fish Biology, 67(3):652–668, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vitale:2005:ETD

[1830] F. Vitale, M. Cardinale, and H. Svedäng. Evaluation of the temporal development of the ovaries in Gadus morhua from the Sound and Kattegat, North Sea. Journal of Fish Biology, 67(3):669–683, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawamura:2005:ETF

[1831] K. Kawamura and K. Uehara. Effects of temperature on free-embryonic diapause in the autumn-spawning bitterling Acheilognathus rhombeus (Teleostei: Cyprinidae). Journal of Fish Biology, 67(3):684–695, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Govoni:2005:OHM

[1832] J. J. Govoni, M. A. West, J. Bonaventura, G. Goddette, and T. E. Jenkins. The ontogeny of haematopoiesis in the marine teleost Leiostomus xanthurus and a comparison of the site of initial haematopoiesis with Opsanus tau. Journal of Fish Biology, 67(3):696–712, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freyhof:2005:HCO

[1833] J. Freyhof and C. Schöter. The houting Coregonus oxyrinchus (L.) (Salmoniformes: Coregonidae), a globally extinct species from the North Sea basin. Journal of Fish Biology, 67(3):713–729, September 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Trotter:2005:FII

[1834] A. J. Trotter, P. M. Pankhurst, and S. C. Battaglene. A finite interval of initial swimbladder inflation in Latris lineata revealed by sequential removal of water-surface films. Journal of Fish Biology, 67(3):730–741, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moura:2005:NDW

[1835] T. Moura, I. Figueiredo, P. Bordalo-Machado, C. Almeida, and L. S. Gordo. A new deep-water chimaerid species, Hydrolagus lusitanicus n. sp., from off mainland Portugal with a proposal of a new identification key for the genus Hydrolagus (Holocephali: Chimaeridae) in the northeast Atlantic. Journal of Fish Biology, 67(3):742–751, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lowe:2005:POG

[1836] C. J. Lowe and W. Davison. Plasma osmolarity, glucose concentration and erythrocyte responses of two Antarctic nototheniid fishes to acute and chronic thermal change. Journal of Fish Biology, 67(3):752–766, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zakhartsev:2005:GHS

[1837] M. Zakhartsev, B. De Wachter, T. Johansen, H. O. Pörtner, and R. Blust. Hsp70 is not a sensitive indicator of thermal limitation in Gadus morhua. Journal of Fish Biology, 67(3):767–778, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2005:EII

[1838] W. B. Wang, A. H. Li, T. Z. Cai, and J. G. Wang. Effects of intraperitoneal injection of cortisol on non-specific immune functions of Ctenopharyngodon idella. Journal of Fish Biology, 67(3):779–793, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simonsen:2005:WHA

[1839] V. Simonsen, M. M. Hansen, K.-L. D. Mensberg, R. I. Sarder, and S. Alam. Widespread hybridization among species of Indian major carps in hatcheries, but not in the wild. Journal of Fish Biology, 67(3):794–808, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bendall:2005:PSM

[1840] B. Bendall, A. Moore, and V. Quayle. The post-spawning movements of migratory brown trout Salmo trutta L. Journal of Fish Biology, 67(3): 809–822, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raine:2005:TTJ

[1841] J. C. Raine, U. Strelive, and J. F. Leatherland. The thyroid tissue of juvenile Oncorhynchus mykiss is tubular, not follicular. Journal of Fish Biology, 67(3):823–833, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2005:CPS

[1842] R. W. Blake, T. C. Law, K. H. S. Chan, and J. F. Z. Li. Comparison of the prolonged swimming performances of closely related, morphologically distinct three-spined sticklebacks Gasterosteus spp. Journal of Fish Biology, 67(3):834–848, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DAnatro:2005:GVA

[1843] A. D’Anatro and M. Loureiro. Geographic variation in Austrolebias luteoflamulatus Vaz-Ferreira, Sierra & Scaglia (Cyprinodontiformes, Rivulidae). Journal of Fish Biology, 67(3):849–865, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krause:2005:IDS

[1844] J. Krause, A. J. W. Ward, A. L. Jackson, G. D. Ruxton, R. James, and S. Currie. The influence of differential swimming speeds on composition of multi-species fish shoals. Journal of Fish Biology, 67(3):866–872, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heins:2005:CCT

[1845] D. C. Heins, J. A. Baker, G. De Silva, and E. L. Birden. Clutch characteristics of two populations of nine-spined stickleback from south-central Alaska. Journal of Fish Biology, 67(3):873–878, September 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Susnik:2005:GVN

[1846] S. Sušnik, J. Schöffmann, and S. Weiss. Genetic verification of native brown trout from the Persian Gulf (Catak Cay River, Tigris basin). Journal of Fish Biology, 67(3):879–884, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bauer:2005:PPE

[1847] C. Bauer, G. Unfer, and G. Loupal. Potential problems with external trailing antennae: antenna migration and ingrowth of epithelial tissue, a case study from a recaptured Chondrostoma nasus. Journal of Fish Biology, 67(3):885–889, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Avery:2005:IRA

[1848] T. S. Avery and J. A. Brown. Investigating the relationship among abnormal patterns of cell cleavage, egg mortality and early larval condition in Limanda ferruginea. Journal of Fish Biology, 67(3):890–896, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VanLiefferinge:2005:MUP

[1849] C. Van Liefferinge, P. Seeuws, P. Meire, and R. F. Verheyen. Microhabitat use and preferences of the endangered Cottus gobio in the River Voer, Belgium. Journal of Fish Biology, 67(4):897–909, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Deegan:2005:SPM

[1850] L. A. Deegan, H. E. Golden, J. Harrison, and K. Kracko. Swimming performance and metabolism of 0+ year Thymallus arcticus. Journal of Fish Biology, 67(4):910–918, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Finstad:2005:CUS

[1851] A. G. Finstad, F. Økland, E. B. Thorstad, and T. G. Heggberget. Comparing upriver spawning migration of Atlantic salmon Salmo salar and sea trout Salmo trutta. Journal of Fish Biology, 67(4):919–930, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aparicio:2005:BPP

[1852] E. Aparicio, E. García-Berthou, R. M. Araguas, P. Martínez, and J. L. García-Marín. Body pigmentation pattern to assess introgression by hatchery stocks in native Salmo trutta from Mediterranean streams. Journal of Fish Biology, 67(4):931–949, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tullis:2005:MBR

[1853] A. Tullis and M. Baillie. The metabolic and biochemical responses of tropical whitespotted bamboo shark Chiloscyllium plagiosum to alterations in environmental temperature. Journal of Fish Biology, 67(4): 950–968, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skarstein:2005:SCP

[1854] F. Skarstein, I. Folstad, and H. P. Rønning. Spawning colouration, parasites and habitat selection in Salvelinus alpinus: initiating speciation by sexual selection? Journal of Fish Biology, 67(4):969–980, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Popple:2005:MPC

[1855] I. D. Popple and W. Hunte. Movement patterns of Cephalopholis cruentata in a marine reserve in St Lucia, W.I., obtained from ultrasonic telemetry. Journal of Fish Biology, 67(4):981–992, October 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Briand:2005:SSM

[1856] C. Briand, D. Fatin, E. Ciccotti, and P. Lambert. A stage-structured model to predict the effect of temperature and salinity on glass eel Anguilla anguilla pigmentation development. Journal of Fish Biology, 67 (4):993–1009, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lauritzen:2005:KEW

[1857] D. V. Lauritzen, F. Hertel, and M. S. Gordon. A kinematic examination of wild sockeye salmon jumping up natural waterfalls. Journal of Fish Biology, 67(4):1010–1020, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguilar:2005:AMH

[1858] A. Aguilar and J. C. Garza. Analysis of major histocompatibility complex class II Beta genes from rockfishes (genus Sebastes). Journal of Fish Biology, 67(4):1021–1028, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almazan-Rueda:2005:PAG

[1859] P. Almazán-Rueda, A. T. M. Van Helmond, J. A. J. Verreth, and J. W. Schrama. Photoperiod affects growth, behaviour and stress variables in Clarias gariepinus. Journal of Fish Biology, 67(4):1029–1039, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Froufe:2005:CMS

[1860] E. Froufe, S. Alekseyev, I. Knizhin, and S. Weiss. Comparative mtDNA sequence (control region, ATPase 6 and NADH-1) divergence in Hucho taimen (Pallas) across four Siberian river basins. Journal of Fish Biology, 67(4):1040–1053, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hedger:2005:HSJ

[1861] R. D. Hedger, J. J. Dodson, N. E. Bergeron, and F. Caron. Habitat selection by juvenile Atlantic salmon: the interaction between physical habitat and abundance. Journal of Fish Biology, 67(4):1054–1071, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tobler:2005:CPC

[1862] M. Tobler, T. Wahli, and I. Schlupp. Comparison of parasite communities in native and introduced populations of sexual and asexual mollies of the genus Poecilia. Journal of Fish Biology, 67(4):1072–1082, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ehrenberg:2005:SAF

[1863] S. Z. Ehrenberg, S. Hansson, and R. Elmgren. Sublittoral abundance and food consumption of Baltic gobies. Journal of Fish Biology, 67(4):

1083–1093, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mai:2005:HSD

[1864] K. Mai, H. Yu, H. Ma, Q. Duan, E. Gisbert, J. L. Zambonino Infante, and C. L. Cahu. A histological study on the development of the digestive system of Pseudosciaena crocea larvae and juveniles. Journal of Fish Biology, 67(4):1094–1106, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shoji:2005:DRP

[1865] J. Shoji and M. Tanaka. Daily ration and prey size of juvenile piscivore Japanese Spanish mackerel. Journal of Fish Biology, 67(4):1107–1118, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turner:2005:EDF

[1866] J. P. Turner and J. R. Rooker. Effect of diet on fatty acid compositions in Sciaenops ocellatus. Journal of Fish Biology, 67(4):1119–1138, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hong:2005:PRT

[1867] Y. J. Hong, Z. J. Yu, L. Zhou, and J. F. Gui. A population of redtransparent, triploid Carassius auratus. Journal of Fish Biology, 67(4): 1139–1143, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Souissi:2005:OCF

[1868] J. Ben Souissi, D. Golani, H. Mejri, and C. Capapé. On the occurrence of cheilopogon furcatus in the Mediterranean Sea. Journal of Fish Biology, 67(4):1144–1149, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:2005:RLT

[1869] H. A. Campbell, C. M. Bishop, D. A. Davies, and S. Egginton. Recording long-term heart rate in Paranotothenia angustata using an electronic datalogger. Journal of Fish Biology, 67(4):1150–1156, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rinchard:2005:OFS

[1870] J. Rinchard, K. Dabrowski, J. J. Van Tassell, and R. A. Stein. Optimization of fertilization success in Sander vitreus is influenced by the sperm : egg ratio and ova storage. Journal of Fish Biology, 67(4):1157–1161, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnes:2005:TCS

[1871] M. E. Barnes, M. H. Zehfus, C. A. Soupir, and W. A. Sayler. Temporal changes in Sander vitreus egg thiamine levels. Journal of Fish Biology, 67 (4):1162–1168, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Binet:2005:GIM

[1872] M.-C. Binet and B. Angers. Genetic identification of members of the Phoxinus eos–neogaeus hybrid complex. Journal of Fish Biology, 67(4): 1169–1177, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2005:DIG

[1873] P. J. Smith, S. M. Mcveagh, V. Allain, and C. Sanchez. DNA identification of gut contents of large pelagic fishes. Journal of Fish Biology, 67 (4):1178–1183, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ojanguren:2005:MMI

[1874] A. F. Ojanguren, J. P. Evans, and A. E. Magurran. Multiple mating influences offspring size in guppies. Journal of Fish Biology, 67(4):1184– 1188, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bird:2005:BR

[1875] David Bird, Fred Utter, and Peter S. Maitland. Book reviews. Journal of Fish Biology, 67(4):1189–1192, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:BRb

[1876] Anonymous. Books received. Journal of Fish Biology, 67(4):1193, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:Ee

[1877] Anonymous. Erratum. Journal of Fish Biology, 67(4):1194, October 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skov:2005:EPT

[1878] C. Skov, J. Brodersen, C. Brönmark, L.-A. Hansson, P. Hertonsson, and P. A. Nilsson. Evaluation of PIT-tagging in cyprinids. Journal of Fish Biology, 67(5):1195–1201, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lower:2005:NIM

[1879] N. Lower, A. Moore, A. P. Scott, T. Ellis, J. D. James, and I. C. Russell. A non-invasive method to assess the impact of electronic tag insertion on stress levels in fishes. Journal of Fish Biology, 67(5):1202–1212, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nishi:2005:AJA

[1880] T. Nishi and G. Kawamura. Anguilla japonica is already magnetosensitive at the glass eel phase. Journal of Fish Biology, 67(5):1213–1224, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoneda:2005:EVT

[1881] M. Yoneda and P. J. Wright. Effects of varying temperature and food availability on growth and reproduction in first-time spawning female Atlantic cod. Journal of Fish Biology, 67(5):1225–1241, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:2005:MGA

[1882] W. S. Grant, R. W. Leslie, and B. W. Bowen. Molecular genetic assessment of bipolarity in the anchovy genus Engraulis. Journal of Fish Biology, 67(5):1242–1265, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward-Campbell:2005:MCR

[1883] B. M. S. Ward-Campbell, F. W. H. Beamish, and C. Kongchaiya. Morphological characteristics in relation to diet in five coexisting Thai fish species. Journal of Fish Biology, 67(5):1266–1279, November 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sundstrom:2005:HRR

[1884] L. F. Sundström, E. Petersson, J. I. Johnsson, J. Dannewitz, J. Höjesjö, and T. Järvi. Heart rate responses to predation risk in Salmo trutta are affected by the rearing environment. Journal of Fish Biology, 67(5):

1280–1286, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aguiar:2005:GCS

[1885] D. H. Aguiar, M. M. Barros, C. R. Padovani, L. E. Pezzato, and M. Dal Pai-Silva. Growth characteristics of skeletal muscle tissue in Oreochromis niloticus larvae fed on a lysine supplemented diet. Journal of Fish Biology, 67(5):1287–1298, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simonsen:2005:ODG

[1886] C. S. Simonsen and A. C. Gundersen. Ovary development in Greenland halibut Reinhardtius hippoglossoides in west Greenland waters. Journal of Fish Biology, 67(5):1299–1317, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruhl:2005:ESS

[1887] N. Ruhl and S. P. McRobert. The effect of sex and shoal size on shoaling behaviour in Danio rerio. Journal of Fish Biology, 67(5):1318–1326, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abdoli:2005:IFA

[1888] A. Abdoli, D. Pont, and P. Sagnes. Influence of female age, body size and environmental conditions on annual egg production of the bullhead. Journal of Fish Biology, 67(5):1327–1341, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2005:CNI

[1889] S. J. Cooke, G. T. Crossin, D. A. Patterson, K. K. English, S. G. Hinch, J. L. Young, R. F. Alexander, M. C. Healey, G. Van Der Kraak, and A. P. Farrell. Coupling non-invasive physiological assessments with telemetry to understand inter-individual variation in behaviour and survivorship of sockeye salmon: development and validation of a technique. Journal of Fish Biology, 67(5):1342–1358, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gilbey:2005:ILA

[1890] J. Gilbey, A. McLay, D. Houlihan, and E. Verspoor. Individual-level analysis of pre- and post first-feed growth and development in Atlantic salmon. Journal of Fish Biology, 67(5):1359–1369, November 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baum:2005:ETG

[1891] D. Baum, R. Laughton, J. D. Armstrong, and N. B. Metcalfe. The effect of temperature on growth and early maturation in a wild population of Atlantic salmon parr. Journal of Fish Biology, 67(5):1370–1380, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kemp:2005:SMS

[1892] P. S. Kemp, M. H. Gessel, and J. G. Williams. Seaward migrating subyearling Chinook salmon avoid overhead cover. Journal of Fish Biology, 67(5):1381–1391, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2005:MMG

[1893] N. G. Andersen and J. E. Beyer. Mechanistic modelling of gastric evacuation in predatory gadoids applying the square root model to describe surface-dependent evacuation. Journal of Fish Biology, 67(5):1392–1412, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2005:GEM

[1894] N. G. Andersen and J. E. Beyer. Gastric evacuation of mixed stomach contents in predatory gadoids: an expanded application of the square root model to estimate food rations. Journal of Fish Biology, 67(5): 1413–1433, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2005:FSS

[1895] R. W. Blake, K. H. S. Chan, and E. W. Y. Kwok. Finlets and the steady swimming performance of Thunnus albacares. Journal of Fish Biology, 67(5):1434–1445, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2005:ORA

[1896] H. B. Santos, E. Rizzo, N. Bazzoli, Y. Sato, and L. Moro. Ovarian regression and apoptosis in the South American teleost Leporinus taeniatus Lütken (Characiformes, Anostomidae) from the São Francisco Basin. Journal of Fish Biology, 67(5):1446–1459, November 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poulet:2005:PHU

[1897] N. Poulet, S. Lek, and C. Argillier. Pikeperch habitat use within a canal network in spring. Journal of Fish Biology, 67(5):1460–1474, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monteiro:2005:TPB

[1898] N. M. Monteiro, V. C. Almada, and M. N. Vieira. Temporal patterns of breeding and recruitment in Nerophis lumbriciformis (Pisces; Syngnathidae) related to seawater temperatures. Journal of Fish Biology, 67(5): 1475–1480, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moku:2005:VDO

[1899] M. Moku, A. Hayashi, K. Mori, and Y. Watanabe. Validation of daily otolith increment formation in the larval myctophid fish Diaphus slendertype spp. Journal of Fish Biology, 67(5):1481–1485, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2005:MEE

[1900] J. Simon and H. Dörner. Marking the European eel with oxytetracycline, alizarin red and coded wire tags: an evaluation of methods. Journal of Fish Biology, 67(5):1486–1491, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:WAM

[1901] Anonymous. Workshop on Advancements in Modelling Physicalbiological Interactions in Fish Early-life History: Recommended Practices and Future Directions: 3–5 April, 2006 in Nantes, France Co-Chairs: Alejandro Gallego (U.K.), Elizabeth North (U.S.A.) and Pierre Petitgas (France). Journal of Fish Biology, 67(5):1492, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:FSBa

[1902] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 67(5):1493, November 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:IVb

[1903] Anonymous. Index, volume 67. Journal of Fish Biology, 67(6):i–xxvi, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eddy:2005:AEE

[1904] F. B. Eddy. Ammonia in estuaries and effects on fish. Journal of Fish Biology, 67(6):1495–1513, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pottinger:2005:EBM

[1905] T. G. Pottinger, K. G. T. Pulman, T. R. Carrick, and A. P. Scott. Evaluation of biochemical methods for the non-destructive identification of sex in upstream migrating salmon and sea trout. Journal of Fish Biology, 67(6):1514–1533, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aoyama:2005:FOB

[1906] J. Aoyama, A. Shinoda, S. Sasai, M. J. Miller, and K. Tsukamoto. First observations of the burrows of Anguilla japonica. Journal of Fish Biology, 67(6):1534–1543, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ciuhandu:2005:EOG

[1907] C. S. Ciuhandu, E. D. Stevens, and P. A. Wright. The effect of oxygen on the growth of Oncorhynchus mykiss embryos with and without a chorion. Journal of Fish Biology, 67(6):1544–1551, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balcombe:2005:VFD

[1908] S. R. Balcombe, S. E. Bunn, F. J. McKenzie-Smith, and P. M. Davies. Variability of fish diets between dry and flood periods in an arid zone floodplain river. Journal of Fish Biology, 67(6):1552–1567, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ribeiro:2005:FAC

[1909] C. Ribeiro, A. J. Almeida, R. Araújo, M. Biscoito, and M. Freitas. Fish assemblages of Cais do Carvão Bay (Madeira Island) determined by the visual census technique. Journal of Fish Biology, 67(6):1568–1584, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strothotte:2005:SGW

[1910] E. Strothotte, G. J. Chaput, and H. Rosenthal. Seasonal growth of wild Atlantic salmon juveniles and implications on age at smoltification. Journal of Fish Biology, 67(6):1585–1602, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Timi:2005:PCG

[1911] J. T. Timi, J. L. Luque, and N. H. Sardella. Parasites of Cynoscion guatucupa along South American Atlantic coasts: evidence for stock discrimination. Journal of Fish Biology, 67(6):1603–1618, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2005:AGS

[1912] T. Ma, M. J. Miller, A. Shinoda, G. Minagawa, J. Aoyama, and K. Tsukamoto. Age and growth of Saurenchelys (Nettastomatidae) and Dysomma (Synaphobranchidae) leptocephali in the East China Sea. Journal of Fish Biology, 67(6):1619–1630, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aas-Hansen:2005:FMM

[1913] Ø. Aas-Hansen, E. H. Jørgensen, and M. M. Vijayan. Fasting modulates metabolic responses to cortisol, GH and IGF-I in Arctic charr hepatocytes. Journal of Fish Biology, 67(6):1631–1645, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharma:2005:USS

[1914] J. G. Sharma, R. Masuda, and M. Tanaka. Ultrastructural study of skin and eye of UV-B irradiated ayu Plecoglossus altivelis. Journal of Fish Biology, 67(6):1646–1652, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaur:2005:HSI

[1915] M. Kaur, F. Atif, M. Ali, H. Rehman, and S. Raisuddin. Heat stressinduced alterations of antioxidants in the freshwater fish Channa punctata Bloch. Journal of Fish Biology, 67(6):1653–1665, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

North:2005:MIS

[1916] A. W. North. Mackerel icefish size and age differences and long-term change at South Georgia and shag rocks. Journal of Fish Biology, 67(6): 1666–1685, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2005:CAA

[1917] F. C. P. Silva, M. F. G. Brito, L. M. Farias, and J. R. Nicoli. Composition and antagonistic activity of the indigenous intestinal microbiota of Prochilodus argenteus Agassiz. Journal of Fish Biology, 67(6):1686–1698, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merrett:2005:NGS

[1918] N. R. Merrett and J. A. Moore. A new genus and species of deep demersal fish (Teleostei: Stephanoberycidae) from the tropical eastern North Atlantic. Journal of Fish Biology, 67(6):1699–1710, December 2005. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefrancois:2005:EHL

[1919] C. Lefrançois, A. Shingles, and P. Domenici. The effect of hypoxia on locomotor performance and behaviour during escape in Liza aurata. Journal of Fish Biology, 67(6):1711–1729, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beyer:2005:PRI

[1920] K. Beyer, D. Kochanowska, M. Longshaw, S. W. Feist, and R. E. Gozlan. A potential role for invasive sunbleak in the further dissemination of a non-native parasite. Journal of Fish Biology, 67(6):1730–1733, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korner:2005:IFB

[1921] O. Körner, E. L. M. Vermeirssen, and P. Burkhardt-Holm. Intersex in feral brown trout from Swiss midland rivers. Journal of Fish Biology, 67(6):1734–1740, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irigoyen:2005:ODG

[1922] A. J. Irigoyen, D. E. Galván, and L. A. Venerus. Occurrence of dusky grouper Epinephelus marginatus (Lowe, 1834) in gulfs of northern Patagonia, Argentina. Journal of Fish Biology, 67(6):1741–1745, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castilho:2005:GDB

[1923] R. Castilho and Y. Ciftci. Genetic differentiation between close eastern Mediterranean Dicentrarchus labrax (L.) populations. Journal of Fish Biology, 67(6):1746–1752, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Copp:2005:BR

[1924] Gordon H. Copp. Book reviews. Journal of Fish Biology, 67(6):1753– 1754, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lorenzen:2005:BR

[1925] Kai Lorenzen. Book reviews. Journal of Fish Biology, 67(6):1754–1755, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lazar:2005:BR

[1926] Thomas Lazar. Book reviews. Journal of Fish Biology, 67(6):1755–1756, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:BRc

[1927] Anonymous. Books received. Journal of Fish Biology, 67(6):1757, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:2005:F

[1928] E. Verspoor. Foreword. Journal of Fish Biology, 67(s1):1–2, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:2005:PSA

[1929] E. Verspoor, J. A. Beardmore, S. Consuegra, C. García de Leániz, K. Hindar, W. C. Jordan, M.-L. Koljonen, A. A. Mahkrov, T. Paaver, J. A. Sánchez, Ø. Skaala, S. Titov, and T. F. Cross. Population structure in the Atlantic salmon: insights from 40 years of research into genetic protein variation. Journal of Fish Biology, 67(s1):3–54, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moller:2005:GSS

[1930] D. Møller. Genetic studies on serum transferrins in Atlantic salmon. Journal of Fish Biology, 67(s1):55–67, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makhrov:2005:ASC

[1931] A. A. Makhrov, E. Verspoor, V. S. Artamonova, and M. O’Sullivan. Atlantic salmon colonization of the Russian Arctic coast: pioneers from North America. Journal of Fish Biology, 67(s1):68–79, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:2005:RDN

[1932] E. Verspoor. Regional differentiation of North American Atlantic salmon at allozyme loci. Journal of Fish Biology, 67(s1):80–103, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cordes:2005:GAF

[1933] J. F. Cordes, D. L. Perkins, H. L. Kincaid, and B. May. Genetic analysis of fish genomes and populations: allozyme variation within and among Atlantic salmon from Downeast rivers of Maine. Journal of Fish Biology, 67(s1):104–117, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skaala:2005:GDB

[1934] Ø. Skaala, J. B. Taggart, and K. Gunnes. Genetic differences between five major domesticated strains of Atlantic salmon and wild salmon. Journal of Fish Biology, 67(s1):118–128, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Consuegra:2005:SEE

[1935] S. Consuegra, C. García De Leániz, A. Serdio, and E. Verspoor. Selective exploitation of early running fish may induce genetic and phenotypic changes in Atlantic salmon. Journal of Fish Biology, 67(s1):129–145, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:2005:AVA

[1936] W. C. Jordan, T. F. Cross, W. W. Crozier, A. Ferguson, P. Galvin, R. H. Hurrell, P. McGinnity, S. A. M. Martin, I. J. J. Moffett, D. J. Price, A. F. Youngson, and E. Verspoor. Allozyme variation in Atlantic salmon from the British Isles: associations with geography and the environment. Journal of Fish Biology, 67(s1):146–168, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blanco:2005:ATS

[1937] G. Blanco, M. D. Ramos, E. Vázquez, and J. A. Sánchez. Assessing temporal and spatial variation in wild populations of Atlantic salmon with particular reference to Asturias (Northern Spain) rivers. Journal of Fish Biology, 67(s1):169–184, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2005:SRP

[1938] P. Morán, J. Perez, J. Dumas, E. Beall, and E. García-Vázquez. Stockingrelated patterns of genetic variation at enzymatic loci in south European Atlantic salmon populations. Journal of Fish Biology, 67(s1):185–199, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:2005:GEL

[1939] E. Verspoor and L. J. Cole. Genetic evidence for lacustrine spawning of the non-anadromous Atlantic salmon population of Little Gull Lake, Newfoundland. Journal of Fish Biology, 67(s1):200–205, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran:2005:GVE

[1940] P. Morán, J. Pérez, and E. Garcia-Vazquez. Genetic variation at enzyme loci in the southernmost European populations of Atlantic salmon. Journal of Fish Biology, 67(s1):206–212, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verspoor:2005:ECD

[1941] E. Verspoor and C. D. Moyes. Evidence for co-dominant allelic expression of a phosphoglucomutase regulatory locus polymorphism in the Atlantic salmon. Journal of Fish Biology, 67(s1):213–218, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:RSa

[1942] Anonymous. Referees for this supplement. Journal of Fish Biology, 67 (s1):219, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:ISa

[1943] Anonymous. Index, 2005 supplement. Journal of Fish Biology, 67(s1): 220–222, September 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rice:2005:UFH

[1944] J. C. Rice. Understanding fish habitat ecology to achieve conservation. Journal of Fish Biology, 67(sB):1–22, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amilhat:2005:HUM

[1945] E. Amilhat and K. Lorenzen. Habitat use, migration pattern and population dynamics of chevron snakehead Channa striata in a rainfed rice farming landscape. Journal of Fish Biology, 67(sB):23–34, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Armstrong:2005:SVP

[1946] J. D. Armstrong. Spatial variation in population dynamics of juvenile Atlantic salmon: implications for conservation and management. Journal of Fish Biology, 67(sB):35–52, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brickhill:2005:FAA

[1947] M. J. Brickhill, S. Y. Lee, and R. M. Connolly. Fishes associated with artificial reefs: attributing changes to attraction or production using novel approaches. Journal of Fish Biology, 67(sB):53–71, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clavero:2005:SRD

[1948] M. Clavero, F. Blanco-Garrido, L. Zamora, and J. Prenda. Size-related and diel variations in microhabitat use of three endangered small fishes in a Mediterranean coastal stream. Journal of Fish Biology, 67(sB):72–85, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gozlan:2005:EOE

[1949] R. E. Gozlan and G. H. Copp. Early ontogenetic ecology of sofie Chondrostoma toxostoma: an integrated approach to organism-environment relationships. Journal of Fish Biology, 67(sB):86–99, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grant:2005:PDD

[1950] J. W. A. Grant and I. Imre. Patterns of density-dependent growth in juvenile stream-dwelling salmonids. Journal of Fish Biology, 67(sB): 100–110, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hiddink:2005:DHS

[1951] J. G. Hiddink, S. Jennings, and M. J. Kaiser. Do haddock select habitats to maximize condition? Journal of Fish Biology, 67(sB):111–124, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hinz:2005:FSD

[1952] H. Hinz, I. Kröncke, and S. Ehrich. The feeding strategy of dab Limanda limanda in the southern North Sea: linking stomach contents to prey availability in the environment. Journal of Fish Biology, 67(sB):125–145, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malavasi:2005:SWG

[1953] S. Malavasi, A. Franco, R. Fiorin, P. Franzoi, P. Torricelli, and D. Mainardi. The shallow water gobiid assemblage of the Venice lagoon: abundance, seasonal variation and habitat partitioning. Journal of Fish Biology, 67(sB):146–165, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Naiman:2005:PLF

[1954] R. J. Naiman and J. J. Latterell. Principles for linking fish habitat to fisheries management and conservation. Journal of Fish Biology, 67(sB): 166–185, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nislow:2005:FCS

[1955] K. H. Nislow. Forest change and stream fish habitat: lessons from ‘Olde’ and New England. Journal of Fish Biology, 67(sB):186–204, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinder:2005:OIS

[1956] A. C. Pinder, R. E. Gozlan, K. Beyer, and J. A. B. Bass. Ontogenetic induced shifts in the ecology of sunbleak Leucaspius delineatus during early development. Journal of Fish Biology, 67(sB):205–217, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sayer:2005:SBI

[1957] M. D. J. Sayer, S. H. Magill, T. J. Pitcher, L. Morissette, and C. Ainsworth. Simulation-based investigations of fishery changes as affected by the scale and design of artificial habitats. Journal of Fish Biology, 67(sB):218–243, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scruton:2005:CSH

[1958] D. A. Scruton, K. D. Clarke, M. M. Roberge, J. F. Kelly, and M. B. Dawe. A case study of habitat compensation to ameliorate impacts of hydroelectric development: effectiveness of re-watering and habitat enhancement of an intermittent flood overflow channel. Journal of Fish Biology, 67(sB):244–260, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:AOP

[1959] Anonymous. Abstracts of oral papers not published in the supplement. Journal of Fish Biology, 67(sB):261–278, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:APP

[1960] Anonymous. Abstracts of posters not published in the supplement. Journal of Fish Biology, 67(sB):279–294, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:RSb

[1961] Anonymous. Referees for this supplement. Journal of Fish Biology, 67 (sB):295, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:ISb

[1962] Anonymous. Index, 2005 supplement. Journal of Fish Biology, 67(sB): 297–300, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2005:FSBb

[1963] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10-14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 67(sB):301, December 2005. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:EJU

[1964] Anonymous. Ethical justification for the use and treatment of fishes in research. Journal of Fish Biology, 68(1):1–2, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crimmen:2006:AWO

[1965] Oliver Crimmen. Alwyne Wheeler 5 October 1929–19 June 2005. Journal of Fish Biology, 68(1):3–6, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordaan:2006:ITS

[1966] A. Jordaan, S. E. Hayhurst, and L. J. Kling. The influence of temperature on the stage at hatch of laboratory reared Gadus morhua and implications for comparisons of length and morphology. Journal of Fish Biology, 68(1):7–24, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vilhunen:2006:RAC

[1967] S. Vilhunen. Repeated antipredator conditioning: a pathway to habituation or to better avoidance? Journal of Fish Biology, 68(1):25–43, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henshel:2006:ASG

[1968] D. S. Henshel, D. W. Sparks, T. P. Simon, and M. J. Tosick. Age structure and growth of Semotilus atromaculatus (Mitchill) in PCBcontaminated streams. Journal of Fish Biology, 68(1):44–62, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kahilainen:2006:MDR

[1969] K. Kahilainen and K. Østbye. Morphological differentiation and resource polymorphism in three sympatric whitefish Coregonus lavaretus (L.) forms in a subarctic lake. Journal of Fish Biology, 68(1):63–79, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2006:ELH

[1970] K. Silva, N. M. Monteiro, V. C. Almada, and M. N. Vieira. Early life history of Syngnathus abaster. Journal of Fish Biology, 68(1):80–86, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hyvarinen:2006:DBT

[1971] P. Hyvärinen and A. Huusko. Diet of brown trout in relation to variation in abundance and size of pelagic fish prey. Journal of Fish Biology, 68 (1):87–98, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correia:2006:EOC

[1972] A. T. Correia, C. Antunes, J. M. Wilson, and J. Coimbra. An evaluation of the otolith characteristics of Conger conger during metamorphosis.

Journal of Fish Biology, 68(1):99–119, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Suski:2006:IET

[1973] C. D. Suski, S. S. Killen, J. D. Kieffer, and B. L. Tufts. The influence of environmental temperature and oxygen concentration on the recovery of largemouth bass from exercise: implications for live–release angling tournaments. Journal of Fish Biology, 68(1):120–136, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pouilly:2006:CTT

[1974] M. Pouilly, S. Barrera, and C. Rosales. Changes of taxonomic and trophic structure of fish assemblages along an environmental gradient in the Upper Beni watershed (Bolivia). Journal of Fish Biology, 68(1):137–156, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Erisman:2006:RBT

[1975] B. E. Erisman and L. G. Allen. Reproductive behaviour of a temperate serranid fish, Paralabrax clathratus (Girard), from Santa Catalina Island, California, U.S.A. Journal of Fish Biology, 68(1):157–184, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ito:2006:DMM

[1976] S. Ito and Y. Yanagisawa. Determinants of male mating success in a natural population of a stream goby of the genus Rhinogobius. Journal of Fish Biology, 68(1):185–195, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bervian:2006:SMV

[1977] G. Bervian, N. F. Fontoura, and M. Haimovici. Statistical model of variable allometric growth: otolith growth in Micropogonias furnieri (Actinopterygii, Sciaenidae). Journal of Fish Biology, 68(1):196–208, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morato:2006:VSF

[1978] T. Morato, W. W. L. Cheung, and T. J. Pitcher. Vulnerability of seamount fish to fishing: fuzzy analysis of life-history attributes. Journal of Fish Biology, 68(1):209–221, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott–Holland:2006:DAC

[1979] T. B. Scott-Holland, S. M. Bennett, and M. B. Bennett. Distribution of an asymmetrical copepod, Hatschekia plectropomi, on the gills of Plectropomus leopardus. Journal of Fish Biology, 68(1):222–235, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Komatsu:2006:SCL

[1980] T. Komatsu, S. Nakamura, and M. Nakamura. A sex cord-like structure and some remarkable features in early gonadal sex differentiation in the marine teleost Siganus guttatus (Bloch). Journal of Fish Biology, 68 (1):236–250, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tintos:2006:DMP

[1981] A. Tintos, J. M. Míguez, J. M. Mancera, and J. L. Soengas. Development of a microtitre plate indirect ELISA for measuring cortisol in teleosts, and evaluation of stress responses in rainbow trout and gilthead sea bream. Journal of Fish Biology, 68(1):251–263, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cuttitta:2006:FSR

[1982] A. Cuttitta, C. Guisande, I. Riveiro, I. Maneiro, B. Patti, A. R. Vergara, G. Basilone, A. Bonanno, and S. Mazzola. Factors structuring reproductive habitat suitability of Engraulis encrasicolus in the south coast of Sicily. Journal of Fish Biology, 68(1):264–275, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boylan:2006:IBS

[1983] P. Boylan and C. E. Adams. The influence of broad scale climatic phenomena on long term trends in Atlantic salmon population size: an example from the River Foyle, Ireland. Journal of Fish Biology, 68(1):276– 283, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dumas:2006:NME

[1984] J. Dumas and S. Marty. A new method to evaluate egg-to-fry survival in salmonids, trials with Atlantic salmon. Journal of Fish Biology, 68 (1):284–304, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raventos:2006:NSC

[1985] N. Raventos. Nest site characteristics and nesting success of the fivespotted wrasse Symphodus roissali in the north-western Mediterranean Sea. Journal of Fish Biology, 68(1):305–309, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takeuchi:2006:PSB

[1986] N. Takeuchi and K. Gushima. Promiscuous spawning behaviour of the tropical herring Spratelloides gracilis. Journal of Fish Biology, 68(1):310– 317, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lagrue:2006:ETP

[1987] C. Lagrue and L. Bollache. Effects of temperature on persistence times of native and invasive gammarid species in the stomachs of Cottus gobio. Journal of Fish Biology, 68(1):318–322, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:R

[1988] Anonymous. Referees 2005. Journal of Fish Biology, 68(1):323–327, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:ICC

[1989] Anonymous. International Conference on Coastal Ecosystems: Towards an integrated knowledge for an Ecosystem Approach for Fisheries: Campeche, Mexico, 26–29 June 2006. Journal of Fish Biology, 68(1): 328, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBa

[1990] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 68(1):329, January 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:GJF

[1991] Anonymous. Growth of the Journal of Fish Biology. Journal of Fish Biology, 68(2):331, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2006:CIF

[1992] F. A. Huntingford, C. Adams, V. A. Braithwaite, S. Kadri, T. G. Pottinger, P. Sandøe, and J. F. Turnbull. Current issues in fish welfare. Journal of Fish Biology, 68(2):332–372, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kidd:2006:DAF

[1993] M. R. Kidd, P. D. Danley, and T. D. Kocher. A direct assay of female choice in cichlids: all the eggs in one basket. Journal of Fish Biology, 68 (2):373–384, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Righton:2006:FEE

[1994] D. Righton, O. S. Kjesbu, and J. Metcalfe. A field and experimental evaluation of the effect of data storage tags on the growth of cod. Journal of Fish Biology, 68(2):385–400, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cal:2006:GGD

[1995] R. M. Cal, S. Vidal, P. Martínez, B. Álvarez-Blázquez, C. Gómez, and F. Piferrer. Growth and gonadal development of gynogenetic diploid Scophthalmus maximus. Journal of Fish Biology, 68(2):401–413, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reis:2006:SVS

[1996] R. E. Reis, E. Trajano, and E. Hingst-Zaher. Shape variation in surface and cave populations of the armoured catfishes Ancistrus (Siluriformes: Loricariidae) from the São Domingos karst area, upper Tocantins River, Brazil. Journal of Fish Biology, 68(2):414–429, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jokikokko:2006:SSW

[1997] E. Jokikokko, I. Kallio-Nyberg, I. Saloniemi, and E. Jutila. The survival of semi-wild, wild and hatchery-reared Atlantic salmon smolts of the Simojoki River in the Baltic Sea. Journal of Fish Biology, 68(2):430– 442, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunbrack:2006:MLG

[1998] R. Dunbrack, J. M. Green, and C. M. Mlewa. Marbled lungfish growth rates in Lake Baringo, Kenya, estimated by mark-recapture. Journal of Fish Biology, 68(2):443–449, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lovy:2006:UMS

[1999] J. Lovy, G. M. Wright, D. W. Wadowska, and D. J. Speare. Ultrastructural morphology suggesting a new hypothesis for development of microsporidians seen in Loma salmonae infecting the gills of rainbow and brook trout. Journal of Fish Biology, 68(2):450–457, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ristovska:2006:EDC

[2000] M. Ristovska, B. Karaman, W. Verraes, and D. Adriaens. Early development of the chondrocranium in Salmo letnica (Karaman, 1924) (Teleostei: Salmonidae). Journal of Fish Biology, 68(2):458–480, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paxton:2006:RVM

[2001] H. Paxton, R. H. C. Bonser, and K. Winwood. Regional variation in the microhardness and mineralization of vertebrae from brown and rainbow trout. Journal of Fish Biology, 68(2):481–487, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Licandeo:2006:AGS

[2002] R. R. Licandeo, J. G. Lamilla, P. G. Rubilar, and R. M. Vega. Age, growth, and sexual maturity of the yellownose skate Dipturus chilensis in the south-eastern Pacific. Journal of Fish Biology, 68(2):488–506, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skoglund:2006:FPD

[2003] H. Skoglund and B. T. Barlaup. Feeding pattern and diet of first feeding brown trout fry under natural conditions. Journal of Fish Biology, 68 (2):507–521, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Louette:2006:ACN

[2004] G. Louette and S. Declerck. Assessment and control of non-indigenous brown bullhead Ameiurus nebulosus populations using fyke nets in shallow ponds. Journal of Fish Biology, 68(2):522–531, February 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lisney:2006:BML

[2005] T. J. Lisney and S. P. Collin. Brain morphology in large pelagic fishes: a comparison between sharks and teleosts. Journal of Fish Biology, 68 (2):532–554, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heubel:2006:TAA

[2006] K. U. Heubel and I. Schlupp. Turbidity affects association behaviour in male Poecilia latipinna. Journal of Fish Biology, 68(2):555–568, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Islam:2006:DDR

[2007] M. S. Islam, M. Hibino, and M. Tanaka. Distribution and dietary relationships of the Japanese temperate bass Lateolabrax japonicus juveniles with two contrasting copepod assemblages in estuarine nursery grounds in the Ariake Sea, Japan. Journal of Fish Biology, 68(2):569–593, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitamura:2006:SCS

[2008] J. Kitamura. Seasonal change in the spatial utilization of host mussels in relation to ovipositor length by female rosy bitterling Rhodeus ocellatus kurumeus. Journal of Fish Biology, 68(2):594–607, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beldade:2006:FRC

[2009] R. Beldade, J. L. Van Tassell, and E. J. Gonçalves. First record of Chromogobius britoi (Teleostei: Gobiidae) on the mainland European coast. Journal of Fish Biology, 68(2):608–612, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2006:HND

[2010] R. A. Martin and E. Rekdal. ‘Hedgehog’: a novel defensive posture in juvenile Amblyraja radiata. Journal of Fish Biology, 68(2):613–617, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Auster:2006:AHE

[2011] P. J. Auster and K. Barber. Atlantic hagfish exploit prey captured by other taxa. Journal of Fish Biology, 68(2):618–621, February 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Summers:2006:RVI

[2012] D. W. Summers, D. E. Roberts, N. Giles, and D. N. Stubbing. Retention of visible implant and visible implant elastomer tags in brown trout in an English chalk stream. Journal of Fish Biology, 68(2):622–627, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goodwin:2006:FFL

[2013] C. E. Goodwin, D. Griffiths, J. T. A. Dick, and R. W. Elwood. A freshwater-feeding Lampetra fluviatilis L. population in Lough Neagh, Northern Ireland. Journal of Fish Biology, 68(2):628–633, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gagliano:2006:AMP

[2014] M. Gagliano, S. Kowalewsky, and M. I. McCormick. An alternative method for the preservation of tropical fish larvae. Journal of Fish Biology, 68(2):634–639, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaudron:2006:FEA

[2015] S. M. Gaudron and M. C. Lucas. First evidence of attraction of adult river lamprey in the migratory phase to larval odour. Journal of Fish Biology, 68(2):640–644, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffmayer:2006:SIV

[2016] E. R. Hoffmayer, G. R. Parsons, and J. Horton. Seasonal and interannual variation in the energetic condition of adult male Atlantic sharpnose shark Rhizoprionodon terraenovae in the northern Gulf of Mexico. Journal of Fish Biology, 68(2):645–653, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahnesjo:2006:BR

[2017] Ingrid Ahnesjö, Lynda Rodwell, David Bruno, and Audrey Geffen. Book reviews. Journal of Fish Biology, 68(2):654–657, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRa

[2018] Anonymous. Books received. Journal of Fish Biology, 68(2):658–659, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBb

[2019] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 68(2):660, February 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cole:2006:ORS

[2020] T. B. Cole and N. E. Stacey. Olfactory responses to steroids in an African mouth-brooding cichlid, Haplochromis burtoni (Günther). Journal of Fish Biology, 68(3):661–680, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Danylchuk:2006:NDL

[2021] A. J. Danylchuk and W. M. Tonn. Natural disturbance and life history: consequences of winterkill on fathead minnow in boreal lakes. Journal of Fish Biology, 68(3):681–694, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jackson:2006:DSP

[2022] M. W. Jackson, D. L. Nieland, and J. H. Cowan, Jr. Diel spawning periodicity of red snapper Lutjanus campechanus in the northern Gulf of Mexico. Journal of Fish Biology, 68(3):695–706, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curtis:2006:LHU

[2023] J. M. R. Curtis and A. C. J. Vincent. Life history of an unusual marine fish: survival, growth and movement patterns of Hippocampus guttulatus Cuvier 1829. Journal of Fish Biology, 68(3):707–733, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lekve:2006:IVA

[2024] K. Lekve, K. Enersen, S. E. Enersen, J. Gjøsæter, and N. Chr. Stenseth. Interannual variability in abundance and length of young coastal cod in the subtidal zone. Journal of Fish Biology, 68(3):734–746, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henry:2006:EIM

[2025] T. B. Henry and J. M. Grizzle. Electric-induced mortality of newly transformed juvenile fishes in waters of different conductivity. Journal of Fish Biology, 68(3):747–758, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bhat:2006:BFP

[2026] A. Bhat and A. E. Magurran. Benefits of familiarity persist after prolonged isolation in guppies. Journal of Fish Biology, 68(3):759–766, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Katano:2006:CFC

[2027] O. Katano, T. Nakamura, S. Abe, S. Yamamoto, and Y. Baba. Comparison of fish communities between above- and below-dam sections of small streams; barrier effect to diadromous fishes. Journal of Fish Biology, 68(3):767–782, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamashita:2006:GRF

[2028] S. Yamashita, T. Yamada, and T. J. Hara. Gustatory responses to feeding- and non-feeding-stimulant chemicals, with an emphasis on amino acids, in rainbow trout. Journal of Fish Biology, 68(3):783–800, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2006:ESS

[2029] H. A. Wright, R. J. Wootton, and I. Barber. The effect of Schistocephalus solidus infection on meal size of three-spined stickleback. Journal of Fish Biology, 68(3):801–809, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hara:2006:FBS

[2030] T. J. Hara. Feeding behaviour in some teleosts is triggered by single amino acids primarily through olfaction. Journal of Fish Biology, 68 (3):810–825, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoie:2006:ETG

[2031] H. Høie and A. Folkvord. Estimating the timing of growth rings in Atlantic cod otoliths using stable oxygen isotopes. Journal of Fish Biology, 68(3):826–837, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2006:PCR

[2032] H. Jensen, P.-A. Amundsen, J. M. Elliott, T. Bøhn, and P. E. Aspholm. Prey consumption rates and growth of piscivorous brown trout in a subarctic watercourse. Journal of Fish Biology, 68(3):838–848, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laptikhovsky:2006:DRP

[2033] V. Laptikhovsky, A. Arkhipkin, and P. Brickle. Distribution and reproduction of the Patagonian toothfish Dissostichus eleginoides Smitt around the Falkland Islands. Journal of Fish Biology, 68(3):849–861, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gerstner:2006:EFA

[2034] C. L. Gerstner, H. Ortega, H. Sanchez, and D. L. Graham. Effects of the freshwater aquarium trade on wild fish populations in differentiallyfished areas of the Peruvian Amazon. Journal of Fish Biology, 68(3): 862–875, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melia:2006:AGA

[2035] P. Melià, D. Bevacqua, A. J. Crivelli, G. A. De Leo, J. Panfili, and M. Gatto. Age and growth of Anguilla anguilla in the Camargue lagoons. Journal of Fish Biology, 68(3):876–890, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Swan:2006:EHS

[2036] S. C. Swan, A. J. Geffen, J. D. M. Gordon, B. Morales-Nin, and T. Shimmield. Effects of handling and storage methods on the concentrations of elements in deep-water fish otoliths. Journal of Fish Biology, 68(3):891– 904, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hurst:2006:TSV

[2037] T. P. Hurst and A. A. Abookire. Temporal and spatial variation in potential and realized growth rates of age-0 year northern rock sole. Journal of Fish Biology, 68(3):905–919, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schjolden:2006:DLA

[2038] J. Schjolden, K. G. T. Pulman, T. G. Pottinger, N. B. Metcalfe, and S. Winberg. Divergence in locomotor activity between two strains of rainbow trout Oncorhynchus mykiss with contrasting stress responsiveness. Journal of Fish Biology, 68(3):920–924, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krajewski:2006:PPG

[2039] J. P. Krajewski and R. M. Bonaldo. Plankton-picking by the goatfish Pseudupeneus maculatus (Mullidae), a specialized bottom forager. Journal of Fish Biology, 68(3):925–930, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohji:2006:DTA

[2040] M. Ohji, T. Arai, and N. Miyazaki. Differences of tributyltin accumulation in the masu salmon Oncorhynchus masou between sea-run and freshwater-resident types. Journal of Fish Biology, 68(3):931–939, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golani:2006:JTB

[2041] D. Golani and O. Sonin. The Japanese threadfin bream Nemipterus japonicus, a new Indo–Pacific fish in the Mediterranean Sea. Journal of Fish Biology, 68(3):940–943, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keefer:2006:LDD

[2042] M. L. Keefer, C. A. Peery, and C. C. Caudill. Long-distance downstream movements by homing adult Chinook salmon. Journal of Fish Biology, 68(3):944–950, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blanchet:2006:IHS

[2043] S. Blanchet, J. J. Dodson, and S. Brosse. Influence of habitat structure and fish density on Atlantic salmon Salmo salar L. territorial behaviour. Journal of Fish Biology, 68(3):951–957, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ah-King:2006:WTN

[2044] M. Ah-King, H. Elofsson, C. Kvarnemo, G. Rosenqvist, and A. Berglund. Why is there no sperm competition in a pipefish with externally brooding males? Insights from sperm activation and morphology. Journal of Fish Biology, 68(3):958–962, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2006:EAD

[2045] A. H. Rikardsen and E. B. Thorstad. External attachment of data storage tags increases probability of being recaptured in nets compared to internal tagging. Journal of Fish Biology, 68(3):963–968, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCormick:2006:PEP

[2046] M. I. McCormick and T. H. Holmes. Prey experience of predation influences mortality rates at settlement in a coral reef fish, Pomacentrus amboinensis. Journal of Fish Biology, 68(3):969–974, March 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pulgar:2006:IGS

[2047] J. M. Pulgar, F. P. Ojeda, and F. Bozinovic. Intraspecific geographic and seasonal physiological variability in an intertidal fish, Girella laevifrons, along a climatic gradient. Journal of Fish Biology, 68(3):975–981, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tirelli:2006:EDA

[2048] V. Tirelli, D. Borme, F. Tulli, M. Cigar, S. Fonda Umani, and S. B. Brandt. Energy density of anchovy Engraulis encrasicolus L. in the Adriatic Sea. Journal of Fish Biology, 68(3):982–989, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBc

[2049] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 68(3):990, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neave:2006:EPP

[2050] F. B. Neave, N. E. Mandrak, M. F. Docker, and D. L. Noakes. Effects of preservation on pigmentation and length measurements in larval lampreys. Journal of Fish Biology, 68(4):991–1001, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kolluru:2006:ECF

[2051] G. R. Kolluru, N. C. Ruiz, N. Del Cid, E. Dunlop, and G. F. Grether. The effects of carotenoid and food intake on caudal fin regeneration in male guppies. Journal of Fish Biology, 68(4):1002–1012, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frisch:2006:FES

[2052] A. Frisch and L. Van Herwerden. Field and experimental studies of hybridization between coral trouts, Plectropomus leopardus and Plectropomus maculatus (Serranidae), on the Great Barrier Reef, Australia.

Journal of Fish Biology, 68(4):1013–1025, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chara:2006:FEH

[2053] J. D. Chará, D. J. Baird, T. C. Telfer, and E. A. Rubio. Feeding ecology and habitat preferences of the catfish genus Trichomycterus in low-order streams of the Colombian Andes. Journal of Fish Biology, 68(4):1026– 1040, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tanaka:2006:FHG

[2054] H. Tanaka, I. Aoki, and S. Ohshimo. Feeding habits and gill raker morphology of three planktivorous pelagic fish species off the coast of northern and western Kyushu in summer. Journal of Fish Biology, 68(4): 1041–1061, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arnekleiv:2006:TSV

[2055] J. V. Arnekleiv, A. G. Finstad, and L. Rønning. Temporal and spatial variation in growth of juvenile Atlantic salmon. Journal of Fish Biology, 68(4):1062–1076, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaminski:2006:EDI

[2056] R. Kamiński, E. Kamler, M. Korwin-Kossakowski, L. Myszkowski, and J. Wolnicki. Effects of different incubation temperatures on the yolkfeeding stage of Eupallasella percnurus (Pallas). Journal of Fish Biology, 68(4):1077–1090, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reglero:2006:OMC

[2057] P. Reglero and H. Mosegaard. Onset of maturity and cohort composition at spawning of Baltic sprat Sprattus sprattus on the basis of otolith macrostructure analysis. Journal of Fish Biology, 68(4):1091–1106, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsland:2006:ETF

[2058] A. K. Imsland, A. Foss, L. O. Sparboe, and S. Sigurdsson. The effect of temperature and fish size on growth and feed efficiency ratio of juvenile spotted wolffish Anarhichas minor. Journal of Fish Biology, 68(4):1107– 1122, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2006:EIS

[2059] K. Hasegawa and K. Maekawa. The effects of introduced salmonids on two native stream-dwelling salmonids through interspecific competition. Journal of Fish Biology, 68(4):1123–1132, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillis:2006:CBS

[2060] J. A. Gillis, P. E. Witten, and B. K. Hall. Chondroid bone and secondary cartilage contribute to apical dentary growth in juvenile Atlantic salmon. Journal of Fish Biology, 68(4):1133–1143, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2006:CBT

[2061] X.-F. Garcia, M. Diekmann, U. Brämick, R. Lemcke, and T. Mehner. Correlations between type-indicator fish species and lake productivity in German lowland lakes. Journal of Fish Biology, 68(4):1144–1157, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gunnarsson:2006:GMF

[2062] Á. Gunnarsson, E. Hjörleifsson, K. Thórarinsson, and G. Marteinsdóttir. Growth, maturity and fecundity of wolffish Anarhichas lupus L. in Icelandic waters. Journal of Fish Biology, 68(4):1158–1176, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paglianti:2006:EST

[2063] A. Paglianti and P. Domenici. The effect of size on the timing of visually mediated escape behaviour in staghorn sculpin Leptocottus armatus. Journal of Fish Biology, 68(4):1177–1191, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watanabe:2006:USI

[2064] S. Watanabe, T. Isshiki, T. Kudo, A. Yamada, S. Katayama, and M. Fukuda. Using stable isotope ratios as a tracer of feeding adaptation in released Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 68(4):1192–1205, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bay:2006:IVP

[2065] L. K. Bay, K. Buechler, M. Gagliano, and M. J. Caley. Intraspecific variation in the pelagic larval duration of tropical reef fishes. Journal of Fish Biology, 68(4):1206–1214, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sasal:2006:NGD

[2066] P. Sasal. Nest guarding in a damselfish: evidence of a role for parasites. Journal of Fish Biology, 68(4):1215–1221, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niemela:2006:TVA

[2067] E. Niemelä, J. Erkinaro, M. Julkunen, E. Hassinen, M. Länsman, and S. Brørs. Temporal variation in abundance, return rate and life histories of previously spawned Atlantic salmon in a large subarctic river. Journal of Fish Biology, 68(4):1222–1240, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Podolska:2006:DBH

[2068] M. Podolska, J. Horbowy, and M. Wyszyński. Discrimination of Baltic herring populations with respect to Anisakis simplex larvae infection. Journal of Fish Biology, 68(4):1241–1256, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovac:2006:OVE

[2069] V. Kováč, S. Katina, G. H. Copp, and S. Siryová. Ontogenetic variability in external morphology and microhabitat use of spirlin Alburnoides bipunctatus from the River Rudava (Danube catchment). Journal of Fish Biology, 68(4):1257–1270, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Borcherding:2006:TDM

[2070] J. Borcherding, A. Scharbert, and R. Urbatzka. Timing of downstream migration and food uptake of juvenile North Sea houting stocked in the Lower Rhine and the Lippe (Germany). Journal of Fish Biology, 68 (4):1271–1286, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurenson:2006:NDE

[2071] C. H. Laurenson. A note on the development of the embryos of anglerfish Lophius piscatorius. Journal of Fish Biology, 68(4):1287–1290, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jordan:2006:PSA

[2072] R. Jordan, K. Kellogg, D. Howe, F. Juanes, J. Stauffer, Jr., and E. Loew. Photopigment spectral absorbance of Lake Malaŵi cichlids. Journal of Fish Biology, 68(4):1291–1299, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Damasceno-Oliveira:2006:ECC

[2073] A. Damasceno-Oliveira, B. Fernández-Durán, J. Gonçalves, P. Serrão, P. Soares-Da-Silva, M. A. Reis-Henriques, and J. Coimbra. Effects of cyclic and constant hydrostatic pressure on norepinephrine and epinephrine levels in the brain of flounder. Journal of Fish Biology, 68 (4):1300–1307, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:2006:BR

[2074] R. J. Wootton, Peter Miller, and Inigo Everson. Book reviews. Journal of Fish Biology, 68(4):1308–1311, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRb

[2075] Anonymous. Books received. Journal of Fish Biology, 68(4):1312, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:Ea

[2076] Anonymous. Erratum. Journal of Fish Biology, 68(4):1313–1316, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:ICG

[2077] Anonymous. 1st International Conference on the Grayling: Organised by The Grayling Research Trust, 11-13 July 2006, University of York, York, U.K. Journal of Fish Biology, 68(4):1317, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBd

[2078] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 68(4):1318, April 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Freitas:2006:EGJ

[2079] R. H. A. Freitas, R. S. Rosa, S. H. Gruber, and B. M. Wetherbee. Early growth and juvenile population structure of lemon sharks Negaprion brevirostris in the Atol das Rocas Biological Reserve, off north-east Brazil. Journal of Fish Biology, 68(5):1319–1332, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zimmer:2006:BTS

[2080] M. P. Zimmer and M. Power. Brown trout spawning habitat selection preferences and redd characteristics in the Credit River, Ontario. Journal of Fish Biology, 68(5):1333–1346, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arrington:2006:ILH

[2081] D. A. Arrington, B. K. Davidson, K. O. Winemiller, and C. A. Layman. Influence of life history and seasonal hydrology on lipid storage in three neotropical fish species. Journal of Fish Biology, 68(5):1347–1361, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawasaki:2006:LKP

[2082] M. Kawasaki, Y. Watanabe, N. Shirafuji, M. Chimura, M. Moku, O. Funaki, T. Saruwatari, and T. Kawamura. Larval Konosirus punctatus (Clupeidae) in a brackish river mouth on the Pacific coast of central Japan. Journal of Fish Biology, 68(5):1362–1375, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaushik:2006:FAP

[2083] S. J. Kaushik, G. Corraze, J. Radunz-Neto, L. Larroquet, and J. Dumas. Fatty acid profiles of wild brown trout and Atlantic salmon juveniles in the Nivelle basin. Journal of Fish Biology, 68(5):1376–1387, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arai:2006:MFB

[2084] T. Arai, J. Yang, and N. Miyazaki. Migration flexibility between freshwater and marine habitats of the pond smelt Hypomesus nipponensis. Journal of Fish Biology, 68(5):1388–1398, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulsey:2006:FSH

[2085] C. D. Hulsey, J. Marks, D. A. Hendrickson, C. A. Williamson, A. E. Cohen, and M. J. Stephens. Feeding specialization in Herichthys minckleyi: a trophically polymorphic fish. Journal of Fish Biology, 68(5):1399–1410, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bos:2006:ISM

[2086] A. R. Bos and R. Thiel. Influence of salinity on the migration of postlarval and juvenile flounder Pleuronectes flesus L. in a gradient experiment.

Journal of Fish Biology, 68(5):1411–1420, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaza:2006:PGE

[2087] G. Plaza, H. Honda, H. Sakaji, and K. Nashida. Patterns of growth in the early life history of the round herring Etrumeus teres. Journal of Fish Biology, 68(5):1421–1435, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watanabe:2006:SFH

[2088] H. Watanabe, T. Kubodera, and S. Kawahara. Summer feeding habits of the Pacific pomfret Brama japonica in the transitional and subarctic waters of the central North Pacific. Journal of Fish Biology, 68(5):1436– 1450, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoffmann:2006:SDP

[2089] E. Hoffmann and B. Borg. Sex differences in pectoral muscles but not in pectoral fins in the three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 68(5):1451–1459, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roy:2006:MNH

[2090] P. K. Roy and S. P. Lall. Mineral nutrition of haddock Melanogrammus aeglefinus (L.): a comparison of wild and cultured stock. Journal of Fish Biology, 68(5):1460–1472, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rogowski:2006:FHA

[2091] D. L. Rogowski, H. Reiser, and C. A. Stockwell. Fish habitat associations in a spatially variable desert stream. Journal of Fish Biology, 68(5):1473– 1483, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Balcombe:2006:DWC

[2092] S. R. Balcombe and P. Humphries. Diet of the western carp gudgeon (Hypseleotris klunzingeri Ogilby) in an Australian floodplain lake: the role of water level stability. Journal of Fish Biology, 68(5):1484–1493, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillette:2006:HUM

[2093] D. P. Gillette, J. S. Tiemann, D. R. Edds, and M. L. Wildhaber. Habitat use by a Midwestern U.S.A. riverine fish assemblage: effects of season, water temperature and river discharge. Journal of Fish Biology, 68(5): 1494–1512, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koizumi:2006:FBM

[2094] I. Koizumi, S. Yamamoto, and K. Maekawa. Female-biased migration of stream-dwelling Dolly Varden in the Shiisorapuchi River, Hokkaido, Japan. Journal of Fish Biology, 68(5):1513–1529, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welker:2006:RAN

[2095] T. L. Welker and D. L. Scarnecchia. River alteration and niche overlap among three native minnows (Cyprinidae) in the Missouri River hydrosystem. Journal of Fish Biology, 68(5):1530–1550, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fonseca:2006:GVJ

[2096] V. F. Fonseca, C. Vinagre, and H. N. Cabral. Growth variability of juvenile soles Solea solea and Solea senegalensis, and comparison with RNA : DNA ratios in the Tagus estuary, Portugal. Journal of Fish Biology, 68(5):1551–1562, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osmundson:2006:PCS

[2097] D. B. Osmundson. Proximate causes of sexual size dimorphism in Colorado pikeminnow, a long-lived cyprinid. Journal of Fish Biology, 68 (5):1563–1588, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Capper:2006:FPD

[2098] A. Capper, I. R. Tibbetts, J. M. O’Neil, and G. R. Shaw. Feeding preference and deterrence in rabbitfish Siganus fuscescens for the cyanobacterium Lyngbya majuscula in Moreton Bay, south-east Queensland, Australia. Journal of Fish Biology, 68(5):1589–1609, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hankison:2006:MDW

[2099] S. J. Hankison, M. J. Childress, J. J. Schmitter-Soto, and M. B. Ptacek. Morphological divergence within and between the Mexican sailfin mollies, Poecilia velifera and Poecilia petenensis. Journal of Fish Biology, 68(5): 1610–1630, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gherardi:2006:IWC

[2100] Francesca Gherardi. International Workshop ‘Chemical ecology in aquatic systems’, Florence, Italy, 16–18 October 2006. Journal of Fish Biology, 68(5):1631, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBe

[2101] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 68(5):1632, May 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:IVa

[2102] Anonymous. Index, volume 68. Journal of Fish Biology, 68(6):i–xxviii, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Limburg:2006:LW

[2103] K. Limburg, R. Gydemo, G. Aneer, L. Nyman, A. Nissling, P. Landergren, L. Vallin, and H. Wickström. Lars Westin 1934–2006. Journal of Fish Biology, 68(6):1633–1635, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barbosa:2006:FMD

[2104] M. Barbosa and A. E. Magurran. Female mating decisions: maximizing fitness? Journal of Fish Biology, 68(6):1636–1661, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henderson:2006:RBM

[2105] A. C. Henderson, J. L. McIlwain, H. S. Al-Oufi, and A. Ambu-Ali. Reproductive biology of the milk shark Rhizoprionodon acutus and the bigeye houndshark Iago omanensis in the coastal waters of Oman. Journal of Fish Biology, 68(6):1662–1678, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Able:2006:SFS

[2106] K. W. Able and J. C. Lamonaca. Scale formation in selected western North Atlantic flatfishes. Journal of Fish Biology, 68(6):1679–1692, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walsh:2006:VGL

[2107] C. T. Walsh, B. C. Pease, S. D. Hoyle, and D. J. Booth. Variability in growth of longfinned eels among coastal catchments of south-eastern Australia. Journal of Fish Biology, 68(6):1693–1706, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2006:SSS

[2108] R. E. Williams and R. J. G. Lester. Stock structure of Spanish mackerel Scomberomorus commerson along the Australian east coast deduced from parasite data. Journal of Fish Biology, 68(6):1707–1712, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quinn:2006:LTD

[2109] T. P. Quinn, P. McGinnity, and T. F. Cross. Long-term declines in body size and shifts in run timing of Atlantic salmon in Ireland. Journal of Fish Biology, 68(6):1713–1730, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goto:2006:DMH

[2110] A. Goto and T. Arai. Diverse migratory histories of Japanese Trachidermus and Cottus species (Cottidae) as inferred from otolith microchemistry. Journal of Fish Biology, 68(6):1731–1741, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peake:2006:FBR

[2111] S. J. Peake and A. P. Farrell. Fatigue is a behavioural response in respirometer-confined smallmouth bass. Journal of Fish Biology, 68(6): 1742–1755, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vaudo:2006:MPR

[2112] J. J. Vaudo and C. G. Lowe. Movement patterns of the round stingray Urobatis halleri (Cooper) near a thermal outfall. Journal of Fish Biology, 68(6):1756–1766, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Song:2006:SIE

[2113] J. Song, A. Mathieu, R. F. Soper, and A. N. Popper. Structure of the inner ear of bluefin tuna Thunnus thynnus. Journal of Fish Biology, 68 (6):1767–1781, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ross:2006:DFM

[2114] L. G. Ross, C. A. Martínez-Palacios, Ma. del C. Aguilar Valdez, M. C. M. Beveridge, and Ma. C. Chavez Sanchez. Determination of feeding mode in fishes: the importance of using structural and functional feeding studies in conjunction with gut analysis in a selective zooplanktivore Chirostoma estor estor Jordan 1880. Journal of Fish Biology, 68(6):1782–1794, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Glover:2006:TSS

[2115] K. A. Glover and Ø. Skaala. Temporal stability of sea louse Lepeophtheirus salmonis Krøyer populations on Atlantic salmon Salmo salar L. of wild, farm and hybrid parentage. Journal of Fish Biology, 68(6):1795– 1807, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marchetti:2006:HCH

[2116] L. Marchetti, M. Capacchietti, M. G. Sabbieti, D. Accili, G. Materazzi, and G. Menghi. Histology and carbohydrate histochemistry of the alimentary canal in the rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 68(6):1808–1821, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tavares-Dias:2006:MCS

[2117] M. Tavares-Dias. A morphological and cytochemical study of erythrocytes, thrombocytes and leukocytes in four freshwater teleosts. Journal of Fish Biology, 68(6):1822–1833, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pakhomov:2006:SDS

[2118] E. A. Pakhomov, T. Bushula, S. Kaehler, B. P. Watkins, and R. W. Leslie. Structure and distribution of the slope fish community in the vicinity of the sub-Antarctic Prince Edward Archipelago. Journal of Fish Biology, 68(6):1834–1866, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Megalofonou:2006:COG

[2119] P. Megalofonou. Comparison of otolith growth and morphology with somatic growth and age in young-of-the-year bluefin tuna. Journal of Fish Biology, 68(6):1867–1878, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gagnon:2006:SBS

[2120] M. M. Gagnon. Serum biliverdin as source of colouration upon sexual maturation in male blue-throated wrasse Notolabrus tetricus. Journal of Fish Biology, 68(6):1879–1882, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brian:2006:EDE

[2121] J. V. Brian, J. J. Augley, and V. A. Braithwaite. Endocrine disrupting effects on the nesting behaviour of male three-spined stickleback Gasterosteus aculeatus L. Journal of Fish Biology, 68(6):1883–1890, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cragg-Hine:2006:CSS

[2122] D. Cragg-Hine, D. C. Bradley, and K. Hendry. Changes in salmon smolt ages in the Welsh River Dee over a 66 year period. Journal of Fish Biology, 68(6):1891–1895, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wirtz:2006:SCP

[2123] S. Wirtz and P. Steinmann. Sperm characteristics in perch Perca fluviatilis L. Journal of Fish Biology, 68(6):1896–1902, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2006:CSS

[2124] E. I. Hasegawa. Comparison of the spectral sensitivity of three species of juvenile salmonids. Journal of Fish Biology, 68(6):1903–1908, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chittaro:2006:DSP

[2125] P. M. Chittaro, J. Gagnon, and B. J. Fryer. The differentiation of Stegastes partitus populations using lapillar and sagittal otolith chemistry. Journal of Fish Biology, 68(6):1909–1917, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johansen:2006:BPR

[2126] J. L. Johansen, N. A. Herbert, and J. F. Steffensen. The behavioural and physiological response of Atlantic cod Gadus morhua L. to shortterm acute hypoxia. Journal of Fish Biology, 68(6):1918–1924, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See erratum [2178].

Anonymous:2006:FSBf

[2127] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2006: Fish Population Structure: Implications to Conservation 10–14 July 2006 University of Aberdeen, Aberdeen, Scotland, UK Convenors: Peter Wright, Gudrun Marteinsdottir and Daniel Ruzzante. Journal of Fish Biology, 68(6):1925, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maclean:2006:E

[2128] Norman Maclean. Editorial. Journal of Fish Biology, 68(A):1, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Susnik:2006:GMC

[2129] S. Sušnik, I. Knizhin, A. Snoj, and S. Weiss. Genetic and morphological characterization of a Lake Ohrid endemic, Salmo (Acantholingua) ohridanus with a comparison to sympatric Salmo trutta. Journal of Fish Biology, 68(A):2–23, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chow:2006:PRA

[2130] S. Chow, T. Nakagawa, N. Suzuki, H. Takeyama, and T. Matsunaga. Phylogenetic relationships among Thunnus species inferred from rDNA ITS1 sequence. Journal of Fish Biology, 68(A):24–35, March 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bardakci:2006:PTB

[2131] F. Bardakci, N. Degerli, O. Ozdemir, and H. H. Basibuyuk. Phylogeography of the Turkish brown trout Salmo trutta L.: mitochondrial DNA PCR-RFLP variation. Journal of Fish Biology, 68(A):36–55, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2006:TSP

[2132] T. Takahashi, J. Snoeks, and K. Nakaya. Trematochromis schreyeni Poll, 1987, a junior synonym of ‘Ctenochromis’ benthicola (Matthes, 1962) (Perciformes: Cichlidae) from Lake Tanganyika. Journal of Fish Biology, 68(A):56–67, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2006:IHB

[2133] S. Yamamoto, S. Kitano, K. Maekawa, I. Koizumi, and K. Morita. Introgressive hybridization between Dolly Varden Salvelinus malma and white-spotted charr Salvelinus leucomaenis on Hokkaido Island, Japan. Journal of Fish Biology, 68(A):68–85, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bosi:2006:CIS

[2134] G. Bosi, A. P. Shinn, E. Simoni, S. Arrighi, and C. Domeneghini. A comparative immunohistochemical study on a galanin-like peptide in the neuroendocrine system of the alimentary canal of three species of siluriform catfishes. Journal of Fish Biology, 68(A):86–100, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ng:2006:IBT

[2135] H. H. Ng. The identity of Batasio tengana (Hamilton, 1822), with the description of two new species of Batasio from north-eastern India (Teleostei: Bagridae). Journal of Fish Biology, 68(A):101–118, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schulz:2006:EIO

[2136] M. Schulz, J. Freyhof, R. Saint-Laurent, K. Østbye, T. Mehner, and L. Bernatchez. Evidence for independent origin of two spring-spawning ciscoes (Salmoniformes: Coregonidae) in Germany. Journal of Fish Biology, 68(A):119–135, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalf:2006:SGM

[2137] V. J. Metcalf and N. J. Gemmell. Sexual genotype markers absent from small numbers of male New Zealand Oncorhynchus tshawytscha. Journal of Fish Biology, 68(A):136–143, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ardia:2006:NAI

[2138] D. R. Ardia and E. D. Clotfelter. The novel application of an immunological technique reveals the immunosuppressive effect of phytoestrogens in Betta splendens. Journal of Fish Biology, 68(A):144–149, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bianco:2006:OSR

[2139] P. G. Bianco, V. Zupo, and V. Ketmaier. Occurrence of the scalloped ribbonfish Zu cristatus (Lampridiformes) in coastal waters of the central Tyrrhenian Sea, Italy. Journal of Fish Biology, 68(A):150–155, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menezes:2006:GVS

[2140] M. R. Menezes, M. Ikeda, and N. Taniguchi. Genetic variation in skipjack tuna Katsuwonus pelamis (L.) using PCR–RFLP analysis of the mitochondrial DNA D-loop region. Journal of Fish Biology, 68(A):156– 161, March 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yuan:2006:ASN

[2141] L. Y. Yuan, Z. Q. Wu, and E. Zhang. Acrossocheilus spinifer, a new species of barred cyprinid fish from south China (Pisces: Teleostei). Journal of Fish Biology, 68(SB):163–173, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2006:DNL

[2142] T. Takahashi and M. Hori. Description of a new Lake Tanganyikan cichlid fish of the genus Cyprichromis (Perciformes: Cichlidae) with a note on sexual dimorphism. Journal of Fish Biology, 68(SB):174–192, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Launey:2006:MGV

[2143] S. Launey, J. Morin, S. Minery, and J. Laroche. Microsatellite genetic variation reveals extensive introgression between wild and introduced stocks, and a new evolutionary unit in French pike Esox lucius L. Journal of Fish Biology, 68(SB):193–216, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salini:2006:GPS

[2144] J. P. Salini, J. R. Ovenden, R. Street, R. Pendrey, Haryanti, and Ngurah. Genetic population structure of red snappers (Lutjanus malabaricus Bloch & Schneider, 1801 and Lutjanus erythropterus Bloch, 1790) in central and eastern Indonesia and northern Australia. Journal of Fish Biology, 68(SB):217–234, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cinquetti:2006:ILH

[2145] R. Cinquetti and L. Dramis. Identification and localization of hormoneproducing cells in the pituitary of male Padogobius martensi (Pisces, Gobiidae): a histochemical and immunocytochemical study. Journal of Fish Biology, 68(SB):235–250, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamazaki:2006:TMP

[2146] Y. Yamazaki, R. Yokoyama, M. Nishida, and A. Goto. Taxonomy and molecular phylogeny of Lethenteron lampreys in eastern Eurasia. Journal of Fish Biology, 68(SB):251–269, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hughes:2006:MDA

[2147] J. M. Hughes and M. J. Hillyer. Mitochondrial DNA and allozymes reveal high dispersal abilities and historical movement across drainage boundaries in two species of freshwater fishes from inland rivers in Queensland, Australia. Journal of Fish Biology, 68(SB):270–291, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sousa-Santos:2006:EEM

[2148] C. Sousa-Santos, M. J. Collares-Pereira, and V. C. Almada. Evidence of extensive mitochondrial introgression with nearly complete substitution of the typical Squalius pyrenaicus-like mtDNA of the Squalius alburnoides complex (Cyprinidae) in an independent Iberian drainage. Journal of Fish Biology, 68(SB):292–301, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:IVS

[2149] Anonymous. Index, volume 68 supplements. Journal of Fish Biology, 68 (SB):i–v, June 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mank:2006:ERG

[2150] J. E. Mank and J. C. Avise. The evolution of reproductive and genomic diversity in ray-finned fishes: insights from phylogeny and comparative analysis. Journal of Fish Biology, 69(1):1–27, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jimenez-Rosenberg:2006:LPJ

[2151] S. P. A. Jiménez-Rosenberg, E. A. González-Navarro, and R. J. Saldierna-Martínez. Larval, prejuvenile and juvenile development of Eucinostomus currani. Journal of Fish Biology, 69(1):28–37, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pothin:2006:DBM

[2152] K. Pothin, C. Gonzalez-Salas, P. Chabanet, and R. Lecomte-Finiger. Distinction between Mulloidichthys flavolineatus juveniles from Reunion Island and Mauritius Island (south-west Indian Ocean) based on otolith morphometrics. Journal of Fish Biology, 69(1):38–53, July 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Korner:2006:SSM

[2153] K. E. Körner, I. Schlupp, M. Plath, and E. R. Loew. Spectral sensitivity of mollies: comparing surface- and cave-dwelling Atlantic mollies, Poecilia mexicana. Journal of Fish Biology, 69(1):54–65, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monteiro:2006:RMS

[2154] N. M. Monteiro, A. Berglund, M. N. Vieira, and V. C. Almada. Reproductive migrations of the sex role reversed pipefish Nerophis lumbriciformis (Pisces; Syngnathidae). Journal of Fish Biology, 69(1):66–74, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weber:2006:DDE

[2155] D. N. Weber. Dose-dependent effects of developmental mercury exposure on C-start escape responses of larval zebrafish Danio rerio. Journal of Fish Biology, 69(1):75–94, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Panfili:2006:ISS

[2156] J. Panfili, D. Thior, J.-M. Ecoutin, P. Ndiaye, and J.-J. Albaret. Influence of salinity on the size at maturity for fish species reproducing in contrasting West African estuaries. Journal of Fish Biology, 69(1): 95–113, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eriksen:2006:PSF

[2157] M. S. Eriksen, M. Bakken, Å. Espmark, B. O. Braastad, and R. Salte. Prespawning stress in farmed Atlantic salmon Salmo salar: maternal cortisol exposure and hyperthermia during embryonic development affect offspring survival, growth and incidence of malformations. Journal of Fish Biology, 69(1):114–129, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papoutsoglou:2006:DEAa

[2158] E. S. Papoutsoglou and A. R. Lyndon. Digestive enzymes along the alimentary tract of the parrotfish Sparisoma cretense. Journal of Fish Biology, 69(1):130–140, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quevedo:2006:ESS

[2159] M. Quevedo and J. Olsson. The effect of small-scale resource origin on trophic position estimates in Perca fluviatilis. Journal of Fish Biology, 69(1):141–150, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazur:2006:LPS

[2160] M. M. Mazur and D. A. Beauchamp. Linking piscivory to spatial– temporal distributions of pelagic prey fishes with a visual foraging model. Journal of Fish Biology, 69(1):151–175, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mackie:2006:ACG

[2161] M. C. Mackie. Anatomical changes to the gonad during protogynous sex change in the half-moon grouper Epinephelus rivulatus (Valenciennes). Journal of Fish Biology, 69(1):176–186, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slotte:2006:ECB

[2162] A. Slotte, N. Mikkelsen, and H. Gjøsæter. Egg cannibalism in Barents Sea capelin in relation to a narrow spawning distribution. Journal of Fish Biology, 69(1):187–202, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shibata:2006:SSR

[2163] J. Shibata and M. Kohda. Seasonal sex role changes in the blenniid Petroscirtes breviceps, a nest brooder with paternal care. Journal of Fish Biology, 69(1):203–214, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jardine:2006:UPI

[2164] T. D. Jardine and R. A. Curry. Unique perspectives on the influence of size and age on consumer δ ^{15}N from a rainbow smelt complex. Journal of Fish Biology, 69(1):215–223, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibson:2006:OBR

[2165] A. K. Gibson and A. Mathis. Opercular beat rate for rainbow darters Etheostoma caeruleum exposed to chemical stimuli from conspecific and heterospecific fishes. Journal of Fish Biology, 69(1):224–232, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Osorio:2006:TDB

[2166] R. Osório, I. L. Rosa, and H. Cabral. Territorial defence by the Brazilian damsel Stegastes fuscus (Teleostei: Pomacentridae). Journal of Fish Biology, 69(1):233–242, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Platell:2006:IIP

[2167] M. E. Platell, P. A. Orr, and I. C. Potter. Inter- and intraspecific partitioning of food resources by six large and abundant fish species in a seasonally open estuary. Journal of Fish Biology, 69(1):243–262, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitamura:2006:ASU

[2168] J. Kitamura. Adaptive spatial utilization of host mussels by the Japanese rosy bitterling Rhodeus ocellatus kurumeus. Journal of Fish Biology, 69 (1):263–271, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petry:2006:LCI

[2169] A. C. Petry and U. H. Schulz. Longitudinal changes and indicator species of the fish fauna in the subtropical Sinos River, Brazil. Journal of Fish Biology, 69(1):272–290, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karakatsouli:2006:GAD

[2170] N. Karakatsouli, K. Tarnaris, C. Balaskas, and S. E. Papoutsoglou. Gill area and dimensions of gilthead sea bream Sparus aurata L. Journal of Fish Biology, 69(1):291–299, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dulcic:2006:FRC

[2171] J. Dulčić and D. Golani. First record of Cyclopterus lumpus L., 1758 (Osteichthyes: Cyclopteridae) in the Mediterranean Sea. Journal of Fish Biology, 69(1):300–303, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rivinoja:2006:MSM

[2172] P. Rivinoja, K. Leonardsson, and H. Lundqvist. Migration success and migration time of gastrically radio-tagged v. PIT-tagged adult Atlantic salmon. Journal of Fish Biology, 69(1):304–311, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tavares-Dias:2006:CMS

[2173] M. Tavares-Dias. Cytochemical method for staining fish basophils. Journal of Fish Biology, 69(1):312–317, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mancini:2006:ECT

[2174] P. L. Mancini and A. F. Amorim. Embryos of common thresher shark Alopias vulpinus in southern Brazil, South Atlantic Ocean. Journal of Fish Biology, 69(1):318–321, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRc

[2175] Anonymous. Book reviews. Journal of Fish Biology, 69(1):322–323, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRd

[2176] Anonymous. Books received. Journal of Fish Biology, 69(1):326–327, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:Eb

[2177] Anonymous. Erratum. Journal of Fish Biology, 69(1):328, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welker:2006:E

[2178] T. L. Welker and D. L. Scarnecchia. Erratum. Journal of Fish Biology, 69 (1):329, July 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [2126].

Dembski:2006:CET

[2179] S. Dembski, G. Masson, D. Monnier, P. Wagner, and J. C. Pihan. Consequences of elevated temperatures on life-history traits of an introduced fish, pumpkinseed Lepomis gibbosus. Journal of Fish Biology, 69(2):331– 346, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zaboukas:2006:BCA

[2180] N. Zaboukas, H. Miliou, P. Megalofonou, and M. Moraitou-Apostolopoulou. Biochemical composition of the Atlantic bonito Sarda sarda from the Aegean Sea (eastern Mediterranean Sea) in different stages of sexual maturity. Journal of Fish Biology, 69(2):347–362, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopes:2006:ESZ

[2181] M. Lopes, A. G. Murta, and H. N. Cabral. The ecological significance of the zooplanktivores, snipefish Macroramphosus spp. and boarfish Capros aper, in the food web of the south-east North Atlantic. Journal of Fish Biology, 69(2):363–378, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robins:2006:VGR

[2182] J. Robins, D. Mayer, J. Staunton-Smith, I. Halliday, B. Sawynok, and M. Sellin. Variable growth rates of the tropical estuarine fish barramundi Lates calcarifer (Bloch) under different freshwater flow conditions. Journal of Fish Biology, 69(2):379–391, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sigler:2006:DPS

[2183] M. F. Sigler, L. B. Hulbert, C. R. Lunsford, N. H. Thompson, K. Burek, G. O’Corry-Crowe, and A. C. Hirons. Diet of Pacific sleeper shark, a potential Steller sea lion predator, in the north-east Pacific Ocean. Journal of Fish Biology, 69(2):392–405, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hulbert:2006:DMB

[2184] L. B. Hulbert, M. F. Sigler, and C. R. Lunsford. Depth and movement behaviour of the Pacific sleeper shark in the north-east Pacific Ocean. Journal of Fish Biology, 69(2):406–425, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Choo:2006:MDA

[2185] C. K. Choo and H. C. Liew. Morphological development and allometric growth patterns in the juvenile seahorse Hippocampus kuda Bleeker. Journal of Fish Biology, 69(2):426–445, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papoutsoglou:2006:DEAb

[2186] E. S. Papoutsoglou and A. R. Lyndon. Digestive enzymes of Anarhichas minor and the effect of diet composition on their performance. Journal of Fish Biology, 69(2):446–460, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Persson:2006:TSS

[2187] L. Persson, A. Bertolo, and A. M. De Roos. Temporal stability in size distributions and growth rates of three Esox lucius L. populations. A result of cannibalism? Journal of Fish Biology, 69(2):461–472, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Congleton:2006:BCI

[2188] J. L. Congleton and T. Wagner. Blood-chemistry indicators of nutritional status in juvenile salmonids. Journal of Fish Biology, 69(2):473– 490, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parmentier:2006:SPT

[2189] E. Parmentier, P. Vandewalle, B. Frédérich, and M. L. Fine. Sound production in two species of damselfishes (Pomacentridae): Plectroglyphidodon lacrymatus and Dascyllus aruanus. Journal of Fish Biology, 69 (2):491–503, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stubhaug:2006:OCL

[2190] I. Stubhaug, Ø. Lie, and B. E. Torstensen. β-oxidation capacity in liver increases during parr-smolt transformation of Atlantic salmon fed vegetable oil and fish oil. Journal of Fish Biology, 69(2):504–517, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillet:2006:ETC

[2191] C. Gillet and P. QuéTin. Effect of temperature changes on the reproductive cycle of roach in Lake Geneva from 1983 to 2001. Journal of Fish Biology, 69(2):518–534, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blann:2006:ESC

[2192] C. A. Blann and M. C. Healey. Effects of species, culture history, size and residency on relative competitive ability of salmonids. Journal of Fish Biology, 69(2):535–552, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McClure:2006:NND

[2193] M. M. McClure, P. B. McIntyre, and A. R. McCune. Notes on the natural diet and habitat of eight danionin fishes, including the zebrafish Danio rerio. Journal of Fish Biology, 69(2):553–570, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Herder:2006:RPT

[2194] F. Herder and J. Freyhof. Resource partitioning in a tropical stream fish assemblage. Journal of Fish Biology, 69(2):571–589, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giari:2006:CCR

[2195] L. Giari, M. Manera, E. Simoni, and B. S. Dezfuli. Changes to chloride and rodlet cells in gills, kidney and intestine of Dicentrarchus labrax (L.) exposed to reduced salinities. Journal of Fish Biology, 69(2):590– 600, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomas:2006:AAG

[2196] J. Tomás. The appearance of accessory growth centres in adult whiting Merlangius merlangus otoliths. Journal of Fish Biology, 69(2):601–607, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chagas:2006:AAR

[2197] E. C. Chagas and A. L. Val. Ascorbic acid reduces the effects of hypoxia on the Amazon fish tambaqui. Journal of Fish Biology, 69(2):608–612, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bystriansky:2006:AAC

[2198] J. S. Bystriansky, P. J. LeBlanc, and J. S. Ballantyne. Anaesthetization of Arctic charr Salvelinus alpinus (L.) with tricaine methanesulphonate or 2-phenoxyethanol for immediate blood sampling. Journal of Fish Biology, 69(2):613–621, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Genner:2006:TRH

[2199] M. J. Genner, A. Botha, and G. F. Turner. Translocations of rocky habitat cichlid fishes to Nkhata Bay, Lake Malaŵi. Journal of Fish Biology, 69(2):622–628, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martins:2006:CFI

[2200] C. I. M. Martins, M. Trenovski, J. W. Schrama, and J. A. J. Verreth. Comparison of feed intake behaviour and stress response in isolated and non-isolated African catfish. Journal of Fish Biology, 69(2):629–636, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sreeraj:2006:DHB

[2201] N. Sreeraj, R. Raghavan, and G. Prasad. The diet of Horabagrus brachysoma (Günther), an endangered bagrid catfish from Lake Vembanad (South India). Journal of Fish Biology, 69(2):637–642, August 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Elsdon:2006:IMC

[2202] T. S. Elsdon and B. M. Gillanders. Identifying migratory contingents of fish by combining otolith Sr:Ca with temporal collections of ambient Sr:Ca concentrations. Journal of Fish Biology, 69(3):643–657, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baldisserotto:2006:VAI

[2203] B. Baldisserotto, M. J. Chowdhury, and C. M. Wood. In vitro analysis of intestinal absorption of cadmium and calcium in rainbow trout fed with calcium- and cadmium-supplemented diets. Journal of Fish Biology, 69 (3):658–667, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baker:2006:UPI

[2204] C. F. Baker and J. A. T. Boubée. Upstream passage of inanga Galaxias maculatus and redfin bullies Gobiomorphus huttoni over artificial ramps. Journal of Fish Biology, 69(3):668–681, September 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuya:2006:CDW

[2205] F. Tuya, P. Sanchez-Jerez, T. Dempster, A. Boyra, and R. J. Haroun. Changes in demersal wild fish aggregations beneath a sea-cage fish farm after the cessation of farming. Journal of Fish Biology, 69(3):682–697, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nobah:2006:CPC

[2206] C. S. K. Nobah, E. P. Kouamelan, V. N’Douba, J. Snoeks, G. G. Teugels, G. Goore-Bi, T. Kone, and T. M. Falk. The colour pattern of the caudal fin, a useful criterion for identification of two species of Tilapia and their hybrids. Journal of Fish Biology, 69(3):698–707, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ashford:2006:CCO

[2207] J. R. Ashford, A. I. Arkhipkin, and C. M. Jones. Can the chemistry of otolith nuclei determine population structure of Patagonian toothfish Dissostichus eleginoides? Journal of Fish Biology, 69(3):708–721, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schaefer:2006:DPT

[2208] J. Schaefer and A. Ryan. Developmental plasticity in the thermal tolerance of zebrafish Danio rerio. Journal of Fish Biology, 69(3):722–734, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sonny:2006:RCI

[2209] D. Sonny, F. R. Knudsen, P. S. Enger, T. Kvernstuen, and O. Sand. Reactions of cyprinids to infrasound in a lake and at the cooling water inlet of a nuclear power plant. Journal of Fish Biology, 69(3):735–748, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shiao:2006:MPC

[2210] J. C. Shiao, L. Ložys, Y. Iizuka, and W. N. Tzeng. Migratory patterns and contribution of stocking to the population of European eel in Lithuanian waters as indicated by otolith Sr : Ca ratios. Journal of Fish Biology, 69(3):749–769, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wong:2006:RCR

[2211] M. K. S. Wong and N. Y. S. Woo. Rapid changes in renal morphometrics in silver sea bream Sparus sarba on exposure to different salinities. Journal of Fish Biology, 69(3):770–782, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Siepker:2006:EAF

[2212] M. J. Siepker, K. G. Ostrand, and D. H. Wahl. Effects of angling on feeding by largemouth bass. Journal of Fish Biology, 69(3):783–793, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fey:2006:ETS

[2213] D. P. Fey. The effect of temperature and somatic growth on otolith growth: the discrepancy between two clupeid species from a similar environment. Journal of Fish Biology, 69(3):794–806, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2006:GCC

[2214] R. W. Blake, S. D. Inglis, and K. H. S. Chan. Growth, carcass composition and plasma growth hormone levels in cyclically fed rainbow trout. Journal of Fish Biology, 69(3):807–817, September 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2006:CFS

[2215] R. W. Blake and K. H. S. Chan. Cyclic feeding and subsequent compensatory growth do not significantly impact standard metabolic rate or critical swimming speed in rainbow trout. Journal of Fish Biology, 69 (3):818–827, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Potthoff:2006:GDG

[2216] M. T. Potthoff and M. C. Christman. Growth depensation and group behaviour in juvenile hybrid striped bass Morone chrysops × Morone saxatilis: effects of group membership, feeding method, ration size and size disparity. Journal of Fish Biology, 69(3):828–845, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kurita:2006:RIV

[2217] Y. Kurita. Regional and interannual variations in spawning activity of Pacific saury Cololabis saira during northward migration in spring in the north-western Pacific. Journal of Fish Biology, 69(3):846–859, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2006:LHE

[2218] B. Jonsson and N. Jonsson. Life-history effects of migratory costs in anadromous brown trout. Journal of Fish Biology, 69(3):860–869, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rushbrook:2006:NCK

[2219] B. J. Rushbrook and I. Barber. Nesting, courtship and kidney hypertrophy in Schistocephalus-infected male three-spined stickleback from an upland lake. Journal of Fish Biology, 69(3):870–882, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sazima:2006:GPM

[2220] C. Sazima, J. P. Krajewski, R. M. Bonaldo, and P. R. Guimarães, Jr. The goatfish Pseudupeneus maculatus and its follower fishes at an oceanic island in the tropical west Atlantic. Journal of Fish Biology, 69(3):883– 891, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scholz:2006:SPH

[2221] K. Scholz and F. Ladich. Sound production, hearing and possible interception under ambient noise conditions in the topmouth minnow Pseudorasbora parva. Journal of Fish Biology, 69(3):892–906, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hobbs:2006:ANH

[2222] J. A. Hobbs, W. A. Bennett, and J. E. Burton. Assessing nursery habitat quality for native smelts (Osmeridae) in the low-salinity zone of the San Francisco estuary. Journal of Fish Biology, 69(3):907–922, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2006:HME

[2223] F. Lahnsteiner and M. Kletzl. High mortality of embryos of a distinct autochthonous Austrian coregonid population during Zuger jar incubation. Journal of Fish Biology, 69(3):923–927, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2006:IIG

[2224] S. K. Wilson, D. T. Wilson, C. Lamont, and M. Evans. Identifying individual great barracuda Sphyraena barracuda using natural body marks. Journal of Fish Biology, 69(3):928–932, September 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Semmens:2006:IIP

[2225] B. X. Semmens, K. E. Luke, P. G. Bush, C. M. R. McCoy, and B. C. Johnson. Isopod infestation of postspawning Nassau grouper around Little Cayman Island. Journal of Fish Biology, 69(3):933–937, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ishikawa:2006:IDP

[2226] M. Ishikawa, S. Mori, and Y. Nagata. Intraspecific differences in patterns of courtship behaviours between the Pacific Ocean and Japan Sea forms of the three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 69(3):938–944, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bekkevold:2006:MSC

[2227] D. Bekkevold. Male size composition affects male reproductive variance in Atlantic cod Gadus morhua L. spawning aggregations. Journal of Fish Biology, 69(3):945–950, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Steinhausen:2006:EED

[2228] M. F. Steinhausen, N. G. Andersen, and J. F. Steffensen. The effect of external dummy transmitters on oxygen consumption and performance of swimming Atlantic cod. Journal of Fish Biology, 69(3):951–956, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petrie:2006:HLL

[2229] M. E. Petrie and C. H. Ryer. Hunger, light level and body size affect refuge use by post-settlement lingcod Ophiodon elongatus. Journal of Fish Biology, 69(4):957–969, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stein:2006:SCC

[2230] D. L. Stein, J. C. Drazen, K. L. Schlining, J. P. Barry, and L. Kuhnz. Snailfishes of the central California coast: video, photographic and morphological observations. Journal of Fish Biology, 69(4):970–986, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Travers:2006:ILW

[2231] M. J. Travers, S. J. Newman, and I. C. Potter. Influence of latitude, water depth, day v. night and wet v. dry periods on the species composition of reef fish communities in tropical Western Australia. Journal of Fish Biology, 69(4):987–1017, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munoz:2006:BBD

[2232] H. Muñoz, P. A. Van Damme, and F. Duponchelle. Breeding behaviour and distribution of the tucunaré Cichla aff. monoculus in a clear water river of the Bolivian Amazon. Journal of Fish Biology, 69(4):1018–1030, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feary:2006:HUT

[2233] D. A. Feary and K. D. Clements. Habitat use by triplefin species (Tripterygiidae) on rocky reefs in New Zealand. Journal of Fish Biology, 69(4):1031–1046, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davey:2006:RUS

[2234] A. J. H. Davey, D. J. Kelly, and B. J. F. Biggs. Refuge-use strategies of stream fishes in response to extreme low flows. Journal of Fish Biology, 69 (4):1047–1059, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geffen:2006:TDD

[2235] A. J. Geffen, C. J. Fox, and R. D. M. Nash. Temperature-dependent development rates of cod Gadus morhua eggs. Journal of Fish Biology, 69 (4):1060–1080, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sullivan:2006:ARG

[2236] M. C. Sullivan, K. W. Able, J. A. Hare, and H. J. Walsh. Anguilla rostrata glass eel ingress into two, U.S. east coast estuaries: patterns, processes and implications for adult abundance. Journal of Fish Biology, 69(4): 1081–1101, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

David:2006:MPH

[2237] C. J. David and T. J. Pandian. Maternal and paternal hybrid triploids of tetras. Journal of Fish Biology, 69(4):1102–1119, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2006:RHU

[2238] J. Yang, T. Arai, H. Liu, N. Miyazaki, and K. Tsukamoto. Reconstructing habitat use of Coilia mystus and Coilia ectenes of the Yangtze River estuary, and of Coilia ectenes of Taihu Lake, based on otolith strontium and calcium. Journal of Fish Biology, 69(4):1120–1135, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schulze:2006:DHU

[2239] T. Schulze, H. Dörner, F. Hölker, and T. Mehner. Determinants of habitat use in large roach. Journal of Fish Biology, 69(4):1136–1150, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niemela:2006:PSA

[2240] E. Niemelä, P. Orell, J. Erkinaro, J. B. Dempson, S. BrØrs, M. A. Svenning, and E. Hassinen. Previously spawned Atlantic salmon ascend a large subarctic river earlier than their maiden counterparts. Journal of Fish Biology, 69(4):1151–1163, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bakker:2006:BAT

[2241] T. C. M. Bakker, D. Mazzi, and S. B. M. Kraak. Broods of attractive three-spined stickleback males require greater paternal care. Journal of Fish Biology, 69(4):1164–1177, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Furuita:2006:LFA

[2242] H. Furuita, T. Unuma, K. Nomura, H. Tanaka, K. Okuzawa, T. Sugita, and T. Yamamoto. Lipid and fatty acid composition of eggs producing larvae with high survival rate in the Japanese eel. Journal of Fish Biology, 69(4):1178–1189, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Werneyer:2006:ODE

[2243] M. Werneyer and B. Kramer. Ontogenetic development of electric-organ discharges in a mormyrid fish, the bulldog Marcusenius macrolepidotus (South African form). Journal of Fish Biology, 69(4):1190–1201, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vassilopoulou:2006:DHD

[2244] V. Vassilopoulou. Dietary habits of the deep-sea flatfish Lepidorhombus boscii in north-eastern Mediterranean waters. Journal of Fish Biology, 69 (4):1202–1220, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2006:EET

[2245] C. I. Martin and I. A. Johnston. Endurance exercise training in common carp Cyprinus carpio L. induces proliferation of myonuclei in fast muscle fibres and slow muscle fibre hypertrophy. Journal of Fish Biology, 69(4): 1221–1227, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melia:2006:SDE

[2246] P. Melià, D. Bevacqua, A. J. Crivelli, J. Panfili, G. A. De Leo, and M. Gatto. Sex differentiation of the European eel in brackish and freshwater environments: a comparative analysis. Journal of Fish Biology, 69 (4):1228–1235, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2006:FRW

[2247] J. R. Britton and G. D. Davies. First record of the white catfish Ameiurus catus in Great Britain. Journal of Fish Biology, 69(4):1236–1238, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gagliardi-Seeley:2006:MSP

[2248] J. L. Gagliardi-Seeley and M. Itzkowitz. Male size predicts the ability to defend offspring in the biparental convict cichlid Archocentrus nigrofasciatus. Journal of Fish Biology, 69(4):1239–1244, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitamura:2006:JBL

[2249] T. Kitamura, M. Kume, H. Takahashi, and A. Goto. Juvenile bimodal length distribution and sea-run migration of the lower modal group in the Pacific Ocean form of three-spined stickleback. Journal of Fish Biology, 69(4):1245–1250, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eryilmaz:2006:FRA

[2250] L. Eryilmaz and C. Dalyan. First record of Apogon queketti Gilchrist (Osteichthyes: Apogonidae) in the Mediterranean Sea. Journal of Fish Biology, 69(4):1251–1254, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorner:2006:ETM

[2251] H. Dörner, T. Schulze, T. Mehner, and F. Hölker. The effect of temperature on mortality in small perch marked with coded wire tags. Journal of Fish Biology, 69(4):1255–1260, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRe

[2252] Anonymous. Book reviews. Journal of Fish Biology, 69(4):1261–1264, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRf

[2253] Anonymous. Books received. Journal of Fish Biology, 69(4):1265, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:C

[2254] Anonymous. Corrigendum. Journal of Fish Biology, 69(4):1266, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schaefer:2006:XEC

[2255] J. Schaefer and A. Ryan. XII European Congress of Ichthyology — ECI XII: 9–13 September, 2007 Dubrovnik (Cavtat), Croatia First Announcement. Journal of Fish Biology, 69(4):1267, October 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pratchett:2006:DAC

[2256] M. S. Pratchett, S. K. Wilson, and A. H. Baird. Declines in the abundance of Chaetodon butterflyfishes following extensive coral depletion. Journal of Fish Biology, 69(5):1269–1280, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lysne:2006:RBG

[2257] D. A. Lysne, W. Hemmingsen, and A. Skorping. Is reduced body growth of cod exposed to the gill parasite Lernaeocera branchialis a cost of resistance? Journal of Fish Biology, 69(5):1281–1287, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murua:2006:RSS

[2258] H. Murua and L. Motos. Reproductive strategy and spawning activity of the European hake Merluccius merluccius (L.) in the Bay of Biscay. Journal of Fish Biology, 69(5):1288–1303, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murua:2006:SVE

[2259] H. Murua, P. Lucio, M. Santurtún, and L. Motos. Seasonal variation in egg production and batch fecundity of European hake Merluccius merluccius (L.) in the Bay of Biscay. Journal of Fish Biology, 69(5):1304–1316, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paradis:2006:DBP

[2260] Y. Paradis, A. Bertolo, A. Perot, S. Débas, and P. Magnan. Do benthivory and piscivory result in similar growth in walleye? Journal of Fish Biology, 69(5):1317–1329, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crook:2006:UOC

[2261] D. A. Crook, J. I. Macdonald, J. P. O’Connor, and B. Barry. Use of otolith chemistry to examine patterns of diadromy in the threatened Australian grayling Prototroctes maraena. Journal of Fish Biology, 69(5): 1330–1344, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2006:ETP

[2262] R. W. Blake, P. Y. L. Kwok, and K. H. S. Chan. Effects of two parasites, Schistocephalus solidus (Cestoda) and Bunodera spp. (Trematoda), on the escape fast-start performance of three-spined sticklebacks. Journal of Fish Biology, 69(5):1345–1355, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dejen:2006:PTR

[2263] E. Dejen, J. Vijverberg, M. De Graaf, and F. A. Sibbing. Predicting and testing resource partitioning in a tropical fish assemblage of zooplanktivorous ‘barbs’: an ecomorphological approach. Journal of Fish Biology, 69(5):1356–1378, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowlands:2006:GMC

[2264] W. Ll. Rowlands, M. Dickey-Collas, A. J. Geffen, and R. D. M. Nash. Gape morphology of cod Gadus morhua L., haddock Melanogrammus aeglefinus (L.) and whiting Merlangius merlangus (L.) through metamorphosis from larvae to juveniles in the western Irish Sea. Journal of Fish Biology, 69(5):1379–1395, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bakke-McKellep:2006:EDS

[2265] A. M. Bakke-McKellep, S. Refstie, S. O. Stefansson, V. Vanthanouvong, G. Roomans, G.-I. Hemre, and Å. Krogdahl. Effects of dietary soybean meal and photoperiod cycle on osmoregulation following seawater exposure in Atlantic salmon smolts. Journal of Fish Biology, 69(5):1396–1426, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manera:2006:ABC

[2266] M. Manera and D. Britti. Assessment of blood chemistry normal ranges in rainbow trout. Journal of Fish Biology, 69(5):1427–1434, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spence:2006:DHP

[2267] R. Spence, M. K. Fatema, M. Reichard, K. A. Huq, M. A. Wahab, Z. F. Ahmed, and C. Smith. The distribution and habitat preferences of the zebrafish in Bangladesh. Journal of Fish Biology, 69(5):1435–1448, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sulikowski:2006:UCV

[2268] J. A. Sulikowski, J. Kneebone, S. Elzey, J. Jurek, W. H. Howell, and P. C. W. Tsang. Using the composite variables of reproductive morphology, histology and steroid hormones to determine age and size at sexual maturity for the thorny skate Amblyraja radiata in the western Gulf of Maine. Journal of Fish Biology, 69(5):1449–1465, November 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solem:2006:IIP

[2269] Ø. Solem, O. K. Berg, and A. J. Kjøsnes. Inter- and intra-population morphological differences between wild and farmed Atlantic salmon juveniles. Journal of Fish Biology, 69(5):1466–1481, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cook:2006:TTN

[2270] A. M. Cook, J. Duston, and R. G. Bradford. Thermal tolerance of a northern population of striped bass Morone saxatilis. Journal of Fish Biology, 69(5):1482–1490, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Benton:2006:ISC

[2271] C. B. Benton and D. L. Berlinsky. Induced sex change in black sea bass. Journal of Fish Biology, 69(5):1491–1503, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cunha:2006:AEC

[2272] F. E. A. Cunha and I. L. Rosa. Anaesthetic effects of clove oil on seven species of tropical reef teleosts. Journal of Fish Biology, 69(5):1504–1512, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hickey:2006:EPN

[2273] M. A. Hickey and G. P. Closs. Evaluating the potential of night spotlighting as a method for assessing species composition and brown trout abundance: a comparison with electrofishing in small streams. Journal of Fish Biology, 69(5):1513–1523, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neto:2006:HPC

[2274] F. Filipak Neto, S. M. Zanata, M. A. F. Randi, É. Pelletier, and C. A. Oliveira Ribeiro. Hepatocytes primary culture from the Neotropical fish, trahira Hoplias malabaricus (Bloch). Journal of Fish Biology, 69(5):1524–1532, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferreira:2006:CSD

[2275] C. E. L. Ferreira and J. E. A. Gonçalves. Community structure and diet of roving herbivorous reef fishes in the Abrolhos Archipelago, southwestern Atlantic. Journal of Fish Biology, 69(5):1533–1551, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moran-Lopez:2006:ABF

[2276] R. Morán-López, E. Da Silva, J. L. Pérez-Bote, and C. Corbacho Amado. Associations between fish assemblages and environmental factors for Mediterranean-type rivers during summer. Journal of Fish Biology, 69(5):1552–1569, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sellin:2006:MDC

[2277] M. K. Sellin and A. S. Kolok. Maternally derived Cu tolerance in larval fathead minnows: how long does it persist? Journal of Fish Biology, 69(5):1570–1574, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pollux:2006:OAC

[2278] B. J. A. Pollux and A. Korosi. On the occurrence of the Asiatic cyprinid Pseudorasbora parva in The Netherlands. Journal of Fish Biology, 69(5): 1575–1580, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBg

[2279] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 69(5):1581, November 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zamzow:2006:UAM

[2280] J. P. Zamzow and U. E. Siebeck. Ultraviolet absorbance of the mucus of a tropical damselfish: effects of ontogeny, captivity and disease. Journal of Fish Biology, 69(6):1583–1594, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kelly:2006:EPF

[2281] B. Kelly, J. B. Dempson, and M. Power. The effects of preservation on fish tissue stable isotope signatures. Journal of Fish Biology, 69(6): 1595–1611, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cannas:2006:GTO

[2282] M. Cannas, J. Schaefer, P. Domenici, and J. F. Steffensen. Gait transition and oxygen consumption in swimming striped surfperch Embiotoca lateralis Agassiz. Journal of Fish Biology, 69(6):1612–1625, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2006:ESI

[2283] R. S. Brown, D. R. Geist, K. A. Deters, and A. Grassell. Effects of surgically implanted acoustic transmitters >2% of body mass on the swimming performance, survival and growth of juvenile sockeye and Chinook salmon. Journal of Fish Biology, 69(6):1626–1638, December 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lugendo:2006:IMM

[2284] B. R. Lugendo, I. Nagelkerken, G. Van Der Velde, and Y. D. Mgaya. The importance of mangroves, mud and sand flats, and seagrass beds as feeding areas for juvenile fishes in Chwaka Bay, Zanzibar: gut content and stable isotope analyses. Journal of Fish Biology, 69(6):1639–1661, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shoji:2006:ERF

[2285] J. Shoji, T. Ohta, and M. Tanaka. Effects of river flow on larval growth and survival of Japanese seaperch Lateolabrax japonicus (Pisces) in the Chikugo River estuary, upper Ariake Bay. Journal of Fish Biology, 69(6): 1662–1674, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morbey:2006:WPH

[2286] Y. E. Morbey, P. Addison, B. J. Shuter, and K. Vascotto. Withinpopulation heterogeneity of habitat use by lake trout Salvelinus namaycush. Journal of Fish Biology, 69(6):1675–1696, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lagardere:2006:SSM

[2287] J. P. Lagardère and A. Mariani. Spawning sounds in meagre Argyrosomus regius recorded in the Gironde estuary, France. Journal of Fish Biology, 69(6):1697–1708, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thorsen:2006:CVP

[2288] A. Thorsen, C. T. Marshall, and O. S. Kjesbu. Comparison of various potential fecundity models for north-east Arctic cod Gadus morhua, L. using oocyte diameter as a standardizing factor. Journal of Fish Biology, 69(6):1709–1730, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morbey:2006:ENR

[2289] Y. E. Morbey and C. G. Guglielmo. Evidence of a novel reproductive tactic in female Kokanee Oncorhynchus nerka. Journal of Fish Biology, 69(6):1731–1743, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tuz-Sulub:2006:ESD

[2290] A. Tuz-Sulub, T. Brulé, K. Cervera-Cervera, and J. C. Espinoza-Mendez. Evidence for sexual dichromatisms in spawning aggregations of yellowfin grouper Mycteroperca venenosa and tiger grouper Mycteroperca tigris from the southern Gulf of Mexico. Journal of Fish Biology, 69(6):1744– 1755, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hwang:2006:GLP

[2291] S.-D. Hwang, M.-H. Song, T.-W. Lee, G. A. McFarlane, and J. R. King. Growth of larval Pacific anchovy Engraulis japonicus in the Yellow Sea as indicated by otolith microstructure analysis. Journal of Fish Biology, 69(6):1756–1769, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mello:2006:OAC

[2292] F. Teixeirade Mello, C. Iglesias, A. I. Borthagaray, N. Mazzeo, J. Vilches, D. Larrea, and R. Ballabio. Ontogenetic allometric coefficient changes: implications of diet shift and morphometric traits in Hoplias malabaricus (Bloch) (Characiforme, Erythrinidae). Journal of Fish Biology, 69(6): 1770–1778, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ricchiari:2006:FFS

[2293] L. Ricchiari, A. Liguoro, P. Andreuccetti, F. Angelini, and M. Prisco. The Fas/FasL system is a possible regulator of germ cell apoptosis in the testis of Torpedo marmorata. Journal of Fish Biology, 69(6):1779–1787, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galloway:2006:ISC

[2294] B. J. Galloway and K. R. Munkittrick. Influence of seasonal changes in relative liver size, condition, relative gonad size and variability in ovarian development in multiple spawning fish species used in environmental monitoring programmes. Journal of Fish Biology, 69(6):1788–1806, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Godiksen:2006:EAP

[2295] J. A. Godiksen, E. H. Hallfredsson, and T. Pedersen. Effects of alternative prey on predation intensity from herring Clupea harengus and sandeel Ammodytes marinus on capelin Mallotus villosus larvae in the Barents Sea. Journal of Fish Biology, 69(6):1807–1823, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2006:MTF

[2296] R. W. Blake and K. H. S. Chan. Models of the turning and fast-start swimming dynamics of aquatic vertebrates. Journal of Fish Biology, 69(6):1824–1836, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lengkeek:2006:BCR

[2297] W. Lengkeek and K. Didderen. Breeding cycles and reproductive behaviour in the river blenny Salaria fluviatilis. Journal of Fish Biology, 69(6):1837–1844, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meisfjord:2006:VDI

[2298] J. Meisfjord, F. Midtøy, and A. Folkvord. Validation of daily increment deposition in otoliths of juvenile Limnothrissa miodon (Clupeidae). Journal of Fish Biology, 69(6):1845–1848, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sato:2006:DDP

[2299] T. Sato. Dramatic decline in population abundance of Salvelinus leucomaenis after a severe flood and debris flow in a high gradient stream. Journal of Fish Biology, 69(6):1849–1854, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Curtis:2006:CMI

[2300] J. M. R. Curtis. A case of mistaken identity: skin filaments are unreliable for identifying Hippocampus guttulatus and Hippocampus hippocampus. Journal of Fish Biology, 69(6):1855–1859, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2006:RBB

[2301] K. Silva, N. M. Monteiro, M. N. Vieira, and V. C. Almada. Reproductive behaviour of the black-striped pipefish Syngnathus abaster (Pisces; Syngnathidae). Journal of Fish Biology, 69(6):1860–1869, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flajshans:2006:DBW

[2302] M. Flajšhans and V. Piačková. Difference in blood and water diffusion distance in gill lamellae of diploid and triploid tench Tinca tinca (L.). Journal of Fish Biology, 69(6):1870–1873, December 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hutchinson:2006:SIA

[2303] J. J. Hutchinson and C. N. Trueman. Stable isotope analyses of collagen in fish scales: limitations set by scale architecture. Journal of Fish Biology, 69(6):1874–1880, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mancini:2006:MAB

[2304] P. L. Mancini, A. L. Casas, and A. F. Amorim. Morphological abnormalities in a blue shark Prionace glauca (Chondrichthyes: Carcharhinidae) foetus from southern Brazil. Journal of Fish Biology, 69(6):1881–1884, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Widmer:2006:ECD

[2305] S. Widmer, F. B.-G. Moore, and B. Bagatto. The effects of chronic developmental hypoxia on swimming performance in zebrafish. Journal of Fish Biology, 69(6):1885–1891, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2006:SVS

[2306] M. C. P. Amorim, R. O. Vasconcelos, J. F. Marques, and F. Almada. Seasonal variation of sound production in the Lusitanian toadfish Halobatrachus didactylus. Journal of Fish Biology, 69(6):1892–1899, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRg

[2307] Anonymous. Book reviews. Journal of Fish Biology, 69(6):1900–1904, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:BRh

[2308] Anonymous. Books received. Journal of Fish Biology, 69(6):1905–1906, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:FSBh

[2309] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 69(6):1907, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:CV

[2310] Anonymous. Contents of volume 69. Journal of Fish Biology, 69(6): xxvii–xxxiv, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:IVb

[2311] Anonymous. Index, volume 69. Journal of Fish Biology, 69(6):i–xxvi, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olwell:2006:PCH

[2312] P. M. Olwell, F. M. Lyng, S. Ni Shuilleabhain, P. Maderna, C. Mothersill, and D. C. Cottell. Primary culture and histological characterization of phagocytic cells from rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 69(sa):1–19, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yashuvi:2006:OGR

[2313] Y. Yashuvi, U. Klenke, O. Palevitch, E. Abraham, Y. Zohar, and Y. Gothilf. Ontogeny of gonadotropin-releasing hormone (GnRH) neurons in hybrid striped bass Morone sp.: whole-mount in situ hybridization analysis. Journal of Fish Biology, 69(sa):20–30, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubert:2006:ERI

[2314] N. Hubert, F. Duponchelle, J. Nuñez, R. Rivera, and J.-F. Renno. Evidence of reproductive isolation among closely related sympatric species of Serrasalmus (Ostariophysii, Characidae) from the Upper Madeira River, Amazon, Bolivia. Journal of Fish Biology, 69(sa):31–51, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wyatt:2006:MAD

[2315] P. M. W. Wyatt, C. S. Pitts, and R. K. Butlin. A molecular approach to detect hybridization between bream Abramis brama, roach rutlius rutilus and rudd Scardinius erythrophthalmus. Journal of Fish Biology, 69(sa): 52–71, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peterman:2006:OAP

[2316] A. E. Peterman and L. Petrie-Hanson. Ontogeny of American paddlefish lymphoid tissues. Journal of Fish Biology, 69(sa):72–88, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duguid:2006:RIG

[2317] R. A. Duguid, A. Ferguson, and P. Prodöhl. Reproductive isolation and genetic differentiation of ferox trout from sympatric brown trout in Loch Awe and Loch Laggan, Scotland. Journal of Fish Biology, 69(sa):89–114, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ye:2006:DCC

[2318] H.-Q. Ye, S.-L. Chen, Z.-X. Sha, and M.-Y. Xu. Development and characterization of cell lines from heart, liver, spleen and head kidney of sea perch Lateolabrax japonicus. Journal of Fish Biology, 69(sa):115–126, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caudron:2006:ARS

[2319] A. Caudron, A. Champigneulle, and R. Guyomard. Assessment of restocking as a strategy for rehabilitating a native population of brown trout Salmo trutta L. in a fast-flowing mountain stream in the northern French Alps. Journal of Fish Biology, 69(sa):127–139, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2006:CSR

[2320] J. B. Olsen, S. J. Miller, K. Harper, J. J. Nagler, and J. K. Wenburg. Contrasting sex ratios in juvenile and adult Chinook salmon Oncorhynchus tshawytscha (Walbaum) from south-west Alaska: sex reversal or differential survival? Journal of Fish Biology, 69(sa):140–144, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OLeary:2006:ARM

[2321] D. B. O’Leary, J. Coughlan, T. V. McCarthy, and T. F. Cross. Application of a rapid method of SNP analysis (glycosylase mediated polymorphism detection) to mtDNA and nuclear DNA of cod Gadus morhua. Journal of Fish Biology, 69(sa):145–153, September 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsuruta:2006:FSG

[2322] T. Tsuruta and A. Goto. Fine scale genetic population structure of the freshwater and Omono types of nine-spined stickleback Pungitius pungitius (L.) within the Omono River system, Japan. Journal of Fish Biology, 69(sb):155–176, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schaschl:2006:PSS

[2323] H. Schaschl and K. M. Wegner. Polymorphism and signature of selection in the MHC class I genes of the three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 69(sb):177–188, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garner:2006:GVM

[2324] S. Garner, T. P. Birt, C. M. Mlewa, J. M. Green, A. Seifert, and V. L. Friesen. Genetic variation in the marbled lungfish Protopterus aethiopicus in Lake Victoria and introduction to Lake Baringo, Kenya. Journal of Fish Biology, 69(sb):189–199, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ready:2006:DFM

[2325] J. S. Ready, E. J. G. Ferreira, and S. O. Kullander. Discus fishes: mitochondrial DNA evidence for a phylogeographic barrier in the Amazonian genus Symphysodon (Teleostei: Cichlidae). Journal of Fish Biology, 69 (sb):200–211, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

April:2006:PBK

[2326] J. April and J. Turgeon. Phylogeography of the banded killifish (Fundulus diaphanus): glacial races and secondary contact. Journal of Fish Biology, 69(sb):212–228, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:2006:LGV

[2327] M. M. Hansen, V. Simonsen, K.-L. D. Mensberg, Md. R. I. Sarder, and Md. S. Alam. Loss of genetic variation in hatchery-reared Indian major carp, Catla catla. Journal of Fish Biology, 69(sb):229–241, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alexandrino:2006:IDI

[2328] P. Alexandrino, R. Faria, D. Linhares, F. Castro, M. Le Corre, R. Sabatié, J. L. Baglinière, and S. Weiss. Interspecific differentiation and intraspecific substructure in two closely related clupeids with extensive hybridization, Alosa alosa and Alosa fallax. Journal of Fish Biology, 69 (sb):242–259, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2006:GDI

[2329] A. V. Oliveira, A. J. Prioli, S. M. A. P. Prioli, T. S. Bignotto, H. F. Júlio, Jr., H. Carrer, C. S. Agostinho, and L. M. Prioli. Genetic diversity of invasive and native Cichla (Pisces: Perciformes) populations in Brazil with evidence of interspecific hybridization. Journal of Fish Biology, 69 (sb):260–277, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferrando:2006:KAI

[2330] S. Ferrando, M. Bottaro, L. Gallus, L. Girosi, M. Vacchi, and G. Tagliafierro. Na^{+} /K^{+} ATPase immunoreactivity in olfactory epithelium of small-spotted catshark Scyliorhinus canicula (L.): possible presence of ion exchanging cells? Journal of Fish Biology, 69(sb):278–282, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spies:2006:DBI

[2331] I. B. Spies, S. Gaichas, D. E. Stevenson, J. W. Orr, and M. F. Canino. DNA-based identification of Alaska skates (Amblyraja, Bathyraja and Raja: Rajidae) using cytochrome c oxidase subunit I (coI) variation. Journal of Fish Biology, 69(sb):283–292, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gross:2006:VIB

[2332] K. Gross, S. Alcorn, A. Murray, R. Morrison, and B. Nowak. In vitro interactions between Neoparamoeba spp. and salmonid leucocytes; the effect of parasite sonicate on anterior kidney leucocyte function. Journal of Fish Biology, 69(sb):293–300, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:CVS

[2333] Anonymous. Contents of volume 69 supplements. Journal of Fish Biology, 69(sb):vii–viii, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:IVc

[2334] Anonymous. Index, volume 69. Journal of Fish Biology, 69(sb):i–vi, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Persson:2006:FDI

[2335] L. Persson and A. M. De Roos. Food-dependent individual growth and population dynamics in fishes. Journal of Fish Biology, 69(sc):1–20, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conover:2006:STS

[2336] D. O. Conover, L. M. Clarke, S. B. Munch, and G. N. Wagner. Spatial and temporal scales of adaptive divergence in marine fishes and the implications for conservation. Journal of Fish Biology, 69(sc):21–47, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2006:FPS

[2337] J. D. Metcalfe. Fish population structuring in the North Sea: understanding processes and mechanisms from studies of the movements of adults. Journal of Fish Biology, 69(sc):48–65, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stefansson:2006:DMP

[2338] G. Stefansson and A. A. Rosenberg. Designing marine protected areas for migrating fish stocks. Journal of Fish Biology, 69(sc):66–78, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bhagat:2006:MDI

[2339] Y. Bhagat, M. G. Fox, and M. T. Ferreira. Morphological differentiation in introduced pumpkinseed Lepomis gibbosus (L.) occupying different habitat zones in Portuguese reservoirs. Journal of Fish Biology, 69(sc): 79–94, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradbury:2006:MGD

[2340] I. R. Bradbury, M. W. Coulson, S. E. Campana, and P. Bentzen. Morphological and genetic differentiation in anadromous smelt Osmerus mordax (Mitchill): disentangling the effects of geography and morphology on gene flow. Journal of Fish Biology, 69(sc):95–114, December 2006. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gum:2006:DIR

[2341] B. Gum, R. Gross, and R. Kuehn. Discriminating the impact of recent human mediated stock transfer from historical gene flow on genetic structure of European grayling Thymallus thymallus L. Journal of Fish Biology, 69(sc):115–135, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsdottir:2006:SSI

[2342] I. G. Jónsdóttir, S. E. Campana, and G. Marteinsdottir. Stock structure of Icelandic cod Gadus morhua L. based on otolith chemistry. Journal of Fish Biology, 69(sc):136–150, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svedang:2006:CGM

[2343] H. Svedäng and A. Svenson. Cod Gadus morhua L. populations as behavioural units: inference from time series on juvenile abundance in the eastern Skagerrak. Journal of Fish Biology, 69(sc):151–164, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turan:2006:UOS

[2344] C. Turan. The use of otolith shape and chemistry to determine stock structure of Mediterranean horse mackerel Trachurus mediterraneus (Steindachner). Journal of Fish Biology, 69(sc):165–180, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2006:EMS

[2345] P. J. Wright, F. C. Neat, F. M. Gibb, I. M. Gibb, and H. Thordarson. Evidence for metapopulation structuring in cod from the west of Scotland and North Sea. Journal of Fish Biology, 69(sc):181–199, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bevacqua:2006:TRS

[2346] D. Bevacqua, P. Melià, A. J. Crivelli, G. A. De Leo, and M. Gatto. Timing and rate of sexual maturation of European eel in brackish and freshwater environments. Journal of Fish Biology, 69(sc):200–208, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coulson:2006:TGD

[2347] M. W. Coulson, I. R. Bradbury, and P. Bentzen. Temporal genetic differentiation: continuous v. discontinuous spawning runs in anadromous rainbow smelt Osmerus mordax (Mitchill). Journal of Fish Biology, 69 (sc):209–216, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jue:2006:ESG

[2348] N. K. Jue. Exploring the structure of genetic variation and the influences of demography on effective population size in the gag grouper Mycteroperca microlepi (Goode & Bean). Journal of Fish Biology, 69(sc):217–224, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:AOP

[2349] Anonymous. Abstracts of oral papers not published in the supplement. Journal of Fish Biology, 69(sc):225–245, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:APP

[2350] Anonymous. Abstracts of posters not published in the supplement. Journal of Fish Biology, 69(sc):246–265, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:RS

[2351] Anonymous. Referees for this supplement. Journal of Fish Biology, 69 (sc):266, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2006:IS

[2352] Anonymous. Index, 2006 supplement. Journal of Fish Biology, 69(sc): 267–270, December 2006. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ferrito:2007:MVA

[2353] V. Ferrito, M. C. Mannino, A. M. Pappalardo, and C. Tigano. Morphological variation among populations of Aphanius fasciatus Nardo, 1827 (Teleostei, Cyprinodontidae) from the Mediterranean. Journal of Fish Biology, 70(1):1–20, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shoup:2007:SSG

[2354] D. E. Shoup, S. P. Callahan, D. H. Wahl, and C. L. Pierce. Size-specific growth of bluegill, largemouth bass and channel catfish in relation to prey availability and limnological variables. Journal of Fish Biology, 70 (1):21–34, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gonzalez-Salas:2007:IVI

[2355] C. Gonzalez-Salas and P. Lenfant. Interannual variability and intraannual stability of the otolith shape in European anchovy Engraulis encrasicolus (L.) in the Bay of Biscay. Journal of Fish Biology, 70(1):35–49, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolner:2007:WPU

[2356] K. C. S. Bolner and B. Baldisserotto. Water pH and urinary excretion in silver catfish Rhamdia quelen. Journal of Fish Biology, 70(1):50–64, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stewart:2007:AVG

[2357] J. Stewart and J. M. Hughes. Age validation and growth of three commercially important hemiramphids in south-eastern Australia. Journal of Fish Biology, 70(1):65–82, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corriero:2007:GCP

[2358] A. Corriero, S. Desantis, C. R. Bridges, D. E. Kime, P. Megalofonou, N. Santamaria, F. Cirillo, G. Ventriglia, A. Di Summa, M. Deflorio, F. Campobasso, and G. De Metrio. Germ cell proliferation and apoptosis during different phases of swordfish (Xiphias gladius L.) spermatogenetic cycle. Journal of Fish Biology, 70(1):83–99, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dantas:2007:NPC

[2359] M. C. Dantas and J. L. Attayde. Nitrogen and phosphorus content of some temperate and tropical freshwater fishes. Journal of Fish Biology, 70(1):100–108, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mann:2007:HES

[2360] D. A. Mann, P. A. Cott, B. W. Hanna, and A. N. Popper. Hearing in eight species of northern Canadian freshwater fishes. Journal of Fish Biology, 70(1):109–120, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hendrixson:2007:ESF

[2361] H. A. Hendrixson, R. W. Sterner, and A. D. Kay. Elemental stoichiometry of freshwater fishes in relation to phylogeny, allometry and ecology. Journal of Fish Biology, 70(1):121–140, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cech:2007:DBP

[2362] M. Čech, J. Kubečka, J. Frouzová, V. Draštík, M. Kratochvíl, J. Matěna, and J. Hejzlar. Distribution of the bathypelagic perch fry layer along the longitudinal profile of two large canyon-shaped reservoirs. Journal of Fish Biology, 70(1):141–154, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eklov:2007:PPI

[2363] P. Eklöv and P. Jonsson. Pike predators induce morphological changes in young perch and roach. Journal of Fish Biology, 70(1):155–164, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:2007:SCS

[2364] T. E. Pitcher, F. H. Rodd, and L. Rowe. Sexual colouration and sperm traits in guppies. Journal of Fish Biology, 70(1):165–177, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zuccarelli:2007:IEO

[2365] M. D. Zuccarelli, N. Jensen, and R. L. Ingermann. Inhibitory effect of osmotic concentration, potassium and pH on motility of the sperm of the North American burbot Lota lota maculosa. Journal of Fish Biology, 70 (1):178–189, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

King:2007:RSE

[2366] H. R. King, N. W. Pankhurst, and M. Watts. Reproductive sensitivity to elevated water temperatures in female Atlantic salmon is heightened at certain stages of vitellogenesis. Journal of Fish Biology, 70(1):190–205, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dolinsek:2007:EHH

[2367] I. J. Dolinsek, J. W. A. Grant, and P. M. Biron. The effect of habitat heterogeneity on the population density of juvenile Atlantic salmon Salmo salar L. Journal of Fish Biology, 70(1):206–214, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buria:2007:MSA

[2368] L. Buria, S. J. Walde, M. Battini, P. J. Macchi, M. Alonso, D. E. Ruzzante, and V. E. Cussac. Movement of a South American perch Percichthys trucha in a mountain Patagonian lake during spawning and prespawning periods. Journal of Fish Biology, 70(1):215–230, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogue:2007:HQE

[2369] C. Hogue and B. Swig. Habitat quality and endoparasitism in the Pacific sanddab Citharichthys sordidus from Santa Monica Bay, southern California. Journal of Fish Biology, 70(1):231–242, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazzoldi:2007:VMR

[2370] C. Mazzoldi, V. Lorenzi, and M. B. Rasotto. Variation of male reproductive apparatus in relation to fertilization modalities in the catfish families Auchenipteridae and Callichthyidae (Teleostei: Siluriformes). Journal of Fish Biology, 70(1):243–256, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hojesjo:2007:MHR

[2371] J. Höjesjö, F. Økland, L. F. Sundström, J. Pettersson, and J. I. Johnsson. Movement and home range in relation to dominance; a telemetry study on brown trout Salmo trutta. Journal of Fish Biology, 70(1):257–268, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pais:2007:WRE

[2372] A. Pais, P. Merella, M. C. Follesa, and G. Garippa. Westward range expansion of the Lessepsian migrant Fistularia commersonii (Fistulariidae) in the Mediterranean Sea, with notes on its parasites. Journal of Fish Biology, 70(1):269–277, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Georgakopoulou:2007:ITD

[2373] E. Georgakopoulou, D. G. Sfakianakis, S. Kouttouki, P. Divanach, M. Kentouri, and G. Koumoundouros. The influence of temperature during early life on phenotypic expression at later ontogenetic stages in sea bass. Journal of Fish Biology, 70(1):278–291, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bauer:2007:CCT

[2374] C. Bauer and G. Loupal. Common carp tissue reactions to surgically implanted radio tags with external antennas. Journal of Fish Biology, 70 (1):292–297, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skov:2007:HCI

[2375] C. Skov, P. A. Nilsson, L. Jacobsen, and C. Brönmark. Habitat-choice interactions between pike predators and perch prey depend on water transparency. Journal of Fish Biology, 70(1):298–302, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2007:EME

[2376] J. Simon. Evaluation of marking European silver eels with visible implant elastomer tags and alcian blue. Journal of Fish Biology, 70(1):303–309, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thangaraj:2007:OJS

[2377] M. Thangaraj and A. P. Lipton. Occurrence of the Japanese seahorse Hippocampus mohnikei Bleeker 1854 from the Palk Bay coast of southeastern India. Journal of Fish Biology, 70(1):310–312, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tonkin:2007:DSD

[2378] Z. Tonkin, A. King, J. Mahoney, and J. Morrongiello. Diel and spatial drifting patterns of silver perch Bidyanus bidyanus eggs in an Australian lowland river. Journal of Fish Biology, 70(1):313–317, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeates:2007:TSD

[2379] S. Yeates, J. Searle, R. G. Ward, and M. J. G. Gage. A two-second delay confers first-male fertilization precedence within in vitro sperm competition experiments in Atlantic salmon. Journal of Fish Biology, 70 (1):318–322, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:R

[2380] Anonymous. Referees 2006. Journal of Fish Biology, 70(1):323–328, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:FSBa

[2381] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 70(1):329, January 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carvalho:2007:YPA

[2382] G. Carvalho and J. Thorpe. Yurii Petrovich Altukhov 1936–2006. Journal of Fish Biology, 70(2):331–333, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chernoff:2007:FSG

[2383] E. Chernoff and R. A. Curry. First summer growth predetermined in anadromous and resident brook charr. Journal of Fish Biology, 70(2): 334–346, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2007:GEB

[2384] C. Fu, D. Li, W. Hu, Y. Wang, and Z. Zhu. Growth and energy budget of F_{2} ‘all-fish’ growth hormone gene transgenic common carp. Journal of Fish Biology, 70(2):347–361, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lovell:2007:PHC

[2385] J. M. Lovell, M. M. Findlay, G. M. Harper, and R. M. Moate. The polarization of hair cells from the inner ear of the lesser spotted dogfish Scyliorhinus canicula. Journal of Fish Biology, 70(2):362–373, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yano:2007:DRF

[2386] K. Yano, J. D. Stevens, and L. J. V. Compagno. Distribution, reproduction and feeding of the Greenland shark Somniosus (Somniosus) microcephalus, with notes on two other sleeper sharks, Somniosus (Somniosus) pacificus and Somniosus (Somniosus) antarcticus. Journal of Fish Biology, 70(2):374–390, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Catalan:2007:RMB

[2387] I. A. Catalán, E. Berdalet, M. P. Olivar, and C. Roldán. Response of muscle-based biochemical condition indices to short-term variations in food availability in post-flexion reared sea bass Dicentrarchus labrax (L.) larvae. Journal of Fish Biology, 70(2):391–405, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brenner:2007:TMP

[2388] M. Brenner and U. Krumme. Tidal migration and patterns in feeding of the four-eyed fish Anableps anableps L. in a north Brazilian mangrove. Journal of Fish Biology, 70(2):406–427, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sulikowski:2007:RCB

[2389] J. A. Sulikowski, W. B. Driggers III, T. S. Ford, R. K. Boonstra, and J. K. Carlson. Reproductive cycle of the blacknose shark Carcharhinus acronotus in the Gulf of Mexico. Journal of Fish Biology, 70(2):428–440, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lappalainen:2007:LGO

[2390] J. Lappalainen and A. S. Tarkan. Latitudinal gradients in onset date, onset temperature and duration of spawning of roach. Journal of Fish Biology, 70(2):441–450, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2007:CSF

[2391] W.-M. Chen, N. Umeda, T. Mitsuboshi, and N. Hirazawa. Circadian selffeeding rhythms in greater amberjack Seriola dumerili (Risso). Journal of Fish Biology, 70(2):451–461, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eriksen:2007:LTE

[2392] M. S. Eriksen, Å. Espmark, B. O. Braastad, R. Salte, and M. Bakken. Long-term effects of maternal cortisol exposure and mild hyperthermia during embryogeny on survival, growth and morphological anomalies in farmed Atlantic salmon Salmo salar offspring. Journal of Fish Biology, 70(2):462–473, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pekcan-Hekim:2007:FEW

[2393] Z. Pekcan-Hekim and J. Horppila. Feeding efficiency of white bream at different inorganic turbidities and light climates. Journal of Fish Biology, 70(2):474–482, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fiorin:2007:CEA

[2394] R. Fiorin, S. Malavasi, A. Franco, and P. Franzoi. Comparative energy allocation in two sympatric, closely related gobies: the black goby Gobius niger and the grass goby Zosterisessor ophiocephalus. Journal of Fish Biology, 70(2):483–496, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nagel:2007:HPS

[2395] L. Nagel and A. S. Grutter. Host preference and specialization in Gnathia sp., a common parasitic isopod of coral reef fishes. Journal of Fish Biology, 70(2):497–508, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abascal:2007:CSM

[2396] F. J. Abascal, J. Cosson, and C. Fauvel. Characterization of sperm motility in sea bass: the effect of heavy metals and physicochemical variables on sperm motility. Journal of Fish Biology, 70(2):509–522, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mallela:2007:DPC

[2397] J. Mallela, C. Roberts, C. Harrod, and C. R. Goldspink. Distributional patterns and community structure of Caribbean coral reef fishes within a river-impacted bay. Journal of Fish Biology, 70(2):523–537, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuroki:2007:AGA

[2398] M. Kuroki, J. Aoyama, S. Wouthuyzen, K. Sumardhiharga, M. J. Miller, and K. Tsukamoto. Age and growth of Anguilla bicolor bicolor leptocephali in the eastern Indian Ocean. Journal of Fish Biology, 70(2): 538–550, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2007:TSF

[2399] D. M. Scott and R. W. Wilson. Three species of fishes from an eutrophic, seasonally alkaline lake are not more tolerant to acute exposure to high pH in the laboratory. Journal of Fish Biology, 70(2):551–566, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Faulk:2007:OGT

[2400] C. K. Faulk, A. D. Benninghoff, and G. J. Holt. Ontogeny of the gastrointestinal tract and selected digestive enzymes in cobia Rachycentron canadum (L.). Journal of Fish Biology, 70(2):567–583, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pilling:2007:PNS

[2401] G. M. Pilling, R. S. Millner, M. W. Easey, D. L. Maxwell, and A. N. Tidd. Phenology and North Sea cod Gadus morhua L.: has climate change affected otolith annulus formation and growth? Journal of Fish Biology, 70(2):584–599, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manabe:2007:FEB

[2402] H. Manabe, M. Ishimura, A. Shinomiya, and T. Sunobe. Field evidence for bi-directional sex change in the polygynous gobiid fish Trimma okinawae. Journal of Fish Biology, 70(2):600–609, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moss:2007:FRJ

[2403] J. H. Moss and D. A. Beauchamp. Functional response of juvenile pink and chum salmon: effects of consumer size and two types of zooplankton prey. Journal of Fish Biology, 70(2):610–622, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bhat:2007:DDA

[2404] A. Bhat and A. E. Magurran. Does disturbance affect the structure of tropical fish assemblages? A test using null models. Journal of Fish Biology, 70(2):623–629, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leguen:2007:DHA

[2405] I. Leguen, V. Véron, C. Sevellec, D. Azam, R. Sabatié, P. Prunet, and J. L. Baglinière. Development of hypoosmoregulatory ability in allis shad Alosa alosa. Journal of Fish Biology, 70(2):630–637, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hallfredsson:2007:EDR

[2406] E. H. Hallfredsson, T. Pedersen, and K. Præbel. Estimation of digestion rates for herring Clupea harengus L. feeding on fish larvae. Journal of Fish Biology, 70(2):638–643, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carseldine:2007:ACM

[2407] L. Carseldine and I. R. Tibbetts. Autoradiographic confirmation of mucus-mediated absorption in the halfbeak Hyporhamphus regularis ardelio (Whitley, 1931). Journal of Fish Biology, 70(2):644–647, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janez:2007:SHI

[2408] J. A. Jañez and M. C. Sueiro. Size at hatching and incubation period of Sympterygia bonapartii (Müller & Henle, 1841) (Chondrichthyes, Rajidae) bred in captivity at the Temaiken Aquarium. Journal of Fish Biology, 70(2):648–650, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McKenzie:2007:SAR

[2409] D. J. McKenzie, J. F. Steffensen, K. Korsmeyer, N. M. Whiteley, P. Bronzi, and E. W. Taylor. Swimming alters responses to hypoxia in the Adriatic sturgeon Acipenser naccarii. Journal of Fish Biology, 70 (2):651–658, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Ea

[2410] Anonymous. Errata. Journal of Fish Biology, 70(2):659, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jue:2007:FSB

[2411] N. K. Jue. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 70(2):660, February 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Juanes:2007:RHM

[2412] F. Juanes. Role of habitat in mediating mortality during the postsettlement transition phase of temperate marine fishes. Journal of Fish Biology, 70(3):661–677, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maunder:2007:RBS

[2413] R. J. Maunder, P. Matthiessen, J. P. Sumpter, and T. G. Pottinger. Rapid bioconcentration of steroids in the plasma of three-spined stickleback Gasterosteus aculeatus exposed to waterborne testosterone and 17β-oestradiol. Journal of Fish Biology, 70(3):678–690, March 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poos:2007:ETC

[2414] M. S. Poos, N. E. Mandrak, and R. L. McLaughlin. The effectiveness of two common sampling methods for assessing imperilled freshwater fishes. Journal of Fish Biology, 70(3):691–708, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reichard:2007:PEL

[2415] M. Reichard, M. Przybylski, P. Kaniewska, H. Liu, and C. Smith. A possible evolutionary lag in the relationship between freshwater mussels and European bitterling. Journal of Fish Biology, 70(3):709–725, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2007:FFRa

[2416] A. D. Nunn, J. P. Harvey, and I. G. Cowx. The food and feeding relationships of larval and 0+ year juvenile fishes in lowland rivers and connected waterbodies. I. Ontogenetic shifts and interspecific diet similarity. Journal of Fish Biology, 70(3):726–742, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2007:FFRb

[2417] A. D. Nunn, J. P. Harvey, and I. G. Cowx. The food and feeding relationships of larval and 0+ year juvenile fishes in lowland rivers and connected waterbodies. II. Prey selection and the influence of gape. Journal of Fish Biology, 70(3):743–757, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Piko:2007:EHC

[2418] A. A. Piko and S. T. Szedlmayer. Effects of habitat complexity and predator exclusion on the abundance of juvenile red snapper. Journal of Fish Biology, 70(3):758–769, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dantagnan:2007:LFA

[2419] H. Dantagnan, A. S. Bórquez, I. N. Valdebenito, I. A. Salgado, E. A. Serrano, and M. S. Izquierdo. Lipid and fatty acid composition during embryo and larval development of puye Galaxias maculatus Jenyns, 1842, obtained from estuarine, freshwater and cultured populations. Journal of Fish Biology, 70(3):770–781, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grabowski:2007:STS

[2420] T. B. Grabowski and J. J. Isely. Spatial and temporal segregation of spawning habitat by catostomids in the Savannah River, Georgia and South Carolina, U.S.A. Journal of Fish Biology, 70(3):782–798, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsen:2007:SSI

[2421] E. M. Olsen, W. Melle, S. Kaartvedt, J. C. Holst, and K. A. Mork. Spatially structured interactions between a migratory pelagic predator, the Norwegian spring-spawning herring Clupea harengus L., and its zooplankton prey. Journal of Fish Biology, 70(3):799–815, March 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mazur:2007:TPQ

[2422] M. M. Mazur, M. T. Wilson, A. B. Dougherty, A. Buchheister, and D. A. Beauchamp. Temperature and prey quality effects on growth of juvenile walleye pollock Theragra chalcogramma (Pallas): a spatially explicit bioenergetics approach. Journal of Fish Biology, 70(3):816–836, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rikardsen:2007:TVM

[2423] A. H. Rikardsen, J. B. Dempson, P.-A. Amundsen, P. A. Bjørn, B. Finstad, and A. J. Jensen. Temporal variability in marine feeding of sympatric Arctic charr and sea trout. Journal of Fish Biology, 70(3):837–852, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baumann:2007:SEJ

[2424] H. Baumann, M. A. Peck, H.-E. Götze, and A. Temming. Starving early juvenile sprat Sprattus sprattus (L.) in western Baltic coastal waters: evidence from combined field and laboratory observations in August and September 2003. Journal of Fish Biology, 70(3):853–866, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smiley:2007:TLC

[2425] J. E. Smiley and M. A. Drawbridge. Techniques for live capture of deepwater fishes with special emphasis on the design and application of a low-cost hyperbaric chamber. Journal of Fish Biology, 70(4):867–878, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baumgartner:2007:DFH

[2426] L. J. Baumgartner. Diet and feeding habits of predatory fishes upstream and downstream of a low-level weir. Journal of Fish Biology, 70(4):879– 894, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendiola:2007:TEG

[2427] D. Mendiola, L. Ibaibarriaga, and P. Alvarez. Thermal effects on growth and time to starvation during the yolk-sac larval period of Atlantic mackerel Scomber scombrus L. Journal of Fish Biology, 70(4):895–910, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendiola:2007:EDG

[2428] D. Mendiola, P. Alvarez, U. Cotano, and A. Martínezde Murguía. Early development and growth of the laboratory reared north-east Atlantic mackerel Scomber scombrus L. Journal of Fish Biology, 70(4):911–933, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purchase:2007:RBF

[2429] C. F. Purchase, D. J. Hasselman, and L. K. Weir. Relationship between fertilization success and the number of milt donors in rainbow smelt Osmerus mordax (Mitchell): implications for population growth rates. Journal of Fish Biology, 70(4):934–946, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carriquiriborde:2007:NVR

[2430] P. Carriquiriborde, J. C. De Luca, F. N. Dulout, and A. E. Ronco. Nucleolar variation in response to nutritional condition in juvenile pejerrey Odontesthes bonariensis (Valenciennes). Journal of Fish Biology, 70(4): 947–958, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yan:2007:CLH

[2431] Y. Z. Yan and Y. F. Chen. Changes in the life history of Abbottina rivularis in Lake Fuxian. Journal of Fish Biology, 70(4):959–964, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Quesenberry:2007:ITT

[2432] N. J. Quesenberry, P. J. Allen, and J. J. Cech, Jr. The influence of turbidity on three-spined stickleback foraging. Journal of Fish Biology, 70(4):965–972, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shigdar:2007:BCM

[2433] S. Shigdar, D. Cook, P. Jones, A. Harford, and A. C. Ward. Blood cells of Murray cod Maccullochella peelii peelii (Mitchell). Journal of Fish Biology, 70(4):973–980, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:BRa

[2434] Anonymous. Book reviews. Journal of Fish Biology, 70(4):981–986, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:BRb

[2435] Anonymous. Books received. Journal of Fish Biology, 70(4):987, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:FSBb

[2436] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 70(4):988, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsuo:2007:AHS

[2437] A. Y. O. Matsuo and A. L. Val. Acclimation to humic substances prevents whole body sodium loss and stimulates branchial calcium uptake capacity in cardinal tetras Paracheirodon axelrodi (Schultz) subjected to extremely low pH. Journal of Fish Biology, 70(4):989–1000, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gillet:2007:EWT

[2438] C. Gillet and J. P. Dubois. Effect of water temperature and size of females on the timing of spawning of perch Perca fluviatilis L. in Lake Geneva from 1984 to 2003. Journal of Fish Biology, 70(4):1001–1014, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caruso:2007:DWC

[2439] J. H. Caruso, S. W. Ross, K. J. Sulak, and G. R. Sedberry. Deep-water chaunacid and lophiid anglerfishes (Pisces: Lophiiformes) off the southeastern United States. Journal of Fish Biology, 70(4):1015–1026, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flynn:2007:SDA

[2440] S. R. Flynn and T. J. Benfey. Sex differentiation and aspects of gametogenesis in shortnose sturgeon Acipenser brevirostrum Lesueur. Journal of Fish Biology, 70(4):1027–1044, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mingist:2007:RBE

[2441] M. Mingist, T. Kitani, N. Koide, and H. Ueda. Relationship between eyed-egg percentage and levels of cortisol and thyroid hormone in masu salmon Oncorhynchus masou. Journal of Fish Biology, 70(4):1045–1056, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scheurer:2007:IPR

[2442] J. A. Scheurer, B. A. Berejikian, F. P. Thrower, E. R. Ammann, and T. A. Flagg. Innate predator recognition and fright response in related populations of Oncorhynchus mykiss under different predation pressure. Journal of Fish Biology, 70(4):1057–1069, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Butts:2007:SPE

[2443] I. A. E. Butts and M. K. Litvak. Stock and parental effects on embryonic and early larval development of winter flounder Pseudopleuronectes americanus (Walbaum). Journal of Fish Biology, 70(4):1070–1087, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinyoles:2007:LHT

[2444] D. Vinyoles and A. De Sostoa. Life-history traits of the endangered river blenny Salaria fluviatilis (Asso) and their implications for conservation. Journal of Fish Biology, 70(4):1088–1108, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cech:2007:IFD

[2445] M. Čech, J. Kubečka, J. Frouzová, V. Draštík, M. Kratochvíl, and J. Jarošík. Impact of flood on distribution of bathypelagic perch fry layer along the longitudinal profile of large canyon-shaped reservoir. Journal of Fish Biology, 70(4):1109–1119, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eilertsen:2007:CUR

[2446] H. C. Eilertsen, T. Wyatt, and E. Hansen. Can ultraviolet radiation influence cod Gadus morhua L. year class strength: a model study. Journal of Fish Biology, 70(4):1120–1133, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2007:MSP

[2447] D. Cote. Measurements of salmonid population performance in relation to habitat in eastern Newfoundland streams. Journal of Fish Biology, 70(4):1134–1147, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bassett:2007:SSL

[2448] D. K. Bassett, A. G. Carton, and J. C. Montgomery. Saltatory search in a lateral line predator. Journal of Fish Biology, 70(4):1148–1160, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kruesi:2007:DSS

[2449] K. Kruesi and G. Alcaraz. Does a sexually selected trait represent a burden in locomotion? Journal of Fish Biology, 70(4):1161–1170, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haakana:2007:PPV

[2450] H. Haakana, H. Huuskonen, and J. Karjalainen. Predation of perch on vendace larvae: diet composition in an oligotrophic lake and digestion time of the larvae. Journal of Fish Biology, 70(4):1171–1184, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jawad:2007:SMG

[2451] L. A. Jawad and S. M. Al-Jufaili. Scale morphology of greater lizardfish Saurida tumbil (Bloch, 1795) (Pisces: Synodontidae). Journal of Fish Biology, 70(4):1185–1212, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Massee:2007:PSE

[2452] K. C. Massee, J. Kim, B. A. Berejikian, and R. W. Hardy. Prey selection and efficiency of naïve and experienced juvenile sockeye salmon. Journal of Fish Biology, 70(4):1213–1223, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abdoli:2007:IVA

[2453] A. Abdoli, D. Pont, and P. Sagnes. Intrabasin variations in age and growth of bullhead: the effects of temperature. Journal of Fish Biology, 70(4):1224–1238, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2007:SCI

[2454] J. M. Wilson, P. Reis-Santos, A.-V. Fonseca, J. C. Antunes, P. D. Bouça, and J. Coimbra. Seasonal changes in ionoregulatory variables of the glass eel Anguilla anguilla following estuarine entry: comparison with resident elvers. Journal of Fish Biology, 70(4):1239–1253, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2007:VSP

[2455] A. D. Nunn, J. P. Harvey, and I. G. Cowx. Variations in the spawning periodicity of eight fish species in three English lowland rivers over a 6 year period, inferred from 0+ year fish length distributions. Journal of Fish Biology, 70(4):1254–1267, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tytell:2007:BLM

[2456] E. D. Tytell and J. K. Alexander. Bluegill Lepomis macrochirus synchronize pectoral fin motion and opercular pumping. Journal of Fish Biology, 70(4):1268–1279, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2007:FUK

[2457] J. R. Britton and G. D. Davies. First U.K. recording in the wild of the bighead carp Hypophthalmichthys nobilis. Journal of Fish Biology, 70 (4):1280–1282, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luiz:2007:RAS

[2458] O. J. Luiz, Jr., J.-C. Joyeux, and J. L. Gasparini. Rediscovery of Anthias salmopunctatus Lubbock & Edwards, 1981, with comments on its natural history and conservation. Journal of Fish Biology, 70(4):1283–1286, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardie:2007:SRF

[2459] S. A. Hardie, S. B. Pyecroft, L. A. Barmuta, and R. W. G. White. Spawning-related fungal infection of golden galaxias Galaxias auratus (Galaxiidae). Journal of Fish Biology, 70(4):1287–1294, April 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frommen:2007:TBS

[2460] J. G. Frommen, C. Brendler, and T. C. M. Bakker. The tale of the bad stepfather: male three-spined sticklebacks Gasterosteus aculeatus L. recognize foreign eggs in their manipulated nest by egg cues alone. Journal of Fish Biology, 70(4):1295–1301, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:2007:CDR

[2461] C. S. Allen, H. B. Rich, Jr., and T. P. Quinn. Condition-dependent reproductive tactics by large and small anadromous male sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 70(4):1302–1307, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Eb

[2462] Anonymous. Erratum. Journal of Fish Biology, 70(4):1308–1310, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baumann:2007:C

[2463] H. Baumann, M. A. Peck, H.-E. Götze, and A. Temming. Corrigendum. Journal of Fish Biology, 70(4):1311–1316, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2007:A

[2464] F. A. Huntingford, C. Adams, V. A. Braithwaite, S. Kadri, T. G. Pottinger, P. Sandøe, and J. F. Turnbull. Announcement. Journal of Fish Biology, 70(4):1317, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:FSBc

[2465] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 70(4):1318, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ljunggren:2007:IVC

[2466] L. Ljunggren and A. Sandström. Influence of visual conditions on foraging and growth of juvenile fishes with dissimilar sensory physiology. Journal of Fish Biology, 70(5):1319–1334, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tammaro:2007:ILS

[2467] S. Tammaro, S. Filosa, V. Frezza, C. M. Motta, P. Simoniello, and R. Carotenuto. Identification and localization of α— and β-spectrins in oocytes of three Antarctic teleosts: Trematomus bernacchii, Trematomus newnesi (Nototheniidae) and Chionodraco hamatus (Channichthyidae). Journal of Fish Biology, 70(5):1335–1349, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robbins:2007:EVA

[2468] R. L. Robbins. Environmental variables affecting the sexual segregation of great white sharks Carcharodon carcharias at the Neptune Islands, South Australia. Journal of Fish Biology, 70(5):1350–1364, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heermann:2007:TSC

[2469] L. Heermann, P. Beeck, and J. Borcherding. Two size classes of 0+ year perch: is phenotypic plasticity based on food resources? Journal of Fish Biology, 70(5):1365–1377, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Natsumeda:2007:VAF

[2470] T. Natsumeda. Variation in age at first reproduction of male Japanese fluvial sculpin induced by the timing of parental reproduction. Journal of Fish Biology, 70(5):1378–1391, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Juchno:2007:CMH

[2471] D. Juchno, A. Boroń, and J. Gołaszewski. Comparative morphology and histology of the ovaries of the spined loach Cobitis taenia L. and natural allopolyploids of Cobitis (Cobitidae). Journal of Fish Biology, 70(5):1392–1411, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes:2007:COM

[2472] B. V. C. Gomes, R. S. Scarpelli, F. P. Arantes, Y. Sato, N. Bazzoli, and E. Rizzo. Comparative oocyte morphology and early development in three species of trahiras from the São Francisco River basin, Brazil. Journal of Fish Biology, 70(5):1412–1429, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ojwang:2007:ESA

[2473] W. Oweke Ojwang, L. Kaufman, E. Soule, and A. A. Asila. Evidence of stenotopy and anthropogenic influence on carbon source for two major riverine fishes of the Lake Victoria watershed. Journal of Fish Biology, 70(5):1430–1446, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wright:2007:RAC

[2474] K. A. Wright, C. M. C. Woods, B. E. Gray, and P. M. Lokman. Recovery from acute, chronic and transport stress in the pot-bellied seahorse Hippocampus abdominalis. Journal of Fish Biology, 70(5):1447–1457, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2007:FBF

[2475] R. Takahashi, T. Moriwaki, and M. Hori. Foraging behaviour and functional morphology of two scale-eating cichlids from Lake Tanganyika. Journal of Fish Biology, 70(5):1458–1469, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takeda:2007:FAD

[2476] Y. Takeda and M. Tanaka. Freshwater adaptation during larval, juvenile and immature periods of starry flounder Platichthys stellatus, stone flounder Kareius bicoloratus and their reciprocal hybrids. Journal of Fish Biology, 70(5):1470–1483, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spoljaric:2007:YLE

[2477] M. A. Spoljaric and T. E. Reimchen. 10 000 years later: evolution of body shape in Haida Gwaii three-spined stickleback. Journal of Fish Biology, 70(5):1484–1503, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valdivia:2007:MPE

[2478] I. M. Valdivia, R. A. Chávez, and M. E. Oliva. Metazoan parasites of Engraulis ringens as tools for stock discrimination along the Chilean coast. Journal of Fish Biology, 70(5):1504–1511, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pisingan:2007:ASL

[2479] R. S. Pisingan and A. Takemura. Apparent semi-lunar spawning rhythmicity in a brackish cardinalfish. Journal of Fish Biology, 70(5):1512– 1522, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neuhaus:2007:BHS

[2480] H. Neuhaus, M. Van der Marel, N. Caspari, W. Meyer, M.-L. Enss, and D. Steinhagen. Biochemical and histochemical study on the intestinal mucosa of the common carp Cyprinus carpio L., with special consideration of mucin glycoproteins. Journal of Fish Biology, 70(5):1523–1534, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chadwick:2007:LTS

[2481] S. Chadwick, B. Knights, J. L. Thorley, and A. Bark. A long-term study of population characteristics and downstream migrations of the European eel Anguilla anguilla (L.) and the effects of a migration barrier in the Girnock Burn, north-east Scotland. Journal of Fish Biology, 70(5):1535– 1553, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huxham:2007:SIR

[2482] M. Huxham, E. Kimani, J. Newton, and J. Augley. Stable isotope records from otoliths as tracers of fish migration in a mangrove system. Journal of Fish Biology, 70(5):1554–1567, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez-Cabello:2007:DPS

[2483] C. Rodríguez-Cabello, F. Sánchez, and I. Olaso. Distribution patterns and sexual segregations of Scyliorhinus canicula (L.) in the Cantabrian Sea. Journal of Fish Biology, 70(5):1568–1586, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gronkjaer:2007:EHS

[2484] P. Grønkjær, J. D. Carl, T. H. Rasmussen, and K. W. Hansen. Effect of habitat shifts on feeding behaviour and growth of 0 year-group flounder Platichthys flesus (L.) transferred between macroalgae and bare sand habitats. Journal of Fish Biology, 70(5):1587–1605, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2007:OAI

[2485] D. M. Scott, R. W. Wilson, and J. A. Brown. The osmoregulatory ability of the invasive species sunbleak Leucaspius delineatus and topmouth gudgeon Pseudorasbora parva at elevated salinities, and their likely dispersal via brackish waters. Journal of Fish Biology, 70(5):1606–1614, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gregson:2007:SEB

[2486] J. N. S. Gregson and T. Burt de Perera. Shoaling in eyed and blind morphs of the characin Astyanax fasciatus under light and dark conditions. Journal of Fish Biology, 70(5):1615–1619, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caruso:2007:ECM

[2487] J. Caruso and P. H. F. Bor. Egg capsule morphology of Parascyllium variolatum (Duméril, 1853) (Chondrichthyes; Parascylliidae), with notes on oviposition rate in captivity. Journal of Fish Biology, 70(5):1620– 1625, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Temple:2007:ESR

[2488] S. E. Temple. Effect of salinity on the refractive index of water: considerations for archer fish aerial vision. Journal of Fish Biology, 70(5): 1626–1629, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sinnatamby:2007:ACP

[2489] R. N. Sinnatamby, J. E. Bowman, J. B. Dempson, and M. Power. An assessment of de-calcification procedures for δ ^{13}C and δ ^{15}N analysis of yellow perch, walleye and Atlantic salmon scales. Journal of Fish Biology, 70(5):1630–1635, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sharma:2007:ADB

[2490] C. M. Sharma and R. Borgstrøm. Age determination and backcalculation of pike length through use of the metapterygoid bone. Journal of Fish Biology, 70(5):1636–1641, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Roustaian:2007:EYR

[2491] P. Roustaian and M. K. Litvak. Effect of yolk redistribution at different developmental stages on survivorship of zebrafish embryos. Journal of Fish Biology, 70(5):1642–1646, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:FSBd

[2492] Anonymous. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 70(5):1647, May 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopes:2007:VCI

[2493] C. A. Lopes, E. Benedito-Cecilio, and L. A. Martinelli. Variability in the carbon isotope signature of Prochilodus lineatus (Prochilodontidae, Characiformes) a bottom-feeding fish of the Neotropical region. Journal of Fish Biology, 70(6):1649–1659, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodgers:2007:SSD

[2494] E. W. Rodgers, R. L. Earley, and M. S. Grober. Social status determines sexual phenotype in the bi-directional sex changing bluebanded goby Lythrypnus dalli. Journal of Fish Biology, 70(6):1660–1668, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Methratta:2007:OVH

[2495] E. T. Methratta and J. S. Link. Ontogenetic variation in habitat associations for four flatfish species in the Gulf of Maine—Georges Bank region. Journal of Fish Biology, 70(6):1669–1688, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Drake:2007:STD

[2496] P. Drake, A. Borlán, E. González-Ortegón, F. Baldó, C. Vilas, and C. Fernández-Delgado. Spatio-temporal distribution of early life stages of the European anchovy Engraulis encrasicolus L. within a European temperate estuary with regulated freshwater inflow: effects of environmental variables. Journal of Fish Biology, 70(6):1689–1709, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Buston:2007:ELS

[2497] P. M. Buston and M. B. García. An extraordinary life span estimate for the clown anemonefish Amphiprion percula. Journal of Fish Biology, 70 (6):1710–1719, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sepulveda:2007:CEP

[2498] C. A. Sepulveda, K. A. Dickson, L. R. Frank, and J. B. Graham. Cranial endothermy and a putative brain heater in the most basal tuna species, Allothunnus fallai. Journal of Fish Biology, 70(6):1720–1733, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dumas:2007:EFF

[2499] J. Dumas, J. G. Bassenave, M. Jarry, L. Barrière, and S. Glise. Effects of fish farm effluents on egg-to-fry development and survival of brown trout in artificial redds. Journal of Fish Biology, 70(6):1734–1758, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rideout:2007:MCF

[2500] R. M. Rideout and M. J. Morgan. Major changes in fecundity and the effect on population egg production for three species of north-west Atlantic flatfishes. Journal of Fish Biology, 70(6):1759–1779, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Riva-Rossi:2007:IPV

[2501] C. Riva-Rossi, M. A. Pascual, J. A. Babaluk, M. García-Asorey, and N. M. Halden. Intra-population variation in anadromy and reproductive life span in rainbow trout introduced in the Santa Cruz River, Argentina. Journal of Fish Biology, 70(6):1780–1797, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cara:2007:TCI

[2502] B. Cara, F. J. Moyano, J. L. Zambonino, and C. Fauvel. Trypsin and chymotrypsin as indicators of nutritional status of post-weaned sea bass larvae. Journal of Fish Biology, 70(6):1798–1808, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2007:SSC

[2503] W. T. White. Species and size compositions and reproductive biology of rays (Chondrichthyes, Batoidea) caught in target and non-target fisheries in eastern Indonesia. Journal of Fish Biology, 70(6):1809–1837, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bunnell:2007:PRB

[2504] D. B. Bunnell, S. E. Thomas, and R. A. Stein. Prey resources before spawning influence gonadal investment of female, but not male, white crappie. Journal of Fish Biology, 70(6):1838–1854, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petry:2007:ETP

[2505] A. C. Petry, A. A. Agostinho, P. A. Piana, and L. C. Gomes. Effects of temperature on prey consumption and growth in mass of juvenile trahira Hoplias aff. malabaricus (Bloch, 1794). Journal of Fish Biology, 70(6): 1855–1864, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seppala:2007:IBC

[2506] T. Seppälä, J. C. Chubb, E. Niemelä, and E. T. Valtonen. Introduced bullheads Cottus gobio and infection with plerocercoids of Schistocephalus cotti in the Utsjoki, an Arctic river in Finland. Journal of Fish Biology, 70(6):1865–1876, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nissling:2007:FEJ

[2507] A. Nissling, M. Jacobsson, and N. Hallberg. Feeding ecology of juvenile turbot Scophthalmus maximus and flounder Pleuronectes flesus at Gotland, Central Baltic Sea. Journal of Fish Biology, 70(6):1877–1897, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuwamura:2007:SCP

[2508] T. Kuwamura, S. Suzuki, N. Tanaka, E. Ouchi, K. Karino, and Y. Nakashima. Sex change of primary males in a diandric labrid Halichoeres trimaculatus: coexistence of protandry and protogyny within a species. Journal of Fish Biology, 70(6):1898–1906, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huertas:2007:OSC

[2509] M. Huertas, P. C. Hubbard, A. V. M. Canário, and J. Cerdà. Olfactory sensitivity to conspecific bile fluid and skin mucus in the European eel Anguilla anguilla (L.). Journal of Fish Biology, 70(6):1907–1920, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Durieux:2007:CSS

[2510] E. D. H. Durieux, J. F. Marques, P. Sasal, M-L. Bégout, and H. N. Cabral. Comparison of Solea solea macroparasites between two nurserycontinental shelf systems in the Bay of Biscay and the Portuguese coast. Journal of Fish Biology, 70(6):1921–1930, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khoshbavar-Rostami:2007:IRG

[2511] H. A. Khoshbavar-Rostami, M. Soltani, and H. M. D. Hassan. Immune responses of great sturgeon Huso huso to Aeromonas hydrophila bacterin. Journal of Fish Biology, 70(6):1931–1938, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Ec

[2512] Anonymous. Errata. Journal of Fish Biology, 70(6):1939–1941, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wilson:2007:FSB

[2513] J. M. Wilson, P. Reis-Santos, A.-V. Fonseca, J. C. Antunes, P. D. Bouça, and J. Coimbra. The Fisheries Society of the British Isles Annual Symposium 2007: Non-Native Fishes: Integrated Biology of Establishment Success and Dispersal 23–27 July 2007 University of Exeter, Exeter, U.K. Convenor: Anne Brown. Journal of Fish Biology, 70(6):1942, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:IVa

[2514] Anonymous. Index, volume 70. Journal of Fish Biology, 70(6):i–xxiv, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:CVa

[2515] Anonymous. Contents of volume 70. Journal of Fish Biology, 70(6):xxv– xxxii, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Azevedo:2007:CCS

[2516] M. F. C. Azevedo, C. Oliveira, B. G. Pardo, P. Martínez, and F. Foresti. Cytogenetic characterization of six species of flatfishes with comments to karyotype differentiation patterns in pleuronectiformes (Teleostei). Journal of Fish Biology, 70(sa):1–15, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamill:2007:CGD

[2517] R. M. Hamill, S. A. Webb, C. Macías Garcia, J. A. Graves, A. E. Magurran, and M. G. Ritchie. Comparison of genetic diversity at microsatellite loci in near-extinct and non-endangered species of Mexican goodeine fishes and prediction of cross-amplification within the family. Journal of Fish Biology, 70(sa):16–32, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2007:TEJ

[2518] X. G. Wang, R. Bartfai, I. Sleptsova-Freidrich, and L. Orban. The timing and extent of ‘juvenile ovary’ phase are highly variable during zebrafish testis differentiation. Journal of Fish Biology, 70(sa):33–44, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wynne:2007:AHD

[2519] J. W. Wynne, M. T. Cook, B. H. Holmes, and N. G. Elliott. Allelic and haplotypic diversity at the major histocompatibility class II within domesticated Australian Atlantic salmon (Salmo salar L.). Journal of Fish Biology, 70(sa):45–59, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Qi:2007:MDP

[2520] D. Qi, S. Guo, J. Tang, X. Zhao, and J. Liu. Mitochondrial DNA phylogeny of two morphologically enigmatic fishes in the subfamily Schizothoracinae (Teleostei: Cyprinidae) in the Qinghai–Tibetan Plateau. Journal of Fish Biology, 70(sa):60–74, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2007:MDP

[2521] X.-L. Chen, T.-Y. Chiang, H.-D. Lin, H.-S. Zheng, K.-T. Shao, Q. Zhang, and K.-C. Hsu. Mitochondrial DNA phylogeography of Glyptothorax fokiensis and Glyptothorax hainanensis in Asia. Journal of Fish Biology, 70(sa):75–93, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Razpet:2007:GVA

[2522] A. Razpet, S. Sušnik, T. Jug, and A. Snoj. Genetic variation among trout in the River Neretva basin, Bosnia and Herzegovina. Journal of Fish Biology, 70(sa):94–110, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaffney:2007:GSD

[2523] P. M. Gaffney, J. Rupnow, and M. L. Domeier. Genetic similarity of disjunct populations of the giant sea bass Stereolepis gigas. Journal of Fish Biology, 70(sa):111–124, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olafsdottir:2007:PEM

[2524] G. Á. Ólafsdóttir, S. S. Snorrason, and M. G. Ritchie. Parallel evolution? Microsatellite variation of recently isolated marine and freshwater threespined stickleback. Journal of Fish Biology, 70(sa):125–131, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalous:2007:FER

[2525] L. Kalous, V. Šlechtová, Jr., J. Bohlen, M. Petrtýl, and M. Švátora. First European record of Carassius langsdorfii from the Elbe basin. Journal of Fish Biology, 70(sa):132–138, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2007:MGI

[2526] J. D. Anderson and D. L. McDonald. Morphological and genetic investigations of two western Gulf of Mexico menhadens (Brevoortia spp.). Journal of Fish Biology, 70(sa):139–147, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nyingi:2007:RIH

[2527] D. W. Nyingi and J.-F. Agnèse. Recent introgressive hybridization revealed by exclusive mtDNA transfer from Oreochromis leucostictus (Trewavas, 1933) to Oreochromis niloticus (Linnaeus, 1758) in Lake Baringo, Kenya. Journal of Fish Biology, 70(sa):148–154, March 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Margarido:2007:CAT

[2528] V. Pavan Margarido, E. Bellafronte, and O. Moreira-Filho. Cytogenetic analysis of three sympatric Gymnotus (Gymnotiformes, Gymnotidae) species verifies invasive species in the Upper Paraná River basin, Brazil. Journal of Fish Biology, 70(sb):155–164, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tzeng:2007:AMM

[2529] C. H. Tzeng, C. S. Chen, and T. S. Chiu. Analysis of morphometry and mitochondrial DNA sequences from two Trichiurus species in waters of the western North Pacific: taxonomic assessment and population structure. Journal of Fish Biology, 70(sb):165–176, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shen:2007:GDA

[2530] K. N. Shen and W. N. Tzeng. Genetic differentiation among populations of the shortfinned eel Anguilla australis from East Australia and New Zealand. Journal of Fish Biology, 70(sb):177–190, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2007:PGS

[2531] S. Yamamoto, K. Uchida, T. Sato, K. Katsura, and T. Takasawa. Population genetic structure and effective population size of ayu (Plecoglossus altivelis), an amphidromous fish. Journal of Fish Biology, 70(sb):191– 201, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2007:RAF

[2532] W.-J. Chen, R. Ruiz-Carus, and G. Ortí. Relationships among four genera of mojarras (Teleostei: Perciformes: Gerreidae) from the western Atlantic and their tentative placement among percomorph fishes. Journal of Fish Biology, 70(sb):202–218, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Terova:2007:CEA

[2533] G. Terova, S. Rimoldi, V. Chini, R. Gornati, G. Bernardini, and M. Saroglia. Cloning and expression analysis of insulin-like growth factor I and II in liver and muscle of sea bass (Dicentrarchus labrax, L.) during long-term fasting and refeeding. Journal of Fish Biology, 70(sb):219– 233, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brooks:2007:TAF

[2534] C. Brooks, G. Hwang, and N. Maclean. Transgene activity following somatic transgenesis in Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 70(sb):234–247, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Almada:2007:MDC

[2535] V. C. Almada, V. S. Domingues, N. M. Monteiro, F. Almada, and R. S. Santos. Molecular data confirm the validity of the Portuguese blenny (Parablennius ruber, Valenciennes, 1836) and its presence in Western Europe. Journal of Fish Biology, 70(sb):248–254, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Demandt:2007:LGV

[2536] M. Demandt and M. Björklund. Loss of genetic variability in reintroduced roach (Rutilus rutilus) populations. Journal of Fish Biology, 70 (sb):255–261, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

VonDerHeyden:2007:SSG

[2537] S. Von Der Heyden, M. R. Lipinski, and C. A. Matthee. Species-specific genetic markers for identification of early life-history stages of Cape hakes, Merluccius capensis and Merluccius paradoxus in the southern Benguela Current. Journal of Fish Biology, 70(sb):262–268, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Ed

[2538] Anonymous. Erratum. Journal of Fish Biology, 70(sb):269, April 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Skarstein:2007:CMV

[2539] T. H. Skarstein, J.-I. Westgaard, and S.-E. Fevolden. Comparing microsatellite variation in north-east Atlantic cod (Gadus morhua L.) to genetic structuring as revealed by the pantophysin (Pan i) locus. Journal of Fish Biology, 70(sc):271–290, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sun:2007:GVB

[2540] Y. H. Sun, C. X. Xie, W. M. Wang, S. Y. Liu, T. Treer, and M. M. Chang. The genetic variation and biogeography of catostomid fishes based on mitochondrial and nucleic DNA sequences. Journal of Fish Biology, 70(sc):291–309, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OLeary:2007:MVC

[2541] D. B. O’Leary, J. Coughlan, E. Dillane, T. V. McCarthy, and T. F. Cross. Microsatellite variation in cod Gadus morhua throughout its geographic range. Journal of Fish Biology, 70(sc):310–335, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murray:2007:ILP

[2542] H. M. Murray, C. T. Leggiadro, and S. E. Douglas. Immunocytochemical localization of pleurocidin to the cytoplasmic granules of eosinophilic granular cells from the winter flounder gill. Journal of Fish Biology, 70(sc):336–345, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castilho:2007:MMD

[2543] R. Castilho, M. Freitas, G. Silva, J. Fernandez-Carvalho, and R. Coelho. Morphological and mitochondrial DNA divergence validates blackmouth, Galeus melastomus, and Atlantic sawtail catsharks, Galeus atlanticus, as separate species. Journal of Fish Biology, 70(sc):346–358, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simonovic:2007:TSS

[2544] P. Simonović, S. Marić, and V. Nikolić. Trout Salmo spp. complex in Serbia and adjacent regions of the western Balkans: reconstruction of evolutionary history from external morphology. Journal of Fish Biology, 70(sc):359–380, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tonteri:2007:EMB

[2545] A. Tonteri, A. Je. Veselov, S. Titov, J. Lumme, and C. R. Primmer. The effect of migratory behaviour on genetic diversity and population divergence: a comparison of anadromous and freshwater Atlantic salmon Salmo salar. Journal of Fish Biology, 70(sc):381–398, June 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cenadelli:2007:INS

[2546] S. Cenadelli, V. Maran, G. Bongioni, L. Fusetti, P. Parma, and R. Aleandri. Identification of nuclear SNPs in gilthead seabream. Journal of Fish Biology, 70(sc):399–405, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konopinski:2007:PDP

[2547] M. K. Konopiński, A. Amirowicz, and K. Kukuła. Probable direction of the postglacial colonization of rivers on northern slopes of the Carpathian Ridge by Barbus carpathicus (Teleostei: Cyprinidae) evidenced by cline of genetic variation. Journal of Fish Biology, 70(sc):406–415, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:IVb

[2548] Anonymous. Index, volume 70. Journal of Fish Biology, 70(sc):i–vi, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:CVSa

[2549] Anonymous. Contents of volume 70 supplements. Journal of Fish Biology, 70(sc):vii–viii, June 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fitzsimons:2007:IEP

[2550] J. D. Fitzsimons, J. L. Jonas, R. M. Claramunt, B. Williston, G. Williston, J. E. Marsden, B. J. Ellrott, and D. C. Honeyfield. Influence of egg predation and physical disturbance on lake trout Salvelinus namaycush egg mortality and implications for life-history theory. Journal of Fish Biology, 71(1):1–16, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dineen:2007:DPS

[2551] G. Dineen, S. S. C. Harrison, and P. S. Giller. Diet partitioning in sympatric Atlantic salmon and brown trout in streams with contrasting riparian vegetation. Journal of Fish Biology, 71(1):17–38, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kruitwagen:2007:IMA

[2552] G. Kruitwagen, I. Nagelkerken, B. R. Lugendo, H. B. Pratap, and S. E. Wendelaar Bonga. Influence of morphology and amphibious lifestyle on the feeding ecology of the mudskipper Periophthalmus argentilineatus. Journal of Fish Biology, 71(1):39–52, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lombarte:2007:OST

[2553] A. Lombarte and A. Cruz. Otolith size trends in marine fish communities from different depth strata. Journal of Fish Biology, 71(1):53–76, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fagundes:2007:FMC

[2554] T. Fagundes, D. M. Gonçalves, and R. F. Oliveira. Female mate choice and mate search tactics in a sex role reversed population of the peacock blenny Salaria pavo (Risso, 1810). Journal of Fish Biology, 71(1):77– 89, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gaye-Siessegger:2007:SLF

[2555] J. Gaye-Siessegger, U. Focken, Hj Abel, and K. Becker. Starvation and low feeding levels result in an enrichment of ^{13}C in lipids and ^{15}N in protein of Nile tilapia Oreochromis niloticus L. Journal of Fish Biology, 71(1):90–100, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Slavik:2007:DSB

[2556] O. Slavík, P. Horký, L. Bartoš, J. Kolářová, and T. Randák. Diurnal and seasonal behaviour of adult and juvenile European catfish as determined by radio-telemetry in the River Berounka, Czech Republic. Journal of Fish Biology, 71(1):101–114, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooper:2007:GIU

[2557] C. A. Cooper and N. R. Bury. The gills as an important uptake route for the essential nutrient iron in freshwater rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 71(1):115–128, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brinker:2007:ENI

[2558] A. Brinker and R. Hamers. Evidence for negative impact of plerocercoid infection of Triaenophorus nodulosus on Perca fluviatilis L. stock in Upper Lake Constance, a water body undergoing rapid reoligotrophication. Journal of Fish Biology, 71(1):129–147, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raine:2007:IRL

[2559] J. C. Raine, K. Hua, D. P. Bureau, M. M. Vijayan, and J. F. Leatherland. Influence of ration level and rearing temperature on hepatic GHR1 and 2, and hepatic and intestinal TRα and TRβ gene expression in late stages of rainbow trout embryos. Journal of Fish Biology, 71(1):148–162, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vila-Gispert:2007:MVP

[2560] A. Vila-Gispert, M. G. Fox, L. Zamora, and R. Moreno-Amich. Morphological variation in pumpkinseed Lepomis gibbosus introduced into Iberian lakes and reservoirs; adaptations to habitat type and diet? Journal of Fish Biology, 71(1):163–181, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ebner:2007:FYO

[2561] B. C. Ebner, J. D. Thiem, and M. Lintermans. Fate of 2 year-old, hatchery-reared trout cod Maccullochella macquariensis (Percichthyidae) stocked into two upland rivers. Journal of Fish Biology, 71(1):182–199, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:2007:PSB

[2562] C. J. Robinson, S. Gómez-Aguirre, and J. Gómez-Gutiérrez. Pacific sardine behaviour related to tidal current dynamics in Bahía magdalena, México. Journal of Fish Biology, 71(1):200–218, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andersen:2007:HPF

[2563] N. G. Andersen and J. E. Beyer. How are prey fishes of multiple meals evacuated from the stomach of a piscivorous fish? Journal of Fish Biology, 71(1):219–234, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Whiteman:2007:EDB

[2564] E. A. Whiteman, I. M. Côté, and J. D. Reynolds. Ecological differences between hamlet (Hypoplectrus: Serranidae) colour morphs: betweenmorph variation in diet. Journal of Fish Biology, 71(1):235–244, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2007:SGS

[2565] N. Jonsson and B. Jonsson. Sea growth, smolt age and age at sexual maturation in Atlantic salmon. Journal of Fish Biology, 71(1):245–252, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Turesson:2007:ELS

[2566] H. Turesson and P. Domenici. Escape latency is size independent in grey mullet. Journal of Fish Biology, 71(1):253–259, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pinder:2007:EAD

[2567] A. C. Pinder, W. D. Riley, A. T. Ibbotson, and W. R. C. Beaumont. Evidence for an autumn downstream migration and the subsequent estuarine residence of 0+ year juvenile Atlantic salmon Salmo salar L., in England. Journal of Fish Biology, 71(1):260–264, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yanase:2007:QAB

[2568] K. Yanase and T. Arimoto. Quantitative assessment of behavioural response to the netting panel in the jack mackerel Trachurus japonicus. Journal of Fish Biology, 71(1):265–271, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wressnig:2007:FPT

[2569] A. Wressnig and D. J. Booth. Feeding preferences of two seagrass grazing monacanthid fishes. Journal of Fish Biology, 71(1):272–278, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moritz:2007:ABB

[2570] T. Moritz and K. E. Linsenmair. The air-breathing behaviour of Brevimyrus niger (Osteoglossomorpha, Mormyridae). Journal of Fish Biology, 71(1):279–283, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Figenschou:2007:FAC

[2571] L. Figenschou, G. Rudolfsen, and I. Folstad. Female Arctic charr do not show apparent benefits from exposing their eggs to sperm from dominant males. Journal of Fish Biology, 71(1):284–289, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Eliason:2007:VHP

[2572] E. J. Eliason, A. Kiessling, A. Karlsson, B. Djordjevic, and A. P. Farrell. Validation of the hepatic portal vein cannulation technique using Atlantic salmon Salmo salar L. Journal of Fish Biology, 71(1):290–297, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rushbrook:2007:SLR

[2573] B. J. Rushbrook, I. Katsiadaki, and I. Barber. Spiggin levels are reduced in male sticklebacks infected with Schistocephalus solidus. Journal of Fish Biology, 71(1):298–303, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Spence:2007:DGR

[2574] R. Spence, M. K. Fatema, S. Ellis, Z. F. Ahmed, and C. Smith. Diet, growth and recruitment of wild zebrafish in Bangladesh. Journal of Fish Biology, 71(1):304–309, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:BRc

[2575] Anonymous. Book reviews. Journal of Fish Biology, 71(1):310–313, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:BRd

[2576] Anonymous. Books received. Journal of Fish Biology, 71(1):314, July 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hurst:2007:CCW

[2577] T. P. Hurst. Causes and consequences of winter mortality in fishes. Journal of Fish Biology, 71(2):315–345, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Renones:2007:AGD

[2578] O. Reñones, C. Piñeiro, X. Mas, and R. Goñi. Age and growth of the dusky grouper Epinephelus marginatus (Lowe 1834) in an exploited population of the western Mediterranean Sea. Journal of Fish Biology, 71 (2):346–362, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dutil:2007:BCU

[2579] J.-D. Dutil, E.-L. Sylvestre, L. Gamache, R. Larocque, and H. Guderley. Burst and coast use, swimming performance and metabolism of Atlantic cod Gadus morhua in sub-lethal hypoxic conditions. Journal of Fish Biology, 71(2):363–375, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beldade:2007:PLD

[2580] R. Beldade, T. Pedro, and E. J. Gonçalves. Pelagic larval duration of 10 temperate cryptobenthic fishes. Journal of Fish Biology, 71(2):376– 382, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tavares-Dias:2007:HBR

[2581] M. Tavares-Dias and F. R. Moraes. Haematological and biochemical reference intervals for farmed channel catfish. Journal of Fish Biology, 71(2):383–388, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2007:NTE

[2582] A. P. Farrell, D. L. Simonot, R. S. Seymour, and T. D. Clark. A novel technique for estimating the compact myocardium in fishes reveals surprising results for an athletic air-breathing fish, the Pacific tarpon. Journal of Fish Biology, 71(2):389–398, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ochi:2007:AHG

[2583] H. Ochi, H. Hata, and M. Hori. Advantage of harmlessness in gaining preferential access to a heterospecific territory. Journal of Fish Biology, 71(2):399–410, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Torisawa:2007:SBR

[2584] S. Torisawa, T. Takagi, H. Fukuda, Y. Ishibashi, Y. Sawada, T. Okada, S. Miyashita, K. Suzuki, and T. Yamane. Schooling behaviour and retinomotor response of juvenile Pacific bluefin tuna Thunnus orientalis under different light intensities. Journal of Fish Biology, 71(2):411–420, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Licandeo:2007:GVL

[2585] R. Licandeo and F. T. Cerna. Geographic variation in life-history traits of the endemic kite skate Dipturus chilensis (Batoidea: Rajidae), along its distribution in the fjords and channels of southern Chile. Journal of Fish Biology, 71(2):421–440, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Persson:2007:FAS

[2586] M. E. Persson, P. Larsson, and P. Stenroth. Fractionation of δ ^{15}N and δ 13C for Atlantic salmon and its intestinal cestode Eubothrium crassum. Journal of Fish Biology, 71(2):441–452, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magnussen:2007:ICG

[2587] E. Magnussen. Interpopulation comparison of growth patterns of 14 fish species on Faroe Bank: are all fishes on the bank fast-growing? Journal of Fish Biology, 71(2):453–475, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Henning:2007:USF

[2588] J. A. Henning, R. E. Gresswell, and I. A. Fleming. Use of seasonal freshwater wetlands by fishes in a temperate river floodplain. Journal of Fish Biology, 71(2):476–492, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petersson:2007:CBT

[2589] E. Petersson and T. Järvi. Characteristics of brown trout males influence growth and survival of their offspring. Journal of Fish Biology, 71(2): 493–509, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fregeneda-Grandes:2007:FIC

[2590] J. M. Fregeneda-Grandes, F. Rodríguez-Cadenas, and J. M. Aller-Gancedo. Fungi isolated from cultured eggs, alevins and broodfish of brown trout in a hatchery affected by saprolegniosis. Journal of Fish Biology, 71(2):510–518, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thilakaratne:2007:EPP

[2591] I. D. S. I. P. Thilakaratne, J. D. McLaughlin, and D. J. Marcogliese. Effects of pollution and parasites on biomarkers of fish health in spottail shiners Notropis hudsonius (Clinton). Journal of Fish Biology, 71(2): 519–538, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fuda:2007:EEF

[2592] K. M. Fuda, B. M. Smith, M. P. Lesser, B. J. Legare, H. R. Breig, R. B. Stack, and D. L. Berlinsky. The effects of environmental factors on rainbow smelt Osmerus mordax embryos and larvae. Journal of Fish Biology, 71(2):539–549, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Olsvik:2007:ECR

[2593] P. A. Olsvik, B. E. Torstensen, and M. H. G. Berntssen. Effects of complete replacement of fish oil with plant oil on gastrointestinal cell death, proliferation and transcription of eight genes’ encoding proteins responding to cellular stress in Atlantic salmon Salmo salar L. Journal of Fish Biology, 71(2):550–568, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Murphy:2007:LDA

[2594] B. F. Murphy, J. M. Leis, and K. D. Kavanagh. Larval development of the Ambon damselfish Pomacentrus amboinensis, with a summary of pomacentrid development. Journal of Fish Biology, 71(2):569–584, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sogabe:2007:FDG

[2595] A. Sogabe and Y. Yanagisawa. The function of daily greetings in a monogamous pipefish Corythoichthys haematopterus. Journal of Fish Biology, 71(2):585–595, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vila:2007:ACS

[2596] S. Vila, M. Munoz, M. Sabat, and M. Casadevall. Annual cycle of stored spermatozoa within the ovaries of Helicolenus dactylopterus dactylopterus (Teleostei, Scorpaenidae). Journal of Fish Biology, 71(2):596–609, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abdallah:2007:FRC

[2597] A. Ben Abdallah, J. Ben Souissi, H. Méjri, C. Capapé, and D. Golani. First record of Cephalopholis taeniops (Valenciennes) in the Mediterranean Sea. Journal of Fish Biology, 71(2):610–614, August 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strand:2007:EGS

[2598] Å. Strand, A. Alanärä, and C. Magnhagen. Effect of group size on feed intake, growth and feed efficiency of juvenile perch. Journal of Fish Biology, 71(2):615–619, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caprioli:2007:FRF

[2599] R. Caprioli, C. Costa, E. Cataldi, C. Boglione, and S. Cataudella. First record of female germ cells in the testis of a wild Mediterranean northern bluefin tuna Thunnus thynnus L. 1758. Journal of Fish Biology, 71(2): 620–622, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2007:ERB

[2600] R. W. Blake, P. Y. L. Kwok, and K. H. S. Chan. The energetics of rheotactic behaviour in Pterygoplichthys spp. (Teleostei: Loricariidae). Journal of Fish Biology, 71(2):623–627, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:ICE

[2601] Anonymous. International Conference on Ecohydrological Processes and Sustainability: Floodplain Management: Opportunities and Concepts for Water Hazard Mitigation, and Ecological and Socioeconomic Sustainability 19–23 May 2008 Łodz, Poland. Journal of Fish Biology, 71 (2):628, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russo:2007:CBS

[2602] T. Russo, C. Costa, and S. Cataudella. Correspondence between shape and feeding habit changes throughout ontogeny of gilthead sea bream Sparus aurata L., 1758. Journal of Fish Biology, 71(3):629–656, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:2007:DFT

[2603] R. Johnston, M. Sheaves, and B. Molony. Are distributions of fishes in tropical estuaries influenced by turbidity over small spatial scales? Journal of Fish Biology, 71(3):657–671, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Urbach:2007:SVA

[2604] D. Urbach, D. Bittner, T. L. Lenz, D. Bernet, T. Wahli, and C. Wedekind. Sperm velocity in an Alpine whitefish: effects of age, size, condition, fluctuating asymmetry and gonad abnormalities. Journal of Fish Biology, 71(3):672–683, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peyronnet:2007:LBP

[2605] A. Peyronnet, K. D. Friedland, N. Ó Maoileidigh, M. Manning, and W. R. Poole. Links between patterns of marine growth and survival of Atlantic salmon Salmo salar, L. Journal of Fish Biology, 71(3):684–700, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morgan:2007:OHA

[2606] S. K. Morgan and A. C. J. Vincent. The ontogeny of habitat associations in the tropical tiger tail seahorse Hippocampus comes Cantor, 1850. Journal of Fish Biology, 71(3):701–724, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacIntosh:2007:ELI

[2607] K. E. MacIntosh and J. Duston. Effect of light intensity and eye development on prey capture by larval striped bass Morone saxatilis. Journal of Fish Biology, 71(3):725–736, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergstrom:2007:WWT

[2608] C. A. Bergstrom and A. R. Palmer. Which way to turn? Effect of direction of body asymmetry on turning and prey strike orientation in starry flounder Platichthys stellatus (Pallas) (Pleuronectidae). Journal of Fish Biology, 71(3):737–748, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2007:TEA

[2609] B. E. Smith, T. J. Ligenza, F. P. Almeida, and J. S. Link. The trophic ecology of Atlantic cod: insights from tri-monthly, localized scales of sampling. Journal of Fish Biology, 71(3):749–762, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dougherty:2007:IDG

[2610] A. B. Dougherty, K. M. Bailey, and K. L. Mier. Interannual differences in growth and hatch date distributions of age-0 year walleye pollock Theragra chalcogramma (Pallas) sampled from the Shumagin Islands region of the Gulf of Alaska, 1985 — 2001. Journal of Fish Biology, 71(3):763–780, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Canavate:2007:ELI

[2611] J. P. Cañavate, A. Prieto, R. Zerolo, M. Sole, C. Sarasquete, and C. Fernández-Díaz. Effects of light intensity and addition of carotene rich Dunaliella salina live cells on growth and antioxidant activity of Solea senegalensis Kaup (1858) larval and metamorphic stages. Journal of Fish Biology, 71(3):781–794, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pears:2007:RBL

[2612] R. J. Pears, J. H. Choat, B. D. Mapstone, and G. A. Begg. Reproductive biology of a large, aggregation-spawning serranid, Epinephelus fuscoguttatus (Forsskål): management implications. Journal of Fish Biology, 71 (3):795–817, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hogan:2007:LDM

[2613] Z. Hogan, I. G. Baird, R. Radtke, and M. J. Vander Zanden. Long distance migration and marine habitation in the tropical Asian catfish, Pangasius krempfi. Journal of Fish Biology, 71(3):818–832, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reber:2007:ITP

[2614] C. M. Reber and W. A. Bennett. The influence of thermal parameters on the acclimation responses of pinfish Lagodon rhomboides exposed to static and decreasing low temperatures. Journal of Fish Biology, 71(3): 833–841, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Webster:2007:EDS

[2615] S. J. Webster, L. M. Dill, and J. S. Korstrom. The effects of depth and salinity on juvenile Chinook salmon Oncorhynchus tshawytscha (Walbaum) habitat choice in an artificial estuary. Journal of Fish Biology, 71 (3):842–851, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farley:2007:HCJ

[2616] J. H. Farley and J. S. Gunn. Historical changes in juvenile southern bluefin tuna Thunnus maccoyii growth rates based on otolith measurements. Journal of Fish Biology, 71(3):852–867, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flajshans:2007:ATT

[2617] M. Flajšhans, K. Kohlmann, and P. Ráb. Autotriploid tench Tinca tinca (L.) larvae obtained by fertilization of eggs previously subjected to postovulatory ageing in vitro and in vivo. Journal of Fish Biology, 71(3): 868–876, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2007:SRS

[2618] J. N. S. Santos, F. G. Araújo, M. A. Silva, and R. M. Vasconcellos. Sex ratio and sexual dimorphism of the anchovy Anchoa januaria (Actinopterygii, Engraulidae) in a tropical bay in south-eastern Brazil. Journal of Fish Biology, 71(3):877–888, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fava:2007:ESS

[2619] D. Fava and M. Toledo-Piza. Egg surface structure in the annual fishes Simpsonichthys (subgenera Ophthalmolebias and Xenurolebias) and Nematolebias (Teleostei: Cyprinodontiformes: Rivulidae): variability and phylogenetic significance. Journal of Fish Biology, 71(3):889–907, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franco:2007:NRA

[2620] M. A. L. Franco, P. A. S. Costa, and A. C. Braga. New records of Aphyonidae (Teleostei: Ophidiiformes) from the south-western Atlantic. Journal of Fish Biology, 71(3):908–912, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2007:CMI

[2621] C. H. Chen, C. C. Wu, K. T. Shao, and J. S. Yang. Chorion microstructure for identifying five fish eggs of Apogonidae. Journal of Fish Biology, 71(3):913–919, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bass:2007:MCG

[2622] C. Santiago Bass, S. Khan, and J. S. Weis. Morphological changes to the gills of killifish associated with severe parasite infection. Journal of Fish Biology, 71(3):920–925, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanson:2007:HTM

[2623] L. M. Hanson and A. P. Farrell. The hypoxic threshold for maximum cardiac performance in rainbow trout Oncorhynchus mykiss (Walbaum) during simulated exercise conditions at 18^{o}C. Journal of Fish Biology, 71 (3):926–932, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ellis:2007:OPS

[2624] J. R. Ellis. Occurrence of pelagic stingray Pteroplatytrygon violacea (Bonaparte, 1832) in the North Sea. Journal of Fish Biology, 71(3): 933–937, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Northcote:2007:SSS

[2625] T. G. Northcote and C. J. Bull. Successful shoreline spawning of rainbow trout in two Canadian alpine lakes. Journal of Fish Biology, 71(3):938– 941, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Ee

[2626] Anonymous. Erratum. Journal of Fish Biology, 71(3):942, September 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2007:FAR

[2627] B. Jonsson, N. Jonsson, and L. P. Hansen. Factors affecting river entry of adult Atlantic salmon in a small river. Journal of Fish Biology, 71 (4):943–956, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Steinhausen:2007:RBC

[2628] M. F. Steinhausen, J. F. Steffensen, and N. G. Andersen. The relationship between caudal differential pressure and activity of Atlantic cod: a potential method to predict oxygen consumption of free-swimming fish. Journal of Fish Biology, 71(4):957–969, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williams:2007:SIP

[2629] A. J. Williams, B. D. Mapstone, and C. R. Davies. Spatial and interannual patterns in growth of an exploited coral-reef fish. Journal of Fish Biology, 71(4):970–992, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gregory:2007:DFP

[2630] M. P. M. Gregory, A. R. Davis, and D. J. Ayre. Diet and feeding periodicity of Cox’s gudgeon Gobiomorphus coxii (Krefft) in a south-eastern Australian stream. Journal of Fish Biology, 71(4):993–1006, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harahush:2007:CBE

[2631] B. K. Harahush, A. B. P. Fischer, and S. P. Collin. Captive breeding and embryonic development of Chiloscyllium punctatum Muller & Henle, 1838 (Elasmobranchii: Hemiscyllidae). Journal of Fish Biology, 71(4): 1007–1022, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karplus:2007:SFL

[2632] I. Karplus, B. Zion, L. Rosenfeld, Y. Grinshpun, T. Slosman, Z. Goshen, and A. Barki. Social facilitation of learning in mixed-species schools of common carp Cyprinus carpio L. and Nile tilapia Oreochromis niloticus (L.). Journal of Fish Biology, 71(4):1023–1034, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamer:2007:CSV

[2633] P. A. Hamer and G. P. Jenkins. Comparison of spatial variation in otolith chemistry of two fish species and relationships with water chemistry and otolith growth. Journal of Fish Biology, 71(4):1035–1055, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lee:2007:SSV

[2634] O. Lee, B. S. Danilowicz, and R. D. M. Nash. Small-scale variability in growth and condition of dab Limanda limanda (L.) larvae in relation to an Irish Sea tidal-mixing front off the west coast of the Isle of Man. Journal of Fish Biology, 71(4):1056–1068, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heggenes:2007:SSH

[2635] J. Heggenes and S. J. Saltveit. Summer stream habitat partitioning by sympatric Arctic charr, Atlantic salmon and brown trout in two subarctic rivers. Journal of Fish Biology, 71(4):1069–1081, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2007:ICN

[2636] Y. Wang, W. Hu, W. y. Liu, Y. p. Wang, and Z. y. Zhu. Identification and characterization of a novel splice variant of gonadotropin α subunit in the common carp Cyprinus carpio. Journal of Fish Biology, 71(4): 1082–1094, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thibault:2007:YSA

[2637] I. Thibault, J. J. Dodson, and F. Caron. Yellow-stage American eel movements determined by microtagging and acoustic telemetry in the st Jean River watershed, Gaspé, Quebec, Canada. Journal of Fish Biology, 71(4):1095–1112, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duponchelle:2007:ERL

[2638] F. Duponchelle, F. Lino, N. Hubert, J. Panfili, J.-F. Renno, E. Baras, J. P. Torrico, R. Dugue, and J. Nuñez. Environment-related life-history trait variations of the red-bellied piranha Pygocentrus nattereri in two river basins of the Bolivian Amazon. Journal of Fish Biology, 71(4): 1113–1134, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Meyer:2007:STD

[2639] W. Meyer, U. Seegers, and R. Stelzer. Sulphur, thiols, and disulphides in the fish epidermis, with remarks on keratinization. Journal of Fish Biology, 71(4):1135–1144, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farley:2007:JSS

[2640] E. V. Farley, Jr., J. M. Murphy, M. Adkison, and L. Eisner. Juvenile sockeye salmon distribution, size, condition and diet during years with warm and cool spring sea temperatures along the eastern Bering Sea shelf. Journal of Fish Biology, 71(4):1145–1158, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collins:2007:RMP

[2641] A. B. Collins, M. R. Heupel, and P. J. Motta. Residence and movement patterns of cownose rays Rhinoptera bonasus within a south-west Florida estuary. Journal of Fish Biology, 71(4):1159–1178, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breau:2007:ASA

[2642] C. Breau, R. A. Cunjak, and G. Bremset. Age-specific aggregation of wild juvenile Atlantic salmon Salmo salar at cool water sources during high temperature events. Journal of Fish Biology, 71(4):1179–1191, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alam:2007:EDD

[2643] M. A. Alam and M. Nakamura. Efferent duct differentiation during female-to-male sex change in honeycomb grouper Epinephelus merra. Journal of Fish Biology, 71(4):1192–1202, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2007:EUM

[2644] J. Li, K. H. S. Chan, R. W. Blake, and P. Y. L. Kwok. Evaluation using multilayer perception neural networks: a case study of undulatory median fin swimming in the knifefish Xenomystus nigri. Journal of Fish Biology, 71(4):1203–1207, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Izzo:2007:ITO

[2645] C. Izzo, K. Rodda, and T. Bolton. Incorporation time of oxytetracycline into calcified structures of the elasmobranch Heterodontus portusjacksoni. Journal of Fish Biology, 71(4):1208–1214, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vacchi:2007:PGP

[2646] M. Vacchi, S. Bussotti, A. M. Miglietta, and P. Guidetti. Presence of the Guinean puffer Sphoeroides marmoratus (Lowe, 1838) in the Mediterranean Sea. Journal of Fish Biology, 71(4):1215–1219, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gisbert:2007:FRP

[2647] E. Gisbert and M. A. López. First record of a population of the exotic mummichog Fundulus heteroclitus (L., 1766) in the Mediterranean Sea basin (Ebro River delta). Journal of Fish Biology, 71(4):1220–1224, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parzefall:2007:PBS

[2648] J. Parzefall, C. Kraus, M. Tobler, and M. Plath. Photophilic behaviour in surface- and cave-dwelling Atlantic mollies Poecilia mexicana (Poeciliidae). Journal of Fish Biology, 71(4):1225–1231, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scenna:2007:AHM

[2649] L. B. Scenna, J. M. Díazde Astarloa, and M. B. Cousseau. Abnormal hermaphroditism in the multispine skate Bathyraja multispinis (Chondrichthyes, Rajidae). Journal of Fish Biology, 71(4):1232–1237, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mincarone:2007:OPP

[2650] M. M. Mincarone, R. A. Bernardes, and F. V. Peppes. Occurrence of Pleuroscopus pseudodorsalis Barnard, 1927 (Uranoscopidae) near Rio Grande Plateau, western South Atlantic. Journal of Fish Biology, 71(4): 1238–1240, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2007:IFS

[2651] J. Simon. Identification of fish species by an egg identification key: an alternative to the incubation of eggs and rearing of larvae for species recognition. Journal of Fish Biology, 71(4):1241–1248, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:BRe

[2652] Anonymous. Book reviews. Journal of Fish Biology, 71(4):1249–1253, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:BRf

[2653] Anonymous. Books received. Journal of Fish Biology, 71(4):1254, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Ef

[2654] Anonymous. Erratum. Journal of Fish Biology, 71(4):1255, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Soria:2007:SPO

[2655] M. Soria, P. Freon, and P. Chabanet. Schooling properties of an obligate and a facultative fish species. Journal of Fish Biology, 71(5):1257–1269, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McBride:2007:SDR

[2656] R. S. McBride and M. R. Johnson. Sexual development and reproductive seasonality of hogfish (Labridae: Lachnolaimus maximus), an hermaphroditic reef fish. Journal of Fish Biology, 71(5):1270–1292, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hofstede:2007:LBM

[2657] R. Ter Hofstede, M. Dickey-Collas, I. T. Mantingh, and A. Wague. The link between migration, the reproductive cycle and condition of Sardinella aurita off Mauritania, north-west Africa. Journal of Fish Biology, 71(5):1293–1302, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernandez:2007:CGT

[2658] L. Fernández and G. Miranda. A catfish of the genus Trichomycterus from a thermal stream in southern South America (Teleostei, Siluriformes, Trichomycteridae), with comments on relationships within the genus. Journal of Fish Biology, 71(5):1303–1316, November 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kristoffersen:2007:GRR

[2659] J. B. Kristoffersen. Growth rate and relative otolith size in populations of adult Müller’s pearlside Maurolicus muelleri. Journal of Fish Biology, 71(5):1317–1330, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hindell:2007:DPU

[2660] J. S. Hindell. Determining patterns of use by black bream Acanthopagrus butcheri (Munro, 1949) of re-established habitat in a south-eastern Australian estuary. Journal of Fish Biology, 71(5):1331–1346, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walker:2007:RLG

[2661] S. P. W. Walker, C. A. Ryen, and M. I. McCormick. Rapid larval growth predisposes sex change and sexual size dimorphism in a protogynous hermaphrodite, Parapercis snyderi Jordan & Starks 1905. Journal of Fish Biology, 71(5):1347–1357, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pecinkova:2007:RBD

[2662] M. Pečínková, L. A. Vøllestad, B. Koubková, J. Huml, P. Jurajda, and M. Gelnar. The relationship between developmental instability of gudgeon Gobio gobio and abundance or morphology of its ectoparasite Paradiplozoon homoion (Monogenea). Journal of Fish Biology, 71(5): 1358–1370, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imsland:2007:PGE

[2663] A. K. Imsland, A. Foss, R. Koedijk, A. Folkvord, S. O. Stefansson, and T. M. Jonassen. Persistent growth effects of temperature and photoperiod in Atlantic cod Gadus morhua. Journal of Fish Biology, 71(5): 1371–1382, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2007:ABS

[2664] S. B. Martin. Association behaviour of the self-fertilizing Kryptolebias marmoratus (Poey): the influence of microhabitat use on the potential for a complex mating system. Journal of Fish Biology, 71(5):1383–1392, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rios:2007:LHA

[2665] F. S. Rios, L. Donatti, M. N. Fernandes, A. L. Kalinin, and F. T. Rantin. Liver histopathology and accumulation of melano-macrophage centres in Hoplias malabaricus after long-term food deprivation and re-feeding. Journal of Fish Biology, 71(5):1393–1406, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2007:ASC

[2666] S. E. Moore, S. A. Hesp, N. G. Hall, and I. C. Potter. Age and size compositions, growth and reproductive biology of the breaksea cod Epinephelides armatus, a gonochoristic serranid. Journal of Fish Biology, 71(5):1407–1429, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlson:2007:GRD

[2667] S. M. Carlson, A. P. Hendry, and B. H. Letcher. Growth rate differences between resident native brook trout and non-native brown trout. Journal of Fish Biology, 71(5):1430–1447, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tudorache:2007:SCE

[2668] C. Tudorache, R. Blust, and G. De Boeck. Swimming capacity and energetics of migrating and non-migrating morphs of three-spined stickleback Gasterosteus aculeatus L. and their ecological implications. Journal of Fish Biology, 71(5):1448–1456, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shepherd:2007:DOS

[2669] S. A. Shepherd and J. B. Brook. Distribution and ontogenetic shifts in habitat and abundance of the temperate western blue groper, Achoerodus gouldii (Richardson). Journal of Fish Biology, 71(5):1457–1478, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:2007:LHA

[2670] A. D. Clarke, A. Lewis, K. H. Telmer, and J. M. Shrimpton. Life history and age at maturity of an anadromous smelt, the eulachon Thaleichthys pacificus (Richardson). Journal of Fish Biology, 71(5):1479–1493, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knight:2007:RBE

[2671] J. T. Knight, G. L. Butler, P. S. Smith, and R. N. E. Wager. Reproductive biology of the endangered Oxleyan pygmy perch Nannoperca oxleyana Whitley. Journal of Fish Biology, 71(6):1494–1511, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

White:2007:CCR

[2672] W. T. White. Catch composition and reproductive biology of whaler sharks (Carcharhiniformes: Carcharhinidae) caught by fisheries in Indonesia. Journal of Fish Biology, 71(6):1512–1540, November 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Solis:2007:CSP

[2673] J. Casas Solis, A. Santerre, M. I. Girón Pérez, R. Reynoso Orozco, and G. Zaitseva. A comparative study of phagocytic activity and lymphoproliferative response in five varieties of tilapia Oreochromis spp. Journal of Fish Biology, 71(6):1541–1545, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Imbert:2007:EVI

[2674] H. Imbert, L. Beaulaton, C. Rigaud, and P. Elie. Evaluation of visible implant elastomer as a method for tagging small European eels. Journal of Fish Biology, 71(6):1546–1554, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Belmaker:2007:GDR

[2675] J. Belmaker, O. Polak, N. Shashar, and Y. Ziv. Geographic divergence in the relationship between Paragobiodon echinocephalus and its obligate coral host. Journal of Fish Biology, 71(6):1555–1561, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conway:2007:SDW

[2676] K. W. Conway and R. Britz. Sexual dimorphism of the Weberian apparatus and pectoral girdle in Sundadanio axelrodi (Ostariophysi: Cyprinidae). Journal of Fish Biology, 71(6):1562–1570, November 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miranda:2007:CFL

[2677] L. A. Miranda, C. A. Strüssmann, L. G. Guilgur, P. H. Strobl-Mazzulla, and G. M. Somoza. Cloning of FSH-β, LH-β and glycoprotein hormone α subunits in pejerrey Odontesthes bonariensis (Valenciennes): expression profile and relationship with GnRH expression and plasma sex steroid levels in male fish. Journal of Fish Biology, 71(6):1571–1589, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2007:CBB

[2678] C. Brown, F. Jones, and V. A. Braithwaite. Correlation between boldness and body mass in natural populations of the poeciliid Brachyrhaphis episcopi. Journal of Fish Biology, 71(6):1590–1601, December 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2007:FBP

[2679] S. W. Griffiths, A. F. Ojanguren, J. E. Orpwood, A. E. Magurran, and J. D. Armstrong. Familiarity-biased patterns of association shift with time among European minnows. Journal of Fish Biology, 71(6):1602– 1612, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Waltham:2007:AWD

[2680] N. J. Waltham and R. M. Connolly. Artificial waterway design affects fish assemblages in urban estuaries. Journal of Fish Biology, 71(6):1613–1629, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tudorache:2007:LFI

[2681] C. Tudorache, P. Viaenen, R. Blust, and G. De Boeck. Longer flumes increase critical swimming speeds by increasing burst–glide swimming duration in carp Cyprinus carpio, L. Journal of Fish Biology, 71(6): 1630–1638, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:2007:DIA

[2682] N. A. Hvidsten, B. Finstad, F. Kroglund, B. O. Johnsen, R. Strand, J. V. Arnekleiv, and P. A. Bjørn. Does increased abundance of sea lice influence survival of wild Atlantic salmon post-smolt? Journal of Fish Biology, 71(6):1639–1648, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wells:2007:ABM

[2683] R. M. G. Wells, J. Baldwin, R. S. Seymour, K. A. Christian, and A. P. Farrell. Air breathing minimizes post-exercise lactate load in the tropical Pacific tarpon, Megalops cyprinoides Broussonet 1782 but oxygen debt is repaid by aquatic breathing. Journal of Fish Biology, 71(6):1649–1661, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feeley:2007:EOU

[2684] M. W. Feeley, D. D. Benetti, and J. S. Ault. Elevated oxygen uptake and high rates of nitrogen excretion in early life stages of the cobia Rachycentron canadum (L.), a fast-growing subtropical fish. Journal of Fish Biology, 71(6):1662–1678, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wood:2007:ESS

[2685] A. D. Wood, J. S. Collie, and N. E. Kohler. Estimating survival of the shortfin mako Isurus oxyrinchus (Rafinesque) in the north-west Atlantic from tag-recapture data. Journal of Fish Biology, 71(6):1679–1695, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McVeigh:2007:EED

[2686] B. R. McVeigh, M. C. Healey, and F. Wolfe. Energy expenditures during spawning by chum salmon Oncorhynchus keta (Walbaum) in British Columbia. Journal of Fish Biology, 71(6):1696–1713, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kooka:2007:RED

[2687] K. Kooka, O. Yamamura, and T. Andoh. Rate of energy depletion and overwintering mortality of juvenile walleye pollock in cold water. Journal of Fish Biology, 71(6):1714–1734, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Orell:2007:SDM

[2688] P. Orell, J. Erkinaro, M. A. Svenning, J. G. Davidsen, and E. Niemelä. Synchrony in the downstream migration of smolts and upstream migration of adult Atlantic salmon in the subarctic River Utsjoki. Journal of Fish Biology, 71(6):1735–1750, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yanase:2007:HMA

[2689] K. Yanase and T. Arimoto. A hydro-mechanical approach to the scaling of swimming performance in the sand flathead Platycephalus bassensis Cuvier: effects of changes in morphological features based on fish size. Journal of Fish Biology, 71(6):1751–1772, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2007:GDB

[2690] Y.-J. Lin, L. Ložys, J.-C. Shiao, Y. Iizuka, and W.-N. Tzeng. Growth differences between naturally recruited and stocked European eel Anguilla anguilla from different habitats in Lithuania. Journal of Fish Biology, 71(6):1773–1787, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2007:ECI

[2691] J.-S. Li, J.-L. Li, and T.-T. Wu. The effects of copper, iron and zinc on digestive enzyme activity in the hybrid tilapia Oreochromis niloticus (L.) × Oreochromis aureus (Steindachner). Journal of Fish Biology, 71(6): 1788–1798, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomes:2007:SFV

[2692] J. L. Gomes, Jr. and L. R. Monteiro. Size and fecundity variation in populations of Poecilia vivipara Block & Schneider (Teleostei; Poeciliidae) inhabiting an environmental gradient. Journal of Fish Biology, 71(6): 1799–1809, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolland:2007:LCV

[2693] J. D. Bolland, J. R. Britton, and I. G. Cowx. Lifetime consequences of variable 0 year group length in riverine populations of chub Leuciscus cephalus (L.). Journal of Fish Biology, 71(6):1810–1819, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardie:2007:RBT

[2694] S. A. Hardie, R. W. G. White, and L. A. Barmuta. Reproductive biology of the threatened golden galaxias Galaxias auratus Johnston and the influence of lake hydrology. Journal of Fish Biology, 71(6):1820–1840, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2007:DFO

[2695] C. Smith, Y. Zhu, H. Liu, and M. Reichard. Deceptive female oviposition behaviour elicits male ejaculation in the European bitterling. Journal of Fish Biology, 71(6):1841–1846, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moser:2007:ESI

[2696] M. L. Moser, D. A. Ogden, and B. P. Sandford. Effects of surgically implanted transmitters on anguilliform fishes: lessons from lamprey. Journal of Fish Biology, 71(6):1847–1852, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ragonese:2007:ZCL

[2697] S. Ragonese and G. B. Giusto. Zenopsis conchifera (Lowe, 1852) (Pisces, Actinopterygii, Zeidae): a new alien fish in the Mediterranean Sea. Journal of Fish Biology, 71(6):1853–1857, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salles:2007:MOA

[2698] C. M. C. Salles, J. B. Salles, V. L. F. Cunha Bastos, R. A. Dias, and J. Cunha Bastos. Monoamine oxidase activity in kidney and heart of Piaractus mesopotamicus (Holmberg). Journal of Fish Biology, 71(6): 1858–1863, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magellan:2007:MCS

[2699] K. Magellan and A. E. Magurran. Mate choice, sexual coercion and gene flow in guppy populations. Journal of Fish Biology, 71(6):1864–1872, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawaguchi:2007:SCG

[2700] Y. Kawaguchi, H. Miyasaka, M. Genkai-Kato, Y. Taniguchi, and S. Nakano. Seasonal change in the gastric evacuation rate of rainbow trout feeding on natural prey. Journal of Fish Biology, 71(6):1873–1878, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tokeshi:2007:EDS

[2701] M. Tokeshi and S. Arakaki. On estimating diversity in species-rich fish assemblages: an example with coral-reef fishes of eastern Indonesia. Journal of Fish Biology, 71(6):1879–1887, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:Eg

[2702] Anonymous. Erratum. Journal of Fish Biology, 71(6):1888, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clarke:2007:IV

[2703] A. D. Clarke, A. Lewis, K. H. Telmer, and J. M. Shrimpton. Index, volume 71. Journal of Fish Biology, 71(6):i–xxiv, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:CVb

[2704] Anonymous. Contents of volume 71. Journal of Fish Biology, 71(6): xxv–xxxii, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Laurent:2007:IDS

[2705] V. Laurent, B. Caneco, A. Magoulas, and S. Planes. Isolation by distance and selection effects on genetic structure of sardines Sardina pilchardus Walbaum. Journal of Fish Biology, 71(sa):1–17, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2007:NIM

[2706] H.-Y. Wang, C.-F. Wang, S.-Y. Du, and S.-C. Lee. New insights on molecular systematics of opsariichthines based on cytochrome b sequencing. Journal of Fish Biology, 71(sa):18–32, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2007:HLG

[2707] M. C. F. Santos, M. L. Ruffino, and I. P. Farias. High levels of genetic variability and panmixia of the tambaqui Colossoma macropomum (Cuvier, 1816) in the main channel of the Amazon River. Journal of Fish Biology, 71(sa):33–44, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mesquita:2007:CPE

[2708] N. Mesquita, C. Cunha, G. R. Carvalho, and M. M. Coelho. Comparative phylogeography of endemic cyprinids in the south-west Iberian Peninsula: evidence for a new ichthyogeographic area. Journal of Fish Biology, 71 (sa):45–75, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2007:NSM

[2709] I. J. Harrison, M. Nirchio, C. Oliveira, E. Ron, and J. Gaviria. A new species of mullet (Teleostei: Mugilidae) from Venezuela, with a discussion on the taxonomy of Mugil gaimardianus. Journal of Fish Biology, 71(sa): 76–97, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carlsson:2007:EFS

[2710] J. Carlsson. The effect of family structure on the likelihood for kin-biased distribution: an empirical study of brown trout populations. Journal of Fish Biology, 71(sa):98–110, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrkjedal:2007:LLM

[2711] I. Byrkjedal, D. J. Rees, and E. Willassen. Lumping lumpsuckers: molecular and morphological insights into the taxonomic status of Eumicrotremus spinosus (Fabricius, 1776) and Eumicrotremus eggvinii Koefoed, 1956 (Teleostei: Cyclopteridae). Journal of Fish Biology, 71(sa): 111–131, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2007:EHV

[2712] G. M. Cooke and L. B. Beheregaray. Extremely high variability in the S72 intron of the Amazonian cardinal tetra (Paracheirodon axelrodi). Journal of Fish Biology, 71(sa):132–140, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cervelli:2007:LSV

[2713] M. Cervelli, M. Bianchi, M. Scalici, G. Gibertini, M. Oliverio, and P. Mariottini. Length and sequence variation in the mitochondrial DNA control region of the Etruscan freshwater goby Padogobius nigricans (Teleostei, Gobiidae). Journal of Fish Biology, 71(sa):141–147, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costa:2007:GAF

[2714] P. M. Costa and M. H. Costa. Genotoxicity assessment in fish peripheral blood: a method for a more efficient analysis of micronuclei. Journal of Fish Biology, 71(sa):148–151, August 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Richard:2007:CGT

[2715] M. L. Lennard Richard, E. Bengtén, M. R. Wilson, N. W. Miller, G. W. Warr, and J. Hikima. Comparative genomics of transcription factors driving expression of the immunoglobulin heavy chain locus in teleost fish. Journal of Fish Biology, 71(sb):153–173, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2007:FGT

[2716] C. Fu, D. Li, W. Hu, Y. Wang, and Z. Zhu. Fast-growing transgenic common carp mounting compensatory growth. Journal of Fish Biology, 71 (sb):174–185, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bohlen:2007:LMD

[2717] J. Bohlen, V. Šlechtová, I. Doadrio, and P. Ráb. Low mitochondrial divergence indicates a rapid expansion across Europe in the weather loach, Misgurnus fossilis (L.). Journal of Fish Biology, 71(sb):186–194, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martinez:2007:ASC

[2718] P. Martínez, C. Bouza, J. Castro, M. Hermida, B. G. Pardo, and L. Sánchez. Analysis of a secondary contact between divergent lineages of brown trout Salmo trutta L. from Duero basin using microsatellites and mtDNA RFLPs. Journal of Fish Biology, 71(sb):195–213, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamner:2007:MCA

[2719] R. M. Hamner, D. W. Freshwater, and P. E. Whitfield. Mitochondrial cytochrome b analysis reveals two invasive lionfish species with strong founder effects in the western Atlantic. Journal of Fish Biology, 71(sb): 214–222, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Porta:2007:SLG

[2720] J. Porta, J. Maria Porta, P. Cañavate, G. Martínez-Rodríguez, and M. Carmen Alvarez. Substantial loss of genetic variation in a single generation of Senegalese sole (Solea senegalensis) culture: implications in the domestication process. Journal of Fish Biology, 71(sb):223–234, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anderson:2007:GEA

[2721] J. D. Anderson and W. J. Karel. Genetic evidence for asymmetric hybridization between menhadens (Brevoortia spp.) from peninsular Florida. Journal of Fish Biology, 71(sb):235–249, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoshikawa:2007:SFH

[2722] H. Yoshikawa, K. Morishima, T. Fujimoto, E. Yamaha, and K. Arai. Simultaneous formation of haploid, diploid and triploid eggs in diploid– triploid mosaic loaches. Journal of Fish Biology, 71(sb):250–263, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bjorklund:2007:GDP

[2723] M. Björklund, T. Aho, and L. C. Larsson. Genetic differentiation in pikeperch (Sander lucioperca): the relative importance of gene flow, drift and common history. Journal of Fish Biology, 71(sb):264–278, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gagnaire:2007:RSM

[2724] P. A. Gagnaire, K. Tsukamoto, J. Aoyama, Y. Minegishi, P. Valade, and P. Berrebi. RFLP and semi-multiplex PCR-based identification of four eel species from the south-western Indian Ocean region. Journal of Fish Biology, 71(sb):279–287, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Close:2007:CEG

[2725] P. G. Close and G. Gouws. Cryptic estuarine gobiid fish larvae (Pisces: Gobiidae) identified as Pseudogobius olorum using allozyme markers. Journal of Fish Biology, 71(sb):288–293, October 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moghadam:2007:LVS

[2726] H. K. Moghadam, M. M. Ferguson, and R. G. Danzmann. Linkage variation at the sex-determining locus within Fraser strain Arctic charr Salvelinus alpinus. Journal of Fish Biology, 71(sb):294–301, October 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heath:2007:EDE

[2727] D. Heath, D. C. Lou, J. Ovenden, and G. A. Begg. Effect of DNA extraction on ageing success in coral trout (Plectropomus leopardus) otoliths. Journal of Fish Biology, 71(sb):302–307, October 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kajimura:2007:ILG

[2728] S. Kajimura and C. Duan. Insulin-like growth factor-binding protein-1: an evolutionarily conserved fine tuner of insulin-like growth factor action under catabolic and stressful conditions. Journal of Fish Biology, 71(sc): 309–325, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takle:2007:CAF

[2729] H. Takle and Ø. Andersen. Caspases and apoptosis in fish. Journal of Fish Biology, 71(sc):326–349, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kellison:2007:DIH

[2730] G. T. Kellison and J. C. Taylor. Demonstration and implications of habitat-specific chemical signatures in otoliths of juvenile summer flounder (Paralichthys dentatus Linnaeus) in North Carolina. Journal of Fish Biology, 71(sc):350–359, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ohayon:2007:GBR

[2731] J. L. Ohayon and C. A. Stepien. Genetic and biogeographic relationships of the racer goby Neogobius gymnotrachelus (Gobiidae: Teleostei) from introduced and native Eurasian locations. Journal of Fish Biology, 71 (sc):360–370, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weiss:2007:SCB

[2732] S. Weiss, I. Knizhin, V. Romanov, and T. Kopun. Secondary contact between two divergent lineages of grayling Thymallus in the lower Enisey basin and its taxonomic implications. Journal of Fish Biology, 71(sc): 371–386, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Janko:2007:DES

[2733] K. Janko, M. Flajšhans, L. Choleva, J. Bohlen, V. Šlechtová, M. Rábová, Z. Lajbner, V. Šlechta, P. Ivanova, I. Dobrovolov, M. Culling, H. Persat, J. Kotusz, and P. Ráb. Diversity of European spined loaches (genus Cobitis L.): an update of the geographic distribution of the Cobitis taenia hybrid complex with a description of new molecular tools for species and hybrid determination. Journal of Fish Biology, 71(sc):387–408, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rengmark:2007:IMG

[2734] A. H. Rengmark, A. Slettan, W. J. Lee, Ø. Lie, and F. Lingaas. Identification and mapping of genes associated with salt tolerance in tilapia. Journal of Fish Biology, 71(sc):409–422, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crespo-Lopez:2007:GSD

[2735] M. E. Crespo-López, I. Pala, T. L. Duarte, T. E. Dowling, and M. M. Coelho. Genetic structure of the diploid–polyploid fish Squalius alburnoides in southern Iberian basins Tejo and Guadiana, based on microsatellites. Journal of Fish Biology, 71(sc):423–436, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitanishi:2007:GSM

[2736] S. Kitanishi, K. Edo, T. Yamamoto, N. Azuma, O. Hasegawa, and S. Higashi. Genetic structure of masu salmon (Oncorhynchus masou) populations in Hokkaido, northernmost Japan, inferred from mitochondrial DNA variation. Journal of Fish Biology, 71(sc):437–452, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sczepanski:2007:CMC

[2737] T. S. Sczepanski, R. B. Noleto, D. L. Z. Kantek, M. C. S. Cortinhas, and M. M. Cestari. Classical and molecular cytogenetics of Atherinella brasiliensis (Teleostei, Atheriniformes) from South coast of Brazil. Journal of Fish Biology, 71(sc):453–460, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Volk:2007:WPD

[2738] J. Volk, D. Bekkevold, and V. Loeschcke. Weak population differentiation in northern European populations of the endangered anadromous clupeid Alosa fallax. Journal of Fish Biology, 71(sc):461–469, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aubry:2007:AYY

[2739] E. Aubry, C. Cagnon, Y. Lalanne, and C. Mouchès. Assessment of young yellow European eel Anguilla anguilla L. exposure to a CYP1A1 inducer by the quantification of increase in hepatic CYP1A1 mRNA using realtime RT-PCR. Journal of Fish Biology, 71(sc):470–477, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fopp-Bayat:2007:VMG

[2740] D. Fopp-Bayat. Verification of meiotic gynogenesis in Siberian sturgeon (Acipenser baeri Brandt) using microsatellite DNA and cytogenetical markers. Journal of Fish Biology, 71(sc):478–485, December 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:IVc

[2741] Anonymous. Index, volume 71. Journal of Fish Biology, 71(sc):i–vii, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:CVSb

[2742] Anonymous. Contents of volume 71 supplements. Journal of Fish Biology, 71(sc):viii–x, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fausch:2007:IEE

[2743] K. D. Fausch. Introduction, establishment and effects of non-native salmonids: considering the risk of rainbow trout invasion in the United Kingdom. Journal of Fish Biology, 71(sd):1–32, December 2007. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia-Berthou:2007:CIF

[2744] E. García-Berthou. The characteristics of invasive fishes: what has been learned so far? Journal of Fish Biology, 71(sd):33–55, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fox:2007:LHT

[2745] M. G. Fox, A. Vila-Gispert, and G. H. Copp. Life-history traits of introduced Iberian pumpkinseed Lepomis gibbosus relative to native populations. Can differences explain colonization success? Journal of Fish Biology, 71(sd):56–69, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Scott:2007:CSL

[2746] D. M. Scott, R. W. Wilson, and J. A. Brown. Can sunbleak Leucaspius delineatus or topmouth gudgeon Pseudorasbora parva disperse through saline waters? Journal of Fish Biology, 71(sd):70–86, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schofield:2007:HTT

[2747] P. J. Schofield, W. F. Loftus, and M. E. Brown. Hypoxia tolerance of two centrarchid sunfishes and an introduced cichlid from karstic Everglades wetlands of southern Florida, U.S.A. Journal of Fish Biology, 71(sd): 87–99, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sorensen:2007:CRD

[2748] P. W. Sorensen and T. R. Hoye. A critical review of the discovery and application of a migratory pheromone in an invasive fish, the sea lamprey Petromyzon marinus L. Journal of Fish Biology, 71(sd):100–114, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanfling:2007:UES

[2749] B. Hänfling. Understanding the establishment success of non-indigenous fishes: lessons from population genetics. Journal of Fish Biology, 71(sd): 115–135, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valiente:2007:GVI

[2750] A. G. Valiente, F. Juanes, P. Nuñez, and E. Garcia-Vazquez. Is genetic variability so important? Non-native salmonids in South America. Journal of Fish Biology, 71(sd):136–147, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Copp:2007:PPI

[2751] G. H. Copp, M. Templeton, and R. E. Gozlan. Propagule pressure and the invasion risks of non-native freshwater fishes: a case study in England. Journal of Fish Biology, 71(sd):148–159, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Verreycken:2007:NIF

[2752] H. Verreycken, D. Anseeuw, G. Van Thuyne, P. Quataert, and C. Belpaire. The non-indigenous freshwater fishes of Flanders (Belgium): review, status and trends over the last decade. Journal of Fish Biology, 71 (sd):160–172, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hesthagen:2007:NNF

[2753] T. Hesthagen and O. T. Sandlund. Non-native freshwater fishes in Norway: history, consequences and perspectives. Journal of Fish Biology, 71 (sd):173–183, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Museth:2007:HMP

[2754] J. Museth, T. Hesthagen, O. T. Sandlund, E. B. Thorstad, and O. Ugedal. The history of the minnow Phoxinus phoxinus (L.) in Norway: from harmless species to pest. Journal of Fish Biology, 71(sd): 184–195, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gjelland:2007:CMM

[2755] K. Ø. Gjelland, T. Bøhn, and P.-A. Amundsen. Is coexistence mediated by microhabitat segregation? An in-depth exploration of a fish invasion. Journal of Fish Biology, 71(sd):196–209, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cucherousset:2007:SIE

[2756] J. Cucherousset, J. C. Aymes, F. Santoul, and R. Céréghino. Stable isotope evidence of trophic interactions between introduced brook trout Salvelinus fontinalis and native brown trout Salmo trutta in a mountain stream of south-west France. Journal of Fish Biology, 71(sd):210–223, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beyer:2007:MUI

[2757] K. Beyer, G. H. Copp, and R. E. Gozlan. Microhabitat use and interspecific associations of introduced topmouth gudgeon Pseudorasbora parva and native fishes in a small stream. Journal of Fish Biology, 71(sd): 224–238, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Britton:2007:IFD

[2758] J. R. Britton, R. R. Boar, J. Grey, J. Foster, J. Lugonzo, and D. M. Harper. From introduction to fishery dominance: the initial impacts of the invasive carp Cyprinus carpio in Lake Naivasha, Kenya, 1999 to 2006. Journal of Fish Biology, 71(sd):239–257, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Irons:2007:RCF

[2759] K. S. Irons, G. G. Sass, M. A. McClelland, and J. D. Stafford. Reduced condition factor of two native fish species coincident with invasion of non-native Asian carps in the Illinois River, U.S.A.. Is this evidence for competition and reduced fitness? Journal of Fish Biology, 71(sd):258– 273, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carpentier:2007:TGP

[2760] A. Carpentier, R. E. Gozlan, J. Cucherousset, J.-M. Paillisson, and L. Marion. Is topmouth gudgeon Pseudorasbora parva responsible for the decline in sunbleak Leucaspius delineatus populations? Journal of Fish Biology, 71(sd):274–278, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:RS

[2761] Anonymous. Referees for this supplement. Journal of Fish Biology, 71 (sd):279, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2007:IS

[2762] Anonymous. Index, 2007 supplement. Journal of Fish Biology, 71(sd): 280–283, December 2007. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nam:2008:AAM

[2763] Y. K. Nam, N. Maclean, G. Hwang, and D. S. Kim. Autotransgenic and allotransgenic manipulation of growth traits in fish for aquaculture: a review. Journal of Fish Biology, 72(1):1–26, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keefer:2008:NDH

[2764] M. L. Keefer, C. C. Caudill, C. A. Peery, and C. T. Boggs. Non-direct homing behaviours by adult Chinook salmon in a large, multi-stock river system. Journal of Fish Biology, 72(1):27–44, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Narum:2008:ICM

[2765] S. R. Narum, D. Hatch, A. J. Talbot, P. Moran, and M. S. Powell. Iteroparity in complex mating systems of steelhead Oncorhynchus mykiss (Walbaum). Journal of Fish Biology, 72(1):45–60, January 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kyne:2008:ARD

[2766] P. M. Kyne, A. J. Courtney, and M. B. Bennett. Aspects of reproduction and diet of the Australian endemic skate Dipturus polyommata (Ogilby) (Elasmobranchii: Rajidae), by-catch of a commercial prawn trawl fishery. Journal of Fish Biology, 72(1):61–77, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kennedy:2008:FPP

[2767] J. Kennedy, P. R. Witthames, R. D. M. Nash, and C. J. Fox. Is fecundity in plaice (Pleuronectes platessa L.) down-regulated in response to reduced food intake during autumn? Journal of Fish Biology, 72(1):78– 92, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takegaki:2008:SAN

[2768] T. Takegaki, Y. Matsumoto, A. Tawa, T. Miyano, and Y. Natsukari. Sizeassortative nest preference in a paternal brooding blenny Rhabdoblennius ellipes (Jordan & Starks). Journal of Fish Biology, 72(1):93–102, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marcalo:2008:SES

[2769] A. Marçalo, P. Pousão-Ferreira, L. Mateus, J. H. Duarte Correia, and Y. Stratoudakis. Sardine early survival, physical condition and stress after introduction to captivity. Journal of Fish Biology, 72(1):103–120, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fey:2008:FAO

[2770] D. P. Fey and J. A. Hare. Fluctuating asymmetry in the otoliths of larval Atlantic menhaden Brevoortia tyrannus (Latrobe) — a condition indicator? Journal of Fish Biology, 72(1):121–130, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bergstad:2008:SFA

[2771] O. A. Bergstad, T. Jørgensen, J. A. Knutsen, and J. A. Berge. Site fidelity of Atlantic cod Gadus morhua L. as deduced from telemetry and stable isotope studies. Journal of Fish Biology, 72(1):131–142, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2008:IHA

[2772] P. G. Jellyman and A. R. McIntosh. The influence of habitat availability and adult density on non-diadromous galaxiid fry settlement in New Zealand. Journal of Fish Biology, 72(1):143–156, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Afonso:2008:LGG

[2773] L. O. B. Afonso, S. Hosoya, J. Osborne, A. K. Gamperl, and S. Johnson. Lack of glucose and hsp70 responses in haddock Melanogrammus aeglefinus (L.) subjected to handling and heat shock. Journal of Fish Biology, 72(1):157–167, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Samson:2008:ETD

[2774] J. C. Samson, S. E. Shumway, and J. S. Weis. Effects of the toxic dinoflagellate, Alexandrium fundyense on three species of larval fish: a food-chain approach. Journal of Fish Biology, 72(1):168–188, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luque:2008:LEP

[2775] J. L. Luque and R. Poulin. Linking ecology with parasite diversity in Neotropical fishes. Journal of Fish Biology, 72(1):189–204, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moreau:2008:OPS

[2776] S. Moreau, C. Péron, K. A. Pitt, R. M. Connolly, S. Y. Lee, and T. Meziane. Opportunistic predation by small fishes on epibiota of jetty pilings in urban waterways. Journal of Fish Biology, 72(1):205–217, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baumgartner:2008:DDV

[2777] L. J. Baumgartner, I. G. Stuart, and B. P. Zampatti. Determining diel variation in fish assemblages downstream of three weirs in a regulated lowland river. Journal of Fish Biology, 72(1):218–232, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loffler:2008:EDS

[2778] J. Löffler, A. Ott, H. Ahnelt, and H. Keckeis. Early development of the skull of Sander lucioperca (L.) (Teleostei: Percidae) relating to growth and mortality. Journal of Fish Biology, 72(1):233–258, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2008:NWB

[2779] P. J. Miller and R. Šanda. A new West Balkanian sand-goby (Teleostei: Gobiidae). Journal of Fish Biology, 72(1):259–270, January 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Collins:2008:DGD

[2780] M. A. Collins, R. S. Shreeve, S. Fielding, and M. H. Thurston. Distribution, growth, diet and foraging behaviour of the yellow-fin notothen Patagonotothen guntheri (Norman) on the Shag Rocks shelf (Southern Ocean). Journal of Fish Biology, 72(1):271–286, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Craig:2008:SGS

[2781] G. Craig, D. Paynter, I. Coscia, and S. Mariani. Settlement of gilthead sea bream Sparus aurata L. in a southern Irish Sea coastal habitat. Journal of Fish Biology, 72(1):287–291, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fernandez:2008:TCN

[2782] A. A. Fernandez and P. R. Bowser. Two cases of non-hybrid melanoma formation in Xiphophorus nezahualcoyotl Rauchenberger, Kallmann & Morizot. Journal of Fish Biology, 72(1):292–300, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guelinckx:2008:EDF

[2783] J. Guelinckx, F. Dehairs, and F. Ollevier. Effect of digestion on the δ 13C and δ ^{15}N of fish-gut contents. Journal of Fish Biology, 72(1):301– 309, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liousia:2008:VCS

[2784] V. Liousia, R. Liasko, E. Koutrakis, and I. D. Leonardos. Variation in clones of the sperm-dependent parthenogenetic Carassius gibelio (Bloch) in Lake Pamvotis (north-west Greece). Journal of Fish Biology, 72(1): 310–314, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:BRa

[2785] Anonymous. Book reviews. Journal of Fish Biology, 72(1):315–321, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:BRb

[2786] Anonymous. Books received. Journal of Fish Biology, 72(1):322, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:R

[2787] Anonymous. Referees 2007. Journal of Fish Biology, 72(1):323–328, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:ISF

[2788] Anonymous. 7th International Symposium on Flatfish Ecology: The impact of environmental changes on flatfish productivity. Journal of Fish Biology, 72(1):329, January 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albert:2008:CSG

[2789] J. S. Albert, H. H. Zakon, P. K. Stoddard, G. A. Unguez, S. K. S. Holmberg-Albert, and M. R. Sussman. The case for sequencing the genome of the electric eel Electrophorus electricus. Journal of Fish Biology, 72(2):331–354, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2008:FMG

[2790] K. H. Brown. Fish mitochondrial genomics: sequence, inheritance and functional variation. Journal of Fish Biology, 72(2):355–374, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McCusker:2008:MMD

[2791] M. R. McCusker, I. G. Paterson, and P. Bentzen. Microsatellite markers discriminate three species of North Atlantic wolffishes (Anarhichas spp.). Journal of Fish Biology, 72(2):375–385, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shikano:2008:CGD

[2792] T. Shikano, Y. Shimada, and H. Suzuki. Comparison of genetic diversity at microsatellite loci and quantitative traits in hatchery populations of Japanese flounder Paralichthys olivaceus. Journal of Fish Biology, 72 (2):386–399, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goodman:2008:MDE

[2793] D. H. Goodman, S. B. Reid, M. F. Docker, G. R. Haas, and A. P. Kinziger. Mitochondrial DNA evidence for high levels of gene flow among populations of a widely distributed anadromous lamprey Entosphenus tridentatus (Petromyzontidae). Journal of Fish Biology, 72(2):400–417, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ray:2008:HRR

[2794] J. M. Ray, N. J. Lang, R. M. Wood, and R. L. Mayden. History repeated: recent and historical mitochondrial introgression between the current darter Etheostoma uniporum and rainbow darter Etheostoma caeruleum (Teleostei: Percidae). Journal of Fish Biology, 72(2):418–434, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schwartz:2008:UGS

[2795] T. S. Schwartz and L. B. Beheregaray. Using genotype simulations and Bayesian analyses to identify individuals of hybrid origin in Australian bass: lessons for fisheries management. Journal of Fish Biology, 72(2): 435–450, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Froukh:2008:SBE

[2796] T. Froukh and M. Kochzius. Species boundaries and evolutionary lineages in the blue green damselfishes Chromis viridis and Chromis atripectoralis (Pomacentridae). Journal of Fish Biology, 72(2):451–457, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:2008:DBR

[2797] R. D. Ward, B. H. Holmes, and G. K. Yearsley. DNA barcoding reveals a likely second species of Asian sea bass (barramundi) (Lates calcarifer). Journal of Fish Biology, 72(2):458–463, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2008:DBI

[2798] P. J. Smith, S. M. McVeagh, and D. Steinke. DNA barcoding for the identification of smoked fish products. Journal of Fish Biology, 72(2): 464–471, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:Ea

[2799] Anonymous. Erratum. Journal of Fish Biology, 72(2):472, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Andrade:2008:RBP

[2800] A. C. Andrade, L. C. Silva-Junior, and M. Vianna. Reproductive biology and population variables of the Brazilian sharpnose shark Rhizoprionodon lalandii (Müller & Henle, 1839) captured in coastal waters of south-eastern Brazil. Journal of Fish Biology, 72(3):473–484, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paakkonen:2008:PEJ

[2801] J.-P. Pääkkönen, S. Rönkkönen, M. Karjalainen, and M. Viitasalo. Physiological effects in juvenile three-spined sticklebacks feeding on toxic cyanobacterium Nodularia spumigena-exposed zooplankton. Journal of Fish Biology, 72(3):485–499, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rowe:2008:LBS

[2802] S. Rowe and J. A. Hutchings. A link between sound producing musculature and mating success in Atlantic cod. Journal of Fish Biology, 72 (3):500–511, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alvarez:2008:TDO

[2803] A. Álvarez, B. Morales-Nin, M. Palmer, J. Tomás, and J. Sastre. A two-dimension otolith growth inverse model. Journal of Fish Biology, 72 (3):512–522, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tobler:2008:TEE

[2804] M. Tobler, R. Riesch, F. J. Garcíade León, I. Schlupp, and M. Plath. Two endemic and endangered fishes, Poecilia sulphuraria (Alvarez, 1948) and Gambusia eurystoma Miller, 1975 (Poeciliidae, Teleostei) as only survivors in a small sulphidic habitat. Journal of Fish Biology, 72(3): 523–533, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yokoi:2008:SMC

[2805] K. Yokoi, H. Ohta, and K. Hosoya. Sperm motility and cryopreservation of spermatozoa in freshwater gobies. Journal of Fish Biology, 72(3): 534–544, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliva:2008:WCM

[2806] M. E. Oliva, I. M. Valdivia, G. Costa, N. Freitas, M. A. Pinheiro De Carvalho, L. Sánchez, and J. L. Luque. What can metazoan parasites reveal about the taxonomy of Scomber japonicus Houttuyn in the coast of South America and Madeira Islands? Journal of Fish Biology, 72 (3):545–554, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bertrand:2008:TPB

[2807] M. Bertrand, D. J. Marcogliese, and P. Magnan. Trophic polymorphism in brook charr revealed by diet, parasites and morphometrics. Journal of Fish Biology, 72(3):555–572, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powter:2008:EMP

[2808] D. M. Powter and W. Gladstone. Embryonic mortality and predation on egg capsules of the Port Jackson shark Heterodontus portusjacksoni (Meyer). Journal of Fish Biology, 72(3):573–584, February 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2008:MSE

[2809] S. X. Chen, W. S. Hong, Y. Q. Su, and Q. Y. Zhang. Microhabitat selection in the early juvenile mudskipper Boleophthalmus pectinirostris (L.). Journal of Fish Biology, 72(3):585–593, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jepsen:2008:ETS

[2810] N. Jepsen, J. S. Mikkelsen, and A. Koed. Effects of tag and suture type on survival and growth of brown trout with surgically implanted telemetry tags in the wild. Journal of Fish Biology, 72(3):594–602, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Takahashi:2008:DNC

[2811] T. Takahashi. Description of a new cichlid fish species of the genus Benthochromis (Perciformes: Cichlidae) from Lake Tanganyika. Journal of Fish Biology, 72(3):603–613, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chittenden:2008:REM

[2812] C. M. Chittenden, S. Sura, K. G. Butterworth, K. F. Cubitt, N. Plantalech Manel-la, S. Balfry, F. Økland, and R. S. McKinley. Riverine, estuarine and marine migratory behaviour and physiology of wild and hatchery-reared Coho salmon Oncorhynchus kisutch (Walbaum) smolts descending the Campbell River, BC, Canada. Journal of Fish Biology, 72 (3):614–628, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Correa:2008:CFF

[2813] S. B. Correa, W. G. R. Crampton, L. J. Chapman, and J. S. Albert. A comparison of flooded forest and floating meadow fish assemblages in an upper Amazon floodplain. Journal of Fish Biology, 72(3):629–644, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peake:2008:GTS

[2814] S. J. Peake. Gait transition speed as an alternate measure of maximum aerobic capacity in fishes. Journal of Fish Biology, 72(3):645–655, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kajimura:2008:SDS

[2815] M. Kajimura, P. J. Walsh, and C. M. Wood. The spiny dogfish Squalus acanthias L. maintains osmolyte balance during long-term starvation. Journal of Fish Biology, 72(3):656–670, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bills:2008:OCV

[2816] M. L. Bills, S. B. Callaway, A. T. Kessler, and M. H. Taylor. Ovarian contraction in various teleost families. Journal of Fish Biology, 72(3): 671–680, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maruyama:2008:VBR

[2817] A. Maruyama, Y. Onoda, and M. Yuma. Variation in behavioural response to oxygen stress by egg-tending males of parapatric fluvial and lacustrine populations of a landlocked goby. Journal of Fish Biology, 72 (3):681–692, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Farrell:2008:CSP

[2818] A. P. Farrell. Comparisons of swimming performance in rainbow trout using constant acceleration and critical swimming speed tests. Journal of Fish Biology, 72(3):693–710, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simpfendorfer:2008:GRJ

[2819] C. A. Simpfendorfer, G. R. Poulakis, P. M. O’Donnell, and T. R. Wiley. Growth rates of juvenile smalltooth sawfish Pristis pectinata Latham in the western Atlantic. Journal of Fish Biology, 72(3):711–723, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vallisneri:2008:SSS

[2820] M. Vallisneri, V. Trotta, S. Cavicchi, and C. Piccinetti. Sex-specific somatic–otolith growth relationship in two Gadidae. Journal of Fish Biology, 72(3):724–730, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Florin:2008:SFP

[2821] A.-B. Florin and A. Lingman. Shrinkage of flounder Platichthys flesus (L.) and turbot Psetta maxima (L.) following freezing. Journal of Fish Biology, 72(3):731–736, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koumoundouros:2008:FRS

[2822] G. Koumoundouros. First record of saddleback syndrome in wild parrotfish Sparisoma cretense (L., 1758) (Perciformes, Scaridae). Journal of Fish Biology, 72(3):737–741, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jatteau:2008:PAS

[2823] P. Jatteau and A. Bardonnet. Photoresponse in allis shad larvae. Journal of Fish Biology, 72(3):742–746, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simontacchi:2008:ASI

[2824] C. Simontacchi, C. Poltronieri, C. Carraro, D. Bertotto, G. Xiccato, A. Trocino, and G. Radaelli. Alternative stress indicators in sea bass Dicentrarchus labrax, L. Journal of Fish Biology, 72(3):747–752, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Acou:2008:ESE

[2825] A. Acou, T. Robinet, E. Lance, C. Gerard, B. Mounaix, L. Brient, B. Le Rouzic, and E. Feunteun. Evidence of silver eels contamination by microcystin-LR at the onset of their seaward migration: what consequences for breeding potential? Journal of Fish Biology, 72(3):753–762, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Killen:2008:LMT

[2826] S. S. Killen, J. A. Brown, and A. K. Gamperl. Lack of metabolic thermal compensation during the early life stages of ocean pout Zoarces americanus (Bloch & Schneider): a benthic, cold-water marine species. Journal of Fish Biology, 72(3):763–772, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giarrizzo:2008:HIF

[2827] T. Giarrizzo and U. Krumme. Heterogeneity in intertidal fish fauna assemblages along the world’s longest mangrove area in northern Brazil. Journal of Fish Biology, 72(3):773–779, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haukenes:2008:CRP

[2828] A. H. Haukenes, B. A. Barton, and H. Bollig. Cortisol responses of pallid sturgeon and yellow perch following challenge with lipopolysaccharide. Journal of Fish Biology, 72(3):780–784, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:Eb

[2829] Anonymous. Errata. Journal of Fish Biology, 72(3):785–786, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hardie:2008:E

[2830] S. A. Hardie, R. W. G. White, and L. A. Barmuta. Errata. Journal of Fish Biology, 72(3):786, February 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wistbacka:2008:TAB

[2831] S. Wistbacka and G. Bylund. Thiaminase activity of Baltic salmon prey species: a comparision of net- and predator-caught samples. Journal of Fish Biology, 72(4):787–802, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geerinckx:2008:OAS

[2832] T. Geerinckx, Y. Verhaegen, and D. Adriaens. Ontogenetic allometries and shape changes in the suckermouth armoured catfish Ancistrus cf. triradiatus Eigenmann (Loricariidae, Siluriformes), related to suckermouth attachment and yolk-sac size. Journal of Fish Biology, 72(4):803–814, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mintenbeck:2008:HAL

[2833] K. Mintenbeck, T. Brey, U. Jacob, R. Knust, and U. Struck. How to account for the lipid effect on carbon stable-isotope ratio (δ ^{13}C): sample treatment effects and model bias. Journal of Fish Biology, 72(4):815– 830, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kraus:2008:SPI

[2834] G. Kraus, J. Tomkiewicz, R. Diekmann, and F. W. Köster. Seasonal prevalence and intensity of follicular atresia in Baltic cod Gadus morhua callarias L. Journal of Fish Biology, 72(4):831–847, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stradmeyer:2008:CBB

[2835] L. Stradmeyer, J. Höjesjö, S. W. Griffiths, D. J. Gilvear, and J. D. Armstrong. Competition between brown trout and Atlantic salmon parr over pool refuges during rapid dewatering. Journal of Fish Biology, 72(4):848– 860, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Filer:2008:AGR

[2836] K. R. Filer and G. R. Sedberry. Age, growth and reproduction of the barrelfish Hyperoglyphe perciformis (Mitchill) in the western North Atlantic. Journal of Fish Biology, 72(4):861–882, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Marshall:2008:CDT

[2837] A. D. Marshall, P. M. Kyne, and M. B. Bennett. Comparing the diet of two sympatric urolophid elasmobranchs (Trygonoptera testacea Müller & Henle and Urolophus kapalensis Yearsley & Last): evidence of ontogenetic shifts and possible resource partitioning. Journal of Fish Biology, 72(4):883–898, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez-Casanova:2008:SMR

[2838] J. C. Pérez-Casanova, L. O. B. Afonso, S. C. Johnson, S. Currie, and A. K. Gamperl. The stress and metabolic responses of juvenile Atlantic cod Gadus morhua L. to an acute thermal challenge. Journal of Fish Biology, 72(4):899–916, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Taylor:2008:CDS

[2839] S. M. Taylor and M. B. Bennett. Cephalopod dietary specialization and ontogenetic partitioning of the Australian weasel shark Hemigaleus australiensis White, Last & Compagno. Journal of Fish Biology, 72(4): 917–936, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2008:VPD

[2840] E. I. Hasegawa, K. Sawada, K. Abe, K. Watanabe, K. Uchikawa, Y. Okazaki, M. Toyama, and R. H. Douglas. The visual pigments of a deep-sea myctophid fish Myctophum nitidulum Garman; an HPLC and spectroscopic description of a non-paired rhodopsin–porphyropsin system. Journal of Fish Biology, 72(4):937–945, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lochet:2008:RAI

[2841] A. Lochet, P. Jatteau, J. Tomás, and E. Rochard. Retrospective approach to investigating the early life history of a diadromous fish: allis shad Alosa alosa (L.) in the Gironde–Garonne–Dordogne watershed. Journal of Fish Biology, 72(4):946–960, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodda:2008:FME

[2842] K. R. Rodda and R. S. Seymour. Functional morphology of embryonic development in the Port Jackson shark Heterodontus portusjacksoni (Meyer). Journal of Fish Biology, 72(4):961–984, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Syrjanen:2008:SGB

[2843] J. Syrjänen, M. Kiljunen, J. Karjalainen, A. Eloranta, and T. Muotka. Survival and growth of brown trout Salmo trutta L. embryos and the timing of hatching and emergence in two boreal lake outlet streams. Journal of Fish Biology, 72(4):985–1000, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stoll:2008:EWD

[2844] S. Stoll, P. Fischer, P. Klahold, N. Scheifhacken, H. Hofmann, and K.-O. Rothhaupt. Effects of water depth and hydrodynamics on the growth and distribution of juvenile cyprinids in the littoral zone of a large pre-alpine lake. Journal of Fish Biology, 72(4):1001–1022, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ibanez:2008:RBS

[2845] A. L. Ibáñez, J. R. Britton, and I. G. Cowx. Relationship between scale growth checks, circuli formation rate and somatic growth in Rutilus rutilus (L.) a fish farm-reared cyprinid. Journal of Fish Biology, 72(4): 1023–1034, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guedes:2008:TRP

[2846] A. P. P. Guedes and F. G. Araújo. Trophic resource partitioning among five flatfish species (Actinopterygii, Pleuronectiformes) in a tropical bay in south-eastern Brazil. Journal of Fish Biology, 72(4):1035–1054, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2008:SCT

[2847] W. T. Liu, Y. Zhang, G. Y. Li, Y. Q. Miao, and X. H. Wu. Structure and composition of teleost scales from snakehead Channa argus (Cantor) (Perciformes: Channidae). Journal of Fish Biology, 72(4):1055– 1067, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanna:2008:TEM

[2848] S. K. Hanna, A. H. Haukenes, R. J. Foy, and C. L. Buck. Temperature effects on metabolic rate, swimming performance and condition of Pacific cod Gadus macrocephalus Tilesius. Journal of Fish Biology, 72(4):1068– 1078, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hamilton:2008:SDE

[2849] I. M. Hamilton and D. Heg. Sex differences in the effect of social status on the growth of subordinates in a co-operatively breeding cichlid. Journal of Fish Biology, 72(4):1079–1088, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gante:2008:FRB

[2850] H. F. Gante, L. Moreira da Costa, J. Micael, and M. J. Alves. First record of Barbonymus schwanenfeldii (Bleeker) in the Iberian Peninsula. Journal of Fish Biology, 72(4):1089–1094, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kneer:2008:RAW

[2851] D. Kneer, H. Asmus, H. Ahnelt, and J. A. Vonk. Records of Austrolethops wardi Whitley (Teleostei: Gobiidae) as an inhabitant of burrows of the thalassinid shrimp Neaxius acanthus in tropical seagrass beds of the Spermonde Archipelago, Sulawesi, Indonesia. Journal of Fish Biology, 72(4):1095–1099, March 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rocha:2008:SCR

[2852] L. A. Rocha and B. W. Bowen. Speciation in coral-reef fishes. Journal of Fish Biology, 72(5):1101–1121, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2008:PSE

[2853] L. Zhang, Q. Y. Tang, and H. Z. Liu. Phylogeny and speciation of the eastern Asian cyprinid genus Sarcocheilichthys. Journal of Fish Biology, 72(5):1122–1137, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tseng:2008:CGV

[2854] M. C. Tseng, T.-Y. Chiang, and J.-P. Wang. Characterization and genetic variations of satellite DNAs in Acrossocheilus paradoxus (Günther, 1868) (Cyprinidae) indicate population expansion. Journal of Fish Biology, 72(5):1138–1153, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pappalardo:2008:GSK

[2855] A. M. Pappalardo, V. Ferrito, A. Messina, F. Guarino, T. Patarnello, V. De Pinto, and C. Tigano. Genetic structure of the killifish Aphanius fasciatus, Nardo 1827 (Teleostei, Cyprinodontidae), results of mitochondrial DNA analysis. Journal of Fish Biology, 72(5):1154–1173, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thuesen:2008:GDS

[2856] P. A. Thuesen, B. J. Pusey, D. R. Peck, R. Pearson, and B. C. Congdon. Genetic differentiation over small spatial scales in the absence of physical barriers in an Australian rainforest stream fish. Journal of Fish Biology, 72(5):1174–1187, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bowen:2008:PRA

[2857] B. R. Bowen, B. R. Kreiser, P. F. Mickle, J. F. Schaefer, and S. B. Adams. Phylogenetic relationships among North American Alosa species (Clupeidae). Journal of Fish Biology, 72(5):1188–1201, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2008:GSB

[2858] S. Kim, C.-G. Kim, J. Oh, B.-J. Kim, H.-S. Seo, W.-S. Kim, and Y.- H. Lee. Genetic similarity between the South Atlantic and the western North Pacific Maurolicus (Stomiiformes: Actinopterygii) taxa, M. walvisensis Parin & Kobyliansky and M. japonicus Ishikawa: evidence for synonymy? Journal of Fish Biology, 72(5):1202–1214, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2008:CSD

[2859] Y. Zhang, L. Wang, M. Shao, J. Li, B. Li, and H. Zhang. A comparative study on the developmental expression of hadh2 in amphioxus and zebrafish. Journal of Fish Biology, 72(5):1215–1222, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [3217].

Saura:2008:TAS

[2860] M. Saura, P. Caballero, and P. Morán. Are there Atlantic salmon in the River Tambre? Journal of Fish Biology, 72(5):1223–1229, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papousek:2008:INH

[2861] I. Papoušek, L. Vetešník, K. Halačka, V. Lusková, M. Humpl, and J. Mendel. Identification of natural hybrids of gibel carp Carassius auratus gibelio (Bloch) and crucian carp Carassius carassius (L.) from lower Dyje River floodplain (Czech Republic). Journal of Fish Biology, 72 (5):1230–1235, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santi-Rampazzo:2008:IPN

[2862] A. P. Santi-Rampazzo, P. B. Nishiyama, P. E. B. Ferreira, and I. C. Martins-Santos. Intrapopulational polymorphism of nucleolus organizer regions in Serrapinnus notomelas (Characidae, Cheirodontinae) from the Paraná River. Journal of Fish Biology, 72(5):1236–1243, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:Ec

[2863] Anonymous. Erratum. Journal of Fish Biology, 72(5):1244, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2008:BJW

[2864] W.-K. Chen, K.-M. Liu, and Y.-Y. Liao. Bioenergetics of juvenile whitespotted bamboo shark Chiloscyllium plagiosum [anonymous (Bennett)]. Journal of Fish Biology, 72(6):1245–1258, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitsos:2008:DCS

[2865] M.-S. Kitsos, Th. Tzomos, L. Anagnostopoulou, and A. Koukouras. Diet composition of the seahorses, Hippocampus guttulatus Cuvier, 1829 and Hippocampus hippocampus (L., 1758) (Teleostei, Syngnathidae) in the Aegean Sea. Journal of Fish Biology, 72(6):1259–1267, April 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baldwin:2008:EFH

[2866] R. E. Baldwin, T. W. Miller, R. D. Brodeur, and K. C. Jacobson. Expanding the foraging history of juvenile Pacific salmon: combining stomach-content and macroparasite- community analyses for studying marine diets. Journal of Fish Biology, 72(6):1268–1294, April 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Klug:2008:OSR

[2867] H. Klug, C. I. Kontax, S. Annis, and N. Vasudevan. Operational sex ratio affects nest tending and aggression in male flagfish Jordanella floridae Goode & Bean. Journal of Fish Biology, 72(6):1295–1305, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lattuca:2008:TIA

[2868] M. E. Lattuca, M. A. Battini, and P. J. Macchi. Trophic interactions among native and introduced fishes in a northern Patagonian oligotrophic lake. Journal of Fish Biology, 72(6):1306–1320, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mercaldo-Allen:2008:RDB

[2869] R. Mercaldo-Allen, C. Kuropat, and E. M. Caldarone. An RNA:DNAbased growth model for young-of-the-year winter flounder Pseudopleuronectes americanus (Walbaum). Journal of Fish Biology, 72(6):1321– 1331, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Timi:2008:PBS

[2870] J. T. Timi, A. L. Lanfranchi, J. A. Etchegoin, and F. Cremonte. Parasites of the Brazilian sandperch Pinguipes brasilianus Cuvier: a tool for stock discrimination in the Argentine Sea. Journal of Fish Biology, 72(6):1332– 1342, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saidi:2008:RBS

[2871] B. Saïdi, M. N. Bradaï, and A. Bouaïn. Reproductive biology of the smooth-hound shark Mustelus mustelus (L.) in the Gulf of Gabès (southcentral Mediterranean Sea). Journal of Fish Biology, 72(6):1343–1354, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2008:SDC

[2872] M. C. P. Amorim, J. M. Simões, P. J. Fonseca, and G. F. Turner. Species differences in courtship acoustic signals among five Lake Malaŵi cichlid species (Pseudotropheus spp.). Journal of Fish Biology, 72(6):1355–1368, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ripley:2008:IDP

[2873] J. L. Ripley and C. M. Foran. Interspecific differences of parental polychlorinated biphenyl exposure on nutrient availability, egg production and brooding in two Syngnathus species. Journal of Fish Biology, 72 (6):1369–1385, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

dePerera:2008:RLS

[2874] T. Burt de Perera and T. C. Guilford. Rapid learning of shelter position in an intertidal fish, the shanny Lipophrys pholis L. Journal of Fish Biology, 72(6):1386–1392, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Studebaker:2008:TVF

[2875] R. S. Studebaker and T. J. Mulligan. Temporal variation and feeding ecology of juvenile Sebastes in rocky intertidal tidepools of northern California, with emphasis on Sebastes melanops Girard. Journal of Fish Biology, 72(6):1393–1405, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berens:2008:DDE

[2876] E. J. Berens and D. J. Murie. Differential digestion and evacuation rates of prey in a warm-temperate grouper, gag Mycteroperca microlepis (Goode & Bean). Journal of Fish Biology, 72(6):1406–1426, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bogetveit:2008:AGT

[2877] F. R. Bogetveit, A. Slotte, and A. Johannessen. The ability of gadoids to take advantage of a short-term high availability of forage fish: the example of spawning aggregations in Barents Sea capelin. Journal of Fish Biology, 72(6):1427–1449, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Filer:2008:EEE

[2878] J. L. Filer, C. G. Booker, and D. W. Sims. Effects of environment on electric field detection by small spotted catshark Scyliorhinus canicula (L.). Journal of Fish Biology, 72(6):1450–1462, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karachle:2008:ESS

[2879] P. K. Karachle and K. I. Stergiou. The effect of season and sex on trophic levels of marine fishes. Journal of Fish Biology, 72(6):1463–1487, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Speed:2008:SPR

[2880] C. W. Speed, M. G. Meekan, D. Rowat, S. J. Pierce, A. D. Marshall, and C. J. A. Bradshaw. Scarring patterns and relative mortality rates of Indian Ocean whale sharks. Journal of Fish Biology, 72(6):1488–1503, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Luchiari:2008:EAC

[2881] A. C. Luchiari and J. Pirhonen. Effects of ambient colour on colour preference and growth of juvenile rainbow trout Oncorhynchus mykiss (Walbaum). Journal of Fish Biology, 72(6):1504–1514, April 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oddone:2008:SMA

[2882] M. C. Oddone and A. F. Amorim. Size at maturity of Atlantoraja platana (Günther, 1880) (Chondrichthyes: Rajidae: Arhynchobatinae) in the south-west Atlantic Ocean. Journal of Fish Biology, 72(6):1515– 1519, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Castello:2008:NHA

[2883] L. Castello. Nesting habitat of Arapaima gigas (Schinz) in Amazonian floodplains. Journal of Fish Biology, 72(6):1520–1528, April 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2008:WDC

[2884] A. D. Nunn, J. P. Harvey, and I. G. Cowx. Winter diet and condition of two 0+ year cyprinid fish species in the lower River Trent, England. Journal of Fish Biology, 72(6):1529–1533, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Golani:2008:ASK

[2885] D. Golani, B. Appelbaum-Golani, and O. Gon. Apogon smithi (Kotthaus, 1970) (Teleostei: Apogonidae), a Red Sea cardinalfish colonizing the Mediterranean Sea. Journal of Fish Biology, 72(6):1534–1538, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

El-Regal:2008:FRP

[2886] M. Abu El-Regal and T. Kon. First record of the paedomorphic fish schindleria (Gobioidei, Schindleriidae) from the Red Sea. Journal of Fish Biology, 72(6):1539–1543, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Raine:2008:DIT

[2887] J. C. Raine and J. F. Leatherland. Distribution of immunoreactive thyroxine in the endostyle and thyroid tissue of the sea lamprey Petromyzon marinus L., 1758. Journal of Fish Biology, 72(6):1544–1552, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fahle:2008:MJV

[2888] S. R. Fahle and J. C. Thomason. Measurement of jaw viscoelasticity in newborn and adult lesser spotted dogfish Scyliorhinus canicula (L., 1758). Journal of Fish Biology, 72(6):1553–1557, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:Ed

[2889] Anonymous. Erratum. Journal of Fish Biology, 72(6):1558, April 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burnard:2008:RPF

[2890] D. Burnard, R. E. Gozlan, and S. W. Griffiths. The role of pheromones in freshwater fishes. Journal of Fish Biology, 73(1):1–16, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nakaya:2008:FSM

[2891] K. Nakaya, R. Matsumoto, and K. Suda. Feeding strategy of the megamouth shark Megachasma pelagios (Lamniformes: Megachasmidae). Journal of Fish Biology, 73(1):17–34, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chizinski:2008:BMZ

[2892] C. J. Chizinski, B. Sharma, K. L. Pope, and R. Patiño. A bioenergetic model for zebrafish Danio rerio (Hamilton). Journal of Fish Biology, 73 (1):35–43, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Etheridge:2008:CVP

[2893] E. C. Etheridge, C. Harrod, C. Bean, and C. E. Adams. Continuous variation in the pattern of marine v. freshwater foraging in brown trout Salmo trutta L. from Loch Lomond, Scotland. Journal of Fish Biology, 73(1):44–53, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prchalova:2008:DPF

[2894] M. Prchalová, J. Kubečka, M. Vašek, J. Peterka, J. Sed’a, T. Jůza, M. Říha, O. Jarolím, M. Tušer, M. Kratochvíl, M. Čech, V. Draštík, J. Frouzová, and E. Hohausová. Distribution patterns of fishes in a canyon-shaped reservoir. Journal of Fish Biology, 73(1):54–78, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neregaard:2008:WAS

[2895] L. Neregård, L. Sundt-Hansen, K. Hindar, S. Einum, J. I. Johnsson, R. H. Devlin, I. A. Fleming, and B. Th. Björnsson. Wild Atlantic salmon Salmo salar L. strains have greater growth potential than a domesticated strain selected for fast growth. Journal of Fish Biology, 73(1):79–95, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leunda:2008:FHE

[2896] P. M. Leunda, J. Oscoz, B. Elvira, A. Agorreta, S. Perea, and R. Miranda. Feeding habits of the exotic black bullhead Ameiurus melas (Rafinesque) in the Iberian Peninsula: first evidence of direct predation on native fish species. Journal of Fish Biology, 73(1):96–114, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caipang:2008:AAS

[2897] C. M. A. Caipang, N. Hynes, M. F. Brinchmann, K. Korsnes, and V. Kiron. Antimicrobial activity in serum of Atlantic cod Gadus morhua L.: comparison between cultured and wild fish. Journal of Fish Biology, 73(1):115–122, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monteiro:2008:LDG

[2898] J. Monteiro, R. Borges, J. Robalo, V. C. Almada, S. Henriques, and E. J. Gonçalves. Larval development of Gobius xanthocephalus with genetic validation of larval identification. Journal of Fish Biology, 73(1):123– 138, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hagen:2008:MPF

[2899] Ø Hagen, V. L. A. Vieira, C. Solberg, and I. A. Johnston. Myotube production in fast myotomal muscle is switched-off at shorter body lengths in male than female Atlantic halibut Hippoglossus hippoglossus (L.) resulting in a lower final fibre number. Journal of Fish Biology, 73(1): 139–152, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gagne:2008:MSI

[2900] S. Gagné and M. A. Rodríguez. Modelling seasonal increments in size to determine the onset of annual growth in fishes. Journal of Fish Biology, 73(1):153–168, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garner:2008:JGA

[2901] S. R. Garner, B. N. Madison, N. J. Bernier, and B. D. Neff. Juvenile growth and aggression in diploid and triploid Chinook salmon Oncorhynchus tshawytscha (Walbaum). Journal of Fish Biology, 73(1):169– 185, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Couturier:2008:IBE

[2902] C. S. Couturier, L. Nonnotte, G. Nonnotte, and G. Claireaux. Interaction between exopolysaccharide and oxygenation levels on habitat selection in the sole Solea solea (L.). Journal of Fish Biology, 73(1):186–195, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kreutzenberger:2008:IIB

[2903] K. Kreutzenberger, F. Leprieur, and S. Brosse. The influence of the invasive black bullhead Ameiurus melas on the predatory efficiency of pike Esox lucius L. Journal of Fish Biology, 73(1):196–205, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Figueiredo:2008:RSL

[2904] I. Figueiredo, T. Moura, A. Neves, and L. S. Gordo. Reproductive strategy of leafscale gulper shark Centrophorus squamosus and the Portuguese dogfish Centroscymnus coelolepis on the Portuguese continental slope. Journal of Fish Biology, 73(1):206–225, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Biro:2008:EMU

[2905] P. A. Biro, C. Beckmann, and M. S. Ridgway. Early microhabitat use by age 0 year brook charr Salvelinus fontinalis in lakes. Journal of Fish Biology, 73(1):226–240, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sepulveda:2008:ERM

[2906] C. A. Sepulveda, K. A. Dickson, D. Bernal, and J. B. Graham. Elevated red myotomal muscle temperatures in the most basal tuna species, Allothunnus fallai. Journal of Fish Biology, 73(1):241–249, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amundsen:2008:FCG

[2907] P.-A. Amundsen and H.-M. Gabler. Food consumption and growth of Atlantic salmon Salmo salar parr in sub-Arctic rivers: empirical support for food limitation and competition. Journal of Fish Biology, 73(1):250– 261, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fitzpatrick:2008:EET

[2908] J. L. Fitzpatrick and N. R. Liley. Ejaculate expenditure and timing of gamete release in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 73(1):262–274, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knudsen:2008:DPC

[2909] R. Knudsen, P.-A. Amundsen, M. Jobling, and A. Klemetsen. Differences in pyloric caeca morphology between Arctic charr Salvelinus alpinus ecotypes: adaptation to trophic specialization or parasite-induced phenotypic modifications? Journal of Fish Biology, 73(1):275–287, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aburto-Oropeza:2008:PSA

[2910] O. Aburto-Oropeza and P. M. Hull. A probable spawning aggregation of the leather bass Dermatolepis dermatolepis (Boulenger) in the Revillagigedo Archipelago, México. Journal of Fish Biology, 73(1):288–295, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braga:2008:FRF

[2911] A. C. Braga, P. A. S. Costa, and G. W. Nunan. First record of the firebrow lanternfish Diaphus adenomus (Myctophiformes: Myctophidae) from the South Atlantic. Journal of Fish Biology, 73(1):296–301, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Landaeta:2008:LDP

[2912] M. F. Landaeta, F. Balbontin, and G. A. Herrera. Larval development of the Patagonian brotula Cataetyx messieri (Pisces, Bythitidae) from fjords of southern Chile. Journal of Fish Biology, 73(1):302–310, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zanella:2008:RTU

[2913] D. Zanella, M. Mrakovčić, P. Mustafić, and M. Ćaleta. Recovery of Telestes ukliva, an endemic species from the Cetina River, Croatia (Cypriniformes, Cyprinidae). Journal of Fish Biology, 73(1):311–316, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miranda:2008:FRG

[2914] R. Miranda, D. Galicia, G. Pulido-Flores, and S. Monks. First record of Girardinichthys viviparus in Lake Tecocomulco, Mexico. Journal of Fish Biology, 73(1):317–322, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Flajshans:2008:CSP

[2915] M. Flajšhans, M. Rodina, K. Halačka, L. Vetešník, D. Gela, V. Lusková, and S. Lusk. Characteristics of sperm of polyploid Prussian carp Carassius gibelio. Journal of Fish Biology, 73(1):323–328, July 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zou:2008:RGD

[2916] S. M. Zou and X. Y. Jiang. Retracted: Gene duplication and functional evolution of Hox genes in fishes. Journal of Fish Biology, 73(2):329– 354, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jiraungkoorskul:2008:PIA

[2917] W. Jiraungkoorskul, S. Sahaphong, N. Kangwanrangsan, and S. Zakaria. The protective influence of ascorbic acid against the genotoxicity of waterborne lead exposure in Nile tilapia Oreochromis niloticus (L.). Journal of Fish Biology, 73(2):355–366, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stet:2008:CMH

[2918] R. J. M. Stet, K. Mudde, J. W. Wynne, A. Nooijen, T. G. Dahlgren, D. E. Ruzzante, and C. André. Characterization of a major histocompatibility class II A gene (Clha-DAA) with an embedded microsatellite marker in Atlantic herring (Clupea harengus L.). Journal of Fish Biology, 73(2): 367–381, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pfeiler:2008:SDH

[2919] E. Pfeiler, T. Watts, J. Pugh, and A. M. Van DerHeiden. Speciation and demographic history of the Cortez bonefish, Albula sp. A (Albuliformes: Albulidae), in the Gulf of California inferred from mitochondrial DNA. Journal of Fish Biology, 73(2):382–394, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goanvec:2008:FCE

[2920] C. Goanvec, M. Theron, T. Lacoue-Labarthe, E. Poirier, J. Guyomarch, S. Le-Floch, J. Laroche, L. Nonnotte, and G. Nonnotte. Flow cytometry for the evaluation of chromosomal damage in turbot Psetta maxima (L.) exposed to the dissolved fraction of heavy fuel oil in sea water: a comparison with classical biomarkers. Journal of Fish Biology, 73(2):395–413, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ki:2008:UMG

[2921] J.-S. Ki, S.-O. Jung, D.-S. Hwang, Y.-M. Lee, and J.-S. Lee. Unusual mitochondrial genome structure of the freshwater goby Odontobutis platycephala: rearrangement of tRNAs and an additional non-coding region. Journal of Fish Biology, 73(2):414–428, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vincent:2008:DSA

[2922] B. N. Vincent, B. F. Nowak, and R. N. Morrison. Detection of serum anti-Neoparamoeba spp. antibodies in amoebic gill disease-affected Atlantic salmon Salmo salar. Journal of Fish Biology, 73(2):429–435, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

OKeeffe:2008:SME

[2923] A. M. O’Keeffe, S. Hubert, M. Voisin, B. Houeix, D. Cotter, and M. T. Cairns. Somatolactin mRNA expression during the parr–smolt transformation in hatchery-reared Atlantic salmon Salmo salar smolts. Journal of Fish Biology, 73(2):436–443, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Delghandi:2008:TTN

[2924] M. Delghandi, M. S. Wesmajervi, T. Tafese, and F. Nilsen. Twentythree novel microsatellite markers developed from Atlantic cod Gadus morhua L. expressed sequence tags. Journal of Fish Biology, 73(2):444– 449, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahonen:2008:SJT

[2925] H. Ahonen and A. J. Stow. Shark jaws and teeth: an unexploited resource for population genetic studies. Journal of Fish Biology, 73(2): 450–455, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ocalewicz:2008:HSC

[2926] K. Ocalewicz, D. Fopp-Bayat, P. Woznicki, and M. Jankun. Heteromorphic sex chromosomes in the ninespine stickleback Pungitius pungitius. Journal of Fish Biology, 73(2):456–462, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Menezes:2008:MDS

[2927] M. R. Menezes, D. Noguchi, M. Nakajima, and N. Taniguchi. Microsatellite development and survey of genetic variation in skipjack tuna Katsuwonus pelamis. Journal of Fish Biology, 73(2):463–473, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:BRc

[2928] Anonymous. Book review. Journal of Fish Biology, 73(2):474–475, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Desta:2008:HUT

[2929] Z. Desta, R. Borgstrøm, Z. Gebremariam, and B. O. Rosseland. Habitat use and trophic position determine mercury concentration in the straight fin barb Barbus paludinosus, a small fish species in Lake Awassa, Ethiopia. Journal of Fish Biology, 73(3):477–497, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wormald:2008:TAM

[2930] C. L. Wormald and M. A. Steele. Testing assumptions of mark–recapture theory in the coral reef fish Lutjanus apodus. Journal of Fish Biology, 73(3):498–509, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hale:2008:AGR

[2931] L. F. Hale and C. G. Lowe. Age and growth of the round stingray Urobatis halleri at Seal Beach, California. Journal of Fish Biology, 73 (3):510–523, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brannas:2008:TRP

[2932] E. Brännäs. Temporal resource partitioning varies with individual competitive ability: a test with Arctic charr Salvelinus alpinus visiting a feeding site from a refuge. Journal of Fish Biology, 73(3):524–535, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tarvainen:2008:VBM

[2933] M. Tarvainen, A. Anttalainen, H. Helminen, T. Keskinen, J. Sarvala, I. Vaahto, and J. Karjalainen. A validated bioenergetics model for ruffe Gymnocephalus cernuus and its application to a northern lake. Journal of Fish Biology, 73(3):536–556, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braicovich:2008:PBT

[2934] P. E. Braicovich and J. T. Timi. Parasites as biological tags for stock discrimination of the Brazilian flathead Percophis brasiliensis in the southwest Atlantic. Journal of Fish Biology, 73(3):557–571, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Guerra:2008:HSW

[2935] R. R. Guerra, N. P. Santos, P. Cecarelli, J. R. M. C. Silva, and F. J. Hernandez-Blazquez. Healing of skin wounds in the African catfish Clarias gariepinus. Journal of Fish Biology, 73(3):572–583, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kobler:2008:SFS

[2936] A. Kobler, T. Klefoth, and R. Arlinghaus. Site fidelity and seasonal changes in activity centre size of female pike Esox lucius in a small lake. Journal of Fish Biology, 73(3):584–596, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2008:STV

[2937] N. G. Smith, C. M. Jones, and J. Van Montfrans. Spatial and temporal variability of juvenile spotted seatrout Cynoscion nebulosus growth in Chesapeake Bay. Journal of Fish Biology, 73(3):597–607, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Staples:2008:EMR

[2938] J. F. Staples, M. Kajimura, C. M. Wood, M. Patel, Y. K. Ip, and G. B. McClelland. Enzymatic and mitochondrial responses to 5 months of aerial exposure in the slender lungfish Protopterus dolloi. Journal of Fish Biology, 73(3):608–622, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valentin:2008:AEF

[2939] A. E. Valentin, X. Penin, J.-P. Chanut, J.-M. Sévigny, and F. J. Rohlf. Arching effect on fish body shape in geometric morphometric studies. Journal of Fish Biology, 73(3):623–638, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burton:2008:PIP

[2940] D. Burton. A physiological interpretation of pattern changes in a flatfish. Journal of Fish Biology, 73(3):639–649, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barry:2008:FEG

[2941] K. P. Barry, R. E. Condrey, W. B. Driggers III, and C. M. Jones. Feeding ecology and growth of neonate and juvenile blacktip sharks Carcharhinus limbatus in the Timbalier–Terrebone Bay complex, LA, U.S.A. Journal of Fish Biology, 73(3):650–662, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Green:2008:EDP

[2942] J. A. Green and R. W. Hardy. The effects of dietary protein : energy ratio and amino acid pattern on nitrogen utilization and excretion of rainbow trout Oncorhynchus mykiss (Walbaum). Journal of Fish Biology, 73(3): 663–682, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Salinas:2008:MPT

[2943] X. Salinas, M. T. González, and E. Acuña. Metazoan parasites of the thumb grenadier Nezumia pulchella, from the south-eastern Pacific, off Chile, and their use for discrimination of host populations. Journal of Fish Biology, 73(3):683–691, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fu:2008:PPO

[2944] S. J. Fu, Z. D. Cao, J. L. Peng, and Y. X. Wang. Is peak postprandial oxygen consumption positively related to growth rate and resting oxygen consumption in a sedentary catfish Silurus meridionalis? Journal of Fish Biology, 73(3):692–701, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lizardo-Daudt:2008:ECC

[2945] H. M. Lizardo-Daudt and C. Kennedy. Effects of cadmium chloride on the development of rainbow trout Oncorhynchus mykiss early life stages. Journal of Fish Biology, 73(3):702–718, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karam:2008:ATA

[2946] A. P. Karam, B. R. Kesner, and P. C. Marsh. Acoustic telemetry to assess post-stocking dispersal and mortality of razorback sucker Xyrauchen texanus. Journal of Fish Biology, 73(3):719–727, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gray:2008:SEE

[2947] S. M. Gray, J. S. McKinnon, F. Y. Tantu, and L. M. Dill. Sneaky egg-eating in Telmatherina sarasinorum, an endemic fish from Sulawesi. Journal of Fish Biology, 73(3):728–731, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gallucci:2008:IPS

[2948] V. F. Gallucci, R. J. Foy, S. M. O’Brien, A. Aires da Silva, H. Nesse, B. Langseth, N. Vega, I. Taylor, and K. J. Goldman. Information from a pregnant salmon shark Lamna ditropis in the eastern North Pacific with observations on oophagous reproduction. Journal of Fish Biology, 73(3):732–739, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lelli:2008:TBN

[2949] S. Lelli, F. Colloca, P. Carpentieri, and B. C. Russell. The threadfin bream Nemipterus randalli (Perciformes: Nemipteridae) in the eastern Mediterranean Sea. Journal of Fish Biology, 73(3):740–745, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rushbrook:2008:CNB

[2950] B. J. Rushbrook and I. Barber. A comparison of nest building by threespined sticklebacks Gasterosteus aculeatus from still and flowing waters. Journal of Fish Biology, 73(3):746–752, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jones:2008:LNB

[2951] A. M. Jones, S. Gardner, and W. Sinclair. Losing ‘Nemo’: bleaching and collection appear to reduce inshore populations of anemonefishes. Journal of Fish Biology, 73(3):753–761, August 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chassot:2008:ITL

[2952] E. Chassot, T. Rouyer, V. M. Trenkel, and D. Gascuel. Investigating trophic-level variability in Celtic Sea fish predators. Journal of Fish Biology, 73(4):763–781, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perez-Jimenez:2008:RBB

[2953] J. C. Pérez-Jiménez and O. Sosa-Nishizaki. Reproductive biology of the brown smoothhound shark Mustelus henlei, in the northern Gulf of California, México. Journal of Fish Biology, 73(4):782–792, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Persson:2008:PSL

[2954] A. Persson and C. Brönmark. Pikeperch Sander lucioperca trapped between niches: foraging performance and prey selection in a piscivore on a planktivore diet. Journal of Fish Biology, 73(4):793–808, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Russo:2008:RBB

[2955] T. Russo, D. Pulcini, Á. O’Leary, S. Cataudella, and S. Mariani. Relationship between body shape and trophic niche segregation in two closely related sympatric fishes. Journal of Fish Biology, 73(4):809–828, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barthel:2008:DLH

[2956] B. L. Barthel, S. J. Cooke, J. H. Svec, C. D. Suski, C. M. Bunt, F. J. S. Phelan, and D. P. Philipp. Divergent life histories among smallmouth bass Micropterus dolomieu inhabiting a connected river–lake system. Journal of Fish Biology, 73(4):829–852, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tonkin:2008:OIW

[2957] Z. Tonkin, A. J. King, and D. S. L. Ramsey. Otolith increment width responses of juvenile Australian smelt Retropinna semoni to sudden changes in food levels: the importance of feeding history. Journal of Fish Biology, 73(4):853–860, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamamoto:2008:IBH

[2958] S. Yamamoto, H. Nakamura, and K. Koga. Interaction between hatchery and wild juvenile white-spotted charr Salvelinus leucomaenis in a stream enclosure experiment. Journal of Fish Biology, 73(4):861–869, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loher:2008:ECL

[2959] T. Loher, S. Wischniowski, and G. B. Martin. Elemental chemistry of left and right sagittal otoliths in a marine fish Hippoglossus stenolepis displaying cranial asymmetry. Journal of Fish Biology, 73(4):870–887, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kenaley:2008:DVM

[2960] C. P. Kenaley. Diel vertical migration of the loosejaw dragonfishes (Stomiiformes: Stomiidae: Malacosteinae): a new analysis for rare pelagic taxa. Journal of Fish Biology, 73(4):888–901, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medina:2008:HFB

[2961] A. Medina, J.-C. Brêthes, and J.-M. Sévigny. Habitat fragmentation and body-shape variation of African hind Cephalopholis taeniops (Valenciennes) in an archipelago system (Cape Verde, eastern Atlantic Ocean). Journal of Fish Biology, 73(4):902–925, September 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stickler:2008:HUA

[2962] M. Stickler, E. C. Enders, C. J. Pennell, D. Cote, K. T. Alfredsen, and D. A. Scruton. Habitat use of Atlantic salmon Salmo salar parr in a dynamic winter environment: the influence of anchor-ice dams. Journal of Fish Biology, 73(4):926–944, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Peyronnet:2008:DOC

[2963] A. Peyronnet, K. D. Friedland, and N. Ó Maoileidigh. Different ocean and climate factors control the marine survival of wild and hatchery Atlantic salmon Salmo salar in the north-east Atlantic Ocean. Journal of Fish Biology, 73(4):945–962, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kendall:2008:RBT

[2964] B. W. Kendall and C. A. Gray. Reproductive biology of two co-occurring mugilids, Liza argentea and Myxus elongatus, in south-eastern Australia. Journal of Fish Biology, 73(4):963–979, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papoutsoglou:2008:EMS

[2965] S. E. Papoutsoglou, N. Karakatsouli, A. Batzina, E. S. Papoutsoglou, and A. Tsopelakos. Effect of music stimulus on gilthead seabream Sparus aurata physiology under different light intensity in a re-circulating water system. Journal of Fish Biology, 73(4):980–1004, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ben-Tzvi:2008:HSC

[2966] O. Ben-Tzvi, A. Abelson, O. Polak, and M. Kiflawi. Habitat selection and the colonization of new territories by Chromis viridis. Journal of Fish Biology, 73(4):1005–1018, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kovacic:2008:LCM

[2967] M. Kovačić. Live colouration, morphology and habitat of Vanneaugobius dollfusi (Gobiidae) in the northern Adriatic Sea. Journal of Fish Biology, 73(4):1019–1023, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Medeiros:2008:DVF

[2968] E. S. F. Medeiros and A. H. Arthington. Diet variation in food intake and diet composition of three native fish species in floodplain lagoons of the Macintyre River, Australia. Journal of Fish Biology, 73(4):1024–1032, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weyl:2008:VAF

[2969] O. L. F. Weyl and A. J. Booth. Validation of annulus formation in otoliths of a temperate population of adult African sharptooth catfish Clarias gariepinus using fluorochrome marking of wild fish. Journal of Fish Biology, 73(4):1033–1038, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ahnesjo:2008:BTP

[2970] I. Ahnesjö. Behavioural temperature preference in a brooding male pipefish Syngnathus typhle. Journal of Fish Biology, 73(4):1039–1045, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sato:2008:STR

[2971] T. Sato and Y. Kano. A simple technique for remote identification of individual spawning fishes. Journal of Fish Biology, 73(4):1046–1052, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Binoy:2008:IHF

[2972] V. V. Binoy and K. J. Thomas. The influence of hunger on food-stocking behaviour of climbing perch Anabas testudineus. Journal of Fish Biology, 73(4):1053–1057, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Knight:2008:CSH

[2973] C. M. Knight, R. E. Gozlan, and M. C. Lucas. Can seasonal homerange size in pike Esox lucius predict excursion distance? Journal of Fish Biology, 73(4):1058–1064, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2008:CSD

[2974] J. G. Davidsen, N. Plantalech Manel-la, F. Økland, O. H. Diserud, E. B. Thorstad, B. Finstad, R. Sivertsgård, R. S. McKinley, and A. H. Rikardsen. Changes in swimming depths of Atlantic salmon Salmo salar postsmolts relative to light intensity. Journal of Fish Biology, 73(4):1065– 1074, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claeson:2008:ORB

[2975] K. M. Claeson and J. W. Hagadorn. The occipital region in the basal bony fish Erpetoichthys calabaricus (Actinopterygii: Cladistia). Journal of Fish Biology, 73(4):1075–1082, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:BRd

[2976] Anonymous. Book reviews. Journal of Fish Biology, 73(4):1083–1089, September 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maclean:2008:EMC

[2977] N. Maclean. Editorial for Molecular and Cell Science JFB. Journal of Fish Biology, 73(5):1091–1095, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasselman:2008:CWP

[2978] D. J. Hasselman. Comments on ‘Weak population differentiation in northern European populations of the endangered anadromous clupeid Alosa fallax’. Journal of Fish Biology, 73(5):1096–1098, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2008:GGR

[2979] C.-C. Chen and R. D. Fernald. GnRH and GnRH receptors: distribution, function and evolution. Journal of Fish Biology, 73(5):1099–1120, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2008:UZD

[2980] J. Chen and J. D. Griffin. Using a zebrafish Danio rerio model system to study NOTCH1-induced T-cell leukaemia. Journal of Fish Biology, 73 (5):1121–1138, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mwita:2008:MPC

[2981] C. J. Mwita and G. Nkwengulila. Molecular phylogeny of the clariid fishes of Lake Victoria, Tanzania, inferred from cytochrome b DNA sequences. Journal of Fish Biology, 73(5):1139–1148, October 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xie:2008:CIR

[2982] F. J. Xie, Z. P. Zhang, P. Lin, S. H. Wang, Z. H. Zou, and Y. L. Wang. Cloning and infection response of tumour-necrosis factor alpha in large yellow croaker Pseudosciaena crocea (Richardson). Journal of Fish Biology, 73(5):1149–1160, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jia:2008:IMD

[2983] Z. Y. Jia, L. Y. Shi, X. W. Sun, and Q. Q. Lei. Inheritance of microsatellite DNA in bisexual gynogenesis complex of Fangzheng silver crucian carp, Carassius auratus gibelio (Bloch). Journal of Fish Biology, 73(5):1161–1169, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2008:MAS

[2984] P. J. Smith, D. Steinke, S. M. Mcveagh, A. L. Stewart, C. D. Struthers, and C. D. Roberts. Molecular analysis of Southern Ocean skates (Bathyraja) reveals a new species of Antarctic skate. Journal of Fish Biology, 73(5):1170–1182, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Byrkjedal:2008:TST

[2985] I. Byrkjedal, D. J. Rees, J. S. Christiansen, and S.-E. Fevolden. The taxonomic status of Theragra finnmarchica Koefoed, 1956 (Teleostei: Gadidae): perspectives from morphological and molecular data. Journal of Fish Biology, 73(5):1183–1200, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nilsson:2008:GAA

[2986] J. Nilsson, J. Östergren, H. Lundqvist, and U. Carlsson. Genetic assessment of Atlantic salmon Salmo salar and sea trout Salmo trutta stocking in a Baltic Sea river. Journal of Fish Biology, 73(5):1201–1215, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Milana:2008:MDA

[2987] V. Milana, L. Sola, L. Congiu, and A. R. Rossi. Mitochondrial DNA in Atherina (Teleostei, Atheriniformes): differential distribution of an intergenic spacer in lagoon and marine forms of Atherina boyeri. Journal of Fish Biology, 73(5):1216–1227, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bektas:2008:MPT

[2988] Y. Bektas and A. O. Belduz. Molecular phylogeny of Turkish Trachurus species (Perciformes: Carangidae) inferred from mitochondrial DNA analyses. Journal of Fish Biology, 73(5):1228–1248, October 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhao:2008:MDC

[2989] Y. Zhao, X. Zhu, and X. Sun. Microsatellite diversity in cultured populations of rainbow trout Oncorhynchus mykiss in China. Journal of Fish Biology, 73(5):1249–1255, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoon:2008:MDV

[2990] M. Yoon, S. Sato, J. E. Seeb, V. Brykov, L. W. Seeb, N. V. Varnavskaya, R. L. Wilmot, D. H. Jin, S. Urawa, A. Urano, and S. Abe. Mitochondrial DNA variation and genetic population structure of chum salmon Oncorhynchus keta around the Pacific rim. Journal of Fish Biology, 73 (5):1256–1266, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Amorim:2008:VMC

[2991] M. C. P. Amorim and R. O. Vasconcelos. Variability in the mating calls of the Lusitanian toadfish Halobatrachus didactylus: cues for potential individual recognition. Journal of Fish Biology, 73(6):1267–1283, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaza:2008:TPG

[2992] G. Plaza, M. Ishida, and D. Aoyama. Temporal patterns of growth in larval cohorts of the Japanese sardine Sardinops melanostictus in a coastal nursery area. Journal of Fish Biology, 73(6):1284–1300, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Good:2008:MBV

[2993] J. P. Good, A. Wells, and N. Hazon. Measurement of blood volume in the elasmobranch fish Scyliorhinus canicula following acute and long-term salinity transfers. Journal of Fish Biology, 73(6):1301–1313, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barletta:2008:FAS

[2994] M. Barletta, C. S. Amaral, M. F. M. Corrêa, F. Guebert, D. V. Dantas, L. Lorenzi, and U. Saint-Paul. Factors affecting seasonal variations in demersal fish assemblages at an ecocline in a tropical–subtropical estuary. Journal of Fish Biology, 73(6):1314–1336, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Claes:2008:EDB

[2995] J. M. Claes and J. Mallefet. Early development of bioluminescence suggests camouflage by counter-illumination in the velvet belly lantern shark Etmopterus spinax (Squaloidea: Etmopteridae). Journal of Fish Biology, 73(6):1337–1350, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cooke:2008:EDC

[2996] S. J. Cooke, C. D. Suski, S. E. Danylchuk, A. J. Danylchuk, M. R. Donaldson, C. Pullen, G. Bulté, A. O’Toole, K. J. Murchie, J. B. Koppelman, A. D. Shultz, E. Brooks, and T. L. Goldberg. Effects of different capture techniques on the physiological condition of bonefish Albula vulpes evaluated using field diagnostic tools. Journal of Fish Biology, 73(6): 1351–1375, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Monteiro:2008:QHO

[2997] S. M. Monteiro, E. Rocha, A. Fontaínhas-Fernandes, and M. Sousa. Quantitative histopathology of Oreochromis niloticus gills after copper exposure. Journal of Fish Biology, 73(6):1376–1392, October 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gunnarsson:2008:GMF

[2998] Á. Gunnarsson, E. Hjörleifsson, K. Thórarinsson, and G. Marteinsdóttir. Growth, maturity and fecundity of female spotted wolffish Anarhichas minor in Icelandic waters. Journal of Fish Biology, 73(6):1393–1406, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hautala:2008:ASD

[2999] A. Hautala. Autumnal shift from diurnal to nocturnal peaking feeding activity of Rutilus rutilus in boreal lake littoral zones. Journal of Fish Biology, 73(6):1407–1418, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Coelho:2008:LHW

[3000] R. Coelho and K. Erzini. Life history of a wide-ranging deepwater lantern shark in the north-east Atlantic, Etmopterus spinax (Chondrichthyes: Etmopteridae), with implications for conservation. Journal of Fish Biology, 73(6):1419–1443, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horne:2008:AOT

[3001] J. K. Horne. Acoustic ontogeny of a teleost. Journal of Fish Biology, 73 (6):1444–1463, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adams:2008:ETS

[3002] R. D. Adams and U. G. Reinhardt. Effects of texture on surface attachment of spawning-run sea lampreys Petromyzon marinus: a quantitative analysis. Journal of Fish Biology, 73(6):1464–1472, October 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chapman:2008:PLB

[3003] D. D. Chapman, B. Firchau, and M. S. Shivji. Parthenogenesis in a largebodied requiem shark, the blacktip Carcharhinus limbatus. Journal of Fish Biology, 73(6):1473–1477, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kristensen:2008:OFM

[3004] P. B. Kristensen, G. P. Closs, P. M. Lokman, and P. Grønkjær. Otolith formation, microstructure and daily increment validation in juvenile perch Perca fluviatilis. Journal of Fish Biology, 73(6):1478–1483, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nanami:2008:FRS

[3005] A. Nanami and H. Yamada. Foraging rates and substratum selection in foraging activity of checkered snapper Lutjanus decussatus (Lutjanidae) in an Okinawan coral reef. Journal of Fish Biology, 73(6):1484–1488, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kamler:2008:TB

[3006] E. Kamler, K. Goryczko, E. D. Le Cren, and J. E. Thorpe. Tadeusz Backiel 1922–2008. Journal of Fish Biology, 73(6):1489–1490, October 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Donaldson:2008:CSF

[3007] M. R. Donaldson, S. J. Cooke, D. A. Patterson, and J. S. Macdonald. Cold shock and fish. Journal of Fish Biology, 73(7):1491–1530, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Plaza:2008:GMR

[3008] G. Plaza and M. Ishida. The growth–mortality relationship in larval cohorts of Sardinops melanostictus, revealed by using two new approaches to analyse longitudinal data from otoliths. Journal of Fish Biology, 73(7): 1531–1553, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tillett:2008:FBP

[3009] B. J. Tillett, I. R. Tibbetts, and D. L. Whithead. Foraging behaviour and prey discrimination in the bluespotted maskray Dasyatis kuhlii. Journal of Fish Biology, 73(7):1554–1561, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gibbs:2008:RIE

[3010] M. A. Gibbs, J. H. Shields, D. W. Lock, K. M. Talmadge, and T. M. Farrell. Reproduction in an invasive exotic catfish Pterygoplichthys disjunctivus in Volusia Blue Spring, Florida, U.S.A. Journal of Fish Biology, 73(7):1562–1572, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoysak:2008:LPE

[3011] D. J. Hoysak and N. E. Stacey. Large and persistent effect of a female steroid pheromone on ejaculate size in goldfish Carassius auratus. Journal of Fish Biology, 73(7):1573–1584, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watzin:2008:SFP

[3012] M. C. Watzin, K. Joppe-Mercure, J. Rowder, B. Lancaster, and L. Bronson. Significant fish predation on zebra mussels Dreissena polymorpha in Lake Champlain, U.S.A. Journal of Fish Biology, 73(7):1585–1599, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Locatello:2008:ETR

[3013] L. Locatello, M. B. Rasotto, B. Adriaenssens, and A. Pilastro. Ejaculate traits in relation to male body size in the eastern mosquitofish Gambusia holbrooki. Journal of Fish Biology, 73(7):1600–1611, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shaw:2008:ADC

[3014] M. Shaw, R. Diekmann, J. Van DerKooij, S. Milligan, P. Bromley, and D. Righton. Assessment of the diets of cod Gadus morhua and whiting Merlangius merlangus juveniles in a frontal region close to the Norwegian trench: co-existence or competition? Journal of Fish Biology, 73(7): 1612–1634, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rodriguez:2008:RIL

[3015] M. Rodriguez and A. Miquelarena. Rineloricaria isaaci (Loricariidae: Loricariinae), a new species of loricariid catfish from the Uruguay River basin. Journal of Fish Biology, 73(7):1635–1647, November 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mukai:2008:DSO

[3016] Y. Mukai, A. D. Tuzan, L. S. Lim, N. Wahid, M. S. Sitti Raehanah, and S. Senoo. Development of sensory organs in larvae of African catfish Clarias gariepinus. Journal of Fish Biology, 73(7):1648–1661, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Corrigan:2008:GRE

[3017] S. Corrigan, C. Huveneers, T. S. Schwartz, R. G. Harcourt, and L. B. Beheregaray. Genetic and reproductive evidence for two species of ornate wobbegong shark Orectolobus spp. on the Australian east coast. Journal of Fish Biology, 73(7):1662–1675, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Robinson:2008:AFC

[3018] E. Robinson and W. Davison. Antarctic fish can survive prolonged exposure to elevated temperatures. Journal of Fish Biology, 73(7):1676–1689, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lessa:2008:OMA

[3019] R. Lessa, P. Duarte-Neto, E. Morize, and R. Maciel. Otolith microstructure analysis with OTC validation confirms age overestimation in Atlantic thread herring Opisthonema oglinum from north-eastern Brazil. Journal of Fish Biology, 73(7):1690–1700, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manju:2008:CDP

[3020] M. Manju, T. G. Sherin, K. N. Rajeesha, P. Sreejith, K. N. Rajasekharan, and O. V. Oommen. Curcumin and its derivatives prevent hepatocyte lipid peroxidation in Anabas testudineus. Journal of Fish Biology, 73(7): 1701–1713, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Keller:2008:BIB

[3021] K. Keller and C. Brown. Behavioural interactions between the introduced plague minnow Gambusia holbrooki and the vulnerable native Australian ornate rainbowfish Rhadinocentrus ornatus, under experimental conditions. Journal of Fish Biology, 73(7):1714–1729, November 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pratchett:2008:IVD

[3022] M. S. Pratchett and M. L. Berumen. Interspecific variation in distributions and diets of coral reef butterflyfishes (Teleostei: Chaetodontidae). Journal of Fish Biology, 73(7):1730–1747, November 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

MacNeil:2008:DHU

[3023] M. A. MacNeil, N. A. J. Graham, M. J. Conroy, C. J. Fonnesbeck, N. V. C. Polunin, S. P. Rushton, P. Chabanet, and T. R. McClanahan. Detection heterogeneity in underwater visual-census data. Journal of Fish Biology, 73(7):1748–1763, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoglund:2008:PSC

[3024] E. Höglund, H.-M. Gjøen, T. G. Pottinger, and Ø. Øverli. Parental stress-coping styles affect the behaviour of rainbow trout Oncorhynchus mykiss at early developmental stages. Journal of Fish Biology, 73(7): 1764–1769, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kawabata:2008:PRP

[3025] Y. Kawabata, J. Okuyama, K. Asami, K. Yoseda, and N. Arai. The post-release process of establishing stable home ranges and diel movement patterns of hatchery-reared black-spot tuskfish Choerodon schoenleinii. Journal of Fish Biology, 73(7):1770–1782, November 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mateus:2008:CSS

[3026] C. S. Mateus, B. R. Quintella, and P. R. Almeida. The critical swimming speed of Iberian barbel Barbus bocagei in relation to size and sex. Journal of Fish Biology, 73(7):1783–1789, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Serrano:2008:SCD

[3027] R. M. Serrano, O. Lopes, P. C. Hubbard, J. Araújo, A. V. M. Canário, and E. N. Barata. Seasonal cell differentiation and olfactory potency of secretions by the anal glands of male peacock blenny Salaria pavo. Journal of Fish Biology, 73(7):1790–1798, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lassalle:2008:LVA

[3028] G. Lassalle, T. Trancart, P. Lambert, and E. Rochard. Latitudinal variations in age and size at maturity among allis shad Alosa alosa populations. Journal of Fish Biology, 73(7):1799–1809, November 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jensen:2008:DNR

[3029] J. L. A. Jensen and A. H. Rikardsen. Do northern riverine anadromous Arctic charr Salvelinus alpinus and sea trout Salmo trutta overwinter in estuarine and marine waters? Journal of Fish Biology, 73(7):1810–1818, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lipej:2008:LTT

[3030] L. Lipej, B. Mavrič, V. Žiža, and J. Dulčić. The largescaled terapon Terapon theraps: a new Indo–Pacific fish in the Mediterranean Sea. Journal of Fish Biology, 73(7):1819–1822, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Giacomello:2008:SDG

[3031] E. Giacomello, L. Marri, D. Marchini, C. Mazzoldi, and M. B. Rasotto. Sperm-duct gland secretion of the grass goby Zosterisessor ophiocephalus exhibits antimicrobial activity. Journal of Fish Biology, 73(7):1823–1828, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:BRe

[3032] Anonymous. Book reviews. Journal of Fish Biology, 73(7):1829–1831, November 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:2008:LIE

[3033] C. J. Hall, M. V. Flores, K. E. Crosier, and P. S. Crosier. Live imaging early immune cell ontogeny and function in zebrafish Danio rerio. Journal of Fish Biology, 73(8):1833–1871, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2008:MRF

[3034] Y. H. Chen and H. J. Tsai. Myogenic regulatory factors Myf5 and Mrf4 of fish: current status and perspective. Journal of Fish Biology, 73(8): 1872–1890, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Braasch:2008:ETP

[3035] I. Braasch, J.-N. Volff, and M. Schartl. The evolution of teleost pigmentation and the fish-specific genome duplication. Journal of Fish Biology, 73(8):1891–1918, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barrientos-Villalobos:2008:GVR

[3036] J. Barrientos-Villalobos and A. Espinosa de los Monteros. Genetic variation and recent population history of the tropical gar Atractosteus tropicus Gill (Pisces: Lepisosteidae). Journal of Fish Biology, 73(8):1919– 1936, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saito:2008:HGF

[3037] T. Saito, S. Washio, K. Dairiki, M. Shimojo, S. Itoi, and H. Sugita. High gene flow in Girella punctata (Perciformes, Kyphosidae) among the Japanese Islands inferred from partial sequence of the control region in mitochondrial DNA. Journal of Fish Biology, 73(8):1937–1945, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ma:2008:ESE

[3038] L. Ma, X. F. Pan, Y. H. Wei, Z. Y. Li, C. C. Li, J. X. Yang, and B. Y. Mao. Embryonic stages and eye-specific gene expression of the local cyprinoid fish Anabarilius grahami in Fuxian Lake, China. Journal of Fish Biology, 73(8):1946–1959, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feng:2008:IEC

[3039] S. Z. Feng, W. S. Li, and H. R. Lin. Identification and expression characterization of pepsinogen a in orange-spotted grouper, Epinephelus coioides. Journal of Fish Biology, 73(8):1960–1978, December 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Xia:2008:SPG

[3040] J.-H. Xia, J.-H. Huang, J.-B. Gong, and S.-G. Jiang. Significant population genetic structure of yellowfin seabream Acanthopagrus latus in China. Journal of Fish Biology, 73(8):1979–1992, December 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2008:CPM

[3041] J. Lin, E. Ziegler, T. P. Quinn, and L. Hauser. Contrasting patterns of morphological and neutral genetic divergence among geographically proximate populations of sockeye salmon Oncorhynchus nerka in Lake Aleknagik, Alaska. Journal of Fish Biology, 73(8):1993–2004, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yang:2008:NSG

[3042] J. Yang, X. Y. Chen, and J. X. Yang. A new species of the genus Mekongina Fowler, 1937 (Cypriniformes: Cyprinidae) from South China. Journal of Fish Biology, 73(8):2005–2011, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kantek:2008:CAC

[3043] D. L. Z. Kantek, R. B. Noleto, F. A. M. Maurutto, L. A. C. Bertollo, O. Moreira-Filho, and M. M. Cestari. Cytotaxonomy of Astyanax (Characiformes, Characidae) from the Upper Iguaçu River Basin: confirmation of the occurrence of distinct evolutionary units. Journal of Fish Biology, 73(8):2012–2020, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berendzen:2008:PBC

[3044] P. B. Berendzen, T. Gamble, and A. M. Simons. Phylogeography of the bigeye chub Hybopsis amblops (Teleostei: Cypriniformes): early Pleistocene diversification and post-glacial range expansion. Journal of Fish Biology, 73(8):2021–2039, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gomez-Uchida:2008:GDB

[3045] D. Gomez-Uchida, K. P. Dunphy, M. F. O’Connell, and D. E. Ruzzante. Genetic divergence between sympatric Arctic charr Salvelinus alpinus morphs in Gander Lake, Newfoundland: roles of migration, mutation and unequal effective population sizes. Journal of Fish Biology, 73(8): 2040–2057, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhou:2008:ECS

[3046] G. Z. Zhou, L. Gui, Z. Q. Li, X. P. Yuan, and Q. Y. Zhang. Establishment of a Chinese sturgeon Acipenser sinensis tail-fin cell line and its susceptibility to frog iridovirus. Journal of Fish Biology, 73(8):2058–2067, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lage:2008:EMP

[3047] C. R. Lage, C. W. Petersen, D. Forest, D. Barnes, I. Kornfield, and C. Wray. Evidence of multiple paternity in spiny dogfish (Squalus acanthias) broods based on microsatellite analysis. Journal of Fish Biology, 73(8):2068–2074, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Prodocimo:2008:LGD

[3048] V. Prodocimo, M. K. Tschá, M. R. Pie, J. F. Oliveira-Neto, A. Ostrensky, and W. A. Boeger. Lack of genetic differentiation in the fat snook Centropomus parallelus (Teleostei: Centropomidae) along the Brazilian coast. Journal of Fish Biology, 73(8):2075–2082, December 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lafferty:2008:ECF

[3049] K. D. Lafferty. Ecosystem consequences of fish parasites. Journal of Fish Biology, 73(9):2083–2093, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grutter:2008:IBG

[3050] A. S. Grutter. Interactions between gnathiid isopods, cleaner fish and other fishes on lizard Island, Great Barrier Reef. Journal of Fish Biology, 73(9):2094–2109, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wegner:2008:HCS

[3051] K. M. Wegner. Historical and contemporary selection of teleost MHC genes: did we leave the past behind? Journal of Fish Biology, 73(9): 2110–2132, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sures:2008:HPI

[3052] B. Sures. Host–parasite interactions in polluted environments. Journal of Fish Biology, 73(9):2133–2142, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feist:2008:HFP

[3053] S. W. Feist and M. Longshaw. Histopathology of fish parasite infections — importance for populations. Journal of Fish Biology, 73(9):2143–2160, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powell:2008:NGI

[3054] M. D. Powell, M. J. Leef, S. D. Roberts, and M. A. Jones. Neoparamoebic gill infections: host response and physiology in salmonids. Journal of Fish Biology, 73(9):2161–2183, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wahli:2008:RAW

[3055] T. Wahli, D. Bernet, H. Segner, and H. Schmidt-Posthaus. Role of altitude and water temperature as regulating factors for the geographical distribution of Tetracapsuloides bryosalmonae infected fishes in Switzerland. Journal of Fish Biology, 73(9):2184–2197, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winger:2008:GSI

[3056] A. C. Winger, R. Primicerio, R. Kristoffersen, S. I. Siikavuopio, and R. Knudsen. Gyrodactylus salaris infecting allopatric Arctic charr Salvelinus alpinus fry: an experimental study of host survival. Journal of Fish Biology, 73(9):2198–2209, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simkova:2008:CEP

[3057] A. Šimková and S. Morand. Co-evolutionary patterns in congeneric monogeneans: a review of Dactylogyrus species and their cyprinid hosts. Journal of Fish Biology, 73(9):2210–2227, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Voutilainen:2008:EEF

[3058] A. Voutilainen, K. Figueiredo, and H. Huuskonen. Effects of the eye fluke Diplostomum spathaceum on the energetics and feeding of Arctic charr Salvelinus alpinus. Journal of Fish Biology, 73(9):2228–2237, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

James:2008:CET

[3059] C. T. James, K. J. Noyes, A. D. Stumbo, B. D. Wisenden, and C. P. Goater. Cost of exposure to trematode cercariae and learned recognition and avoidance of parasitism risk by fathead minnows Pimephales promelas. Journal of Fish Biology, 73(9):2238–2248, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cowx:2008:ELI

[3060] I. G. Cowx, D. Rollins, and R. Tumwebaze. Effect of Ligula intestinalis on the reproductive capacity of Rastrineobola argentea in Lake Victoria. Journal of Fish Biology, 73(9):2249–2260, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Unger:2008:TNA

[3061] J. Unger, A. Brinker, and H. B. Stich. Transmission of neozoic Anguillicoloides crassus and established Camallanus lacustris in ruffe Gymnocephalus cernuus. Journal of Fish Biology, 73(9):2261–2273, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Garcia:2008:MPS

[3062] A. Garcia, M. N. Santos, S. Damiano, G. Nascetti, and S. Mattiucci. The metazoan parasites of swordfish from Atlantic tropical–equatorial waters. Journal of Fish Biology, 73(9):2274–2287, December 2008. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saraiva:2008:FRP

[3063] A. Saraiva, M. Hermida, M. J. Costa, C. Maia, A. R. Reis, C. Cruz, and A. Valente. First record of Philometra ovata (Nematoda) infection in Gobio lozanoi in Portugal. Journal of Fish Biology, 73(9):2288–2292, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sebastian:2008:CPS

[3064] H. Sebastian, P. A. Haye, and M. S. Shivji. Characterization of the pelagic shark-fin trade in north-central Chile by genetic identification and trader surveys. Journal of Fish Biology, 73(10):2293–2304, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vignon:2008:ESL

[3065] M. Vignon, F. Morat, R. Galzin, and P. Sasal. Evidence for spatial limitation of the bluestripe snapper Lutjanus kasmira in French Polynesia from parasite and otolith shape analysis. Journal of Fish Biology, 73(10): 2305–2320, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hubley:2008:CSC

[3066] P. B. Hubley, P. G. Amiro, and A. J. F. Gibson. Changes in scale circulus spacings of an endangered Atlantic salmon Salmo salar population: evidence of a shift in marine migration? Journal of Fish Biology, 73(10): 2321–2340, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neregaard:2008:GHA

[3067] L. Neregård, L. Sundt-Hansen, B. Th. Björnsson, and J. I. Johnsson. Growth hormone affects behaviour of wild brown trout Salmo trutta in territorial owner–intruder conflicts. Journal of Fish Biology, 73(10):2341– 2351, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartman:2008:CRA

[3068] K. J. Hartman and F. J. Margraf. Common relationships among proximate composition components in fishes. Journal of Fish Biology, 73(10): 2352–2360, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brodeur:2008:NEK

[3069] N. N. Brodeur, M. V. Noël, O. Venter, L. Bernatchez, S. Dayanandan, and J. W. A. Grant. No evidence of kin bias in dispersion of young-of-the-year Atlantic salmon Salmo salar L. in a natural stream. Journal of Fish Biology, 73(10):2361–2370, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parravicini:2008:DMH

[3070] V. Parravicini, M. Donato, C. Morri, E. Villa, and C. N. Bianchi. Date mussel harvesting favours some blennioids. Journal of Fish Biology, 73 (10):2371–2379, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Little:2008:MTS

[3071] T. J. Little, M. Perutz, M. Palmer, C. Crossan, and V. A. Braithwaite. Male three-spined sticklebacks Gasterosteus aculeatus make antibiotic nests: a novel form of parental protection? Journal of Fish Biology, 73(10):2380–2389, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carreau-Green:2008:OCM

[3072] N. D. Carreau-Green, R. S. Mirza, M. L. Martínez, and G. G. Pyle. The ontogeny of chemically mediated antipredator responses of fathead minnows Pimephales promelas. Journal of Fish Biology, 73(10):2390– 2401, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bautista-Vega:2008:DDS

[3073] A. A. Bautista-Vega, Y. Letourneur, M. Harmelin-Vivien, and C. Salen-Picard. Difference in diet and size-related trophic level in two sympatric fish species, the red mullets Mullus barbatus and Mullus surmuletus, in the Gulf of Lions (north-west Mediterranean Sea). Journal of Fish Biology, 73(10):2402–2420, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsunagawa:2008:FMJ

[3074] T. Tsunagawa and T. Arai. Flexible migration of Japanese freshwater gobies Rhinogobius spp. as revealed by otolith Sr:Ca ratios. Journal of Fish Biology, 73(10):2421–2433, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dufour:2008:HRI

[3075] E. Dufour, T. O. Höök, W. P. Patterson, and E. S. Rutherford. High-resolution isotope analysis of young alewife Alosa pseudoharengus otoliths: assessment of temporal resolution and reconstruction of habitat occupancy and thermal history. Journal of Fish Biology, 73(10): 2434–2451, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baulier:2008:NSS

[3076] L. Baulier and M. Heino. Norwegian spring-spawning herring as the test case of piecewise linear regression method for detecting maturation from growth patterns. Journal of Fish Biology, 73(10):2452–2467, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Beaudouin:2008:GCE

[3077] R. Beaudouin, V. Ginot, and G. Monod. Growth characteristics of eastern mosquitofish Gambusia holbrooki in a northern habitat (Brittany, France). Journal of Fish Biology, 73(10):2468–2484, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Petursdottir:2008:LCF

[3078] H. Petursdottir, A. Gislason, and S. Falk-Petersen. Lipid classes and fatty acid compositions of muscle, liver and skull oil in deep-sea redfish Sebastes mentella over the reykjanes ridge. Journal of Fish Biology, 73(10):2485–2496, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shen:2008:RSR

[3079] K. N. Shen and W. N. Tzeng. Reproductive strategy and recruitment dynamics of amphidromous goby Sicyopterus japonicus as revealed by otolith microstructure. Journal of Fish Biology, 73(10):2497–2512, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wu:2008:CMM

[3080] C. L. Wu, C. W. Liu, H. W. Sun, H. C. Chang, C. J. Huang, C. F. Hui, and J. L. Wu. The carp M1 muscle-specific creatine kinase subisoform is adaptive to the synchronized changes in body temperature and intracellular pH that occur in the common carp Cyprinus carpio. Journal of Fish Biology, 73(10):2513–2526, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2008:TEO

[3081] F. C. Neat, P. J. Wright, and R. J. Fryer. Temperature effects on otolith pattern formation in Atlantic cod Gadus morhua. Journal of Fish Biology, 73(10):2527–2541, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Harrison:2008:GTC

[3082] T. D. Harrison and A. K. Whitfield. Geographical and typological changes in fish guilds of South African estuaries. Journal of Fish Biology, 73(10):2542–2570, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Uran:2008:SMI

[3083] P. A. Urán, R. Aydin, J. W. Schrama, J. A. J. Verreth, and J. H. W. M. Rombout. Soybean meal-induced uptake block in Atlantic salmon Salmo salar distal enterocytes. Journal of Fish Biology, 73(10):2571–2579, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brydges:2008:HDD

[3084] N. M. Brydges, A. M. Holmes, and V. A. Braithwaite. How do different populations of three-spined sticklebacks Gasterosteus aculeatus combine spatial information? Journal of Fish Biology, 73(10):2580–2586, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Clark:2008:CHH

[3085] T. D. Clark, E. J. Eliason, E. Sandblom, S. G. Hinch, and A. P. Farrell. Calibration of a hand-held haemoglobin analyser for use on fish blood. Journal of Fish Biology, 73(10):2587–2595, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:CV

[3086] Anonymous. Contents of volume 73. Journal of Fish Biology, 73 (10):xxxiv–xxxxiv, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2008:IV

[3087] Anonymous. Index, volume 73. Journal of Fish Biology, 73(10):i–xxxiii, December 2008. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Ea

[3088] Anonymous. Editorial. Journal of Fish Biology, 74(1):1, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Perga:2009:OCS

[3089] M.-E. Perga, A. Bec, and O. Anneville. Origins of carbon sustaining the growth of whitefish Coregonus lavaretus early larval stages in Lake Annecy: insights from fatty-acid biomarkers. Journal of Fish Biology, 74(1):2–17, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sisneros:2009:MCA

[3090] J. A. Sisneros, P. W. Alderks, K. Leon, and B. Sniffen. Morphometric changes associated with the reproductive cycle and behaviour of the intertidal-nesting, male plainfin midshipman Porichthys notatus. Journal of Fish Biology, 74(1):18–36, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Silva:2009:HHG

[3091] P. Silva, L. M. P. Valente, M. Olmedo, M. H. Galante, R. A. F. Monteiro, and E. Rocha. Hyperplastic and hypertrophic growth of lateral muscle in blackspot seabream Pagellus bogaraveo from hatching to juvenile. Journal of Fish Biology, 74(1):37–53, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moritz:2009:EOD

[3092] T. Moritz, J. Engelmann, K. E. Linsenmair, and G. Von DerEmde. The electric organ discharges of the Petrocephalus species (Teleostei: Mormyridae) of the Upper Volta System. Journal of Fish Biology, 74 (1):54–76, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carey:2009:EIT

[3093] G. R. Carey, P. G. Kraft, R. L. Cramp, and C. E. Franklin. Effect of incubation temperature on muscle growth of barramundi Lates calcarifer at hatch and post-exogenous feeding. Journal of Fish Biology, 74(1):77– 89, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Allen:2009:CSB

[3094] P. J. Allen, C. C. Barth, S. J. Peake, M. V. Abrahams, and W. G. Anderson. Cohesive social behaviour shortens the stress response: the effects of conspecifics on the stress response in lake sturgeon Acipenser fulvescens. Journal of Fish Biology, 74(1):90–104, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeBrabandere:2009:SIE

[3095] L. De Brabandere, M. J. Catalano, T. K. Frazer, and M. S. Allen. Stable isotope evidence of ontogenetic changes in the diet of gizzard shad Dorosoma cepedianum. Journal of Fish Biology, 74(1):105–119, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2009:EAS

[3096] C. T. Phillips and C. E. Johnston. Evolution of acoustic signals in Cyprinella: degree of similarity in sister species. Journal of Fish Biology, 74(1):120–132, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Forseth:2009:TGP

[3097] T. Forseth, S. Larsson, A. J. Jensen, B. Jonsson, I. Näslund, and I. Berglund. Thermal growth performance of juvenile brown trout Salmo trutta: no support for thermal adaptation hypotheses. Journal of Fish Biology, 74(1):133–149, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Probst:2009:ILD

[3098] W. N. Probst and R. Eckmann. The influence of light on the diel vertical migration of young-of-the-year burbot Lota lota in Lake Constance. Journal of Fish Biology, 74(1):150–166, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zahorska:2009:MVA

[3099] E. Záhorská, V. Kováč, I. Falka, K. Beyer, S. Katina, G. H. Copp, and R. E. Gozlan. Morphological variability of the Asiatic cyprinid, topmouth gudgeon Pseudorasbora parva, in its introduced European range. Journal of Fish Biology, 74(1):167–185, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2009:RSA

[3100] D. Li, W. Hu, Y. Wang, Z. Zhu, and C. Fu. Reduced swimming abilities in fast-growing transgenic common carp Cyprinus carpio associated with their morphological variations. Journal of Fish Biology, 74(1):186–197, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reichard:2009:DCP

[3101] M. Reichard, M. Polačik, and O. Sedláček. Distribution, colour polymorphism and habitat use of the African killifish Nothobranchius furzeri, the vertebrate with the shortest life span. Journal of Fish Biology, 74(1):198– 212, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hasegawa:2009:ECD

[3102] K. Hasegawa and S. Yamamoto. Effects of competitor density and physical habitat structure on the competitive intensity of territorial white spotted charr Salvelinus leucomaenis. Journal of Fish Biology, 74(1): 213–219, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McGrath:2009:ESA

[3103] S. P. McGrath, P. A. Butcher, and M. K. Broadhurst. Effects of salinity and anatomical hook location on the mortality and physiological response of angled-and-released sand whiting Sillago ciliata. Journal of Fish Biology, 74(1):220–234, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Betti:2009:LDA

[3104] P. Betti, L. Machinandiarena, and M. D. Ehrlich. Larval development of Argentine hake Merluccius hubbsi. Journal of Fish Biology, 74(1):235– 249, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoff:2009:SBS

[3105] G. R. Hoff. Skate Bathyraja spp. egg predation in the eastern Bering Sea. Journal of Fish Biology, 74(1):250–269, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kreiner:2009:SSS

[3106] A. Kreiner, E. K. Stenevik, and W. Ekau. Sardine Sardinops sagax and anchovy Engraulis encrasicolus larvae avoid regions with low dissolved oxygen concentration in the northern Benguela Current system. Journal of Fish Biology, 74(1):270–277, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aedo:2009:GVA

[3107] J. R. Aedo, M. C. Belk, and E. M. Habit. Geographic variation in age, growth and size structure of Percilia irwini from south-central Chile. Journal of Fish Biology, 74(1):278–284, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Park:2009:LTE

[3108] M. O. Park and I.-S. Park. Long-term effects of passive integrated transponder (PIT) tagging on the kelp grouper Epinephelus bruneus. Journal of Fish Biology, 74(1):285–288, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simon:2009:GME

[3109] J. Simon, H. Dörner, and C. Richter. Growth and mortality of European glass eel Anguilla anguilla marked with oxytetracycline and alizarin red. Journal of Fish Biology, 74(1):289–295, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fassler:2009:DDS

[3110] S. M. M. Fässler, P. G. Fernandes, S. I. K. Semple, and A. S. Brierley. Depth-dependent swimbladder compression in herring Clupea harengus observed using magnetic resonance imaging. Journal of Fish Biology, 74 (1):296–303, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Makiguchi:2009:EES

[3111] Y. Makiguchi and H. Ueda. Effects of external and surgically implanted dummy radio transmitters on mortality, swimming performance and physiological status of juvenile masu salmon Oncorhynchus masou. Journal of Fish Biology, 74(1):304–311, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tedesco:2009:VGR

[3112] P. A. Tedesco, P. Sagnes, and J. Laroche. Variability in the growth rate of chub Leuciscus cephalus along a longitudinal river gradient. Journal of Fish Biology, 74(1):312–319, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:R

[3113] Anonymous. Referees 2008. Journal of Fish Biology, 74(1):320–328, January 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ward:2009:CDB

[3114] R. D. Ward, R. Hanner, and P. D. N. Hebert. The campaign to DNA barcode all fishes, FISH-BOL. Journal of Fish Biology, 74(2):329–356, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tseng:2009:DBT

[3115] M. C. Tseng, C. T. Jean, W. L. Tsai, and N. C. Chen. Distinguishing between two sympatric Acanthopagrus species from Dapeng Bay, Taiwan, using morphometric and genetic characters. Journal of Fish Biology, 74 (2):357–376, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Valdez-Moreno:2009:PDF

[3116] M. Valdez-Moreno, N. V. Ivanova, M. Elías-Gutiérrez, S. Contreras-Balderas, and P. D. N. Hebert. Probing diversity in freshwater fishes from Mexico and Guatemala with DNA barcodes. Journal of Fish Biology, 74(2):377–402, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Purcell:2009:PSC

[3117] J. F. H. Purcell, R. K. Cowen, C. R. Hughes, and D. A. Williams. Population structure in a common Caribbean coral-reef fish: implications for larval dispersal and early life-history traits. Journal of Fish Biology, 74 (2):403–417, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lajbner:2009:RAH

[3118] Z. Lajbner, V. Šlechtová, V. Šlechta, M. Švátora, P. Berrebi, and P. Kotlík. Rare and asymmetrical hybridization of the endemic Barbus carpathicus with its widespread congener Barbus barbus. Journal of Fish Biology, 74(2):418–436, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phillips:2009:GSM

[3119] R. D. Phillips, A. W. Storey, and M. S. Johnson. Genetic structure of Melanotaenia australis at local and regional scales in the east Kimberley, Western Australia. Journal of Fish Biology, 74(2):437–451, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Babaryka:2009:VSM

[3120] A. Babaryka, E. Kühn, and R. W. Köster. In vivo synthesis of meganuclease for generating transgenic zebrafish Danio rerio. Journal of Fish Biology, 74(2):452–457, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nikolic:2009:SMD

[3121] N. Nikolic, K. Fève, C. Chevalet, B. Høyheim, and J. Riquet. A set of 37 microsatellite DNA markers for genetic diversity and structure analysis of Atlantic salmon Salmo salar populations. Journal of Fish Biology, 74 (2):458–466, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kitano:2009:EIH

[3122] S. Kitano, K. Hasegawa, and K. Maekawa. Evidence for interspecific hybridization between native white-spotted charr Salvelinus leucomaenis and non-native brown trout Salmo trutta on Hokkaido Island, Japan. Journal of Fish Biology, 74(2):467–473, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Manera:2009:RCB

[3123] M. Manera, L. Giari, and B. S. Dezfuli. Rodlet cell biometry: interspecific and intraspecific variability. Journal of Fish Biology, 74(2):474–481, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hoff:2009:EDE

[3124] G. R. Hoff. Embryo developmental events and the egg case of the Aleutian skate Bathyraja aleutica (Gilbert) and the Alaska skate Bathyraja parmifera (Bean). Journal of Fish Biology, 74(3):483–501, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maldonado:2009:MDR

[3125] E. Maldonado, N. Hubert, P. Sagnes, and B. De Mérona. Morphology– diet relationships in four killifishes (Teleostei, Cyprinodontidae, Orestias) from Lake Titicaca. Journal of Fish Biology, 74(3):502–520, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dubreuil:2009:EDA

[3126] J. Dubreuil and P. Petitgas. Energy density of anchovy Engraulis encrasicolus in the Bay of Biscay. Journal of Fish Biology, 74(3):521–534, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Campbell:2009:HRV

[3127] H. Campbell, W. Davison, K. P. P. Fraser, L. S. Peck, and S. Egginton. Heart rate and ventilation in Antarctic fishes are largely determined by ecotype. Journal of Fish Biology, 74(3):535–552, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seppanen:2009:MDS

[3128] E. Seppänen, H. Kuukka, A. Voutilainen, H. Huuskonen, and N. Peuhkuri. Metabolic depression and spleen and liver enlargement in juvenile Arctic charr Salvelinus alpinus exposed to chronic parasite infection. Journal of Fish Biology, 74(3):553–561, February 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seppanen:2009:MRG

[3129] E. Seppänen, K. Tiira, H. Huuskonen, and J. Piironen. Metabolic rate, growth and aggressiveness in three Atlantic salmon Salmo salar populations. Journal of Fish Biology, 74(3):562–575, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnes:2009:EFC

[3130] M. E. Barnes, D. Bergmann, J. Jacobs, and M. Gabel. Effect of Flavobacterium columnare inoculation, antibiotic treatments and resident bacteria on rainbow trout Oncorhynchus mykiss eyed egg survival and external membrane structure. Journal of Fish Biology, 74(3):576–590, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kieffer:2009:OCA

[3131] J. D. Kieffer and A. M. Wakefield. Oxygen consumption, ammonia excretion and protein use in response to thermal changes in juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 74(3):591–603, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chimura:2009:HPD

[3132] M. Chimura, Y. Watanabe, H. Okouchi, N. Shirafuji, and T. Kawamura. Hatch-period-dependent early growth and survival of Pacific herring Clupea pallasii in Miyako Bay, Japan. Journal of Fish Biology, 74 (3):604–620, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2009:MTM

[3133] B. Jonsson and N. Jonsson. Migratory timing, marine survival and growth of anadromous brown trout Salmo trutta in the River Imsa, Norway. Journal of Fish Biology, 74(3):621–638, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Middlemas:2009:HUD

[3134] S. J. Middlemas, D. C. Stewart, S. Mackay, and J. D. Armstrong. Habitat use and dispersal of post-smolt sea trout Salmo trutta in a Scottish sea loch system. Journal of Fish Biology, 74(3):639–651, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Teletchea:2009:ROD

[3135] F. Teletchea, J.-N. Gardeur, E. Kamler, and P. Fontaine. The relationship of oocyte diameter and incubation temperature to incubation time in temperate freshwater fish species. Journal of Fish Biology, 74 (3):652–668, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Goto:2009:RTD

[3136] T. Goto. Reversals in two dextral flounder species, Microstomus achne and Cleisthenes pinetorum (Pleuronectida; Teleostei), from Japan. Journal of Fish Biology, 74(3):669–673, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kieffer:2009:BPJ

[3137] J. D. Kieffer, L. M. Arsenault, and M. K. Litvak. Behaviour and performance of juvenile shortnose sturgeon Acipenser brevirostrum at different water velocities. Journal of Fish Biology, 74(3):674–682, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pander:2009:ESM

[3138] J. Pander, J. Schnell, K. Sternecker, and J. Geist. The ‘egg sandwich’: a method for linking spatially resolved salmonid hatching rates with habitat variables in stream ecosystems. Journal of Fish Biology, 74(3):683– 690, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lindberg:2009:PRP

[3139] M. Lindberg, P. Rivinoja, L.-O. Eriksson, and A. Alanärä. Post-release and pre-spawning behaviour of simulated escaped adult rainbow trout Oncorhynhus mykiss in Lake Övre Fryken, Sweden. Journal of Fish Biology, 74(3):691–698, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Morita:2009:LVE

[3140] K. Morita, T. Tamate, Y. Sugimoto, Y. Tago, T. Watanabe, H. Konaka, M. Sato, Y. Miyauchi, K. Ohkuma, and T. Nagasawa. Latitudinal variation in egg size and number in anadromous masu salmon Oncorhynchus masou. Journal of Fish Biology, 74(3):699–705, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brunnschweiler:2009:DDB

[3141] J. M. Brunnschweiler, H. Baensch, S. J. Pierce, and D. W. Sims. Deepdiving behaviour of a whale shark Rhincodon typus during long-distance movement in the western Indian Ocean. Journal of Fish Biology, 74 (3):706–714, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neat:2009:ETR

[3142] F. C. Neat, M. Breen, R. M. Cook, I. M. Gibb, and P. J. Wright. Electronic tags reveal behaviour of captured and discarded fish. Journal of Fish Biology, 74(3):715–721, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:BRa

[3143] Anonymous. Book reviews. Journal of Fish Biology, 74(3):722–723, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Eb

[3144] Anonymous. Erratum. Journal of Fish Biology, 74(3):724–725, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grutter:2009:E

[3145] A. S. Grutter. Erratum. Journal of Fish Biology, 74(3):726, February 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakamaki:2009:CEA

[3146] K. Sakamaki and Y. Satou. Caspases: evolutionary aspects of their functions in vertebrates. Journal of Fish Biology, 74(4):727–753, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Saraiva:2009:ENA

[3147] J. L. Saraiva, E. N. Barata, A. V. M. Canário, and R. F. Oliveira. The effect of nest aggregation on the reproductive behaviour of the peacock blenny Salaria pavo. Journal of Fish Biology, 74(4):754–762, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bass:2009:CBF

[3148] C. Santiago Bass and J. S. Weis. Conspicuous behaviour of Fundulus heteroclitus associated with high digenean metacercariae gill abundances. Journal of Fish Biology, 74(4):763–772, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dapra:2009:RPC

[3149] F. Daprà, F. Gai, M. T. Costanzo, G. Maricchiolo, V. Micale, B. Sicuro, G. Caruso, L. Genovese, and G. B. Palmegiano. Rice protein-concentrate meal as a potential dietary ingredient in practical diets for blackspot seabream Pagellus bogaraveo: a histological and enzymatic investigation. Journal of Fish Biology, 74(4):773–789, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Montero:2009:EDR

[3150] D. Montero, G. Lalumera, M. S. Izquierdo, M. J. Caballero, M. Saroglia, and L. Tort. Establishment of dominance relationships in gilthead sea bream Sparus aurata juveniles during feeding: effects on feeding behaviour, feed utilization and fish health. Journal of Fish Biology, 74 (4):790–805, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barrows:2009:PVL

[3151] A. P. W. Barrows, K. M. Martin-Smith, and M. S. P. Baine. Population variables and life-history characteristics of the alligator pipefish Syngnathoides biaculeatus, in Papua New Guinea. Journal of Fish Biology, 74(4):806–819, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grandcourt:2009:RBI

[3152] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, A. T. Al Shamsi, and S. A. Hartmann. Reproductive biology and implications for management of the orange-spotted grouper Epinephelus coioides in the southern Arabian Gulf. Journal of Fish Biology, 74(4):820–841, March 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aras:2009:SCA

[3153] N. M. Aras, A. Bayir, A. N. Sirkecioglu, M. Bayir, E. Aksakal, and H. I. Haliloglu. Seasonal changes in antioxidant defence system of liver and gills of Salmo trutta caspius, Salmo trutta labrax and Salmo trutta macrostigma. Journal of Fish Biology, 74(4):842–856, March 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yokouchi:2009:DDB

[3154] K. Yokouchi, J. Aoyama, M. J. Miller, T. K. McCarthy, and K. Tsukamoto. Depth distribution and biological characteristics of the European eel Anguilla anguilla in Lough Ennell, Ireland. Journal of Fish Biology, 74(4):857–871, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Madison:2009:ELC

[3155] B. N. Madison, R. S. Dhillon, B. L. Tufts, and Y. S. Wang. Exposure to low concentrations of dissolved ammonia promotes growth rate in walleye Sander vitreus. Journal of Fish Biology, 74(4):872–890, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Williamson:2009:TMM

[3156] D. H. Williamson, G. P. Jones, S. R. Thorrold, and A. J. Frisch. Transgenerational marking of marine fish larvae: stable-isotope retention, physiological effects and health issues. Journal of Fish Biology, 74(4): 891–905, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dunbrack:2009:CGE

[3157] R. Dunbrack, L. A. Giguère, and J.-F. St-Pierre. A comparison of gut evacuation models for larval mackerel (Scomber scombrus) using serial photography. Journal of Fish Biology, 74(4):906–920, March 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Durieux:2009:TSM

[3158] E. D. H. Durieux, M. G. Meekan, D. Ponton, and L. Vigliola. Temperature, selective mortality and early growth in the short-lived clupeid Spratelloides gracilis. Journal of Fish Biology, 74(4):921–938, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schludermann:2009:EIS

[3159] E. Schludermann, H. Keckeis, and H.-L. Nemeschkal. Effect of initial size on daily growth and survival in freshwater Chondrostoma nasus larvae: a field survey. Journal of Fish Biology, 74(4):939–955, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duncan:2009:IRK

[3160] W. P. Duncan, O. T. F. Costa, M. L. G. Araújo, and M. N. Fernandes. Ionic regulation and Na^{+}-K^{+}-ATPase activity in gills and kidney of the freshwater stingray Paratrygon aiereba living in white and blackwaters in the Amazon Basin. Journal of Fish Biology, 74(4):956–960, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Achleitner:2009:GWF

[3161] D. Achleitner, H. Gassner, and R. Schabetsberger. ‘Global worming’: first record of an epidemic of Triaenophorus crassus in a population of Arctic charr Salvelinus umbla. Journal of Fish Biology, 74(4):961– 966, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vinni:2009:SGP

[3162] M. Vinni, J. Lappalainen, T. Malinen, and H. Lehtonen. Stunted growth of pikeperch Sander lucioperca in Lake Sahajärvi, Finland. Journal of Fish Biology, 74(4):967–972, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:BRb

[3163] Anonymous. Book reviews. Journal of Fish Biology, 74(4):973–976, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Ec

[3164] Anonymous. Erratum. Journal of Fish Biology, 74(4):977–978, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kreiner:2009:E

[3165] A. Kreiner, E. K. Stenevik, and W. Ekau. Erratum. Journal of Fish Biology, 74(4):979–981, March 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Grimes:2009:OTH

[3166] A. C. Grimes and M. L. Kirby. The outflow tract of the heart in fishes: anatomy, genes and evolution. Journal of Fish Biology, 74(5):983–1036, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2009:LVS

[3167] C. Li, M. L. Bessert, J. Macrander, and G. Ortí. Low variation but strong population structure in mitochondrial control region of the plains topminnow, Fundulus sciadicus. Journal of Fish Biology, 74(5):1037– 1048, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernardi:2009:NFC

[3168] G. Bernardi. The name of the father: conflict between Louis and Alexander Agassiz and the Embiotoca surfperch radiation. Journal of Fish Biology, 74(5):1049–1055, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wares:2009:EDT

[3169] J. P. Wares. Evolutionary dynamics of transferrin in Notropis. Journal of Fish Biology, 74(5):1056–1069, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caudron:2009:ETC

[3170] A. Caudron, A. Champigneulle, and R. Guyomard. Evidence of two contrasting brown trout Salmo trutta populations spatially separated in the River Borne (France) and shift in management towards conservation of the native lineage. Journal of Fish Biology, 74(5):1070–1085, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franckowiak:2009:TES

[3171] R. P. Franckowiak, B. L. Sloss, M. A. Bozek, and S. P. Newman. Temporal effective size estimates of a managed walleye Sander vitreus population and implications for genetic-based management. Journal of Fish Biology, 74(5):1086–1103, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2009:SGM

[3172] P. J. Smith, C. D. Struthers, C. D. Paulin, S. M. McVeagh, and R. K. Daley. Shallow genetic and morphological divergence among seaperches in the South Pacific (family Scorpaenidae; genus Helicolenus). Journal of Fish Biology, 74(5):1104–1128, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stiver:2009:MPN

[3173] K. A. Stiver, J. L. Fitzpatrick, J. K. Desjardins, and S. Balshine. Mixed parentage in Neolamprologus pulcher groups. Journal of Fish Biology, 74(5):1129–1135, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Griffiths:2009:CMC

[3174] A. M. Griffiths, D. Bright, and J. R. Stevens. Complete mitochondrial control region sequences indicate a distinct variety of brown trout Salmo trutta in the Aral Sea. Journal of Fish Biology, 74(5):1136–1142, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Graham:2009:ICC

[3175] C. T. Graham and C. Harrod. Implications of climate change for the fishes of the British Isles. Journal of Fish Biology, 74(6):1143–1205, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Timi:2009:SSS

[3176] J. T. Timi, A. L. Lanfranchi, and J. A. Etchegoin. Seasonal stability and spatial variability of parasites in Brazilian sandperch Pinguipes brasilianus from the Northern Argentine Sea: evidence for stock discrimination. Journal of Fish Biology, 74(6):1206–1225, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Speranza:2009:BCD

[3177] E. D. Speranza and J. C. Colombo. Biochemical composition of a dominant detritivorous fish Prochilodus lineatus along pollution gradients in the Paraná–Río de la Plata Basin. Journal of Fish Biology, 74(6):1226– 1244, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Schofield:2009:SEB

[3178] P. J. Schofield, W. F. Loftus, and J. A. Fontaine. Salinity effects on behavioural response to hypoxia in the non-native Mayan cichlid Cichlasoma urophthalmus from Florida Everglades wetlands. Journal of Fish Biology, 74(6):1245–1258, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gregersen:2009:SSV

[3179] F. Gregersen, L. A. Vøllestad, E. M. Olsen, and T. O. Haugen. Siblingsize variation in brown trout Salmo trutta in relation to egg size and stream size. Journal of Fish Biology, 74(6):1259–1268, April 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuliskova:2009:FIM

[3180] P. Kulíšková, P. Horký, O. Slavík, and J. I. Jones. Factors influencing movement behaviour and home range size in ide Leuciscus idus. Journal of Fish Biology, 74(6):1269–1279, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kim:2009:CEC

[3181] J.-W. Kim, G. E. Brown, I. J. Dolinsek, N. N. Brodeur, A. O. H. C. Leduc, and J. W. A. Grant. Combined effects of chemical and visual information in eliciting antipredator behaviour in juvenile Atlantic salmon Salmo salar. Journal of Fish Biology, 74(6):1280–1290, April 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pierce:2009:RBS

[3182] S. J. Pierce, S. A. Pardo, and M. B. Bennett. Reproduction of the blue-spotted maskray Neotrygon kuhlii (Myliobatoidei: Dasyatidae) in south-east Queensland, Australia. Journal of Fish Biology, 74(6):1291– 1308, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koumoundouros:2009:TIP

[3183] G. Koumoundouros, C. Ashton, D. G. Sfakianakis, P. Divanach, M. Kentouri, N. Anthwal, and N. C. Stickland. Thermally induced phenotypic plasticity of swimming performance in European sea bass Dicentrarchus labrax juveniles. Journal of Fish Biology, 74(6):1309–1322, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pon:2009:PEB

[3184] L. B. Pon, S. G. Hinch, S. J. Cooke, D. A. Patterson, and A. P. Farrell. Physiological, energetic and behavioural correlates of successful fishway passage of adult sockeye salmon Oncorhynchus nerka in the Seton River, British Columbia. Journal of Fish Biology, 74(6):1323–1336, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Aboim:2009:RGR

[3185] M. A. Aboim, C. Cunha, and M. M. Coelho. Redistribution of the geographical ranges of the Iberian cyprinid genus Pseudochondrostoma based on a phylogenetic analysis: implications for the historical rearrangements of the north-western Iberian drainages. Journal of Fish Biology, 74(6):1337–1346, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Stewart:2009:SBA

[3186] D. C. Stewart, S. J. Middlemas, S. Mackay, and J. D. Armstrong. Oversummering behaviour of Atlantic salmon Salmo salar returning to rivers in the Cromarty Firth, north-east Scotland. Journal of Fish Biology, 74 (6):1347–1352, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Ed

[3187] Anonymous. Erratum. Journal of Fish Biology, 74(6):1353, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Seppanen:2009:E

[3188] E. Seppänen, H. Kuukka, A. Voutilainen, H. Huuskonen, and N. Peuhkuri. Erratum. Journal of Fish Biology, 74(6):1354, April 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hansen:2009:ASA

[3189] S. L. Hansen and I. Hunt Von Herbing. Aerobic scope for activity in age 0 year Atlantic cod Gadus morhua. Journal of Fish Biology, 74(7):1355– 1370, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dupuch:2009:DPR

[3190] A. Dupuch, P. Magnan, A. Bertolo, L. M. Dill, and M. Proulx. Does predation risk influence habitat use by northern redbelly dace Phoxinus eos at different spatial scales? Journal of Fish Biology, 74(7):1371– 1382, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Berg:2009:PWL

[3191] O. K. Berg, A. G. Finstad, Ø. Solem, O. Ugedal, T. Forseth, E. Niemelä, J. V. Arnekleiv, A. Lohrmann, and T. F. Næsje. Pre-winter lipid stores in young-of-year Atlantic salmon along a north–south gradient. Journal of Fish Biology, 74(7):1383–1393, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sullivan:2009:SFC

[3192] S. M. P. Sullivan and M. C. Watzin. Stream–floodplain connectivity and fish assemblage diversity in the Champlain Valley, Vermont, U.S.A. Journal of Fish Biology, 74(7):1394–1418, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bamberger:2009:ENS

[3193] A. Bamberger. Evaluation of a new semi-natural incubation technique for Atlantic salmon Salmo salar fry. Journal of Fish Biology, 74(7):1419– 1433, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weng:2009:HBY

[3194] K. C. Weng, M. J. W. Stokesbury, A. M. Boustany, A. C. Seitz, S. L. H. Teo, S. K. Miller, and B. A. Block. Habitat and behaviour of yellowfin tuna Thunnus albacares in the Gulf of Mexico determined using pop-up satellite archival tags. Journal of Fish Biology, 74(7):1434–1449, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:2009:SSC

[3195] T. E. Pitcher, S. M. Doucet, J.-M. J. Beausoleil, and D. Hanley. Secondary sexual characters and sperm traits in Coho salmon Oncorhynchus kisutch. Journal of Fish Biology, 74(7):1450–1461, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Immanuel:2009:DMP

[3196] G. Immanuel, R. P. Uma, P. Iyapparaj, T. Citarasu, S. M. Punitha Peter, M. Michael Babu, and A. Palavesam. Dietary medicinal plant extracts improve growth, immune activity and survival of tilapia Oreochromis mossambicus. Journal of Fish Biology, 74(7):1462–1475, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winger:2009:ETI

[3197] A. C. Winger, R. Kristoffersen, S. I. Siikavuopio, and R. Knudsen. Experiments to test if allopatric Salvelinus alpinus are suitable year-round hosts of Gyrodactylus salaris (Monogenea). Journal of Fish Biology, 74 (7):1476–1486, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Das:2009:BSR

[3198] T. Das, A. K. Pal, S. K. Chakraborty, S. M. Manush, R. S. Dalvi, S. K. Apte, N. P. Sahu, and K. Baruah. Biochemical and stress responses of rohu Labeo rohita and mrigal Cirrhinus mrigala in relation to acclimation temperatures. Journal of Fish Biology, 74(7):1487–1498, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2009:WBF

[3199] P. J. Miller. A west Balkanian freshwater gobiid fish, Knipowitschia mrakovcici sp. nov. (Teleostei: Gobiidae). Journal of Fish Biology, 74 (7):1499–1507, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wuenschel:2009:SPW

[3200] M. J. Wuenschel, K. W. Able, and D. Byrne. Seasonal patterns of winter flounder Pseudopleuronectes americanus abundance and reproductive condition on the New York Bight continental shelf. Journal of Fish Biology, 74(7):1508–1524, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pursche:2009:TII

[3201] A. R. Pursche, I. M. Suthers, and M. D. Taylor. Tethering induces increased stress artifacts in social fish species. Journal of Fish Biology, 74(7):1525–1531, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hvidsten:2009:IST

[3202] N. A. Hvidsten, A. J. Jensen, A. H. Rikardsen, B. Finstad, J. Aure, S. Stefansson, P. Fiske, and B. O. Johnsen. Influence of sea temperature and initial marine feeding on survival of Atlantic salmon Salmo salar post-smolts from the Rivers Orkla and Hals, Norway. Journal of Fish Biology, 74(7):1532–1548, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Leong:2009:EPS

[3203] H. Leong, A. F. H. Ros, and R. F. Oliveira. Effects of putative stressors in public aquaria on locomotor activity, metabolic rate and cortisol levels in the Mozambique tilapia Oreochromis mossambicus. Journal of Fish Biology, 74(7):1549–1561, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiffan:2009:BTS

[3204] K. F. Tiffan, T. J. Kock, W. P. Connor, R. K. Steinhorst, and D. W. Rondorf. Behavioural thermoregulation by subyearling fall (autumn) Chinook salmon Oncorhynchus tshawytscha in a reservoir. Journal of Fish Biology, 74(7):1562–1579, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sammons:2009:ERF

[3205] S. M. Sammons and M. J. Maceina. Effects of river flows on growth of redbreast sunfish Lepomis auritus (Centrarchidae) in Georgia rivers. Journal of Fish Biology, 74(7):1580–1593, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bittner:2009:HNA

[3206] D. Bittner, D. Bernet, T. Wahli, H. Segner, C. Küng, and C. R. Largiadèr. How normal is abnormal? Discrimination between deformations and natural variation in gonad morphology of European whitefish Coregonus lavaretus. Journal of Fish Biology, 74(7):1594–1614, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dalyan:2009:AST

[3207] C. Dalyan and L. Eryilmaz. The Arabian scad Trachurus indicus, a new Indo–Pacific species in the Mediterranean Sea. Journal of Fish Biology, 74(7):1615–1619, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lopez-Vasquez:2009:DEE

[3208] K. López-Vásquez, C. A. Castro-Pérez, and A. L. Val. Digestive enzymes of eight Amazonian teleosts with different feeding habits. Journal of Fish Biology, 74(7):1620–1628, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Moore:2009:MPA

[3209] A. Moore and W. D. Riley. Magnetic particles associated with the lateral line of the European eel Anguilla anguilla. Journal of Fish Biology, 74 (7):1629–1634, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heflin:2009:STC

[3210] B. Heflin, L. Young, and R. L. Londraville. Short-term cycling of skin colouration in the blackspotted rockskipper Entomacrodus striatus. Journal of Fish Biology, 74(7):1635–1641, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Konecna:2009:RDD

[3211] M. Konečná, P. Jurajda, and M. Reichard. River discharge drives recruitment success of the European bitterling Rhodeus amarus in a regulated river in central Europe. Journal of Fish Biology, 74(7):1642–1650, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Johnston:2009:DFP

[3212] P. Johnston, F. Bérubé, and N. E. Bergeron. Development of a flatbed passive integrated transponder antenna grid for continuous monitoring of fishes in natural streams. Journal of Fish Biology, 74(7):1651–1661, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fischer:2009:DTT

[3213] C. Fischer and I. Schlupp. Differences in thermal tolerance in coexisting sexual and asexual mollies (Poecilia, Poeciliidae, Teleostei). Journal of Fish Biology, 74(7):1662–1668, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vetesnik:2009:FRN

[3214] L. Vetešník, K. Halačka, I. Papoušek, J. Mendel, and A. Šimková. The first record of a natural hybrid of the roach Rutilus rutilus and nase Chondrostoma nasus in the Danube River Basin, Czech Republic: morphological, karyological and molecular characteristics. Journal of Fish Biology, 74(7):1669–1676, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Follesa:2009:FRC

[3215] M. C. Follesa, A. Mulas, C. Porcu, and A. Cau. First record of Chilomycterus reticulatus (Osteichthyes: Diodontidae) in the Mediterranean Sea. Journal of Fish Biology, 74(7):1677–1681, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jorgensen:2009:SHS

[3216] S. J. Jorgensen, A. P. Klimley, and A. F. Muhlia-Melo. Scalloped hammerhead shark Sphyrna lewini, utilizes deep-water, hypoxic zone in the Gulf of California. Journal of Fish Biology, 74(7):1682–1687, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

He:2009:CSD

[3217] X. He and S. Yang. Comments on ‘Significance of developmental expression of amphioxus Branchiostoma belcheri and zebrafish Danio rerio Hsd17b10 in biological and medical research’. Journal of Fish Biology, 74(8):1689–1692, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic). See [2859].

Chong:2009:MMI

[3218] S.-W. Chong, V. Korzh, and Y.-J. Jiang. Myogenesis and molecules — insights from zebrafish Danio rerio. Journal of Fish Biology, 74(8):1693– 1755, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sakai:2009:MMS

[3219] H. Sakai, K. Iguchi, Y. Yamazaki, V. G. Sideleva, and A. Goto. Morphological and mtDNA sequence studies on three crucian carps (Carassius:

Cyprinidae) including a new stock from the Ob River system, Kazakhstan. Journal of Fish Biology, 74(8):1756–1773, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2009:EFS

[3220] S. Y. Chen, R. D. Zhang, J. G. Feng, H. Xiao, W. X. Li, R. G. Zan, and Y. P. Zhang. Exploring factors shaping population genetic structure of the freshwater fish Sinocyclocheilus grahami (Teleostei, Cyprinidae). Journal of Fish Biology, 74(8):1774–1786, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Li:2009:CCM

[3221] S. F. Li, J. W. Xu, Q. L. Yang, C. H. Wang, Q. Chen, D. C. Chapman, and G. Lu. A comparison of complete mitochondrial genomes of silver carp Hypophthalmichthys molitrix and bighead carp Hypophthalmichthys nobilis: implications for their taxonomic relationship and phylogeny. Journal of Fish Biology, 74(8):1787–1803, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Walter:2009:NBG

[3222] R. P. Walter, G. D. Haffner, and D. D. Heath. No barriers to gene flow among sympatric polychromatic ‘small’ Telmatherina antoniae from Lake Matano, Indonesia. Journal of Fish Biology, 74(8):1804–1815, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Karaiskou:2009:MVW

[3223] N. Karaiskou, A. Triantafyllidis, V. Katsares, T. J. Abatzopoulos, and C. Triantaphyllidis. Microsatellite variability of wild and farmed populations of Sparus aurata. Journal of Fish Biology, 74(8):1816–1825, May 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:AE

[3224] Anonymous. Anguillid eels. Journal of Fish Biology, 74(9):1827–1832, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tsukamoto:2009:OMS

[3225] K. Tsukamoto. Oceanic migration and spawning of anguillid eels. Journal of Fish Biology, 74(9):1833–1852, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuroki:2009:SSA

[3226] M. Kuroki, J. Aoyama, M. J. Miller, T. Yoshinaga, A. Shinoda, S. Hagihara, and K. Tsukamoto. Sympatric spawning of Anguilla marmorata and Anguilla japonica in the western North Pacific Ocean. Journal of Fish Biology, 74(9):1853–1865, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reveillac:2009:CLS

[3227] É. Réveillac, T. Robinet, M.-W. Rabenevanana, P. Valade, and É. Feunteun. Clues to the location of the spawning area and larval migration characteristics of Anguilla mossambica as inferred from otolith microstructural analyses. Journal of Fish Biology, 74(9):1866–1877, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zenimoto:2009:ESI

[3228] K. Zenimoto, T. Kitagawa, S. Miyazaki, Y. Sasai, H. Sasaki, and S. Kimura. The effects of seasonal and interannual variability of oceanic structure in the western Pacific North Equatorial Current on larval transport of the Japanese eel Anguilla japonica. Journal of Fish Biology, 74 (9):1878–1890, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bonhommeau:2009:EMD

[3229] S. Bonhommeau, O. Le Pape, D. Gascuel, B. Blanke, A.-M. Tréguier, N. Grima, Y. Vermard, M. Castonguay, and E. Rivot. Estimates of the mortality and the duration of the trans-Atlantic migration of European eel Anguilla anguilla leptocephali using a particle tracking model. Journal of Fish Biology, 74(9):1891–1914, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fukuda:2009:IWT

[3230] N. Fukuda, M. Kuroki, A. Shinoda, Y. Yamada, A. Okamura, J. Aoyama, and K. Tsukamoto. Influence of water temperature and feeding regime on otolith growth in Anguilla japonica glass eels and elvers: does otolith growth cease at low temperatures? Journal of Fish Biology, 74(9):1915– 1933, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cote:2009:DEO

[3231] C. L. Côté, M. Castonguay, G. Verreault, and L. Bernatchez. Differential effects of origin and salinity rearing conditions on growth of glass eels of the American eel Anguilla rostrata: implications for stocking programmes. Journal of Fish Biology, 74(9):1934–1948, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sullivan:2009:IIE

[3232] M. C. Sullivan, M. J. Wuenschel, and K. W. Able. Inter and intra-estuary variability in ingress, condition and settlement of the American eel Anguilla rostrata: implications for estimating and understanding recruitment. Journal of Fish Biology, 74(9):1949–1969, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dutil:2009:ARG

[3233] J.-D. Dutil, P. Dumont, D. K. Cairns, P. S. Galbraith, G. Verreault, M. Castonguay, and S. Proulx. Anguilla rostrata glass eel migration and recruitment in the estuary and Gulf of St Lawrence. Journal of Fish Biology, 74(9):1970–1984, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Acou:2009:EDY

[3234] A. Acou, A. Legault, P. Laffaille, and E. Feunteun. Environmental determinism of year-to-year recruitment variability of European eel Anguilla anguilla in a small coastal catchment, the Frémur River, north-west France. Journal of Fish Biology, 74(9):1985–2001, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

BureauDuColombier:2009:SAB

[3235] S. Bureau Du Colombier, V. Bolliet, and A. Bardonnet. Swimming activity and behaviour of European Anguilla anguilla glass eels in response to photoperiod and flow reversal and the role of energy status. Journal of Fish Biology, 74(9):2002–2013, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jellyman:2009:RAS

[3236] D. J. Jellyman, D. J. Booker, and E. Watene. Recruitment of Anguilla spp. glass eels in the Waikato River, New Zealand. Evidence of declining migrations? Journal of Fish Biology, 74(9):2014–2033, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pujolar:2009:GCA

[3237] J. M. Pujolar, G. A. De Leo, E. Ciccotti, and L. Zane. Genetic composition of Atlantic and Mediterranean recruits of European eel Anguilla anguilla based on EST-linked microsatellite loci. Journal of Fish Biology, 74(9):2034–2046, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Maes:2009:MGS

[3238] G. E. Maes, B. Van Vo, A. J. Crivelli, and F. A. M. Volckaert. Morphological and genetic seasonal dynamics of European eel Anguilla anguilla recruitment in southern France. Journal of Fish Biology, 74(9):2047– 2068, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Watanabe:2009:MME

[3239] S. Watanabe, M. J. Miller, J. Aoyama, and K. Tsukamoto. Morphological and meristic evaluation of the population structure of Anguilla marmorata across its range. Journal of Fish Biology, 74(9):2069–2093, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cairns:2009:SLG

[3240] D. K. Cairns, D. A. Secor, W. E. Morrison, and J. A. Hallett. Salinitylinked growth in anguillid eels and the paradox of temperate-zone catadromy. Journal of Fish Biology, 74(9):2094–2114, June 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dorner:2009:PTP

[3241] H. Dörner, C. Skov, S. Berg, T. Schulze, D. J. Beare, and G. Van der Velde. Piscivory and trophic position of Anguilla anguilla in two lakes: importance of macrozoobenthos density. Journal of Fish Biology, 74(9): 2115–2131, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carpentier:2009:ECF

[3242] A. Carpentier, L. Marion, J.-M. Paillisson, A. Acou, and E. Feunteun. Effects of commercial fishing and predation by cormorants on the Anguilla anguilla stock of a shallow eutrophic lake. Journal of Fish Biology, 74(9):2132–2138, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Breukelaar:2009:RCM

[3243] A. W. Breukelaar, D. Ingendahl, F. T. Vriese, G. De Laak, S. Staas, and J. G. P. Klein Breteler. Route choices, migration speeds and daily migration activity of European silver eels Anguilla anguilla in the River Rhine, north-west Europe. Journal of Fish Biology, 74(9):2139–2157, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sjoberg:2009:ESP

[3244] N. B. Sjöberg, E. Petersson, H. Wickström, and S. Hansson. Effects of the swimbladder parasite Anguillicola crassus on the migration of European silver eels Anguilla anguilla in the Baltic Sea. Journal of Fish Biology, 74(9):2158–2170, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reveillac:2009:DKU

[3245] É. Réveillac, P.-A. Gagnaire, R. Lecomte-Finiger, P. Berrebi, T. Robinet, P. Valade, and É. Feunteun. Development of a key using morphological characters to distinguish south-western Indian Ocean anguillid glass eels. Journal of Fish Biology, 74(9):2171–2177, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bevacqua:2009:SSF

[3246] D. Bevacqua, G. A. De Leo, M. Gatto, and P. Melià. Size selectivity of fyke nets for European eel Anguilla anguilla. Journal of Fish Biology, 74(9):2178–2186, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kydd:2009:LSP

[3247] E. Kydd and C. Brown. Loss of shoaling preference for familiar individuals in captive-reared crimson spotted rainbowfish Melanotaenia duboulayi. Journal of Fish Biology, 74(10):2187–2195, June 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaspersson:2009:DDG

[3248] R. Kaspersson and J. Höjesjö. Density-dependent growth rate in an age-structured population: a field study on stream-dwelling brown trout Salmo trutta. Journal of Fish Biology, 74(10):2196–2215, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haxton:2009:VLB

[3249] T. J. Haxton and C. S. Findlay. Variation in large-bodied fish-community structure and abundance in relation to water-management regime in a large regulated river. Journal of Fish Biology, 74(10):2216–2238, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Strunjak-Perovic:2009:NAM

[3250] I. Strunjak-Perovic, N. Topic Popovic, R. Coz-Rakovac, and M. Jadan. Nuclear abnormalities of marine fish erythrocytes. Journal of Fish Biology, 74(10):2239–2249, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yasue:2009:SVG

[3251] N. Yasue and A. Takasuka. Seasonal variability in growth of larval Japanese anchovy Engraulis japonicus driven by fluctuations in sea temperature in the Kii Channel, Japan. Journal of Fish Biology, 74(10): 2250–2268, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rahman:2009:EFT

[3252] M. M. Rahman and C. G. Meyer. Effects of food type on diel behaviours of common carp Cyprinus carpio in simulated aquaculture pond conditions. Journal of Fish Biology, 74(10):2269–2278, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paxton:2009:IFY

[3253] C. G. M. Paxton, I. J. Winfield, J. M. Fletcher, D. G. George, and D. P. Hewitt. Investigation of first year biotic and abiotic influences on the recruitment of pike Esox lucius over 48 years in Windermere, U.K. Journal of Fish Biology, 74(10):2279–2298, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown-Peterson:2009:IIW

[3254] N. J. Brown-Peterson and D. C. Heins. Interspawning interval of wild female three-spined stickleback Gasterosteus aculeatus in Alaska. Journal of Fish Biology, 74(10):2299–2312, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bolland:2009:DSS

[3255] J. D. Bolland, I. G. Cowx, and M. C. Lucas. Dispersal and survival of stocked cyprinids in a small English river: comparison with wild fishes using a multi-method approach. Journal of Fish Biology, 74(10):2313– 2328, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Magellan:2009:ESE

[3256] K. Magellan and A. E. Magurran. The effect of social environment during ontogeny on life history expression in the guppy Poecilia reticulata.

Journal of Fish Biology, 74(10):2329–2337, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yoda:2009:ARP

[3257] M. Yoda and M. Yoneda. Assessment of reproductive potential in multiple-spawning fish with indeterminate fecundity: a case study of yellow sea bream Dentex hypselosomus in the East China Sea. Journal of Fish Biology, 74(10):2338–2354, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parra:2009:LAG

[3258] I. Parra, A. Almodóvar, G. G. Nicola, and B. Elvira. Latitudinal and altitudinal growth patterns of brown trout Salmo trutta at different spatial scales. Journal of Fish Biology, 74(10):2355–2373, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tiira:2009:STD

[3259] K. Tiira, A. Laurila, K. Enberg, and J. Piironen. Short-term dominance: stability and consequences for subsequent growth. Journal of Fish Biology, 74(10):2374–2385, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Doupe:2009:IPS

[3260] R. G. Doupé, M. J. Knott, J. Schaffer, and D. W. Burrows. Investigational piscivory of some juvenile Australian freshwater fishes by the introduced Mozambique tilapia Oreochromis mossambicus. Journal of Fish Biology, 74(10):2386–2400, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kuparinen:2009:LPT

[3261] A. Kuparinen, R. B. O’Hara, and J. Merilä. Lunar periodicity and the timing of river entry in Atlantic salmon Salmo salar. Journal of Fish Biology, 74(10):2401–2408, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shephard:2009:DIG

[3262] S. Shephard and D. C. Jackson. Density-independent growth of floodplain river channel catfish Ictalurus punctatus. Journal of Fish Biology, 74(10):2409–2414, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feeley:2009:CMP

[3263] M. W. Feeley, O. J. Luiz, Jr., and N. Zurcher. Colour morph of a probable queen angelfish Holacanthus ciliaris from Dry Tortugas, Florida. Journal of Fish Biology, 74(10):2415–2421, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:CV

[3264] Anonymous. Contents of volume 74. Journal of Fish Biology, 74(10): xxxii–xxxxiv, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:IVa

[3265] Anonymous. Index, volume 74. Journal of Fish Biology, 74(10):i–xxxi, June 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sopinka:2009:LSR

[3266] N. M. Sopinka, J. L. Fitzpatrick, J. K. Desjardins, K. A. Stiver, S. E. Marsh-Rollo, and S. Balshine. Liver size reveals social status in the African cichlid Neolamprologus pulcher. Journal of Fish Biology, 75(1): 1–16, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Martin:2009:ESS

[3267] K. L. Martin, C. L. Moravek, and J. A. Flannery. Embryonic staging series for the beach spawning, terrestrially incubating California grunion Leuresthes tenuis with comparisons to other Atherinomorpha. Journal of Fish Biology, 75(1):17–38, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Miller:2009:ETW

[3268] J. A. Miller. The effects of temperature and water concentration on the otolith incorporation of barium and manganese in black rockfish Sebastes melanops. Journal of Fish Biology, 75(1):39–60, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oliveira:2009:LDS

[3269] C. Oliveira, M. T. Dinis, F. Soares, E. Cabrita, P. Pousão-Ferreira, and F. J. Sánchez-Vázquez. Lunar and daily spawning rhythms of Senegal sole Solea senegalensis. Journal of Fish Biology, 75(1):61–74, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cantanhede:2009:VPS

[3270] G. Cantanhêde, R. Fugi, and N. S. Hahn. Variation in prey selection of a piscivorous fish after the impoundment of a neotropical reservoir: prey size and type. Journal of Fish Biology, 75(1):75–86, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chang:2009:SPF

[3271] C.-H. Chang, C.-C. Chiao, and H. Y. Yan. The structure and possible functions of the milkfish Chanos chanos adipose eyelid. Journal of Fish Biology, 75(1):87–99, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2009:MGJ

[3272] Y. J. Lin and W. N. Tzeng. Modelling the growth of Japanese eel Anguilla japonica in the lower reach of the Kao-Ping River, southern Taiwan: an information theory approach. Journal of Fish Biology, 75(1):100– 112, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Koehn:2009:MSH

[3273] J. D. Koehn. Multi-scale habitat selection by Murray cod Maccullochella peelii peelii in two lowland rivers. Journal of Fish Biology, 75(1):113– 129, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cruz:2009:IPG

[3274] E. M. Vera Cruz and C. L. Brown. Influence of the photoperiod on growth rate and insulin-like growth factor-I gene expression in Nile tilapia Oreochromis niloticus. Journal of Fish Biology, 75(1):130–141, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kalamarz:2009:MCD

[3275] H. Kalamarz, M. Nietrzeba, J. Fuentes, G. Martinez-Rodriguez, J. M. Mancera, and E. Kulczykowska. Melatonin concentrations during larval and postlarval development of gilthead sea bream Sparus auratus: more than a time-keeping molecule? Journal of Fish Biology, 75(1):142–155, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zischke:2009:SSB

[3276] M. T. Zischke, T. H. Cribb, D. J. Welch, W. Sawynok, and R. J. G. Lester. Stock structure of blue threadfin Eleutheronema tetradactylum on the Queensland east coast, as determined by parasites and conventional tagging. Journal of Fish Biology, 75(1):156–171, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomiyama:2009:RFI

[3277] T. Tomiyama, M. Watanabe, and Y. Kurita. Rapid fluctuation in infection levels of Neoheterobothrium hirame (Monogenea) in Japanese flounder Paralichthys olivaceus in the Joban area, Japan. Journal of Fish Biology, 75(1):172–185, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kooka:2009:WLD

[3278] K. Kooka, O. Yamamura, N. Ohkubo, and S. Honda. Winter lipid depletion of juvenile walleye pollock Theragra chalcogramma in the Doto area, northern Japan. Journal of Fish Biology, 75(1):186–202, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zydlewski:2009:PMR

[3279] G. B. Zydlewski, J. Zydlewski, and J. Johnson. Patterns of migration and residency in coastal cutthroat trout Oncorhynchus clarkii clarkii from two tributaries of the lower Columbia River. Journal of Fish Biology, 75(1):203–222, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mcphie:2009:RCP

[3280] R. P. Mcphie and S. E. Campana. Reproductive characteristics and population decline of four species of skate (Rajidae) off the eastern coast of Canada. Journal of Fish Biology, 75(1):223–246, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Specziar:2009:FGF

[3281] A. Specziár and E. T. Rezsu. Feeding guilds and food resource partitioning in a lake fish assemblage: an ontogenetic approach. Journal of Fish Biology, 75(1):247–267, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baker:2009:VIP

[3282] D. W. Baker, C. May, and C. J. Brauner. A validation of intracellular pH measurements in fish exposed to hypercarbia: the effect of duration of tissue storage and efficacy of the metabolic inhibitor tissue homogenate method. Journal of Fish Biology, 75(1):268–275, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reeve:2009:PSF

[3283] A. Reeve, R. D. Handy, and S. H. Gruber. Prey selection and functional response of juvenile lemon sharks Negaprion brevirostris. Journal of Fish Biology, 75(1):276–281, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pardoe:2009:CTT

[3284] H. Pardoe and G. Marteinsdóttir. Contrasting trends in two condition indices: bathymetric and spatial variation in autumn condition of Icelandic cod Gadus morhua. Journal of Fish Biology, 75(1):282–289, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Neuenfeldt:2009:SAC

[3285] S. Neuenfeldt, K. H. Andersen, and H.-H. Hinrichsen. Some Atlantic cod Gadus morhua in the Baltic Sea visit hypoxic water briefly but often. Journal of Fish Biology, 75(1):290–294, July 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haynes:2009:PGM

[3286] G. D. Haynes, D. M. Gilligan, P. Grewe, and F. W. Nicholas. Population genetics and management units of invasive common carp Cyprinus carpio in the Murray–Darling Basin, Australia. Journal of Fish Biology, 75(2): 295–320, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pereira:2009:GET

[3287] A. N. Pereira, A. Márquez, M. Marin, and Y. Marin. Genetic evidence of two stocks of the whitemouth croaker Micropogonias furnieri in the Río de la Plata and oceanic front in Uruguay. Journal of Fish Biology, 75(2):321–331, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Larmuseau:2009:MPD

[3288] M. H. D. Larmuseau, J. Freyhof, F. A. M. Volckaert, and J. K. J. Van Houdt. Matrilinear phylogeography and demographical patterns of Rutilus rutilus: implications for taxonomy and conservation. Journal of Fish Biology, 75(2):332–353, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hua:2009:PAE

[3289] X. Hua, W. Wang, W. Yin, Q. He, B. Jin, J. Li, J. Chen, and C. Fu. Phylogeographical analysis of an estuarine fish, Salanx ariakensis (Osmeridae: Salanginae) in the north-western Pacific. Journal of Fish Biology, 75(2):354–367, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Alekseyev:2009:PSD

[3290] S. S. Alekseyev, R. Bajno, N. V. Gordeeva, J. D. Reist, M. Power, A. F. Kirillov, V. P. Samusenok, and A. N. Matveev. Phylogeography and sympatric differentiation of the Arctic charr Salvelinus alpinus (L.) complex in Siberia as revealed by mtDNA sequence analysis. Journal of Fish Biology, 75(2):368–392, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jamandre:2009:PFM

[3291] B. W. Jamandre, J.-D. Durand, and W. N. Tzeng. Phylogeography of the flathead mullet Mugil cephalus in the north-west Pacific as inferred from the mtDNA control region. Journal of Fish Biology, 75(2):393– 407, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kantek:2009:CLD

[3292] D. L. Z. Kantek, M. R. Vicari, W. A. M. Peres, M. M. Cestari, R. F. Artoni, L. A. C. Bertollo, and O. Moreira-Filho. Chromosomal location and distribution of As51 satellite DNA in five species of the genus Astyanax (Teleostei, Characidae, Incertae sedis). Journal of Fish Biology, 75(2):408–421, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ji:2009:MCM

[3293] X.-S. Ji, S.-L. Chen, X.-L. Liao, J.-F. Yang, T.-J. Xu, H.-Y. Ma, Y.- S. Tian, Y.-L. Jiang, and P.-F. Wu. Microsatellite-centromere mapping in Cynoglossus semilaevis using gynogenetic diploid families produced by the use of homologous and non-homologous sperm. Journal of Fish Biology, 75(2):422–434, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Abbas:2009:FRN

[3294] K. Abbas, M. Y. Li, W. M. Wang, and X. Y. Zhou. First record of the natural occurrence of hexaploid loach Misgurnus anguillicaudatus in Hubei Province, China. Journal of Fish Biology, 75(2):435–441, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Liu:2009:CEH

[3295] J.-X. Liu and B. Ely. Complex evolution of a highly conserved microsatellite locus in several fish species. Journal of Fish Biology, 75(2):442–447, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Noleto:2009:UZZ

[3296] R. B. Noleto, A. P. Amorim, M. R. Vicari, R. F. Artoni, and M. M. Cestari. An unusual ZZ/ZW sex chromosome system in Characidium fishes (Crenuchidae, Characiformes) with the presence of rDNA sites. Journal of Fish Biology, 75(2):448–453, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Popper:2009:EAS

[3297] A. N. Popper and M. C. Hastings. The effects of anthropogenic sources of sound on fishes. Journal of Fish Biology, 75(3):455–489, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Khait:2009:GCJ

[3298] V. Khait, E. Tahiraj, N. Seemungal, S. Breakstone, and P. Moller. Group cohesion in juvenile weakly electric fish mormyrus rume proboscirostris. Journal of Fish Biology, 75(3):490–502, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jung:2009:PSM

[3299] C. A. Jung, N. C. Barbee, and S. E. Swearer. Post-settlement migratory behaviour and growth-related costs in two diadromous fish species, Galaxias maculatus and Galaxias brevipinnis. Journal of Fish Biology, 75(3):503–515, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paulsen:2009:EES

[3300] H. Paulsen, O. S. Kjesbu, V. Buehler, R. A. J. Case, C. Clemmesen, G. Carvalho, L. Hauser, W. F. Hutchinson, E. Moksness, H. Otterå, A. Thorsen, and T. SvåSand. Effects of egg size, parental origin and feeding conditions on growth of larval and juvenile cod Gadus morhua. Journal of Fish Biology, 75(3):516–537, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nadeau:2009:PPE

[3301] P. S. Nadeau, S. G. Hinch, L. B. Pon, and D. A. Patterson. Persistent parental effects on the survival and size, but not burst swimming performance of juvenile sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 75(3):538–551, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Brown:2009:SCC

[3302] G. E. Brown, M. C. Harvey, A. O. H. C. Leduc, M. C. O. Ferrari, and D. P. Chivers. Social context, competitive interactions and the dynamic nature of antipredator responses of juvenile rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 75(3):552–562, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yamane:2009:FRT

[3303] H. Yamane, K. Watanabe, and Y. Nagata. Flexibility of reproductive tactics and their consequences in the brood parasitic fish Pungtungia herzi (Teleostei: Cyprinidae). Journal of Fish Biology, 75(3):563–574, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anto:2009:PSF

[3304] J. Anto, J. Majoris, and R. G. Turingan. Prey selection and functional morphology through ontogeny of Amphiprion clarkii with a congeneric comparison. Journal of Fish Biology, 75(3):575–590, August 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Blake:2009:SFG

[3305] R. W. Blake, J. Li, and K. H. S. Chan. Swimming in four goldfish Carassius auratus morphotypes: understanding functional design and performance employing artificially selected forms. Journal of Fish Biology, 75 (3):591–617, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:2009:ESN

[3306] R. J. Wootton and D. A. Fletcher. Effect of spawning number and ration on reproductive performance of the batch-spawning three-spined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 75(3):618–629, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Funes-Rodriguez:2009:IDP

[3307] R. Funes-Rodríguez, J. F. Elorduy-Garay, A. Hinojosa-Medina, and A. Zárate-Villafranco. Interannual distribution of Pacific hake Merluccius productus larvae in the southern part of the California Current. Journal of Fish Biology, 75(3):630–646, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Papastamatiou:2009:DSF

[3308] Y. P. Papastamatiou, J. E. Caselle, A. M. Friedlander, and C. G. Lowe. Distribution, size frequency, and sex ratios of blacktip reef sharks Carcharhinus melanopterus at Palmyra Atoll: a predator-dominated ecosystem. Journal of Fish Biology, 75(3):647–654, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oshiumi:2009:RES

[3309] C. Oshiumi and J. Kitamura. The reproductive ecology of the southern red tabira bitterling Acheilognathus tabira jordani in Japan. Journal of Fish Biology, 75(3):655–667, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Munro:2009:TMF

[3310] A. R. Munro, B. M. Gillanders, S. Thurstan, D. A. Crook, and A. C. Sanger. Transgenerational marking of freshwater fishes with enriched stable isotopes: a tool for fisheries management and research. Journal of Fish Biology, 75(3):668–684, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kroon:2009:BRE

[3311] F. J. Kroon, P. L. Munday, and D. A. Westcott. Book review: Equivalent whole-body concentrations of 11-ketotestosterone in female and male coral goby Gobiodon erythrospilus, a bidirectional sex-changing fish. Journal of Fish Biology, 75(3):685–692, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Weinstein:2009:GCC

[3312] M. P. Weinstein, S. Y. Litvin, V. G. Guida, and R. C. Chambers. Is global climate change influencing the overwintering distribution of weakfish Cynoscion regalis? Journal of Fish Biology, 75(3):693–698, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tavernetti:2009:EBF

[3313] R. Tavernetti, S. Morgan, and Q. Yu. Effect of biological fouling on passive collectors used to estimate fish recruitment. Journal of Fish Biology, 75(3):699–706, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carassou:2009:DPS

[3314] L. Carassou, R. Le borgne, and D. Ponton. Diet of pre-settlement larvae of coral-reef fishes: selection of prey types and sizes. Journal of Fish Biology, 75(3):707–715, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Carter:2009:IAB

[3315] A. B. Carter, A. J. Williams, and G. R. Russ. Increased accuracy of batch fecundity estimates using oocyte stage ratios in Plectropomus leopardus. Journal of Fish Biology, 75(3):716–722, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cavalli:2009:TFH

[3316] L. Cavalli, C. M. Knight, M. Durbec, R. Chappaz, and R. E. Gozlan. Twenty-four hours in the life of Zingel asper. Journal of Fish Biology, 75(3):723–727, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hain:2009:KAI

[3317] T. J. A. Hain and B. D. Neff. Kinship affects innate responses to a predator in bluegill Lepomis macrochirus larvae. Journal of Fish Biology, 75(3):728–737, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baldauf:2009:CAT

[3318] S. A. Baldauf, H. Kullmann, T. Thünken, S. Winter, and T. C. M. Bakker. Computer animation as a tool to study preferences in the cichlid Pelvicachromis taeniatus. Journal of Fish Biology, 75(3):738–746, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Thebault:2009:NMR

[3319] M. T. Thebault, K. Pichavant, and N. Kervarec. ^{31}P nuclear magnetic resonance measurements of phosphate metabolites and intracellular pH in turbot Psetta maxima red blood cells using a novel flow method. Journal of Fish Biology, 75(3):747–754, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Winfield:2009:BRF

[3320] Ian J. Winfield. Book review: Fish Behaviour — Edited by C. Magnhagen, V. A. Braithwaite, E. Forsgren and B. G. Kapoor. Journal of Fish Biology, 75(3):755–756, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jobling:2009:BRF

[3321] Malcolm Jobling. Book review: Feeding and Digestive Functions in Fishes — Edited by J. E. P. Cyrino, D. Bureau and B. G. Kapoor. Journal of Fish Biology, 75(3):756–757, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Feist:2009:BRA

[3322] Stephen W. Feist. Book review: Atlas of Fish Histology — Edited by F. Genten, E. Terwinghe and A. Danguy. Journal of Fish Biology, 75 (3):757–758, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Crean:2009:BRA

[3323] Kevin Crean. Book review: Applied Bioinformatics, Statistics and Economics in Fisheries Research — Edited by A. K. Roy and N. Sarangi. Journal of Fish Biology, 75(3):758–759, August 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Smith:2009:MCM

[3324] W. D. Smith, J. J. Bizzarro, V. P. Richards, J. Nielsen, F. Márquez-Flarías, and M. S. Shivji. Morphometric convergence and molecular divergence: the taxonomic status and evolutionary history of Gymnura crebripunctata and Gymnura marmorata in the eastern Pacific Ocean. Journal of Fish Biology, 75(4):761–783, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Falahatkar:2009:RHC

[3325] B. Falahatkar, S. Poursaeid, M. Shakoorian, and B. Barton. Responses to handling and confinement stressors in juvenile great sturgeon Huso huso. Journal of Fish Biology, 75(4):784–796, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franco:2009:FFI

[3326] M. A. L. Franco, A. C. Braga, G. W. A. Nunan, and P. A. S. Costa. Fishes of the family Ipnopidae (Teleostei: Aulopiformes) collected on the Brazilian continental slope between 11^{o}and 23^{o}S. Journal of Fish Biology, 75(4):797–815, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2009:RFA

[3327] F. Lahnsteiner. The role of free amino acids in semen of rainbow trout Oncorhynchus mykiss and carp Cyprinus carpio. Journal of Fish Biology, 75(4):816–833, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hartviksen:2009:DCI

[3328] M. B. Hartviksen, Y. Kamisaka, A.-E. O. Jordal, R. M. Koedijk, and I. Rønnestad. Distribution of cholecystokinin-immunoreactive cells in the gut of developing Atlantic cod Gadus morhua L. larvae fed zooplankton or rotifers. Journal of Fish Biology, 75(4):834–844, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yan:2009:HSS

[3329] H. Y. Yan. A histochemical study on the snout tentacles and snout skin of bristlenose catfish Ancistrus triradiatus. Journal of Fish Biology, 75 (4):845–861, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Patel:2009:IPT

[3330] M. Patel, F. I. Iftikar, E. M. Leonard, Y. K. Ip, and C. M. Wood. Ionoregulatory physiology of two species of African lungfishes Protopterus dolloi and Protopterus annectens. Journal of Fish Biology, 75(4):862–884, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMullin:2009:LVL

[3331] V. A. McMullin, K. R. Munkittrick, and D. A. Methven. Latitudinal variability in lunar spawning rhythms: absence of a lunar pattern in the northern mummichog Fundulus heteroclitus macrolepidotus. Journal of Fish Biology, 75(4):885–900, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svendsen:2009:VTS

[3332] J. C. Svendsen, K. Aarestrup, J. Dolby, T. C. Svendsen, and R. H. B. Christensen. The volitional travel speed varies with reproductive state in mature female brown trout Salmo trutta. Journal of Fish Biology, 75 (4):901–907, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Colleye:2009:ASS

[3333] O. Colleye, B. Frederich, P. Vandewalle, M. Casadevall, and E. Parmentier. Agonistic sounds in the skunk clownfish Amphiprion akallopisos: size-related variation in acoustic features. Journal of Fish Biology, 75(4): 908–916, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Powell:2009:ROB

[3334] A. Powell, G. S. Parry, J. D. R. Houghton, and D. M. Herdson. Rare occurrence of a bigeye tuna Thunnus obesus in British waters, with notes on other occurrences of sub-tropical tunas. Journal of Fish Biology, 75 (4):917–923, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pitcher:2009:SDF

[3335] T. E. Pitcher, J. J. Beausoleil, J. A. Abbott, and J. L. Vandereerden. Sperm design and function in the redside dace Clinostomus elongatus. Journal of Fish Biology, 75(4):924–931, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zamani:2009:ADE

[3336] A. Zamani, A. Hajimoradloo, R. Madani, and M. Farhangi. Assessment of digestive enzymes activity during the fry development of the endangered Caspian brown trout Salmo caspius. Journal of Fish Biology, 75 (4):932–937, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lim:2009:EAT

[3337] S. M. Lim and D. J. Ellerby. The effects of acute temperature change on cost of transport at maximal labriform speed in bluegill Lepomis macrochirus. Journal of Fish Biology, 75(4):938–943, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Ee

[3338] Anonymous. Erratum. Journal of Fish Biology, 75(4):944, September 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hino:2009:MBR

[3339] H. Hino, N. G. Miles, H. Bandoh, and H. Ueda. Molecular biological research on olfactory chemoreception in fishes. Journal of Fish Biology, 75(5):945–959, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Puebla:2009:ESM

[3340] O. Puebla. Ecological speciation in marine v. freshwater fishes. Journal of Fish Biology, 75(5):960–996, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ketmaier:2009:PBA

[3341] V. Ketmaier, F. Finamore, C. Largiadèr, M. Milone, and P. G. Bianco. Phylogeography of bleaks Alburnus spp. (Cyprinidae) in Italy, based on cytochrome b data. Journal of Fish Biology, 75(5):997–1017, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vreven:2009:NMS

[3342] E. J. Vreven and J. Snoeks. A new mastacembelid species from Lake Tanganyika: a case of complex evolutionary history. Journal of Fish Biology, 75(5):1018–1047, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dhillon:2009:CEM

[3343] R. S. Dhillon, A. J. Esbaugh, Y. S. Wang, and B. L. Tufts. Characterization and expression of a myosin heavy–chain isoform in juvenile walleye Sander vitreus. Journal of Fish Biology, 75(5):1048–1062, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Geist:2009:IGC

[3344] J. Geist, M. Kolahsa, B. Gum, and R. Kuehn. The importance of genetic cluster recognition for the conservation of migratory fish species: the example of the endangered European huchen Hucho hucho (L.). Journal of Fish Biology, 75(5):1063–1078, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mani:2009:KDE

[3345] I. Mani, R. Kumar, M. Singh, B. Kushwaha, N. S. Nagpure, P. K. Srivastava, K. Murmu, D. S. K. Rao, and W. S. Lakra. Karyotypic diversity and evolution of seven mahseer species (Cyprinidae) from India. Journal of Fish Biology, 75(5):1079–1091, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhong:2009:OYI

[3346] S. Zhong, Y.-P. Wang, D.-S. Pei, D.-J. Luo, L.-J. Liao, and Z.-Y. Zhu. A one-year investigation of the relationship between serum GH levels and the growth of F_{4} transgenic and non-transgenic common carp Cyprinus carpio. Journal of Fish Biology, 75(5):1092–1100, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vergara-Chen:2009:IWC

[3347] C. Vergara-Chen, M. González-Wangüemert, E. Bermingham, and L. D’Croz. Identification of weakfish Cynoscion (Gill) in the Bay of Panama with RFLP markers. Journal of Fish Biology, 75(5):1101–1107, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haney:2009:CLW

[3348] R. A. Haney, B. J. Turner, and D. M. Rand. A cryptic lineage within the pupfish Cyprinodon dearborni suggests multiple colonizations of South America. Journal of Fish Biology, 75(5):1108–1114, October 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Merette:2009:DDG

[3349] D. Mérette, A. Bourret, and J. Turgeon. Diversity and distribution of gynogenetic hybrids between Fundulus diaphanus and Fundulus heteroclitus in Porter’s Lake (Nova Scotia) in relation to salinity. Journal of Fish Biology, 75(5):1115–1122, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nagelkerken:2009:CDC

[3350] I. Nagelkerken, G. Van Der Velde, S. L. J. Wartenbergh, M. M. Nugues, and M. S. Pratchett. Cryptic dietary components reduce dietary overlap among sympatric butterflyfishes (chaetodontidae). Journal of Fish Biology, 75(6):1123–1143, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Burns:2009:RPV

[3351] J. G. Burns, P. Di Nardo, and F. H. Rodd. The role of predation in variation in body shape in guppies Poecilia reticulata: a comparison of field and common garden phenotypes. Journal of Fish Biology, 75(6): 1144–1157, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2009:ETI

[3352] L. Zhang, Y. J. Wang, M. H. Hu, Q. X. Fan, S. G. Chenung, P. K. S. Shin, H. Li, and L. Cao. Effects of the timing of initial feeding on growth and survival of spotted mandarin fish Siniperca scherzeri larvae. Journal of Fish Biology, 75(6):1158–1172, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wang:2009:STV

[3353] C. H. Wang, Y. T. Lin, J. C. Shiao, C. F. You, and W. N. Tzeng. Spatiotemporal variation in the elemental compositions of otoliths of southern bluefin tuna Thunnus maccoyii in the Indian Ocean and its ecological implication. Journal of Fish Biology, 75(6):1173–1193, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Parafati:2009:LTH

[3354] M. Parafati, O. Senatori, and A. Nicotra. Localization of tyrosine hydroxylase immunoreactive neurons in the forebrain of the guppy Poecilia reticulata. Journal of Fish Biology, 75(6):1194–1205, October 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Matsuzaki:2009:BMD

[3355] S. S. Matsuzaki, K. Mabuchi, N. Takamura, M. Nishida, and I. Washitani. Behavioural and morphological differences between feral and domesticated strains of common carp Cyprinus carpio. Journal of Fish Biology, 75(6):1206–1220, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zupa:2009:HIO

[3356] R. Zupa, A. Corriero, M. Deflorio, N. Santamaria, D. Spedicato, C. Marano, M. Losurdo, C. R. Bridges, and G. De Metrio. A histological investigation of the occurrence of non-reproductive female bluefin tuna Thunnus thynnus in the Mediterranean Sea. Journal of Fish Biology, 75(6):1221–1229, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Horppila:2009:FNS

[3357] J. Horppila and L. Nurminen. Food niche segregation between two herbivorous cyprinid species in a turbid lake. Journal of Fish Biology, 75 (6):1230–1243, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tierney:2009:IMC

[3358] K. B. Tierney, D. A. Patterson, and C. J. Kennedy. The influence of maternal condition on offspring performance in sockeye salmon Oncorhynchus nerka. Journal of Fish Biology, 75(6):1244–1257, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Barnett:2009:ADF

[3359] L. A. K. Barnett, D. A. Ebert, and G. M. Cailliet. Assessment of the dorsal fin spine for chimaeroid (Holocephali: Chimaeriformes) age estimation. Journal of Fish Biology, 75(6):1258–1270, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fontes:2009:MSR

[3360] J. Fontes, J. E. Caselle, P. Afonso, and R. S. Santos. Multi-scale recruitment patterns and effects on local population size of a temperate reef fish. Journal of Fish Biology, 75(6):1271–1286, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ruggerone:2009:LMF

[3361] G. T. Ruggerone, J. L. Nielsen, and B. A. Agler. Linking marine and freshwater growth in western Alaska Chinook salmon Oncorhynchus tshawytscha. Journal of Fish Biology, 75(6):1287–1301, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sucre:2009:EDD

[3362] E. Sucré, M. Charmantier-Daures, E. Grousset, G. Charmantier, and P. Cucchi-Mouillot. Early development of the digestive tract (pharynx and gut) in the embryos and pre-larvae of the European sea bass Dicentrarchus labrax. Journal of Fish Biology, 75(6):1302–1322, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shohet:2009:FGP

[3363] A. J. Shohet and P. J. Watt. Female guppies Poecilia reticulata prefer males that can learn fast. Journal of Fish Biology, 75(6):1323–1330, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Preciado:2009:RUD

[3364] I. Preciado, J. E. Cartes, A. Serrano, F. Velasco, I. Olaso, F. Sánchez, and I. Frutos. Resource utilization by deep-sea sharks at the Le Danois Bank, Cantabrian Sea, north-east Atlantic Ocean. Journal of Fish Biology, 75 (6):1331–1355, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rakowitz:2009:IHM

[3365] G. Rakowitz, J. Kubečka, C. Fesl, and H. Keckeis. Intercalibration of hydroacoustic and mark–recapture methods for assessing the spawning population size of a threatened fish species. Journal of Fish Biology, 75 (6):1356–1370, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Adamson:2009:PSS

[3366] E. A. S. Adamson, D. A. Hurwood, A. M. Baker, and P. B. Mather. Population subdivision in Siamese mud carp Henicorhynchus siamensis in the Mekong River basin: implications for management. Journal of Fish Biology, 75(6):1371–1392, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bradbury:2009:SFC

[3367] I. R. Bradbury, P. V. R. Snelgrove, P. Bentzen, B. De Young, R. S. Gregory, and C. J. Morris. Structural and functional connectivity of marine fishes within a semi-enclosed Newfoundland fjord. Journal of Fish Biology, 75(6):1393–1409, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conrad:2009:BTN

[3368] J. L. Conrad and A. Sih. Behavioural type in newly emerged steelhead Oncorhynchus mykiss does not predict growth rate in a conventional hatchery rearing environment. Journal of Fish Biology, 75(6):1410–1426, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reyjol:2009:ETB

[3369] Y. Reyjol, J.-P. Léna, F. Hervant, and D. Pont. Effects of temperature on biological and biochemical indicators of the life-history strategy of bullhead Cottus gobio. Journal of Fish Biology, 75(6):1427–1445, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pettersson:2009:ECR

[3370] A. Pettersson, J. Pickova, and E. Brännäs. Effects of crude rapeseed oil on lipid composition in Arctic charr Salvelinus alpinus. Journal of Fish Biology, 75(6):1446–1458, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Duan:2009:EAC

[3371] M. Duan, T. Zhang, W. Hu, L. F. Sundström, Y. Wang, Z. Li, and Z. Zhu. Elevated ability to compete for limited food resources by ‘allfish’ growth hormone transgenic common carp Cyprinus carpio. Journal of Fish Biology, 75(6):1459–1472, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Loewen:2009:ENS

[3372] T. N. Loewen, D. Gillis, and R. F. Tallman. Ecological niche specialization inferred from morphological variation and otolith strontium of Arctic charr Salvelinus alpinus L. found within open lake systems of southern Baffin Island, Nunavut, Canada. Journal of Fish Biology, 75 (6):1473–1495, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Niedermuller:2009:FSC

[3373] S. K. Niedermüller, L. S. Schiemer, and J. Herler. Food selection of a corallivorous butterflyfish Chaetodon austriacus in the Red Sea, with particular emphasis on size and fish symbionts of corals. Journal of Fish Biology, 75(6):1496–1504, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Woodall:2009:FOL

[3374] L. C. Woodall, H. J. Koldewey, S. V. Santos, and P. W. Shaw. First occurrence of the lined seahorse Hippocampus erectus in the eastern Atlantic Ocean. Journal of Fish Biology, 75(6):1505–1512, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Newman:2009:FRR

[3375] S. J. Newman. First record of Randall’s snapper Randallichthys filamentosus (Perciformes: Lujanidae) from the eastern Indian Ocean (northwestern Australia). Journal of Fish Biology, 75(6):1513–1517, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Partridge:2009:FAM

[3376] G. J. Partridge, G. I. Jenkins, R. G. Doupé, S. De Lestang, B. M. Ginbey, and D. French. Factors affecting mark quality of alizarin complexonestained otoliths in juvenile black bream Acanthopagrus butcheri and a prescription for dosage. Journal of Fish Biology, 75(6):1518–1523, October 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bernreuther:2009:ETG

[3377] M. Bernreuther, A. Temming, and J.-P. Herrmann. Effect of temperature on the gastric evacuation in sprat Sprattus sprattus. Journal of Fish Biology, 75(7):1525–1541, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lahnsteiner:2009:EDK

[3378] F. Lahnsteiner. The effect of different kinds of electrolyte and nonelectrolyte solutions on the survival rate and morphology of zebrafish Danio rerio embryos. Journal of Fish Biology, 75(7):1542–1559, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Dimmlich:2009:SDB

[3379] W. F. Dimmlich, T. M. Ward, and W. G. Breed. Spawning dynamics and biomass estimates of an anchovy Engraulis australis population in contrasting gulf and shelf environments. Journal of Fish Biology, 75(7): 1560–1576, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wisenden:2009:FCN

[3380] B. D. Wisenden, A. Unruh, A. Morantes, S. Bury, B. Curry, R. Driscoll, M. Hussein, and S. Markegard. Functional constraints on nest characteristics of pebble mounds of breeding male hornyhead chub Nocomis biguttatus. Journal of Fish Biology, 75(7):1577–1585, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fitzgibbon:2009:PMI

[3381] Q. P. Fitzgibbon and R. S. Seymour. Postprandial metabolic increment of southern bluefin tuna Thunnus maccoyii ingesting high or low-lipid sardines Sardinops sagax. Journal of Fish Biology, 75(7):1586–1600, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huysentruyt:2009:OCM

[3382] F. Huysentruyt, M. Brunain, and D. Adriaens. Ontogeny of the cranial musculature in Corydoras aeneus Callichthyidae, Siluriformes. Journal of Fish Biology, 75(7):1601–1614, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lefrancois:2009:EPH

[3383] C. Lefrançois, R. S. Ferrari, J. Moreira da Silva, and P. Domenici. The effect of progressive hypoxia on spontaneous activity in single and shoaling golden grey mullet Liza aurata. Journal of Fish Biology, 75(7):1615–1625, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nendick:2009:SPA

[3384] L. Nendick, A. Grant, M. Gardner, M. Sackville, C. J. Brauner, and A. P. Farrell. Swimming performance and associated ionic disturbance of juvenile pink salmon Oncorhynchus gorbuscha determined using different acceleration profiles. Journal of Fish Biology, 75(7):1626–1638, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Oshima:2009:ETF

[3385] M. Oshima, Y. Kato, R. Masuda, S. Kimura, and Y. Yamashita. Effect of turbulence on feeding intensity and survival of Japanese flounder Paralichthys olivaceus pelagic larvae. Journal of Fish Biology, 75(7): 1639–1647, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cicia:2009:SAE

[3386] A. M. Cicia, W. B. Driggers III, G. W. Ingram, Jr., J. Kneebone, P. C. W. Tsang, D. M. Koester, and J. A. Sulikowski. Size and age estimates at sexual maturity for the little skate Leucoraja erinacea from the western Gulf of Maine, U.S.A. Journal of Fish Biology, 75(7):1648–1666, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Poisot:2009:IBI

[3387] T. Poisot, A. Šimková, P. Hyršl, and S. Morand. Interactions between immunocompetence, somatic condition and parasitism in the chub Leuciscus cephalus in early spring. Journal of Fish Biology, 75(7):1667–1682, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Binning:2009:SMG

[3388] S. A. Binning, L. J. Chapman, and A. Cosandey-Godin. Specialized morphology for a generalist diet: evidence for Liem’s Paradox in a cichlid fish. Journal of Fish Biology, 75(7):1683–1699, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Davidsen:2009:MBS

[3389] J. G. Davidsen, A. H. Rikardsen, E. Halttunen, E. B. Thorstad, F. Økland, B. H. Letcher, J. Skardhamar, and T. F. Næsje. Migratory behaviour and survival rates of wild northern Atlantic salmon Salmo salar post-smolts: effects of environmental factors. Journal of Fish Biology, 75(7):1700–1718, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wells:2009:FEP

[3390] R. J. D. Wells and J. R. Rooker. Feeding ecology of pelagic fish larvae and juveniles in slope waters of the Gulf of Mexico. Journal of Fish Biology, 75(7):1719–1732, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Millot:2009:RTB

[3391] S. Millot, M.-L. Bégout, and B. Chatain. Risk-taking behaviour variation over time in sea bass Dicentrarchus labrax: effects of day–night alternation, fish phenotypic characteristics and selection for growth. Journal of Fish Biology, 75(7):1733–1749, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malca:2009:AGG

[3392] E. Malca, J. F. Barimo, J. E. Serafy, and P. J. Walsh. Age and growth of the gulf toadfish Opsanus beta based on otolith increment analysis. Journal of Fish Biology, 75(7):1750–1761, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Simonot:2009:CVV

[3393] D. L. Simonot and A. P. Farrell. Coronary vascular volume remodelling in rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 75(7): 1762–1772, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Melnychuk:2009:MED

[3394] M. C. Melnychuk and V. Christensen. Methods for estimating detection efficiency and tracking acoustic tags with mobile transect surveys. Journal of Fish Biology, 75(7):1773–1794, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reardon:2009:HLH

[3395] E. E. Reardon and L. J. Chapman. Hypoxia and life-history traits in a eurytopic African cichlid. Journal of Fish Biology, 75(7):1795–1815, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mendes:2009:DTP

[3396] T. C. Mendes, R. C. Villaça, and C. E. L. Ferreira. Diet and trophic plasticity of an herbivorous blenny Scartella cristata of subtropical rocky shores. Journal of Fish Biology, 75(7):1816–1830, November 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Krog:2009:DGE

[3397] C. Krog and N. G. Andersen. Are the dynamics of gastric evacuation of natural prey in a piscivorous flatfish different from what is going on in a gadoid? Journal of Fish Biology, 75(7):1831–1844, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haesler:2009:RPM

[3398] M. P. Haesler, C. M. Lindeyer, and M. Taborsky. Reproductive parasitism: male and female responses to conspecific and heterospecific intrusions at spawning in a mouth-brooding cichlid Ophthalmotilapia ventralis. Journal of Fish Biology, 75(7):1845–1856, November 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Heenan:2009:RDC

[3399] A. Heenan, S. D. Simpson, M. G. Meekan, S. D. Healy, and V. A. Braithwaite. Restoring depleted coral-reef fish populations through recruitment enhancement: a proof of concept. Journal of Fish Biology, 75(7):1857– 1867, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Webster:2009:ENI

[3400] M. M. Webster and K. N. Laland. Evaluation of a non-invasive tagging system for laboratory studies using three-spined sticklebacks Gasterosteus aculeatus. Journal of Fish Biology, 75(7):1868–1873, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Honorato-Sampaio:2009:OSB

[3401] K. Honorato-Sampaio, G. B. Santos, N. Bazzoli, and E. Rizzo. Observations on the seasonal breeding biology and fine structure of the egg surface in the white piranha Serrasalmus brandtii from the São Francisco River basin, Brazil. Journal of Fish Biology, 75(7):1874–1882, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Malavasi:2009:CSA

[3402] S. Malavasi, C. Valerio, and P. Torricelli. Courtship sounds and associated behaviours in the Canestrini’s goby Pomatoschistus canestrinii. Journal of Fish Biology, 75(7):1883–1887, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Young:2009:EAA

[3403] M. J. Young. The efficacy of the aquatic anaesthetic AQUI-S(R) for anaesthesia of a small freshwater fish, Melanotaenia australis. Journal of Fish Biology, 75(7):1888–1894, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Welker:2009:PEO

[3404] T. L. Welker and J. L. Congleton. Preliminary examination of oxidative stress in juvenile spring Chinook salmon Oncorhynchus tshawytscha of wild origin sampled from transport barges. Journal of Fish Biology, 75(7):1895–1905, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hilborn:2009:BRE

[3405] Ray Hilborn. Book review: Enclosing the Fisheries: People, Places and Power — Edited by M. E. Lowe and C. Carothers. Journal of Fish Biology, 75(7):1906–1908, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Nunn:2009:BRF

[3406] Andy D. Nunn. Book review: Fish larval physiology — Edited by R. N. Finn and B. G. Kapoor. Journal of Fish Biology, 75(7):1908–1909, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Proudlove:2009:BRF

[3407] Graham S. Proudlove. Book review: Fish life in special environments — Edited by P. Sebert, D. W. Onyango and B. G. Kapoor. Journal of Fish Biology, 75(7):1909–1911, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tipping:2009:BRM

[3408] Edward Tipping. Book review: Metal contamination in aquatic environments. science and lateral management — By Samuel N. Luoma and Philip S. Rainbow. Journal of Fish Biology, 75(7):1911–1912, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeomans:2009:BRS

[3409] Willie Yeomans. Book review: Scotland’s freshwater fish: ecology, conservation & folklore — By P. S. Maitland. Journal of Fish Biology, 75(7): 1912–1914, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Axford:2009:BRE

[3410] Stephen Axford. Book review: The Ecosystem Approach to Fisheries — Edited by G. Bianchi & H. R. Skjoldal. Journal of Fish Biology, 75(7): 1914–1916, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wolter:2009:BRT

[3411] Christian Wolter. Book review: Turbinenbedingte Schädigung des Aals (Anguilla anguilla) — By G. Ebel. Journal of Fish Biology, 75(7):1916– 1917, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Wootton:2009:DSG

[3412] R. J. Wootton. The Darwinian stickleback Gasterosteus aculeatus: a history of evolutionary studies. Journal of Fish Biology, 75(8):1919–1942, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2009:TSS

[3413] F. A. Huntingford and M. L. Ruiz-Gomez. Three-spined sticklebacks Gasterosteus aculeatus as a model for exploring behavioural biology. Journal of Fish Biology, 75(8):1943–1976, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bell:2009:IFS

[3414] M. A. Bell. Implications of a fossil stickleback assemblage for Darwinian gradualism. Journal of Fish Biology, 75(8):1977–1999, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hendry:2009:ASC

[3415] A. P. Hendry, D. I. Bolnick, D. Berner, and C. L. Peichel. Along the speciation continuum in sticklebacks. Journal of Fish Biology, 75(8): 2000–2036, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gelmond:2009:RES

[3416] O. Gelmond, F. A. Von Hippel, and M. S. Christy. Rapid ecological speciation in three-spined stickleback Gasterosteus aculeatus from middleton Island, Alaska: the roles of selection and geographic isolation. Journal of Fish Biology, 75(8):2037–2051, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zanella:2009:OFS

[3417] L. N. Zanella, D. Zanella, M. Mraković, M. Miletić, P. Mustafić, and M. Ćaleta. Occurrence of four-spined Gasterosteus aculeatus in an isolated Croatian river population. Journal of Fish Biology, 75(8):2052– 2061, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Myhre:2009:BAL

[3418] F. Myhre and T. Klepaker. Body armour and lateral-plate reduction in freshwater three-spined stickleback Gasterosteus aculeatus: adaptations to a different buoyancy regime? Journal of Fish Biology, 75(8):2062– 2074, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yeates-Burghart:2009:LVP

[3419] Q. S. Yeates-Burghart, C. O’Brien, W. A. Cresko, C. M. Holzapfel, and W. E. Bradshaw. Latitudinal variation in photoperiodic response of the three-spined stickleback Gasterosteus aculeatus in western North America. Journal of Fish Biology, 75(8):2075–2081, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sebire:2009:FRN

[3420] M. Sebire, I. Katsiadaki, and A. P. Scott. Further refinement of the noninvasive procedure for measuring steroid production in the male threespined stickleback Gasterosteus aculeatus. Journal of Fish Biology, 75(8): 2082–2094, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Macnab:2009:RPS

[3421] V. Macnab, I. Katsiadaki, and I. Barber. Reproductive potential of Schistocephalus solidus-infected male three-spined stickleback Gasterosteus aculeatus from two U.K. populations. Journal of Fish Biology, 75(8): 2095–2107, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Candolin:2009:PRA

[3422] U. Candolin. Population responses to anthropogenic disturbance: lessons from three-spined sticklebacks Gasterosteus aculeatus in eutrophic habitats. Journal of Fish Biology, 75(8):2108–2121, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lenz:2009:DRM

[3423] T. L. Lenz, C. Eizaguirre, J. P. Scharsack, M. Kalbe, and M. Milinski. Disentangling the role of MHC-dependent ‘good genes’ and ‘compatible genes’ in mate-choice decisions of three-spined sticklebacks Gasterosteus aculeatus under semi-natural conditions. Journal of Fish Biology, 75(8): 2122–2142, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Frommen:2009:PIB

[3424] J. G. Frommen, M. Mehlis, and T. C. M. Bakker. Predator-inspection behaviour in female three-spined sticklebacks Gasterosteus aculeatus is associated with status of gravidity. Journal of Fish Biology, 75(8):2143– 2153, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Mehlis:2009:CAR

[3425] M. Mehlis, T. C. M. Bakker, K. Langen, and J. G. Frommen. Cain and Abel reloaded? Kin recognition and male–male aggression in threespined sticklebacks Gasterosteus aculeatus L. Journal of Fish Biology, 75(8):2154–2162, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vlieger:2009:HSD

[3426] L. Vlieger and U. Candolin. How not to be seen: does eutrophication influence three-spined stickleback Gasterosteus aculeatus sneaking behaviour? Journal of Fish Biology, 75(8):2163–2174, November 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Cerda:2009:MPD

[3427] J. Cerdà. Molecular pathways during marine fish egg hydration: the role of aquaporins. Journal of Fish Biology, 75(9):2175–2196, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Chen:2009:MPC

[3428] W.-J. Chen, V. Lheknim, and R. L. Mayden. Molecular phylogeny of the cobitoidea (Teleostei: Cypriniformes) revisited: position of enigmatic loach ellopostoma resolved with six nuclear genes. Journal of Fish Biology, 75(9):2197–2208, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

DeOliveira:2009:MCE

[3429] R. R. De Oliveira, E. Feldberg, M. B. Dos Anjos, and J. Zuanon. Mechanisms of chromosomal evolution and its possible relation to natural history characteristics in Ancistrus catfishes (Siluriformes: Loricariidae). Journal of Fish Biology, 75(9):2209–2225, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Franca:2009:SCC

[3430] G. F. França, C. Oliveira, and I. Quagio-Grassiotto. Spermatic cell characteristics in gymnotiformes (Teleostei: Ostariophysi) and their phylogenetic meaning. Journal of Fish Biology, 75(9):2226–2243, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Haponski:2009:LGP

[3431] A. E. Haponski, T. L. Bollin, M. A. Jedlicka, and C. A. Stepien. Landscape genetic patterns of the rainbow darter Etheostoma caeruleum: a catchment analysis of mitochondrial DNA sequences and nuclear microsatellites. Journal of Fish Biology, 75(9):2244–2268, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hanfling:2009:SPS

[3432] B. Hänfling, C. Dümpelmann, N. G. Bogutskaya, R. Brandl, and M. Brändle. Shallow phylogeographic structuring of Vimba vimba across Europe suggests two distinct refugia during the last glaciation. Journal of Fish Biology, 75(9):2269–2286, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ghosh:2009:NDT

[3433] J. Ghosh, R. W. Wilson, and T. Kudoh. Normal development of the tomato clownfish Amphiprion frenatus: live imaging and in situ hybridization analyses of mesodermal and neurectodermal development. Journal of Fish Biology, 75(9):2287–2298, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Fraser:2009:MCI

[3434] B. A. Fraser and B. D. Neff. MHC class IIB additive and non-additive effects on fitness measures in the guppy Poecilia reticulata. Journal of Fish Biology, 75(9):2299–2312, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bellafronte:2009:DZZ

[3435] E. Bellafronte, M. R. Vicari, R. F. Artoni, V. P. Margarido, and O. Moreira-Filho. Differentiated ZZ/ZW sex chromosomes in Apareiodon ibitiensis (Teleostei, Parodontidae): cytotaxonomy and biogeography. Journal of Fish Biology, 75(9):2313–2325, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Santos:2009:MKP

[3436] U. Santos, C. M. Völcker, F. A. Belei, M. B. Cioffi, L. A. C. Bertollo, S. R. Paiva, and J. A. Dergam. Molecular and karyotypic phylogeography in the Neotropical Hoplias malabaricus (Erythrinidae) fish in eastern Brazil. Journal of Fish Biology, 75(9):2326–2343, December 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Caputo:2009:MDV

[3437] V. Caputo, M. Giovannotti, P. Nisi Cerioni, A. Splendiani, M. Marconi, and J. Tagliavini. Mitochondrial DNA variation of an isolated population of the Adriatic brook lamprey Lampetra zanandreai (Agnatha: Petromyzontidae): phylogeographic and phylogenetic inferences. Journal of Fish Biology, 75(9):2344–2351, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zhang:2009:DCM

[3438] Y. Zhang, P. Jiang, H. Gang, J.-Q. Shi, and X. Sun. Development and characterization of microsatellite markers to study population genetic diversity of Przewalski’s naked carp Gymnocypris przewalskii in China. Journal of Fish Biology, 75(9):2352–2356, December 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pie:2009:MDR

[3439] M. R. Pie, R. A. Baggio, W. A. Boeger, L. A. Patella, A. Ostrensky, J. R. S. Vitule, and V. Abilhoa. Molecular data reveal a diverse Astyanax species complex in the upper Iguaçu River. Journal of Fish Biology, 75(9):2357–2362, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Bennett:2009:MDD

[3440] M. G. Bennett, J. Heath Howell, B. R. Kuhajda, and R. M. Wood. Mitochondrial DNA divergence in the critically imperilled pygmy madtom, Noturus stanauli (Siluriformes: Ictaluridae). Journal of Fish Biology, 75(9):2363–2372, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ocalewicz:2009:FCA

[3441] K. Ocalewicz, S. Dobosz, H. Kuzminski, and K. Goryczko. Formation of chromosome aberrations in androgenetic rainbow trout Oncorhynchus mykiss. Journal of Fish Biology, 75(9):2373–2379, December 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Ef

[3442] Anonymous. Erratum. Journal of Fish Biology, 75(9):2380, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jonsson:2009:RLE

[3443] B. Jonsson and N. Jonsson. A review of the likely effects of climate change on anadromous Atlantic salmon Salmo salar and brown trout Salmo trutta, with particular reference to water temperature and flow. Journal of Fish Biology, 75(10):2381–2447, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Arlinghaus:2009:CPS

[3444] R. Arlinghaus, A. Schwab, S. J. Cooke, and I. G. Cowx. Contrasting pragmatic and suffering-centred approaches to fish welfare in recreational angling. Journal of Fish Biology, 75(10):2448–2463, December 2009. CO-DEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

McMeans:2009:SIC

[3445] B. C. McMeans, J. A. Olin, and G. W. Benz. Stable-isotope comparisons between embryos and mothers of a placentatrophic shark species. Journal of Fish Biology, 75(10):2464–2474, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Jacobsen:2009:DRA

[3446] I. P. Jacobsen, J. W. Johnson, and M. B. Bennett. Diet and reproduction in the Australian butterfly ray Gymnura australis from northern and north-eastern Australia. Journal of Fish Biology, 75(10):2475–2489, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Pierce:2009:VAB

[3447] S. J. Pierce and M. B. Bennett. Validated annual band-pair periodicity and growth parameters of blue-spotted maskray Neotrygon kuhlii from south-east Queensland, Australia. Journal of Fish Biology, 75(10):2490– 2508, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Paddack:2009:SSD

[3448] M. J. Paddack, S. Sponaugle, and R. K. Cowen. Small-scale demographic variation in the stoplight parrotfish Sparisoma viride. Journal of Fish Biology, 75(10):2509–2526, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Vasquez:2009:LHC

[3449] A. García Vásquez, J.-C. Alonso, F. Carvajal, J. Moreau, J. Nuñez, J.-F. Renno, S. Tello, V. Montreuil, and F. Duponchelle. Life-history characteristics of the large Amazonian migratory catfish Brachyplatystoma rousseauxii in the Iquitos region, Peru. Journal of Fish Biology, 75(10): 2527–2551, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Costas:2009:CAS

[3450] N. Costas, M. Álvarez, and I. Pardo. Characterization of an Atlantic salmon Salmo salar stream at the southern limit of its eastern Atlantic distribution. Journal of Fish Biology, 75(10):2552–2570, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Baensch:2009:CHT

[3451] F. U. Baensch and C. S. Tamaru. Captive hybridization of two geographically isolated pygmy angelfish species, Centropyge fisheri and Centropyge resplendens. Journal of Fish Biology, 75(10):2571–2584, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Galvan:2009:FPS

[3452] D. E. Galván, F. Botto, A. M. Parma, L. Bandieri, N. Mohamed, and O. O. Iribarne. Food partitioning and spatial subsidy in shelter-limited fishes inhabiting patchy reefs of Patagonia. Journal of Fish Biology, 75(10):2585–2605, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Phelps:2009:OCP

[3453] Q. E. Phelps, M. R. Noatch, H. A. Lewis, D. J. Myers, J. M. Zeigler, J. S. Eichelberger, M. J. Saltzgiver, and G. W. Whitledge. Otolith chemistry of prey fish consumed by a fish predator: does digestion hinder Russian doll techniques? Journal of Fish Biology, 75(10):2606–2614, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Zanata:2009:NSM

[3454] A. M. Zanata, J. L. O. Birindelli, and C. R. Moreira. New species of Moenkhausia eigenmann (Characiformes: Characidae) from Rio Xingu and Rio Tapajós basins, Brazil, with comments on a putative case of polymorphic Batesian mimicry. Journal of Fish Biology, 75(10):2615– 2628, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Tomiyama:2009:PRP

[3455] T. Tomiyama, K. Ebe, G. Kawata, and T. Fujii. Post-release predation on hatchery-reared Japanese flounder Paralichthys olivaceus in the coast of Fukushima, Japan. Journal of Fish Biology, 75(10):2629–2641, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gheorghiu:2009:TDE

[3456] C. Gheorghiu, D. J. Marcogliese, and M. E. Scott. Temporal dynamics of epidermal responses of guppies Poecilia reticulata to a sublethal range of waterborne zinc concentrations. Journal of Fish Biology, 75(10):2642– 2656, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Felip:2009:OTF

[3457] A. Felip, M. Carrillo, and S. Zanuy. Older triploid fish retain impaired reproductive endocrinology in the European sea bass Dicentrarchus labrax. Journal of Fish Biology, 75(10):2657–2669, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Boyle:2009:CSS

[3458] K. S. Boyle and T. E. Cox. Courtship and spawning sounds in bird wrasse Gomphosus varius and saddle wrasse Thalassoma duperrey. Journal of Fish Biology, 75(10):2670–2681, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Rosa:2009:CDP

[3459] R. Rosa, M. Caetano-Filho, O. A. Shibatta, and L. Giuliano-Caetano. Cytotaxonomy in distinct populations of Hoplias aff. malabaricus (Characiformes, Erythrinidae) from lower Paranapanema River basin. Journal of Fish Biology, 75(10):2682–2694, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Ditlecadet:2009:DCI

[3460] D. Ditlecadet, P. U. Blier, N. R. Le François, and F. Dufresne. Digestive capacities, inbreeding and growth capacities in juvenile Arctic charr Salvelinus alpinus. Journal of Fish Biology, 75(10):2695–2708, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Lin:2009:DOA

[3461] Y.-J. Lin, J.-C. Shiao, L. Ložys, M. Plikšs, A. Minde, Y. Iizuka, I. Rašals, and W.-N. Tzeng. Do otolith annular structures correspond to the first freshwater entry for yellow European eels Anguilla anguilla in the Baltic countries? Journal of Fish Biology, 75(10):2709–2722, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kristoffersen:2009:FAF

[3462] J. B. Kristoffersen and A. Magoulas. Fluctuating asymmetry and fitness correlations in two Engraulis encrasicolus populations. Journal of Fish Biology, 75(10):2723–2736, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hall:2009:EAP

[3463] K. C. Hall, M. K. Broadhurst, P. A. Butcher, and S. J. Rowland. Effects of angling on post-release mortality, gonadal development and somatic condition of Australian bass Macquaria novemaculeata. Journal of Fish Biology, 75(10):2737–2755, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Helland:2009:MDB

[3464] I. P. Helland, L. A. Vøllestad, J. Freyhof, and T. Mehner. Morphological differences between two ecologically similar sympatric fishes. Journal of Fish Biology, 75(10):2756–2767, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Casper:2009:FHM

[3465] B. M. Casper and D. A. Mann. Field hearing measurements of the Atlantic sharpnose shark Rhizoprionodon terraenovae. Journal of Fish Biology, 75(10):2768–2776, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Svensson:2009:FOE

[3466] P. A. Svensson, J. D. Blount, E. Forsgren, and T. Amundsen. Female ornamentation and egg carotenoids of six sympatric gobies. Journal of Fish Biology, 75(10):2777–2787, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kendall:2009:AVV

[3467] B. W. Kendall, C. A. Gray, and D. Bucher. Age validation and variation in growth, mortality and population structure of Liza argentea and Myxus elongatus (Mugilidae) in two temperate Australian estuaries. Journal of Fish Biology, 75(10):2788–2804, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Yokoo:2009:MCO

[3468] T. Yokoo, T. Sakamoto, K. Kanou, M. Moteki, H. Kohno, P. Tongnunui, and H. Kurokura. Morphological characters and occurrence patterns of juveniles of two estuarine gobies, Acentrogobius kranjiensis and Acentrogobius malayanus, verified by molecular identification. Journal of Fish Biology, 75(10):2805–2819, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Gilbey:2009:SDG

[3469] J. Gilbey, E. Cauwelier, C. S. Jones, A. Mclay, L. R. Noble, and E. Verspoor. Size-dependent growth of individual Atlantic salmon Salmo salar alevins from hatch to first feeding. Journal of Fish Biology, 75(10): 2820–2831, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sulikowski:2009:ASS

[3470] J. A. Sulikowski, A. M. Cicia, J. R. Kneebone, L. J. Natanson, and P. C. W. Tsang. Age and size at sexual maturity of the smooth skate Malacoraja senta from the western Gulf of Maine. Journal of Fish Biology, 75(10):2832–2838, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Conway:2009:GFO

[3471] K. W. Conway and R. L. Mayden. Gill-filament ossifications: a possible morphological synapomorphy uniting the families Balitoridae and Cobitidae (Ostariophysi: Cypriniformes). Journal of Fish Biology, 75(10): 2839–2844, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kume:2009:SRM

[3472] M. Kume and S. Mori. Sea-run migratory behaviour in the Japan Sea form of three-spined stickleback Gasterosteus aculeatus in the tidal pool of eastern Hokkaido Island, Japan. Journal of Fish Biology, 75(10):2845– 2850, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Kaiser:2009:IPF

[3473] M. J. Kaiser and F. A. Huntingford. Introduction to papers on fish welfare in commercial fisheries. Journal of Fish Biology, 75(10):2852– 2854, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Metcalfe:2009:WWC

[3474] J. D. Metcalfe. Welfare in wild-capture marine fisheries. Journal of Fish Biology, 75(10):2855–2861, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Huntingford:2009:TAF

[3475] F. A. Huntingford and S. Kadri. Taking account of fish welfare: lessons from aquaculture. Journal of Fish Biology, 75(10):2862–2867, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Sandoe:2009:BNP

[3476] P. Sandøe, C. Gamborg, S. Kadri, and K. Millar. Balancing the needs and preferences of humans against concerns for fishes: how to handle the emerging ethical discussions regarding capture fisheries? Journal of Fish Biology, 75(10):2868–2871, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Evans:2009:EFW

[3477] J. C. Evans. The ethics of fish welfare. Journal of Fish Biology, 75(10): 2872–2874, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Hart:2009:BRF

[3478] Paul J. B. Hart. Book review: Fisheries for Global Welfare and Environment: Memorial Book of the 5th World Fisheries Congress 2008 — Edited by K. Tsukamoto, T. Kawamura, T. Takeuchi, T. D. Beard and M. J. Kaiser. Journal of Fish Biology, 75(10):2875–2876, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Albanese:2009:BRF

[3479] Brett Albanese. Book review: Freshwater Fishes of South Carolina — Edited by F. C. Rodhe, R. G. Arndt, J. W. Foltz and J. M. Quattro. Journal of Fish Biology, 75(10):2876–2878, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Shields:2009:BRS

[3480] Brian Shields. Book review: Salmonid Spawning Habitat in Rivers: Physical Controls, Biological Responses, and Approaches to Remediation — Edited by D. Sear and P. DeVries. Journal of Fish Biology, 75(10): 2878–2879, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:Eg

[3481] Anonymous. Erratum. Journal of Fish Biology, 75(10):2880, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Reyjol:2009:CV

[3482] Y. Reyjol, J.-P. Léna, F. Hervant, and D. Pont. Contents of volume 75. Journal of Fish Biology, 75(10):2881–2894, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).

Anonymous:2009:IVb

[3483] Anonymous. Index, volume 75. Journal of Fish Biology, 75(10):i–xxxii, December 2009. CODEN JFIBA9. ISSN 0022-1112 (print), 1095-8649 (electronic).